A Complete Bibliography of Publications in The Fibonacci Quarterly

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

25 April 2024 Version 1.166

Title word cross-reference

# [Blo53a].

(−1)^{n} [GZ22]. (2, 1) [Hau93b]. (2, F) [Dan98, SJC92]. (2, T) [SC97]. (3, F) [SCJ93, SCJ95]. (4k − 1) [Fer73]. (8r + 1) [Hun78]. (a, b) [HS18, Hei79]. (a − 1)^{n} + (b − 1)^{n} − (c − 1)^{n} = 0 [Lay75]. (a^{k} − b^{k})/(a − b) [Sch13a]. (d, 10) = 1 [BAP73]. (H − L)/15 [SH80b]. (m, F) [AH97]. (N + 1) [Adl89]. (n, ϕ(m)) = 1 [DeL82]. (N − 1) [Adl89]. (nr − s) [Kim81b]. (±1) [CHKS05]. (q, r) [CDM^{+}06]. (r + 1)f_{r+1} = xf′_{r} + (K − r + 1)x^{2}f_{r−1} [Say79]. (x(x − 1)/2)^{2} = (y(y − 1)/2)^{2} [Luo96]. (x, ay) = 1 [Zel92]. (x_{1} + x_{2} + ⋯ + x_{n})^{2} = x_{1}^{3} + x_{2}^{3} + ⋯ + x_{n}^{3} [Utz77]. + [HCG03]. 0.5k(ak + 1) [Ark70a]. 1 [Ant91, BP07, GC90, Hir13d, Kui69, Raw89]. 1 + 2^{p} + 3^{p} + ⋯ + n^{p} [dBdV94]. {1, 2, 5} [ZG14]. {1, 5, 10} [ZG14]. 1.13198824 … [Vis00]. 1/89 [Lon81a]. 1/d [BAP73]. 10^{2^{n}} + 1 [Wil88]. 10^{n} ≡ 1 (mod n) [KC89]. 10 × 10 × 10 [Ark73c, Ark74]. 150 [Cla08]. 151645121320 [Reb75]. 153 = 1^{3} + 5^{3} + 3^{3} [MW92]. $189.00 [Mos08]. 1 ≤ C ≤ 1000 [Rob90]. 2 [BD11, Cai96a, CKMR11, Coh99, DFLT14, How01a, KA03, Kar04b, Len94a, Len94b, Pla96b, Rei79, SS91, Sob17]. 2 ∗ n [Wil97]. 2, 3 [Min81b]. 2178 [Slo14]. 24 [HT86]. 25 · 10^{9} [PSW80]. 2^{a} ± 2^{b} + 1 [Her18]. 2^{k} [CJ96, Jar06]. 2^{m} ± 1 [BLS75]. 2^{n}k [HB76d]. 2K [Jac92]. 2S^{2} [Esw79a, Esw79b]. 2 × 2 [Gal68a, Gal68b, LF15, Lar15b, LO14]. 2x^{2} = y^{2}(3y − 1)^{2} ± 2 [PR02]. 2x^{2} = y^{2}(5y − 3)^{2} ± 2 [Rao05b]. 3 [AS76, AS81a, AP21, Cai96a, CS92, DL20, Kos20, Lev68a, Mel99c, MT17, ÖAD03a, Özk03, Rei79, Rob14]. 3/2 [Moo94]. 38 [Wol99]. 3^{k} [BB11, SC05]. 3k + 1 [Rob05b]. 3 × 3 [BH80a, Car72c]. 3X + 1 [LaD18, And00, And02, Hol11, Jos98, Zar09]. 3x + d [Hol11]. 3z^{2} + 1 [Ant85a]. 4 [DW91, Fil97a, How93, KA03, Mel99d, Sch97, Wat88]. 4^{m} [AH81]. 4k + 1 [Rob94]. 4 × 4 [Pad94]. 4 × n [Lig73]. 4x^{2} = 5y^{2}(5y − 3)^{2} ± 16 [Rao03]. 4|ϕ(|u_{n}|) [LS06]. 5 [You98]. 5^{k} [Nie72]. 5 log ϕ [Ade14b]. 5x^{2} [KK15]. 5x^{2} + 6x + 1 = y^{2} [Eme66]. 6 [Ker82]. 6000 < p < 7000 [KB64]. 6^{2^{n}} + 1 [Wil88]. 7 [Lud88b]. [0, 1] [SS91]. [an + 1/2] [Kim91a]. [x/c] [Str76]. ^{e} [AV03]. A!B! = C! [Hab19]. A [Alp18, Alp19a, Gue90]. a + a^{2}2^{p} + a^{3}3^{p} + ⋯ + a^{n}n^{p} [dB95]. A5^{n} − 1 [Wil87]. A7^{n} − 1 [Wil87]. a^{2} ≡ −1 (mod b) [Owi87]. A_{n+k} = ∑_{i=1}^{k} a_{i}A_{n+i−1}^{a_{i}} [Shi95]. A_{n}^{2}(x) [Bru75a]. ab + 1 [Hei79]. ai [HW81]. α [fC03, DRS11b, Fil98c, SS01]. α^{i} [SS01]. an^{2} + bn [And84]. ax + b [Str76]. ax + by = n [Tri00b]. B [CKR98, LL87a, Bal13, GGL21, Har78, Man75]. b^{2} ≡ −1 (mod a) [Owi87]. β [TR01]. binom(x, p) = binom(y, 2) [Kis88]. binom(y + 1, x) = binom(y, x + 1) [Owi79]. modp^{n} [Sim98]. C [CP84, Sha12]. C^{n} [Tau76]. C_{n+2} = C_{n+1} + C_{n} + n^{m} [WH70]. C_{n,3} = C_{n−1,3} + C_{n−3,3} + r [Djo05b]. C_{n} = C_{n−1} + C_{n−2} + k [BJ90]. C_{n} = C_{n−1} + C_{n−2} + r [Zha97c]. cosh x/ cos x [GT72]. ∪_{n∈N}N^{n} [Mai70]. CX^{2} [Rob90]. d [FF89, KS02a]. D(4) [FHT10, Fil15, RHT18]. d_{k}(n) [EK69]. DQ^{2} = R^{2} + N [Eme69]. E [TR04, AHS87]. {E^{2} + (λp − 2)E + (1 − λp − λ^{2}q)}^{m}G_{n} = n^{k} [Mal83]. e^{x} [Bre74a]. ℓ_{p}(F) [Ogu20]. exp(1 − e^{x}) [UC69]. F [FF89, Ter96a]. F (2, 9) [Hol95]. F (F (n)) [Par77a]. f(m) − f(n) [Cal97]. f(A) [BR82]. f(x) = (x + 1)^{−1} [Bra95]. F_{n}^{2} + F_{n+1}^{2} = F_{2n+1} [Mel99e]. {F_{n}^{2}/d} [FF89]. F_{12n} [Goo71]. F_{18} [Pon19]. F_{184} [AL19]. {F_{2k}, 5F_{2k}} [Fil15]. F_{2k}, F_{2k+6}, F_{2k+4} [FHT10]. F_{3n}/F_{n} [Jar68b]. F_{3qr}(x) [Bei78]. F_{5n}/F_{n} [Jar68b]. F_{5n} ≡ 0 (mod 5) [Hir13c]. F_{ai} [HW81]. F_{F_{n}} [Cam22a, For67a]. F_{k+1}/F_{k} [Som75]. F_{kn}/F_{n} [FP96]. F_{mn} [ES16e]. F_{mn+r} [Des86]. F_{n} [DD65, ES16e, Hag64, HS99, Luc03d, Mor83]. F_{n} = P (x) [TU20]. F_{n}^{2} ± F_{k}^{2} = K^{2} [BJ79a, BJ79b]. F_{n}^{3} [Ozb12]. F_{n_{1}} + F_{n_{2}} + F_{n_{3}} + F_{n_{4}} + F_{n_{5}} = 2^{a} [TD22]. F_{nm} [Gla78]. F_{n} ± F_{m} = y^{a} [KKLL21]. F_{p} [Tao17]. F_{p−(5/p)} ≡ 0 (mod p) [Som72]. F_{p−ε}/p [Wil82]. F_{u} [Hal65a]. F_{x}^{2} ± F_{y}^{2} [Sha65]. \frac{1}{2}(F_{n} − 1)(F_{n+1} − 1) [Chu20]. \frac{1}{2}(F_{n} − 1)(F_{n+2} − 1) [Chu20]. g [Ten09, ZJ14]. G_{α} [Bre74b]. G_{j,k} [Gre78b]. G_{n} = G_{n−1} + G_{n−c} [BJS96]. GCD(3^{n} − 2, 2^{n} − 3) [Izo05]. Gl_{2}(C) [Dam89]. H [HB75d]. H_{n} [GD09, WH74a, BJE95, Kom10]. H_{n+1}/H_{n} [GD09, Kom10]. H_{n} = H_{n−1} + H_{n−2} [BJE95]. ∞ [BMRS04, CMR02, MRS99, MRS00, MR00]. J_{1}(z) [How85]. J_{k}(z) [How87].

K [AC11, Cha91, Cul17, Djo05a, Er83b, Ker82, KK01, Kla66a, Lee87, LLS97, PM82, Phi83, PGP85, PG89, PA91, SJ84, Tau73b, TS89b, Zah82, Zho96, Ark73a, AS74, AAB^{+}90, BL93, Bol84, BL13e, Cai96b, Cha86c, CM14, DP16, DL20, DOU08a, DOU08b, Dem08, DGMS14, Dun69b, ESS20, EMS10, Er83a, EA12, ES17, FP09, Flo67, FF99b, Gab70, GGL24, GPS13, HS12, Hau00, Hir12b, IA11, Kah05, KŞ13, KS04, Kil07, KK89, Lee00, LLKS01, LK03b, Luc21e, LS18b, aMS22a, McD87a, McD87b, NKW23, ÖTA05, OS05, PP94, PW23, Pet80, Phi14, Puc01, RL18, Sha74d, Sir97, Som89, Som07, Ste78a, Sza17, TK04, Tót02, Wan95, WS80, XPP87, lY08, YK03, You92a, You92b, Zho03, ZH03]. k^{2} + 1 [Wil75a]. k · 2^{m} ± 1 [Sun06c]. L [LY13]. {L_{n}^{2}/d} [FF89]. L^{p} [KS88]. l_{1} [DL92]. L_{18} [Pon19]. L_{2n+1} [Aga91]. L_{5n}/L_{n} [Jar68b]. L_{5p} [Jar68a]. L_{m} [Mor83]. L_{m}^{r} [ES10e]. λ [HM83]. λ × N [HM83]. ⌊n^{2}α⌋ − n⌊nα⌋ [Hen11b]. ⌊nα⌋ [Bun06]. ⌊nΦ^{m}⌋ [Len06b]. LU [Pro22]. M [BCOR95, BB69, BCHR91, Cat74a, Cha86a, DF92, Kwo89b, Kwo89a, PB89, Ryd96, SK12, Woo79, And74a, BPT14, Coh67, Col83, Des78b, Des86, Ehr89, Ehr94, FRU15, For67b, Ful81c, Gau04, JMNRS22, Lev85, Mam61, ÖAD03a, Rob63, Sch19, Sha68, SK17, The71, TS00, Vin63, Wad78, Wad92a, Wal60]. M(a) [Gue90]. m − n [Cal97]. m = −1, 1, 2 [Len06b]. m^{2} − 4k = 5n^{2} [Kim07]. Q [Kir22]. Q( ¯ 5pF_{p}) [Tao17]. N = ⟨ · ∈^{\} − ∞ [Rie69]. M × N [Eng90]. N! [Kun87]. N [Buc64a, Buc64c, Adl89, BJF99, Boa79, Buc64b, CST05, CVZ17, Coh91, Edg72, Eng01, Jon96, Kla66b, Kla68b, Mai70, Oli96, Sil77a, Ste78b, Tur79, Vin78, Vor09, Vos98, WJ76, All81, Bai83, Ber68, BHM92, Bok84, CS89, Coh67, CDW13, Col83, DeL84, DB66, DB67, Ehr90, Gla78, Goo75, Gru94, Gru01, Hah72, Hen94, Her78, Her82, Hir12b, Hir13d, Hir14a, HJ68a, Igu94, Kis14, LL87a, Luc03d, LS06, LY90, Ran66, Rus12, Sch20, SK12, Ste71, Tau73a, Wil86, Wil97, Wil98d]. n(n − 1) − 1 [FB99]. n(x, y) [Aga87]. N + 1 [Edg72]. N + 2 [Edg72]. n > 100 [DD65]. n > 4 [HE80]. N^{2} ± 1 [WJ76]. N^{n} [Mai70, Som81]. N_{n} = x^{a} ± x^{b} + 1 [RB22]. Nb^{2} = c^{2} + N + 1 [AL79b]. nF_{n} [HS99]. N × N [BB69]. n|Q(n) [Du00, Lin98]. O [AJ96]. O(k^{2} log n) [Pet80]. O(k^{3} log n) [Pet80]. O(log n) [Sho78, WS80]. Ω(n) [All81, Luc03d]. P [AJ96, How93, JO24, Kar96, Rob83b, Rob84, AS00, Bal19b, Bal19a, BJ02, BSY05, BYD^{+}19, Bou20, Dav14, DeL77, DK18, GMP91, GKRS01, KA03, KP77, Len17, Li00a, Lon81c, LMR96, MR15, MZ92, Nyb23, PT18, San16, Sla00, Som75, Som81, Som91, SC04, Som06a, SK15a, SK16, SK21, Sun78, TKS11, Wil98c, YK77, You94, Zho99]. p + 1 [DK18]. p > 5 [Jar68a]. p^{3} [KW95]. p^{α}M^{2β} [MH75a]. p^{a} ± p^{b} + 1 [LS07]. P^{e} [CJS94]. p^{k} [CM07]. p^{r} [Som07]. P_{2k}, 2P_{2k+2} [AFT22]. P_{2n+1}(F_{k}) = F_{(2n+1)k} [Bev76]. P_{2n+1}(x) [Bev76]. P_{m} × P_{n} [Har94]. P_{r,k} [MR85]. P_{t} [MB89]. p ≡ 1 (mod 3) [KS10c]. p ≡ 3 (mod 4) [Bru70]. Φ [Hir15b, Tót87, Cha06a, Cha09, GL83, HK78, Jon90, Jon91b, Leu23, Wan93, WW06]. ϕ(F_{n}) ≡ 0 mod 4 [HE80]. Φ(n) + 1 [CS89]. Φ(x) = n [SG85]. Φ_{2}(x) [The71]. π [Cha06a, Cha09, Dra69, Fre92a, Hau94b, Hir15b]. π = 2 ∑ arccotf_{2k+1} [And78b]. Π_{j≠i}n_{j} ≡ 1 mod n_{i} [BJ95]. ±1 [Gal68a, Gal68b]. (mod m)

[Cav77a]. (mod n) [Bro97]. (mod p) [Sun06a]. \binom{∏_{k=0}^{n} n}{k} [Hir13a]. P X^{2} [Rob83b, Rob84, Zho99]. Q [Gou81, Bic65, Hen15b, How93, Ivi72, Lev01, Lin67c, MS95g, Ste81a, Ste81b, Tau68c, Zha99a, Aga91, AS84, AJ92, AB20, BKV22, Car74c, Car74d, Car75a, Cig03, CHT06, CP09, FT14, Gou67a, JLW07, KL93, Mun05, Smi09, Yan88, Yan91]. Q(\sqrt{5}) [Dod84, Lin68]. Q_{k}^{∗} [LLS97]. QR [Mat95]. QX + 1 [Ste81a, Ste81b]. R [Tau68b, AP92, AB85, BH14a, Dub89, DMRS97, ER00, EMR02, HP75a, Hor86a, HC11, LP03a, MR99, SC10a, Som87, Swa77, TR01, TMR02a, TMR02b, TR03, WR01]. R(3, 3, 3) [SC84]. R(N) [Eng01, Sto09a]. r, s [Ful80b]. ℜ^{3} [Nor19]. S [Esw79a, MKL^{+}23, DMP07, ESW13, Fin91, Gou74c]. S(n, k) [Kwo89a]. S^{n} [Gle81]. sec [Boy07]. sech [Boy07]. σ(m)σ(n) = (m + n)^{2} [Kis81]. σ(n) = 2n + t [Rob80]. σ(n) = σ(n + 1) [GS74]. σ(θ(n)) [Bal94]. Sl_{2}(C) [Dam89]. ∑ 10^{−k(i+1)}F_{ai} [HW81]. ∑_{1}^{n} F_{k}(x)y^{n−k} [Swa77]. ∑_{a=1}^{p−1}(a/p)a [Mon90b]. ∑_{j=1}^{k} jF_{j}^{p} = F_{n}^{q} [AL20]. ∑_{j=i}^{n+i−1} F_{aj−b} [Hen07]. ∑_{k=1}^{n} k^{m}F_{k+r} [Bro67f]. ∑_{k=1}^{n} k^{r} [Kha81]. ∑ a^{k}k^{p} [HT00]. ∑ r^{k}x^{r} [Gau89]. T [Bol84, Bol86, Bol89, Sel83, Tur79, Tur88b, GKH77, KK01, KS16, Lyn70, Rob00]. T_{3} [Bol89]. T_{n} = T_{n−1} + T_{n−2} + hn + k [FF99a]. tan [Boy07]. tanh [Boy07]. τ [EH07, LM19b]. u [Hal65a]. u(3, 1) [DN96]. u^{2} − Dv^{2} = C [DeL81]. u_{0} = u_{1} = 1 [DN96]. U_{2p} [Som06a]. U_{k} [Som06a]. u_{n} [Man68]. u_{n+1} ≡ u_{n} + u_{n}^{−1} (mod m) [Nag84]. u_{n+1}u_{n−1} − u_{n}^{2} = ±1 [HBJ78c]. u_{n+2} = u_{n+1} + u_{n} [Zöl93]. U_{n} = binom(x, 3) [Sza02]. U_{n} = Uqx^{2} [AJ92]. U_{p} [Som06a]. v [Hu02]. V_{n} = binom(x, 3) [Sza02]. V_{n} = V qx^{2} [AJ92]. w^{2} − 1, w^{3} ± 1 [Rob81]. {W_{n}} [Gau98]. W_{n}(a, b; p, q) [Hor67, LL87d, Hor69, SH79a]. W_{n+2} = dW_{n+1} − cW_{n} [Zei72]. W_{n+2} = dW_{n+1} − dW_{n} [Zei65]. X [FL18, KLM^{+}19, Ano16a, Dda20, DD03]. X, X + 1, Z [FL68]. x^{2} [Kar69]. x^{2}(3x − 1)^{2} = 8y^{2} ± 4 [PR01]. x^{2}(5x − 3)^{2} = 20y^{2} ± 16 [Rao02]. x^{2}(5x − 3)^{2} = 8y^{2} ± 4 [Rao05a]. x^{2} + (x + 1)^{2} = y^{2} [Kar69]. x^{2} + 2^{a} · 11^{b} = y^{n} [CDL^{+}10]. x^{2} + 7^{2k} = y^{n} [LT07]. x^{2} + a^{2}y^{m} = z^{2}n [Zel92]. x^{2} + xy − y^{2} = ±D [PH75]. x^{2} − Dy^{2} = ±1 [Wul75]. x^{2} − k = T_{n}(a^{2} ± 1) [Udr98]. x^{2} = 5y^{2} − 4 [Per16]. x^{5} ± p^{2}x − k [FE99]. x^{5} ± x^{3} + n [SW98]. x^{m} [ISTY88]. x^{n} ≡ a (mod m) [DeL82]. x_{n+1} = (x_{n}^{2} + A)/(x_{n} − 1) [Alp11]. xa ± xb ± 1 [LL14]. y^{n} + (y + 1)^{n} + ⋯ + x^{n} [WB80]. Z [BA98, Mat73]. z(p^{2}) = z(p) [Hee84]. Z[\sqrt{2}] [Weg77]. Z[ζ_{12}] [EI03]. Z^{k} − Z^{k−1} − 1, k > 0 [Fer76]. Z_{2} [Ava11]. Z_{2}[i] [Jos98]. Z_{p} [Mul80b, Mul80a].

-Additive [Fin91]. -Adic [BSY05, Coh99, GMP91, Har78, Len94a, Man75, MR15, San16, You94, AS00, Bal19b, Bal19a, EH07, Len17, MZ92, PT18, Sob17, Ter96a, Dav14]. -Algorithm [TR01, TR04]. -analogs [JLW07]. -Analogue [Smi09, CHT06]. -Analogues [Aga91]. -Ary [GKH77, Cha86c, Hau00]. -Bell [HM83]. -Binomial [AB20, Gou67a, How93]. -bonacci [BPT14, JMNRS22, Swa77, LL87a, PW23, Sir97, Kir22]. -Bonacci-Like [AC11]. -by- [Fil97a, Mel99c, Mel99d, Pla96b]. -Canonical [Sil77a]. -Cassini [Alp19a, Alp18]. -circulant [SC10a, ZJ14]. -Color [CDW13, Sha12]. -convex [DRS11b]. -Convolution [HB75d]. -Coordinates [Dda20, FL18, KLM^{+}19]. -Core [Rob00]. -Cube [AS76, AS81a]. -Cubes [AS74, Ker82]. -Decomposition [Pro22]. -decreasing [BKV22]. -Deficient-Perfect [AP21]. -Densities [BA98]. -Determinants [Yan91]. -Digit [Puc01]. -Dimensional [Bai83, Kis14, BP07, Jon96]. -Diophantine [RHT18]. -Discrepancy [KS88]. -Enumeration [CP09]. -Expansions [Leu23, SS01]. -extensions [OS05]. -Factors [GC90]. -Fibonacci [BH14a, BYD^{+}19, BL13e, Car74c, Car75a, Cig03, CDM^{+}06, DGMS14, ESS20, EMS10, FP09, HC11, LK03b, Lyn70, Mun05, NKW23, RL18, Tau73a, Tau68c]. -Fibonomial [Gou74c]. -Free [Dun69b, PB89]. -Function [GL83]. -Functions [LY13]. -Generalized [BMRS04, CMR02, CM14, Dub89, DMRS97, ER00, Flo67, KS04, LLKS01, MRS99, MRS00, MR99, MR00, TR01, TMR02a, WR01, DL20, GGL24, Luc21e, ÖTA05, TR03, lY08, LLS97]. -Gonal [ES17, Ten09]. -Identity [Car74d]. -In-a-Row [Bol84]. -inner [Zah82]. -Invariant [CHKS05]. -Laguerre [AS84]. -Latin [Kar96]. -Legal [MKL^{+}23]. -Lucas [Kah05, DOU08a, DOU08b, Dem08, DGMS14, Lee00, NKW23, YK03]. -manifold [MT17]. -manifolds [CS92]. -Matrices [MS95g]. -Matrix [Gou81, Ivi72, Zha99a]. -Metrics [DL92]. -Multiperfect [AHS87]. -Niven [Cai96a, Gru94, Wil97]. -Normal [HP75a]. -Number [Bok84, Ehr90, LY90, Lud88b, CST05]. -Numbers [CP84, Tau73b]. -Order [Sza17]. -packing [CKMR11]. -Pair [Fil15]. -Partitions [Jon91b, Wan93, WW06]. -Periodic [BD11]. -Polynomial [Yan88]. -polynomials [TKS11, Tau68b]. -Positions [JO24]. -Power [LL87a]. -rational [Bou20]. -Reflected [Aga87]. -Regular [Rob14]. -Regularity [MR15]. -Representation [FF89]. -repunits [GGL21]. -Reverse [You92a, You92b]. -Riven [Gru01]. -Seidel [FT14]. -Self-Numbers [Cai96b]. -Separation [KL93]. -Sequences [Bol84, Gab70, Lev68a, Edg72]. -Series [Nyb23]. -Sets [MB89]. -Shuffle [PP94]. -sion [KS16]. -Smith [McD87a, McD87b]. -Spheres [Fer73]. -Square [TS00]. -Stability [SC04]. -Step [ÖAD03a, Özk03, Sch19, KA03]. -Strong [DF92]. -Subcomplete [LP03a]. -th [Buc64a, Buc64c, DB66, Gau04, Lev85, Buc64b]. -th-order [LS18b]. -tic [Coh67]. -TLB [KS02a]. -Triangles [Hu02]. -Triple [FHT10]. -Triples [Hah72]. -Tuples [Sch97, You98, Oli96, Woo79]. -Values [Fil98c]. -Words [fC03]. -Zeckendorf [EA12, AC11].

0 [Gal08, Sch94]. 0-7923-2491-9 [Sch94]. 01021a [Ano88c].

1 [Kih00]. 1-4020-4245-0 [Mos08]. 1-4020-4246-9 [Mos08]. 10018a [Bri83a, Bri83b]. 11th [Ano03a]. 133 [BL13d]. 147 [Sub07]. 18th [BJ18]. 1992 [BHP92, BPH93, Swa92]. 1994 [Bro96, NS95]. 19th [Joh20]. 1st [Di 93].

2 [Boo79, Ove67, PHJ67]. 2-adic [And20]. 2-Class [Tao17]. 2-Dimensional [EH22]. 2008/2009 [Cat09]. 20th [Joh22b]. 21 [Bri83a, Bri83b]. 218 [FS14]. 21st [Jef08]. 25th [IEE84]. 2n [Rob87].

3 [Gou68, Pet94]. 315 [Wer83].

45th [WK09]. 46th [WK09]. 49th [IEE08].

5 [Høy09, Kai22, Kih00]. 536 [Bro68a]. 557 [Pic79]. 57 [GP93]. 5th [IEE81].

7 [PHJ67].

806 [MS00a, MS00b]. 84f [Bri83a, Bri83b]. 88m [Ano88c]. ’89 [Gon89].

9 [Sch94]. 92h [GP93]. ’96 [ACM96]. 976 [Gal08]. 976-1-59102-475-0 [Gal08]. 978 [Høy09]. 978-0-387-72930-5 [Høy09]. 978-0-691-17486-0 [Tol17b, Tol17a]. 978-1-912827-03-9 [Hal22].

= [Jon90].

A. [Kar70, Puc62, Sha87, Sla00]. A000975 [Sto17]. A013583 [FBJ01]. abū [Sha88a]. Abaci [Cio18, Hor04, Ano93i, Sig02, Eis51, Geg93, Miu81, Ago49, Fra02, Pis02]. Abacus [Tor19, BF90, Fra03, Uli11, PB15]. Abbaci [FGD20, Lün92, Lün92, Fre22, Han11, Dev12, Han07, Spi00a, Spi00b]. Abbaco [Høy05]. ABC [Pen20, Yab07]. ABC-conjecture [Yab07]. Abel [Sto70]. Abelian [DMSS16, Smi68, Yap70]. ABFIA [EGS22]. abridgement [Hug10]. Absolute [Moo92, Tue02, DH17]. Absorption [Ste79]. Abstract [HP75a, WW75, Kel08, Sug93, Zah82]. abu [Sha90]. Abundant [Wal84a]. Academic [Ano88c, Sch94]. Accelerated [GFMR^{+}24]. Acceleration [BL86, Gri92, Phi07a]. Accelerations [ABCM11]. Achieving [Wlo67]. Acknowledgement [Hog68c]. Acknowledgements [Ano86a]. Acknowledgment [HKN11, Wil98b]. Acknowledgments [Ano65a, Ano82a, Ano83a, Ano84a, Ano85a]. ACM [ACM12, Gon89]. ACM-SIGSAM [Gon89]. Across [BYD^{+}19, HBJ79a, LF15, JH73]. activity [Cla08]. Acute [HJ68a]. Acyclic [EG89, QM90]. Ad [Sha87]. Adam [Hal22]. adaptive [PC13]. Add [Dek21, TLC93]. add-with-carry [TLC93]. Addenda [Ano78a, Ber84, BJ79a]. Addendum [Ano77a, CSH72a, FH94, Goo86b]. Addition [BH14b, Day70, LN84]. Additional [Alf63a, GG79c, Kos18b, Kos23b, Kos23c, Kos23d, Kos23e, Kos23a, Kos23f, Pro16a]. Additions [Bra71, Cap89]. Additive [Alu96, Ewe07a, Fin91, Hig72, Hog77a, Hog77b, Hog80, HBJ84, Kir77, MRPC95a, MRPC95b, RS11a, Tri69, Krá72a, Krá72b, Mas94b]. Adic [BSY05, Bro03, Coh99, GMP91, Har78, Len94a, Man75, MR15, San16, You94, And20, AS00, Bal19b, Bal19a, EH07, Len17, MZ92, PT18, Sob17, Ter96a, Dav14]. Adjacent [D’S89]. Adler [Rob78b]. Advanced [And90, Ano78a, Ano79b, Bro96, FL22, Fro23a, Fro23b, Fro23c, Fro23d, Fro24, Hog63a, Hog63b, Hog63c, Hog63d, Hog64a, Hog64b, Hog64c, Hog64d, Hog65a, Hog65b, Hog65c, Hog65d, Hog66a, Hog66b, Hog66c, Hog66d, Hog67b, Hog67c, Hog67d, Luc03a, Luc03b, Luc03c, Luc04a, Luc04b, Luc04c, Luc04d, Luc05a, Luc05b, Luc05c, Luc05d, Luc06a, Luc06b, Luc06c, Luc06d, Luc07a, Luc07b, Luc07c, Luc07d, Luc09a, Luc09b, Luc09c, Luc09d, Luc10a, Luc10b, Luc10c, Luc10d, Luc11a, Luc11b, Luc11c, Luc11d, Luc12a, Luc12b, Luc12c, Luc12d, Luc13d, Luc13a, Luc13b, Luc13c, Luc14d, Luc14a, Luc14b, Luc14c, Luc15a, Luc15b, Luc15c, Luc15d, Luc16a, Luc16b, Luc16c, Luc16d, Luc17d, Luc17a, Luc17b, Luc17c, Luc18a, Luc18b, Luc18c, Luc18d, Luc19a, Luc19b, Luc19c, Luc19d, Luc20a, Luc20b, Luc20c, Luc20d, Luc21a, Luc21b, Luc21c, Luc21d, Luc22a]. Advanced [Luc22b, Luc22c, Whi67a, Whi68a, Whi68b, Whi68c, Whi68d, Whi69a, Whi69b, Whi69c, Whi69d, Whi70a, Whi70b, Whi70c, Whi70d, Whi71a, Whi71b, Whi71c, Whi71d, Whi72a, Whi72b, Whi72c, Whi72d, Whi73a, Whi73b, Whi73c, Whi73d, Whi74a, Whi74b, Whi74c, Whi74d, Whi75a, Whi75b, Whi75c, Whi75d, Whi76a, Whi76b, Whi76c, Whi76d, Whi77b, Whi77c, Whi77d, Whi77e, Whi78a, Whi78b, Whi78c, Whi78d, Whi79a, Whi79b, Whi79c, Whi79d, Whi80a, Whi80b, Whi80c, Whi80d, Whi81a, Whi81b, Whi81c, Whi81d, Whi82a, Whi82b, Whi82c, Whi82d, Whi83a, Whi83b, Whi83c, Whi83d, Whi84a, Whi84b, Whi84c, Whi84d, Whi85a, Whi85b, Whi85c, Whi85d, Whi86b, Whi86c, Whi86d, Whi86e, Whi87a, Whi87b, Whi87c, Whi87d, Whi88a, Whi88b, Whi88c, Whi88d, Whi89a, Whi89b, Whi89c, Whi89d, Whi90a]. Advanced [Whi90b, Whi90c, Whi90d, Whi91a, Whi91b, Whi91c, Whi91d, Whi92a, Whi92b, Whi92c, Whi92d, Whi93a, Whi93b, Whi93c, Whi93d, Whi94a, Whi94b, Whi94c, Whi94d, Whi95a, Whi95b, Whi95c, Whi95d, Whi96a, Whi96b, Whi96c, Whi96d, Whi97a, Whi97b, Whi97c, Whi97d, Whi98a, Whi98b, Whi98c, Whi98d, Whi99a, Whi99b, Whi99c, Whi99d, Whi00a, Whi00b, Whi00c, Whi00d, Whi01a, Whi01b, Whi01c, Whi01d, Whi02a, Whi02b, Whi02c, Whi02d, Whi03]. Advantage [NN87]. Adventures [Tor19]. Affects [LHWS96]. Again [Car77a, Hin77, Hun66a, Hun72]. Ages [GG69, GGA83, Bro70a]. Agrees [AC11]. Agronomof [Tue23a]. Aha [Gin02]. Ahead [Nod90]. aided [Rol03]. Aitken [Jam90, LMV09, MP85, Phi84, Phi07a, Zha98b]. al [Sha88a, Sha90]. al-Karajī [Sha88a]. al-Karaji [Sha90]. Alan [CS12, Teu04]. Alfred [Ano88f, Har09, Gal08]. Algebra [Bic69, Erc76, Goo68a, Goo68b, HA66, HS13, LK03a, Miu81, SG76, Fra40, FR85, LK03b, Sha88a, Sha90, Alf66a]. Algebraic [Alf64f, BPGC97, BPIK07, CZZ11, Chu90, Coo14b, Erc73, Gon89, Hsu73, Kli81, Law83, Mel16b, Mel16g, She72, TB91, Wiȩ80, Par17]. Algebras [JK94, Par17]. Algorithm [BY16, Boy97, Boy98, Bro72a, BD71, Cha01, Cre89, Day70, Dun66, Eng01, Epa85, Er83b, Har70a, Har74, HF91, Jen97, Kar69, Kel65, Kno92a, May87, MM73, Moo89, Moo92, Pih92, Pih01, Pih10, PP79, Rie76, Swa94, TR01, TR04, TE81, Wei79, DM08, EGS22, Joh04, MSPF95, MR84, Pet80, Tak00, WS80, Boo79, MS00a, MS00b, Ove67, PHJ67]. Algorithmic [Ewe01, Rab96a, Rab99a, Rab99b]. Algorithms [AUFP13, Ani95, AG87, BH69a, Bro74a, CLX08, DMSS16, FT84, FT87, Hir73, Kar74, Knu97, SY89, SK81, CLX13, DMR^{+}16, FLP08]. aliquot [IL07]. Alladi [HCG03]. Alley [Boy97, Boy98, Jen97, Swa94]. Allied [Joh40, Sin80]. Allocation [Fer76, Hil81e, Hir73, McI82]. Almost [Hen11a, Jac89, JRdlR10, Kim81a, Kim81b, Knu66a, MW22, Nyb98, Nyb09, Som79, Sub99, TA02, Fan09, Knu66b, LS06, Som06a]. Almost-Isosceles [Nyb98]. Almost-Recursiveness [Hen11a]. Almost-sure [JRdlR10]. Along [MW22]. Alongside [Ber11]. Alpha [Tur89a]. alphabet [GSS19]. Altered [DT71]. Alternate [Hin80a, Kwo10, Kwo14, Rob02a]. Alternating [Gau12, GL01, Goo94b, Len17, Mel00a, Mel16g, Mos93, Rac19, Tan75]. Alternative [DGST88, Hor94a, McL79, Mel03b, Smi93, Cox87, Hol95]. Alternatives [KBN08]. AM1 [SBM97]. Amateur [Man64, Man67, Man68, Man77]. Amazing [Che12, Lau10]. Ambrosia [Pic83b]. Ambrosiana [Pic83b, Pic83a]. American [Bur92]. Amicable [BN77b, BN93, Hag87a, McC85, Naj89, PS21]. Among [HBJ77a, AKO^{+}03]. Analog [Hod74, LR85, Lee87]. Analogous [Can03a, Can03b]. Analogs [Mel99e, Wal87a, JLW07]. Analogue [AK74, Bro70d, Hey80, Mak12, Shi20, Smi09, CHT06, KA19, SS95a, SS95b]. Analogues [Aga91, Bha16, Hor86c, Nor14]. Analysis [And74b, Ark65b, CFG19, Goo91, Gru01, Har02, RH82, Sen73, Pic79, Rob78b]. Analytic [Byr63, Byr65, CR93, Dav77, DP78b, Hag64, Nav01, Yoo21, Rol03]. Analytical [HF73]. Analyzing [Bro70c, Bro72a, Wei73]. Ancient [Moo70, RT02, Sin85]. André [Swa00c]. Andrews [BHP92, BPH93, BAS93, Bri06, Cha07, Dob14, Gup78a, Hir81, Pad89]. Angle [DeT81, Gla70, Hog68b]. Angled [Nyb98]. angular [BPIK07]. Ann [IEE81]. Anniversary [Ano71c, Ano72g]. Annos [Sha87]. Announcement [Ano66a, Ano66b, Ano72a, Ano80a, Ano82b, Ano83b, Ano84b, Ano84c, Ano85b, Ano85c, Ano86b, Ano86c, Ano87b, Ano87c, Ano87d, Ano88d, Ano88e, Ano89d, Ano89a, Ano89b, Ano89c, Ano90b, Ano90d, Ano90c, Ano90a, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano93a, Ano93b, Ano93c, Ano93f, Ano93g, Ano93e, Ano94c, Ano94a, Ano94b, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano99a, Ano99b, Ano00a, Ano00b, Ano00c, Ano01a, Ano01b, Ano02a, Ano09a, Ano10a, Ano11a, Ano11b, Ano11c, Ano12a, Ano13a, Ano14b, Ano17d, Ano18b, Ano19a, Ano19b, Ano19c, Ano19d, Ano20a, Ano20b, Ano21a, Ano21b, Ano22a, Ano22b, Ano24]. Announcements [Ano17c, Ano18a, Ano17b]. Annual [Ano71b, Ano87b, Ano97b, IEE08, IEE84, Jef08]. Anomalies [Iak81a]. Ansatz [DB22]. Anthesis [Mun71]. Anti [AJS22, LF15, Lar15b]. Anti-Diagonals [LF15, Lar15b]. Anti-Palindromic [AJS22]. Antimagic [Tri74a]. Antiprisms [Mab19]. Antisocial [Lew95]. Any [AML85, Jar68a, KK01, Man68]. Apery [Cha84a, Hir12a, Hir13b]. apex [Rob78b]. Apollonius [Bau80, Hor73, Tri74b]. Apparition [DK18, Des78b, KP77, Sal75, Vin63, Wil75b]. Appear [BDD^{+}17]. Appearance [CVZ17, Fer97, Hor63a, Mar12a, Mar12b, Mar12d, Mar12c, Mar13a, Mar13b, SK13b, LP14]. Appearing [BJ89, BJ93, Hen20, Kou90, Lin90]. Appears [Bas63e]. Appending [JM14]. Appl [Sol07, UNS07]. Application [BH77a, Ber68, BH69a, BD77, BAS93, Car84, CMR02, CLRS86, Coo84, Dun67b, Erc76, Fer78, FKL70, Goo91, Gou69, Hog70a, Hsu73, mJ10, Ma98, Mig77, MH16, MR99, Rya75, Swe64, TR01, TMR02b, Tho64a, TJ92, Weg81a, DÖ03, Kar04a, PW95, Sob17, Sug93]. Applications [Alp19a, AKP19, Bai83, BPH93, Bro64b, BG76, CP84, Cha84b, DS20, DGST88, HCD97, Joh22b, Kra90, Mun17, ND87, ÖAD03b, Pha84, PGP85, Puc62, Ris05, ZZ06, Zho03, AK05, BHP92, EB88, IA11, KM18, Kos01, Kos09, Ran95, Vaj89, Vaj08, YZ17, Ano89d, Ano89a, Ano89b, Ano89c, Ano90b, Ano90d, Ano90c, Ano93k, Ano94c, Ano97a, Ano99a, Ano99b, Ano01a, Ano01b, Ano02a, Ano02e, Ano03b, Ano03a, Ano09a, AKP20, BHP92, BPH93, BJ18, Coo10, Coo16, DMY13, Fre87, Fre88, Fre90, Joh09, Joh12b, Joh20, Phi04, dBF85, Ano93f, Ano93g, Gau02, Gau10, Gra89c, Sch94]. Applied [BJ03, GR93, Sha74d, Sub07, Flo18]. Approach [BL17a, BP09, Ber76, Bro67f, FH88, Gal86, Hen14, Hsu95, JV19, Kwo10, Lar16, LO18, LN07, Mel03b, Mil75a, PS80, PS19, Sin90, BCQ07, BP10b, Cox87, GTU13, Spi00a, Spi00b]. Approaches [GD09, Hen94, Kom10]. approche [Spi00a, Spi00b]. Approximating [Crî14, Hir11a]. Approximation [All75a, BMRS01, BT00, Els96, Els98, Law83, PVB98, Rie99b, TD16, TD19, Ton97, Glu76, Nea07]. Approximations [CR93, Els00, Els03, Kom01a]. Arabic [Sha90]. Arbitrarily [Jar67a, SK21]. Arbitrary [AJ91d, Bun75c, DeC70, Lud87, Sha74e, Wit72]. Arbor [IEE81]. Arch [DeT74]. Archimède [Lan81]. Architecture [Pre68]. Architectures [And90, Mat95]. Archives [Mos08]. Area [Gou73, PS92, Tre96b, Tre96a]. Area-Bisecting [PS92]. Areas [Her19, JC17]. Argand [WSH74]. Arguments [Mey05]. Arises [KH82]. Arising [CFH^{+}16, CFH^{+}17, Gri11d, HB76a, HS84, JH73, Kis14, Tue22, Wil00, Gri15a]. Arithmetic [AUFP13, Alo76, Ani95, AESS04, Bac81, Bal90, BN77a, Bre74b, Bro97, CSV73, CS95, Dam89, DJ67, Fen03, GKL^{+}21, GG15, Gri74, Gua78a, Gua78b, Gua78c, Gue78, Han70, HS98, Hau88, Hau93b, HPV95, Hsu97, IEE81, JP19, Jar67a, KK05, Kim81a, Kim81b, LY02, LJ67, LJ70, Lon81b, Luc99, Luc00, Mac00b, Mah91, PT17, RP19, SP18, Sch84, Ste73a, Vau76, Xie02, BBB06, CHL19, Dev11, FP09, Lün91a, Lün91b, PZ12, SE20a, Sho78]. Arithmetic-Geometric [Hsu97, Xie02]. Arithmetical [Ark70a, Els00, Hau90, Hau00, Lon85, MS90, MR95, SK20, Hol06, PT89]. arithmétique [MS15b]. Arlington [Swa92]. Armstrong [MW92]. Arndt [HT22]. Aroids [DB71]. Arose [Zei89]. Array [AC11, And14, BH76, BJ03, Car63, EH22, HBJ78d, HBJ80, Kim95b, Wal63]. Arrays [All76, Car72c, CR72, Car73, CH74, Chu70, EA12, GP16, Gou74c, HB75a, HB76b, HB76e, HBJ77d, HBJ78a, HBJ82a, Kim91b, LY02, Lon73, Ros67, Sul89, YL04, YG18, ZD02, AB00]. Art [And05, Nay65, BQ03, B^{+}96]. Arthur [And05]. Articles [Ano97f]. artificial [EGS22]. Artist [Hed76]. Ary [GKH77, Cha86c, Hau00, Igu94]. ascending [Lün91a, Lün91b]. Aslam [Sha88a, Sha90]. Aspect [Lyn73]. Aspects [BTD22, FM95, HS82, WH74b, BHS06]. assignment [PC13]. Associate [LPS16]. Associated [AE79, BF91, BK96, Byr63, Cus68, Ehr83, EH22, Fra67, GKH77, HS15, Hal67, HS66, HL67a, Hor69, Hor77b, HP81, Hor84a, Hor86c, Hor93, Hor96d, How80, How84, Hsu87, Kah09, KK11, Kim91d, Kim01b, Kol65, Kom03, Kou94, Lee01b, Mil60, ND87, PP14, PR01, SH79a, Sir97, SŠŠ22b, Sma77, SG85, Swa99a, Swa00b, Tri69, Lee00, Rao05a]. Associates [Weg77]. Association [Ano69a, Ano70a, Ano73d, Ano74b, Ano74c, Ano77b, Ano71b, Ano02c, BJ14, Cat09]. Associativity [Gou64a]. Asymptotes [WS91]. Asymptotic [BMRS01, BZ07, Bro70b, BW81, Cha05, CK88, Hir13a, Hoi22, MS15a, MNPX17, NS90, Ped71, Pro96, Rie98, Sha84b, St.84, TS89a, Cha06b, DP24]. Asymptotically [TRE73]. ATM [SMSM01]. Atom [Hun69]. Atoms [Wlo63, Wlo68b]. Auctions [Gil05, Gil07]. Audio [FM15]. aufsteigende [Lün91a, Lün91b]. Augmented [BN93, GFJT19]. August [B^{+}96, Bro96, Bur92]. Aurifeuille [Mel16d]. Austria [Jef08]. Author [Ano93d]. Autobiography [Gri73b]. Autocorrelation [FF94, RW81]. automata [Sal12]. Automated [LO18]. Automatic [DMSS16, DMR^{+}16, Sha88b]. automaton [Cla08]. Automorphic [Cal72]. autonomous [LO14]. Average [BBG^{+}13, BS96, Coo06]. avoidability [DMR^{+}16]. Avoiding [HZ17, Sed14, Ben06, Zei96]. Away [EF80, Fla82, Mar12b, Pon17a, Sch70, Tri67f].

B [Put76]. Babbage [FZZ08]. Back [HM72]. Back-to-Back [HM72]. Bailey [FS01]. Baillie [CG03]. Baker [Lin96]. Balanced [Hor92d, Hor02d, LF16, DP24, HS22]. Balanced-Power [LF16]. Balances [BPT14]. Balancing [BLPS11, BLP10, Lip04, LPS16, LHWS96, PR13, PP14, PD15, PP17, PP20, PPK21, RP18, RP19, SP18, LJ97, Pan07, Rou13]. Balancing-like [PP14]. Ballot [Car72f]. Balmer [Wlo73]. Band [HML01]. Bands [SBM97]. Bargaining [BFT15, Rac19]. Barnabas [Høy09, Oak09]. Barthélemy [Spi00a, Spi00b]. Bartók [Low71a]. Base [Bal13, CKR98, DE18, Dek20a, Dek21, SS91, EZ04]. Based [BL17a, BBL95, Bom20, BH17, Cam22a, DRT11, FM15, Gou77b, GFJT19, Knu64, SCY^{+}23, Tan98, TF92, BYD^{+}19, EE10, EMG17, EGS22, Len06a, SS95a, SS95b, Spi00a, Spi00b]. basée [Spi00a, Spi00b]. Bases [Arn95, Bro71b, DES17, DH67, DH68, Fri66, GH18, HM72, Lud87, Wit16, ESW07, SR12]. Basic [Hor65, Lee97, Lee01a, Woo75, Woo79]. basis [BSS08, LLP^{+}18]. BBP [Ade14a, Ade14b]. BBP-Type [Ade14a, Ade14b]. be [HW81, LL87a, Sun06a]. Beatty [Dek20b, Gri15b, Kim11, NKW23, RS11b]. Beatty-Related [Gri15b]. beauty [Hun70b]. Beds [Zen78]. bee [EGS22]. Beginner [Tho63a, Tho63b, Tho63c, Tho63e, Tho63g, Tho64b, Tho64c]. behave [Di 13]. Behavior [BDE^{+}14a, BDE^{+}14b, BW81, CCT05, CFH^{+}17, CCG^{+}19, DK18, DDK^{+}14, Gor12a, Hir13a, MS15a, MNPX17, Pro96, Win68, Alt88, BZ07, KS02a, LP14]. Béjaïa [AV03]. Bela [Low71a]. Belated [Hog68c]. Bell [HM83, Car80e, EA76, Fra77, Ges81]. Belong [Adl89, Ful78]. Belongs [Sto77]. Below [BT00]. Bemerkung [Geg93]. benchmarks [KS02a]. Benford [BGTM^{+}11, BDE^{+}14a, Bra78, CHL19, Gil05, Gil07, Gor12a, Pet81, RM11, Sen73, SB93, Was81, Wlo71a]. Benjamin [And05]. Benzenoid [BPGC97, TB91, EBK89]. Bergman [BDD^{+}22]. Bergum [Sch94, Ano21c, Gra89c, Sin90]. Bernoulli [Sla00, AD10, BLM20, BRS04b, BN09, Byr75a, Car68a, Cas91b, Cer21, CH06, Fer73, Fox01, Her78, Her82, How94, JV19, KY09, Kel57, Kel58, KY20, Lin21, Liu01, LL05, Mag09, MV15, Mom01, SP02, Sim98, SR83, Ste86b, Tos78, WSP04, WP07, You06, You17, ZL95, Zie19]. Bernoullianlike [Kun90]. Bernstein [LL85, LB91]. Bessel [DH76, Hin77, How85, How87]. Beta [Dek20a, Har98]. Beta-Expansions [Dek20a, Har98]. Betting [RH82]. Between [AS05, Byr75a, Byr75b, Daf20, Dei66a, Dif12, EA12, ES17, GMP91, Hew77, Hin77, Hir15b, HM72, KKLS18, KP77, Kuh12, Mai70, MS15a, Raa63, RP71, Ste76, Ste21, Str76, WB12, Yoo21, AS00, Bra00, Can51, Car39, CHF85, D’S89, Gin54, JV08, Rem11, Sub19, Van86, Wat86, ZQZ20]. Bezuszka [Moh81]. Bi [BTD22, Gou67b, Hag87a, MT17, CT18]. Bi-Colored [Gou67b]. Bi-orderability [MT17]. Bi-periodic [BTD22, CT18]. Bi-Unitary [Hag87a]. biased [LLP^{+}18]. Bibliography [Bee11]. Biblioteca [Pic83a, Arr69, Pic83b]. Bicknell [Ano10b]. Bicknell-Johnson [Ano10b]. Bienaymé [Hey81]. Bijection [Dif12, KW21, Mun17]. Bijective [BO17, Eli21, GZ22, Woo07]. Binary [Ade14b, AN79, All17, AG78, BS96, Cha86b, Cha90, CDW13, Hor92d, JS19, KR89, LPS16, Min81b, Mos93, NN87, ORSW84, RW81, Rob91b, Rob02b, Tur88a, Boy97, BP07, Gol06, Rou13, Yor14]. Binet [BMRS01, BRS04a, Bro69g, Cas89, ES12e, Giu71, Gri11a, Iak81b, LLKS01, SE20a, SJ84, Whi77a]. Binet-Like [Gri11a]. Binets [Spi82]. Binomial [AK74, AB20, Bai92, BS11, Bos82, Bos85, Bru73, CK24, Car65, Car66a, Car76d, Car81b, CK97, Com79, CK16, Edg86, Els05, Fil92, Fil95, Fra80, Fra98, Fre85, Gal98, GS81, Gau12, GL01, GMP91, Gou64b, Gou67a, Gou69, Gou71b, Gou72a, Gou72b, Gou74a, Gou95, Hau93a, Hau96, Hir15a, Hod74, Hod92, Hog67a, Hol00, How93, KK11, KI10, Kis14, Kom14, Kou90, Kwo89b, Lay77, LL87b, Len17, Lin65, Lin71, LP96, Lon73, LH74, LAL16, Luc02, Man67, MC00, Nak14, Nyb99a, OS05, PG89, PA91, Pla97b, Pro94, Roc81, SW10, She11, Sin73a, Sin75, SSS24, Str73, Sun01a, TF64, Tov85, Tri00a, Wan05, WD76, Wen02, Wes66, Yoo21, GTU13, Hol06]. Binomial-Catalan [Gau12]. Binomial-Truncated [Kou90]. Binomials [Cal00]. Biographies [Wei88]. Biophysics [Zei89]. Bipartite [HK95, LG91, Lee00]. Bipartitie [DD87]. Bipyramids [Ala89]. Biquadratic [BB04]. Bisecting [PS92]. Bisection [CDW13]. Bit [Alt88, VKSAN93]. Bit-wise [Alt88]. Bits [Mos93]. Bivariate [Ess03, IA11, TKS11, YL97]. BK [EB17]. BK-space [EB17]. Blackwell [Chr12]. Blithe [Eve81]. Blocks [BW97, DN96]. Blockwalking [Jon96]. blossoms [Han12]. Board [Bea64, Bea68, TL08, Tha19]. Bode [Gre76, Gre78a, McL79]. Body [DA79]. Bollettino [Cio18]. Bombelli [RS98b]. bonacci [BPT14, JMNRS22, Swa77, AC11, LL87a, PW23, Sir97, Kir22]. Bondarenko [Ano93f, Ano93g, Ano93e]. Bones [Wil83]. Book [Alf66a, And05, Ano80a, Ano88a, Ano90e, Ano93f, Ano93g, Ano93e, Ano93i, Ano16a, Ber11, BJ99, Bic72a, Bic73a, Bro66, Bro67b, Bro68a, Bro69a, Bro70a, Cio18, CG96, Eve68, Gal08, Gau02, Gau10, Gra89c, Hal22, Har09, Hor04, Høy09, Mos08, Sch94, Sim18, Tol17b, Tol17a, Ano88c, FS87, Sig02, Pic79]. Books [Gal08, Hal22, Ano22d]. Boolean [MN76, Wan00]. Borel [SS91]. Boris [Ano93f, Ano93g, Ano93e]. Born [Lev88]. borrow [TLC93]. Boston [Ano88c, Sch94]. Both [Bau71, Kel57, Kel58]. Bound [Cat74a, Col87b, EKS88, Fos66, Hoi22, Jav22, Sto09a, DP24, Kir94, TD16, TD19]. Bounded [BK00, Coo06]. Bounds [Ala98, Boy92, CDH^{+}22, Hab19, Hag84, Kut02, Mar13b, SK13b, SC10a]. Boustrophedon [CBH17]. Box [Tri89]. Bracelet [Tri67b]. Bracelets [Tri69]. Bracket [Gou64b, Gou67a, Gou69, Gou94, Hau97a, Sho99b]. Brahmagupta [RS02, Sur96, Sur98, Swa98]. Brain [Alv69]. Branching [BYD^{+}19, Hey81]. Breast [HST94]. Bridged [Sta17]. Brief [Mit03]. Bring [EF98]. Brother [Ano88f]. Brousseau [Ano88f, Mel03b, OS20]. Brown [KsS73, Röd94]. Bruckman [Ano16b, Ano16c, Ano16f, Ano17j, Ano17g, Ano17h, Ano18c, Ano18d, Ano18e, Ano18f, Ano19e, Ano19f, Ano19g, Ano19h, Ano20c, Ano20d, Ano20e, Ano20f, Ano17i, Car75c, Gou72d, GQ09, SK22b]. Bruxelles [ABN^{+}19]. Bruyn [HT00]. BSTs [Rou13]. Bubbling [Zen78]. Buck [Tor19]. Buddy [CT75, PN77]. Bugia [AV03]. Built [KR22]. Bulgarian [EJ17]. Bunder [LB13, LBF14]. Bunny [Hil21]. Bunny-pocalypse [Hil21]. Bünyamin [Sub07]. Burgstahler [Gre02]. Busy [Bri64a]. Byproducts [Kos15b]. Byrd [Gou68]. Byte [Win11].

C [Ano90e, Har74, LM71, Wes80, DB22]. C. [LM71]. C399 [Goo93a]. C416 [Goo94a]. Caen [Ano17c, Ano18a, Ano17b, Coo16]. Cal [Ano02b]. Calcolo [RS98b]. calculate [Sho78]. Calculating [Wil83]. Calculation [Flo67, Hor04, Man67, Man77, Pet96, SBM97, TF93, FLP08, RS98b, Sig02]. calculations [Sha88a, Sha90]. Calculator [Alf64d, Ano63b]. California [Bro68c]. Calkin [Pau23]. called [Sin85]. calls [Man05a, Rob99]. Camouflage [Ano04a, KC04]. Can [HW81, LL87a]. Cancer [Blu72, HST94]. Candidates [Goo75]. Candido [Cer17]. Canon [Nor76]. Canonical [Sil77a]. Cantor [Nag01, Süt89, Wal90]. Capitulum [Bom20]. Cardinalities [Jos79]. Cardinality [ESW13]. Cardioides [RS98a]. Cards [Bro72b]. Carlitz [CS79, CK95, DeL88, DV89, Gau04, How00, Jen66, LR85, Lee87, Mel03b, Sha74d, SM93, Wal63]. Carlo [NS95, LN07]. Carlson [Sin73b]. Carmichael [Arn95, EV04, Yab01]. Carry [Nod90, VKSAN93, TLC93]. cars [Wir08]. Case [Ata89, BCOR95, Bun75c, CFG19, Gal79, Gou72a, Jag21, Pow77, TR03]. Cases [KH72, TMR02b]. Cassini [Alp18, Alp19a, Bea22a, Hen17, PW15, SR19, Vol10]. Catalan [Hen17, BGW14, Boy92, Ege83, FS01, Gau12, HB76a, HB76b, HB76e, HBJ77d, IL07, Ima98, Kos09, Lan00, Lan02, LOF14, Lar15a, Lar16, LO18, LM19b, MS95b, Oll22, Pro22, Sul89, Tue22, Gau10]. Cayley [Hen14, Mur82a]. cedar [Fra84]. cells [DMP07]. cellular [Cla08]. Centered [ES17]. Centers [Ste78a]. Central [BH78c, Cha81, FJL^{+}20, Hau94b, LM17, LM19a, SH79b, Zha98a]. centuries [Fra03]. century [Arr69, Fre92a, Pep03, Uli11]. Certain [AJ90, AJ91c, AJ91f, AJ97, AJ98, Ark70a, Bas81, BD81, BN77a, BNOS03, BD70, BD72, Bro75c, Bru73, Bru77c, But73, CJ96, Car65, Car72a, Car78b, Cha82, Chu70, Coh82, Col88a, CBH17, CKR98, Cus68, Dil91, DRS11a, Fie67a, Fie73a, FH88, FB94, FH98, Fil00, FP86, Ful77, Gil05, Gir81, Gri74, Gua78a, Gua78b, Gua78c, HPV95, HS84, Hor65, Hor88a, Hor99, How01a, Kar74, KI10, Kla73b, KsS73, KS88, Kun90, LOF14, Lay77, Led67, Lin68, LAL16, MST13, McD94a, McG09, McG10, McN88, McN92, Mel97, Mel99b, Mel99f, MC00, Mel00d, Mel02, Mel04a, Mel10b, Mel10c, Mel15, Mel16f, Mel17a, Mel17d, Mid79, MSS08, Mon90a, NB89, Pop77, Pop84, Ros67, ST90, Say76, Ske65, Sol76, Som91, SK13a, SK17, SK20, Sto67, Sty66]. Certain [Swa97b, TS04, Tur74, WH71, Wal76b, WD76, Zha01b, ZW04b, BZ07, EP07, Gil07, JH05, ST06, Sob17]. Certificates [Cal00]. Certification [Boo79]. Cesàro [FG05, Mel03a, Hin80a]. Chains [CMR02, Dun67a, Far86, Kut02, MR99, MR00, TB91, TS92, BPIK07]. Challenge [BL71]. Chance [Cla86]. change [BV05a, BSS08]. changed [Dev17, Sim18]. Changing [PB15]. Chaos [Sch91]. Chaotic [Fra94, Jos98]. Chapter [Cio18, Ano90e, ST89]. Character [KS10a, Leh66]. Characteristic [Adl71, AH95, BH77a, Bro72d, Bru00, Car65, Eng91, Fer78, HH65, Lev68b, PH75, WP19]. Characteristics [vRWT90]. Characterization [Ava11, BL13a, Blu72, Bro65a, CST05, Cho91, DeL81, ES14e, HS85, HW81, LL87a, May68, McH82, Nor05, IMS97, PZ12]. Characterizations [fC03, MB89, Sch89]. Characterize [Mel97]. Characterizing [KS02a, LF15, Len94a, McN92]. Characters [KS10b]. Charge [SBM97]. Charge-Transfer [SBM97]. Charles [Eis51]. Chebychev [Pet84, Pet85]. Chebyshev [ABD05, Alp20, BWH75, Bos85, Bus63, CR93, FZ01, Gru14, Hor77b, Hor79a, Hor05, Kim79a, Kim80a, Kim80d, KM05, Lan92, MS95e, ZZ06, Zha02, ZW04a]. Checkerboards [And70]. Checks [Jan02]. Cheese [Coo16]. Chemistry [Hos73]. Cheo [KC91a]. Chess [HB72a]. Chessboard [DeF14, Fra70]. Ching [Bic73a, LM71]. choice [LLP^{+}18, Wat86]. Choix [ABN^{+}19]. Cholesky [HF91]. Chord [Pri05]. Chou [LM71]. Choudhry [Bia95].

Chromotology [BJ78a]. Cipher [SCY^{+}23]. Circle [DMY13, vR89]. Circles [Hor84a, Hor84b, HK89, Sch81, Sel83, SHB84, Spo22b]. Circuits [EH22, Sta17]. Circulant [Fra80, Hen14, Lin70, ACOM21, BT12, DGMS14, İpe11, JH15, SC10a, SCH11, Sol05, Sol07, UNS07, ZJ14]. Circulants [Goo86a, Goo86b, Wat88]. Circular [Cha90, Cha91, KL93, Kon95, PP17, Wal76b]. Circularly [Smi68]. Circumscribes [DeT92]. Clarification [Iak81a]. Class [AJ94b, AJ95a, Arp94, Arp04, CY91, Djo98, EI03, Fil91, Gau98, Geo75, Gra89b, HS19a, Hir73, Hor96a, HPL93, Kos24b, KG24, Kyr85, Kyr94, LF14a, LF14b, LF15, LF19, Lin65, Mag81, Mol88, Pla94, Pla95, Rob80, Swa97a, Tao17, Tur86, Vse11, WH04, Whi70e, Zei78, Zel97, Zha98b, ZW02, AK05, AI92, CHL19, DK17, DRS11a, EMG17, Flo18, Igu94, KA03, MZ91, Tra88]. Classes [BJS96, Car78f, Dif12, Gry03, Kos23p, Kos23q, Kou90, Mel04a, Pla97a, SHB84, JV08]. Classic [Gor12b, Mel16h]. Classical [Ark73b, Ark73c, Col88b, Mel16d, AK05, İpe11, Kos19a, Sun05, ZJ14]. classification [CS92]. Classroom [Gan59, Mea56, Tho63d, Zei63]. Clausen [Kun90]. Clavius [Par17]. clock [Dun09]. Close [CC14, Kos16, LS13]. Closed [BL92, Col83, Dan98, JMNRS22, Lar15a, LF17, Mel16a, Mel17a, Mel17b, Mel17c, Mel18d, Mel18b, Mel18c, Mel18e, Mel18f, SP02, KM18]. Clucas [Mos08]. cluster [PC13]. Clusters [LaD18]. Coaxal [Hor84a, Hor84b]. cobalancing [Pan07]. Code [Arr69, HKW15, BP10a, NO15]. Codes [Abr95, Kim10c, KBN08, KBN10, Rit23, AAB^{+}16, BKV22, EMG17, Mol18]. Codex [Pic79, Pic83b]. codice [Pic83b, Arr69, Pic83a]. Coding [Abr97, Abr00, GKH77, MS03, EE10, LLP^{+}18, SMS99]. Coefficient [Bai92, Bal17, BH76, Car66a, GL01, Hen15a, HBJ77d, Hol94, How93, OS05]. Coefficients [AK74, Ard04, Ark69, Ark73a, AB20, Bei78, BP09, BS11, BLL71, Bol86, Bos82, Bos85, Bru77b, Car65, Car76e, Cas86, DMY13, Edg86, Els05, Fah16, Fra98, FZ01, Fre85, GT72, GS81, GMP91, Gou64b, Gou67a, Gou69, Gou71b, Gou72b, Gou95, GQ11, Hod74, Hod92, Hog67a, HA71, Hol00, How01a, ISTY88, JV19, KI10, KPAO11, KW21, KW95, Kin73, Koh66, Kom14, Kos15b, Kwo89b, LL87b, LL83, Len17, Lin71, Liu92, Lon73, LH74, Luc02, MT12, MST13, MC00, NR15, Phi83, Pla97b, PS76, Roc81, Sha79c, She11, Sin73a, Sin75, Str73, Sty66, Tau65, Tee94, TF64, Tov85, Tri00a, Vol94, WD76, Whi86a, Yan88, Yoo21, Zha09, BB19, Bri83a, Bri83b, GTU13, Hol06]. Coin [FW78, Gri11d, Hir12b]. Coin-Tossing [Gri11d]. Coincidence [Gre02]. Coincidences [BL13d, BL13c, BBL14, Dei66a, Lin67a]. Collapse [ES15a]. Collatz [Kan15, Sha06]. Collecting [Ber11]. Collection [LM17, Jar66]. Collections [CR10]. Collision [TCC^{+}04]. Collocation [KYS13, MH16]. colony [EGS22]. Color [CDW13, Sha12, EB87]. Colored [Gou67b, TS89b]. Colorful [DX22]. Colorings [HW09]. colouring [DP24]. Column [Beh12, BH78c, Cal97, Kim94, PS02, PS76]. Column-Finite [Cal97]. Column-Justified [PS02]. Columns [HA73]. Combination [Ozb12, Tep76]. Combinations [Abr71, Abr76, BK00, BCSS11, Chu71, Lar15a, Lar16, LO18, Mel10b, Mel10a, Mun07, Tep74]. Combinatorial [Aga91, And05, And74b, Asc74, BTD22, BP09, BR14, BH78a, Ber78, BG16, Bro68g, Car78e, Car78a, Car78h, Cha84b, Chu85, Cre86, EG16, EA20b, EA20a, ES10e, Fil98a, Gal86, Ger09, Gou67b, Gou72d, HH00, Han76, Hod74, Hog70b, HS82, Hsu73, JV08, Kle91, KL91a, Lan92, LL85, Lew80, Poo88, Rob91b, SO02, Sha13, Tau76, Ted19a, Tri00a, Web95b, Woa02, Zei89, ZG17, BQ03, BCQ07, BP10b, Bra06, EA19, LaG13, de 81b, dLM94]. Combinatoric [KS04]. Combinatorics [DS20, Reb75, RS08]. combined [Kum21]. Combining [Bro69f]. Combs [AE23]. Comment [Kah05, Kah06, Wer83, Sub07]. Comments [Cap89, CG03, GG15, Hei64, KTW84, Win65a, Pic79]. commerce [AV03]. Common [BHL21, Day70, DT71, Edg86, Gou72b, HS15, Har70a, Har70b, Kim79a, Luc73, She69, She73, Dij90, Dij96, Mor95, Str73, YZ17]. Comp [GP93]. compact [DH17]. Companion [Swa00c, TR03]. Comparing [BWWW16]. comparison [ORSW84]. Compasses [Puc62]. Compendium [Spi00a, Spi00b]. Compendy [Spi00a, Spi00b]. Compiling [IK87]. Complement [LN81]. Complementary [AS00, BHK07, Hol66a, Hol66b, Kim81a, KM19, Mor10]. Complements [Wei79]. Complete [AC13, Bee11, Bri92, Bur71, Coh70, GWZ07, HB63, HW81, HW73, LL87a, Li00a, LM79, Nor16, Par98, Ral85, SK17, GCMP14a, Tre96b, Tre96a]. Completeness [BHL^{+}20, Bro63b, Bro78c, Day69a, HC72, Sch89]. Completion [Hor00a]. Complex [AJ91b, Ano04a, BL13a, BL13b, BN03, BN19, CDS03, Goo93b, Goo93c, Har81, Hor63b, KC04, Pla96b, RS02, Sur98]. Complexes [SBM97]. Complexification [Kai22]. Complexity [Bra90, Hay92, Sha78a, BV05b, MSPF95]. Component [CM07, Wal87b]. Components [Iak77a, KST20, LS18a, Wal88]. Composer [Nor72]. Compositae [MC76]. Composite [Bru77b, DJ67, DJ68, Fra87, Hor97a, Izo02, JM14, Ral85, SW02, Šiu11, SK13c, Sun75b, Gri21, Knu90, Som06b]. Composites [CFL^{+}22, HBJ77a]. Composition [Gou64b, HBJ82a, Kim80a, The71, Whi66a, BSS08]. Compositions [Abr71, Abr76, AH75a, AJS22, Car76a, Car79a, Car79b, Cha82, CDW13, Dif12, Fuc84, HZ17, HCG03, Hic78b, HL69a, HB75b, HA77a, HT22, Kim01a, KR10, Mab19, Moh67, Mun20, Mye75, RM04, Rob14, Sha12, Sho06, Sil11, WH14, Win65a, Win65b]. Compound [Col87a, FJL^{+}20, Sug93]. Compression [KBN08, KBN10, SC07, SM08]. Comprising [Lar15a]. Comput [FS14, Sol07, UNS07]. Computation [DGST88, Gon89, Jef08, Lan11, Rok96, Sub07, CCL89, GPS87]. Computational [FZ03, ZW02, BHS06, EB88, Liu07]. Computations [Lin67a]. compute [KŞ13]. Computer [AG87, Gel63, IEE81, IEE84, IEE08, LN81, Rok96, Ola00, SO09]. Computers [Alu97, Mak94, Wol92]. Computing [ACM12, AC87, CH89, Er83a, Er83b, GL80, NS95, Pet80, Tak00, WCL99, WS80]. Comrade [Tue23a]. Concatenations [AFRT23]. Concave [BRGGGS20]. Concavity [AE79, Car72a]. Concepts [Hen67]. Conceptual [PB15]. Concerning [Adl89, And82, BBHM85, But78, Car72b, Chi86, CW85, CS89, Els98, Hin80b, How90, Kel65, KC91a, Kim95a, Kla65, Kui69, MH75a, Mel00b, Mel09, O’D79b, SH82, Shi95, Ste86a, Sto65, SH70, TS89a, Veg74, Jun61, Sun06b]. Conclusion [Zho99]. Concrete [GKP89, GKP94]. Condition [LF15, Sul89, TMR02a]. Conditional [ST18]. Conditions [CR72, Els00, LF14a, Lay75, Li00b, Pla95]. Conference [ACM96, Ano70e, Ano71b, Ano73d, Ano77b, Ano83b, Ano86c, Ano89d, Ano89a, Ano89b, Ano89c, Ano90b, Ano90d, Ano90c, Ano93k, Ano94c, Ano97a, Ano99a, Ano99b, Ano01a, Ano01b, Ano02a, Ano02e, Ano03b, Ano03a, Ano09a, Ano12a, Ano13a, Ano14b, Ano17a, Ano17c, Ano18a, Ano19c, Ano20a, Ano20b, Ano21a, Ano21b, Ano22a, Ano22b, Ano24, BHP92, BPH93, BJ18, BJ19, BJ20, BJ22, Coo10, Coo16, Fre87, Fre88, Fre90, Joh09, Joh12b, Joh20, Joh22b, Ano17b, Phi04, SC07, SM08, Swa92, dBF85, NS95, Ano74c, Ano82b, Ano84b, Ano84c, Ano87b, Ano87c, Ano87d, Ano88d, Ano88e, Ano90a, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano93a, Ano93b, Ano93c, Ano95e, Ano95f, Ano95g, Ano96d, Ano97b, Ano97c]. Conference [Ano97d, Ano98a, Ano00a, Ano00b, Ano00c, Ano10a, Ano11a, Ano11b, Ano11c, Ano17d, Ano18b, Ano19a, Ano19b, Ano19d, Fre92b, Fre94, Fre97, Fre99, Phi01, Phi07b]. Configuration [SCY^{+}23]. Confirmations [Kos19c, Kos21e, Kos21f, Kos23d, Kos23l]. Congruence [Bis84, DeL82, Dob14, Du89, Els96, ES12e, Fox21, FP86, HMT23, HB74a, KW95, Kyr85, MS94, PRW93, Sin90, Som89, Tur74, Zar09, Zho03, FZZ08, SE20b]. Congruences [ASD15, Bai92, BJ95, Che04, CD13, Fox01, Ges81, GMP91, Gic22, Gil92a, HL09, Hir15a, Hir18, HH89, How90, How94, Kyr94, Liu01, Liu05, McI20, MS95c, Met14, Mid79, Mon76b, NS99, Sbu02, SHC74, Sim98, Sla00, Sun02, Wal79, Yam19, You94, You17]. Congruent [DeL84, Hun78, Kom12, Ste11a, Zel97, Pic81, Sun06a]. congruenti [Pic81]. Congruential [McI92]. congruo [Pic81]. congruo-congruenti [Pic81]. congruous [Pic81]. congruous-congruent [Pic81]. Conics [Mel97]. Conifers [Bro69c]. Conjecture [AJ94c, Beh12, BBHM85, Bev70b, Bru94c, But78, Car72b, DV88, FA89, HM94, Hen18, HH72, Hoc79, HJ68b, HsSW95, Kan15, KAP10, Kla18, Sin80, Sin73a, Tay76, SR12, Yab07, Hen19]. Conjectured [Car81b]. Conjectures [BA98, Fin91, GS81, HKN10, Hol11, Kim79b, Kim95a, Mel09, Sha23, Tho65, HKN11]. conjugacy [DK17]. Conjugate [Eng91, Iak81a, Kin68, Mun20]. Conjugated [BPGC97]. Conjugates [RS02]. conjugation [Séé98]. Connected [Vau76, DRS11b, DRS11a]. Connecting [Ste73b]. Connection [BB69, BJ89, BJ93, Bra00, CDM00, GZ19, Hir15b, JV19, KKLS18, KP77, Raa63, Yoo21]. Connections [Hor03a, Hor05, Kuh12, Mar93]. Connectivity [Ord93]. Consecutive [AS77, CM14, CK93, DE18, Gau04, Gio94, Gri11d, Gru94, Gru07, GH18, Guy82, HS15, Hir12b, Jar67a, Kon95, Luc02, Mar12d, Mar13a, aMS22a, Mon76c, Mon85, PP14, PM82, Pri18, RP18, SD18, Shu20, Wal76a, Wil96, Wil97, CM15, HS12]. Consequences [Cas89, Ewe82, MS95b, Rob02d, BV05a]. Considered [DeL70]. Consisting [AH95, Kar70, Kar72]. Constant [BJS96, BT00, BW97, Crî14, GQ11, Hir11a, Liu92, Nak20, NR15, Ton97, WTvR89, BCD^{+}20, Pet99]. Constantly [Bru77a]. Constants [CDM00, PLL81]. Constellation [MH87a]. Constraints [CS81a, AC87]. Constructed [Goo68a, Goo68b, GS74, Kon10, RW21, HS22]. Constructing [Arn95, Rot73, tR84]. Construction [Gro97, Nyb04, Nyb09, Wil96, Wil97, BB79, Man17]. Constructions [Ped71, Puc62, Puc62]. Constructive [Gal72]. contact [Rob78b]. Contain [Aus23, AC13, SK17]. Containing [BJ79a, BJ79b, BG21, Car66b, Car77c, CM17, Gic22, Pro16b, Say76, Tri09]. Contains [Bur71, Zöl93]. contemporary [SO09]. Context [MV91, Moo93b]. Context-Free [MV91]. Context-Sensitive [Moo93b]. Continuants [She72]. Continuation [Nav01]. Continued [Adl78, Alb73, Alf64a, And17, And20, BCOR95, BD11, BCFX09, Bow14, BB65, Bro71a, CK10, Coo06, Dra64, Dra69, Els98, Fil86a, Goo93b, Goo93c, HK91, Her20, HA64, HB77b, Jon91a, Kom01a, Kom12, Len03b, LM10, LJ67, LJ70, Lon81b, MC04, Moo93a, Rat93, Rie93, RT04, Rob95, San23, Sha82, SH88, Sma77, Stă14, Wat92, Wil00, WTvR89, Wul76, Car07, CCL89, Fra84, Kom09, Lün91a, Lün91b]. Continuous [BGTM^{+}11, Sco68a, Süt89]. Contour [GFJT19]. contraction [PRS96]. Contractions [Bow14]. Control [Dei72, McC72]. Controlled [Fie68b]. Conventions [Han11]. Convergence [Ark69, Ark73a, CS80, CS81d, CS82, Hen15a, Hud87, Jor64, LOF14, Len17, TMR02a, Yam80]. Convergent [Ger77, Ger78, Hol66a, Hol66b, MRS00]. Convergents [BCOR95, BCHR91, BB65, VWT85]. Converging [Ava11, Cas86]. Converse [OP18]. Conversion [LN84]. Convex [CD87, CR10, DRS11b, DRS11a]. Convoluted [FZ03]. Convolution [BG16, Hau00, Hog70c, HB72b, HB75d, HB75a, HB76b, HB76e, HBJ77c, HBJ77d, HBJ78a, HBJ78d, HBJ80, Hor00b, Kou90, Kra90, LY02, Liu02, Rob91b, Sta82, Tur88b, TS89b, Wen02, BBB06]. Convolution-Type [Rob91b]. Convolutions [BGW14, Chu90, CZ10, CH83, Dor93, Gue78, Hor02a, Hor02b, Hor03b, KY20, KG18b, PG89]. Convolved [BH76, BH77b, Cor89, Gou78a, HA77b, Ram14, YZ17]. Convolving [Gau04]. Conway [Bra90, CVZ17, Nor20]. Cooper [Pro97]. Coordinates [Dda20, FL18, HP86, JK93, KLM^{+}19]. Copies [Gue90]. Core [Rob00]. Corner [Bro67e, Bro68d, Tho63a, Tho63b, Tho63c, Tho63e, Tho63g, Tho64b, Tho64c]. Corollary [Ste85]. Corput [Pil12]. Corrected [Blo64]. correcting [EMG17]. Correction [Ano65b, Ano93h, Knu66b, Sta00, EE10]. Corrections [Ano64a, Ano64b, Ano65a, Ano65c, Ano65d, Ano67a, Ano69b, Ano70b, Ano70c, Ano72b, Ano80b, Hor61a, Sin73b]. Correlation [SBM97]. Correspondences [Mai70]. Corresponding [Rit23, Rus12]. Corrigenda [Ano72c, NW83]. Corrigendum [Ano99c, AKP20, BL13d, CH74, FS14, GP93, Hol66a, MS00b, Zar98]. cosecant [GB06]. Cosine [Mel17c, Mel18c, Mel18f, Mel18g, Mel18a, mJ10]. Cospan [BK00]. cost [ORSW84]. Costs [Abr97, Abr00]. Cotangent [Jam02]. cotangents [Goo93a]. Count [And05, BPGC97, Sla77, BQ03, BPIK07]. Counted [BJF99]. Counterexample [Lin67c]. Counterexamples [CPH70, Sin70]. Counting [AH97, BH14a, BLM20, BE22, Bro75c, BP07, CM01, Chu21c, Def13, EKS88, Far86, Fie73a, Fie73b, Hah72, Hoi22, Kah80, Kah87, Kim02, Lor73, LS13, Raw87, Ros72, Sch17, Sch22, Wu83, ZS85, Flo18, Len06a, Nor19]. Counts [BBS81, Fie74, NC85]. Coupled [Hir06a, Hir06b]. Couples [Abr97, Pla96a]. course [Bur92]. Cousins [AMS^{+}22b, Smi96]. Cover [Ano63a, Ano14a]. Covering [AH95, CVZ17, Vor09]. Covolutions [Hau94a]. CPU2000 [KS02a]. Create [GLMS22]. Crisis [Cha73]. Criteria [Rie69, SK13a, BLS75]. Criterion [Bol89, Gou72c, Gou74c, GG85, Gou89, Ste86b, Sun06a]. Critically [Hor02d]. Cross [Bri64a, EG16, EG17, LF14a, Pul94]. Cross-Family [LF14a]. Cross-Jump [Pul94]. Crossword [Gou66a, Gou66b]. Crostic [BJ78b, Bic71a]. Crowds [Sil98]. crypto [Mül06]. Cryptography [Smi94, SMS99]. cryptosystem [SS95a, SS95b, XLP99]. Cryptosystems [BBL95]. Crystal [Swa92]. CS1 [MSPF95]. Cubature [Gir83, Squ81, Tem93]. Cube [Ark73b, Ark73c, AS76, AS81a, DeL95, Hen82, Kis14, Car39, RS98b]. Cubed [AE22, EA20a]. Cubes [AH81, Ark73c, AS74, Ark74, ESS20, HPL93, Hsu93c, JHK97, Ker82, KŽ05, Kos19c, LW81, LM85, Mel03c, MCS01, Ozb12, QW96, Sca99, TV07, Ves15, Wu97, AAB^{+}16, CKMR11, ESS21, Gin54, Mol18]. Cubic [ABCM12, Cha00, DB67, Eli01, GT03, KS10a, KS10b, Now86, RW21, Miu02]. cubica [Car39]. Cubiche [RS98b]. Cuboid [Col88b, Col94, Hun64c, Wal85a]. Cuboids [BH69b, Col88a]. culture [Høy05]. Cumulative [CDM00]. Curiosa [Tri67a]. Curiosity [Ban76a, Bro72h, Bro73c, Hin78, Mel04b, Sch50]. Curiosum [Blo53a]. Curious [AL19, Bic68a, Bro68a, BAP73, KL87, Wlo68a, DG67]. curve [DRS11a]. Curves [HS88, Hor88b, JQ97, Kum74, LY13, RR14, CDM^{+}06]. Cut [Goo91]. Cut-Grow [Goo91]. Cuts [Erc72, TL08]. Cycle [BBP^{+}85, Chr84, Kan15, Sha11, Sha79d, WP19]. Cycles [Ava13, CST05, Dul20, Ehr94, Pag74, ESS21]. Cyclic [BW97, DES17, Daz02, Goo74b, HZ17, JK94, Kah87, KR10, Moh81, Mos93, SR12]. Cyclomatic [Fuc84]. Cyclotomic [Bei78, Kim80b, Kos19a]. Cyclotomy [Dil91]. Cyclotomy-Generated [Dil91].

D [Bri83a, Bri83b]. D. [Har74]. Daisies [Swi04]. dal [Pic83b, Pic83a]. dans [MS15b, Spi00a, Spi00b]. Darwin [PKS18]. Data [Boy98, SC07, SM08, Car39, Kum21]. Databases [AGO95]. date [CHF85]. dati [Car39]. Davenport [Jon78]. David [HS75, Kor94, LA92, LSA07, LK10]. Davis [Hal22, Izo99]. Davison [Bow88]. Daykin [Har74]. days [Dun09]. DCC [SC07, SM08]. Debate [KTW84].

Decadic [PLL81]. decay [ET99]. December [Swa92]. Decimal [Ber66, HW81, Hud87, Köh85, Lin84, Lon81a, Tri69]. Decimals [BJ89, BJ93, Goo74b, Lin90, Wal90]. Decimation [Gol95, Gol06]. Decision [BS96, DMR^{+}16, DMSS16]. Decoding [Kle07]. Decomposing [Jar68b]. Decomposition [FH98, HS99, Pro22, TMR06]. Decompositions [BBG^{+}13, BDE^{+}14a, BDE^{+}14b, CFH^{+}16, CFH^{+}17, CCG^{+}19, Chu21a, GCK^{+}19, KKMW11, LY02, LM17, LM19a, MNPX17, YL04]. Decreased [HBJ77a]. decreasing [BKV22]. Dedekind [Car84, DM10, Mey05, Rob04a, San23, WY00, ZY02, ZW04c]. Dedication [Ano96a, Ano02b]. Deductive [Bro96, Bro96]. Dee [Mos08]. Defect [SK21]. Deficient [AP21]. Defined [AH95, And95b, BHM92, Bro97, Hal83, Hen15a, Hol66a, Hol66b, Kim10c, Mil75b, Moo93a, Nag84, Swa66b, Zei65, Zei72, GL80]. Defining [SCJ93]. Definition [Bun81, Gri12]. Degeneracy [HM83, LM79, McQ76]. Degenerate [KY20, SC04]. Degree [Bru00, Bur90, Coo19, HT86, Kir77, LF19, Rus12]. Degrees [BYD^{+}19]. del [Arr69]. Delannoy [EG17]. Deleting [Nyb23]. Delight [KG17a, KG17b, KG18a]. della [Arr69, Pic83b, Pic83a]. DeMorgan [BG76]. Denominator [Mel16a]. Dense [KBN08]. Densities [BA98]. Density [Dun69b, Gua78a, Gua78b, Gua78c, PB89, She68, Ste73a, Str76]. Department [Ano63f, Ano63g]. Dependent [BGTM^{+}11, Fla82]. Derangement [Abe18, BO17, Eli21]. Derangements [Car78d, Ula23]. Derivation [Gau89, Kha81, Pet80]. Derivations [Bed13, Tre16a]. Derivative [Boy07, Djo00a, FH93, FH94, HSF94, HF97, HF01, Wan95, Zho96, Djo05a, PG13]. Derivatives [Djo01b, Fra87, Lan02, Sta82, YL97]. Derive [GBC16]. Derived [Ahu69, AE22, Bru72, Eli01, Mel16b, Ozb12, TJ88, Tri74a, EB17, Ogu20]. Described [FL68]. describes [Boy97]. Descriptions [Sha82]. Design [Bea68, Bot82, Gou74a, Pre68, SCY^{+}23, SMSM01, Bro96]. Desk [Ano63b]. Desnanot [TT16]. detection [EE10]. Determinant [BFMR20, CFL^{+}22, Fra80, Gal68a, Gal68b, Lin71, Mea56, Pro22, Sha13, Spi06, Wal87a, Zei70, Chu10, Fen11a]. Determinantal [Hor86a]. Determinants [Alf64i, Alf65c, AJ91d, BL92, BJS96, Bic72b, BH73b, BT12, CDNN02, CN04, Car66b, fC03, Con76, Fos66, Gra89b, HB73b, Jai69, Kla66a, Mac10, Pro16b, TT16, Yan91, BCQ07, BP10b, IA14, Kar05, Kwo07, SCH11]. Determination [Bac66, Bas81, BH69a, Car70c, CY91, Ewe01, McD11, PLL81, Sha94, Sin73b, WJ76, Fre92a]. Determine [Kim91c, SK13a]. Determining [Col83, Eng01, Rei93a]. Deterministic [LSS12, LLM^{+}20]. Dethroning [Kel08]. Deutsch [Pro21]. developed [THKH23]. Development [Bro78a, LaG13]. Developments [She72]. Deviations [PP14]. Devlin [Tol17b, Tol17a, Sim18]. Diagnosis [HST94]. Diagonal [BJ81, BJ93, HB69, HB74b, Hor78a, Hor79a, Hor80, Hor03b, MH90, SP77, Sul89, Swa99a, Swa00b, Yoo21]. Diagonalization [LP96]. Diagonals [CH94, Gri11c, LF15, Lar15b]. Diagrams [BS96, Ehr94, HSG63, WSH74]. diameter [EG89, QM90]. Diatomic [BJ03, GG79a, GG79b, GG79c, Nor14]. Diatoms [Bro77, Par69]. Dickson [FMH94, Fil97b, SK22a, SK22b, You02]. Differ [Tri10]. Difference [Bab76, Bro67f, Bus63, Bus70, But78, CC03, Car64b, DDK^{+}14, FG05, Giu72, Gre09, Hsu95, Jar68b, Jen66, Jen71, Kla68a, Kla73a, KYS13, Lay66, LHLT95, Mad69, McD82, Moh81, MKL^{+}23, Nyb01b, Nyb02, She72, SCJ93, TS04, Weg75, Whi70e, BB19, DH17, mJ10, LO14, RT72]. Difference-Operational [Hsu95]. Differences [BK00, Bea03, Bro78a, Coh90, Ehr90, EA12, Gir81, Hir86, Kim79a, Kos15a, Kos18a, Sta82, Van86]. Different [BYD^{+}19, Des78a, Gar04, Hir14a, Jar67a, Kar72]. Differentiable [Fer73]. Differential [AJ95a, Asv89, Gry03, Hog64e, KYS13, SSS24, mJ10]. Difficult [MT02]. Difficult-to-Pass [MT02]. Digit [Alf64b, Bal13, Dei66a, Dun69a, Gor12a, GTNP96, Gri75, Gri10, KC86, Kun87, Lin67a, Lin75, ML84, Puc01, Tri65, Tri71, Whi78e, You93, BH05]. Digital [Arc84, Bro94, KC91a, KC93, Tri67b, Tri69, SS95a, SS95b]. Digits [BS97, Dav76, DG98, FM95, Hei64, Her73, Jar63b, JM14, Kno81, KG19, Puc02b, Rie98, Shu20, Sli98, Ter96b, Wag81, Web75, BHS11, CV86, LS14b]. Digraph [BHM92]. Digraphs [GKRS01, LMR96, SK12, Wil98d, ZS85, Sug93]. Dimension [Ani95, Rei93a, Hol06]. Dimensional [Bai83, EH22, FJL^{+}20, Gou81, Jon96, Kis14, McQ76, Sch85, AB00, BDL^{+}16, BP07]. Dimensions [BCH12, BYD^{+}19, CD87, GM11, Jag21, Ran66]. Diminnie [KG17a, KG17b, KG18a]. Dines [Kla18]. Dinner [Lew95]. Diophantine [AFT22, Adl68, Adl69b, AL79b, And82, Ant85b, Arp94, Arp04, Ber68, BT00, But73, CDL^{+}10, Cha00, CM14, CM15, Coh70, DB01, Drm93, Els00, Eva77, FKL70, Gre68, Hab19, Hog71a, Izo99, Jon75, Jon76b, Kis88, Lin68, LT07, LM19b, Luo96, McD95, Mel16h, MS79, NS01, PP20, PH75, PR01, PR02, Rao02, Rao03, Rao05a, Rao05b, RB22, RHT18, Sha78c, Sin89, Sob17, TU20, TD22, Ton97, Udr98, Utz77, Wal76a, Wiȩ80, Zel92, ZG14]. Diophantus [AHS79, AHS80, Duj96, Duj99, Duj02, Jon78]. Direct [ES16e, Flo67, Gup76]. Directed [And02, Bro97, EG89, QM90]. Direction [Ata95]. Dirichlet [CK10, Dav82, Kom09, Nav01, Pow77]. Disc [MC76]. Disc-Florets [MC76]. discoveries [HD19]. Discovering [HS19b, Laf71, Lon86a]. Discovery [TR90, Wol99, Ola00, WK09]. discrepancies [TF93]. Discrepancy [KS88, Pil12, TF92]. Discrete [BVB95, GR93, HLW94, ND87, SS95a, SS95b]. Discriminant [Lef82, Li21, Luc21e]. Discriminator [MR95]. Discriminators [HS19a]. Discussion [Fie73b]. Disease [DS96]. Dishon [Pro04]. Disjoint [AH95, CR10, Vor09, Zöl93]. Dissection [HJ68a]. Distance [KS02b]. Distances [Gle81]. Distinct [BJF99, BJF01, Bro69i, DeF14, Fer65, Fil86b, Hic73, Jar68a, Kla66b, Kla68b, Lu70, Szi17]. Distributed [Alu97, CV88, HCD97, SK15a, SK16]. Distribution [Ahu69, Ant91, BL93, BBG^{+}13, BD70, BD72, Bru86b, BB11, CJS94, Cav77b, Cav77a, DL11, Ess03, Fra98, GPS13, Geo75, Hen78a, Hen95, Hin92, Jac89, Jac90, Jac92, JC22, KC86, Kui69, KsS72, Mon85, Nag84, Nie72, Phi14, Puc02a, SC05, Som91, Ste05, WP19, Web75, GP91, GP93, Hly17, LLP^{+}18, TP86]. Distributional [BRS04a]. Distributions [BVB95, Dun67b, GCK^{+}19, Hor88d, Mit03, ND87, PG89, PA91, Pro95, XPP87]. Distributive [HH85]. Divergent [Nyb23, SC01]. Divide [Cal97]. Dividends [Kos18b, KG18b, Kos19d]. Divides [CS89, ES16e, Luc03d, SK13a]. Dividing [Boa01, Du00, How01a, Lin98]. divine [Hun70b]. Divisibility [AJ91a, AZ20, Bal15, Bal17, BJ90, Bru79, BF95, Erd81, ES16e, FH20, FZ01, Gou74c, Hal66, HB74a, HL74, HBJ77b, HBJ78b, HLS79, HS85, Hor88c, Hor92c, Kim76, Kim78, Kim79b, Kim80c, Kos19a, Lás82, Len94b, Len03a, Li21, Lon81d, MST13, McN87, McN88, Mon90a, Nor05, OP14, OP18, PTW13, PPK21, Pla94, Pla95, Rob82a, Rob02b, Ros83, SE13, ST18, Sch13a, SY84, Sin80, Sin90, Sol76, Som80, Van12, Vos98, WP69, Wei66a, Yua02, Zho97]. Divisible [Jar67a, Vol94]. Division [Bun81, FF99b, May87, CHF85, de 95]. Divisions [Har70b, She69, She73]. Divisor [Day70, Gou72b, Har70a, Har70b, Hoc74, Mil75b, She69, She73, Str73, Dij90, Dij96, Ewe07a]. Divisors [Adl89, Cat74b, DJ68, DT71, EV04, Izo05, Jar63a, Jar68a, Kim79a, LW87, Lu70, McN92, Mon88, Som79, Win11, Yab01, Ewe07b, Mor95, Vou95]. Dmitri [Bev70b]. Do [Ste11a, BDD^{+}17, JRdlR08]. D’Ocagne [MS95a, Mel00b]. Does [Cal97, PKS18, Rob99]. Domains [Gri12]. Domination [BCJ^{+}05, Har94, CKMR11]. Domino [BHT10, Tha19, Wu83]. Dominoes [Rea80]. Donald [Kla18]. Dordrecht [Mos08, Sch94]. Double [BJ78b, CS81b, Ess03, Kim10a]. Doubling [Ehr94, Phi07a]. Doubly [AS73, Kim10a]. Dov [Bro66, Bro67b]. Down [CP09]. Downs [Fif70]. Drainage [Sha72b]. Drawing [Bea64, Bea68, Sug93, Tre96b, Tre96a]. Drobot [Som02]. Drops [HCG03]. Dual [GG85, Gou89, HH00, TR90, WB12]. Duality [AS05]. Duals [Chu71, ZT88]. Ducci [Ava11, Ava12, Ava13, BZ07, BM07a, CST05, CZZ11, CZZ14, Ehr90, GS95, LY99, Mak12, Sch97, Won82, You98]. Ducci-Processes [Sch97, You98]. Ducci-Sequences [GS95, LY99]. due [Car39]. Duke [LM71]. Duncan [Kui69, KsS73]. Dunlap [BJ99, Cad16]. DVD [BJ09]. Dying [Alf63b, Fen11b, HL69b]. Dynamic [Hir73, HRR03, LHWS96, LJ97, McI82, Sir97, LSS12]. dynamic-multithreading [LSS12]. Dynamical [PW01, DK17]. Dynamics [Bra95, HSR97, JK97, Rei04].

e-book [Mos08]. E. [Mac88]. E.75.P. [Pic83b]. Each [Ewe86b, Gou71a, Sto77]. Earliest [Lar78]. early [Hug10, THKH23]. Easy [McD11, SK13a, SC84]. eBay [Gil07, Gil05]. Economics [Fau64, Fau65]. Ed [Høy09, Cio18, Mos08]. EDA [SBM97]. Edge [WB12]. Edge-Length [WB12]. Edited [Gla70, Ped71]. Edition [Fre22]. Editor [Abe75, Ano70f, Ano73e, Ano75a, Ano76a, Ano76b, Ano76c, Ano77c, Ano91e, Ano97e, Ano97f, Ano98d, Bau64, Ber85, Ber86, Bev70a, Bev77, BJ86, Bro63a, Coh64a, Coh73, Dre80, Erd74, FR66, Gou76, HL84, Hal65b, HF86, Hil87a, Hin97, HL90, Hsu93a, Jam77, Knu74, Kui76, McC86, Nao65, Rob82b, Rus80, Sha89, Sti00, Wei66b, Zei71b, Zei73, Zie63, Bri83a, Bri83b, Jun61]. Editorial [Ano94d, Bro67c]. Editors [Ano73f, Ano00e, Tri67d]. Edouard [Kim98a, KTW84, Wil98a]. eds [Sch94]. Education [PB15]. Educational [Lew70]. Effective [Wil87]. Effectively [Gry03]. effectiveness [Bur92]. effects [Sha88a, Sha90]. Efficiency [Ove73]. Efficient [AUFP13, LLP^{+}18, AM15]. Efficiently [Duv91]. Egg [Lev84]. Egypt [RT02]. Egyptian [DG66, Pih01, Pih10]. Egyptians [LaB71]. Eigenshape [Goo91]. Eigenvalues [Fer78, LKL02]. Eigenvectors [CP03, MC00]. Eight [AS76, Bro70b, Hor75]. eighteenth [Pep03]. Eighth [Fre99, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a]. Eigth [Ano71c]. Either [FF89]. Elchatayn [Mar74]. Electric [EH22, Zuk76]. Electrical [Fer97, Wlo71b]. elegant [Lün91a, Lün91b]. elegantes [Lün91a, Lün91b]. Elemental [Ral85]. Elementary [Ano91e, Ano00e, Bas63a, Bas63b, Bas63c, Bas63d, ES00a, ES00b, ES01a, ES01b, ES01c, ES01d, ES02a, ES02b, ES02c, ES02d, ES03a, ES03b, ES03c, ES03d, ES04a, ES04b, ES04c, ES04d, ES05a, ES05b, ES05c, ES05d, ES06a, ES06b, ES06c, ES06d, ES07a, ES07b, ES07c, ES07d, ES09a, ES09b, ES09c, ES09d, ES10a, ES10b, ES10c, ES10d, ES11a, ES11b, ES11c, ES11d, ES12d, ES12a, ES12b, ES12c, ES13d, ES13a, ES13b, ES13c, ES14d, ES14a, ES14b, ES14c, ES15b, ES15c, ES15d, ES15e, ES16a, ES16b, ES16c, ES16d, Ewe03, Hil64a, Hil64b, Hil64c, Hil64d, Hil65a, Hil65b, Hil65c, Hil65d, Hil66a, Hil66b, Hil66c, Hil66d, Hil67a, Hil67b, Hil67c, Hil67d, Hil68a, Hil68b, Hil68c, Hil68d, Hil69a, Hil69b, Hil69c, Hil69d, Hil70a, Hil70b, Hil70c, Hil70d, Hil71a]. Elementary [Hil71b, Hil71c, Hil71d, Hil72a, Hil72b, Hil72c, Hil72d, Hil73a, Hil73b, Hil73c, Hil73d, Hil74a, Hil74b, Hil74c, Hil74d, Hil75a, Hil75b, Hil75c, Hil75d, Hil76c, Hil76d, Hil76e, Hil76f, Hil77b, Hil77c, Hil77d, Hil77e, Hil78a, Hil78b, Hil78c, Hil78d, Hil79a, Hil79b, Hil79c, Hil79d, Hil80a, Hil80b, Hil80c, Hil80d, Hil81a, Hil81b, Hil81c, Hil81d, Hil82a, Hil82b, Hil82c, Hil82d, Hil83a, Hil83b, HP83, HPW83, HPW84a, HPW84b, HPW84c, HP84, HPW85a, HPW85b, HPW85c, HPW85d, HPW86a, HPW86b, HPW86c, HPW86d, HPW87a, HPW87b, HPW87c, Hil87b, Hil88a, Hil88b, Hil88c, Hil88d, Hil89a, Hil89b, Hil89c, Hil89d, Hil90a, Hil90b, Hil90c, Hil90d, Hil91a, Hil91b, HS13, Kno92a, Kwo17a, Kwo17b, Kwo17c, Kwo17d, Kwo18a, Kwo18b, Kwo18c, Kwo18d, Kwo19a]. Elementary [Kwo19b, Kwo19c, Kwo19d, Kwo20a, Kwo20b, Kwo20c, Kwo20d, Kwo21a, Kwo21b, Kwo21c, Kwo21d, Kwo22d, Kwo22a, Kwo22b, Kwo22c, Kwo23a, Kwo23b, Kwo23c, Kwo23d, Kwo24, Mea65, Nyb03, Rab91a, Rab91b, Rab92a, Rab92b, Rab92c, Rab92d, Rab93a, Rab93b, Rab93c, Rab93d, Rab94a, Rab94b, Rab94c, Rab94d, Rab95a, Rab95b, Rab95c, Rab95d, Rab96b, Rab96c, Rab96d, Rab96e, Rab97a, Rab97b, Rab97c, Rab97d, Rab98a, Rab98b, Rab98c, Rab98d, Rab99c, Rab99d, Rab99e, Rab99f, Rab00a, Rab00b, Spe77, Sun73, Wlo67, MM06]. Elements [AJ91d, Ful78, KK01, Kla66b, Kon95, LM10, Lin75, Mor77, MT17, Pih92, Sha79b, Wlo71d, Wul76, Zei70, Fol04]. Elephant [PKS18]. Eleventh [Ano03b, Phi04]. Ellipse [Hun74a, Mon76a]. Ellipses [RS98a]. Elliptic [Bru77c, Hor88a, JQ97, LY13, LS14a, Mel99b, CDM^{+}06, GP91, GP93]. Elliptical [Pri05]. Embedded [Hen20]. Embedding [HP92, JHK97, Sun75a, Sun78, Ves15]. Encoded [Kle07]. Encoding [AH15]. Encounter [Des94]. Encountered [Fis76]. Encryption [Smi93]. End [Bro73b, MH75b, Wlo71d, BH05]. End-Labeled [MH75b]. Endings [Hun66b]. Engel [Sha93]. Engel-Type [Sha93]. English [Hor04, Mos08, Sig02]. Enhanced [QW96]. Enrico [Cio18]. Entire [Goo86c]. entre [CHF85]. Entries [KK05, SS17, TT16, İpe11, Kar05, Kwo07, MSS08, ZJ14]. Entropy [Hor82c, Hor88d]. Entry [AL79a, Eng91, HK78]. Enumerating [DeF14]. Enumeration [Car72c, Car73, CH74, Car78b, Car78c, Car81a, Cha82, Chu70, CP09, DMSS16, Fie64, Fie66, Kim01a, Lig73, Lig79, MH75b, Ros67]. Enumerator [Ewe86a, Ewe01]. Equal [Ber89, Gou71b, LaD18, Lan90, LH74, Nyb12, Weg75]. Equal-Sum-And-Product [Nyb12]. equalities [Gri13, Hen06]. Equality [Des78a, LF14a]. Equals [Kim95b]. Equation [Alb81, AL79b, AH95, BCM10, Ber74, Ber68, BG91, But78, But73, CDL^{+}10, Car64b, Cha00, CM14, Cor89, DeL76, DeL81, DB66, DB67, EES82, Eme66, Eme69, ES14e, Hab19, Hog64e, HT86, Izo99, Jen66, Jon90, KST20, Kis81, Kis88, Kla68a, KS11, Lin68, LS18a, Luo96, Mak88, Mel16e, Mel16h, MK87, MS79, Per16, PH75, Rob80, Sin89, Sza02, TS04, TS00, Utz77, Weg81a, Zel92, AL20, BB19, CM15, mJ10, KKLL21, Kim07, LT07, RB22, TU20, TD22, Mil91]. Equations [Adl68, Adl69b, And82, AJ92, Ant85b, Arp94, Arp04, Asv89, Bab76, BY16, Bre74b, Bus63, CC03, CM01, Dda20, DB67, DB70, Drm93, EF98, EM15, Eva77, FL18, FKL70, Giu72, Gre68, Gre09, Gry03, Hog71a, Jen71, KM19, Kla68a, Kla73a, KYS13, Lay66, Luc00, LM19b, NS01, PP20, PR01, PR02, Rao02, Rao03, Rao05b, She72, SH80b, Swe64, Tan98, Tsu21, Udr98, Wal76a, Whi70e, KLM^{+}19, KSSS16, LO14, MH16, Miu02, RT72, Rao05a, Sob17, Wiȩ80]. equiareal [Rai54]. equidistant [BM07b]. Equiprobability [EV83]. Equivalence [AH97, Gou75, Gou89, Lee92, Veg74, RS08, ZQZ20]. Equivalent [Dun67a, Pet81, WTvR87]. Equivalents [Ris82]. Eratosthenes [Led68, Sai17]. Errata [Ano68a, Ano68b, Ano69c, Ano70d, Ano71a, Ano72d, Ano72e, Ano73a, Ano73b, Ano73c, Ano74a, Ano79a, Ano82c, Ano04a, Bal19a, BP10b, CS01, Gal68a, ZW01a, ZW06, Ano88c]. Erratum [Goo93c, Krá72a, OS21, Sol07, UNS07]. error [EE10, EMG17]. Escalator [Gru09]. especially [Sal12]. Espresso [SCY^{+}23]. Essays [KTW84]. Estimating [Hir12a, Hir13b]. Estimation [Rie98]. estrazione [Car39]. Etc [Coh64b]. Euclid [DS19, Epa85, Fol04, Hor66, JO24, Moo89, SO09]. Euclidean [Bun81, Cha01, Dun66, HP86, Kel65, Kno92a, MM73, Moo92, Ped71, Rie76]. Euler [AM75b, AML85, Ark70b, AHS79, AHS80, BLM20, Ben02, BD11, BRS04b, Byr75b, Cak91, CH06, Che04, CJCK17, Crî14, DFLT14, DL11, Duj99, Duj02, Ehr84, Fox01, GL83, HL09, HK78, Hir11a, Hir15a, JV19, Lin21, Liu01, Liu05, LS14b, McI20, Mel16e, Met14, MR95, SP19, Sho99a, SR83, TZ96, Tos78, Tót87, TS89a, WSP04, Zha98a]. Eulerian [ASV71, Ark73b, CS75a, CH78, Car78g, CH02, Gou78a, Hen82, Kou94, NS99, Rob98, Rui17, WH03]. EUR144.00 [Mos08]. Europe [Par17]. EusLisp [Mat90]. Evaluated [MS95f]. Evaluating [Sty66]. Evaluation [Bru77c, Fil00, Gou78a, Gri74, Lar16, LO18, SSS24, SC84, WD76, Wit72, Ewe07a, Ron91, SMSM01]. Evaluations [Lar15a]. Even [AM75a, AJ92, AJ96, Beh12, Di 93, Gri75, Hic78a, LY99, MW22, Pri68, RT04, Rob02c, Rot73, RZ95, GB06]. Even-Oddness [Gri75]. Eventually [AC11]. Every [AC11, Hin78, McD82, Zöl93, Boy97]. evolution [Cla08]. Ewell [Rob02a]. Exact [OP18, PTW13, PPK21, Wal85b, FS99, FS14]. Exactly [AP21, Pon19, Sto77]. Examination [Woo97]. Example [Low71b]. Exceptional [Nyb12]. exceptionally [Boy97]. Exclusion [Kon10]. execution [Wat86]. Existence [AHS87, AJ96, CS03, Des78b, DJ68, Jar67a, Ker82, Lev68a, Li00b, MH75a, McD87a, NR15, Pla96a, Wun63, AAB^{+}16, Mol18]. expanding [Rob78b]. Expansion [Ahu69, Bre74a, Byr63, Byr65, DeL95, Des86, Dra64, Gla78, Gri11a, Hil76a, Hil76b, Hil74e, ISTY88, Kra90, Lon81a, Sha74c, Sma77, Smi78, VPB99]. Expansions [Bro75c, CK10, Cas86, Cha81, CS81b, CH83, CKR98, Dek20a, Fre73, GQ09, GTNP96, Har98, Hud87, KM95, Köh85, LOF14, Leu23, Mal83, PVB98, SS01, Sha84a, Sha86, Sha93, Sha94, Ste05, Sun06b, Ter96a, Van96, WH70, Wul76, Kom09]. Expected [Cla86]. Experience [Fre87]. Experimental [SBM97]. experimentation [Rol03]. Explicit [BH69a, BR82, CH02, Hab19, How86, JH15, Sch20, Sha82, Sha74b, Boy07]. Exploration [SCY^{+}23, Sto17]. Exploring [Alf63c, Alf63d, Alf63f, Alf63e, Alf64c, Alf64d, Alf64e, Alf64f, Alf65a, Alf65d, Alf66b, Bro65b, Laf64, Pec65, RR14, RSHL07, Ryd96, TE81]. Explosion [Bro67e, Bro68d]. exponent [KA03]. Exponential [AS73, CB73, CH83, HS19a, Hop95, KC86, Kyr85, Lin75, MH87b, EG89, TR03, TMR06]. exponentially [MM06]. Exponentials [BR85, DH76, Fra91, Gou63a]. Exponentiated [Lar15a, Lar16, LO18]. Exponentiation [CS80, CS81d, CS82, Sha11, Ste85, Ste88, Smi94]. Exponentiations [Bun14]. Exponents [Adl89, BH05]. Expressing [Cha09]. Expression [Fer66, TB91]. Expressions [Alf65e, AESS04, Bre74b, Fil98a, Gri11d, KW13, Mel18c, Sha74b, Gri09, Gri21]. Extended [BVB95, BB90, ES15f, ES16e, FMH94, Iak81b, Kos19d, Kos20, KG20, Kos21a, Kos21b, Kos22a, Kos22b, Kyr93, Lin67a, Sca99, SR83, Tsu21, Wu97, WSH74, dBL83, BL86]. Extendibility [Fil15, Kih00, MB89, ZG14]. Extending [BCH12, BRS04b, BCD^{+}20, Edg16, Gri12, Mon74]. Extensible [MBBB21]. Extension [Ark65a, AH70, Ark70b, Bra96, CG67, Duj99, Duj02, ES15f, Gou72c, Hor79b, Hor96a, Jos98, KC91a, LK10, MOP71, RW21, TMR02a, Wie88, Wul81, deB74, KC63, LSA07, SB93]. Extensions [And17, BH80b, Boh79, Che12, CS79, CS81c, CW85, FA89, FZZ08, Gou95, HCG03, Hor80, How90, JK98, Kor94, Kos14b, KG17a, KG17b, Kos17, KG18a, Kos18b, Sco68a, Stă14, Whi66b, Kro68, OS05]. Extraction [fC95a, CY00, HM94, Car39, CY07]. Extraordinary [Gri17b]. Extremal [BC77, Erd81, Lar88]. Extremalities [All17]. Extreme [Ahu69, CHF85]. extremely [Krá72a, Krá72b]. Extremum [MK02]. eyes [Eis51].

F [Gou68, Gra89c, Sch94, Sha87, Zho03, ZH03]. F-L [Zho03, ZH03]. Fabulous [Gal08, PL07, Har09]. Fact [Bol77]. facteurs [Ber96]. Factor [BC10, Yor14, Zha01a, BL13e, RL15]. Factor-set [Yor14]. Factorable [Con76]. Factorial [BRS04a, Cha81, DE18, Hau94b, TMR02b, Zha98a]. Factorials [LM19b, MS15a, IL07, RG14]. Factoring [Gru14]. Factorisation [Mil75a]. Factorization [DD70, DD65, FE99, Gal68a, Gal68b, Hir11b, KS16, Lev01, McD01, PW93, SW98, Goo93a, JMNRS22, Mat95]. Factorizations [Alf65d, Blo64, BMS88a, BMS88b, CDS03, Hen07, LKL02, BFS19, BLS75]. Factors [Alf63a, Alf63g, Alf64h, Ely65, EG76, Fra80, Gar04, GLMS22, GC90, HS15, HS81, Kim14, Li21, Pon19, Ber96, CH05, RL18, WJ76]. Fahr [Kwo10]. Failures [JQ97]. Fall [All77, Ano71b, Ano70e]. False [Han07]. Falsi [Tho63d]. Familiar [Eli21]. Families [Cha00, Gri11b, Hen20, LF14a, LF14b, LF15, Mel99a, Mel15, Mel16f, Mel18b, Mel18a, Pro15, Ter12]. Family [AFT22, AN83, Alp19b, Asv87a, Asv87b, BCS19, BH78b, CCT05, DD87, DG84, ESS20, ES15f, Fil97a, Ger09, Gre09, Gri18, Hen22a, KS11, Kos15c, Kos21c, Kos21d, Kos24c, LF14a, MS20, NS01, Nyb04, Nyb09, Tur89b, EMS10, HP92, KSSS16, Kos14a, Mol18, Wag07]. Famous [Wlo68d]. Fantasy [Bic65]. Far [DDK^{+}14]. Far-Difference [DDK^{+}14]. Farey [All75a, All75b, All75c, AS77, Dev99, Gup76, Pra82, ZY02]. Farkas [Big21]. Fascinating [HB97]. Fast [CCL89, Er83b, Kle07, Mae92a, Mae92b, MCPR95, PKS18, Rei92, TV07, TF92, Boy97, Joh04, Mas94a, Tak00]. Faster [DD95]. Faulhaber [Cer21, Zie19]. Faulty [JHK97]. Favorite [Rib05]. FCRC [ACM96]. FDTD [Rem11]. Features [Pet96]. Fence [AE22]. Fennessey [PW15]. Ferguson [Jun61, KS04]. Fermat [Ber74, BG91, CS81a, Dor93, Fox21, Gil89, Hew77, HB77a, Hor79a, LJ82, Mil91, Zei74]. Fern [Hil74e, Sha74c]. Feudominoes [Tur88c]. Few [BG21]. FFF [Rib05]. FFT [Pet96]. Fib [BF23]. Fibbinary [LSW14]. Fibonaccenes [KŽ05]. Fibonacci [Ano71b, Ano89d, Ano89a, Ano89b, Ano89c, Ano90b, Ano90d, Ano90c, Ano93k, Ano94c, Ano97a, Ano99a, Ano99b, Ano01a, Ano01b, Ano02a, Ano02e, Ano03b, Ano03a, Ano09a, Bee11, BP10b, BHP92, BPH93, BJ79a, BJ87, BJ99, BJ18, BL13d, Bri83a, Bri83b, Bro72h, Buc64c, Cad16, CSH72a, Cio18, Coo10, Coo16, FH94, FS14, Fre87, Fre88, Fre90, Gal08, Goo93c, Gra89c, HKN11, Har09, Hor04, Joh09, Joh12b, Joh20, Kah05, Kah06, Phi04, Puc62, Rib05, Sch94, Sim18, Sol07, Sub07, UNS07, Win65a, ZW01a, ZW06, dBF85, Abr95, Abr97, Abr00, Abr09, ABCM11, AK94, AAF15, AFRT23, Ade14a, Ade17, Ago49, Ago53, ABD05, AV03, ASV71, Ala89, Ala98, AGO95, AK74, Alf63a, Alf63c, Alf63d, Alf63f, Alf63g, Alf63h, Alf63i]. Fibonacci [Alf64a, Alf64c, Alf64d, Alf64e, Alf64f, Alf64j, Alf64h, Alf64i, Alf65a, Alf65d, Alf65e, Alf66b, All75a, All75b, All75c, AS77, AL79a, All17, AH16, Alu96, Alu97, Alv69, ACF20, AM15, And70, And95a, And20, AP92, ACOM21, AJ90, AJ91a, AJ91b, AJ91f, AJ96, And74a, And69, And74b, Ani95, Ano63b, Ano63c, Ano72f, Ano73d, Ano74b, Ano74c, Ano77b, Ano82b, Ano84b, Ano84c, Ano88c, Ano95a, Ano95b, Anoxx, Ano02c, Ano04a, Ano10a, Ano11a, Ano11b, Ano11c, Ano12a, Ano13a, Ano14b, Ano17d, Ano17c, Ano18a, Ano18b, Ano19a, Ano19b, Ano19c, Ano19d, Ano20a, Ano20b, Ano21a, Ano21b, Ano22a, Ano22b, Ano24, Ant85a, Ant85b, AK05, AF87, AB00, Arc84, AGJ^{+}09, Ard04, Ark65a, AH70, AH75b, AAB^{+}89b, Art04, AAB^{+}16, AESS04]. Fibonacci [Asv87a, Asv87b, Asv89, AAS85, Ata86, Ata89, AHM92, Ata95, AKO^{+}03, ADS14, AG87, AS20, Aus23, AW66, AD98, dAA79, BH80a, Bac66, Bac81, BCH12, BY16, BF16, Bai83, BCOR95, BE07, BBS81, BB83, Ban76a, Ban76b, BKV22, Bar71, BHS11, BB85, Bas63e, Bas63f, BH63a, BH63b, BP10a, BRGGGS20, Bea50, Bea64, Bea68, Bea18, Bea22a, Bea22b, Bec90, Beč01, Bed13, Bee04, BLPS11, BB19, Ben93, BQ99, BQS00, BNOS03, Ben06, BCQ07, BH14a, BCS19, Ber66, BD11, Ber11, BH76, BH77a, BH77b, BH79, BBHM85, BMRS01, BRS04a, BMRS04, Ber68, BH69a, Ber76, Ber75, Ber77, Bev71, BK90, Bha16, BJ78a, BJ79b, BJ85, BJ89, BJ90, BJE95, BJS96, BJF99, BJF01, BJ02, BJ03, BJ14, BH63c, Bic65]. Fibonacci [Bic69, Bic70, Bic71a, Bic71b, BH71, Bic72b, BH73a, Bic73b, Big21, BSY05, Blo53a, Blo53b, Blo64, Blo65, Blu72, Boa01, BHK07, BPGC97, Bol77, Bol84, BDD^{+}17, BYD^{+}19, Bom20, BFS19, BCFX09, Boo79, BK16, Bot82, Bou20, Bow14, Boy97, Bra90, Bra00, Bra71, BL13c, BL13e, BHL21, Bre64b, Bre64a, Bre94, BL86, Bri64a, BB65, BCJ^{+}05, BMS88a, BMS88b, Bri98, Bri92, BLT12, BDL^{+}16, Bro63a, Bro64a, Bro67d, Bro69b, Bro69c, Bro69h, Bro71a, Bro72b, Bro72c, Bro72d, Bro72e, Bro72f, Bro72g, Bro73a, Bro73c, Bro74b, Bro77, Bro65a, Bro65b, Bro68g, BD72, BH78c, BB08, BC77, BG76, Bru79, Bru96, BA98, BM07b, Bru70, BH63d, BV05b, Buc64a, Buc64b, Buc67, Bud80, BL17b, Bun75a, Bun75c, Bun78, Bun92, BB11, BK96, Bur71, Bur90]. Fibonacci [Bus63, Bus65, But78, BS96, Byr63, Byr75a, CDNN02, CDS03, CN04, CLX08, CLX13, CCZ15, CK10, CCT05, CR75, Cam22b, Cam23, Can51, Cap88, Cap89, CC03, CW03, Car63, Car64a, Car64b, Car66b, Car66c, Car68b, CH69, Car70a, CF70, Car70b, Car71a, CSH72b, CSH72c, CSH72d, CHS72, CSV73, Car74c, Car74b, Car74a, Car75a, Car77b, Car77a, CS77, Car78f, Car78g, Car80a, Car78i, CE66, CH94, Cas86, CKMR11, Cat09, Cat74a, CS92, CS04, Cha05, Cha06b, Cha07, CC16, Cha86c, Cha86b, Cha86a, Cha90, CMR02, Cha01, CM14, CM15, CZ81, CK12, CC14, CHKS05, Cho91, CS21, CJ24, Chr12, Chu10, CZ10, CL12, CMX20, Chu21b, Chu23, Chu92, fC93, fC95b, CCC00, CH05, CLHY12, CGS20, CG67, CB73, Chu74, Cig03, Cla86, Cla08, Coh64b, Coh79, CDM^{+}06, CDW13]. Fibonacci [CDM00, CK16, Coo84, Coo19, CV86, Cor89, CT18, Cox87, CM17, CT75, Cre70, Cro75, Cro76, CH89, CHF85, CM10, Cur68, CSS07, CMS12, Cus68, DP16, Daf20, DRT11, DMY13, Dan98, DN96, DH17, DB71, Dav71, Dav72, DGNW85, DD67, Day69b, DD70, Daz99, Daz02, DLR16, DL20, De 81a, DeC70, DeL70, DeL77, DeT81, De 78b, Dei66b, Dei66a, Dei72, DK17, Dek20b, DOU08b, DDK^{+}14, Den87, DL11, Des94, DD02, Des71, Des78a, Des86, Dev99, Dev11, Dev17, Dez77, DL92, Di 13, DF89, DFM90, DF92, Di 93, Dij78, Dij79, Dij81, Dij90, Dij96, DÖ03, DM08, Dil87, Dil00, DM10, DGMS14, Djo01b, Djo04, Djo05a, Dob14, Dod84, DX22, DD65, DP78a, DP78b, DGST88, Drm93, Dro00, Dro95, DMR^{+}16, DMSS16, Dub89, DS96, DMRS97, DK99]. Fibonacci [DT71, Duj96, Dun67b, Dun67a, Dun97, Dun09, DG66, DP24, DRS11b, DRS11a, Edg16, EZ04, EA19, EA20b, EA20a, ESS20, ESS21, Ehr83, Ehr89, EBKT85, EB87, EB88, EBK89, EMS10, ER00, EF98, EI03, Elm67, ET99, Eng90, Eng91, EBJ00, Epa85, ES15a, Er83a, Er83b, Er84a, Er84b, EB17, Erc76, EG72, EV83, EE10, EMG17, Ess03, Esw78a, Esw79a, EGS22, EM01, ES14e, ES15f, Eve75, FG05, FP09, FM15, Far86, Fau64, Fau65, Fed64, Fee67, Fei63, FZ03, Fen11a, Fer66, Fer76, Fer78, Fer65, Fer68, Fer69, Fer01, FFD91, FFD92, Fer97, FH08, Fie73b, Fie74, Fil86b, Fil92, Fil93, FH93, FF94, FB94, Fil95, FM95, Fil96, Fil97a, FH98, FE99, FW78, Fis76, FOR89, FK72, Fis96, FRU15, Fla78]. Fibonacci [Flo67, Fol04, FLP08, FPC08, FKM^{+}14, Fra84, FS99, Fra94, Fra02, Fra03, Fra70, FT84, FT87, Fre92a, Fre22, Fre68, Fre73, FF89, Fre92b, Fre94, FF99b, Fre06, Fro18, FZ08, Ful78, Gab70, Gal79, Gal86, Gam95, Gan59, GCMP14a, GCMP14b, GL16, GGL24, GR09, Gau04, Gau11, Geb67, Gel81, Gel63, Ger77, Gic09, Gic22, GM81, Gin54, Gir83, Giu04, GM20, Gla78, GZ15, Gla95, GSS19, Glu76, GR83, GR93, Gol06, GGL21, Goo71, Goo74a, Goo93b, Goo93a, Goo94a, Goo86c, Goo68a, Goo68b, Gou63a, Gou63b, Gou63c, Gou65a, Gou66a, Gou66b, Gou77a, GKH77, Gou78b, Gou81, Gou99, GQ07, GKD94, GZ19, GP94, Gra64, GS17, GLMS22, Gre64, GM78, Gre77a, Gre77d, Gre77b, Gri73a, GL80, Gri92, Gri09, Gri11c, Gri11d, Gri11b, Gri12, GG13]. Fibonacci [Gri13, GWT15, Gri15b, Gri17a, Gri18, Gri21, Gro97, Gru07, GFJT19, Gui20, GTU13, Gup76, Gup78a, GC90, Haa93, Hal68, Hal83, Hal65a, Hal64, Hal65d, Hal65c, Hal66, Hal67, HKN10, Han11, Han72, Han78, Har94, HS18, HS22, Har81, HS64, Har65, HS66, HD19, Har98, Has81, HB97, Hau96, Hau94b, HHJM15, Hay92, HSR97, HMT23, HK75, HK78, Hei64, Hei67, Hen94, Hen06, Hen78b, Her18, Her73, Her03, HLW94, Hey80, Hey81, Hig87, HG81, HW09, Hil21, HSG63, Hil74e, Hil74f, Hil74g, HP94, HPV95, Hin80a, HS19b, Hir81, Hly17, HT93, HH85, HB63, Hog64e, HB64a, HB64c, HB64b, Hog67a, HL67a, HL67b, Hog68a, HL68, HL69a, Hog69, Hog70b, Hog70c, Hog71a, Hog71b, HB72a, HBK72, Hog72a, HB73c, HA73, HCB73, HB73a, HL74, HB74c, HP75a]. Fibonacci [HA75, HB76c, HB76d, HBJ77a, HBJ77b, HBJ77c, HB77b, HB77a, HBJ78a, HBJ78c, HBJ80, HBJ82a, HBJ82b, HBJ82c, HBJ84, Hol75, Hol95, HR96, Hon15, Hop95, HS84, HT18, Hor61a, Hor61b, Hor61c, Hor63a, Hor63b, Hor77a, HLS79, HP86, HS88, HF91, Hor91, Hor82c, Hor83, Hor88d, Hor92d, Hor02d, Hor64, Hor66, Hor73, Hos73, Hos76, How03, HC11, Høy05, HSS03, HPL93, Hsu93c, HW15, HW81, Hug10, HW73, Hun66a, Hun70a, Hun72, Hun64b, Hun69, IR12, Igu94, IMS97, IK87, IA11, IT91, İpe11, IA14, Ism09, Ivi69, Iye69a, Iye69b, Iye69d, Iye69c, Jac89, Jac92, Jag21, JP19, Jai69, Jai74b, Jam18, Jam90, Jam02, JRdlR08, JRdlR10, Jar46, Jar63a, Jar63b, Jar64, JK97, JK98, JC22, Jav22, Jen93, Jen94, JR98, JLW07, JHK97, mJ10, JK94, Joh04]. Fibonacci [Joh12a, JS16, Joh22b, Joh40, Jon75, Jor64, Jor65, JV08, KM18, KLM^{+}19, Kai22, Kal82, Kan14, KI16, KA03, KY03, Kar04a, Kar05, Kar04b, KŞ13, KISS15, Kel57, Kel58, Kel92, Kel08, KK15, KS04, KK14, KH82, Kil07, KPAO11, KA13, KN08, Kim79a, Kim80b, Kim81b, Kim90, Kim01a, Kim02, Kim04, KM14, KKLS18, Kin63, Kin68, KP69, Kin73, KMT22, KR22, Kir77, KPT83, KP77, Kla68b, KŽ05, Kle91, Kle07, KBN08, KBN10, Kli81, KG02, KM95, Kno92b, Knu64, Knu88, Knu90, KC04, KSSS16, Koh66, Köh85, Kol65, Kom03, Kom09, KMP14, Kom14, KL91b, KP10b, Kos01, Kos14a, Kos15c, KH72, Kuh12, KT19, Kui69, KsS72, Kui82, Kum21, KYS13, Kwo07, Kwo14, LaB71, LaG13, Laf64, LW81, LLP^{+}18, LL14, Lan11, Lan92, Lan04]. Fibonacci [LL13, Lar88, Lay77, Led67, LL87c, LL88, Lee92, LL95, LLS97, LLKS01, LKL02, LKC03, LK03b, LK03a, Lee10, Leg15, Leh64, Leh66, Leh75, LL83, Len95, Lev68b, Lev01, Lew91, LSS76, Li99a, Li00b, Li12, Li21, LN81, LN84, Lin84, Lin67b, Lin70, Lip04, Liu02, LX21, LF69, Lon85, Lon86a, Lor29, Low71b, Luc99, Luc00, Luc02, LS07, LH09, LPW11, LS13, LY13, LS14b, LS14a, LP14, LS18a, LP18, LM19b, Luc21e, LOT23, LO14, Lün91a, Lün91b, Lyn70, Lyn73, LS18b, Ma98, MZ07, Ma11, Mab19, Mac00a, Mac00b, MG02, Mac10, Mae92a, Mae92b, Mah91, Mak94, MM06, Mak88, Mam61, Man64, Man67, Man68, Man77, MOP71, Man05b, Man17, Mar93, MS03, MB19, Mar12a, Mar12b, Mar12d, Mar12c, MBR^{+}13, Mar13b, Mar15, MBBB21, MSPF95, MR84, Mar09, Mas94a, Mas94b]. Fibonacci [MCPR95, MRPC95a, MRPC95b, MS04, Mat73, MMYA10, May68, May82, aMS22a, MP85, MS90, McC72, McD98a, McD02, McD11, MR07, MP66, Mea56, MR15, MS95d, Mel99a, Mel99f, Mel00a, Mel00d, Mel00c, Mel01, Mel03a, Mel03c, Mel04a, Mel04b, Mel09, Mel10b, Mel10a, Mel15, Mel16d, Mel16b, Mel16c, Mel16g, Mel16h, Ano17b, Mel17a, Mel17b, Mel17d, Mel18d, Mel18b, Mel18e, Men89, Mid79, Mil11, Mil60, Mil71, MK02, MN20, MPPW22, Min89, MH16, MSS08, Moh68, MV91, Mol18, Mol88, MC04, MN76, Mon74, Moo93a, Moo93b, Mor83, Mor89, Mor77, Mor95, MW22, MRS99, MRS00, MR99, MR00, MS15b, Mun07, Mun17, Mun05, Mur82a, Mur82b, Mus93, Mye75, NO15, Nav01, Nea07, Nef67, NW79, Neu17, NS05a, Nie72, NS94, NN87, NKW23, Nyb99b, Nyb01a, Nyb03, O’C72]. Fibonacci [OR99, ÖTA05, Ogu20, ON09, Ond70, OP14, OP18, Osl07, Ove65, Ove67, Ove73, Owi79, Oze05, ÖAD03b, ÖAD03a, Özk03, PTW13, PT17, Par70, Par68, Par17, Pas10, Pec65, PS02, Pel72, Pen20, Pep03, Pet94, PH75, Pet99, Pet12, Pet80, Pha84, PM82, Phi83, PGP85, PM85, PG89, PA91, PD18, Phi84, Pih88, PHJ67, PP79, Pil12, Pin02, PZ12, Pis02, PB15, Pla94, Pla95, Pla96a, PRS96, PS80, PH65, Pon68, Pon17a, Pon17b, Pon19, Pop84, Pop85, PL07, Pra82, Pri05, Pri18, PS76, Pro81, PT82, Pro83, Pro16b, PT89, Pru67, Puc01, QW96, Raa63, Rab99b, RS11a, Rai48, Rai54, RG14, Ram14, RRD14, RR14, RMA96, Ran66, Ran95, Rao53, Rao55, Rao03, Rea70, Rec75, Red94, Rei54, Rei92, Rei93b, Rei04, Rem11, Rie99a]. Fibonacci [RHT18, Ris05, Rit23, RS98b, Rob78a, Rob81, Rob83a, Rob83b, Rob83c, Rob90, Rob94, Rob96, Rob78b, Rob63, Rok96, Rol03, Ros85, Ros83, RT02, Rou13, Rou11, RSHL07, RH63, Rug63, RL15, RL18, RP71, Rus11, Rus82, RS11b, Rya75, Ryd96, Ryt06, SE13, SE20a, SE20b, Saf92, Sal12, San90a, San90b, SI03, SI05, SBM97, Sbu02, Sca99, Sch81, Sch82, Sch84, Sch50, Sch19, Sco68a, Sco68b, Sco68c, Sco18, Sed14, See91, SR12, Sha68, Sha88a, Sha90, Sha73, Sha72a, ST73, SH71b, SH73, Sha74c, Sha74d, SHC74, Sha78a, Sha79b, Sha84b, SH88, Sha60, Sha76, Sha72b, Sha64a, Sha13, SC10a, SCH11, She68, SCY^{+}23, Shi20, SC05, Sho78, SD18, SMSM01, Sig02, SG76, Sil76, Sil77a, Sil77b, Sil63, Sil11, SH80a, Sin85, Sin75, Sla16, Sla77, SH76]. Fibonacci [Smi09, SK81, Sob17, SC98, Sol05, Som72, Som75, Som77, Som82, STL22, Spi70, SJC92, SCJ93, SCJ95, SC97, Spi06, Spo22b, Spo22a, SR19, Squ81, Sri20, St.84, Sta00, SMS99, Sta17, Sta75, Sta76, Sta80a, Sta80b, Ste78b, Ste05, Ste11a, Ste95, Ste21, Ste23, Sto09a, Sto65, Sto77, Sto75, Str99a, Str99b, SYY04, Sub19, SB93, Sur14, Süt89, Swa66a, Swa73, Swa97b, Swa99a, Swe64, Swi04, TR01, TMR02a, TR03, TMR06, TP86, Tak00, TD16, TD19, Tan00, Tao17, TV07, Tau68c, Tau68a, Tau73a, TL08, Tay67, Ted19b, Tem93, Ter96a, Ter96b, TF92, TF93, Tho63d, Tho63f, Tho65, TJ88, Ton24, Tor19, TS92, Tra88, Tre16a, Tre16b, Tre96b, Tre96a, Tru73, Trz94, Trz96, Tue22, Tue23b, TKS11, Tur74, Tur79, Tur86, Tur88a, Tur89b, TR90]. Fibonacci [TS00, THKH23, Uli11, UP83, Vaj89, Vaj08, Van12, Var89, Vau76, Ves15, Vin78, Vin63, Vis00, Vor61, Vor83, Vse11, WS67, Wad74, Wad78, Wag06, Wag07, WR01, Wal10, Wal60, Wal63, Wal79, Wal84a, Wal85c, Wal80, WH71, WH74b, WH74a, Wan95, Wan00, WH04, WW17, Wan20, Was81, Was99, WP69, Web75, Web95b, Wei66a, Wei16, Wel67, Wel94, WY00, Wes66, Wes80, Whi63, Whi86a, Wil86, Wil72, Wil75a, Wil75b, Wil82, WS80, Wil98c, Win65b, Wit72, Wlo13, Wlo63, Wlo65, Wlo71a, Wlo73, Wol98, Woo07, Wu97, Wul81, Wun63, WSH74, YK77, Yal77, YS08, lY08, YZ18, Yay11, YZ17, YK03, Yoo21, Yor14, You94, YL97, YZ02, Zar70, Zec72a, Zec72b, Zei77, Zei96, Zei63, Zei64, Zei67, Zha97a, Zha97c, ZJ98, Zha99a, Zha02, Zha04]. Fibonacci [ZG17, Zha99b, ZW01b, Zha01b, ZW01c, ZW03, ZW04c, ZL06, Zha09, ZQZ20, Zho96, ZJ14, Zuk76, ZT88, dF11, dF14, dBL83, deB74, de 81b, de 95, Tol17b, Tol17a, Bro69a, Gau02, Hal22, Høy09, Oak09]. Fibonacci- [Tur79]. Fibonacci-Automatic [DMSS16, DMR^{+}16]. Fibonacci-Free [Cho91]. Fibonacci-Like [Asv87a, Asv87b, Asv89, BCH12, BBHM85, BCFX09, CC03, CZ10, HG81, JK97, Köh85, Mor89, Pha84, Pla94, Pla95, Som77, Som82, Tay67, Tur89b, Vse11, Wal84a, BV05b, IK87, IA14, Knu90, MMYA10, Mor95, Rit23]. Fibonacci-norm [HKN11, HKN10]. Fibonacci-number [LLP^{+}18]. Fibonacci-polynomial [EE10]. Fibonacci-Producing [HT18]. Fibonacci-Related [BD72, GG13, Lyn73, PS76]. Fibonacci-Rowed [BH80a]. Fibonacci-Sum [FKM^{+}14]. Fibonacci-Type [DGNW85, Gre77b, HLW94, HB77b, KK14, PH75, PGP85, PM85, PG89, Rug63, WH04, GQ07, HLS79, KM18, Ma11]. Fibonacci-Wise [Dun67a]. Fibonacci-Word [Gri18]. Fibonaccian [Fer60, Jun61, KC63, Kro68, Kum74]. Fibonaccious [Ely65]. Fibonacci’s [NW83, Lün91a, Lün91b, Gri15a]. Fibonacci’s-rabbits [Gri15a]. Fibonomial [BP09, BR14, DK18, Gou69, Gou74c, Hol94, KPAO11, KW21, KW95, MT12, MST13]. Fibonomial-Coefficient [Hol94]. Fibonomials [Bal15, Gou72c]. Field [BN03, BN19, DES17, Dod84, Lin68, Pla96b, Sin74, ESW07]. Fields [dOGC96, Bas81, Bou20, Bri92, BF95, Mac00a, Mak12, Mol88, Stă14, WP19, BN06, GWZ07]. fifteenth [Fra03, Fre92a, Uli11, Joh12b]. fifteenth-century [Fre92a]. Fifth [BHP92, BPH93, Fre06, Ano90a, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b]. Fifty [Joh12a]. Fig [Sch94]. Figurate [Hin78]. Figuring [Ano16a]. Filbert [KP10a, KP13, KA19, Ric01]. Filipponi [AJ94c, Bru94c, PP00]. Filling [Her03, Tri89]. Filtering [Arc84]. Filters [Bot82, Zuk76, AAF15]. Final [Ano18b, BH05]. Find [Fis81, Her19]. Finding [Ark73c, BRS04b, Dev17, Gou90, Har70a, Har70b, Har74, MS79, She69, She73, Ten09, PT89, Sim18, Tol17b, Tol17a]. Finite [dOGC96, AD98, Bri92, Bro67f, Bro78a, BF95, Cal97, Cha12, DB22, GQ11, Jos79, Kno92b, KÖS22k, Maš09, Mel04a, Mel15, Mel16a, Mel17b, Mel17c, Mel18d, Mel18b, Mel18c, Mel18e, Mel18f, MM75, Sin74, Sof12, Ton97, WP19, Wei79, Yap70, BN06, GWZ07, Gou06, ÖAD03b]. Firard [Gou99]. Firenze [Arr69]. Firing [Tor19]. First [Ano09a, Ark73b, Car80c, Car80d, FBJ01, Gor12a, Hor77b, Hor92a, KC86, Kim94, Kun90, Kun87, Lee02a, Par77b, PRW93, Sch20, Ste71, Web75, Wit72, CHF85, LS14b, dBF85]. Fission [Kim14]. Fissions [KKLS18]. Five [Bro69d, Gar04, Sun02]. Fivefold [Man17]. Fiven [Hoi22]. Fixed [BL13b, BN77a, Gri14a, Mar12a, Som82, SK13b, Tri09, Tri10]. Fixed-Term [Gri14a]. Florence [Arr69, Pic79]. Florets [MC76]. Florida [IEE84]. Flos [Pic83a, Pic83b]. Flow [Sch22]. flower [Han12]. Flowers [De 78b, Rib05]. Fluid [Zen78]. Focusing [Lar20]. Folded [Gre76, Gre78a]. Folyominoes [Tur88c]. fondements [MS15b]. Fontené [Gou69]. Forbidden [Pro95]. Force [Can03a, Can03b]. Ford [Put76]. Foreword [Ano14c, Ano17e]. Forgeries [Ano95a, Ste95]. Forgotten [Akr93, BB08, Mar09, Dev17, Sim18]. fork [Sar90]. fork-join [Sar90]. Form [Akr93, Alf63g, All75c, And84, Ant85a, BAP73, Dan98, Dra69, EF98, Esw79a, FB99, FL68, FP96, Goo71, HH00, Her18, Hun78, Izo99, LL14, Lar15a, LF17, MH75a, McD02, Now86, Pet81, Rob83b, Rob90, Rob94, Rob05b, Wil75a, Zho99, BJ85, Boy07, Esw79b, EP07, KK15, LS07, Rob84, Sun06c]. Formal [Gou75, Hau93a]. Formed [Ark73c]. Forms [Ani95, Coh82, Eli01, ES12e, Gau12, Giu71, Iak81b, Mel16a, Mel17a, Mel17b, Mel17c, Mel18b, Mel18e, Mel18f, Pas10, Pin02, Rob81, Wil87]. Formula [Ade14b, AM75b, And78b, Bas64, Ben02, Ber84, BMRS01, BRS04a, Ber76, BK90, Bro78a, Bru75a, BR82, Cas89, Col83, CK88, Dei66b, Fil92, For67b, Gau89, Gou77a, Gri18, Gup78a, Hag64, Hen94, Hir81, Hor82a, Hor86c, HT86, Jar67b, KS04, Kha81, Kra90, LLKS01, Lef82, McC81, MV15, Mom01, Muw81, Nor22, PP00, Pel72, Rob91a, SP02, Spi82, SJ84, Swa77, TS89a, WH03, Wan20, Whi77a, Cha07, CV86, KM18, Rei54, Rob07, Rol03, SE20a, ZZ06]. Formulae [And69, Els05, Hic73, Tau65, Tru73, Whi66a, Blo53b, CY07, CSS07, Rao55]. Formulas [Ade14a, AS84, AP70, AAB^{+}89b, AHM92, BJF99, Bro63a, Bro68f, Bro69g, Bro73a, Car71a, Car71b, Car72e, Cha09, CH02, CB14, Du00, ES10e, FZ03, Fil97c, Fil97d, Gla95, Gou67a, Gou78b, Gou99, Gri11a, GZ22, Hor97c, How86, Hsu95, HT00, Jar68b, KPAO11, Lee01b, Lin98, Liu02, MH85, MS95d, Mel02, Mel18d, Mel18c, Nas76, NS99, Pet84, Rei93b, Sch20, Sha13, Str99b, Tre16b, Wen02, Zei64, Zei67, ZW02, dBdV94, dB95, Kel57, Kel58, Liu07, Tem93]. Formulating [Hen18]. Formulation [BRS04a, LF17]. FORTRAN [KTW84]. Found [HB73b]. Foundations [IEE84, IEE08]. Four [Ark74, Bro69e, DeL88, Ewe86b, Ewe03, GG15, Gou74a, Hen22b, Hir18, HPL71, Kar69, Kim80a, Kos21e, Kos21f, Lar15b, MC76, Rob83a, Vol13, Web82]. Four-Number [Web82]. Four-Part [Ewe86b]. Four-Square [Ewe03]. Four-Step [Kar69]. Four-Tuples [DeL88]. Fourier [DP78a]. fourteenth [Fra03, Ano09a, Coo10]. Fourth [BS11, HB64a, Mel00a, Can51, Ano89d, Ano89a, Ano89b, Ano89c, Ano90b, Ano90d, Ano90c, Fre90]. fra [Car39]. Fractal [DK18, Gro97, HS19b, Hol06, Kim10a, RR14, RRD14]. Fractals [Ano93f, Ano93g, Rei93a, Sch91, Tur03]. Fraction [Adl78, Alb73, And17, And20, BCOR95, Bic68a, BB65, CK10, Dra64, Dra69, GQ09, Her20, HB77b, Lin84, Pih01, Pih10, Sma77, Spo22b, WTvR89, Wlo68a, Wul76, Kom09]. Fractional [Har94, Kim81b, Mel18d, Mel18b, Mel18c, Gau11]. Fractional-power [Gau11]. Fractions [Alf64a, All75a, All75c, AS77, BD11, BCFX09, Bow14, Bro71a, BAP73, DG66, Els98, Fil86a, Goo93b, Goo93c, HK91, HA64, HW81, Köh85, Kom01a, Kom12, LL87a, Len03b, LM10, LJ67, LJ70, Lon81b, MC04, Moo93a, Rat93, Rie93, RT04, Rob95, San23, Sha82, SH88, Stă14, WH03, Wat92, Wil00, Car07, CCL89, Coo06, Fra84, Fre06, Lün91a, Lün91b, MS15b, Nea07]. Frances [Bro70a]. Franklin [Nor16]. Fredholm [MH16]. Fredric [Joh22a]. Free [Cho91, Dun69b, Kim83, MV91, Nyb09, PB89, SK12, Wal80, PG13, Sla16]. Freitag [Ano96a, BCH12, Rib00]. French [Ber96, CHF85, Spi00a, Spi00b, Zec72b]. Frequences [Ber96]. Frequencies [YS08, Ber96]. Frequency [DN96, SS01]. Friendly [BS87, Bal90]. Friendly-Pairs [BS87, Bal90]. Friends [Ano02c]. Frobenius [BS93, Pad89, Röd94, SK22a]. Fun [Bee04, HB72a]. Function [AC72, BL93, Bra72, Bro78b, Bru86b, Bun75a, Bus65, CVZ17, Car72d, CSV73, Car77b, Car77c, Cop18, DFLT14, DG98, EKS88, Eva77, Ewe80, Fra94, Fra67, Fuc84, GL83, GP83, GKRS01, Goo86c, Goo68b, Gou64b, Gou67a, Gou69, Gou94, Hah72, Hal83, Hau97a, HK78, Hei67, Hen11b, Hen15b, Hin78, HL67a, Hon15, How85, How87, Jos98, KP69, Kol65, Lan00, Lan02, LOF14, LO18, LS13, LM19b, Mak12, McD94b, MP66, Mil75b, Mil91, MS92, MR95, Par68, PS19, Rob04b, Sha73, Sha74d, Sho99b, Sin90, SŠŠ22a, SŠŠ22b, Sta80a, Sto09a, Tót87, TS89a, UC69, Whi86a, Wil00, ZG17, CJ24, Ewe07a, Ewe07b, mJ10, LTT95, Len06b, LS14b, Rob99, SS95a, SS95b, Goo68b]. Functional [Bre74b, Bur90, Gry03, Rat93, Web95a]. Functions [Ain77, AH77, AJ91b, BS87, Bal90, BCM10, BN77a, BRS04b, BH63c, BSY05, BDM12, Bru72, Bru75b, Bru77c, Bru77b, BR82, BL17b, Bun78, Bun12, Bur90, BS96, Byr63, Byr65, Car69, Car74b, Car75b, CR93, CB73, CS81b, CS81c, CKR98, DH76, DeL95, DL11, Dil00, Djo04, Elm67, Fra87, Ful81a, Gir81, Gou63b, Gou99, Gua78a, Gua78b, Gua78c, Gue78, Has81, Hau88, Hau90, HR91, Hau93b, Hau00, Hin77, Hog71b, HA73, HBJ78d, HH79, HBJ79b, Hor69, Hor74a, Hor78a, Hor79a, Hor80, Hor88a, How96, JP19, Jen97, JQ97, Kel92, Kim80c, Kim10b, Köh85, LKC03, LW87, Luc99, Luc00, LY13, Mac00b, MH87b, MH87c, MH90, MS90, ML84, Mel99b, Mel99c, Mel99d, Mel16a, Mel17c, Mel18c, Mel18f, Mel18g, Mel18a, Nyb03, Pet85, PH88, Pon17b]. Functions [Pop77, Pop86, RW21, SP18, SP19, Say76, Sho99a, SH79b, SK20, Spi70, Sta03, Sto09b, Tsu21, Vau76, Wal76b, Wil77, ZW01c, CHS72, DRS11b, DRS11a, GL80, LM93, Man05a, Mül06, NS94, RL89, Wan00, You19]. Fundamental [DeL81, Knu97, Mel00b, Pet85, Sha74e, SR19, TMR06]. Fundamentals [Goo68a]. fundraiser [Ano22d]. Further [DS20, Hei64, Hor63a, Kro68, Mel18e, Mel18f, OS05, Sen73, SGBS86, Swa68, Swa00c, Tre16a, WH74a, ZQZ20]. Fusion [Kim14]. Fuss [BGW14].

G [Ano90e, Gra89c, Sch94, SH82, AFRT23]. g-Repdigits [AFRT23]. G. [Har74, Pro04]. G.C.D. [Gou95, McD91]. Gained [Bri64a]. Galois [SCY^{+}23, ZQZ20]. Galton [Hey81]. Gambler [KH82]. Game [BDD^{+}22, BSEFM19, Bok84, BCD^{+}20, CST05, Cla86, Cre87, Cre88, Ehr90, ES15a, EJ17, GFMR^{+}24, Hoc79, HBJS79, JO24, Len03b, LY90, Lud88b, MN20, Sch85, Web82, Wlo72, BP07]. Games [CDH^{+}21, CDH^{+}22, EF80, Fla82, HLW94, LLM^{+}20, MSY22, Sch70, Fra84, LM71, Bic73a]. Gap [BBG^{+}13, PR13, vRWT90]. Gaps [BCC^{+}20, LM19a, MS15a]. Gate [Nod90]. Gates [CLX08, CLX13]. Gateway [Swa92]. Gauss [Rob02d]. Gaussian [AK74, Ber77, BDE^{+}14b, CCG^{+}19, Col94, DDK^{+}14, HJ68b, Jor65, PH88]. GCD [AS05, Bes91, BL92, Chu96, FJ12, HS12, Kor94, LSA07, LK10, aMS22a, Sin73a, TJ92]. GCD-Closed [BL92]. GCD-s [HS12]. GCDs [LK10]. Gegenbauer [Dil87, HP81, Hor85a]. Gelin [FG05, Mel03a]. General [AJ94b, AJ95a, AJ95b, AD98, Bal15, Bro89, Bru75b, Col83, Col87b, Cre70, Cro76, DD13, Ful81b, Gau04, Gou75, Hal65d, Han78, Hei67, HP72, HB76b, Hor93, Ivi72, KA03, Lar15b, LF17, Lay77, LL87a, Lee01b, Lin84, MJ95, MS79, Par64, Pha84, Pla96b, Sto75, TMR02b, Tay75, Tru73, Wit16, Zei71a, Zho99, Kar05, WS80]. Generalised [HS22, Mil11, Rou11]. Generalization [AJ94a, AAS85, Ata86, Ata89, Ata95, BCOR95, Bow88, Bri06, Bro67d, Bro63b, BG76, Cas89, Cas91a, Com79, Dav82, DDK^{+}14, Dil87, Djo98, Fer76, Fow82, FS01, GL83, Gbu81, Gil89, Gil92a, GZ15, Gou74c, Gri73a, Gri17b, HS64, Hen17, Hen22b, HBJS79, Hor86a, KS04, KZ92, Koh66, LB13, LL87c, Lee02a, Liu96, LA92, Ma98, Mac88, McC88, McI92, MS95b, MS95a, Nyb99a, Raa63, Ran66, Rec75, Sch85, Sha74c, Sha79a, Sho99a, Sin89, Som02, SK22b, Swa77, Swa99b, Tót87, Yu98, Zha99a, Zho97, KA19, Krč06, RRD14, Sha88b, YZ17]. Generalizations [AM75b, AH97, And69, AKP19, Bru73, Bus63, Car81b, Dan98, EG17, FG05, Gou63a, GQ10, Hor97a, Ism09, Kel72, Kos23g, Kos23h, Kos23k, Kos24b, Kos24c, KT19, Lan92, Man05b, McG10, MS95c, Mel99g, Mur82b, Nor22, RR14, SM93, Sla16, SJC92, SCJ95, Sto70, Swa00a, TR90, Wad92b, Wil72, ZL98, ZW01a, ZW01b, Bil14, Bra06, SI05, AKP20]. Generalized [Abr95, Abr97, Abr00, ABCM11, Alf65a, AC72, AP92, AJ91a, AJ91b, Ant85b, AH75b, AHM92, AS20, BSEFM19, BB69, Bas63f, BCFS22, Bea22b, BBG^{+}13, BQS00, BLP10, BH74, BH77a, Ber84, BMRS01, BRS04a, BMRS04, Ber68, BH69a, Ber76, Ber75, BJ85, BJ90, BJ93, BJE95, BJS96, BH73a, BH73b, Bis84, Blo65, BH14b, Boy09, Bra00, BBB06, BBL14, Bro73a, Bro71b, Bru79, Bru89, BR82, Bun75c, BHS06, BK96, Bur90, Car70b, CH78, CMR02, CM14, Chr12, CMX20, Chu21a, CG67, CS85, CK16, Coo84, Coo19, CJCK17, Cus68, DN96, Day69b, DeC70, Def13, Den87, DFM90, DM10, Djo96, Djo00a, Djo00b, Djo01a, Djo01b, Djo04, Dre97, DG98, Dub89, DMRS97, Duj96, ER00, Eng91, EBJ00, EA76, ES14e, FZ03, Fer66]. Generalized [Fil92, Fil97c, Fil98c, Fil00, FOR89, FK72, Fla78, Flo67, Fra67, Ful78, Gab70, GP16, Gau98, Ger77, Ger78, GW19, GU66, Gla95, Gou69, Gou71b, Gre78b, GT01, GCK^{+}19, Gup74, Hal68, HPS14, HS66, Hau90, aHZ02, Hil74f, Hod74, HH65, Hog67a, Hog68a, HB69, Hog70b, Hog70c, Hog72a, Hog72b, HB73c, HB73b, HL74, HB75a, HA75, HA76a, HBJ77e, HBJ80, HBJ82c, HBJ84, Hol00, Hor61a, Hor61c, Hor65, Hor74a, Hor77a, Hor78c, Hor82a, HP86, Hor86b, HF91, Hor96d, Hor98, Hor00a, How03, Hsu87, Hsu97, Iak77b, Iak77a, Iak81b, Iye69a, Jai69, Jai74b, JHK97, Jon96, Kal82, KK15, KS04, KK11, KP10a, KPAO11, Kim80b, Kim11, Kin68, Kin73, Kle91, Kol65, Kom14, KY20, Kuh12, Kyr93, Lan11, LK09, Lar16, LO18, LF19, LL88, LL95, LLS97]. Generalized [LLKS01, LKC03, LK03a, Lev68c, Li00b, LM17, LM19a, Li21, LJ82, Lin71, Lin75, Lud88a, Mac00b, MH87a, aMS22a, MS90, MS94, MS95d, MS95g, Mel98, Mel00c, Mel01, Mel04a, Mel15, Mel16e, Mel17b, Mel18d, Mel18b, Mel18e, Mil60, Mil71, MSY22, Moh68, MV15, Mol88, MT17, MRS99, MRS00, MR99, MR00, Mun05, Mus93, ND87, Nyb99b, Nyb01a, Nyb03, OS98, OS05, OS21, Pad89, Pag74, PH88, Pet12, PA91, Pon68, Pro15, Ram16, RFLT20, Rui17, Rus82, RT18, Sbu02, Sch84, SH71a, SH71b, SH73, Sha79b, Sha79c, Sha84b, SH88, SH91, She72, Sho99b, SH80a, SŠŠ22a, SH76, SH79b, SH80b, SCJ93, SC97, SR19, Sto77, Str99b, Sul89, Swa73, Swa97a, Swa99a, Swa99c, TR01, TMR02a, Tho63f, TF64, Tos78, TS89a, TS89b, Ula23, UP83, WS67, Wad74]. Generalized [Wad78, WR01, Wal79, WH71, WH74b, WH74a, WH84, WW17, Wan20, Whi77a, Wil75b, Wol98, WM75, Wul81, Xie02, YK77, Yal77, You94, YWH96, Yua02, Zak83, Zec72a, Zha97b, Zha97c, ZL98, ZJ98, Zha99a, Zha01b, ZW01c, ZW03, ZD04, ZW04c, dBL83, BL13d, BL13c, BHL21, Bre94, BM07b, CM15, CHT06, CM10, DMP07, DL20, Djo05a, EM01, Flo18, GGL24, GTU13, Hol06, Ila71, JRdlR10, JLW07, KI16, KŞ13, Kil07, Len06a, Luc21e, Ma11, MP85, Mol18, ÖTA05, RT72, Ram14, Ric93, RL15, Sob17, TR03, TK04, Wag06, Wan00, Wlo13, lY08, YZ18, Yay11, YZ17, ZW06, HBJ78c, Ano93f, Ano93g, Ano93e]. Generalizes [Mel16h]. Generalizing [CFH^{+}14, CK95, FS01, HT22, MMMS22, MMM^{+}22, PD18]. Generate [Kar69, Low71b]. Generated [All76, Alp11, Bru86a, Cai96b, DvT83, Dil91, DT20, Gou74b, HBJ82a, HS81, HH82, Hor84a, Hor88a, KM14, LP03b, LS18b, NKW23, Rei93a, SHB84, Smi68, Sun79, UC69, Win65a, Win65b, Zar70, Zöl93, Sha06]. Generates [Du89]. Generating [Ain77, Ano63b, Arp94, Arp04, BL93, BLL71, Bra90, Bru72, Bru75b, Bru77b, BR82, BL17b, Bus65, Car69, Car72d, Car74b, Car75b, Car77b, fC95b, CB73, CS81b, CS81c, Cre89, DF92, Djo04, Fra67, Ful81a, Geb67, Gou63b, Gra82, Han72, HR91, Hen18, Hog71b, HA73, HBJ78d, HH80, Hon15, Hor74a, Hor77a, How96, JM14, Kim80c, Kim10b, Köh85, Kol65, Lan00, Lan02, LO18, Lau77, MH87b, MH87c, Oli96, Pon17b, Pop77, PS19, Ser74, Sha74d, SH79b, Sta03, Tan00, Tsu21, ZG17, Len06b, RL89]. Generation [CM17, EH95, Fie67a, Fie68a, Hor92a, Jan02, MOP71, MRPC95a, MRPC95b, MN76, Som84, TF92, L’E94, LSS12, Mas94b, MS00a, MS00b]. Generations [DRT11]. Generative [Mit03]. Generator [BL17a, MCPR95, Krá72a, Krá72b, MNZ90, MZT90, Mas94a, Pet91, Pry94, Wol92]. Generator-Based [BL17a]. Generators [Alu96, Alu97, And90, Col87a, Erc79, HK75, MM65, MZ93, McC85, Naj89, Pet94, PS76, Ral85, VKSAN93, Alt88, AM15, And95a, AI92, CMP94, Edd90, Mak94, MZ91, Mar92, MS04, PW95, TLC93]. genius [Dev17, Sim18]. Genocchi [Hor92a, Hor92b, ZJ98]. Genome [YS08]. gentle [MSPF95]. geometria [Arr69]. geometriae [Fre92a]. Geometric [Alf64f, BL93, BYD^{+}19, Bom20, Bra00, CL12, Erc73, Ful80a, Gri14b, Hsu97, Pro95, Sch84, Tri73, Whi78e, Xie02, YL04, EB17, Fra40]. geometrica [Fra40]. Geometrical [HP94, Jai74b, Puc62]. Geometrically [Ste71]. geometrie [Hug10, Høy09, Oak09, FH08]. Geometry [Bea66b, Ber84, Bro72c, DeT81, Hor82a, Hun64b, TS00, Tur03, Arr69]. Gerald [Ano21c]. Germain [Nyb12]. German [Geg93, Lün91a, Lün91b, Lün92]. Germany [Bro96]. Gerneralized [Fri66]. Gessel [Kyr94]. Getting [Tri67c]. GF [Wie88]. GFSR [TF93]. Gibonacci [Kos15a, Kos18a, KG18b, Kos19c, Kos19d, Kos20, KG20, Kos21a, Kos21b, Kos21c, Kos21d, Kos21e, Kos21g, Kos21h, Kos22c, Kos22d, Kos22a, Kos22b, Kos22h, Kos22j, Kos23b, Kos23c, Kos23d, Kos23e, Kos23a, Kos23g, Kos23i, Kos23k, Kos23j, Kos23l, Kos23m, Kos23p, Kos23q, Kos24a, Kos24b, Kos24c]. Gibonomial [Kos15b]. Gies [Bro70a, Bro70a]. Ginsburg [Kos14b, Kos17, Kos18b]. Girard [Wan20]. Giusti [Cio18]. Given [Car78d, Car80b, DB66, DB67, HK91, Kla68a, KS02b, PH75, Tsu21, BM07a, Rob06b]. Gives [Goo86c]. Gizeh [Bea68]. Glaisher [How90, MS92]. Glaser [Bre98].

Glass [AJ95b, Bro89, HBJ79a]. Gleason [SC84]. Global [You19]. Goka [AN79]. Gold [Alf65d]. Goldbach [HJ68b]. Golden [AAF15, All77, Bea66a, Ber66, BFT15, BJ78b, BJD95, BJ99, BH69b, Bra00, BF91, Bro17, Bru94a, BH17, Cad16, Cha73, Chu21a, fC95a, DA79, DvL23, Eng87, Erc72, Erc73, Erc76, Fis81, Fow82, GD09, GM11, Goo93b, Goo93c, Gou64a, HPS14, Hed76, Hin77, HBJ79b, Hun64c, Hun74a, Lar78, Lee92, Lev84, Lin96, LL84, Mel10c, Met97, Mig77, Mon76a, Moo93a, Moo94, Pet86, Pet96, Pro96, Pro15, Pru67, Rac19, Rie99b, Rus12, RS11b, Sch82, Shu72, SH77, Sza17, VWT85, Wal85a, WTvR87, Wlo63, Wlo67, Wlo68b, Wlo71b, Wlo71d, YWH96, Bea50, BL65, Dun97, Fis96, Gri13, KN08, Krč06, Nea07, Vaj89, Vaj08, Ade14a, BT91, CVZ17, Dek20a, EF98, Her14, LSW14, Ove65, RT02, WB12]. Golden-Fibonacci [Lee92]. golden-ratio [Gri13]. Golden-Ratio-Based [BH17]. Goldpoint [Tur03]. Gonal [ES17, Ten09]. Good [Raw87]. Goose [Lev84]. Gould [Bru73, GS81, Gou89, Hen17, HH72, LA92, Sin73a]. Governing [LF14a]. GPU [AM15]. Graceful [BB83]. Grammar [HR96, Web95a]. Grand [Ram16]. Graph [Ala98, AH15, HW09, Kos14a, Kos15c, Kos19c, Kos21e, Kos21f, Kos23d, Kos23l, Ord93, San90b, Sha78a, BV05a, LaG13]. Graph-Theoretic [AH15, Kos15c, Kos19c, Kos21e, Kos21f, Kos23d, Kos14a, Kos23l]. Graphics [CLRS86]. Graphs [And02, Ano93f, Ano93g, BBS81, BB83, BCFS22, Bro97, CK97, DD87, Def13, DQ83, ESW13, Fie74, FKM^{+}14, Gou67b, HS84, HPL93, KR22, KPT83, KŽ05, NC85, NR17, PT82, RM04, Rit23, SBM97, SGBS86, Ted19b, CM07, EBKT85, EB87, EB88, EBK89, HP92, Lee00, Ryt06, Sta17, Wag06]. Gratitude [Ano00d]. Gray [BKV22, Kim10c, Rit23]. Great [Acz11, Teu04, Bea68]. Greatest [AC72, BC10, BJE95, Bro78b, Car77c, Day70, DT71, Eva77, Eve75, Gou72b, Har70a, Har70b, HBJ79b, Jar63a, Kim79a, McD94b, She69, She73, Str73, Dij90, Dij96, Mor95]. Greedy [Cre89, Pih01, Pih10, Ste05]. Greek [Cha73]. Greenwood [SC84]. Greig [Gou89]. Grids [Hor99, MBBB21]. Group [Lin67c, LP03b, Man77, Som72, Sun73, Sun78, Tao17, Wal80, Zha01a, Hol95, ÖAD03b]. Group-Theoretical [Sun73]. Grouping [Wlo67]. Groups [AD98, CM01, De 81a, EV89, Kno92b, KÖS22k, MT17, Sha76, Smi68, Som77, Sun79, Wil86, Yap70, CR75, CS92, DÖ03, KA03, KY03, Men89, Özk03, Sla16]. Grow [Goo91, PKS18, Swi04, JRdlR08, MM06]. Growth [BDD^{+}17, BWWW16, Cha05, Coh79, ET99, EK69, Hop95, Kla76, Lan11, CM07, Cha06b, IT91, JRdlR10]. Grundy [Sch85]. Guessing [Wal85b].

H [Bic73a, Bri83a, Bri83b, Pro04, Wer83]. H-315 [Wer83]. Hagenberg [Jef08]. Hail [Eve81]. Halsey [Bun75a, Car77b]. Hamilton [Hen14]. Hamiltonian [Süt89]. Hanging [Boa79]. Hanoi [HS19b, Rit23]. Happy [GT01, GT03, GH18]. Haradam [Hil74f]. hardback [Mos08]. Hardness [CLX08]. Harmonic [BJ81, Cer77, CK12, CS98, CS01, GP83, HMT23, Pre68, Rie99a, Sch84, Ste21, Wal83]. harmony [SO09]. Harris [Rie76]. Hart [Hal22]. Hart-Davis [Hal22]. Hash [Jen97]. Hasse [HSG63]. Hausdorff [WR01]. Having [BJF01, BJ02, Pih92, Rie93, SK21, Bri83a, Bri83b, LL83, MSS08]. Head [Ber11]. Heads [Gri11d, Hir12b]. Heaps [AK94, Boy97, DGST88, FT84, FT87, GFJT19, HKTZ17, BLT12, KT08]. Heights [GT03, HL67a]. held [Bro96]. Help [Fie73b]. Heptagonal [Hin80b, Rao02, Rao03, Rao05a, Rao05b]. Here [Boy97]. Hermite [AKP20, AKP19, Djo96, Gou63a, Gou72c, Gou74c, Gou95, MB19, Sub90, Sub95]. Heron [Nea07, PS21]. Heronian [Sin73b, Car70c, Hun67]. Herta [Ano96a, Rib00]. Hessenberg [IA14, Li12]. Hexagon [Gup74, HH72, HH71, Moo73]. Hexagonal [JD22, O’D79b, Tan00, TS92, BPIK07]. hexagonal-square [BPIK07]. Hidden [Gup74, HH71, Moo73, Tre16b, Jag21]. Hiding [Hun74b]. Hierarchies [De 81a]. High [Ani95, Bic69, Coo19, MCPR95, Mas94a, lY08]. high-order [lY08]. Higher [Alo76, Bro69d, CSH72c, CSH72d, Car74a, Cha86a, CMX20, CBH17, EV04, Fox21, GM11, Gou81, How94, Iak77b, Iak81a, Iak81b, JV19, KY09, Kir77, KMP14, Lás82, Liu01, Mel10b, Mel10a, Moo93a, RMA96, Shi20, Som84, ZL95, KA13]. Higher-Dimensional [Gou81]. Higher-Order [Cha86a, EV04, Iak81a, KMP14, Liu01, Moo93a, Som84, JV19, RMA96]. Higher-Ordered [Lás82]. Hilbert [BJ09, TA02]. Hilton [Sha74c, Sha79a]. Hints [Tho63e]. Histogram [Hen19]. Historical [BJ14, Tue23b]. History [Ano63c, BJ87, Bro64a, Gou81, Høy09, Mit03, Mos08, Tue23a, FR85]. Hockey [Jon96]. Hofstadter [HM94, Hen15b, Rus11, Sto09b]. Hoggatt [BJ80, Bro69a, Car81b, FA89, HCG03, Sin90]. Hollow [HKTZ17]. Holographic [CLX08, CLX13]. Holomorphic [BRS04b]. Homeomorphic [Hol66b]. Hommage [Lan81]. Homogeneous [Giu72, Kla73a, Lee01b, MMMS22, MMM^{+}22, Maš09, Spi97, Vor09]. Homomorphic [Hol66a]. Homomorphism [Maš09]. Homomorphism-Homogeneous [Maš09]. Honour [Dij79, Dij78]. Horadam [Gra89c, Sch94, ACOM21, BL13a, BL13b, BL17a, BTD22, Bun12, Cas91a, CMM^{+}17, DV88, Gau98, Hau02, LB13, LBF14, Sha79a, Sha87]. Horner [GPS87, TMR06]. Hosoya [BE22, BFM19, BFMR20, CFL^{+}22, FJ12, NR17]. Hotel [Swa92]. Howard [CMM^{+}17, Joh22a]. Huffman [Kir94]. Hughes [Høy09, Oak09]. Human [DA79, YS08]. Humbert [MD87, WW17]. hundred [Hor75]. Hurwitz [SŠŠ22b, WTvR89]. Hurwitz-Type [SŠŠ22b]. hybrid [EGS22]. hydrocarbons [EBK89]. Hyper [KKLS18]. Hyper-Fibonacci [KKLS18]. Hyperbolas [CMS12, Hen78b, Kim90, Mel16f, Spo22a]. Hyperbolic [Ehr83, MS03, MS95d, Say76, Trz96]. Hypercube [Joh90]. Hypercubes [JHK97, LM85, EG89, QM90]. Hypergeometric [BSY05, Dil00, Lar16, Mil91]. hyperharmonic [DM08]. Hyperperfect [BN85, Gar04, Hag87b, Min81a, tR84]. Hyperspaces [Haa93]. Hypersurfaces [Hor86a, Hor86c]. Hypotenuse [Sha75b].

I. [Puc62]. ibn [Sha88a, Sha90]. idea [Sug93]. Ideal [EI03]. Ideals [Sha79b]. Ideas [Mos08, Cri07]. Identical [Lev68b, EBKT85]. Identically [SK15a, SK16]. Identities [Abd96, Abe18, Ade17, AM75a, AE22, AE23, ACF20, Art04, BQ99, BQS00, BR14, BCS19, BLM20, Ber78, BJE95, BJS96, Bic72b, Boy09, BG16, BB65, Bru73, Bum87, Car66a, CF70, Car70b, Car75c, Car78a, Cer17, Cha12, CB73, CK16, CMM^{+}17, Coo19, DD67, Du89, Ehr83, Fah16, FG05, FH88, Fil95, Fro18, GR09, Gau98, Ger09, Gir81, Gou72a, Gou72d, Gou74a, GBC16, Gri12, Gri14b, Han72, Han78, Har65, Hen17, Hen22a, Hic78a, HW09, Hir15a, HB64a, HB64b, HPL71, Hol66a, Hol66b, HS84, Hor71, Hor77a, HC11, Iye69a, Jen93, JD22, Joh40, Jon96, Kel92, KK11, Kim80a, Kis14, Kli81, KP78, KMP14, Kom14, Kos14b, Kos17, Kos18b, Kra90, LL13, LL85, Len06b, Lew80, Lin88, Liu01, LL05, Lon86a, Ma11, Mac10, Mag09, MS95g]. Identities [Mel99a, Mel99g, Mel16b, Mel16c, Mel16g, Pet88, PD18, PW15, Rob82a, Rob91b, Sco68a, Ser74, Sha74a, SW10, She11, Shi20, Spi06, Sto75, Swa66a, Swa97b, Ted19a, TJ88, Tue22, Vol13, Vse11, WH74a, WSP04, Wul81, YL97, YZ02, Zei63, Zei65, Zei67, Zei71a, Zha97a, Zha97b, ZL98, Zha98a, ZJ98, Zha04, ZW04a, ZW01a, ZW01b, ZW01c, ZW03, Bra06, Cha07, CH06, Fen11a, Gau11, Gri13, HHJM15, MZ07, Man05b, Sun06b, Wlo13, Woo07, YZ18, ZZ06, ZW06, dF14]. Identitites [Bed13]. Identity [Alb73, AML85, Asc74, AKP19, AKP20, BAS93, Bha16, Bru75b, Bun75c, Car74d, Car78h, Car81b, Che12, Chu90, CC20, Chu85, CS79, Com79, Cul76, EG16, Ewe82, Gou67b, GQ10, Hal65d, Han76, Hen22b, Hic78b, Hin80a, HL66, How01b, Hsu73, Kah09, LF14b, LF19, Lau10, Lin75, Liu09, Ma98, McC88, MS95b, Mel99e, Mel03a, Mel03c, Mel16d, Mun20, Nak14, PD18, Ros85, San90b, Sch17, Smi09, SR19, Sto70, TT16, Tue22, Vol10, Zei72, ZH03, Bea22a, Bri06, Gri17a]. Idiot [Boo75]. IEEE [IEE08]. If [SK12, SK13a]. II [Adl69b, AJ95b, AH70, AHS80, Arp04, BH63b, BJ93, Bre64a, BN19, Car70a, CSH72d, Car79a, Car80d, Car81a, Cha06b, CS81d, DMY13, Day69b, DMR^{+}16, Duj02, EA20b, Fau65, GG79b, Goo68b, Gua78b, Hic78a, Hil76b, Hir13b, Hog77b, HA77a, KPT83, Kom01a, KG18a, LJ70, Man67, MMM^{+}22, Mel00d, Mel01, Nyb01a, Pih10, Rob83c, Som91, SK16, Sta80b, Ste81b]. III [CS82, DMSS16, GG79c, Gua78c, Jes63a, Jes64, Lon81b, Man68, Mel17d, RH63]. illuminating [Gin02]. Illumination [MBR^{+}13]. Illustration [Sco68b]. Image [Gua78a, Gua78b, Gua78c, Kum21]. images [Dek20b]. Impedance [FFD92]. Imperfect [Fra77]. implement [Boy97]. Implementation [Pry94, VKSAN93, Mat90]. implementations [KTW84]. implemented [Boy97]. Implicit [Log81]. implies [Pen20]. Important [Jar67b]. Improved [FLP08, FT84, FT87, KM18]. improvement [Kah06, SYY04, Sub07]. Impulses [BH69a]. In-Winding [HA76b]. Including [Coh91, Kom14, Wiȩ80]. Inclusion [Kon10]. Inclusion-Exclusion [Kon10]. Incomplete [BTD22, Djo04]. Incongruent [Her81]. Increased [HBJ77a, KK01]. Increasing [Lar15b]. Incredible [Bum87, Sha74a]. Independence [BRGGGS20, DT20]. Independent [DCB93, ESW13, KC86, Ste01, Zel93]. indeterminate [Pic79]. Indeterminates [Hig87]. Index [Ano63d, Ano64d, Ano65f, Ano66c, Ano67b, Ano68c, Ano69d, Ano70g, Ano71d, Ano72h, Ano73g, Ano74d, Ano75b, Ano76d, Ano77d, Ano78b, Ano79b, Ano79c, Ano80c, Ano81, Ano82d, Ano83c, Ano84d, Ano85d, Ano86d, Ano87e, Ano88g, Ano89f, Ano90g, Ano91g, Ano92d, Ano93d, Ano93l, Ano94f, Ano95h, Ano96e, Ano97i, Ano99f, Ano00g, Ano01d, Ano02f, Ano03e, Ano04c, Ano05b, Ano06b, Ano07c, Ano09b, Ano10d, Ano11e, Ano12b, Ano13b, Ano14e, Ano15b, Ano16e, Ano17l, Ano18g, Ano19i, Ano20g, Ano21d, Ano23, Cam22b, Hos73, MKL^{+}23, NR17, Phi07a, Van12, Jar06]. Index-doubling [Phi07a]. Indexed [Cam23, Ess03, Pil12]. indexes [FP09]. India [Sin85]. Indicator [Chr84, EGS22]. Indices [Bro72g, Dei66a, GKL^{+}21, RP19, SK13c, GGL24]. Induced [De 81a]. Induction [Hay92, Puc62]. Inductively [Hol66a, Hol66b]. Inequalities [AKO^{+}03, DL92, Jar64, ST90, Puc62]. Inequality [But73, GR93, Zei78]. Infectious [DS96]. Infinite [Ade17, AJ91c, AJ98, BR85, BH79, BMRS01, Bro69b, Bro69h, Bru77c, DP16, DH67, DeL95, Dra64, Dra69, DT20, EH22, Fil00, GCMP14b, Gri18, Hil76a, Hil76b, HS82, KM14, Kos21g, Kos21h, Kos21i, Kos22e, Kos22c, Kos22d, Kos22f, Kos22g, Kos22a, Kos22b, Kos23g, Kos23h, MS95f, MS92, Moh78, Nyb04, Pro16a, Red94, SHB84, Sil91, Ter12, Tri74c, Zha00, CH05, CLHY12, GCMP14a, GSS19, Hol95, KA13]. Infinitely [Du89, McD87a, Sto70, ESW07, Pen20]. Infinitude [Bru94d, DJ68, Leh64, Rob05b, SW02, Weg81b]. Infinity [MPPW22, Sha05]. Influence [VKSAN93]. Information [Lyn73, Pro94]. Ingmar [Gal08, Har09]. Inhomogeneous [Lee97]. Initial [Dav76, Dun69a, Jen71, Kim78, Kno81, Sli98, Spi97, Whi78e, BHS11]. Initial-Value [Jen71]. Injectivity [dBL83]. inner [Zah82]. inradius [Rob06b]. Inscribed [Hun63]. Insects [De 78b]. inserting [Car39]. inserzione [Car39]. Instability [Ove65]. Instance [Bru94a]. Institute [Bro96]. Instruction [Sar90, PC13]. Instruments [Wil83]. Integer [Alp11, AC72, AZ20, BJE95, BJ02, Bro75a, Bro63b, Bro78b, BL17b, Cad71, Cam23, Car77c, Col83, Cus68, Di 13, Dil87, Eme66, Esw79a, Eva77, Eve75, Fen03, Fie68b, HZ17, Hal65a, HKN10, Her19, Hin78, HBJ79b, Hon15, JS16, Mab19, McD82, McD94b, ML84, MS94, Mel16f, Nyb03, PP20, Tee94, Tri09, Tri10, Tsu21, TS89b, TR90, Wal76a, Zha97b, ZL98, Dek20b, Duv91, HKN11, KKLL21, WW06]. Integers [Alf63f, AJ91e, ABN^{+}19, Bro03, Bro72f, Bro69i, Bro75c, CW22, Cap88, Col83, Col94, DH67, Dul20, Dun69b, Ehr84, Erd81, FG23, Fer65, Fil86b, FB99, Fox21, Gal72, Gau04, Goo86c, Gou71a, Gre09, Guy82, HW12, Hei79, Hen78a, Hig72, Hil81e, HBJ79b, Hog80, HBJ84, HJ68b, Hor78c, HS81, HS82, Hor94b, Hor94d, How85, How96, How98, HM72, Ibs90, JM14, Kim95a, Kim01a, KK89, Kyr94, LaD18, Laf64, Lud87, Lud88a, Mai70, Mar12b, McN92, Moh81, Mon74, Mon88, Nyb01b, Nyb02, O’C72, Pla94, Pla97a, PB89, Puc02b, Rap74, Sal75, SS01, Ste76, Sto67, Tan75, Ter96a, Tót02, Tri65, Tri71, Wan93, Wei79, Zha01a, EH07, EZ04, Kim07, LPW11, Mag09, Wil87]. Integrable [Gry03]. Integral [Bis84, CY91, DES17, EV89, FF89, Gal68a, Gal68b, GZ15, Gle81, Gou73, Hah72, HK89, PP14, Pon17b, Som79, SSS24, Wat88, Yam80, ESW07, GB06, MH16, Ste23]. Integrality [Cal00, LAL16]. Integrals [GM20, MBR^{+}13]. Integration [HSF94, NS94]. Integrity [AJ98, FB94, Zha00]. interconnection [Hsu93c]. Interdisciplinary [Mos08, B^{+}96]. Interesting [Bru72, Bru96, DD02, HK75, HP75b, HM72, Kos15b, Lay75, Moh76, Ter12]. Interests [Man64, Man67, Man68, Man77]. Interferometers [Bot82]. Interlacing [Kim10c]. International [ACM96, ACM12, Ano87c, Ano87d, Ano88d, Ano88e, Ano89d, Ano89a, Ano89b, Ano89c, Ano90b, Ano90d, Ano90c, Ano90a, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano93a, Ano93b, Ano93c, Ano93k, Ano94c, Ano95e, Ano95f, Ano95g, Ano96d, Ano97a, Ano97c, Ano97d, Ano98a, Ano99a, Ano99b, Ano00a, Ano00b, Ano00c, Ano01a, Ano01b, Ano02a, Ano02e, Ano03b, Ano03a, Ano09a, BHP92, BPH93, BJ18, Coo10, Coo16, Fre87, Fre88, Fre90, Fre97, Fre99, Gon89, Jef08, Joh09, Joh12b, Joh20, Joh22b, Mos08, Phi01, Phi04, Phi07b, dBF85, BB79]. interpolation [LMV09, BSY05, DP78a]. Interpretation [AB20, EA20b, EA20a, Hod74, KL91a, Sir97, Bra06, EA19]. Interpretations [Aga91, BG16, Goo68a]. Intersecting [BHT10, Oll22]. Intersection [CD87]. Intersections [Sha83, Ste73b]. Interspersed [Kim10a]. Interspersion [Kim94]. Interspersions [Kim10a]. intertwined [Hir05]. Interval [Har02, Her03, Kom03]. Interval-Filling [Her03]. Intervals [Adl78, BDE^{+}14b, DH67, DH68]. Intervenes [Gou74a]. Intrinsic [DS04]. Introducing [Wir08]. Introduction [Ano14d, BHL^{+}20, RF72, SMS99, MSPF95, RS08, Mar93, Pic79]. Invariance [LF15, Lar15b]. Invariant [Ber78, CHKS05, HS19a, KK95, Sun01a, Tur86, AAF15, Igu94]. Invariants [Bas81, Bra90]. Inverse [DD13, GK76, Gou94, Hsu87, MH85, MS95d, MS92, Mun20, Sho99b, Tue21, Wan05, Gou06, JH15, LMV09]. Inverse-Conjugate [Mun20]. Inverses [HML01, HB76a, HB76e, Pla97b, Roc81, Tri00a, BT12, SCH11]. Inversion [Hau90, Hsu95, Kra90, NS99, BHS06]. invert [Høy05]. Inverting [GQ11]. Investigation [Bro75a, Gel63]. Invitation [Øre67, Bro69a]. Involutions [KP78]. Involutory [CP03]. Involve [Mel00c, Mel01, Mel04a, Mel15, Mel17a, Mel18c]. Involving [Ade17, Alf64i, AE22, AE23, AGJ^{+}09, BH78a, Bic72b, Bro74b, Bro68g, Byr65, CK24, CD13, Els05, Eva77, Fah16, Fil00, Fre68, Gau98, Gla95, Gou63c, Gri11d, Har65, HL09, Her03, Iye69a, Iye69d, Jai69, Kla66a, Kos21g, Kos21h, Kos21i, Kos22e, Kos22c, Kos22d, Kos22f, Kos22g, Kos22a, Kos22b, Kos22h, Kos22i, Kos22j, Kos23b, Kos23c, Kos23d, Kos23e, Kos23a, Kos23f, Kos23g, Kos23h, Kos23i, Kos23k, Kos23j, Kos23l, Kos23m, Kos23o, Kos23n, Kos23p, Kos23q, Kos24a, Kos24b, KG24, Kos24c, LB13, LBF14, Lin71, Liu01, LL05, Liu09, Luc00, Mac10, MH85, McI20, MS95d, Mel99a, Mel03c, MS92, MC04, Mon88, Nef67, Nyb99b, Nyb01a, PP20, PRW93, Sch20, Ste88, Ste21, Sto67, Swa97b, Tri00a, Wal79, WB12, WD76, Wel67, YL97, YZ02, Zei77, Zha97a, Zha97b, ZL98, Zha98a, ZJ98]. Involving [ZW02, Zha04, ZW04a, ZW01a, ZW01b, ZW03, ACOM21, Blo53b, Bra06, BM07b, CH06, CV86, Kel57, Kel58, MZ07, Ma11, Mag09, Man05b, Rao55, Rei54, ZW06]. Irrational [Els96, HS81, Kim95a, Kim98b, Kim01b, Kom01b, Kom10, Nyb99b, Nyb01a, Pau23, Sha75a, Sub95]. Irrationalities [ABCM12]. Irrationality [AJ91c, McD94a]. Irrationals [All75a, BCHR91, Bun12, BK96, HB77b, LM10, PVB98]. Irreducibility [BH74]. Irreducible [Lev01]. ISBN [Gal08, Hal22, Høy09, Mos08, Sch94, Tol17b, Tol17a]. Ising [Tra88]. Island [IEE84]. Isodecimal [Man06]. Isolated [AG78, GM96]. isomorphism [JV08]. Isomorphisms [Ste76]. Isosceles [HJ68a, Nyb98]. ISSAC [Gon89, Jef08]. Issues [Min81a]. Italia [RS98b]. Italian [Ago49, Arr69, Car39, Fra40, Mar74, Pic81, Pic83b, RS98b]. Italy [Fra03, PB15, RS98b]. item [Boy97]. Iterated [CS80, CS81d, CS82, GG13, Lin98, Lin67b, MV84, Ste85, Ste88, Wie88, Sal12]. Iterates [Kim91a]. Iterating [May87, Wag81]. Iteration [And00, Bur90, Kar69, Kar74, PT18, Rat93, SK20, CM07, Cla08, Joh04]. Iterations [Cop18, Fra91]. Iterative [Des71, Tan98, KSSS16, Sho78]. IV [BH63d, Man77, Sin73a]. IX [BH71].

J [And05, Ano90e, BL13d, Moh81]. J. [Rob02a]. Jacobi [Ark66, BH69a, Ewe03, JH05, TT16]. Jacobsthal [ACOM21, BBHM85, BT12, Djo00a, GB15, HBJ78a, Hor88b, Hor96b, HF97, Hor97b, Hor97c, Hor02a, KG17c, Kos19b, KG19, Kos21f, Kos21i, Kos22e, Kos22f, Kos22g, Kos22i, Kos23f, Kos23h, Kos23o, Kos23n, KG24, Swa99b]. Jacobsthal-Type [Hor97c, Hor02a]. Jacobthal [Djo00b]. Jarden [Bro67b, Bro66, CK95]. Java [Bee04]. Jeannin [Swa00c]. Jennifer [And05]. Jerrard [EF98]. JoAnn [Ano03d]. John [Mos08, Ano15a, Boy97, Slo15]. Johnson [Ano10b]. join [Sar90]. Joint [Ste05]. Jordan [ACF20, Sho99a]. Joseph [Bro70a, Gla70, Ped71]. Journey [Ber11]. Jr. [BJ80, Bro69a]. Julia [BJ09]. July [BHP92, BPH93, Bro96, Gon89, Jef08]. Jump [Hen14, Pul94]. June [NS95]. Justified [PS02]. Juxtaposition [Pad94].

Kāmil [Sha88a]. Kahya [Sub07]. Kalman [AAF15]. Kamil [Sha90]. Kaprekar [PLL81, Tri71, You93]. Karajī [Sha88a]. Karaji [Sha90]. Kátai [Pih92]. Keith [Sim18, Tol17b, Tol17a]. Kekulé [TB91]. Kennedy [Pro97]. Kenney [Moh81]. Kentucky [CFH^{+}14]. Kettenbrüche [Lün91a, Lün91b]. key [LLP^{+}18, Mül06, Smi93, SS95a, SS95b, XLP99]. keystream [And95a]. Killer [GP94]. Kimberling [Beh12, Hen11b, Szi17]. Kind [Aga90, Cak98, Car80c, Car80d, Dav14, Di 93, Fra91, Gau04, Hor77b, Hsu93b, Hsu97, Jai74a, KK95, Kun90, Led67, Lee02a, Len94b, Par77c, Par77b, PRW93, Sit70, AD10, WP07]. kinds [DMP07]. Kluwer [Sch94].

Knapsack [HKW15]. Knapsack-Like [HKW15]. kneading [BV05b]. Knife [Erc72]. Knight [CD78]. Knots [KP10b, Tur86]. know [Cri07]. knowledge [Fol04]. Known [Bol77, CBH17, Du00, RT02, Wol99, WK09]. Korovkin [Puc62]. Korselt [LS14a]. Koshy [Gau02, Gau10]. Kostovskii [Puc62]. Kraft [GR93]. Kronecker [Law83, Spe77, Wil76]. Kummer [Bal15, Gil92a, Ma98, Met14, Ste86b]. kung [Bic72a]. kx [McD02].

L [Ano88c, Hor04, Kui69, Zho03, ZH03]. L. [Jen66, LR85, Lee87, SH82]. Labeled [Leg15, MH75b]. Lacunary [How98, MS20, Sha80, You03]. Ladder [Ark65b, FFD92, Ris82]. Ladders [Ste01]. Laden [Fre87]. Lady [Bri64a]. Lagged [Alu96, Alu97, Hey81, Mak94, MCPR95, MRPC95a, MRPC95b, Pet94, AM15, MS04, Mas94a, Mas94b]. Lagged-Fibonacci [Mak94, MCPR95, MRPC95a, MRPC95b, AM15, MS04, Mas94a, Mas94b]. Lagrange [BT00, Kra90]. Lagrange-Type [Kra90]. Laguerre [AS84, Djo01a, Moh76]. Lah [AE79, ND87, Tau68a, Tau68b]. Laid [Lev84]. Lainiotis [AAF15]. Lambda [BH63c, MP66, SH02]. Lambert [AJ90, Bra72, Hor88a, Mel99b]. Lamé [Bro67g, Moo89]. Langford [GU66, Lev68c]. Language [AGO95, Dek20b, Zei96]. Languages [HSS03]. Laplace [Giu71, HH00]. Larcombe [PW15]. Large [CP79, DD95, Ibs90, SK21, Wei73, Zha01a, HP92, Tak00, Yan95]. Largest [Cos02, Wal87b, BL13e, RL15]. lasses [AH97]. Last [CS81a, Hir15c, Jar63b, Lin75, Mil91, SW10, Win11, Zei74]. latency [PC13]. latency-adaptive [PC13]. Latin [AS74, Ark74, AS76, AS81a, AS81b, Kar96, Ker82, LM85, Lig73, Lig79]. Lattice [Aga87, BCC^{+}20, CG67, Chu74, Ege83, Eng90, FJL^{+}20, Gre64, HM83, McQ76, Oll22, Sta75, Sto65, Sto67, Woa02, Zar70, NS94, TLC93]. Lattices [CCG^{+}19, EI03, Fuc84, HH85, BDL^{+}16, Igu94]. Law [Bra78, Gre77c, McL79, Mit03, Tay75, IT91, BGTM^{+}11, CHL19, Gil05, Gil07, Pet81, RM11, Sen73, SB93, Was81, Wlo71a]. Laws [Sch91]. Layman [Car98]. Lazy [Ste05, Gru07]. lazy-Fibonacci [Gru07]. LCM [AS05, Bes91, Chu96]. Leaf [Dav71]. leafy [THKH23]. Leap [Sha94]. Learner [Cur68]. learning [Ola00]. Least [BJF01, BJ02, Bro75a, BD71, Edg86, GTNP96, Jar68a, Moo92, Wal88]. Leaves [THKH23]. Lebesgue [Süt89]. Lectures [Puc62]. Ledin [Hen22b, SO21]. Left [BJ89, BJ93, Lin90]. Legacy [CS12, Teu04]. Legal [CFH^{+}16, CFH^{+}17, MKL^{+}23]. Legs [Weg75]. Lehmann [Gal08, Har09]. Lehmer [Bri83a, Bri83b, Bri83a, Bri83b, Bru93, CJS98, Fil97d, HS15, Jar06, Kra70, LP03b, McD91, Ros88, Sch13b, SM93, SK13c, SK15b, Tri73, Vou95]. Lekkerkerker [Pih88]. Lemoine [Pih92]. Lempel [JMNRS22]. Lenard [How00]. Length [CST05, Cha91, Cre87, Cre88, LY90, Lud88b, MM73, WB12, Web82, Krá72a, Krá72b]. Lengths [MC04, Nak20, Pri05, WP19, YK77]. Lens [Hir11b]. Leonard [GG69, GGA83, Spi00a, Spi00b, Bro70a]. Leonardi [Lün92]. Leonardo [Anoxx, Geg93, Hor04, Wes80, Ago49, Ago53, AV03, Arr69, Bab41, BF90, Baš78b, Baš78a, Beč01, Car39, Cho82, Dev12, Dij81, Fol04, Fra40, FR85, Fra02, Fra03, Fre22, Geg93, Gie13, Giu04, Glu76, Gri73b, Han07, Høy05, Kin63, Lor29, Mar74, McC19, Miu81, Miu02, Pep03, Pic79, Pic81, Pic83b, Pic83a, Sha88a, Sha90, Sig02, Spi00a, Spi00b, Uli11]. Lesevergnügen [Lün92]. Less [Kre80]. Lesson [Bro68b, Bro68e, Bro69g, Bro69e, Bro69d, Bro69f, Bro70c, Bro70b]. Letter [Abe75, Ano73e, Ano75a, Ano76a, Ano76b, Ano76c, Ano77c, Bau64, Ber85, Ber86, Bev70a, Bev77, BJ86, Bri83a, Bri83b, Bro63a, Cat09, Coh64a, Coh73, Dre80, Erd74, FR66, Gou76, HL84, Hal65b, HF86, Hil87a, Hin97, HL90, Hor91, Hsu93a, Jam77, Jun61, Knu74, Kui76, McC86, Nao65, Rob82b, Rus80, Sha89, Sti00, Tri67d, Wei66b, Zei71b, Zei73, Zie63]. lettere [Ago49]. Letters [Ano70f, Ano73f, Ago49]. Level [EJ17]. Levels [HCG03]. Lexicographic [HBJ82b, HO87]. L’Hôpital [Sco68b]. Li [Bic72a, Bic73a]. Li-kung [Bic72a]. Liber [Cio18, FGD20, Hor04, Lün92, Ses14, Ano93i, Sig02, Geg93, Ago49, Arr69, Baš78b, Baš78a, Dev12, Eis51, Fra02, Fre22, Han07, Han11, McC19, Miu81, Pis02, Spi00a, Spi00b, Geg93]. Library [Arr69, Pic79, Pic83b, MS00a, MS00b]. Lichtenberg [Hin17]. life [Bic72a, Cho82, kS74, Ste11b, Teu04]. Like [AML85, AC11, Asv87a, Asv87b, Asv89, BCH12, Bas64, BBHM85, BCFX09, CC03, Cer17, CZ10, Edw09, Fie67a, GCMP14b, Gri11a, HKW15, HG81, JK97, JQ97, Köh85, Mor89, Mos93, NW83, PP19, Pha84, Pla94, Pla95, Šiu11, Som77, Som82, Ste01, Tay67, Tur89b, Vse11, Wal84a, ACOM21, BV05b, DRS11b, DRS11a, For67a, GCMP14a, IK87, IA14, Knu90, MMYA10, Mor95, NW79, Osl07, PP14, Rit23, TKS11]. Limerick [Alf65b]. Limit [FJL^{+}20, Hin77, LM17, LM19a, Moo94, Nef67, YWH96, BK16]. Limitations [Fie66]. Limited [LJ67, LJ70, Lon81b]. Limiting [GCK^{+}19, HA77b]. Limits [BH77b, Kim12, Yan88]. Line [Car73, CH74, Lee97, Lee01a, Lee01b, Lee02b, Met97]. Line-Sequence [Lee97]. Line-Sequences [Lee01a, Lee01b, Lee02b]. Line-Sequential [Lee01a]. linéaires [Spi00a, Spi00b]. Linear [Alf65e, Alp19b, AB85, Ani95, Bab76, BKM17, BCFS22, Ber68, Big21, BHL^{+}20, BN03, BN19, Bro68b, Bro69g, Bro69e, Bro69d, Bro69f, Bro70c, Bro70b, Bro72a, Bro76, BsS95, BW81, Cal00, CFH^{+}16, CFH^{+}17, Drm93, DT20, ER00, Far86, Fer78, Fra94, Fre85, Geo89, Giu72, Gol95, Goo68a, Goo68b, Gou71a, Gre68, HW12, Han78, Hen06, HH79, Hor92c, HCD97, Izo02, Jen71, Jes63b, Jes63a, Jes64, JK93, JS19, Kim80c, Kim91c, KM19, KM05, Kla68a, Kla73a, Knu66a, Kos19a, KW13, KYS13, LF14b, Lar15a, Lar16, LO18, LF19, Lás82, Lee97, LK03a, Lev85, Li99b, Lit20, Liu92, LHLT95, MMMS22, MMM^{+}22, MO09, MO10, MJ95, MZ92, MS79, NS90, NR15, Pla96a, Rab96a, Rab99a, Rob82a, Rob91b, Rob02b, Rob04a, Rob76, RM11, Sbu02]. Linear [Sel84, SY84, Sha74b, Sha83, SO02, She72, Sin70, Som79, Som84, Som89, Som91, SK15a, SK16, SK17, SK21, Spi97, Sun01b, Swe64, Sza17, TR04, Tan98, TT16, Ves15, Vor09, Whi67b, Zöl93, Bab41, BB19, BN06, GPS87, Knu66b, Lay66, LK03b, Sla16, Som06b, Som07, Sun06a, lY08, Spi00a, Spi00b].

Linearization [TMR02b]. Linearly [BD81, CK24, Hen11a, Rap74]. Lines [Ben93, Fer97, Ste73b]. link [Sub19]. Links [FFD91]. List [Ano17f, Ano22c, BV05a]. Literature [Gou72d]. Little [Bri64a, Gil89]. Lives [Acz11]. living [kS74, Bic72a]. Load [LHWS96, LJ97]. Local [dOGC96, Mul80b, Mul80a, Stă14, CHL19, LP14]. Locating [CH05]. Log [BRGGGS20, Er83a, GL80]. Log-Concave [BRGGGS20]. Logarithm [Bro17, Len94a, VPB99, LMV09]. Logarithmic [Dei66b, Ron91]. Logarithms [BD70, Kui69, KsS73, SS95a, SS95b]. Lognormal [Mit03]. London [Sch94]. Long [BP10a, Bot82, BW97, Dun09, Jar67a, Mel99g, Mos93, Rie93, Ano02b]. Longest [Cha91, PM85]. Look [Cal97, Gri73a, Nod90]. Look-Ahead [Nod90]. Loops [Maš09]. Lopsided [KR89]. Lost [Alf65d, HS13, Coo14b]. Lottery [Hen95, KC91b]. Low [TF92, TF93]. Lower [Cat74a, EKS88, Fos66, Hag84, Kut02]. LUC [Smi93]. Luca [BF90, FR85]. Lucas [Bal19a, BT12, Bro69a, FH94, Fro18, Gau02, Goo93c, GP93, GS17, Kah05, Sol07, UNS07, PC13, AFRT23, Ade17, AH15, Alf63a, Alf64a, Alf64g, Alf65a, AJ90, AJ91a, AJ91b, AJ91c, AJ91f, Ant85a, Arn97, AC13, BFF11, BF16, BW80, Bal19b, BB69, BV05a, BB85, Bay08, Bed13, BQ99, BH74, BH79, Bha16, Bil14, Bis84, BBL95, BB90, Bra71, BBL14, BB65, BMS88a, BMS88b, BG21, Bro69b, Bro70d, Bro69i, Bru93, Bru94d, Bru94b, Bru96, BR82, BB11, BK96, Byr65, Byr75b, CDS03, CN04, CK10, CJ96, CH69, CF70, Car72b, CSH72e, CHS72, CKMR11, Cha91, CL11, CL12, CLM22, Chu74, CS85, CDM00, CT18, CMS12, Daf20, DFLT14, DD67, Day69a, Daz02, Dda20, DOU08a, DOU08b, Dem08, DD13, Des78b, Des86, Di 13, DM10]. Lucas [DGMS14, Djo00a, Djo01b, Djo04, Djo05a, DX22, Dre97, DK99, DT20, Edg16, EKS88, Esw78b, Esw79b, ES14e, ES17, Fee67, Fei67, FZ03, Fer69, Fer97, Fie67a, Fil91, Fil93, FH93, Fil98a, Fil98c, FH98, Fin75, FH20, FPC08, FF89, Fro18, FZZ08, FZ08, GCMP14a, GCMP14b, Gau11, Gla78, Gla95, Goo93b, Goo93a, Goo94a, GP91, Gou77a, GG85, Gre18, Gri11b, Gri13, Gui20, GTU13, HS12, HS15, Han72, Han78, Har98, Hau96, Hau94b, aHZ02, Hil74e, Hil74f, Hil74g, HPV95, Hin80a, Hog69, Hog70a, Hog71b, HBK72, HP75a, HBJ77b, HBJ77e, HBJ80, Hol94, Hor77a, HM85, Hor86a, HS88, HSF94, HF95, HF97, Hor00b, Hu02, Ila71, İpe11, Ism09, JP19, Jar63a, Jar64, Jar67b, Jar68a, Jen93, Jen94, JLW07, Jon76b, Jor65, Jos83, JQ97, Kap45]. Lucas [KY09, KK15, KPAO11, KA19, Kim79a, Kim80b, KM95, KTW84, Kom09, KL91a, KL93, Kos01, Kos14b, Kos17, Kos18b, Kos19b, KB64, Kra70, KT19, Kut02, Kwo07, Kwo14, LaB71, LW81, LTT95, LMV09, Lee00, Len95, Lew91, LM93, Lin67b, Liu09, LF69, Luc00, LS06, Ma98, MZ07, Mac00b, Mac88, MS20, Mar12c, Mar13a, Mar15, McD91, McD94b, McD94a, McD95, McD98b, McD01, McI92, Mel97, Mel99a, Mel99c, Mel99d, Mel99f, Mel00a, Mel00d, Mel04a, Mel09, Mel16c, Mel17d, Meš13, MK02, MV15, Mor75, MCS01, Mus93, NC85, NKW23, NR17, ÖTA05, OP18, Osl07, Oze05, PG13, PP00, PT18, PP20, PPK21, Pet84, Pet85, PH88, PD18, Pho91, Pih88, Pon17a, Pon17b, Pon19, Pop85, Pru67, PW01, RT72, Ram14, Rao02, RP18, RP19, Ric93, RFLT20].

Lucas [Rob81, Rob83a, Rob91a, Rob05a, RW21, Ros88, Ros72, SE20b, San16, SBM97, ST06, Sha74c, Sha78a, SM93, Sha13, SC10a, SCH11, Shi20, Sin80, SŠŠ22b, SS95a, SS95b, Sol05, Som80, Som06a, Som09, SC10b, SK13b, SK13c, SK15c, SK20, SK22a, Spo22a, Ste78b, Sto75, Str99a, Str99b, Sun06b, Sur14, Swa97b, Swa99a, Szi17, TK04, Tau68a, Tue21, Vaj89, Vaj08, Vou95, Wal76b, Wan95, WCL99, Was81, Wel94, Wil77, Wil83, Wil98a, Wlo13, Wlo71a, Woo97, Wul81, WSH74, XLP99, Yab02, Yab07, Yam19, YZ18, YZ17, YK03, You95b, You06, You17, YL97, Yua02, Zah82, Zec72b, Zei64, ZJ98, Zha01a, Zha04, Zha99b, Zha01b, ZW01c, ZW04c, Zho96, Zho99, ZJ14]. Lucas-Balancing [PP20, PPK21, RP18, RP19]. Lucas-Based [BBL95]. Lucas-Like [Fie67a]. Lucas-sets [Zah82]. Lucas-Type [AJ90, FH20, Hol94]. Lucasian [BB04, HPS97, Rie69]. Lucasnomial [Bal17]. Lucasnomials [Bal15]. Lyndon [Bas20].

M. [Lax74, Moh81, Pro04]. Machin [Cha09]. Machin-Type [Cha09]. MacMahon [CR72]. Madachy [Bus70, Gla70, Ped71]. Madrid [KTW84]. Magic [Alf64c, AH81, Ano16a, Ark73b, AS76, AS81a, BCH12, Bro72b, Bro65b, Chi86, CBH10, Fre68, Hah75, Kar70, Kar72, Nor16, Pad94, Pad97, Rot73, Tri74a, Wlo65]. Magicness [CBH10]. Mahameleth [Ses14]. Mahon [DV88]. Maine [Bur92]. make [Rob99]. Makes [Ano72f]. Making [Erc72]. man [Dev11]. Management [LSS76]. Manager [Ano97h]. Mandelbrot [CC16, Dev99, Rat93]. Manifestations [Gou77a]. manifold [MT17]. manifolds [CS92]. Manipulation [Rab96a, Rab99a]. Mann [Bol89, Gou72c, Gou74c, GG85, Gou89]. manuscript [Fre92a]. Many [Du89, Jar67a, McD87a, Raw89, Sto70, ESW07, HHP02, Pen20, Rob99]. Map [WP19, Zar09, CM07]. Maple [Yan95]. Mapped [HSS03]. Mapping [Bra95, Goo91, JK97, KISS15]. Maps [GKRS01, GG13, Lin96, Lin98, MZ92, Sta82, Igu94]. March [SC07, SM08]. Marjorie [Ano10b]. Markoff [Coh79, KST20, Sri20]. Markov [CMR02, LS18a, MR99, MR00, Nef67]. Marktoberdorf [Bro96]. Married [Sto09b]. Marriott [Swa92]. Marsaglia [PW95]. Martin [Izo99]. Masked [BL17a]. Mass [Ano88c]. Master [HPL71, Hor91]. masters [Uli11]. Match [HB72a]. matching [Bun06, EBKT85]. Matchings [Far86, TS92]. Matematiche [Cio18]. Math [Alf65b, Ano16a, GP93, Sol07, UNS07]. Mathematica [Mae92a, Mae92b]. Mathematical [Alf64j, Ano95c, Bic72a, Buc64a, Buc64c, Buc64b, Bur92, Fed64, HHP02, Hor63b, Hor91, Mam61, Mil71, Puc62, SH73, Sha60, Ste95, Tri67e, Win68, Cri07, Dev17, Fol04, Giu04, Hun70b, Pet99, kS74, Sim18, PB15, Eve68]. mathematician [Lün92]. Mathematicians [Acz11]. Mathematics [Ano22d, Bic68b, Gla70, GKP89, GKP94, Her19, Høy09, Lew70, Mad68a, Mad68b, Mad68c, Mad68d, Mad69, Mar93, Ped71, Puc62, Sha05, Ste11b, Sub07, BF90, Bro70a, Dev12, GG69, GGA83, HD19, Mar92, SO09, Puc62]. Mathematikers [Lün92]. Mathématiques [AV03]. Mathieu [Spi00a, Spi00b]. MATLAB [Har02]. Matrices [AS20, Aus23, BH80a, Bea18, BH78a, BH78b, BBHM85, BH69a, Bes91, BL92, BJS96, BH63c, BFM19, BFMR20, BN06, BC77, Cal97, Chu96, Daz99, Daz02, DeC73, Er84a, Erc76, Fer78, Fil92, Fil97a, Gal68a, Gal68b, GBC16, Gue78, HML01, HB76a, HF91, Kar96, LKL02, LK03a, Leh75, LY02, MH87d, MT85, MS95g, Mil60, Mor89, NS05b, Nor16, Pas10, PW93, SO02, SS17, Sun01b, TA02, Veg74, Wad74, WH71, Wal76b, Was99, YL04, ACOM21, BT12, CC16, DGMS14, Duv91, FZ08, İpe11, JH15, Kar04a, Kar05, Li12, LO14, MSS08, SC10a, SCH11, Sol05, Sol07, UNS07, Yor14, ZJ14]. Matrix [BB69, Bic65, Bre64b, Bre64a, CN04, CP03, Cam22a, Car65, EMR02, Er84b, Erc79, FH88, FT14, Gou81, Gre18, HH65, HB76b, HB76e, Ima98, Ivi72, Kal82, Kal90, KP10a, KP13, KP69, LF15, Lar15b, LLS97, Lin67c, Liu92, LP96, Lun77, MH86, MP66, MS95f, Mel99c, Mel99d, MC00, PS02, Pla96b, PS80, Pro22, Ric01, Rob63, Sma77, Zha99a, Zho03, Chu10, CT18, EB17, Fen11a, Ila71, Kil07, KA19, LK03b, Ogu20, Pet80, TR03, TMR06, TD16, TD19]. Matrix-Based [Cam22a]. Maximal [Ala98, BBS81, EBJ00, ESW13, Hor94b, MZ92, MN76, SH79c, Ste01]. Maximizing [Has81]. Maximum [Fos66, Jos79, OW64, Sal75]. May [ACM96, ACM12, IEE81, EG89]. Maze [Pau02]. MCMC [Hly17]. Mean [All77, Bru77a, DA79, Dek20a, DvL23, HPS14, Lin96, LL84, Sil98, VWT85, WTvR87, Car39, CHF85, Fis96, KN08]. Means [Fie64, Fie68a, GW19, NKW23, Sch81, Sch84, TT16]. Measure [DP78b, Süt89]. Measures [SS91]. Mechanics [Fer78, Puc62, Zen78]. Media [But90]. Mediant [Adl78]. mediante [Car39]. medie [Car39]. medieval [Lün91a, Lün91b, Sin85]. Medium [DD95]. Meeting [Ano69a, Ano70a, Ano71c, Ano72g, Ano73d, Ano74b, Jef08]. Meets [CK12, Gui20, Lan04, Rus11]. Meixner [CS81c]. Melham [Fil98b, KAP10, Oze09, Pro09]. Member [LF14a]. Members [DFLT14, DJ67, Fie66, GM96, Jar67a]. Memoriam [Ano99d, Ano07a, Ano21c, BJ80, How00, Joh22a, Rib00]. Memory [Alu97, AM15, Ano96b, CLRS86, Fer76, Fre87, Hir73, LSS76, Mat95]. Memory-Laden [Fre87]. mergings [AC87]. Mersenne [Kap45, KB64, Kra70, LJ82, Wol99, WK09]. mesh [HP92]. Message [Ano00d]. Meta [CCT05, DRT11, DB22]. Meta-C-Finite [DB22]. Meta-Fibonacci [CCT05, DRT11]. Metallic [GW19, NKW23]. Metallonacci [AE22]. Method [Alf65e, All75c, BRS04b, Bro75c, Col83, Du89, Fil86a, GM81, Gup76, Hen18, Hen22a, KYS13, Kyr94, Liu92, Mea65, Ove73, Puc62, Rot73, Sha74d, SW10, TMR02b, WD76, DOU08b, Dem08, EE10, Kah05, Kah06, KŞ13, MH16, PG13, Pet80, Rou13, SYY04, Sub07, WCL99, YK03]. Methods [EH95, EMR02, Er84b, Gou74b, Kal82, Kal90, Ove65, Glu76, Kil07, LMV09, Mus93, NS95, TR03]. Metric [All77, Els98, Sha86]. Metrics [DL92]. Metrod [McC88]. Michigan [IEE81]. Microcomputer [Moo89]. Middle [Bal17, Bro70a, GG69, GGA83]. Milan [Pic83b]. Milano [Pic83b, Pic83a]. Mine [Alf65d]. Minimal [dOGC96, BV05a, Dez77, HLW94, Hor94d, MN76, Pih92, RW81]. Minimax [OW64]. Minimum [Her78, Her82, KK95, Kwo89b, Kwo89a, Now86, Wul75, Boy97, BDL^{+}16, Tre96b, Tre96a]. minimum-area [Tre96b, Tre96a]. Minmax [Hor96c]. Minoan [Pre68]. minor [Sch50]. Minus [BJ90, Gel81]. MINX [PHJ67]. Miscellanea [Gil92b]. Mitchison [Rob78b]. mittelalterlicher [Lün91a, Lün91b]. Mixed [Dor93, HM87, Kim80d, McQ76]. Mixing [Kim80d]. Mixture [Min81b]. Mixtures [XPP87]. Mnich [Boh79]. Möbius [BHS06, Hsu95, JK97]. Mod [Ant91, CJS94, FRU15, Jac92, Kui69, Nie72, Ryd96, Sla00, Ehr94]. Model [DS96, Kla76, Bic72a, IT91, Rob78b, kS74, Tra88, Var89]. Models [Kos15c, Mit03, Win68, Kos14a]. Modern [Hal22, Hor04, Low71b, Sig02]. Modes [Boa79, GPS13, Phi14]. Modification [AN79, Rie76]. Modified [AS00, AS05, Bru84, Cop18, Cre89, Fil97b, Hin77, Kon10, Sha93, She68, Swa00a, Trz94, Trz96, You02]. Modular [CW03, Lin75, Pas10, SE20a]. Modules [Man68]. Moduli [AC13, Bur71, Cav77b, Des78a, Man77, SK17]. Modulo [Adl89, Alf64j, And74a, BCOR95, BCHR91, BHM92, BD70, BD72, Bru70, BB11, CJ96, Cat74a, Cha86a, Coh67, DW91, DeL77, DK18, Ehr89, Fra98, Gic22, GKRS01, Har78, Her78, Her82, Hir18, Hol00, How93, JC22, Jav22, Kar96, Kar04b, KW95, KS10c, Kwo89b, Kwo89a, Li00a, Liu05, Lon81c, LMR96, Mam61, Meš13, Mon76b, Mor83, NS99, Pla97a, Rob63, Sed14, Sha11, Sha68, SC05, Som75, Som81, Som91, SK12, SK15a, SK16, SK17, SK21, The71, Tur74, Vin78, Vin63, Wad78, Wad92a, Wal60, Wel94, Wil98c, Wil98d, YK77, CM07, Hol06, ÖAD03a, Som07]. Modulus [Sun75b]. Moessner [Lon86b]. Moivre [Eri98, Lin88]. Moivre-Type [Lin88]. Moment [BL93, BD11, BRS04a, ER00, WR01, AV03]. Moments [fC03, Gri11b]. Monic [Tee94]. Monica [Smi96]. Monier [Arn97]. Monkey [MZ93, PW95]. Monte [NS95, LN07, NS95]. Monzingo [McN92]. Morals [Alf65b]. Mordell [KS11]. Morgan [AJ94a, Ano63e, Dor99, Hor96d, Hor97a, Hor98, Hor99, Hor00a, Hor00c, HF01, Hor02b, Hor03a, Lee02a, Swa66b, Swa68, Swa00a, Swa00b]. Moriarty [Gou72a, Gou74a]. Morphic [Dek20b]. morphism [Sal12]. morphisms [Sal12, See91, Séé98, dL97]. Morrill [Bic73a]. Morse [Nyb23]. Mosaic [Moo70]. Mosaics [Moo70]. Most [Hir18, Mar12b]. Motivated [Lau10]. Motivation [HA64]. Motzkin [BGW14, Ram16, Woa02]. moyenne [CHF85]. MR [Ano88c, Bri83a, Bri83b, GP93]. Multi [CJCK17, Cro75, Moh91]. Multi-Multigrades [Cro75]. Multi-Sets [Moh91]. Multialliance [CDH^{+}21]. Multidimensional [ABJ11, Moo93a, PG13]. multifractality [Jag21]. Multigrade [Cro76]. Multigrades [Cro75]. Multilayer [But90]. Multilevel [LN84]. Multinomial [Bol86, Car76e, Com79, Hil76a, Hil76b, Hil77a, HA71, How93, Koh66, Phi83, Sha79c, Tau65, Vol94]. Multiparameter [CP84, Cer77, ED94]. Multipartite [Car79b]. Multiperfect [AHS87, Hag84, Hag87a]. Multiplayer [CDH^{+}21]. Multiple [BYD^{+}19, Bro89, Bru86a, Cam22b, Car74b, Car80f, CZ10, Edg86, aHZ02, Ivi69, JH73, MW73, Tov85, ZW02]. Multiple-Index [Cam22b]. Multiples [HZ17, Lar15a, Lar16, LO18, Ske65, You92a, You92b, Jam18]. Multiplication [BH80a, Cat74c, DD95, Puc02b, Sha78b, TJ88, Knu88, Pet80]. Multiplicative [BS87, HK95, Hau88, LG91, LP03b, Mac00b, She11, MS04]. Multiplicity [Zar97, Zar98]. Multipliers [CS03, Som07]. multiprocessor [LJ97]. Multisecting [Bru75b]. Multisection [Gla70, HBJ80, Len03a, HA73]. Multiset [Kan14, Len06b]. multithreading [LSS12]. Multivariate [Abd96, PA91]. Multos [Sha87]. Muses [Alf64b]. Music [Car98, Lar78, Low71b, Low71a, Nor64]. Musical [JR98]. Mutual [Gle81]. Mutually [Ark73c, CR10, Kah80, SR12]. My [Ber11, Bru77d]. Mystery [Hol65].

N [Gra89c, Puc62, Sch94, EG76, Rob87]. N. [Puc62]. Nagell [EM15]. Naive [Hir13c]. Naming [Hig87]. Napier [Wil83]. Narayana [KÖS22k]. Narcissistic [Mad72]. National [Arr69, Pic79]. NATO [Bro96]. Natural [AH95, BWWW16, Bro17, Day69b, Dek21, DvL23, Dez77, Gou77b, KK11, PP14, Ske65, Sto77, Zöl93, Zec72b]. Naturally [Gil05, Gil07]. Naturals [Ste76]. Nature [Bas63e, BDD^{+}17]. naturels [Zec72b]. Nazionale [Arr69]. Near [BL65, Zha01a, Spi00a, Spi00b]. Near-golden [BL65]. Nearest [McQ76]. Nearly [Bra00, HH79]. NEC [Pet94]. Necessary [Lay75]. Necmettin [Sub07]. need [Cri07]. Needed [Raw89, She69, She73]. Negative [Bun92, Gre68, GH18, Hor92b, Hor98, Lee94, Mon74, PG89, PA91, Puc02b, GGL24]. Negatively [Hor94d]. Neighbor [McQ76]. nel [Ago49]. Nerve [Fis76]. Nest [Adl78]. Nested [DG84, GCMP14b, Osl07]. Nets [Mab19]. Network [Ark65b, FFD92, FT84, FT87, LHWS96, Swa99c, Wlo71b]. Networks [GKD94, HCD97, Mye75, Ris82, Sta00, LJ97]. Nevada [NS95]. Neville [LMV09]. Newton [Fil86a, GM81, Kom01a, Rie99b]. NFSRs [ZQZ20]. Nielsen [How90]. nilpotency [KA03]. nilpotent [KA03, KY03, ÖAD03b, Özk03]. Nim [Fla78, Hal83, HRR03, Len03b, PH65, Sil77b, Whi63]. Nim-Type [Len03b]. Nine [Wal88, MSPF95]. Ninth [Ano72g, Ano71b, Ano99a, Ano99b, Ano00a, Ano00b, Ano00c, Phi01]. Niven [Cai96a, CK88, CK93, DE18, DD03, Gru94, Gru07, KC89, Wil96, Wil97, Wit16]. No [GG15, Bri83a, Bri83b, GP93]. Nodes [BS96, DCB93, JHK97]. Nombres [Zec72b, Spi00a, Spi00b]. Non [BPGC97, BCD^{+}20, CFH^{+}16, CFH^{+}17, CFG19, Fer73, Gou65a, Gre68, HG81, Hir05, Kos19a, LF14b, LF19, Lay66, LW87, LAL16, MH75a, Nyb03, Oll22, RM11, Sha93, Sha75b, Sta03, Woo75, Wun63, Zel97, AAB^{+}16, BM07b, DH17, Kla73a, LO14, Pen20, THKH23]. non- [AAB^{+}16]. non-absolute [DH17]. Non-Basic [Woo75]. Non-Benzenoid [BPGC97]. Non-classical [Kos19a]. Non-Congruent [Zel97]. Non-constant [BCD^{+}20]. non-equidistant [BM07b]. Non-Existence [MH75a, Wun63].

Non-Fibonacci [Gou65a, HG81, THKH23]. non-Homogeneous [Kla73a]. Non-Hypotenuse [Sha75b]. Non-Integer [Nyb03]. Non-Integrality [LAL16]. Non-Intersecting [Oll22]. Non-Linear [LF14b, LF19, RM11, Lay66]. Non-Negative [Gre68]. Non-Periodic [Sha93]. Non-polynomial [CFG19]. Non-Positive [CFH^{+}16, CFH^{+}17]. Non-Pythagorean [Zel97]. Non-Standard [Fer73]. Non-Terminating [Sha93]. Non-trivial [Hir05]. Non-Unitary [LW87]. Non-Weighted [Sta03]. Noncentral [ED94]. Noncommuting [Hig87]. Nonconsecutive [Nyb01b]. Nondecreasing [Pro21]. None [Pla94, Pla95, Zho97]. Nonexhaustive [Abr00]. Nonexistence [Coh82, Di 93, EP07]. Nonhomogeneous [AP92, Asv89, EMR02, TMR02b]. Nonlinear [Alp19b, Alp20, EH95, Lin65, mJ10, KSSS16, Kro68, KA19]. nonnegative [KKLL21]. nonperiodic [NS94]. Nonresidues [Mon85]. Nonsquare [MW87]. Nonunitary [Hag90]. Nonzero [Kim78, Sub95]. Nørgård [Nor76, Sha05]. Nörlund [AD10, Liu07, TZ96]. Norm [HKN10, HKN11]. Normal [Boa79, DES17, HP75a, SS91, ESW07]. Normic [Stă14]. norms [İpe11, SC10a, Sol05, Sol07, UNS07, ZJ14]. Notated [Lar78]. Note [Aga87, AN66, Alf65c, AH16, And81, AJ91c, AJ94b, AJ95b, Ark66, BH80a, BR85, Bas64, Ber74, Bia95, Bic69, Bol86, Bos85, Bre98, Bri98, Bro72d, Bro72e, Cak91, Cak98, Cap88, Car64a, Car72e, Car75b, CS75b, Cas91b, Cha84a, Chi86, Chr84, Chu23, CS89, DF89, DL16, Dun66, Dun69a, EF80, Fil86b, Fil91, Fil98b, Fre97, Gio94, Gir83, Gou67b, Gou68, HK95, Hal65c, Har70b, Hau88, Hau93b, Hau97a, Hau02, HP94, Hog77c, Hor88c, Hor02d, Hsu73, Iye69b, Izo95, Kah09, KS10b, Lar15b, LL88, Lee94, LL95, Lon86b, May82, McL79, McN87, Met14, MK87, MM75, Mon76a, Mor75, Nod90, Nyb98, O’D79a, OS20, Ove67, Phi83, Phi14, Pih91, PHJ67, Pop85, Pra82, Pro04, Put76, Rea80, Rei79, Ric13, Rob95, Röd94, Ros85, Sel84, SO21]. Note [SGBS86, Sli98, Spi70, Tót87, Tri10, Tur89b, Uma72, Vau76, VPB99, Wil88, Woo79, Yua02, Zei74, ZW02, ZW04b, CR75, CT18, EM01, QM90, Tre96b, Tre96a, Wil82, Yay11]. Notebook [HS13, Coo14b]. Notes [Alf64j, And17, Buc64a, Buc64c, Buc64b, Car74c, Car74b, Car74a, Car75a, Car77b, Car77a, Car80a, Fed64, Gan59, Geb67, Hor63b, Hor83, HsSW95, Jam90, KB64, Mam61, Mea56, Mil71, Ros88, Rus82, SH73, Sha60, Sin73a, Tho63d, Tue23b, Wei16, Wun63, Zei63, Zha99b, FPC08]. Notion [Ant91]. nouvelle [Spi00a, Spi00b]. Novel [Lan11, Win68]. NSW [SW02]. Nucleotide [YS08]. Nullspace [KG02]. Nullspace-Primes [KG02]. Number [Adl71, Ala98, AM75a, Alu96, Alu97, And90, And20, AH95, AH97, Ano95b, AH75b, BL13b, BBP^{+}85, BLM20, BDE^{+}14b, BJF99, BJF01, Bok84, BL13d, Bro72d, Bro72e, Bro74b, BB08, But90, BS96, Byr65, CST05, Can03a, Car78d, Car80b, CM01, Chi86, CLRS86, Coh72, Col87a, Col83, CB14, Cos02, Cre87, Cre88, CP09, D’S89, Dav77, DD03, DeL76, Dez77, Dod84, DCB93, Ege83, Ehr90, EH95, Eng90, EV04, Fis81, FF89, Fuc84, Gla78, Goo75, Gou67a, Gru09, Gui77b, Hal65a, Har70b, Hen94, Hic78a, Hin78, Hir12b, Hir13d, Hir14a, Hoc74, Hod92, Hor61b, Hsu87, HW81, Hun66b, Jar68a, KK11, Kim91d, Kis88, Kla72, Kla73b, Kli81, Kno81, KKMW11, Kom03, Kom10, KL91a, KS02b, Kre80, Lan92, Lar15a, Lar16, LO18, Lee10, Lu70, LY90]. Number [Lud88b, MM65, Mag81, MC76, MZ93, Mar09, MCPR95, MRPC95a, MRPC95b, Maš09, McD91, McI82, Mig77, Moh78, OR14, Pag74, Pau02, Pet94, PH75, Pet91, Pro81, Puc01, Ral85, Rob02c, Rob02d, Ros72, ST90, SH73, Sha64a, She69, She73, SH80a, Sma77, Ste01, Sto77, Sun73, SC84, Tao17, TB91, Tri00b, VKSAN93, Vol94, Wal87b, Web82, Wil00, Wlo68c, Wlo68d, Wlo71c, Wlo72, You14, Zöl93, Alt88, AM15, AI92, AC87, BFS19, Bro69a, BM07a, BP07, Bur92, Can03b, CKMR11, CMP94, Edd90, ESS21, ESW07, FS99, FS14, Krá72a, Krá72b, L’E94, LLP^{+}18, LSS12, LX21, Mak94, Man05a, MT85, MNZ90, MZT90, MZ91, Mar92, Mas94a, Mas94b, MS00a, MS00b, Øre67, PG13, PW95, PT89, Pry94, Rob06b, Spi00a, Spi00b, TLC93, Vis00, Wag06]. number [Wag07, Wol92]. Number-Theoretic [ST90]. Numbers [Abe18, AFRT23, Ade14a, Ade17, Aga90, AE79, AHS87, ASV71, Ala89, Alf63c, Alf63g, Alf64d, Alf64i, Alf64g, All75b, AH77, AE22, AE23, Alv69, And81, And82, And84, AH95, AJ91a, AJ91f, And74b, Ano89d, Ano89a, Ano89b, Ano89c, Ano90b, Ano90d, Ano90c, Ano93k, Ano94c, Ano97a, Ano99a, Ano99b, Ano01a, Ano01b, Ano02a, Ano02e, Ano03b, Ano03a, Ano09a, Ant85a, Ant85b, ABBB13, AGJ^{+}09, Ark65a, AH70, Ark73c, Arn95, Art04, AKO^{+}03, AG87, AP21, Aus23, AC13, dAA79, Bac81, BFF11, Bai83, BW79, BBS81, Bau71, Bea22b, BN85, Bec90, BP99, BH14a, Ber66, BD11, Ber68, BH69a, Ber76, Ber75, Ber77, Bev71, BD77, BK90, BJ79a, BJ79b, BJ85, BJ90, BJS96, BJF99, BJF01, BJ18, Bic69, Bic70]. Numbers [Bic71b, BH71, Bic73b, Blo64, Blu72, Boa01, BPGC97, Bol77, BDD^{+}17, Boy09, Boy92, Bra96, BWWW16, Bro64a, Bro72c, Bro72f, Bro77, Bro65a, Bro68g, Bro69i, Bro71b, BH78c, Bru72, Buc67, Bud80, Bun92, BB11, BK96, Bus63, Bus65, Byr63, Byr65, Byr75a, Byr75b, CP84, CDS03, CN04, Cai96a, Cai96b, Cak91, Cak98, Cal72, Cam22b, Cam23, CZZ11, CJS98, Car64a, Car64b, Car66b, Car66c, Car68a, CH69, Car72b, Car72a, Car72f, CSV73, Car74c, Car74a, CS75a, CS77, CH78, Car78g, Car80e, Car80c, Car80d, Car78i, Cha84a, Cha81, Cha82, Cha84b, Cha91, Cha01, CM14, CZ81, Che04, CC14, CG67, Chu74, Cla86, Coh78, Coh82, CW85, Coh64b, CDW13, CDM00, CG96, Coo10, Coo16, CK16, CK88, CK93, CM17, Cur68, DP16, Daf20, DE18, Dav72]. Numbers [Dav14, DD67, DH68, Day69b, DD70, DLR16, Dda20, DD03, DeL70, DeL77, DeT81, Dei66b, Dei66a, Dei72, Dek21, DvL23, DL11, Des94, Des86, Dez77, Dil00, DD95, Djo04, Dod84, DX22, DD65, DP78a, DP78b, DL16, Drm93, Dub89, DP95, Duj96, Dun67b, DT20, Edw09, EG16, EG17, EA20b, EA20a, EES82, Ehr88, ED94, EF98, Ell12, EM15, Els96, Els98, Els03, EA76, Epa85, Er83a, Er83b, Er84a, Er84b, EG72, Esw78a, Esw79a, EV04, ES14e, ES16e, ES17, Ewe86b, Ewe88, Ewe92, Ewe01, FM15, Fee67, FZ03, Fer66, Fer73, Fer76, Fer65, Fer68, Fer69, Fer01, FFD91, FFD92, Fer97, Fie67a, Fie67b, Fie68a, Fie73a, Fie74, Fil86b, Fil93, FM95, Fil96, Fil97d, Fil98a, Fil98c, FE99, Fin75, FW78, Flo67, Fra94, Fra70]. Numbers [Fra77, Fra67, Fre68, Fre73, Fre87, Fre88, Fre90, FF99b, FS01, Fri66, GCMP14b, Gar04, GL16, GR09, Gau98, Gau02, Gau10, Gau12, Geb67, Gel81, Ger78, Ges81, GL01, Gic22, GM81, Gio94, Gla78, GZ15, GR83, Goo71, Goo74a, Goo86a, Goo86b, Goo93b, Goo93c, Goo86c, Gou63b, Gou63c, Gou65a, Gou77b, Gou78a, GZ19, Gra64, Gra89a, Gre64, Gre77a, Gre78b, GB15, Gru94, Gru01, GT01, GT03, GH18, Gup78a, GC90, GM96, Haa93, Hag84, Hag87a, Hag87b, Hag90, HK95, Hal66, Hal67, Han70, Har78, Har94, Har81, HS98, HS64, Har65, Hau93a, Hau94b, Hay92, HL09, Hen82, HML01, Her78, Her18, Her03, Hew77, HSG63, HMM71, Hil74e, Hil74g, HP94, HPV95, HPS97, Hin80b, HS19b, Hir86, Hir81, Hir12a, Hir13b, Hir15a, HT93, HH84, Hog64e, Hog67a]. Numbers [HL67b, Hog68a, HL69a, Hog70b, Hog70c, Hog72a, HA73, HCB73, HB74c, HB74d, HB76c, HB76d, HBJ77a, HBJ78c, HBJ82a, HBJ82c, Hoi22, HR96, Hon15, HS84, Hor63b, Hor65, Hor74a, Hor74b, Hor84b, HP86, Hor94b, Hor94d, HF95, Hor96b, Hor99, Hor00b, Hor00c, Hos73, How80, How84, How86, HH89, How90, How94, How03, HC11, Hsu93b, Hun64a, Ima98, Ism09, Iye69a, Iye69d, Izo95, Izo02, Jac92, JP19, Jai69, Jam90, Jam02, Jan02, Jar63a, Jar63b, Jar64, Jar67b, JC22, Jen93, Jen94, Joh09, Joh12b, JS16, Joh20, Joh22b, Jon75, Jon76b, Jor65, Kal82, Kap45, KY09, KS04, KK95, Kim81b, Kim95a, Kim98b, Kim01a, Kim01b, Kim02, Kim04, KKLS18, Kir22, KPT83, Kla68b, Kle91, KK89, Kno90, KM95, Knu64, Kom01b, Kom12, KMP14, KY20]. Numbers [KL91b, KL93, KG17c, KG19, Kou90, Kou94, KL87, KB64, Kra70, KV92, Kui69, KsS72, Kun90, Kyr85, Laf71, LL14, LG91, Lan00, Lan02, LL13, LOF14, Law83, LL87a, LL88, LL95, LLKS01, LL83, Len94b, Len95, Li21, LJ82, Lin88, LSW14, Lip04, Liu01, Liu02, Liu05, LL05, Liu09, Log81, LF69, Low71b, Luc99, Luc00, LP03b, Luc03d, LS13, LY13, LS14a, LM19b, LOT23, Luc73, Lyn70, Mab19, Mac10, Mad72, MS20, Mak88, Man64, Man67, MOP71, Mar12b, Mar12d, Mar12c, Mar13a, aMS22a, McC81, McC72, McC85, MH75a, McD82, McD87a, McD87b, McD96, McD98a, McD98b, McD01, McD02, McI20, Mea56, MS95d, Mel99a, Mel99f, Mel00a, Mel00d, Mel03c, Mel04a, Mel09, Mel10b, Mel10a, Mel15, Mel16c, Mel17b, Mel17d, Mel18e, Met14, Mid79, Mig77, Mil60, Mil71]. Numbers [MW92, Min89, Min81a, Moh68, MM90, MV91, MV15, MW87, Mol88, MC04, Mom01, Mon74, Moo94, Moo93b, Mun05, Mur82a, Mur82b, Mye75, ND87, Nef67, Nie72, NKW23, Nyb99b, Nyb01a, Nyb01b, Nyb03, NS99, NR15, O’D79b, ON09, Oll22, Ond70, OP14, OP18, Owi79, PR13, PP14, PD15, PP17, PP19, PP00, PTW13, PT17, Par77c, Par77b, PW23, PP20, PPK21, Pau23, PRW93, Pet91, Phi84, Phi04, Pil12, Pin02, PS80, Pon17a, Pon17b, Pon19, Pop84, PL07, PR01, PR02, Pri05, Pri18, Pri68, PT82, Pro96, Pro16b, PS19, Pro22, Pru67, Puc01, Puc02b, Pul94, Rab99b, Rao53, Rao02, Rao03, Rao05b, RP18, RP19, Rec75, Red94, Rei92, Rei93b, Rie99a, RHT18, Rob78a, Rob81, Rob83a, Rob83b, Rob83c, Rob84, Rob90, Rob91a, Rob94, Rob98]. Numbers [Rob14, Rok96, Ros79, Ros83, Rus82, RS11b, Rya75, Ryd96, Saf92, SP18, SP19, Sai17, SP02, SI03, Sch13b, Sch19, Sch20, SW02, Sha93, Sha75b, SH71b, Sha74c, Sha74d, SHC74, Sha77, Sha78a, Sha79b, Sha84b, Sha60, Sha76, Shi20, SC05, SD18, Shu20, Sil11, Sim98, Sin80, Sin75, Sir97, Sit70, Šiu11, Ske65, Sla77, SS91, Smi96, SK81, SK13c, Sri20, Sta76, Sta80a, Sta80b, Ste71, Ste11a, Ste86b, Ste21, Sto65, Sto75, Str99a, Str99b, Sub99, SH70, Swa97b, Swe64, Tan00, Tau68a, Tau68b, Tau73b, Tau76, Tau89, TL08, Ted19a, Ter96b, Tho63d, Tho63f, Tos78, TS92, Tri74c, Tur74, Tur86, UC69, Van86, Van96, Vau76, Vor61, Vor83, Wal76a, Wal79, Wal83, Wal84a, Wal85c, Wal88, WH74b, Was81, Web75, Wei66a, Wel67]. Numbers [WY00, Wiȩ80, Wil72, Wil75a, Wil96, Wil97, WTvR87, Win65a, Win65b, Wlo63, Wlo65, Wlo71a, Wlo73, WSH74, Yab02, Yal72, YS08, YWH96, Yu98, Zak83, Zar70, Zei63, Zei64, Zei67, Zel93, Zel97, Zha97a, Zha98a, Zha02, Zha04, Zha99b, ZW01a, ZW01b, Zha01b, ZW03, ZD04, ZW04c, Zho99, dBF85, tR84, AD10, ACOM21, BHS11, Bea50, Bea22a, Ben06, BHP92, BPH93, Blo53b, Bra06, BL13e, Bri83a, Bri83b, BM07b, CCZ15, Cha07, CM15, Chr12, CHT06, CH89, CSS07, DL20, DOU08a, DOU08b, Dem08, Dev11, Dij81, DM08, Dun97, DRS11b, DRS11a, EA19, EBK89, EMS10, EB17, Esw78b, Esw79b, EM01, EP07, FP09, Fan09, FLP08, FZ08, GCMP14a, Gau11, Gin54, Gol06, Goo93a, Goo94a, GQ07, GL80, Gri09, Gru07, GTU13, Hen06, Hog69, IL07, İpe11]. numbers [IA14, JH05, KLM^{+}19, Kah05, Kan14, Kar04a, Kar05, KŞ13, Kel57, Kel58, KK15, Knu90, Kos01, Kos09, LaG13, Lee00, Li12, Liu07, LS07, LH09, LS14b, LP18, MZ07, Mag09, Man06, Man05b, MP85, Mun07, Nyb05, ÖTA05, Ogu20, Osl07, Pan07, Pet80, Pic81, PRS96, PT89, RG14, RMA96, Ran95, Rao55, Rao05a, Rei54, Rol03, RSHL07, RL15, RL18, SE13, SE20a, ST06, Sha06, SC10a, SCH11, Sho78, Sin85, Sol05, Sol07, Som06b, Sta17, Ste23, Sub19, Sun06c, SB93, Tak00, TK04, Ten09, UNS07, Vaj89, Vaj08, WS80, Wlo13, WP07, Yab07, Yan95, lY08, YK03, Yor14, You06, Zec72b, Zei96, ZW06, ZL06, ZJ14, BT12, Gal08, Sla00, BJ99, Cad16, Gra89c, Puc62, Bro69a, Har09, Sch94]. Numeration [BS97, Zec72a]. Numerator [BH76, HBJ77d]. numeri [Pic81]. Numerical [BY16, CM10, FFD91, Hor00a, Pop86, Ste78a, Trz94, Rol03, Sun05, Trz96]. NY [ACM12].

Obey [Wlo71a, GQ07, SB93]. Obituary [Ano63e]. Object [AGO95, Mat90]. Object-oriented [Mat90]. Objects [CD87]. Observation [Fil97c, Mel00b, ST73]. Observations [Col88b, Eme66, HKN10, HKN11]. Obtained [Wil72]. Obtaining [Du00, Gup76, Lin98]. Occupancy [Abr71, DMP07]. Occupational [HM83]. Occurrence [HS99, SS01]. Occurrences [Far86, Tov85, Sal12]. Occurring [Gil05, Gil07]. October [IEE84, IEE08]. Odd [AM75a, AJ92, Bru94b, Chi86, Coh78, CW85, DL16, Duj02, Ell12, Gra89a, GM96, Hag90, Hic78a, LSW14, Liu05, MH75a, Mel09, OR14, PP00, Pih01, Pih10, RT04, Rob02c, Som82, Ste81a, Ste81b, SH70, Wal87b, Wal88, EP07, Mil11]. Odd-Order [Chi86]. Odd-Subscripted [PP00]. Oddness [Gri75]. Odds [Tad65]. Ode [Jon76a]. oeuvre [MS15b]. off [Bre74a]. Offers [Rac19]. Offset [RW81]. Oh [Bru77d, Wei88]. Old [Bri64a, Duj99, Mic64, Pel72, TJ92]. Olde [Bro72h, Bro73c, Mel04b]. Older [Gou72d]. Ollerton [Hen22b]. Olympiad [Ben02]. Omega [Tur89a]. Omission [Ano65e]. Omissions [Ano64c]. Omitted [Lor73]. Once [Car80a, Hun72]. One [BJ90, Bre74a, BD70, BD72, Bur90, Car74a, Car77a, CS77, Car80a, DeL84, Fin75, Fla82, Gel81, Geo75, HBJ77a, HRR03, Ism09, Kim83, Mai70, Mar12b, McQ76, Par77c, Par77b, Pon17a, RW81, Sto77, Wlo68a, Fos66]. One-Dimensional [McQ76]. One-Free [Kim83]. One-One [Mai70]. One-Pile [Fla82, HRR03]. Ones [AH75a, AG78, Du00, KG19, KV92, St.84, Wul76]. Only [SK12, Kel08, Puc62]. Op [Nay65]. open [dF11]. Operation [ABN^{+}19, Tri65]. Operational [AS84, Gou63c, Gou78b, Hsu95, Nas76]. Operations [Naj89, PT89, Sho78]. Operator [Hal68]. Operators [CS75a, Mal83, SSS24, WH70, DH17, Ogu20, TD16, TD19]. optical [SMSM01]. Optimal [AK94, McC72, Rok96, vR89, Kir94, PG13]. Optimality [AW66, Hor83]. Optimization [DOU08a, FT84, FT87, Tan98, Web95a, Kah06, SYY04, Sub07]. Optimization-Based [Tan98]. Optimum [KR89]. Order [Ade18, Alf65c, Alo76, AH81, AP92, AJ91d, AJ97, AB85, Ark73b, AS76, AS81a, Bab76, BL93, Bas20, Boa79, Bri64b, Bro69e, Bro69d, Bro74a, BsS95, CK24, CS03, CSH72c, CSH72d, Car74a, Car78a, Cat74b, Cha86a, Cha91, Chi86, CMX20, CBH17, DP16, Dil87, DJ68, EMR02, Er83a, Er83b, EV04, For67b, Ful81c, Geo89, GPS13, GL01, Hau93b, HH84, Hor86a, Hor92a, Hor92b, Hor93, Hor94c, HS85, Hor88c, Hor92c, HSL01, Iak77b, Iak81a, Iak81b, Izo02, KY09, Ker82, Kim91a, Kim91b, Kim10c, Kir77, Kla66a, KMP14, Kos15a, Kos20, KG20, Kos21c, Kos21d, Kos21e, Kos21f, KW13, LR85, LK09, LM85, Lee87, Lee97, Lee01b, Len94a, Len95, Lev85, Li99b, Li00a, Liu01, LB91, MH90, Mar12a, Mar12b, Mar12d, Mar12c, Mar13a, Mar13b, MO09, MO10, MS94]. Order [Mel02, Mel17a, Moo93a, Nor05, PP94, PM82, Phi83, PGP85, PG89, PA91, Phi14, Pla96a, Pop77, Pop86, Rab96a, Rab99a, Rob82a, Rob04a, Rus81, SH72, Sha74e, Sha83, Shi20, Sil77a, Som79, Som84, Som87, Som89, Som91, SC04, SK12, SK13a, SK15a, SK16, SK17, SK21, SJ84, Spi97, Sta03, Sza17, TMR02b, TT16, Tri74a, Tur79, TS89b, XPP87, Yal73, Zei71a, Zha97b, ZL98, ZL95, Zho96, Zho03, ZH03, Djo05a, Gou06, Hir05, Hir06a, Hir06b, HL66, HST94, IA11, JH05, JV19, KŞ13, Kil07, KA13, LP14, LS18b, NO15, Pet80, RMA96, Som06b, Som07, TK04, WS80, lY08]. Order- [Er83a, Er83b, KŞ13, Kil07, Pet80, TK04, WS80]. Order-Theoretic [HH84]. orderability [MT17]. Ordered [Cad71, Car72d, CR10, Lás82, Pro83]. Ordering [Alf63h, Cre70, HBJ82b, HO87]. Orderings [Goo75, Kim04]. Orders [Bec90, How94, Kos18a, Kos21a, Kos21b, Zha01a]. Ordinary [MH87c]. Øre [Bro69a]. Oregon [Gon89]. Oresme [GZ19, Hor74b]. Organization [CLRS86]. organizer [Sug93]. oriented [Mat90]. origin [Rob78b]. Originating [BKM17]. Orleans [B^{+}96]. Orono [Bur92]. Orthogonal [ABD05, AN66, Ahu69, AM80, Ark73c, Ark74, AS81b, Car66c, DMY13, Gro97, Ker82, Tau89, AK05, Pet12]. orthogonality [HHJM15]. Oscillator [Boa79]. oscillators [MB19]. Ostrowski [TMR02a]. Other [And90, BCH12, BJ89, Bro73b, Cha84b, Hir86, Hor69, JC17, KM14, MH86, WH84, Fol04, HD19, Men89]. Otherwise [Nor20]. Our [Ano98d]. Outerplanar [Ala98, ESW13]. Outerplane [BBS81]. Overlays [Boi69]. Øystein [Bro69a].

P [Ano88c, Pic83a, Puc62, Bom20, Sch94, Gou68]. P-3 [Gou68]. p-Proportions [Bom20]. P. [Kar70, Puc62]. P.Q.M. [Pag74]. P.Q.M.-Cycles [Pag74]. Pacioli [BF90, FR85]. Packing [Wei79, CKMR11]. Padovan [GWZ07, Ted19a]. Page [Ano63a, Ano72f, Ano14a, Ano17k]. Paint [BJ78a]. Pair [BH80b, Fil15, Swa99c, Wan05]. Pairs [AFT22, AKO^{+}03, BS87, Bal90, BN77b, BN93, CD13, CP79, Gri11d, Hir12b, HH78, Hor78b, Hsu87, Kyr93, Moh81, Naj89, Sho99b, Sil76, Tur89a, Wal76a, Di 13, HS18]. Pairwise [LS13, Tót02, Sho06]. Palatine [Pic79]. Palatino [Arr69, Arr69]. Palindromes [DS04, Dro95, Que93, Sha12]. Palindromic [AJS22, Coh90, HS98, HB75b, Kim98b, Kom01b, San23, Tri74c]. Palm [Dav71]. Paper [All77, Bre98, CSH72a, Gou68, HCG03, Hor80, Kui69, KsS73, Pro04, Röd94, Sel84, SH82, Sin90, Sla00, Hal22, Jun61]. papers [Jar66, Geb67, KB64, Wun63]. Parachute [DL92]. Parallel [Alu96, Alu97, Ani95, DGST88, Gla70, Mas94b, MCPR95, MRPC95a, MRPC95b, Mat95, NC85, PRS96, Ste73b, CMP94, Edd90, LSS12, Mak94, Mas94a, MS04, Pry94, Wol92]. Parallelism [CV88, Sar90]. Parameter [Ism09, Lin96, Mel16f, Mel16h, Mel18a, Var89]. Parameterizing [MS04]. Parameters [AJ96, Fie68b, Gre68]. Parametric [Col88b, YY19]. Parametrized [DOU08b]. Parent [Sur14]. Parities [GS81]. Parity [DG98, Ege83, Hod92, Kun76, Rob04b, Sch00]. parking [Wir08]. Part [Asv89, Ewe86b, Gra89a, Adl68, Adl69b, AJ95b, BH63a, BH63b, Bic70, Bic71b, BH71, Bic73b, Bre64a, Bro72f, BH78c, BH63d, EA20b, Fau64, Fau65, GG79c, Goo68a, Goo68b, HB63, HB64c, HL67b, HA73, HCB73, HB74c, HB76c, Jes63b, Jes63a, Jes64, KG17b, KG18a, Mel00d, Mel01, Mel17d, RH63, She72]. Partial [ASD15, Bec90, Car64b, Chu90, Chu21b, CK88, Djo01b, FH20, GQ09, Hor94c, Jen66, Jen71, Kim10b, RP71, Sto70, Str99a, YL97, Coo06]. Participants [Ano17f]. Particles [HM83, McQ76]. Particular [LF14a, LF14b, LF15, Ord93, SK20]. Parties [Lew95]. Partition [AML85, Cho91, Col87b, Ewe80, Fie64, Gir81, Gou77b, Gup78b, HK95, Hil74f, HA76a, Hog77a, Hog77b, HBJ84, Kis14, KP78, Moh68, Pin02, Rob04b, Sha79a, Whi86a, CJ24]. Partitioning [SS91]. Partitions [AM75b, AM75a, Cad71, Car72d, Car76c, CLM22, Col83, Cre89, El 09, Ewe86b, Fie64, Fie66, Fie67a, Fie68a, Fie73a, Fuc84, Gue90, Hic73, Hic78b, Hic78a, Hir13d, Hir14a, Hir18, HB75c, Hog80, Jon91b, Kim91d, Kla68b, KR04, LG91, LP03a, Man75, Mun17, Muw81, Pad89, Par98, Pow96, Pro81, Rob96, Rob00, Rob02c, de 78a, SI03, Sch82, Sil11, Wan93, Wei16, CS21, Pro83, WW06]. Partly [Cad71, Car72d]. Parts [AM75a, Cam23, El 09, HZ17, Hic73, Hic78a, Hir14a, Hir18, Kim81b, Rob02c, IL07]. Pascal [Ano93f, Ano93g, Ano93e, And78a, AS00, AS05, BVB95, BJ89, BJ93, BH73b, BH14b, Boi69, Bol84, Bol86, Bol89, BB90, BN09, Bre64b, Bre64a, Bud80, CH94, DMP07, DW91, DQ83, Edw09, Eri98, GS95, Gou65b, Gou75, Har78, HS66, Hew77, HH72, HB64a, Hog68b, HB69, Hog72a, HA73, HB73b, HB76a, HB76b, HB76e, HA76a, HBJ78b, Hor64, Jon76a, Kar96, Kar04b, KMST97, Kub04, Kun76, Lee94, LY02, Lon81c, Lon81d, Lun77, MMYA10, Naj93, OS98, OS21, PS02, Raa63, Ram16, Rei93a, Sha77, SGBS86, SP77, SH79b, SH79c, SH80b, Sof12, STL22, Tur88b, Wil72, Wlo68d, Wlo71e, WM75, YL04, YG18, YY19]. Pascal- [Bol84, Bol86, Bol89, Tur88b]. Pascal-Like [Edw09, MMYA10]. Pascal-Type [Naj93]. Pass [MT02]. Passing [Ano02c, Tor19]. Path [Kim02, ZS85, Kir94]. Path-Counting [Kim02]. Pathological [Kum74]. Paths [Aga87, But90, Car72f, Chu74, Ege83, FJL^{+}20, FKM^{+}14, Fra70, Gre64, Kim01a, Oll22, PS92, Pro21, Ram16, Sto65, Sto67, Woa02]. Patricia [Ano07a]. Pattern [And78a, HS82, Moo70]. Patterns [Alf64f, BJD95, BDD^{+}17, BYD^{+}19, Bom20, CCC00, EA12, Har98, Sha72b, Tur88a, TR90, CLHY12]. Patton [RF72]. Paul [Ano16b, Ano16c, Ano16f, Ano17j, Ano17g, Ano17h, Ano18c, Ano18d, Ano18e, Ano18f, Ano19e, Ano19f, Ano19g, Ano19h, Ano20c, Ano20d, Ano20e, Ano20f, Gou68, Ano17i]. PC [MNZ90]. Peculiar [Gor12a]. Peirce [Eis51]. Pell [ACOM21, BCM10, Bic75, BHL21, CL11, CY00, Coh72, DLR16, Dda20, DeL76, DeL81, DB01, DV88, EES82, Erc79, ES14e, FL18, GZ19, Hor71, Hor84b, HM85, HM87, Hor88b, HSF94, Hor94b, Hor94d, HF95, Hor00b, Hor03a, Hor05, KLM^{+}19, KST20, Kil07, KA13, KR04, Li12, MH85, MH86, MH87a, MH87b, MH87c, MH87d, MH90, McD96, McD02, McD11, Mel16e, Mel16h, MK87, NS05b, O’D79a, PR01, PR02, Rao05a, Rao05b, Rob84, SE20b, SP19, SI05, Ser74, Tsu21, WH84, Weg81a, dF14]. Pell-Type [O’D79a]. Pellian [Ber74, CSH72f, Sha78c]. Pendulum [Her20]. Pennsylvania [ACM96, IEE08]. Penrose [RS98a]. pentadiagonal [IA14]. Pentagonal [DeT74, Han70, PR01, PR02, TS92]. Pentagrams [Fei72]. Penultimate [Gri75]. Perfect [AH81, AP21, CCZ15, CJS98, Coh78, CW85, Col88a, Col94, DL16, Edg72, Ell12, EM15, Fed64, Fra77, Gra89a, Hag90, Hun66b, Lev68a, LF69, MH75a, OR14, PP94, PD15, Pri68, Ric13, SH70, TS92, Wal87b, Wal88, Yab02, AAB^{+}16, EP07, LP18, Mol18]. Perimeter [DeT92]. Perimeters [Ber89]. Period [BL13b, BCHR91, Cat74a, DeL77, Gre77c, HK91, MC04, Rob76, Sta80a, Vin63, Wil86, BE07, BM07a, KY03]. Periodic [ABCM12, Alf63i, BL13a, BD11, BMRS04, BH69a, Car07, Fer78, HB77b, Lau77, Leh75, Lew91, MO09, Mor77, Sha93, TS04, Wlo71d, YK77, BTD22, BB19, CCL89, CT18]. Periodicities [CK10]. Periodicity [BL17a, Blo65, DeC73, ES15f, Hei64, Her73, Jar63b, KG19, KH72, McG09, McG10, Poo88, She68, Sta76, Sta80b, Yal73]. Periods [Alf64j, AL79a, BCOR95, Bou20, CDM^{+}06, Des78a, Ehr89, Ehr90, Hen15b, Her78, Her82, KS10c, Kwo89b, Kwo89a, MZ92, Pla97a, Rei93b, RT04, SH76, Som77, Som82, Bre94, Jam18, SE20a]. Permanent [KP69]. Permitted [Goo75]. Permutation [Liu96, Mul80b, Mul80a]. Permutations [Car78c, Car80b, Car81a, KS02b, Kre80, Sta85, Yan91, FPC08]. Permutative [Ske65]. Perrin [Li12]. Perron [BH69a]. Perspective [AAS85, Gin02]. Pertaining [Rob91b]. PETASYS [AI92]. Petersen [Wag06]. Phased [BQS00]. Phi [Ber66, Dek20a, Dek21, Hin77, Hol65, Hun74b, Pow96, Sha75a]. Phi-Partitions [Pow96]. Philadelphia [ACM96, IEE08]. Philippou [Gra89c, Sch94]. Photo [KTW84]. Phyllotaxis [Goo91, Kar65, dSM63, Swi04, Rob78b]. Phyllotaxy [Car78i]. Physical [Høy09, Tre16a]. Physics [Dav72]. Physiology [Fis76]. Pi [Bru77d, Dra64, Ton24]. Pi-Oh-MY [Bru77d]. Picking [Tri67f]. Piecewise [Fra94]. Pierce [KM95, PVB98, Sha84a, Sha86, Sha94, VPB99]. Piero [AJ94c]. Pile [Fla78, Fla82, HRR03]. Pillai [DL20]. Pine [Bro68c]. Pineapple [Car78i]. Pineapples [Ond70]. Pioneers [Wei88]. Pisa [Bro70a, Anoxx, Baš78b, Baš78a, Beč01, Car39, Cho82, Dev12, FR85, GG69, GGA83, Han07, McC19, Miu02, Pic79, Pic81, Spi00a, Spi00b, Uli11]. Pisani [Lün92]. Pisano [Geg93, Hor04, Arr69, Bab41, BF90, Car39, Fra40, Geg93, Gie13, Gri73b, Mar74, Miu81, Pic81, Pic83b, Pic83a, Sig02, Wes80]. Pise [Spi00a, Spi00b]. Pisot [AH95, DeL70]. Piza [Gou89]. Place [Hun74b]. Places [Ber66]. Plan [HG81]. Plane [Her81, MS03]. Plants [Goo91, NS05a, THKH23]. Plastic [Nak20]. Plates [Bro89, HBJ79a, JH73]. platforms [LSS12]. Plato [Fei72]. Platonic [WB12]. Pleasant [Mon83]. Pleasantry [Ban76b]. pleasure [Lün92]. PLRS [MNPX17]. plurality [Par17]. PMO [SBM97]. pocalypse [Hil21]. Point [CG67, Eng91, MBR^{+}13, Pet96]. Points [AL79a, BN77a, FRU15, Gle81, HK78, JK93, Lin21, Mar12a, Mel16f, Nag01, SK13b, Sta85, TF92, vR89, BM07b]. Poisson [GPS13, Kou90, Phi14]. Poisson-Truncated [Kou90]. Policies [Ano94d]. Pollution [Dei72]. Poly [CJCK17, SP02]. Poly-Bernoulli [SP02]. Poly-Euler [CJCK17]. Polygonal [And81, And82, CB14, EES82, Far86, Hir86, HH84, Luc73, PS92]. Polygons [Alf63d, Bud80, Her81, JC17, Pec65, RF72]. Polyhedra [Fei72]. Polyhedrons [Wlo71e]. Polylogarithm [Coh99]. Polymers [BPGC97]. Polynomial [Alp19a, Asv89, BKM17, Bea03, Bei78, BH76, Bev71, BK90, Bra90, Bru00, BF95, Car65, Fah16, Gau12, Goo68a, HH65, HBJ77d, Hor96a, HF01, Jen93, Kim10b, Kim12, KS16, KKLS18, Kin73, Kos14b, KG17a, KG17b, Kos17, Kos18a, KG18a, Kos18b, Kos19d, Kos20, KG20, Kos21a, Kos21b, Kos21c, Kos21d, Kos21e, Kos21f, Kos21g, Kos21h, Kos21i, Kos22e, Kos23b, Kos23a, Kos23f, Kos23i, Kos23k, Kos23j, Kos23l, Kos23m, Kos23n, Kos24a, Kos24b, KG24, Kos24c, LF14a, LF14b, LF15, Lee01b, Lee02a, Luc21e, MH87d, Moo93a, Pet84, Pet85, SI05, Sto75, Sur14, TMR02b, TR04, Yan88, CFG19, EE10, Wag07]. Polynomials [dOGC96, ABD05, AN66, AM80, AN83, ASV71, AS84, All76, Alp20, AJ94a, AJ94b, AJ95a, AJ95b, Ark65b, BY16, BB69, BRGGGS20, Bed13, BH74, BWH75, BH79, BBHM85, Bev76, BH73a, BN09, Bos85, BDM12, BR82, Bur90, Bus63, Byr63, Car66c, Car75a, CH78, Car78g, Cas91b, CH02, Cha91, CL11, CL12, Cig03, CJCK17, DMY13, DD13, Dil87, Dil91, Djo96, Djo98, Djo00a, Djo00b, Djo01a, Djo01b, Dor99, DV88, Fer78, FH93, FH94, FMH94, Fil97b, FH20, Fox01, FZ01, Gau04, Geo75, GM20, Gla95, Gou63a, Gru14, HSR97, Hen15a, Her82, HL67a, HL68, HB73c, HB73a, HL74, HBJ78a, HBJ78b, HT18, Hor77b, Hor79a, HP81, HH82, Hor85a, HM85, Hor85b, Hor86b, HM87, Hor92a, Hor92b, HSF94, Hor96c, Hor96d, HF97]. Polynomials [Hor97c, Hor98, Hor00a, Hor02a, Hor02c, Hor03b, Jai74a, JK98, JV19, JS16, JH73, KH82, Kim79a, Kim80a, Kim80b, Kim80d, Kim10c, KM05, KG02, KM95, KY20, Kos19a, Kos16, Kos22c, Kos22d, Kos22f, Kos22g, Kos22a, Kos22b, Kos22h, Kos22i, Kos22j, Kos23c, Kos23d, Kos23e, Kos23g, Kos23h, Kos23o, Kos23p, Kos23q, Kou94, Lan92, Lan00, Lan02, Lee01a, LL83, Lev01, Lin21, Liu01, Liu02, Mac00b, MH85, MH86, MH87a, MH87b, MH87c, MH90, MD87, Moh76, Mul80b, Mul80a, Nor05, Pet84, PGP85, PM85, PG89, PA91, Pop85, PS76, RS02, Rui17, Rus12, Sha13, Shi20, SR83, Sub90, Sub95, Sur96, Sur98, Swa66b, Swa68, Swa77, Swa97a, Swa99a, Swa99b, Swa99c, Swa00a, Swa00b, Swa00c, Tau68c, Tau68a, Tau68b, Tee94, TF92, TZ96, Vse11, Wal76b, WH84]. Polynomials [Wan95, WH04, WP69, WSP04, You02, YL97, YZ02, Zei78, ZJ98, Zha02, ZW04a, ZL95, Zho96, AK05, BK16, Boy07, Bri83a, Bri83b, CC16, CH06, Chu10, Djo05a, EBKT85, EB87, Gau11, Ila71, IA11, JLW07, Ma11, MB19, Pet12, RT72, Ram14, Rem11, Ric93, RT18, SR12, Sla16, Sun06a, TF93, TKS11, WW17, YZ17, Zah82, ZZ06]. Polyomials [Hor97b]. Polyominoes [And70, KR22, Kla65]. Polyphase [Lyn70]. Polytope [EG16, EG17, dF11]. Polytopes [Ris05]. Polyvibrating [MOP71]. Popes [Hig87]. Popular [Puc62]. Population [Bro67e, Bro68d, Kla76, PKS18]. Populations [Par69]. portable [Pry94, Wol92]. Portland [Gon89]. Porto [Bru94c]. Posamentier [Gal08, Har09]. position [Han07]. Positions [JO24]. Positive [AC11, ABN^{+}19, BHL^{+}20, Cad71, CFH^{+}16, CFH^{+}17, Col83, HZ17, HW12, HKN10, Hei79, Hen78a, Hig72, Hog80, HBJ84, Hor94b, Luc03d, Mai70, Mel16f, PP20, SS01, Ste71, Tót02, Wan93, EZ04, HKN11, Kim07, WW06]. Positiveness [NS90]. possessing [EBKT85]. Possibility [Pla96b]. Possible [Som82, Wlo71d]. POVMs [GS17]. Power [Ard04, Ark69, Ark73a, AAB^{+}90, BB69, BSS08, Boy09, Bro68f, Bro74b, Cav68, CC14, CL11, Chu96, CM89, DFLT14, DeL84, Fer97, Fil96, Fox21, Goo74b, Gou99, Hau93a, HB64a, HL66, How01a, IR12, IT91, Jar63a, KK01, Lar15b, LF16, LF17, LL87a, McI20, MS95f, Mit03, Mon76b, Nak20, Nyb09, Sch91, Sha11, Sha74d, SK12, Ste78a, TCC^{+}04, WH03, Wil98d, Zei65, Zha09, Duv91, Gau11]. Powerful [McD82, McD87b, MW87, Van86, Wal76a, Yab07]. Powers [AN83, Alf64i, Bas64, Bea66a, BLPS11, Bre64b, Bre64a, Bro94, Bru00, Buc64a, Buc64c, Buc64b, Bun75b, Bun12, Bun14, Car66b, CH69, CBH10, Dam89, DvL23, Des86, DB66, DB67, DB70, Fer01, Gau04, Gla78, Gou63b, Gou78a, HB97, HR91, HBJ77a, Hor74a, How96, How98, HW73, Ila71, Jon91a, Kal90, KC93, Kil76, Kla66a, KK89, Kom14, KL93, Lan90, Lan00, LF15, Len17, LF69, LMR96, LS18b, Mar12c, Mel99a, Mel00a, Mel09, Mel10b, Mel10a, Mel16c, Meš13, MW22, NKW23, OP14, OP18, PPK21, PS02, Pop77, Pro16b, RFLT20, Rob78a, Rob83c, de 78a, Sel83, Sha74b, SH77, Sta03, SS17, Sta80a, Ste78b, Sti98, Sun78, Sun01b, TT16, Tee94, Tep74, Wes66, Wil70, ZG17, ZW03, BH05, BM07b, Can51, CM15, DL20, GB06, Goo93a, LP18]. powers [Mag09, TR03, ZW06]. Pp [Mos08, Gal08, Hal22, Tol17b, Tol17a]. Practica [Høy09, Oak09, FH08, Hug10, Fre92a]. practice [Spi00a, Spi00b]. praticque [Spi00a, Spi00b]. Pratt [Bay08]. Preceding [Gou71a]. Precisely [Rob83a]. Predetermined [Mea56]. Predictable [Dra64, Sha84a]. prefixes [BFS19]. Préhoude [Spi00a, Spi00b]. premier [CHF85]. Prescribed [Cav77b]. Present [Mac10]. presentation [MR84]. Preserve [Bro78c]. Preserving [AS20, Aus23]. Press [Ano88c, Tol17b, Tol17a]. pressure [Rob78b]. Prices [Gil05, Gil07]. Primality [Bol89, Bri98, Gou72c, Gou74c, GG85, Gou89, LM93, Mor75, Rie69, RW21, Sun06c, Wil88, Yan95, Bay08, BB04, BLS75, WJ76, Wil87, Wil98a]. Primary [Som80, Som82]. Prime [Alf64j, All81, AH16, BC10, Bal19b, BJ02, Bre74b, BWWW16, Bru86b, Bru70, Buc67, CP79, DK18, DJ67, Ehr88, El 09, EG76, EV04, Gar04, Gic22, Gio94, Gou64b, Her85, HJ68b, Hol00, Izo05, Jar63a, Jar68a, Kar96, KP77, KS10c, Lay66, LM85, Lee10, Man64, MST13, McN92, Meš13, Mon76b, Mon88, NS99, Pau02, Pla97a, Pon19, Rob83b, Rob84, Sai17, Sed14, Sha11, Sha64a, Som82, SK13a, SK13c, Tao17, Tee94, Tót02, Vol94, Weg75, Wol99, Zho99, BL13e, Hol06, Jar06, LX21, RL15, RL18, Sho06, Bal19a, Gri75]. Prime-Power [Jar63a]. Primed [Tri67c]. Primefree [JS19]. Primer [BH63a, BH63b, Bic70, Bic71b, BH71, Bic73b, Bic75, Bro72f, BH78c, BH63d, GG79a, GG79b, GG79c, HB64c, HL67b, HA73, HCB73, HB74c, HB76c, HBJ78c, RH63]. Primes [Alf64h, Boa01, BG21, CZZ14, CFL^{+}22, CD13, Dro00, FZ01, HBJ77a, Hun78, Jar67a, JC22, Jav22, Kar70, Kar72, Kla09, KG02, Kun90, Leh64, Lu70, MG02, Mad69, Moh78, Nyb12, Nyb23, Par70, Rob94, Rob05b, de 78a, SY84, Som79, Som80, SK15c, Sun02, Tri10, TJ92, Van12, Weg81b, Wei73, Zie63, BE07, CCZ15, DK99, Jar06, MR07, Pen20, Som06a, WK09]. Primitive [Alf63g, Ber82, Ber89, CY91, DeL77, DJ68, Jar63a, Jar68a, KP77, Li99a, Li00b, May82, Mol88, Mon76c, Pho91, San90a, Sha72a, ST73, SH76, Tay76, Ter12, Vou95, Weg75, Weg77, Weg81b, Wul81, Yab01, Rob06b]. Primordial [PH88]. Princeton [Tol17b, Tol17a]. Principle [Kon10, Wlo67]. Priority [Wil98b, HKN11]. Prize [Ano15a, Ano16b, Ano16c, Ano16f, Ano17j, Ano17g, Ano17h, Ano18c, Ano18d, Ano18e, Ano18f, Ano19e, Ano19f, Ano19g, Ano19h, Ano20c, Ano20d, Ano20e, Ano20f, Ano17i, Slo15]. Probabilistic [BNOS03, FM95, Gam95, Hey80, Nak14, SY89]. Probability [All81, Ess03, Gor12b, PGP85, Sha73, Tót02, Tur79, AK05]. Problem [Alf63b, Alf66a, Alm86, And02, Ano63f, Ano63g, Ano98b, Ano00e, Ark73b, Ark73c, AHS79, AHS80, Bau80, Ben02, BH78a, BJ09, Blu72, Boh79, Bri64a, Bro68g, BS93, CJS98, Car78e, CY00, CG67, Chu85, Col87b, CM17, Dub93, Duj96, Duj99, Duj02, Dun69a, Mil22, ER00, Fen11b, Gal86, Gbu81, GU66, Gou81, Gre78a, Hir15c, HL69b, HB77a, HJ68b, Jen71, Joh90, Jon78, KH82, Kim17, Kim19, Kla72, KF73, LaD18, Lan92, Lax74, Lev68c, Lin75, LPS16, Mil20, Moh81, Nyb12, OS20, Pag74, Röd94, Ros72, SO21, Shu72, Ste81a, Ste81b, Szi17, Tri74b, Tri89, Web95b, Wer83, Whi78e, Wlo68c, Wlo71c, Zei89, ZS85, DL20, GGL24, HA66, LMV09, RS08, Gou68]. Problèmes [Spi00a, Spi00b]. Problems [Abr97, Abr00, Abr71, Ano78a, Ano79b, Ano91e, Bas63a, Bas63b, Bas63c, Bas63d, Bol84, Erd81, ES00a, ES00b, ES01a, ES01b, ES01c, ES01d, ES02a, ES02b, ES02c, ES02d, ES03a, ES03b, ES03c, ES03d, ES04a, ES04b, ES04c, ES04d, ES05a, ES05b, ES05c, ES05d, ES06a, ES06b, ES06c, ES06d, ES07a, ES07b, ES07c, ES07d, ES09a, ES09b, ES09c, ES09d, ES10a, ES10b, ES10c, ES10d, ES11a, ES11b, ES11c, ES11d, ES12d, ES12a, ES12b, ES12c, ES13d, ES13a, ES13b, ES13c, ES14d, ES14a, ES14b, ES14c, ES15b, ES15c, ES15d, ES15e, ES16a, ES16b, ES16c, ES16d, FL22, Fro23a, Fro23b, Fro23c, Fro23d, Fro24, Gor12b, Gui77a, HH00, Hee84, Hil64a, Hil64b, Hil64c, Hil64d, Hil65a, Hil65b, Hil65c, Hil65d, Hil66a, Hil66b, Hil66c, Hil66d, Hil67a]. Problems [Hil67b, Hil67c, Hil67d, Hil68a, Hil68b, Hil68c, Hil68d, Hil69a, Hil69b, Hil69c, Hil69d, Hil70a, Hil70b, Hil70c, Hil70d, Hil71a, Hil71b, Hil71c, Hil71d, Hil72a, Hil72b, Hil72c, Hil72d, Hil73a, Hil73b, Hil73c, Hil73d, Hil74a, Hil74b, Hil74c, Hil74d, Hil75a, Hil75b, Hil75c, Hil75d, Hil76c, Hil76d, Hil76e, Hil76f, Hil77b, Hil77c, Hil77d, Hil77e, Hil78a, Hil78b, Hil78c, Hil78d, Hil79a, Hil79b, Hil79c, Hil79d, Hil80a, Hil80b, Hil80c, Hil80d, Hil81a, Hil81b, Hil81c, Hil81d, Hil82a, Hil82b, Hil82c, Hil82d, Hil83a, Hil83b, HP83, HPW83, HPW84a, HPW84b, HPW84c, HP84, HPW85a, HPW85b, HPW85c, HPW85d, HPW86a, HPW86b, HPW86c, HPW86d, HPW87a, HPW87b, HPW87c, Hil87b, Hil88a, Hil88b, Hil88c, Hil88d, Hil89a, Hil89b, Hil89c, Hil89d, Hil90a]. Problems [Hil90b, Hil90c, Hil90d, Hil91a, Hil91b, Hog63a, Hog63b, Hog63c, Hog63d, Hog64a, Hog64b, Hog64c, Hog64d, Hog65a, Hog65b, Hog65c, Hog65d, Hog66a, Hog66b, Hog66c, Hog66d, Hog67b, Hog67c, Hog67d, Hog70b, Hor78c, Kwo17a, Kwo17b, Kwo17c, Kwo17d, Kwo18a, Kwo18b, Kwo18c, Kwo18d, Kwo19a, Kwo19b, Kwo19c, Kwo19d, Kwo20a, Kwo20b, Kwo20c, Kwo20d, Kwo21a, Kwo21b, Kwo21c, Kwo21d, Kwo22d, Kwo22a, Kwo22b, Kwo22c, Kwo23a, Kwo23b, Kwo23c, Kwo23d, Kwo24, Luc03a, Luc03b, Luc03c, Luc04a, Luc04b, Luc04c, Luc04d, Luc05a, Luc05b, Luc05c, Luc05d, Luc06a, Luc06b, Luc06c, Luc06d, Luc07a, Luc07b, Luc07c, Luc07d, Luc09a, Luc09b, Luc09c, Luc09d, Luc10a, Luc10b, Luc10c, Luc10d, Luc11a, Luc11b, Luc11c, Luc11d, Luc12a, Luc12b, Luc12c, Luc12d, Luc13d, Luc13a, Luc13b, Luc13c, Luc14d, Luc14a, Luc14b, Luc14c, Luc15a, Luc15b, Luc15c]. Problems [Luc15d, Luc16a, Luc16b, Luc16c, Luc16d, Luc17d, Luc17a, Luc17b, Luc17c, Luc18a, Luc18b, Luc18c, Luc18d, Luc19a, Luc19b, Luc19c, Luc19d, Luc20a, Luc20b, Luc20c, Luc20d, Luc21a, Luc21b, Luc21c, Luc21d, Luc22a, Luc22b, Luc22c, Mur82b, PRW93, Rab91a, Rab91b, Rab92a, Rab92b, Rab92c, Rab92d, Rab93a, Rab93b, Rab93c, Rab93d, Rab94a, Rab94b, Rab94c, Rab94d, Rab95a, Rab95b, Rab95c, Rab95d, Rab96b, Rab96c, Rab96d, Rab96e, Rab97a, Rab97b, Rab97c, Rab97d, Rab98a, Rab98b, Rab98c, Rab98d, Rab99c, Rab99d, Rab99e, Rab99f, Rab00a, Rab00b, Sin70, Tho63e, WR01, Whi67a, Whi68a, Whi68b, Whi68c, Whi68d, Whi69a, Whi69b, Whi69c, Whi69d, Whi70a, Whi70b, Whi70c, Whi70d, Whi71a, Whi71b, Whi71c, Whi71d, Whi72a, Whi72b, Whi72c, Whi72d, Whi73a]. Problems [Whi73b, Whi73c, Whi73d, Whi74a, Whi74b, Whi74c, Whi74d, Whi75a, Whi75b, Whi75c, Whi75d, Whi76a, Whi76b, Whi76c, Whi76d, Whi77b, Whi77c, Whi77d, Whi77e, Whi78a, Whi78b, Whi78c, Whi78d, Whi79a, Whi79b, Whi79c, Whi79d, Whi80a, Whi80b, Whi80c, Whi80d, Whi81a, Whi81b, Whi81c, Whi81d, Whi82a, Whi82b, Whi82c, Whi82d, Whi83a, Whi83b, Whi83c, Whi83d, Whi84a, Whi84b, Whi84c, Whi84d, Whi85a, Whi85b, Whi85c, Whi85d, Whi86b, Whi86c, Whi86d, Whi86e, Whi87a, Whi87b, Whi87c, Whi87d, Whi88a, Whi88b, Whi88c, Whi88d, Whi89a, Whi89b, Whi89c, Whi89d, Whi90a, Whi90b, Whi90c, Whi90d, Whi91a, Whi91b, Whi91c, Whi91d, Whi92a, Whi92b, Whi92c, Whi92d, Whi93a, Whi93b, Whi93c, Whi93d, Whi94a, Whi94b, Whi94c, Whi94d, Whi95a, Whi95b, Whi95c, Whi95d, Whi96a]. Problems [Whi96b, Whi96c, Whi96d, Whi97a, Whi97b, Whi97c, Whi97d, Whi98a, Whi98b, Whi98c, Whi98d, Whi99a, Whi99b, Whi99c, Whi99d, Whi00a, Whi00b, Whi00c, Whi00d, Whi01a, Whi01b, Whi01c, Whi01d, Whi02a, Whi02b, Whi02c, Whi02d, Whi03, DG67, Flo18, Han11, Kro68, Pic79, Spi00a, Spi00b, Bro68a]. Procedure [Lig73]. Proceedings [ACM12, Bro96, IEE81, SC07, ACM96, BHP92, BPH93, B^{+}96, NS95, Ano17c, Ano18a, Gon89, IEE08, Jef08, Ano17b, Swa92, SM08]. Process [Lon86b, Nef67]. Processes [Hey81, Sch97, Won82, You98, KSSS16]. processors [Edd90]. proches [Spi00a, Spi00b]. Produced [Mel16g]. Producing [HT18]. Product [Ade17, Cha12, Dav82, DeL95, Ewe82, FG23, FG05, KK01, Kno90, LF15, LF16, LF17, Lee02b, Lu70, Mar12d, Mar13a, Mel16a, Nyb12, Sta82, Ste73a, Wag81, Zei72, Sun06a, Zah82]. Production [Wel67]. Products [BH75, Ber75, Bro76, Bun75b, Bun14, Bus73a, Cam22b, DD67, DLR16, Dda20, DT20, EES82, Fer69, FH98, GCMP14b, Gau12, Goo94b, Gou63b, Gou71b, GQ07, Gre18, Har98, Hir86, Hor79b, HP86, KK14, Kim04, Kla68a, KL91b, Kos15a, Kos18a, Kos20, KG20, Kos21a, Kos21b, Kos21c, Kos21d, Kos21e, Kos21f, Kos21g, Kos21h, Kos21i, Kos22e, Kos22h, Kos22i, KW13, Lon73, LH74, LS13, LOT23, LO14, Mel98, Mel99f, Mel00d, Mel00c, Mel01, Mel17a, Mel17b, Mel17c, Mel17d, Mel18d, Mel18b, Mel18e, Mel18f, Mel18g, Mel18a, Nyb04, Pon17a, Pri05, Rab99b, RP18, RP19, Red94, Rus82, Ste21, Szi17, Tad65, Tau73a, Tep76, Tri10, Zei77, Zei70, GCMP14a, KLM^{+}19, Osl07, ST06, WP07]. Professor [Ano96b, Gou72a, Woo97]. Profile [Ros79, Zak83]. Profiles [Wu83]. Program [Ano71c, Ano72g, Ano73d, Ano74b, Bro96, Sho78, BB79]. Programmer [Mae92a, Mae92b]. Programming [AGO95, Big21, Pla96b, Bee04, Kro68]. Programs [CFG19, Web95a, B^{+}96]. Progression [Bac81, GKL^{+}21, GG15, Hsu97, RP19, Ste73a]. Progressions [CS95, DJ67, HS98, Jar67a, KK05, PZ12]. Projection [Gro97]. Projective [MZ92]. Prolog [CV88]. Prometheus [Gal08]. promptuary [KTW84]. Pronic [McD98a, McD98b]. Proof [Adl69a, Alb73, AW66, Beh12, BO17, Bol77, Bro78b, Bru93, Chu85, CK95, Eli21, ES16e, Ewe03, FH20, Gou75, Hag64, Hen17, Hen19, HH72, Hir13c, HE80, Hor94a, KS04, KAP10, Kwo14, LB13, LBF14, Mic64, Nak14, PD18, Pow77, Pri18, Rob02a, Ros88, San90b, Sha75a, Sin73a, Som72, Spe77, Sun73, SC84, Tri73, Woa02, Yab01, Yam80, Yam19, ZG17, BQ03, Hol95, Len06a, MM06, And05]. Proofs [BQ03, BR14, BCS19, ES10e, Ger09, GG15, GZ22, Hay92, Lar15b, PP00, SW10, Sha13, Woo07, And05]. Propagation [But90, Win68]. Properties [Aga89, Alf63i, And14, AJ95a, Ark70a, BJ90, Bis84, Bri83a, Bri83b, Bru79, Car72a, CH94, CMR02, DQ83, Djo96, Djo01b, Djo05b, DP78b, DMSS16, Erc73, Erd81, ES16e, FJ12, Hal66, Hal67, HPV95, HL74, HBJ78b, HBJ78d, Hor65, Hor67, HF91, Jai74b, JK98, KK95, Kim76, Kim80d, KS16, Kno92a, KR10, Lás82, LL87b, LL87c, LL87d, Lee97, Lee01a, LL83, Len03a, Li21, Lon81d, MST13, MS94, MD87, Moh76, MR99, OS98, OS05, OS21, PTW13, PT17, Per16, Pla96a, Pon19, RS11a, RR14, Rao53, Rob82a, Ros83, Rou11, SH72, Sha74e, Sha84b, Sha60, SR83, Sit70, Sol76, Som80, STL22, Swa66b, Swa68, Swa99c, Swa00c, Tan00, Tur74, VKSAN93, Wad78, Wad92a, WH71, WP69, Wil77, Yal72, Zha97c, Zho03, ACOM21, EB17, GSS19]. properties [HLS79, KI16, SE13, dLM94]. Property [AE79, AM80, Alb81, AS77, AL19, AZ20, BD81, Bic68a, Bos82, BAP73, fC95a, DD02, FH20, Gam95, Gel63, GR83, Goo94b, Gou72b, Gra64, GWT15, Hah75, HP94, HA71, HH78, HP75b, KC86, KsS72, Kui82, Kun87, Lar88, McI92, Mic64, Nag84, Nyb03, Pad94, Pla94, PW01, Rob02b, Sch13b, Sil76, Sin80, Ske65, Spi06, VWT85, Wei66a, Whi67b, WTvR87, Wlo68a, Zho97, Boy97, Mil11, de 81b, de 95]. Proportion [BS97, Hun70b]. Proportional [Hil81e]. Proportions [Bom20, GM11, Gri10, Low71a, Nor64, Nor72, St.84, BF90, Car39]. proporzionali [Car39]. proposal [Høy05]. Proposals [KE14, Kim17]. Proteins [JR98]. Protocol [JQ97]. Proving [Hen22a]. Pseudo [Esw78a, Esw79a, Fer68, Geb67, Jan02, Leh64, Par70, AI92, Esw78b, Esw79b]. Pseudo-Fibonacci [Esw78a, Esw79a, Fer68]. pseudo-Lucas [Esw78b, Esw79b]. Pseudo-Primes [Leh64, Par70]. Pseudo-Random [Geb67, AI92]. Pseudogroup [Fer78]. Pseudoprimes [AJ96, Arn95, BW80, Bru94d, Bru94b, Bru96, CJS98, CG03, DF89, DFM90, DF92, Di 93, Dre97, EKS88, GLMS22, Hau96, PS80, PSW80, RZ95, Som09, SC10b, SK22a, SK22b, Arn97, GP91, GP93, Som06a]. Pseudorandom [EH95, MCPR95, MRPC95a, MRPC95b, VKSAN93, CMP94, Krá72a, Krá72b, Mas94a, Mas94b, MS00a, MS00b, PW95, Pry94]. Psi [Alb81]. Psi-Type [Alb81]. PSW [CG03]. Public [Smi93, Mül06, SS95a, SS95b, XLP99]. Public-key [Smi93, Mül06, SS95a, SS95b, XLP99]. Publications [Bee11, Puc62].

Publishers [Sch94]. Pullman [Fre94]. Pulsated [ADS14]. Pure [Her19, Sha06]. Purely [MO09]. Putnam [Woo97]. Puzzle [Gou66a, Gou66b]. Puzzled [Win11]. Puzzler [Hol65]. Puzzles [Bro68a, DG67]. Pyramid [AS05, Bea68, Hog72a, Sha77]. Pyramids [Ano93f, Ano93g, Bol86, Ano93e]. Pythagoras [Mar93, Uma71]. Pythagorean [Ano90e, BJ79a, AS20, Aus23, BW79, Ber82, Ber89, BJ79b, Coh70, Els03, FL68, Hen67, Hin92, Hor82b, Luc02, MM90, Oli96, Rai48, Rob06b, ST89, SH71a, Sil98, Ste86a, Tay76, Ter12, Tri09, Uma72, Wat92, Weg75, WW75, Weg77, Weg81b, Zel93, Zel97].

Q [Kum21]. Q-transformation [Kum21]. QRT [GG13]. Quadrate [RS98b]. Quadratic [BD81, BCHR91, Bea03, Bou20, Bro71a, BK96, CD13, DB66, Gic09, Hin78, Leh66, LM10, Lin68, Mol88, Mon85, PVB98, Tay75, Vse11, WP19, Wie88, You95b]. Quadratics [Mah91]. Quadrative [GKRS01]. Quadratorum [McC19, Arr69, Baš78b, Baš78a]. Quadrinomial [SH79c]. Quadruples [Oli96, Rie99a]. Quality [Jan02, MCPR95, CMP94, Mas94a]. quand [CHF85]. Quantum [Fer78, LLP^{+}18]. Quart [Bri83a, Bri83b]. Quarterly [Ano63c, Joh12a, Bee11, BJ87]. Quartic [Tay76]. Quasi [AM80, Daz02, Hen15b, Hor03a, Tau89, NS95]. Quasi-Cyclic [Daz02]. quasi-Monte [NS95]. Quasi-Orthogonal [AM80, Tau89]. Quasi-Periods [Hen15b]. quasicrystal [Jag21]. Quasicrystalline [Man17]. Quasigroup [Kum21]. Quasiperfect [Coh82]. Quaternion [Iak77a]. Quaternions [Abr09, Hor63b, Iak77b, Iak77a, Iak81a, Iak81b, Iye69b, Iye69c, Swa73]. Queries [KTW84, BDL^{+}16]. quest [Dev17, Sim18]. Question [GQ09, Li99a, Pro97]. Questions [GB15, dF11]. Queueing [Mag81]. quickies [Tri67e, Eve68]. quickly [CH89]. Quilt [MN20]. Quinn [And05]. Quintic [DES17]. Quintuple [Ewe82]. Quintuple-Product [Ewe82]. Quotient [Dob14, PT17, Wil82]. Quotients [BH77b, Bun81, GL16, NB89, Coo06].

R. [Kui69]. Rabbit [Alf63b, BJ78a, Dub93, Fen11b, HL69b, Wel67]. Rabbits [GKH77, HD19, Hog68a, Lev88, Gri15a, Hal22]. Rabin [Arn97]. Rabinowitz [Ano00d]. Radical [Ell12, OR14]. Radicals [GCMP14b, Osl07]. radice [Car39]. Radici [RS98b]. Radix [Kno90]. raising [Duv91]. raison [CHF85]. Ramanujan [Art04, Cha07, Che12, CC20, Coo14b, EM15, Gui20, HS13, Hir15c, How86, HH89, Lau10, LM19b]. Ramified [Stă14]. Ramsey [AGJ^{+}09, SC84]. Ramus [Kah09]. Random [Alu96, Alu97, And90, BK00, BGTM^{+}11, Cha05, Col87a, DGNW85, Edd90, Geb67, Gri11b, Gri15a, Hop95, Jan02, JV19, Kou90, Lan11, MM65, MZ93, Pet94, Pet91, Pet99, Sta85, TRE73, Vis00, Alt88, AM15, AI92, Cha06b, CM10, ET99, HS18, JRdlR08, JRdlR10, L’E94, LSS12, Mak94, MM06, MT85, MNZ90, MZT90, MZ91, Mar92, Neu17, TLC93, Wol92]. random-number [LSS12, Pet91]. Randomness [MT02]. range [BP10a, BDL^{+}16]. Rank [BE07, DK18, Des78b, Gup78b, KP77, Rob76, Sal75, SK13b, Vin63, Wil75b].

Rank-Vector [Gup78b]. Rapid [All75c]. Rapidly [Cas86]. rapprochement [CHF85]. rare [THKH23]. Raster [CLRS86]. Raster-Graphics [CLRS86]. Rate [Cha05, Lan11, Len17, Cha06b, JRdlR10]. Ratio [Ber66, BJD95, BJ99, BF91, BH17, Cad16, Cha73, Gal79, Gal86, GD09, Goo93b, Goo93c, Hin77, HBJ79b, Lar88, Lar15b, Mel10c, Pet86, Pet96, Pro15, Rac19, Rus12, RS11b, Sch81, SH77, SG85, Sza17, Wlo63, Wlo67, Wlo68b, Wlo71b, Wlo71d, Zuk76, Dun97, Gri13, Krč06, Nea07, RL15, Ade14a, CVZ17, LSW14, WB12]. Rational [BCM10, Bro03, Bro74b, Bur90, CR93, Els00, Els03, Fox01, Hau93b, HT18, JS16, Kla81, Law83, Lin21, Nag01, Nak20, Sha93, Bou20, Gri09]. Rationally [SSS24]. Rationals [D’S89, Els96, Leu23, Nor19]. Ratios [Alf64a, Bea22b, BFT15, Bro70b, Bro71a, Den87, Fee67, GM81, HA77b, KM95, Moo93a, Som75, TR01, WB12]. RATS [SC01]. Ray [MC76]. Ray- [MC76]. readability [Wat86]. reading [Lün92]. Real [ABBB13, Bou20, Bro71b, D’S89, DH68, Els98, Fil93, Fri66, HF95, KK89, Kno90, Mol88, Van96, Nyb05]. Realizable [MW22]. Really [And05, Bru93, Fis81, BQ03, Cri07]. Reals [Mel10c]. Rearrangement [Gou77b]. Rearrangements [Def13]. reason [CHF85]. Recamán [AMS^{+}22b]. Rechenkunst [Lün91a, Lün91b]. Reciprocal [AJ91f, Bac81, Bes91, Bra71, CD13, Goo74a, Gre77c, Her03, HB76d, How87, HSL01, JK98, LK09, MS95e, Mel99b, Mel03b, Mel15, Mel18a, ON09, Pla96a, Pop86, Zha99b, Zha01b, ZW04b, KA13]. Reciprocals [ASD15, AJ90, AJ97, Buc67, CCD^{+}22, Goo94b, Gre77d, Gre77b, Gri92, Hen11a, Her85, HT93, HB76c, Hor88a, Jen94, Kos22h, Kos22i, McD94a, Mel00c, Mel01, Mel02, Mel18g, Nyb03, Nyb23, Pop84, Rab99b, SH71b, Tho63f, Sco18]. Reciprocity [Car84, Eng91, Tay75, Tay76, You95b, BB04]. Recognition [SK81, TV07, Ves15]. Recombination [CT75]. reconciliation [CHF85]. Recounting [BQ99]. Recreation [Bro67e, Bro68d]. Recreational [Dev12, Gla70, Mad68a, Mad68b, Mad68c, Mad68d, Mad69, Ped71, Han11]. Recreations [Ano95c, Ste95]. Rectangle [BL65, Hun63]. Rectangles [Bas63f, BH69b, Lig73, Lig79, Rea80, Ric13]. Rectangular [BH17, DP95, MS03, Tha19]. Recurrence [AD10, AE79, AM75b, Alp20, AH95, And95b, AP92, AJ91e, AB85, AP70, Bas20, BCFS22, BO17, BCD^{+}20, BHL^{+}20, Bot82, Bro89, BsS95, CS03, Car78e, CJS94, CM89, CCD^{+}22, DD02, EMR02, Eli01, Eli21, Els05, Fer78, For67a, For67b, FL68, Fre85, Ful80a, Ful80b, Ful81a, Ful81b, Ful81c, Ful77, Gou75, HW12, HKW15, HR91, Hau97b, HA77a, HH82, Hor84a, Hor88a, Hor92a, Hor94c, HST94, Jac90, Jes63b, Jes63a, Jes64, JK93, JS19, Kim91a, Kim91c, Kim10c, Knu66a, Kos19c, LK09, LF16, LF17, LF19, Li99b, LPS16, LS18b, MV84, MJ95, Mom01, MKL^{+}23, NS90, NR15, Pop77, Rus81, Say79, Sed14, SH72, Sha79c, SH79a, Sha80, Sin70, Sli98, SH80b, Sta03, Sti98, Sul89, Swa98, TMR02b, TR04, TT16, TZ96, Vol13, Vor09, Whi66b]. Recurrence [Yap70, ZW02, ZL95, Car07, GQ07, Hir05, Knu66b, Ron91, Ten09, Wag07, lY08]. Recurrence-Generated [HH82, Hor84a, Hor88a]. Recurrences [Alp19b, ADS14, BKM17, BW81, CP84, Car78a, Cat74b, Cat74c, CFH^{+}16, CFH^{+}17, CM01, DG84, Ewe80, Geo89, Gou63c, Gou90, Gry03, Hor92c, How94, How98, Izo02, KK14, KM05, KW13, Kwo10, Lev85, Li00a, Lit20, Liu92, LB91, LHLT95, MS20, MMMS22, MMM^{+}22, MO09, MO10, Pla96a, Rab96a, Rab99a, Rob91b, Rob02b, Rob04a, Sel84, SY84, Som79, Som80, Som84, Som87, Som89, Som91, SC04, SK13a, SK15a, SK16, SK17, SK21, Spi97, SS17, WH14, Wol98, You03, Bre94, BN06, GPS87, Hir06a, Hir06b, KA13, PZ12, Som06b, Som07]. Recurrent [Alf63e, AHM92, BC10, BD81, CK24, Cav77a, Coh91, Eme69, Hen11a, HL66, HSL01, Kim76, LU05, Rob76, Zei71a, Gou06]. Recurring [Aga89, AJ91d, AJ97, Ark69, AAB^{+}89a, BH75, Bri64b, BN03, BN19, Bro68e, BF95, DJ68, Gol95, Jar66, Mel02, MZ92, Sol76, Yal73, Zöl93, Bro66, Bro67b]. Recurring-Sequence [AAB^{+}89a]. Recursion [Bro68b, Bro69g, Bro69e, Bro69d, Bro69f, Bro70c, Bro70b, Bro72a, Bro74a, Bro76, Cam22a, Hay92, Hic73, Kah09, Muw81, Par69, Par77a, Pha84, Rob82a, Sti98, Wat86, Whi86a, Whi67b, IK87, RSHL07, Wir08]. Recursion-Type [Hic73, Muw81]. Recursions [DV89, Hen22b, MRPC95a, MRPC95b, Naj93, Mas94b]. Recursive [BH78a, BJF99, BL65, BD70, Bro75c, CFG19, FKL70, FP86, GBC16, Hen20, HA77b, HH80, Hol66a, Hol66b, HCD97, KI10, KZ92, KsS73, Lás82, McG09, McG10, Nag84, Par64, Rap74, RM11, Sal75, Sha74b, Sha74e, Sha78b, Sha83, SO02, Shi95, SJ84, Sun01b, Sza17, Tri65, WS91, Whi66a, Woa02, Zei70, Cox87, Man05a, MSS08, Rob07, Rob99]. Recursively [Hen15a, Mil75b]. Recursiveness [Hen11a]. Rédei [BCM10]. rediscover [Dev17, Sim18]. Rediscoveries [Gou77a]. Reduced [BCHR91, BN77b, BN93, Li00a, Pow96, Sun75b, Wan93, WW06]. Reduction [AB85, Car71a, Mel02, Rei93b]. Reductions [CLX08]. Redundant [BS97]. Redux [Coo14a, DS19]. Referees [Ano87a, Ano88b, Ano89e, Ano90f, Ano91f, Ano92c, Ano93j, Ano94e, Ano95d, Ano96c, Ano97g, Ano98c, Ano99e, Ano00f, Ano01c, Ano02d, Ano03c, Ano04b, Ano05a, Ano06a, Ano07b, Ano10c, Ano11d]. References [OR99]. Refinement [HT00]. Reflected [Aga87]. Reflections [Bro89, Bru86a, HBJ79a, JH73, MW73, SH02]. Regarding [Rob95]. Regeneration [Sta85]. regola [Mar74]. Regula [Tho63d]. Regular [Her81, Kla73b, Kre80, Kun90, LU05, MS03, Rob14, Wlo71e, EB17]. Regularity [MR15]. Related [AH77, Alv69, AJ91f, Arc84, AKO^{+}03, Bac81, BVB95, BBS81, BWH75, BH77b, BL65, Bic75, BB90, Bra90, BD72, BS93, BL17b, Car64b, Car66c, CSV73, Car78a, Car78g, CH02, Cha81, Cha82, Cha84b, CM14, CLM22, DeL95, DD02, Dor99, Dre97, DMR^{+}16, Edw09, EG16, Eme66, ES16e, ES17, Fie67b, Fie74, Gil92a, GG13, Gri15b, GC90, Hen67, Hod92, HB76a, HB76e, HBJ77d, HBJ82c, Hor00a, How85, Jai74a, KMT22, Lan00, Lan02, Leh75, Lon81a, Lyn73, Mac10, MP66, MS95e, Mon85, MR95, NC85, PT17, Par70, PS76, Röd94, Ros83, SC98, Van96, Wal76a, Wan05, Zha99b, Zha01b, ZD02, ZD04, AK05, CHS72, CS92, CM15, DRS11b, DRS11a, EBK89, FZ08, IA14, LO14, Mag09, MSS08, SE13, Wag07, Woo07].

Relating [AM75a, Fox01, Hic78b, Hic78a, Tay76]. Relation [AE79, Alf64j, And95b, AP92, AH75b, BCFS22, BCD^{+}20, Bro89, Bro72g, Bru86b, Cam22a, Daf20, FP86, Ful80a, Ful81c, Ful77, Gou72c, Gou72d, Hau97b, Hos73, HST94, Kwo14, LS18b, Mar15, MJ95, Par69, Pha84, Say79, Sha79c, Sha64a, Sur14, Veg74, Vin63, Whi86a, Woa02, Yap70, Rem11]. Relations [Alf65a, Alf66b, Alp20, AJ91e, Bot82, Bro68b, Bro69g, Bro69e, Bro69d, Bro69f, Bro70c, Bro70b, Bro76, Byr75a, Byr75b, CM89, CCD^{+}22, DD13, Edg16, EMR02, Els96, ES12e, For67a, For67b, Fre85, Ful80b, Ful81a, Ful81b, HB74a, HA77a, Hor92a, Hsu87, Jes63b, Jes63a, Jes64, Kah09, KT19, MKL^{+}23, Par77a, RP71, SH72, SH79a, Sha80, Sin90, Sin70, Som89, Sti98, Sto67, Swa98, TMR02b, TR04, Tue21, Vol13, Whi66b, Whi67b, AD10, Car07, Hir05, Ron91, SE20b, lY08]. Relationship [Gou89, Hew77, Hin77, PB15, Str76, Can51, Gin54]. Relationships [ES17, HM72]. Relatively [All81, El 09, Gou64b, Lay66, Tót02, Sho06]. Relatives [Kos16, Rus12, Vol10]. Relaxed [DGST88]. Reliability [KF73]. Relost [Bri64a]. Remainder [BD71, Gla95, Moo92]. Remark [Ata95, BS93, Ell12, Gal98, Kui69, KsS73, OR14, Röd94, San89, San23, Sch00, Sla00, Veg69, Geg93]. Remarkable [And17, Zar70]. Remarks [Bal94, BBL95, Bri64b, Car78h, Car80e, Cre70, Dav76, Fen11b, Fie67b, Kan15, Mam61, Mor10, Pih01, Pih10, Sta76, Sta80b, Wal63, Was99, Mül06]. Renaissance [Mos08, Giu04]. Rencontres [DD87]. Renegades [BJ89, BJ93]. renewal [Chr12]. Reorder [You95a]. reordering [Sar90]. Repblocks [KLT18]. Repdigits [AFRT23, FL18, RP18, RP19, SP19]. Repeated [Sha11, Sin75]. Repeating [BJ89, BJ93, Lin90]. repetition [DP24]. Repetitions [DN96, AB00]. Replicating [DvT83]. Reply [Bro65b, Sub07]. Report [Ano02e, Ano17a, BJ18, BJ19, BJ20, BJ22, Coo10, Coo16, Fre88, Fre90, Joh09, Joh12b, Joh20, Joh22b, Phi01, Phi04, Phi07b]. Reports [dBF85]. Representable [Bro75c]. Representation [Alf63f, Bev71, Bro72f, BT91, Cap88, Cap89, Chu20, Day69b, DvL23, Dez77, Dil87, DD95, Eng01, Fer65, Fil86b, FF89, GZ15, Gri15b, Hal83, Hin80a, HH84, Hor94a, Hor94d, Hor96b, Hor97b, Hor99, Jon75, Jon76b, Kar04b, Kla66b, Kle91, Kno90, Laf64, LU05, McD95, Mel10c, Nyb01b, Nyb02, Pla96a, Puc02b, Sto09a, TS89b, TR90, Zec72b, ÖTA05, Ste23, Wal10, Zec72b]. Representations [ABCM12, ABJ11, AF87, BF16, BJ85, BJF99, BJF01, BJ02, BJ03, Bro69i, BDM12, Bun92, Bus73a, Cal72, CW22, Car68b, Car70a, CSH72a, CSH72b, CSH72c, CSH72d, CSH72e, CSH72f, CSH72g, Dek20a, DDK^{+}14, DG98, Ewe92, Ful80b, Gri10, Gri14a, HKW15, HB63, HP72, HB77b, HBJ79b, HBJ82b, HBJ82c, HBJ84, Hor94b, Hor00c, HM72, JR98, KK89, LLKS01, LN81, McD82, MN76, O’C72, Pih88, Shu20, SG76, Sil77a, Sil77b, CHS72, EH07, EZ04, Kim07]. Represented [BJ89, BJ93, HW81, LL87a, Lin90]. Representing [Fil98c, Gal72, HW12, Rob87, Ula23]. Represents [Adl78]. Reproducible [MCPR95, Mas94a, Pry94]. Repunits [KC89, Wit16, GGL21]. Required [Har70b]. Rescue [Hun70a]. Research [Ano70e, Ano74c, Ano77b, Ano82b, Ano84b, Ano84c, Ano86c, Ano95e, Ano95f, Ano95g, Ano96d, Blu72, Bro69b]. Residuacity [Tay76]. Residue [AC13, Li00a, Sun75b]. Residues [Alf64e, Bur71, Coh67, DeL84, Gic09, Hal64, Hol94, Hol00, Jac90, Man77, Mon85, Som81, Som91, Tay67, Yal77, Jav22, Man68, SK17]. Resistance [Nod90, IT91]. Resistor [Mye75]. Resizable [Boy98]. Resolution [Sza02]. Resonance [KŽ05]. Respect [Man68, Som80, Som82]. Response [Hog76, Wal84b]. Restricted [Abr76, Car76a, Car79a, Car79b, Car80f, Col87b, DMP07, Dif12, Els96, Els00, Ewe80, Moh67, Sul89, Sho06]. Result [Ava13, Boa01, Cas91a, CK95, Dav77, Els98, GC90, Hun67, LR85, LB13, LBF14, Lee87, MS95a, Ste88, Tri73, WB12, Mac88]. Resultant [Akr93]. Resultants [Rus12]. Resulting [Mad69]. Results [AGJ^{+}09, Bia95, BB90, CW85, Gil92a, GG79c, GB15, HG81, HP72, Hol11, Iye69c, Kla65, LM17, Lon81a, MH75a, Pro16a, Rap70, Rug63, SBM97, Ste86a, Tau89, Tos78, Tsu21, Wel67, ZW04c, CM10, DRS11a, ZQZ20]. Retirement [Ano97h]. Retires [Ano91e]. Retraction [Buc64c]. Retrieval [Lyn73]. Retrograde [BJ89, BJ93]. Return [Neu17, HW15]. Reverse [FG23, You92a, You92b, JH05]. Reversing [Ava12]. Review [Alf66a, And05, Ano88a, Ano90e, Ano16a, BJ99, Bic72a, Bic73a, Bro66, Bro67b, Bro68a, Bro70a, Cad16, Cio18, Eve68, Gal08, Gau02, Gau10, Gra89c, Hal22, Har09, Høy09, Sch94, Sim18, Tol17b, Tol17a]. Reviews [Bro69a, Mos08, Oak09]. Revisited [Boo75, Bum87, CD78, Dub93, Hor85a, Jam90, KG17b, Kos17, KG18a, Kos18b, KG20, Kos21b, Kos21d, Kos21h, Kos22e, Kos22d, Kos22g, Kos22b, Kos23b, Kos23e, Kos23g, Kos23m, Kos23q, Kos24a, Li99b, Mar09, Mel04b, NN87, Pla97b, Rob05a, Uma71, Nyb05, TP86]. Revived [Alf63b]. revolution [Dev11]. revolutionised [HD19]. rewriting [Kan14]. Reznick [Nor22]. Rhombus [KMST97, Ram16, YG18, YY19]. Rhythmic [Low71b, Sha05]. Riccati [Lan04]. Rich [CS12]. Richard [BJ99, Cad16, Ano99d]. Riesel [BFF11, LH09]. Right [BJ89, BJ93, Bru94a, Lin90, Nyb98, Sch81]. Right-Angled [Nyb98]. Ring [Bun75c, DeC70, DeC73, Rob76, Sun75a]. Ringel [Kwo10]. Rings [Bun81]. Riordan [Ano15a, Slo15, NS05b, PW93, YG18, ZD02]. Rises [CP09, HCG03]. Rising [Pro16b, SP77, Swa00b, TT16]. RIT [Coo14a]. Riven [Gru01]. Robinson [BJ09]. Robust [AF87, Cap89]. Rodriques [Hor97c]. Rogers [Art04, Cha07]. role [Ola00]. roles [Høy05]. Romance [Bic68b]. Romanian [Wes80]. Romans [Spi00a, Spi00b]. Rooks [Fra70]. room [HHP02]. Root [Alb81, AH95, Bic65, DeL95, Goo93b, Goo93c, Leh66, Boy97, Car39, Nea07]. rootfinding [Mus93]. Roots [Bas64, Bau71, BP99, BCFX09, Bre74a, Bre74b, Dam89, DeL77, DB66, DB67, DB70, Hau94a, HB73a, HH82, Jon91a, KP77, KS10a, KS10b, Kon95, Li99a, Li00b, MS15a, May82, Mol88, Mon76c, Pet86, Pho91, RT04, San90a, Sha72a, ST73, SH80b, TR04, Fre06, RS98b]. Roulette [Boo75]. Rounding [CC03]. Routine [Tri71, You93, You95a]. Row [AC11, Bol84, SP77]. Rowed [BH80a]. Rows [EA12]. RSA [JQ97, Smi93]. RSA-Like [JQ97]. Ruggles [CMM^{+}17]. Ruin [KH82]. Rule [Bal15, Gre76, Hsu93b, Sco68b, ZD04, Cla08, GPS87, Mar74]. Ruler [Puc62]. Rules [KM14, Pel72, Sha94, tR84, NS94, Ric93]. Run [Cha91]. Runs [Mos93, PM85]. Russian [Baš78b].

S [Bic73a, Eis51, Gal08, Gla70, Ped71, Puc62, HS12]. S. [Moh81]. Saalschützian [Car76b]. Sadek [Mel16e]. Sale [Ano93d, Ano22d]. Same [Hir86, Kos15a, PG89, PA91, Pih92, BP07]. Satisfied [Mel16c]. Satisfy [Fox21]. Satisfying [AJ91e, Bev76, But78, HBJ78c]. Saturated [HM83]. scalable [MS00a, MS00b]. Scaled [LF16, LF17]. Scalene [Pec65]. Scenario [Gri11d, Gri15a]. Schaake [Ano90e]. scheduling [PC13]. Schema [Fer76]. Scheme [CT75, Tan98]. Schinzel [Pih91]. Schmidt [Mig77]. Schnirelmann [PB89]. Schöffl [Bre98]. School [Bic69, BB79]. Schools [PB15]. Schreier [BCFS22, CMX20, Chu23]. Schreier-Type [Chu23]. Schröder [LB91]. Schur [Kel92]. Sci [FS14]. Science [IEE84, IEE08, Ola00, SO09, Zei96]. Sciences [Høy09]. scientific [NS95, Ola00]. Scienze [Cio18]. Scotland [BHP92, BPH93, Fre92b]. Scott [Sco68c]. Scrapbook [Sco68c]. SDR [Kla73b]. Search [AW66, CP79, GR93, Hag87b, Has81, HG81, NN87, OW64, Ove65, Ove73, Shu72, Tue23a, Wit72, Boo79, DOU08b, Kah05, Kah06, KC63, Kro68, MR07, ORSW84, Ove67, PG13, PHJ67, PP79, Rou13, SYY04, Sub07, Var89, YK03]. searches [Fra84]. searching [Fer60, Jun61]. sec [Arr69]. Secant [AN83, GB06]. Second [Ade18, AJ97, Ata86, Bas20, Beh12, Bic68a, Bri64b, Bro69e, BsS95, Cak98, CK24, CS03, Car80c, Car80d, Cat74b, Dav14, DJ68, FH93, FH94, Gau04, Geo89, HL66, Hor94c, HS85, Hsu93b, HSL01, HST94, Izo02, Jai74a, Jon78, KK95, Kim91a, Kim91b, Kim10c, Kla66a, KW13, Kun90, LK09, Lee97, Lee01b, Len94b, Li99b, Li00a, LB91, MO09, MO10, MS94, Mel02, Mel17a, Nor05, Par77c, Pla96a, Pop77, Pop86, Rab99a, Rob82a, Rus81, Sha83, Sit70, Som79, Som84, Som91, SC04, Som06b, SK13a, SK15a, SK16, SK17, SK21, Spi97, Sta03, TT16, Yal73, Zha97b, ZL98, AD10, Hir05, Hir06a, WP07, Fre87]. Second- [Yal73]. Second-Order [Ade18, AJ97, Bas20, Bro69e, BsS95, CK24, CS03, Cat74b, DJ68, Geo89, Hor94c, HS85, HSL01, Izo02, Kim91a, Kim91b, Kim10c, Lee97, Lee01b, Li99b, Li00a, LB91, MS94, Mel02, Nor05, Pla96a, Pop77, Rab99a, Rob82a, Rus81, Sha83, Som79, Som84, Som91, SC04, SK13a, SK15a, SK16, SK17, SK21, Sta03, Zha97b, ZL98, HL66, HST94, Som06b, Hir05, Hir06a]. Section [Bea66a, Bro17, Mil22, Erc73, Fow82, Gou64a, Hed76, HBJ79b, Kim19, Lar78, Mil20, Rie99b, Shu72, AAF15, Bea50, KN08, Vaj89, Vaj08, EF98, Ove65, RT02]. Sections [Bra00, Eng87]. Sector [Fie74]. Secure [Smi93]. security [Kum21, LTT95, XLP99]. Seeds [CS98, CS01]. Seeking [Alf65d]. segmenti [Car39]. segments [Car39]. Seidel [FT14]. séjour [AV03]. selection [Var89]. Self [Ben06, BYD^{+}19, Bun06, Cai96b, DvT83, GP16, Gra82, HH80, HPL93, JO24, Kub04, Lar20, Oli96, Sch17, Sch22, Wan05, Zei96]. Self-avoiding [Ben06, Zei96]. Self-Counting [Sch17, Sch22]. Self-Generating [Gra82, HH80, Oli96]. Self-Inverse [Wan05]. Self-Replicating [DvT83]. Self-Similar [GP16, HPL93, JO24]. Self-Similarity [Kub04, Lar20, BYD^{+}19]. Semi [Bro63b, Col88a, Hor78c, Weg77]. Semi-Associates [Weg77]. Semi-Completeness [Bro63b]. Semi-Perfect [Col88a]. Semi-Solved [Hor78c]. Semicircle [DeT92]. Semigroup [Sir97, Sun75a]. Sensitive [Moo93b]. Sentance [SB93]. Separable [Wan00]. Separation [KL91b, KL93]. Sequence [AN79, AMS^{+}22b, Alf63h, Alf64j, All75b, AL79a, All17, And70, AC11, And69, AAB^{+}89a, AAS85, Ata86, Ata89, Ata95, AZ20, Bac66, Bas63e, BH63a, BH63b, Ben02, BH76, BH77a, BH77b, Bic75, BSY05, Bra90, Bri64b, BN03, BN19, Bro68f, Bro72a, Bro72d, Bro75a, BD70, BH78c, Bro78c, BDM12, BC77, Bru84, Bru72, BA98, Bru70, BH63d, Bun75c, Bur71, CSH72g, Car78d, Car80a, CFH^{+}14, CK12, Chu21b, CLM22, Chu23, fC95a, CY00, Coo84, Cus68, Dan98, DN96, Dav71, Day69a, DeC70, DeL70, DB01, DD02, Des78a, Des78b, Ehr89, EV83, FBJ01, Fis76, FK72, Ful78, Gam95, Gan59, Gau04, Geb67, Gel63, GG79a, GG79b, GG79c, Gou71a, HKW15, Hau02, HK75, HK78, Hei64, Hey80, Hig87, Hin17, HH85, HB64c, HB76b, HBJ77d, HB77a, HBJ82c]. Sequence [Hor61a, Hor61c, Hor63a, Hor65, Hor67, Hor69, Hor74a, HS81, Jac89, Jar46, JK98, Jav22, Joh40, Kel08, KK11, Kim91c, KMT22, KP77, KZ92, Kla81, KS10c, Knu64, Kuh12, Kui69, KsS72, KsS73, Kui82, KÖS22k, LF17, LF19, Lau77, Lay66, LL87c, LL87d, Lee97, Mar12a, Mar13b, Mat73, May68, McC72, McD94a, McD96, McH82, MR15, Mel00b, MPPW22, Min81b, Moo93a, Mor83, NB89, Nor14, Nyb23, O’D79a, Ozb12, Pan07, PT17, Par70, Par64, PM82, Phi83, Pil12, Poo88, Pra82, PR02, Pru67, PW01, Raa63, RS11a, Rao02, Rao03, Rao05b, Reb75, Rob76, RH63, RP71, Sch82, SDH77, Sha68, Sha74e, SH79a, SHB84, Šiu11, SŠŠ22b, Sma77, Spi82, SJ84, SJC92, SCJ95, Sta76, Sta80a, Sta80b, Sto17, TRE73, Tur79, Vin78, Vin63, WS67]. Sequence [Wad78, Wad92a, Wad92b, Wal84a, Wal80, WH74a, Wan20, Woa02, YK77, Yam80, Yoo21, ZG17, deB74, AAF15, BE07, Bil14, Dev99, DP24, EMG17, GGL24, GGL21, Gou06, HW15, KI16, Kil07, Knu90, LP14, Mil11, MSS08, Ogu20, PZ12, Ran66, Rao05a, SI05, Sha06, TD16, TD19, BT91, Her14]. Sequence-Generated [HS81, SHB84]. Sequence-to-Sequence [Bro78c]. Sequences [Abr97, Abr09, ABCM11, Ade18, Adl71, Aga89, AS73, AH15, Alf63e, Alf64h, AH16, Alo76, Alp11, Alp18, Alp19a, AH95, And95b, AH97, AP92, AJ90, AJ91d, AJ91e, AJ97, And74a, Arc84, Asv87a, Asv87b, Ata89, AHM92, AG78, AS20, Ava11, Ava12, Ava13, AD98, BC10, BL13a, BL13b, BL17a, Bal19b, Bal19a, BBB04, BHS11, Bas63f, BD81, BTD22, BLP10, BD11, BWH75, BH75, BH77b, BH78a, BH80b, BMRS01, BRS04a, BMRS04, BJ89, BJ93, BJE95, Bic75, BGW14, Bis84, Blo65, BHL^{+}20, Bol84, BYD^{+}19, Bou20, Bra00, BBL14, Bri92, BG21, Bro68e, Bro72e, Bro74a, Bro75b, Bro76, Bro63b, BD70, BD72, BF91, BsS95, Bru86a, Bru79, Bru00, BF95, BL17b, But78, Byr65, CCT05, CW22, CK24, CZZ11, CZZ14, CJ96, CS03]. Sequences [Car72a, Car72f, Car74a, CS75b, Car77a, CS77, Car78b, Car78c, Car80b, Car81a, CJS94, Cav77a, Cha05, Cha86b, Cha86a, Cha90, CMR02, Cha82, CZ10, CMX20, Chu21c, CS85, Coh91, CK16, CBH17, Cre86, Cre70, DRT11, DFLT14, DvT83, DGNW85, DeL95, Den87, DD13, DM10, Djo00a, Djo05b, DJ68, Dre97, Dro00, DMRS97, DT71, Edg72, ER00, Eli01, Eme69, Eng91, EBJ00, Erc76, Erc79, Erd81, Ess03, Far86, Fie67a, Fie67b, Fie68b, Fil91, FH93, FH94, FF94, Fil97c, FF99a, Fil00, Fin91, FKL70, FW78, FOR89, For67a, FL68, FP86, FF89, Gab70, GP16, GD09, Ger77, Ger78, Gic09, GS95, Gol95, Goo68a, Goo68b, Gou74b, GKH77, Gou90, Gou99, GBC16, Gre18, GM78, Gre76, Gre78a, GG13, Gri18, Gro97, Gru94, Gru09, GH18, Gup76, GM96]. Sequences [HS15, HW12, HS19a, HS66, HR91, Hau93a, Hau94a, aHZ02, Hen11a, Her73, Her03, Her19, HLW94, Hig79, Hil74f, Hir11b, Hir12b, Hod92, HB63, HL66, HB76a, HB76e, HBJ77c, HBJ77e, HA77b, HH80, HO87, Hol66a, Hol66b, Hop95, Hor77a, Hor84a, Hor86a, Hor88a, HF91, Hor93, HSF94, Hor94c, Hor96a, HF97, HF01, HS85, Hor88c, HSL01, HW73, Ibs90, IR12, Jac90, Jai74b, Jam90, Jam02, Jar67a, JK97, JC17, JM14, JS19, JQ97, Kah80, KC86, Kim76, Kim78, Kim79b, Kim80c, Kim81a, Kim81b, Kim98b, Kim10a, Kim10b, Kim11, Kim12, KS16, KKLS18, Kin68, Kir77, Kla66a, Kla66b, Kno92b, Koh66, Köh85, Kol65, Kom01b, Kom10, Kon10, Kos19a, Kou94, KsS73, KS88, Lan11, LK09, LB13, LBF14, Lar20, Lás82, LU05, Lee01a]. Sequences [Lee01b, Lee02b, LKC03, Leh75, Lev68a, Lev68b, Lew91, LSS76, Li99b, Lin90, LM79, Luc02, Luc73, LY99, LS18b, Ma98, Mah91, MH87a, Mak12, MS90, McD91, McD94b, McD95, McG09, McG10, ML84, McN87, McN88, McN92, MS94, MS95e, Mel97, Mel99c, Mel99d, Mel00c, Mel01, Mel02, MZ92, MR85, Mon88, Mon90a, Mor75, Mor10, MRS99, MRS00, MR99, MR00, MKL^{+}23, Nag84, NS90, NC85, NW79, NW83, NKW23, Nor05, Nyb09, NR15, NR17, PP14, PT18, Par77c, Par77b, Pau23, PS02, PH75, Pet84, Pet85, Pet88, Phi84, Pih92, Pla94, Pla95, Pla96a, Pla97a, Pop77, PR01, Rap74, Raw87, RFLT20, Rit23, Rob82a, RM11, Rug63, RS11b, ST18, Sal75, San16, SBM97, Sbu02, Sch00, Sch84, Sha74b, Sha78b, Sha78c, Sha79a, Sha83, SM93, SO02, SC01, She68]. Sequences [Shi95, Sli98, SH76, Sol76, Som77, Som82, SK13b, SK15c, SK20, SCJ93, SC97, St.84, Sta03, Ste79, Sti98, Sto67, Sto77, Sun01a, Sun01b, Sza17, Szi17, TR01, TMR02a, Tay67, TZ96, Tri74c, Trz94, Tsu21, Tue21, Tur88a, Tur89b, TS00, UP83, Van12, Van96, Wad74, WR01, Wal85a, Wan95, Wan05, Whi78e, Wil86, Wil75b, Wlo68d, Yal73, Yal77, Yam19, You94, You95b, Yua02, Zar97, Zar98, Zei77, Zei65, Zei70, Zei71a, Zei72, Zha97b, Zha97c, Zha98b, ZL98, ZJ98, Zha99a, Zha01a, ZW02, ZD02, ZL95, Zho96, Zho03, Zöl93, ZT88, dBL83, BB19, BHK07, BL13d, BL13c, BHL21, BL86, BZ07, BM07a, BV05b, CHL19, Cha06b, CHKS05, CDM^{+}06, CT18, Cox87, CM10, Dek20b, Di 13, DÖ03, Djo05a, DMR^{+}16, DP24, ET99, Flo18, GWZ07, Gol06, HS12, HS18]. sequences [HS22, HLS79, JRdlR08, JRdlR10, Jar66, Jar06, JH05, KM18, KA03, KY03, LS06, MM06, MT85, MMYA10, MP85, Mor95, Mus93, ÖAD03b, Özk03, Pet12, Phi07a, SI05, Sha88b, Sob17, Som06a, Sun05, Sun06b, TR03, Ten09, TF93, Vis00, Vou95, WCL99, Wan00, Yab07, YZ18, Yay11, Bro66, Bro67b]. Sequential [Lee01a, OW64]. Serial [Mag81]. Series [Alf63a, AJ90, AJ91c, AJ91f, AJ98, Ano63b, Ano95b, Ard04, Ark69, Ark73a, AAB^{+}90, Bac81, BLL71, BH79, BL65, Bra71, Bro69b, Bro69h, BDM12, BG76, Bru77c, Buc64a, Buc64c, Buc64b, Bus70, CK10, Cat74a, Cer77, CK12, CS81b, CH83, Coh99, CM89, Fil00, Goo71, Goo74a, Goo74b, Goo68b, Gou77b, Gou78b, Gou90, GQ11, Gri14b, HS15, HPS14, Hau93a, HMT23, Her85, HB76c, HB76d, Hor88a, How01a, LOF14, Lar15a, Lar86, LL87a, Lin84, Lyn73, Mad69, Mam61, Man64, Man67, Man68, Man77, McD94a, MS95f, Mel99b, Mor77, Nav01, NC85, Nyb99b, Nyb01a, Nyb23, Pet94, Pop84, Pop86, Pro15, Puc62, Rea70, RT02, SBM97, Say76, Sil91, SC98, Ste78b, Ste21, Sut78, Tri00a, Wal60, WB80, WH03, Wlo73, Xie02, You14, ZY02, Zha99b, Zha00, Zha01b]. Series [ZW04b, Zha09, Ber96, BSS08, Can51, CHF85, Gou06, Kom09, ÖAD03a, Rai48, Sco18, You19]. Series-Parallel [NC85]. Set [AH95, But78, Car76c, CR10, Hoc74, Kih00, KL87, Mai70, Mun17, Nyb98, Pro81, SCJ93, Ste01, Sto77, Wal90, BK16, Dev99, Süt89, WW06, Yor14]. Sets [BCFS22, BL92, Can03a, CW03, CMX20, Chu23, CM17, DCB93, ESW13, Fie73b, Gua78a, Gua78b, Gua78c, HA76a, Jos79, KK01, Kim91c, LF15, LH74, Mai70, MBR^{+}13, MM75, Moh91, MB89, Nag01, SS91, Smi96, Tan00, Wei79, ZT88, Can03b, Zah82]. setting [Sho06]. Seven [Bro70c, Pri68]. Seventeenth [Coo16]. Seventh [Ano95e, Ano95f, Ano95g, Ano96d, Fre97]. Several [Arn95, MZ07, Sto75, KC63, Sun05]. Sex [Sil91]. Shanks [Bol89, Gou72c, Gou74c, GG85, Gou89]. Shannon [Hen22b, Cas91a, Fil98b]. shared [Mat95]. Sharper [Mar13b]. Shaw [Bic72a]. shell [DRS11b, DRS11a]. shell-like [DRS11b, DRS11a]. Shift [Bro73a, For67b, HS19a, HH65, Tru73]. Shift-Invariant [HS19a]. Shifted [JS19, Wag81]. Shiue [Röd94]. Shoemaker [Erc72]. Shoppe [Bro73c, Mel04b, Bro72h]. Short [All77, BCOR95, BJ87, Geb67, KB64, Rei93b, Wun63, Yam19, Bur92, ESS21, Krá72a, Krá72b]. Shorter [Adl69a]. Showing [FFD91]. Shuffle [PP94]. Shuffled [HSS03]. Shuffling [Geb67]. Shujā [Sha88a]. Shuja [Sha90]. Shuja-ibn [Sha90]. Sibley [Bic73a]. SIC [GS17]. Sided [FOR89]. Sides [Fer97, Gou73, Nak20]. Sidney [Lar86]. siècles [AV03]. Sierpinski [Izo95, LH09, BFF11]. Sieve [Str99b, Sai17]. Sieves [Bus73b]. Sieving [Gou74b, Mad69]. SIGGRAPH [B^{+}96]. Sight [HF73]. Sigler [Hor04]. signature [SS95a, SS95b]. Signed [Man75]. Significance [Gri75]. Significant [Gor12a, GTNP96]. Signs [HCG03]. SIGSAM [Gon89]. Silver [BFT15]. Silverman [HLW94]. Simalarides [Sla00]. Similar [Byr65, Coo14b, GP16, HPL93, JO24, Wil77]. Similarity [Kub04, Lar20, BYD^{+}19]. similarly [GL80]. Simple [Adl78, Bus73b, Du89, Eli21, Fil86a, Gri11a, Kha81, LJ67, LJ70, Lon81b, McC72, MS95c, PD18, Sha75a, SS91, Yab01, Yap70, Joh04, Ste23, Wal10]. Simpler [Fie64, HR96]. Simplices [Kub04]. Simplified [Bro78b, CT75]. Simson [Ber84, FG05, Hor82a, Hor86c, HT86, Mel03a]. Simulation [Fer97, Swa92]. Simultaneous [Bab76, BJ85, DJ67, Hen22b]. Sine [GM78, LOF14, Mel17c, Mel18c, Mel18f, Mel18g, Mel18a, mJ10]. Singer [IEE84]. Singly [Rot73]. Singmaster [HsSW95]. Singular [MP66, Süt89]. sion [KS16]. Sisyphus [Cop18]. Site [Ano98b]. Six [Bro69f, Gar04, Sha79d]. Sixth [Ano93a, Ano93b, Ano93c, Ano93k, Ano94c]. Size [Fie66, Fla82]. Sizes [EJ17, Man77, Moo70, Sca99]. Skew [Goo86a, Goo86b, Wat88, JH15]. Skolem [GGL24]. Slants [Fei64]. Slow [CGS20, Cur68]. Small [BDE^{+}14b, BF23, Col83, Puc02b, Wil96]. Smallest [DeT92, KC91b, Lu70]. Smith [Cos02, McD87a, McD87b, Smi96, TA02, Wal87a]. Smogorzhevskii [Puc62]. Smooth [Sto09a]. Snapshots [BJ14]. Snowbird [SC07, SM08]. snowflake [RRD14]. so-called [Sin85]. Social [Win68]. société [AV03]. Society [Bur92, PB15]. Soddy [Sel83]. Solar [LL84, Rea70]. Solids [WB12]. Solitaire [EJ17]. Solsaa [Ano07a]. Solution [Alm86, Ark73b, Ark73c, Ark74, AHS79, AHS80, BY16, BJ79b, Bro68d, CG67, Coh70, Eme66, Ful77, Gou68, GS74, Lin84, Mal83, MV84, MS79, Owi79, Owi87, Pha84, RT72, Web95b]. Solutions [Alp19b, Ano78a, Arp94, Arp04, Bas63a, Bas63b, Bas63c, Bas63d, BJ79a, CC03, CM01, Cha00, Col88b, DeL81, ES00a, ES00b, ES01a, ES01b, ES01c, ES01d, ES02a, ES02b, ES02c, ES02d, ES03a, ES03b, ES03c, ES03d, ES04a, ES04b, ES04c, ES04d, ES05a, ES05b, ES05c, ES05d, ES06a, ES06b, ES06c, ES06d, ES07a, ES07b, ES07c, ES07d, ES09a, ES09b, ES09c, ES09d, ES10a, ES10b, ES10c, ES10d, ES11a, ES11b, ES11c, ES11d, ES12d, ES12a, ES12b, ES12c, ES13d, ES13a, ES13b, ES13c, ES14d, ES14a, ES14b, ES14c, ES15b, ES15c, ES15d, ES15e, ES16a, ES16b, ES16c, ES16d, FL22, Fro23a, Fro23b, Fro23c, Fro23d, Fro24, Ful81b, Gou75, Gre68, HLW94, Hil64a, Hil64b, Hil64c, Hil64d, Hil65a, Hil65b, Hil65c, Hil65d, Hil66a, Hil66b, Hil66c, Hil66d]. Solutions [Hil67a, Hil67b, Hil67c, Hil67d, Hil68a, Hil68b, Hil68c, Hil68d, Hil69a, Hil69b, Hil69c, Hil69d, Hil70a, Hil70b, Hil70c, Hil70d, Hil71a, Hil71b, Hil71c, Hil71d, Hil72a, Hil72b, Hil72c, Hil72d, Hil73a, Hil73b, Hil73c, Hil73d, Hil74a, Hil74b, Hil74c, Hil74d, Hil75a, Hil75b, Hil75c, Hil75d, Hil76c, Hil76d, Hil76e, Hil76f, Hil77b, Hil77c, Hil77d, Hil77e, Hil78a, Hil78b, Hil78c, Hil78d, Hil79a, Hil79b, Hil79c, Hil79d, Hil80a, Hil80b, Hil80c, Hil80d, Hil81a, Hil81b, Hil81c, Hil81d, Hil82a, Hil82b, Hil82c, Hil82d, Hil83a, Hil83b, HP83, HPW83, HPW84a, HPW84b, HPW84c, HP84, HPW85a, HPW85b, HPW85c, HPW85d, HPW86a, HPW86b, HPW86c, HPW86d, HPW87a, HPW87b, HPW87c, Hil87b, Hil88a, Hil88b, Hil88c, Hil88d, Hil89a, Hil89b, Hil89c, Hil89d]. Solutions [Hil90a, Hil90b, Hil90c, Hil90d, Hil91a, Hil91b, Hog63a, Hog63b, Hog63c, Hog63d, Hog64a, Hog64b, Hog64c, Hog64d, Hog65a, Hog65b, Hog65c, Hog65d, Hog66a, Hog66b, Hog66c, Hog66d, Hog67b, Hog67c, Hog67d, Izo99, Kis88, Kla68a, Kwo17a, Kwo17b, Kwo17c, Kwo17d, Kwo18a, Kwo18b, Kwo18c, Kwo18d, Kwo19a, Kwo19b, Kwo19c, Kwo19d, Kwo20a, Kwo20b, Kwo20c, Kwo20d, Kwo21a, Kwo21b, Kwo21c, Kwo21d, Kwo22d, Kwo22a, Kwo22b, Kwo22c, Kwo23a, Kwo23b, Kwo23c, Kwo23d, Kwo24, Lay66, Luc03a, Luc03b, Luc03c, Luc04a, Luc04b, Luc04c, Luc04d, Luc05a, Luc05b, Luc05c, Luc05d, Luc06a, Luc06b, Luc06c, Luc06d, Luc07a, Luc07b, Luc07c, Luc07d, Luc09a, Luc09b, Luc09c, Luc09d, Luc10a, Luc10b, Luc10c, Luc10d, Luc11a, Luc11b, Luc11c, Luc11d, Luc12a, Luc12b, Luc12c, Luc12d, Luc13d, Luc13a, Luc13b, Luc13c, Luc14d, Luc14a, Luc14b, Luc14c]. Solutions [Luc15a, Luc15b, Luc15c, Luc15d, Luc16a, Luc16b, Luc16c, Luc16d, Luc17d, Luc17a, Luc17b, Luc17c, Luc18a, Luc18b, Luc18c, Luc18d, Luc19a, Luc19b, Luc19c, Luc19d, Luc20a, Luc20b, Luc20c, Luc20d, Luc21a, Luc21b, Luc21c, Luc21d, Luc22a, Luc22b, Luc22c, Mil91, PP20, PH75, Rab91a, Rab91b, Rab92a, Rab92b, Rab92c, Rab92d, Rab93a, Rab93b, Rab93c, Rab93d, Rab94a, Rab94b, Rab94c, Rab94d, Rab95a, Rab95b, Rab95c, Rab95d, Rab96b, Rab96c, Rab96d, Rab96e, Rab97a, Rab97b, Rab97c, Rab97d, Rab98a, Rab98b, Rab98c, Rab98d, Rab99c, Rab99d, Rab99e, Rab99f, Rab00a, Rab00b, Rob80, Sed14, Swe64, Tho63g, TS04, Tri00b, Wal85b, Whi67a, Whi68a, Whi68b, Whi68c, Whi68d, Whi69a, Whi69b, Whi69c, Whi69d, Whi70a, Whi70b, Whi70c, Whi70d, Whi71a, Whi71b]. Solutions [Whi71c, Whi71d, Whi72a, Whi72b, Whi72c, Whi72d, Whi73a, Whi73b, Whi73c, Whi73d, Whi74a, Whi74b, Whi74c, Whi74d, Whi75a, Whi75b, Whi75c, Whi75d, Whi76a, Whi76b, Whi76c, Whi76d, Whi77b, Whi77c, Whi77d, Whi77e, Whi78a, Whi78b, Whi78c, Whi78d, Whi79a, Whi79b, Whi79c, Whi79d, Whi80a, Whi80b, Whi80c, Whi80d, Whi81a, Whi81b, Whi81c, Whi81d, Whi82a, Whi82b, Whi82c, Whi82d, Whi83a, Whi83b, Whi83c, Whi83d, Whi84a, Whi84b, Whi84c, Whi84d, Whi85a, Whi85b, Whi85c, Whi85d, Whi86b, Whi86c, Whi86d, Whi86e, Whi87a, Whi87b, Whi87c, Whi87d, Whi88a, Whi88b, Whi88c, Whi88d, Whi89a, Whi89b, Whi89c, Whi89d, Whi90a, Whi90b, Whi90c, Whi90d, Whi91a, Whi91b, Whi91c, Whi91d, Whi92a, Whi92b, Whi92c, Whi92d, Whi93a, Whi93b, Whi93c, Whi93d, Whi94a, Whi94b]. Solutions [Whi94c, Whi94d, Whi95a, Whi95b, Whi95c, Whi95d, Whi96a, Whi96b, Whi96c, Whi96d, Whi97a, Whi97b, Whi97c, Whi97d, Whi98a, Whi98b, Whi98c, Whi98d, Whi99a, Whi99b, Whi99c, Whi99d, Whi00a, Whi00b, Whi00c, Whi00d, Whi01a, Whi01b, Whi01c, Whi01d, Whi02a, Whi02b, Whi02c, Whi02d, Whi03, Wul75, KKLL21, TMR06]. solvability [NS01]. Solve [Liu92]. Solved [Hor78c]. Solving [Ano16a, BCM10, BB19, EMR02, GGL24, Kla73a, KYS13, TMR02b, Tan98, WH70, Wol98, HA66, KSSS16, MH16, Alf66a]. Some [Abe18, Ade17, Ade18, AS73, AS84, Alf64i, All17, And78a, AP92, And69, Ant85b, Bal94, BVB95, BTD22, BH80b, Bis84, BBL95, BPGC97, BB90, BG16, Bro64b, Bro78c, BF91, Bru73, Bru79, Bru96, BM07b, Buc67, Bus73b, CK10, Car66a, Car66b, Car66c, CF70, Car70b, Car71b, CHS72, CSV73, Car75b, Car75c, Car76d, Car76e, Car77c, Car78f, Car78g, Car78h, Car80e, Car80f, Car81b, CH94, Cas89, Cha84b, CG03, Chu90, CS81a, CS85, CS79, CS81c, CW85, Col88b, Con76, CBH10, CB14, CK16, Coo19, CPH70, Cre86, Cre70, CSS07, DH17, Dav76, DM10, Djo96, Djo01b, Djo05b, Edg86, Edg16, Eme66, EB17, Erc73, Erd81, ES10e, Fah16, Fen11b, Fie73b, Fil95, FM95, Fro18, GS81, Geo89, GK76, Gou72d, Gou74b, Gri12, Gri14b, Gry03]. Some [Hal67, HL09, HML01, Hir86, Hod92, HB64b, Hog71a, Hog71b, HP72, HC72, Hol66a, Hol66b, HS84, HC11, HM72, Ibs90, Iye69c, Jai74b, JK98, Jen93, KI16, KK95, KK11, Kim79b, Kis14, Kla65, KR10, Köh85, Kom14, KG18b, Kos19d, Kuh12, Lar15a, Lás82, Lay75, LL87b, LL87c, LL87d, Lee97, Lee01a, Lee01b, Lew80, Lit20, LL05, Liu07, Lon81d, Low71a, Mad72, MT12, MT02, MH75a, MS94, MS95b, MS95c, MS95f, MS95g, Mel99g, Mel99e, Mel03b, Mel09, MD87, Mül06, Mur82b, Muw81, Nas76, NW79, NW83, Nor22, OS98, OS21, PTW13, PT17, PP20, PW93, Per16, Pet84, Pet88, Pla95, Pla97a, Pon19, Pro94, Pro16a, Puc62, Rao53, Rao55, Rap70, Rob82a, Rob91b, Rob02d, RW21, Ros83, Rug63, Sha82, Sha84a, Sha23, SH72, SHC74, Sha74e]. Some [Sha80, Sha84b, Sha60, SW10, Shi20, SR83, Sin70, Sin73b, Sit70, STL22, SSS24, Sta76, Sta80b, Ste86a, Ste21, Str73, Swa00c, TMR02b, Tau89, Ter12, Tos78, Tru73, Tue23a, Tue23b, Tur03, Vau76, Vol13, Wad78, Wad92a, Wal63, Wal79, WH71, WH74b, WH74a, Was99, Wel67, Wen02, WS91, Wiȩ80, Wil77, Wlo13, Wlo68d, WSP04, Wul81, XPP87, YZ18, YZ02, Zha97a, Zha97b, Zha97c, ZL98, Zha98a, ZJ98, Zha04, ZW04a, Zha00, ZW01a, ZW01b, ZW01c, ZW03, ZW04c, ZW06, dF11, dLM94, BV05a, CM10, FPC08, Gou67a, Gri21, HLS79, KM18, Man05b, Men89, Ran95, Wil87, ZZ06, dF14]. Somer [Sel84]. Sometimes [Kil76]. Sominskii [Puc62]. Somme [Zec72b]. Sommerville [HZ17]. Somos [Kai22]. Somos-5 [Kai22]. Sophie [Nyb12]. Sorting [Chu85, Lyn70]. Source [BB65]. Sources [Høy09, Spi00a, Spi00b]. Space [HM83, KC91b, McQ76, SCY^{+}23, EB17]. Spaces [CE66, Eve75, Hen95, DH17, KI16, Ogu20, TD16, TD19, Zah82]. Spacing [vR89]. Spain [Ano99d]. Spanning [BBP^{+}85, FKM^{+}14, Hil74g, Mye75, Sla77]. Sparse [Ima98, Mat95]. Speakers [Ano22c]. SPEC [KS02a]. Special [Alf66b, Arp94, Arp04, Bal19b, Bal19a, Bun75c, CS03, CSH72g, DB70, Fie68a, Fie68b, Fil97d, Ful81c, Hog71b, HB73b, HB75c, Hor67, Hun64a, JM14, Kla66b, KH72, Pow77, SH79a, Som07, Som09]. Specially [Hau88]. Species [MC76]. Specification [Car78d]. Spectra [CK97, LaG13]. Spectral [HH80, Rya75, İpe11, ZJ14]. Spectrum [Alp18, Nyb05, Süt89]. speed [Wat86]. Spheres [Car98, Fer73]. Spherical [KISS15, MBR^{+}13, MBBB21]. spines [CS92]. Spiral [Gou74b]. Spirals [And70, Dav71, HA76b, THKH23]. Spirit [Eve81, Mel03a]. sporangia [THKH23]. Sports [Ano72f]. Spot [DRT11]. Spot-Based [DRT11]. Spreadsheet [Rol03]. Spreadsheet-aided [Rol03]. Springer [Høy09, Mos08]. Springs [Sub90]. SPRNG [MS00a, MS00b]. Square [Alf64g, BCH12, BBP^{+}85, Bau71, BP99, BAS93, Bic65, BN09, Coh64b, DeL95, EF98, Esw78a, Ewe86b, Ewe03, Fed64, Fin75, Fre68, Gic22, Goo93b, Goo93c, Hal67, Hei79, Her85, HJ68a, KK05, KL91a, Laf71, Liu05, Nor16, Pad94, RT04, Sha64b, SK12, Sub99, TS92, Tri74a, TS00, BPIK07, Esw78b, Nea07, RS98b]. Square-Triangular [Laf71]. Squared [AE22, AE23, Bas63f, EA20b, Fed64, Ton24, EA19]. Squarefree [Gra89a]. Squares [Alf64c, Bic72b, Bro67a, Bro72d, Bro65b, BsS95, CM14, Chi86, CBH10, Ewe86a, GG15, Gup74, Hah75, HH71, HBJ77b, Hog77c, How03, Jar68b, Kar70, Kar72, Kar74, Kos23b, Kos23a, Kos23f, Kos23i, Kos23k, Kos23j, Kos23l, Kos23m, Kos23n, Kos24a, Kos24b, KG24, Kos24c, Lon73, Luc02, aMS22a, McD11, Mon88, Moo73, Nyb02, O’C72, OS98, OS21, Pad97, Pri18, PS19, Rob83a, Rob87, Rot73, Rui17, Rus81, Sch19, Sch20, SW10, Sun79, Tri74a, Ula23, Van86, Wun63, Yab02, FS87, FS99, FS14, IMS97, Rob06a, Rob07, Ano88c, Ano88a, Pic79]. Squaring [BHM92, CM07]. Squence [HBJ80]. St [BHP92, BPH93]. Stability [CJ96, SC04]. Staggered [SH91]. Staircase [GWT15, Met97]. Standard [Fer73, PP14, dL97, Séé98]. Stanley [Ano00d]. Star [Bea66b, Bud80, Kor94, LSA07, LK10, Ste01, HS75, LA92]. Star-Like [Ste01]. Statements [Coo14b]. States [Mag81, McI82]. Static [Nod90]. statistical [IA11]. Statistics [Bro69c, KC91b, Tha19]. Staudt [Kun90]. Steinhaus [BH14b]. Step [Kar69, ÖAD03a, Özk03, Sch19, KA03]. Stephen [Mos08]. Stern [BJ03, GG79a, GG79b, GG79c, Nor14]. Steven [Ano96b]. Stick [Jon96]. Stirling [Ben02, Boy09, Bra96, Bra06, CP84, Cak98, Car80c, Car80d, Cha84b, CHT06, Dav14, ED94, Fie68a, Fra67, Gau04, GL01, Gou67a, Gou78a, HMM71, Hir14b, How80, How84, How90, Hsu87, Hsu93b, KK95, KY20, Kyr93, Len94b, Liu09, Par77c, Par77b, PRW93, Sit70, Tos78, You14, Yu98, ZD04]. STOC’12 [ACM12]. Stochastic [BC77, Tor19]. Stolarsky [Gbu81, Hen78a, Kim91b]. Stopping [And00, CZ81, LaD18, Pel72]. Storage [McI82]. Storia [Cio18]. Storing [Wei73]. Straightedges [Gla70]. Strange [Acz11, Gou72a]. Strategies [MSY22]. Strategy [CDH^{+}21, Hen88]. Stream [SCY^{+}23]. Streams [PB15]. Street [Bri64a]. Strengthened [Jar64]. Strict [BLT12]. Striking [Low71a]. String [JR98, IMS97]. Strings [AF87, BW97, Cap89, Chu21a, Mos93, Wei73, BV05a, BH05, BP07]. Strip [Alf65e]. Stroeker [Mak88]. Strong [Arn95, DF92, FH20, HS85, Hor88c, Hor92c, Kim78, Kim79b, Nor05, ST18]. Strongly [Pad97]. Structural [Min81a]. Structure [BPGC97, BK96, JO24, Man17, Sha11, SCJ93, Sun75b, Zar09, BPIK07, MT85, Ryt06, TLC93]. Structures [Boy98, Fer73, TB91, Hol06]. Studied [Mel16e]. Studies [Høy09, Mos08]. Study [Bro96, Har98, Hei79, SH79c, Win68, Hun70b, Jag21]. Stufe [Sin74]. Sturmian [Ber96, ES15a, See91, dLM94, dL97]. sturmiennes [Ber96]. Subcomplete [LP03a]. Subgraph [NC85]. Subgraphs [ESS20]. Subgroups [BN03, BN19]. Subject [CR72, Fie66]. Submission [Ano97f]. Subprime [CVZ17]. Subscript [Fie73b, Fil96]. Subscripted [PP00, Hor94d]. Subscription [Ano97h, Cat09]. Subscripts [AJ91a, Bac81, CL12, Des71, EBJ00, Eng01, Fil93, HB76d, HF95, Hor98, Jar63a, KG19, Lin67b, Sto75]. Subsemigroups [Hig72]. Subsequences [Bru96, CL12, Hal65c, LU05, OP14, PTW13]. Subsequent [Gou77a]. Subsets [Chu21c, CR10, Gri17b, KL91b, KL93, Kon95, Tan75]. Substitute [HMM71]. Substitutions [HL68]. Subtract [You95a, TLC93]. subtract-with-borrow [TLC93]. Subtractive [Cha01, Kno92a]. subtree [Boy97]. Subword [Sal12, Ryt06]. Subwords [Pro95, Wal10]. Success [Cha91, PM82, PM85]. successions [Mun07]. Successive [Sch82]. Such [Gou71a, Hei79, Mon83, Sto77]. suffix [Ryt06]. Suffixes [CCC00]. Suggested [Blu72, Bru73, Car78e, Mil75a]. Sui [Pic81]. suite [CHF85]. suites [Ber96]. Sum [BJF01, Boy09, Bro67a, Cha12, CM14, Dez77, DG98, FG23, Fil86b, FKM^{+}14, Gou71a, Gou78a, Gou99, Gri92, Gri15b, HW12, Hau96, aHZ02, How03, Jar68b, Kil76, Kla66b, Lee10, LAL16, Mey05, Mon90b, Nyb12, ON09, Oze09, Pla97b, Pri18, Ric93, Rie98, Rob87, Sch19, Shu20, Spi06, Ste71, Tue02, Tue23b, Weg75, Wiȩ80, ZW02, CM15, Ewe07b, KA13, LX21, Pro09, RL18, Zec72b]. Sum-of-Digits-Function [DG98]. sum-of-divisors [Ewe07b]. Summand [Mel16a]. Summands [BDE^{+}14b, BCC^{+}20, Gou64b, KKMW11, Sho06]. Summary [Bro74b]. Summation [AJ90, AJ91f, AJ97, Bra71, Bro67f, Bro69h, Car71b, Car72e, Cer77, CB14, Cul76, DB70, Fau64, Fau65, Fil97c, Fil97d, Hog77c, Hor88a, Hor03a, Hsu93b, MH85, MH86, Mea65, MS95d, MS95g, Mel00c, Mel01, Mel02, OS20, Pet84, Pop86, Rab99b, Rus81, SO21, Sto70, Tau65, Tue22, WB80, WH03, Xie02, Zei64, Zei67, Zha01b, ZD04, ZW04b, CV86, Rao55, RL89, Rei54]. Summations [Alf63i, Bro74b, Car71a, Fre73, GK76, Han78, Hor00b, MS95f, Mel16b, Pon68, Sil63, Tre16a, Blo53a]. summer [BB79]. Summing [Alf65e, HB76c, Sil91]. Sums [Ade18, ASD15, Bal13, Bas64, Bea03, BNOS03, BH75, Ber75, BJ81, BJ93, BJD95, Bro69i, Bro94, Bus73a, Cam23, CK24, CH69, Car76d, Car76e, Car77c, Car78f, Car80f, Car84, Cav68, CL11, Chu21b, CS95, CK16, CKR98, CCD^{+}22, DP16, Day69b, DM10, DB66, DX22, DP78a, Dul20, Els05, Er83a, Er84b, EG72, Ewe86a, Ewe88, Ewe92, Ewe01, Fer65, FA89, FB94, Gau98, GB06, Gau12, GP83, GL01, Gic22, Goo94b, Gre77d, Gre77b, Gri11b, Gri74, Guy82, Hau93a, Hen14, Hir86, HB69, HB74b, HP75a, HA76a, Hor79b, Hor94c, How96, How98, HSL01, Ivi69, Iye69d, Jen94, Kal90, Kar74, KC91a, KC93, KK14, KI10, KPAO11, Kim79a, Kim83, Kim10b, KMT22, Kos19d, Kos20, KG20, Kos21a, Kos21b, Kos21c, Kos21d, Kos21e, Kos21f, Kos21g, Kos21h, Kos21i]. Sums [Kos22e, Kos22c, Kos22d, Kos22f, Kos22g, Kos22a, Kos22b, Kos22j, Kos23b, Kos23c, Kos23d, Kos23e, Kos23a, Kos23f, Kos23g, Kos23h, Kos23i, Kos23k, Kos23j, Kos23l, Kos23m, Kos23o, Kos23n, Kos23p, Kos23q, Kos24a, Kos24b, KG24, Kos24c, Lan90, LK09, Lar15a, Lar16, LO18, Lay77, Led67, Len17, Lin65, Man67, MT12, aMS22a, McI20, MO10, MS95e, Mel99a, Mel99b, Mel99f, Mel00a, Mel00d, Mel03b, Mel04a, Mel09, Mel15, Mel16a, Mel16g, Mel17a, Mel17b, Mel17c, Mel17d, Mel18d, Mel18b, Mel18c, Mel18e, Mel18f, Mel18g, Mel18a, MS92, O’C72, Oze05, PW23, Pop85, Pro16a, PS19, Pro22, Rie93, Rob83a, Rob04a, Roc81, Rui17, RP71, Rus82, San23, Say76, Sch20, SH91, Sha65, SD18, SP77, Sof12, Sta03, SSS24, Tep74, Tep76, Ter96b, Tri00a, Ula23, WD76, WY00, Wes66]. Sums [WP07, Yoo21, You17, Zei77, ZY02, ZW04c, Goo93a, IL07, LP18, Mag09, Rob06a, Rob07, ST06]. sup [Pic83b, Pic83a]. supercomputer [AI92]. Supercomputers [And90]. Supercomputing [ACM96, AM15]. Superimposed [Ark73c, Ark74]. superior [Pet91]. superperfect [Fan09]. Superposition [Ogu20]. Superpseudoprimes [SK15b]. Supplement [BMS88a]. Supplied [Fer97]. supplies [Pet91]. Suppose [Lev88]. sure [JRdlR10]. surfaces [Igu94]. Surprisingly [WB12]. Survey [Hor78c]. Sury [Kwo14]. Suzanne [Smi96]. Switch [You95a, SMSM01]. SX [Pet94]. SX-3 [Pet94]. Sylvester [Akr93, Chu10, LAL16]. Symbiotic [Kim01b]. symbol [JH05]. Symbolic [Blo53b, Gon89, HL68, Jef08, RL89]. Symmetric [Abd96, AW66, Bro75b, fC93, Gou99, Gri09, HK91, HMM71, LKL02, Mey05, Nyb03, OW64, Shi20, SK12, DM08]. Symmetrical [HZ17]. Symmetries [FRU15, Kub04, You14]. Symmetry [Goo94b]. Symposium [ACM12, Gon89, IEE81, IEE84, IEE08, Jef08]. Synthesis [Hor96a]. System [And82, And95b, AC13, BJ95, Bur71, CV88, CT75, DB71, Hen22b, Lan92, LL84, McI82, Owi87, RH82, Rea70, Sha05, SK17, SR19, Sun75b, Zöl93, DK17, SS95a, SS95b, TMR06]. Systematic [Bru73, Gou72d, Hag87b]. Systems [Ark73c, AS76, AS81b, Bab76, BS97, DvT83, Fer78, Gra82, HCD97, Kim81a, Kim91d, KM19, Kla73b, Li00a, Mag81, PN77, SCJ93, Swe64, Wiȩ80, Bab41, Wag07].

T [And05, Joh22a]. Tab [Sch94]. Table [Ano88c, Bro72g, Mor77, TJ88, Wlo71d]. Tables [BMS88a, BMS88b]. Tagiuri [Hen17, FG05, Hen18, Hen19]. Take [EF80, Fla82, JS16, Sch70]. Take-Away [EF80, Fla82, Sch70]. Talks [Ano22c]. Tamely [Stă14]. Tan [LM71]. Tangent [Ges81, GK76, MS92]. Tannenbaum [BB85]. Tantalizing [Alm86]. Task [GFJT19]. Task-Based [GFJT19]. Taussig [Ano96a, Rib00]. Tausworthe [TRE73]. Taxman [Hen88]. Taylor [Goo68b]. Tchebichef [Jai74a]. teaching [Ola00]. Teaser [Alv69]. Technical [Geb67, KB64, Wun63]. Technique [AW66, Var89]. Techniques [CMP94, HH00, MH86]. teh [But90]. Telegraph [BY16]. Telescope [Bro73b]. Tell [Cul17]. Ten [Ark73b, AS81a, Pet96]. Tend [Wlo71a, SB93]. Tenth [BJ09, Pet86, Ano01a, Ano01b, Ano02a, Ano02e]. Term [CJS94, Gou71a, Gri14a, Jac90, Kim78, Lar15a, LF17, Par64, Vol13]. Terminal [Dei66a, FM95, Her73, Hor88d, Lin67a]. Terminating [Sha93, Wal90]. Terms [BDM12, Byr65, Cha09, Dez77, Dra64, Dra69, FBJ01, Fil98c, Fil00, HS15, HBJ79b, JK93, Kim91c, LF19, McC72, McD94a, Mel00c, Mel01, Mil91, Nyb23, Owi79, Rus81, WTvR89, Can51, Cha06a, HS12, Jar06, KŞ13, Mil11]. Ternary [KG17c, Now86, Wal90]. Terquem [CG67]. Test [Bru93, Jor64, KB64, Kra70, Ros88, TCC^{+}04, BB04, PW95]. Testing [Bri98, Mor75, CMP94, LM93, Wil98a, Yan95]. Tests [Kap45, MZ93, MT02, RW21, Vos98, Sun06c, Wil87]. Tetranacci [JD22, McL79, PS19, Sch19, Wad92a, Wad92b]. Texts [Kle07, Fol04]. th [AP92, Ark73a, AAB^{+}90, Buc64a, Buc64c, Buc64b, DeL84, Des86, DB66, DB67, Ful81c, Gau04, Gla78, Hen94, HW81, KK01, KK89, Lev85, LS18b, Sha74d, Som87, Som89, Som07, Ste78a, Sun78, Wan95, Wol99, Zho03, ZH03, BB69, Djo05a, Kla66a, Lee87, PM82, Ste78b, Tur79, TS89b, Zho96]. th-Order [AP92, Ful81c, Som87, Som89, Zho03, ZH03, Som07, Lee87, TS89b, Zho96, Djo05a]. th-Power [Ste78a]. Thanks [Ano16d]. Thee [Eve81]. Their [Alp18, Alp19b, AH97, AJ91a, Ano89d, Ano89a, Ano89b, Ano89c, Ano90b, Ano90d, Ano90c, Ano93f, Ano93g, Ano93k, Ano94c, Ano97a, Ano99a, Ano99b, Ano01a, Ano01b, Ano02a, Ano02e, Ano03b, Ano03a, Ano09a, Bai83, Bau71, BPH93, BMRS01, BJ18, Bro64a, CK97, Chu71, Coo10, Coo16, De 81a, Dei66a, DQ83, FG23, Fie67a, Fil98c, FT84, FT87, Fre87, Fre88, Fre90, Ger78, Gir81, GM20, GZ15, GZ19, Her19, Hor83, HCD97, Joh09, Joh12b, Joh20, Joh22b, Kos16, PVB98, Pla96a, Pon17a, RS02, Ris05, Rus81, Swa99a, Swa00c, ZT88, dBF85, AB00, BHP92, Bra06, LS14b, Phi04]. Them [Mel16b, Mel16g, Pet84]. Theme [Hor92d, MSPF95]. Theodorus [Hor91]. Theorem [AC72, Ark66, Ark70b, AH75b, BAS93, BCD^{+}20, BB90, Bow88, Bro64b, Bro67g, Bro78b, Car84, CFH^{+}14, Cav68, Eve75, Ewe03, Gal72, GKL^{+}21, Gou72c, Hau90, Hil77a, Hil74e, Hin80b, Hog72b, HB73c, Hol94, Hor94a, Kel72, KC91a, Kun90, Law83, LB91, LAL16, MMMS22, MMM^{+}22, McN92, Meš13, Mig77, Moo89, Nyb99a, Pih88, Pih91, Pow77, Put76, Rob98, Rob02a, Rob02d, San89, SH71a, Sha74c, Som02, SK22b, Spe77, Sun73, SH70, TJ92, TS89b, Veg69, Wil76, WTvR89, Yab01, vRWT90, Arn97, Chr12, Len06a, Woo07, CS81a, DS19, Gil89, HS75, Kor94, LA92, LK10, Zei74, LSA07]. Theorems [AS00, Bru89, Car76b, FJL^{+}20, Fil98b, Gou74c, Gou95, HC72, Jar46, LM19a, O’D79b, Swa98, TR90, BM07b]. Theoret [FS14]. Theoretic [AH15, DP78b, HH84, Kos15c, Kos19c, Kos21e, Kos21f, Kos23d, ST90, Wil00, BV05a, Kos14a, Kos23l]. Theoretical [BHS06, Sun73]. Theory [ACM12, BL13d, Can03a, Can03b, CLRS86, Dav77, EB88, Goo86a, Goo86b, Gou67b, Gui77b, Hig79, Hoc79, Kno81, Lin67c, PW95, Rya75, San90b, Sha86, Sun73, Zho03, BF90, Bur92, EBK89, Han12, Kah06, Øre67, Pet91, SYY04, Sub07, Vaj89, Vaj08, Bro69a]. There [AML85, GG15, Pen20]. Thermodynamical [Gal79, Gal86]. These [MOP71]. Theta [BDM12]. Thevenin [Ris82]. thick [KT08]. Thin [KT08]. thinker [Teu04]. Thinning [Web95a]. Third [Alf65c, BS11, Bro74a, Car78a, Hor88c, Hor92c, Kil76, LR85, MH90, Rab96a, SH72, Tri74a, Yal73, Gou06, Hir06b, NO15, Ano87b, Ano87c, Ano87d, Ano88d, Ano88e, Fre88]. Third-Order [Bro74a, Hor88c, LR85, MH90, Rab96a, SH72, Tri74a, Yal73, Gou06, Hir06b]. thirteenth [Uli11, Joh09]. Thomas [Gau02, Gau10]. Thoro [Bev70b, Sin80]. Those [CS89, Coo14b]. Thought [Mos08]. Three [Adl68, Adl69b, AMS^{+}22b, Ano02c, Ark73c, Ata89, BAS93, Bro69g, CD87, CY07, Cre88, El 09, Ewe86a, Ewe88, Ewe01, Fie68b, Fra98, HPS14, HBJ79a, Lar15a, LF19, Mel03c, Mul80a, Nor14, Pon19, Sch89, Ula23, Wil70, vRWT90, Rob06a]. Three-Number [Cre88]. Three-Square [BAS93]. Three-Term [Lar15a]. Three-Variable [Mel03c]. threshold [DP24]. Thue [Nyb23]. tic [Coh67]. Ticket [KC91b]. Tickets [Hen95]. Ties [Goo75]. Tight [Sto09a]. Tiles [Hol75, RS98a]. Tiling [AE22, AAB^{+}89a, AAB^{+}90, AB20, BCS19, BJD95, Hen78b, Her81, HA75, Rea80, Wu83]. Tilings [AE23, BQS00, BCSS11, BH17, BHT10, DS20, DeF14, DX22, JD22, MS03, Ric13, Tha19, Tre16b]. Tilling [TL08]. Time [BH69a, Bra90, Cla08, Er83a, Fla82, Hir86, Win65a, Win65b, AAF15, GL80]. Times [And00, CZ81, Cul17, LaD18, Mon83, UP83]. Title [Ano93d, Ano17k]. titled [Kah05, Kah06]. TLB [KS02a]. Toast [Ano98d]. Toned [BCSS11, DS20]. tool [Lün91a, Lün91b]. Tools [Ped71]. Topic [Bro69b]. Topics [Mon85]. Topographs [Nor20]. Topological [Bol77, Bra90, DP78b, Hos73, EBK89, EB17]. Topologies [Bro03, Jos79, MM75]. Topology [LHWS96, Hsu93c]. Torsion [MT17]. Tosses [Hir12b]. Tossing [FW78, Gri11d]. Total [And00, LaD18]. Totally [CR10, Stă14]. Totient [MR95, Sho99a, Sin90]. Totients [Puc02a]. Tour [CD78]. Tournaments [Maš09]. Tower [HS19b, Rit23]. trace [Igu94]. Traces [Gre18]. Track [Mae92a, Mae92b]. Trail [Bro68c]. Transcendental [Knu64, Mig77, Nyb04]. Transfer [SBM97]. Transform [Gal98, Giu71, Gou94, Hau97a, HB75d, Mat73, Pro94, Sta82]. Transformation [HF73, Sun01a, Zha98b, Kum21]. Transformations [Bro78c, Gou90, Tho64a]. Transformed [LM79]. Transforms [CBH17, DP78a, HBJ82c]. Translation [Hor04, Sig02]. Transmission [AF87, Ben93, Cap89]. Transposable [Lud87, Lud88a]. Trapping [D’S89]. trattato [Arr69]. treatise [Arr69]. treatises [Fra03]. Treatment [Bru73, Erc73, Gou72d]. Trebly [AS76]. Trebly-Magic [AS76]. Tree [BBS81, DCB93, EJ17, Fie74, Hor02d, KM14, Bay08, Dev99, PRS96]. tree-contraction [PRS96]. Trees [Abr95, Abr97, Abr00, BKM17, BBP^{+}85, BRGGGS20, BCJ^{+}05, Cap88, Cha86c, GFJT19, Hil74g, Hor82c, Hor83, Hor88d, Hor92d, KR89, KM14, Leg15, MH75b, Mye75, Rei79, SY89, Sla77, Tur88b, TS89b, TR90, You92b, Boy97, Fra84, Kir94, ORSW84, Rou13, Ryt06, Tre96b, Tre96a, Wag07, ZL06]. Triads [FL68, SBM97]. Triangle [And78a, AS00, BE22, BJ81, BFM19, BFMR20, Boi69, Bol89, Bre64b, Bre64a, Bru94a, Bud80, CH94, CFL^{+}22, DW91, DeT92, DK18, Dun67a, Edw09, Fei67, FFD91, FJ12, GS95, Gou65b, Gou73, GG85, Har78, HS66, HB97, Hew77, HB64a, Hog68b, Hog70a, HA73, HB74b, HP75a, HB76a, HBJ78b, Hos76, Hun63, Jon76a, Jos83, Kar96, Kar04b, Kuh12, Kun76, Lee94, Log81, Lon81c, Lon81d, Nak20, Raa63, Rei93a, Rob05a, Sch82, SH02, SP77, SH79c, STL22, Trz94, Wil72, Wil98c, Wlo68d, Wlo71e, WM75, Sof12]. Triangles [Ano93f, Ano93g, Ano93e, BVB95, Bar71, Ber89, BJ93, BH69b, BH73b, BH14b, Bol84, Bol86, BB90, Car70c, CY91, EV89, Eri98, FA89, Goo91, Hen20, Hen67, HJ68a, HB69, Hog70c, HB72b, HBK72, HB73b, HA76a, HBJ77d, Hu02, Hun67, HK89, JC17, Kir77, Kub04, Kun76, LHLT95, Naj93, Nyb98, OS98, OS21, PS02, PS21, Sch81, Sin73b, SH79b, SH80b, Trz96, Tur88b, WW75, Wel94, BBB06, DMP07, MMYA10, Rai48, Rai54, Rob06b, Sun05]. Triangular [All76, BW79, Bau71, BP99, CR72, Chu70, DGMS14, DP95, Ewe88, Ewe92, Ewe01, Gre77a, HB74d, KLT18, KG19, Laf71, Lin96, McD96, Min89, Nyb01b, PP19, Rob02d, Ros67, Sha76, Sub99, Tri74c, Wal85c, Wiȩ80, Sub19]. Triangular-Like [PP19]. Tribonacci [Fro18, RT18, AH77, AE23, AHM92, BBB04, BD77, BGW14, Bru84, Edw09, Eli01, Fei63, Flo18, Ger78, GR83, How01b, Hud87, JH15, Kan14, Kla09, KS10a, KS10b, KS10c, KS11, Kuh12, Lin88, Lin90, LOT23, McC81, Pet88, Rob14, Sch20, SDH77, Sha23, Sha77, Šiu11, SŠŠ22a, Spi82, Tue23a, Yal72]. Tribonacci-Like [Šiu11]. Tribute [Ano88f, Ano03d]. trichotomy [BE07]. Trick [Lon85]. tricobsthal [RT18]. Tridiagonal [BH78a, BH78b, CN04, Fer78, Leh75, Chu10, Fen11a]. Trigonometric [Cha12, GR09, Joh40, MS95d, Mel16a, Xie02]. Trigonometrical [MH85, Goo93a]. Trigonometry [Rap70, Trz96, Goo94a]. Trinomial [HB74b, Lef82]. Trios [Kah87]. Triple [AS20, Aus23, Coh70, FHT10, LF17, Mel98, PW93, Tue23b]. Triples [Ark74, BW79, Ber82, BJ79a, BJ79b, Coh72, DeL76, DV89, Hah72, Han72, Hin92, Hor61b, Hor82b, Luc02, LS13, Mon85, RHT18, Ser74, SH73, Sil98, Sin73a, Tay76, Ter12, Tri09, Wat92, Weg75, Weg77, Weg81b, Woo75, ZG14, ST89, Ano90e]. Triplets [DB01, Ste86a, Uma72]. Trisection [Dem08]. trivariate [Gri17a]. Trivial [Bru93, Hir05]. True [Cul17]. Truncated [Coh99, Kou90, Lig79, Zha09, CHKS05]. truncation [DGMS14]. Trying [Bri64a, Fis81]. Tschebyscheff [Hor69]. Tumble [Joh04]. T’ung [LM71]. Tuning [TCC^{+}04]. Tuples [DeL88, Oli96, Sch97, Woo79, You98]. Turán [BCFS22]. Turing [CS12, Han12, Swi04, Teu04]. Turner [Ano90e]. Twelfth [Cio18, Phi07b]. Twenty [HPL71]. Twenty-Four [HPL71]. Twin [CP79, HJ68b, MG02, Zie63]. Twist [CC20]. Two [AFRT23, Bab76, BH80a, BCSS11, Bil14, BDL^{+}16, Bro68b, Car73, CH74, CJS94, CD87, CZ10, Cre87, DS20, DLR16, Dda20, Dek21, Dif12, DJ67, El 09, Ewe80, Ewe86b, Ewe92, FJL^{+}20, Fer97, Fie67b, Fil98b, FOR89, Fla78, GGL21, Gou71a, Gou75, GB15, Hau97b, Hig87, Hir86, HBJ79a, How03, Hun64a, Jac90, Jar46, Jar68b, Jar68a, Kan15, Kil76, KS16, Kla68a, Kos23p, Kos23q, Kou90, Kra90, Kwo07, Lan92, Lar15a, LK10, LOT23, Luc73, Mel16f, Mel16g, Mel16h, Mel18a, Mun17, O’D79b, PP00, Par77a, Pro15, RS02, RFLT20, Rob04a, Sch85, Shu20, Sin73a, Ste73b, Sur98, Tho65, Tri65, Tri71, Van12, You93, Zei71a, dBF85, AB00, Car39, CM15, CLHY12, KLM^{+}19, LP18, NS94, RL15, RL18]. Two- [Ewe86b]. Two-Digit [Tri65, Tri71, You93]. Two-Dimensional [FJL^{+}20, Sch85, AB00]. Two-Line [Car73, CH74]. Two-Number [Cre87]. Two-Parameter [Mel18a]. Two-Pile [Fla78]. Two-Sided [FOR89]. Two-Term [CJS94, Jac90, Lar15a]. Two-Toned [BCSS11, DS20]. Two-Variable [Kra90, Lan92]. two-way [CLHY12]. Twos [AH75a]. Type [Ade14a, Ade14b, Alb81, AJ90, AS81a, BH69a, BG76, CS04, Cha05, Cha09, Chu23, DGNW85, EM15, FH20, Gre77b, HLW94, Hic73, HB77b, Hol94, HF91, Hor97c, Hor98, Hor02a, KK14, Kra90, Len03b, LM17, Lin88, MS03, Mel17a, Meš13, Muw81, Naj93, NR15, O’D79a, Ped71, PH75, PGP85, PM85, PG89, Rob91b, Rug63, RP71, Sha93, SŠŠ22b, SP77, WH04, Wan20, BK16, CR75, Cha06b, DH17, GQ07, HLS79, JH15, KM18, Ma11]. Types [Coh79, JM14, Sch89, Som09].

Ubiquitous [Kla81]. Unary [Moo93b]. unbiased [SR12]. Unbounded [AF87, Cap89]. Unboundedness [Gre09]. Uncommon [Win11]. Uncounted [Smi09]. Uncovering [Pet99]. Underlying [KC86]. Unequal [Abr97, Abr00]. Unexpected [CVZ17, Des94, Hor03a]. Unified [BCS19, Gui77b, PC13]. Uniform [Ant91, BD70, BD72, Cav77b, Cav77a, Dun67b, Jac89, Kui69, L’E94, MM65]. Uniformly [Hen22a]. Unifying [dF14, KM18]. Unimodal [GR93]. Unimodular [Tho64a]. Unions [RM04]. Unique [AAB^{+}89b, Bro69i, Hor94a, Hor94d, Pon19, PW01, Kim07]. Uniqueness [Gal72, Hor00c, Pow96]. Unit [BH73b, BAP73, DMY13, Hen82, KL91b, Rie93, Wul81]. Unitary [BN85, Gra89a, Hag84, Hag87a, Hag87b, LW87, McC85, Naj89, Wal83, Wal87b, Wal88]. Units [Kli81, Moo70]. Unity [Kon95]. Univariate [Kos15c]. Universal [Cai96b, CPH70, BP10a, MZT90, NO15]. Universality [Tem93, Tra88]. University [BHP92, BPH93, IEE81, NS95, Tol17b, Tol17a]. Unlike [Lan90]. unreasonable [Bur92]. Unsolved [Hor78c]. Unusual [Gou78b, SE20b]. Up-Down [CP09]. Upper [Col87b, Hoi22, Jav22, Mar13b, Sha76, SK13b, Sto09a, DP24, TD16, TD19]. Ups [Fif70]. upward [Tre96b, Tre96a]. USA [NS95, ACM96, ACM12, IEE08]. usage [Pry94]. usages [MS15b]. Use [Bot82, HST94, Mac10, Wil83, Ago49]. Used [Low71b]. Useful [Fer76]. Usefulness [KBN10]. Uses [FT84, FT87]. Using [Ark65b, BY16, BJF99, Bun92, Cap89, CLM22, Cre89, Gau04, GKL^{+}21, Gou78a, GBC16, GLMS22, Gro97, HKW15, Har98, HBJ82c, Hsu93b, JQ97, KBN08, Kyr94, Lar16, MRPC95a, MRPC95b, Mea56, MS95f, MS95g, Mor75, Nyb23, Puc62, Rug63, SW10, Tan98, WH03, Zha01a, EB17, Kum21, LLP^{+}18, Mas94b, Ogu20, PG13, SE20a, WJ76, AF87]. uso [Ago49]. Uspenskii [Puc62]. usually [Boy97]. Utah [SM08, SC07].

V [HB64c]. V. [Puc62]. Vajda [Ano96b]. valid [BFS19]. Valuation [Bal19b, Bal19a, MR15, PT18, San16, Sob17]. Valuations [Coh99, GM20]. Value [Ahu69, Jen71, Lew70, Lin96, Sal75]. Valued [BJS96, Nyb99b, Nyb01a, Nyb04, Zah82]. Values [Bea03, BL17b, Cla86, Col83, Fil98c, Fox01, Hor00a, JS16, Lev68b, Lin21, Lor73, Low71b, Mea56, Nyb12, Pon17b, SH79b, SH79c, Spi97, Tsu21, BHL21]. Vanishing [Sha64b]. Variable [Gri11b, Kra90, Lan92, Mel03c]. Variables [Bab76, BGTM^{+}11, Ber68, Hau97b, Kou90, Mul80a, RS02, Sto75, Sur98, KC63, NS94]. Variance [MNPX17, CJ24]. Variant [Gou65b, Hal83, KH82]. Variants [KP13, Rit23, Pro09]. Variation [Jon78, MC76, You93]. variations [BP10a, MSPF95, NO15, HB76c]. Varieties [CS04]. Various [Bru72, Chu21c, HM72, JC17]. Varn [Abr95]. Varol [Liu96]. Vector [And90, Gup78b, Zah82]. vector-valued [Zah82]. Vectors [CMS12, Ful78]. Vegas [NS95]. Verblunsky [DMY13]. Verification [HF73]. Verifying [HT22]. Verner [Bro69a, BJ80]. Version [GR93, HMT23, Cio18]. Vertices [JC17, Ste01]. Very [Hun64a]. VI [HL67b]. Via [AE23, BCM10, BSY05, DD95, JD22, SSS24, You95b, Bal15, BB19, Boy97, BBB06, DP24, Ege83, Fen11a, Gri11b, GG13, Gri14b, How96, KK89, Mar93, MS90, Mel10c, Pet84, Pet12, Rob98, Rui17, Spi06, Tur79, WW17]. Vieta [Osl07, Hor02c, Hor03b, Kos16]. Vieta-like [Osl07]. Viète [GCMP14a, GCMP14b]. Viète-Like [GCMP14b, GCMP14a]. View [BNOS03, Car98, Hor82c, GCMP14a]. Views [Rei04]. VII [Bic70]. VIII [Bic71b]. Vindicated [KTW84]. Vine [Ano99d, Ano03d]. Vinogradov [Hau90]. Virginia [Swa92]. Visits [Woo97]. vistas [HHP02]. Visual [B^{+}96]. Visualizing [BJD95]. Vol [Puc62]. volgarizzazione [Arr69]. Volterra [MH16]. Volume [Ano64d, Ano65f, Ano66c, Ano67b, Ano68c, Ano69d, Ano70g, Ano71d, Ano72h, Ano73g, Ano74d, Ano75b, Ano76d, Ano77d, Ano78b, Ano79c, Ano80c, Ano81, Ano82d, Ano83c, Ano84d, Ano85d, Ano86d, Ano87e, Ano88g, Ano89f, Ano90g, Ano91g, Ano92d, Ano93l, Ano94f, Ano95h, Ano96e, Ano97i, Ano99f, Ano00g, Ano01d, Ano02f, Ano03e, Ano04c, Ano05b, Ano06b, Ano07c, Ano09b, Ano10d, Ano11e, Ano12b, Ano13b, Ano14e, Ano15b, Ano16e, Ano17l, Ano18g, Ano19i, Ano20g, Ano21d, Ano23, Ozb12, Sch94, Ano63d]. Vorob’ev [Puc62]. Vosmansky [GQ10]. Voyce [AJ94a, Dor99, Hor96d, Hor97a, Hor98, Hor99, Hor00a, Hor00c, HF01, Hor02b, Hor03a, Lee02a, Swa66b, Swa68, Swa00a, Swa00b]. vulgarization [Arr69].

W. [Boh79, Mig77, Put76]. Waerden [NR15]. Waiting [PM82, UP83]. Waksman [Veg69]. Walk [JV19]. Walking [ABBB13, MPPW22, Tue02]. Walks [Kos19b, NS05b, Ben06, CGS20, Gri15a, Neu17, Zei96]. Wall [Kla18, Li99a, Li99b]. Wallis [Hir14b]. Wang [CS79]. Ward [Ano63e, Gou69, Lax74]. Waring [Bas64, Fil92, Gou99, Wan20, ZZ06]. was [CHF85]. Watching [Swi04]. Water [Dei72]. Watermarking [FM15]. Watson [Ewe82, Hey81]. Wave [Zuk76]. Wavelength [Bot82]. Way [Her19, CLHY12]. Web [Ano98b]. Weighted [Ade18, BNOS03, CS75b, Car78b, Car80c, Car80d, Cha84b, DP16, DX22, DMRS97, Gau04, How84, MS15a, Mel17b, Mel17c, Mel18d, Mel18e, Mel18f, Mel18g, Mye75, Oze05, Rui17, Sta03, SSS24, Sal12, TD16, TD19]. weights [Kir94]. Weinstein [San89]. Weisner [CS81c]. Weiss [Pro04]. Welcome [Ano16d]. Well [Bol77, CBH17]. Well-Known [Bol77, CBH17]. Were [RT02]. Werkzeug [Lün91a, Lün91b]. West [Giu04]. Western [Lar78]. Whack [Ber11]. Wheel [Def13]. Where [AJ92, BJE95, BAP73, DeL82, DK18, Esw79a, FF89, HW81, Rob83b, Rob84, Rob90, Tao17, Zho99, BH05, Hal65a]. Which [Alf64h, Arn95, AC13, CS89, DLR16, Hir86, KH82, KK01, Kom10, LF69, Luc03d, Mel97, Mel01, Mel16a, Pet96, Pon17a, Rob78a, Rob83a, Rob83c, Som79, SK17, Sta82, Zöl93, Bro78c, Bur71, DeT92, Du89, Fin75, GD09, GQ07, KLM^{+}19, LS06, LS14a, Mel00c, Pon19, Som06a]. who [Dev17, Sim18]. Whose [AH95, AJ91d, DFLT14, DB67, FG23, Fil96, Ful78, JK93, Lon73, LS13, McD94a, Wul76, Zei70, Kla68a]. Wide [Kyr94]. Wiedemann [SR12]. Wieferich [Kla09, MR07, Pen20]. Wieferichs [Hee84]. Wilderness [Acz11]. Wilf [Pau23]. Williams [Mül06]. Wilson [AH75b, Rob98, SH82]. Winding [HA76b]. windows [HHP02]. Winners [Ano16f]. Winning [CDH^{+}21, Hen88, MSY22, BP07]. Winter [Swa92]. Wise [Dun67a, Alt88]. within [EBJ00, HB73b, KS02b, Sho06]. Without [Hir14b, KL91b, KV92, Pri18, Sai17, AG78, BW97, Fis81, GM96, KL93, Kon95, Mos93, Ped71, Smi94]. Wlodarski [SB93]. Wolstenholme [HMT23, MR07]. Word [And20, CCC00, Gri18, RR14, Tur88a, TR90, CH05, CLHY12, Dro95, GSS19, Krá72a, Krá72b, RRD14, Wal10, de 95]. word-length [Krá72a, Krá72b]. Words [Bas20, BPT14, Chu92, fC93, fC95b, fC03, CDW13, CP09, DMSS16, HS19b, Kir22, KR22, KG17c, Pri18, RW81, BKV22, CLHY12, Dek20b, DMR^{+}16, FS99, FS14, HS22, HW15, JMNRS22, Man17, Ryt06, See91, de 81b, dLM94]. work [Cho82, Sha88a, Sha90]. World [Ber11, Wlo63, Wlo68b, Dev17, Sim18]. Worst [CFG19]. Worst-Case [CFG19]. Wright [Mil91]. Wyler [Li99b]. Wythoff [BH80b, BJ85, Fra84, GP16, HH78, HBJS79, Hor78b, Kim95b, Kim11, Sil76, Sil77b, Tur89a].

x [Jon90, Bro72f]. XI [HA73]. XII [AV03, HCB73]. XIII [AV03, Bic73b]. XIV [HB74c]. XV [Arr69, HB76c]. XVI [BH78c]. xvii [Mos08, HBJ78c]. xxxiv [Sch94].

Ye [Bro73c, Mel04b, Bro72h]. Years [Joh12a, Sha94, Hor75]. Yield [Win65a, Win65b]. Yien [KC91a]. Yildirim [Sub07]. Yildiz [Sub07]. York [ACM12, Gal08, Høy09]. young [Hor75, CMM^{+}17].

Zeckendorf [AUFP13, AC11, ABJ11, And14, BSEFM19, Bal13, BBG^{+}13, BDE^{+}14a, BDE^{+}14b, BCD^{+}20, BCC^{+}20, Bro64b, Bru89, BT91, Bun92, Cap88, CFH^{+}14, CCG^{+}19, Chu21a, CLM22, CDH^{+}21, CDH^{+}22, DD95, DG98, Eng01, EA12, Fen03, FH98, FP96, GFMR^{+}24, Ger09, GKL^{+}21, GTNP96, Gri10, Gri14a, Gri15b, Gru07, GCK^{+}19, HS99, HML01, Hog72b, HB73c, Kel72, Kim83, Kim91d, Kim95b, Kim98a, KKMW11, Len06a, LM17, LM19a, LLM^{+}20, MMMS22, MMM^{+}22, MSY22, Shu20, TS89b, TR90, Woo07]. Zero [Car74a, Car77a, CS77, Fos66, Par77c, Par77b, Sed14, BK16, Süt89, Car80a]. Zero-Avoiding [Sed14]. Zero-One [Car74a, Car77a, CS77, Par77c, Par77b, Fos66, Car80a]. Zeroes [Fer76, Geo75]. Zeros [Alf64j, Bac66, BRS04b, Dil91, HSR97, Kim10c, St.84, Sub95, Tee94, WH04, Ric93]. Zeta [GP83, SŠŠ22a, SŠŠ22b, You19]. Zigzag [CW22, HSG63, Hor85b, Hor86b]. Ziv [JMNRS22].

References

Arkin:1989:RST

[AAB^{+}89a] J. Arkin, D. C. Arney, G. E. Bergum, S. A. Burr, and B. J. Porter. Recurring-sequence tiling. Fibonacci Quarterly, 27(4): 323–332, August 1989. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/27-4/arkin2.pdf.

Arkin:1989:UFF

[AAB^{+}89b] J. Arkin, D. C. Arney, G. E. Bergum, S. A. Vurr, and B. J. Porter. Unique Fibonacci formulas. Fibonacci Quarterly, 27(4): 296–302, August 1989. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/27-4/arkin1.pdf.

Arkin:1990:TPP

[AAB^{+}90] J. Arkin, D. C. Arney, G. E. Bergum, S. A. Burr, and B. J. Porter. Tiling the k th power of a power series. Fibonacci Quarterly, 28 (3):266–272, August 1990. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/28-3/arkin.pdf.

Ashrafi:2016:NEP

[AAB^{+}16] Ali Reza Ashrafi, Jernej Azarija, Azam Babai, Khadijeh Fathalikhani, and Sandi Klavzar. The (non-)existence of perfect codes in Fibonacci cubes. Information Processing Letters, 116(5):387– 390, May 2016. CODEN IFPLAT. ISSN 0020-0190 (print),

1872-6119 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0020019016000120.

Adam:2015:GSF

[AAF15] Maria Adam, Nicholas Assimakis, and Alfonso Farina. Golden section, Fibonacci sequence and the time invariant Kalman and Lainiotis filters. Applied Mathematics and Computation, 250 (??):817–831, January 1, 2015. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300314015446.

Atanassov:1985:NPG

[AAS85] Krassimir T. Atanassov, Liliya C. Atanassova, and Dimitar D. Sasselov. A new perspective to the generalization of the Fibonacci sequence. Fibonacci Quarterly, 23(1):21–28, February 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/23-1/atanassov.pdf.

Andrica:1985:RLR

[AB85] Dorin Andrica and Serban Buzeteanu. On the reduction of a linear recurrence of order r. Fibonacci Quarterly, 23(1):81–84, February 1985. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/23-1/andrica.pdf.

Apostolico:2000:FAT

[AB00] Alberto Apostolico and Valentin E. Brimkov. Fibonacci arrays and their two-dimensional repetitions. Theoretical Computer Science, 237(1–2):263–273, April 28, 2000. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.elsevier.nl/gej-ng/10/41/16/171/21/34/
abstract.html;   http://www.elsevier.nl/gej-ng/10/41/
16/171/21/34/article.pdf.

Azose:2020:TIB

[AB20] Jonathan J. Azose and Arthur T. Benjamin. A tiling interpretation of the q-binomial coefficients. Fibonacci Quarterly, 58(2): 99–??, May 2020. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/58-2/azose.pdf.

AragonArtacho:2013:WRN

[ABBB13] Francisco Aragón Artacho, David H. Bailey, Jonathan M. Borwein, and Peter B. Borwein. Walking on real numbers. The Mathematical Intelligencer, 35(1):42–60, March 2013. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

URL http://gigapan.com/gigapans/106803;   http://www.
davidhbailey.com/dhbpapers/tools-walk.pdf.

Abrate:2011:AGF

[ABCM11] Marco Abrate, Stefano Barbero, Umberto Cerruti, and Nadir Murru. Accelerations of generalized Fibonacci sequences. Fibonacci Quarterly, 49(3):255–266, August 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/49-3/abrate.pdf.

Abrate:2012:PRC

[ABCM12] Marco Abrate, Stefano Barbero, Umberto Cerruti, and Nadir Murru. Periodic representations for cubic irrationalities. Fibonacci Quarterly, 50(3):252–??, August 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/50-3/abrate.pdf.

Abderrezzak:1996:MSI

[Abd96] Abdelhamid Abderrezzak. Multivariate symmetric identities. Fibonacci Quarterly, 34(5):386–393, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-5/abderrezzak.pdf.

Aharonov:2005:FCO

[ABD05] Dov Aharonov, Alan Beardon, and Kathy Driver. Fibonacci, Chebyshev, and orthogonal polynomials. American Mathematical Monthly, 112(7):612–630, August/September 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://www.jstor.org/stable/30037546.

Abercrombie:1975:LE

[Abe75] Alexander G. Abercrombie. Letter to the editor. Fibonacci Quarterly, 13(2):171–173, April 1975. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/13-2/letter.
pdf.

Abel:2018:SNI

[Abe18] Ulrich Abel. Some new identities for derangement numbers. Fibonacci Quarterly, 56(4):313–??, November 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/56-4/abel.pdf.

Anderson:2011:MZR

[ABJ11] Peter G. Anderson and Marjorie Bicknell-Johnson. Multidimensional Zeckendorf representations. Fibonacci Quarterly, 49(1): 4–9, February 2011. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/49-1/anderson.pdf.

Angelini:2019:C

[ABN^{+}19] Éric Angelini, Lars Blomberg, Charlie Neder, Rémy Sigrist, and N. J. A. Sloane. “Choix de Bruxelles”: A new operation on positive integers. Fibonacci Quarterly, 57(3):195–??, August 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-3/angelini.pdf.

Abramson:1971:CCO

[Abr71] Morton Abramson. Combinations, compositions and occupancy problems. Fibonacci Quarterly, 9(3):225–236, May 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-3/abramson-a.pdf.

Abramson:1976:RCC

[Abr76] Morton Abramson. Restricted combinations and compositions. Fibonacci Quarterly, 14(5):439–451, December 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-5/abramson.pdf.

Abrahams:1995:VCG

[Abr95] Julia Abrahams. Varn codes and generalized Fibonacci trees. Fibonacci Quarterly, 33(1):21–25, February 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-1/abrahams.pdf.

Abrahams:1997:CSG

[Abr97] Julia Abrahams. Couples sequences of generalized Fibonacci trees and unequal costs coding problems. Fibonacci Quarterly, 35(4): 309–316, November 1997. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/35-4/abrahams.pdf.

Abrahams:2000:NGF

[Abr00] Julia Abrahams. Nonexhaustive generalized Fibonacci trees in unequal costs coding problems. Fibonacci Quarterly, 38(2):127– 135, May 2000. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/38-2/abrahams.pdf.

Abrate:2008:FSQ

[Abr09] Marco Abrate. Fibonacci sequences of quaternions. Fibonacci Quarterly, 46/47(4):356–365, November 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/46_47-4/abrate.pdf.

Anaya:1972:GGI

[AC72] Robert Anaya and Janice Crump. A generalized greatest integer function theorem. Fibonacci Quarterly, 10(2):207–212, February 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/10-2/anaya.pdf.

Atkinson:1987:CNM

[AC87] M. D. Atkinson and H. W. Chang. Computing the number of mergings with constraints. Information Processing Letters, 24 (5):289–292, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anderson:2011:EPK

[AC11] Peter G. Anderson and Curtis Cooper. Every positive K-Bonacci-like sequence eventually agrees with a row of the K-Zeckendorf array. Fibonacci Quarterly, 49(4):303–??, November 2011. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/49-4/anderson.pdf.

Avila:2013:MWL

[AC13] Brandon Avila and Yongyi Chen. On moduli for which the Lucas numbers contain a complete residue system. Fibonacci Quarterly, 51(2):151–??, May 2013. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/51-2/avila.pdf.

Alzate:2020:FIJ

[ACF20] Santiago Alzate, Oscar Correa, and Rigoberto Flórez. Fibonacci identities from Jordan identities. Fibonacci Quarterly, 58(5):2– ??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/58-5/alzate.pdf.

ACM:1996:FCP

[ACM96] ACM, editor. FCRC ’96: Conference proceedings of the 1996 International Conference on Supercomputing: Philadelphia, Pennsylvania, USA, May 25–28, 1996. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

ACM:2012:SPA

[ACM12] ACM, editor. STOC’12: Proceedings of the 2012 ACM International Symposium on Theory of Computing: May 19–22, 2012, New York, NY, USA. ACM Press, New York, NY 10036, USA, 2012. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Andrade:2021:CLM

[ACOM21] Enide Andrade, Dante Carrasco-Olivera, and Cristina Manzaneda. On circulant like matrices properties involving Horadam, Fibonacci, Jacobsthal and Pell numbers. Linear Algebra and its Applications, 617(??):100–120, May 15, 2021. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0024379521000355.

Aczel:2011:SWL

[Acz11] Amir D. Aczel. A Strange Wilderness: the Lives of the Great Mathematicians. Sterling, New York, NY, USA, 2011. ISBN 1-4027-8584-4 (hardback), 1-4027-9085-6 (e-book). xix + 284 pp. LCCN QA21 .A29 2011.

Aydin:1998:GFS

[AD98] Hüseyin Aydin and Ramazan Dikici. General Fibonacci sequences in finite groups. Fibonacci Quarterly, 36(3):216–221, June 1998. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/36-3/aydin.pdf.

Agoh:2010:RRN

[AD10] Takashi Agoh and Karl Dilcher. Recurrence relations for Nörlund numbers and Bernoulli numbers of the second kind. Fibonacci Quarterly, 48(1):4–12, February 2010. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/48-1/
agoh.pdf.

Adegoke:2014:GRF

[Ade14a] Kunle Adegoke. The Golden Ratio, Fibonacci numbers and BBPtype formulas. Fibonacci Quarterly, 52(2):129–138, May 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-2/adegoke.pdf.

Adegoke:2014:NBB

[Ade14b] Kunle Adegoke. A new binary BBP-type formula for \sqrt{5} log ϕ. Fibonacci Quarterly, 52(4):357–359, November 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/52-4/adegoke.pdf;   http://www.fq.math.ca/Papers/
52-4/adegoke4282014.pdf.

Adegoke:2017:SIP

[Ade17] Kunle Adegoke. Some infinite product identities involving Fibonacci and Lucas numbers. Fibonacci Quarterly, 55(4):343–??, November 2017. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/55-4/adegoke.pdf.

Adegoke:2018:WSS

[Ade18] Kunle Adegoke. Weighted sums of some second-order sequences. Fibonacci Quarterly, 56(3):252–??, August 2018. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/56-3/adegoke.pdf.

Adler:1968:TDE

[Adl68] Irving Adler. Three Diophantine equations — Part I. Fibonacci Quarterly, 6(6):360–369, December 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-6/
adler-a.pdf.

Adler:1969:SP

[Adl69a] Irving Adler. A shorter proof. Fibonacci Quarterly, 7(5):538– ??, December 1969. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/7-5/adler-a.pdf.

Adler:1969:TDE

[Adl69b] Irving Adler. Three Diophantine equations — Part II. Fibonacci Quarterly, 7(2):181–193, April 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-2/
adler.pdf.

Adler:1971:SCN

[Adl71] Irving Adler. Sequences with a characteristic number. Fibonacci Quarterly, 9(2):147–162, April 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-2/
adler.pdf.

Adler:1978:SCF

[Adl78] Irving Adler. A simple continued fraction represents a mediant nest of intervals. Fibonacci Quarterly, 16(6):527–529, December 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/16-6/adler.pdf.

Adler:1989:CDE

[Adl89] Irving Adler. Concerning the divisors of N and the exponents they belong to modulo (N − 1) or (N + 1). Fibonacci Quarterly, 27(3):259–266, June 1989. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/27-3/adler.pdf.

Atanassov:2014:PFR

[ADS14] Krassimir T. Atanassov, Daryl R. DeFord, and Anthony G. Shannon. Pulsated Fibonacci recurrences. Fibonacci Quarterly, 52(5): 22–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Papers1/52-5/Atanassov.pdf.

Ahuja:1979:CPR

[AE79] J. C. Ahuja and E. A. Enneking. Concavity property and a recurrence relation for associated Lah numbers. Fibonacci Quarterly, 17(2):158–161, April 1979. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/17-2/ahuja.pdf.

Allen:2022:FTDa

[AE22] Michael A. Allen and Kenneth Edwards. Fence tiling derived identities involving the Metallonacci numbers squared or cubed. Fibonacci Quarterly, 60(5):5–??, December 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-5/allen.pdf.

Allen:2023:IIT

[AE23] Michael A. Allen and Kenneth Edwards. Identities involving the Tribonacci numbers squared via tilings with combs. Fibonacci Quarterly, 61(1):21–??, February 2023. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/61-1/
allen.pdf.

Astola:2004:FAE

[AESS04] J. T. Astola, K. Egiazarian, M. Stanković, and R. S. Stanković. Fibonacci arithmetic expressions. Automation and Remote Control, 65(6):842–856, June 2004. CODEN AURCAT. ISSN 0005-1179 (print), 1608-3032 (electronic).

Apostolico:1987:RTU

[AF87] Alberto Apostolico and A. S. Fraenkel. Robust transmission of unbounded strings Using Fibonacci representations. IEEE Transactions on Information Theory, 33(2):238–245, March 1987. CO-DEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Adedji:2023:FLN

[AFRT23] Kouèssi Norbert Adédji, Alan Filipin, Salah Eddine Rihane, and Alain Togbé. Fibonacci or Lucas numbers that are concatenations of two g-repdigits. Fibonacci Quarterly, 61(1):68–??, February 2023. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/61-1/adedji.pdf.

Adedji:2022:FDP

[AFT22] Kouessi Norbert Adédji, Alan Filipin, and Alain Togbé. On the family of Diophantine pairs P_{2k}, 2P_{2k+2}. Fibonacci Quarterly, 60 (1):25–??, February 2022. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/60-1/adedji.pdf.

Austin:1978:BSI

[AG78] Richard Austin and Richard Guy. Binary sequences without isolated ones. Fibonacci Quarterly, 16(1):84–85, February 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-1/austin.pdf.

Atkins:1987:FNC

[AG87] John Atkins and Robert Geist. Fibonacci numbers and computer algorithms. College Mathematics Journal, 18(4):328–336, September 1987. CODEN ???? ISSN 0746-8342 (print), 1931-

1346 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/07468342.1987.11973055.

Agarwal:1987:NRL

[Aga87] A. K. Agarwal. A note on n(x, y)-reflected lattice paths. Fibonacci Quarterly, 25(4):317–319, November 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-4/agarwal.pdf.

Agarwal:1989:PRS

[Aga89] A. K. Agarwal. Properties of a recurring sequences. Fibonacci Quarterly, 27(2):169–175, May 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-2/
agarwal.pdf.

Agarwal:1990:NKN

[Aga90] A. K. Agarwal. On a new kind of numbers. Fibonacci Quarterly, 28(3):194–199, August 1990. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/28-3/agarwal.pdf.

Agarwal:1991:CIA

[Aga91] A. K. Agarwal. Combinatorial interpretations of the q-analogues of L_{2n+1}. Fibonacci Quarterly, 29(2):137–140, May 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-2/agarwal.pdf.

Ardal:2008:RRI

[AGJ^{+}09] Hayri Ardal, David S. Gunderson, Veselin Jungic, Bruce M. Landman, and Kevin Williamson. Ramsey results involving the Fibonacci numbers. Fibonacci Quarterly, 46/47(1):10–17, February 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/46_47-1/landman.pdf.

Agostini:1949:LDL

[Ago49] A. Agostini. L’uso delle lettere nel Liber abaci di Leonardo Fibonacci. (Italian) [The use of the letters in Liber abaci Leonardo Fibonacci]. Boll. Un. Mat. Ital. (3), 4(??):282–287, ???? 1949.

Agostini:1953:LF

[Ago53] Amedeo Agostini. Leonardo Fibonacci. Archimede. Rivista per gli Insegnanti e i Cultori di Matematiche Pure e Applicate, 5(??): 205–206, ???? 1953. ISSN 0390-5543.

Albano:1995:FPL

[AGO95] Antonio Albano, Giorgio Ghelli, and Renzo Orsini. Fibonacci: a programming language for object databases. VLDB Journal: Very Large Data Bases, 4(3):403–444, July 1995. CODEN VLDBFR. ISSN 1066-8888 (print), 0949-877X

(electronic). URL http://ftp.informatik.rwth-aachen.de/
dblp/db/indices/a-tree/a/Albano:Antonio.html;   http:
//ftp.informatik.rwth-aachen.de/dblp/db/indices/a-tree/
g/Ghelli:Giorgio.html;   http://ftp.informatik.rwth-aachen.
de/dblp/db/indices/a-tree/o/Orsini:Renzo.html;   https:
//www.math.utah.edu/pub/tex/bib/fibquart.bib. Electronic

edition.

Arkin:1970:EFN

[AH70] Joseph Arkin and Verner E. Hoggatt, Jr. An extension of Fibonacci numbers — II. Fibonacci Quarterly, 8(2):199–216, March 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-2/arkin.pdf.

Alladi:1975:COT

[AH75a] Krishnaswami Alladi and V. E. Hoggatt, Jr. Compositions with ones and twos. Fibonacci Quarterly, 13(3):233–239, October 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/13-3/alladi1.pdf.

Arkin:1975:GFN

[AH75b] Joseph Arkin and V. E. Hoggatt, Jr. The generalized Fibonacci number and its relation to Wilson’s theorem. Fibonacci Quarterly, 13(2):107–109, April 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-2/
arkin.pdf.

Alladi:1977:TNR

[AH77] Krishnaswami Alladi and V. E. Hoggatt, Jr. On Tribonacci numbers and related functions. Fibonacci Quarterly, 15(1):42–45, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/15-1/alladi.pdf.

Alspach:1981:PMC

[AH81] Brian Alspach and Katherine Heinrich. Perfect magic cubes of order 4^{m}. Fibonacci Quarterly, 19(2):97–105, April 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-2/alspach.pdf.

Ando:1995:DCS

[AH95] Shiro Ando and Teluhiko Hilano. A disjoint covering of the set of natural numbers consisting of sequences defined by a recurrence whose characteristic equation has a Pisot number root. Fibonacci Quarterly, 33(4):363–367, August 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-4/ando.pdf.

Ando:1997:CNE

[AH97] Shiro Ando and Masahumi Hayashi. Counting the number of equivalence lasses of (m, F) sequences and their generalizations. Fibonacci Quarterly, 35(1):3–8, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-1/ando.pdf.

Alexander:2015:GTE

[AH15] James Alexander and Paul Hearding. A graph-theoretic encoding of Lucas sequences. Fibonacci Quarterly, 53(3):237–??, August 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/53-3/alexander.pdf.

Alm:2016:NPF

[AH16] Jeremy F. Alm and Taylor Herald. A note on prime Fibonacci sequences. Fibonacci Quarterly, 54(1):55–??, February 2016. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/54-1/alm.pdf.

Atanassov:1992:RFG

[AHM92] K. Atanassov, J. Hlebarska, and S. Mihov. Recurrent formulas of the generalized Fibonacci and Tribonacci sequences. Fibonacci Quarterly, 30(1):77–79, February 1992. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/30-
1/atanassov.pdf.

Arkin:1979:ESP

[AHS79] Joseph Arkin, V. E. Hoggatt, Jr., and E. G. Strasus. On Euler’s solution of a problem of Diophantus. Fibonacci Quarterly, 17(4): 333–339, December 1979. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/17-4/arkin.pdf.

Arkin:1980:ESP

[AHS80] Joseph Arkin, V. E. Hoggatt, Jr., and E. G. Straus. On Euler’s solution to a problem of Diophantus — II. Fibonacci Quarterly, 18(2):170–176, April 1980. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/18-2/arkin.pdf.

Aiello:1987:EMN

[AHS87] W. Aiello, G. E. Hardy, and M. V. Subbarao. On the existence of e-multiperfect numbers. Fibonacci Quarterly, 25(1):65–71, February 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/25-1/aiello.pdf.

Ahuja:1969:OED

[Ahu69] J. C. Ahuja. Orthogonal expansion derived from the extreme value distribution. Fibonacci Quarterly, 7(5):488–493, December 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/7-5/ahuja-a.pdf.

He:2002:MSG

[aHZ02] Ping an He and Zhizheng Zhang. The multiple sum on the generalized Lucas sequences. Fibonacci Quarterly, 40(2):124–127, May 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/40-2/he.pdf.

Arno:1992:PSC

[AI92] Steven Arno and Ken Iobst. The PETASYS supercomputer and a class of pseudo-random number generators. Technical report SRC-TR-92-069, Supercomputing Research Center: IDA, Lanham, MD, USA, April 1992. 29 pp.

Ainsworth:1977:GF

[Ain77] O. R. Ainsworth. On generating functions. Fibonacci Quarterly, 15(2):161–163, April 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-2/ainsworth.pdf.

Andre-Jeannin:1990:LSS

[AJ90] R. André-Jeannin. Lambert series and the summation of reciprocals in certain Fibonacci–Lucas-type sequences. Fibonacci Quarterly, 28(3):223–226, August 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
3/andre-jeannin.pdf.

Andre-Jeannin:1991:DGF

[AJ91a] Richard André-Jeannin. Divisibility of generalized Fibonacci and Lucas numbers by their subscripts. Fibonacci Quarterly, 29 (4):364–366, November 1991. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/29-4/andre-
jeannin2.pdf.

Andre-Jeannin:1991:GCF

[AJ91b] Richard André-Jeannin. Generalized complex Fibonacci and Lucas functions. Fibonacci Quarterly, 29(1):13–18, February 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-1/andre-jeannin.pdf.

Andre-Jeannin:1991:NIC

[AJ91c] Richard André-Jeannin. A note on the irrationality of certain Lucas infinite series. Fibonacci Quarterly, 29(2):132–135, May 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/29-2/andre-jeannin.pdf.

Andre-Jeannin:1991:DWE

[AJ91d] Richard André-Jeannin. On determinants whose elements are recurring sequences of arbitrary order. Fibonacci Quarterly, 29 (4):304–308, November 1991. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/29-4/andre-
jeannin1.pdf.

Andre-Jeannin:1991:SIS

[AJ91e] Richard André-Jeannin. Sequences of integers satisfying recurrence relations. Fibonacci Quarterly, 29(3):205–208, August 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-3/andre-jeannin2.pdf.

Andre-Jeannin:1991:SCR

[AJ91f] Richard André-Jeannin. Summation of certain reciprocal series related to Fibonacci and Lucas numbers. Fibonacci Quarterly, 29(3):200–204, August 1991. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/29-3/andre-
jeannin1.pdf.

Andre-Jeannin:1992:EWO

[AJ92] Richard André-Jeannin. On the equations U_{n} = Uqx^{2}, where q is odd, and V_{n} = V qx^{2}, where q is even. Fibonacci Quarterly, 30(2): 133–135, May 1992. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/30-2/andre-jeannin.pdf.

Andre-Jeannin:1994:GMV

[AJ94a] Richard André-Jeannin. A generalization of Morgan–Voyce polynomials. Fibonacci Quarterly, 32(3):228–231, June 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-3/andre-jeannin.pdf.

Andre-Jeannin:1994:NGC

[AJ94b] Richard André-Jeannin. A note on a general class of polynomials. Fibonacci Quarterly, 32(5):445–454, November 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-5/andre-jeannin.pdf.

Andre-Jeannin:1994:CPF

[AJ94c] Richard André-Jeannin. On a conjecture of Piero Filipponi. Fibonacci Quarterly, 32(1):11–13, February 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-1/andre-jeannin.pdf.

Andre-Jeannin:1995:DPG

[AJ95a] Richard André-Jeannin. Differential properties of a general class of polynomials. Fibonacci Quarterly, 33(5):453–458, November 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/33-5/andre-jeannin.pdf.

Andre-Jeannin:1995:NGG

[AJ95b] Richard André-Jeannin. A note on a general glass of polynomials, Part II. Fibonacci Quarterly, 33(4):341–351, August 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-4/andre-jeannin.pdf.

Andre-Jeannin:1996:EEF

[AJ96] Richard André-Jeannin. On the existence of even Fibonacci pseudoprimes with parameters P and O. Fibonacci Quarterly, 34(1): 75–78, February 1996. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/34-1/andre-jeannin.pdf.

Andre-Jeannin:1997:SRC

[AJ97] Richard André-Jeannin. Summation of reciprocals in certain second-order recurring sequences. Fibonacci Quarterly, 35(1):68– 74, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/35-1/andre-jeannin.pdf.

Andre-Jeannin:1998:ICI

[AJ98] Richard André-Jeannin. On the integrity of certain infinite series. Fibonacci Quarterly, 36(2):174–180, May 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-2/andre-jeannin.pdf.

Andrews:2022:APC

[AJS22] George E. Andrews, Matthew Just, and Greg Simay. Antipalindromic compositions. Fibonacci Quarterly, 60(2):158–??, May 2022. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/60-2/andrews.pdf.

Alexanderson:1974:FAG

[AK74] G. L. Alexanderson and L. F. Klosinski. A Fibonacci analogue of Gaussian binomial coefficients. Fibonacci Quarterly, 12(2): 129–132, April 1974. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/12-2/alexanderson.pdf.

Abuaiadh:1994:FHO

[AK94] D. Abuaiadh and J. H. Kingston. Are Fibonacci heaps optimal? Lecture Notes in Computer Science, 834:442–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antzoulakos:2005:CPR

[AK05] Demetrios L. Antzoulakos and Markos V. Koutras. On a class of polynomials related to classical orthogonal and Fibonacci polynomials with probability applications. Journal of Statistical Planning and Inference, 135(1):18–39, November 2005. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0378375805000522.

Atanassov:2003:IAR

[AKO^{+}03] K. T. Atanassov, R. Knott, K. Ozeki, A. G. Shannon, and L. Szalay. Inequalities among related pairs of Fibonacci numbers. Fibonacci Quarterly, 41(1):20–22, February 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-1/atanassov.pdf.

Aursukaree:2019:GHI

[AKP19] Saralee Aursukaree, Tammatada Khemaratchatakumthorn, and Prapanpong Pongsriiam. Generalizations of Hermite’s identity and applications. Fibonacci Quarterly, 57(2):126–??, May 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-2/aursukaree.pdf.

Aursukaree:2020:CGH

[AKP20] Saralee Aursukaree, Tammatada Khemaratchatakumthorn, and Prapanpong Pongsriiam. Corrigendum to Generalizations of Hermite’s Identity and Applications. Fibonacci Quarterly, 58(1):80–??, February 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/58-1/
PongsriiamCorrigendum11102019.pdf.

Akritas:1993:SFF

[Akr93] Alkiviadis G. Akritas. Sylvester’s forgotten form of the resultant. Fibonacci Quarterly, 31(4):325–332, November 1993. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-4/akritas.pdf.

Allard:1979:PEP

[AL79a] A. Allard and P. Lecomte. Periods and entry points in Fibonacci sequence. Fibonacci Quarterly, 17(1):51–57, February 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-1/allard.pdf.

Anderson:1979:DE

[AL79b] David A. Anderson and Milton W. Loyer. The Diophantine equation Nb^{2} = c^{2}+N+1. Fibonacci Quarterly, 17(1):69–70, February 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/17-1/anderson.pdf.

Altassan:2019:CPF

[AL19] Alaa Altassan and Florian Luca. On a curious property of F_{184}. Fibonacci Quarterly, 57(4):363–??, November 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-4/altassan.pdf.

Altassan:2020:E

[AL20] Alaa Altassan and Florian Luca. On the equation ∑_{j=1}^{k} jF_{j}^{p} = F_{n}^{q}. Journal of Number Theory, 217(??):256–277, December 2020. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X20301578.

Alameddine:1989:FNB

[Ala89] A. F. Alameddine. Fibonacci numbers and bipyramids. Fibonacci Quarterly, 27(3):247–252, June 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-3/
alameddine.pdf.

Alameddine:1998:BFN

[Ala98] Ahmad Fawzi Alameddine. Bounds on the Fibonacci number of a maximal outerplanar graph. Fibonacci Quarterly, 36(3):206–210, June 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/36-3/alameddine.pdf.

Alberti:1973:APC

[Alb73] Furio Alberti. Another proof for a continued fraction identity. Fibonacci Quarterly, 11(5):533–534, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-5/alberti.pdf.

Alberti:1981:RPP

[Alb81] Furio Alberti. A root property of a psi-type equation. Fibonacci Quarterly, 19(1):56–??, February 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
1/alberti.pdf.

Alfred:1963:AFF

[Alf63a] Brother U. Alfred. Additional factors of the Fibonacci and Lucas series. Fibonacci Quarterly, 1(1):34–42, February 1963. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-1/alfred1.pdf.

Alfred:1963:DRP

[Alf63b] Brother U. Alfred. Dying rabbit problem revived. Fibonacci Quarterly, 1(4):53–56, December 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-4/
alfred2.pdf.

Alfred:1963:EFN

[Alf63c] Brother U. Alfred. Exploring Fibonacci numbers. Fibonacci Quarterly, 1(1):57–63, February 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-1/
alfred3.pdf.

Alfred:1963:EFP

[Alf63d] Brother U. Alfred. Exploring Fibonacci polygons. Fibonacci Quarterly, 1(3):60–??, October 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-3/
alfred2.pdf.

Alfred:1963:ERS

[Alf63e] Brother U. Alfred. Exploring recurrent sequences. Fibonacci Quarterly, 1(2):81–83, April 1963. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/1-2/alfred.
pdf.

Alfred:1963:EFR

[Alf63f] Brother U. Alfred. Exploring the Fibonacci representation of integers. Fibonacci Quarterly, 1(4):72–??, December 1963. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-4/alfred3.pdf.

Alfred:1963:FPF

[Alf63g] Brother U. Alfred. On the form of primitive factors of Fibonacci numbers. Fibonacci Quarterly, 1(1):43–45, February 1963. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-1/alfred2.pdf.

Alfred:1963:OFS

[Alf63h] Brother U. Alfred. On the ordering of the Fibonacci sequence. Fibonacci Quarterly, 1(4):43–46, December 1963. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-4/alfred1.pdf.

Alfred:1963:PPF

[Alf63i] Brother U. Alfred. Periodic properties of Fibonacci summations. Fibonacci Quarterly, 1(3):33–41, October 1963. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-3/alfred1.pdf.

Alfred:1964:CFF

[Alf64a] Brother U. Alfred. Continued fractions and Fibonacci and Lucas ratios. Fibonacci Quarterly, 2(4):269–276, December 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-4/alfred1.pdf.

Alfred:1964:DML

[Alf64b] Brother U. Alfred. A digit muses …. Fibonacci Quarterly, 2(3): 210–??, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/2-3/alfred1.pdf.

Alfred:1964:EFM

[Alf64c] Brother U. Alfred. Exploring Fibonacci magic squares. Fibonacci Quarterly, 2(3):216–??, October 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-3/
alfred2.pdf.

Alfred:1964:EFN

[Alf64d] Brother U. Alfred. Exploring Fibonacci numbers with a calculator. Fibonacci Quarterly, 2(2):138–??, April 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/2-2/alfred2.pdf.

Alfred:1964:EFR

[Alf64e] Brother U. Alfred. Exploring Fibonacci residues. Fibonacci Quarterly, 2(1):42–??, February 1964. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/2-1/alfred3.
pdf.

Alfred:1964:EGA

[Alf64f] Brother U. Alfred. Exploring geometric–algebraic Fibonacci patterns. Fibonacci Quarterly, 2(4):318–319, December 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-4/alfred2.pdf.

Alfred:1964:SLN

[Alf64g] Brother U. Alfred. On square Lucas numbers. Fibonacci Quarterly, 2(1):11–12, February 1964. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/2-1/alfred1.
pdf.

Alfred:1964:PWF

[Alf64h] Brother U. Alfred. Primes which are factors of all Fibonacci sequences. Fibonacci Quarterly, 2(1):33–38, February 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-1/alfred2.pdf.

Alfred:1964:SDI

[Alf64i] Brother U. Alfred. Some determinants involving powers of Fibonacci numbers. Fibonacci Quarterly, 2(2):81–92, April 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-2/alfred1.pdf.

Alfred:1964:MNR

[Alf64j] U. Alfred. Mathematical notes: Relation of zeros to periods in the Fibonacci sequence modulo a prime. American Mathematical Monthly, 71(8):897–899, October 1964. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Alfred:1965:EGF

[Alf65a] Brother U. Alfred. Exploring generalized Fibonacci–Lucas relations. Fibonacci Quarterly, 3(4):319–320, December 1965. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/3-4/alfred.pdf.

Alfred:1965:MML

[Alf65b] Brother U. Alfred. Math morals (a limerick). Fibonacci Quarterly, 3(1):53–??, February 1965. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/3-1/alfred1.
pdf.

Alfred:1965:NTO

[Alf65c] Brother U. Alfred. Note on third order determinants. Fibonacci Quarterly, 3(1):59–60, February 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-1/
alfred2.pdf.

Alfred:1965:SLG

[Alf65d] Brother U. Alfred. Seeking the lost gold mine or exploring for Fibonacci factorizations. Fibonacci Quarterly, 3(2):129–130, April 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/3-2/alfred1.pdf.

Alfred:1965:SMS

[Alf65e] Brother U. Alfred. A strip method of summing linear Fibonacci expressions. Fibonacci Quarterly, 3(3):224–226, October 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/3-3/alfred.pdf.

Alfred:1966:BTP

[Alf66a] Brother U. Alfred. Algebra Through Problem Solving — book review. Fibonacci Quarterly, 4(3):264–??, October 1966. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-3/review2-a.pdf. See [HA66].

Alfred:1966:ESF

[Alf66b] Brother U. Alfred. Exploring special Fibonacci relations. Fibonacci Quarterly, 4(3):262–263, October 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/4-3/alfred.pdf.

Alladi:1975:AIF

[All75a] Krishnaswami Alladi. Approximation of irrationals with Farey Fibonacci fractions. Fibonacci Quarterly, 13(3):255–259, October 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/13-3/alladi2.pdf.

Alladi:1975:FSF

[All75b] Krishnaswami Alladi. A Farey sequence of Fibonacci numbers. Fibonacci Quarterly, 13(1):1–10, February 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-1/alladi1.pdf.

Alladi:1975:RMF

[All75c] Krishnaswami Alladi. A rapid method to form Farey Fibonacci fractions. Fibonacci Quarterly, 13(1):31–??, February 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-1/alladi2.pdf.

Alladi:1976:PGT

[All76] Krishnaswami Alladi. On polynomials generated by triangular arrays. Fibonacci Quarterly, 14(5):461–465, December 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-5/alladi.pdf.

Allen:1977:MPF

[All77] H. D. Allen. Metric paper to fall short of “Golden Mean”. Fibonacci Quarterly, 15(3):220–??, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/15-3/allen.pdf.

Alladi:1981:PRP

[All81] Krishnaswami Alladi. On the probability that n and Ω(n) are relatively prime. Fibonacci Quarterly, 19(3):228–232, August 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-3/alladi.pdf.

Allouche:2017:SEB

[All17] Jean-Paul Allouche. Some extremalities of the binary Fibonacci sequence. Fibonacci Quarterly, 55(5):??, December 2017. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-5/allouche.pdf.

Almkvist:1986:STP

[Alm86] Gert Almkvist. A solution to a tantalizing problem. Fibonacci Quarterly, 24(4):316–322, November 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-
4/almkvist.pdf.

Alonso:1976:ASH

[Alo76] James Alonso. Arithmetic sequences of higher order. Fibonacci Quarterly, 14(2):147–152, April 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-2/
alonso.pdf.

Alperin:2011:ISG

[Alp11] Roger C. Alperin. Integer sequences generated by x_{n+1} = (x_{n}^{2} + A)/(x_{n} − 1). Fibonacci Quarterly, 49(4):362–??, November 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/49-4/alperin.pdf.

Alperin:2018:CST

[Alp18] Roger C. Alperin. A-Cassini sequences and their spectrum. Fibonacci Quarterly, 56(2):153–??, May 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/56-
2/alperin.pdf.

Alperin:2019:CPS

[Alp19a] Roger C. Alperin. A-Cassini polynomial sequences and applications. Fibonacci Quarterly, 57(1):14–??, February 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-1/alperin.pdf.

Alperin:2019:FNR

[Alp19b] Roger C. Alperin. A family of nonlinear recurrences and their linear solutions. Fibonacci Quarterly, 57(4):318–??, November 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/57-4/alperin.pdf.

Alperin:2020:NRR

[Alp20] Roger C. Alperin. A nonlinear recurrence and its relations to Chebyshev polynomials. Fibonacci Quarterly, 58(2):140–??, May 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/58-2/alperin.pdf.

Altman:1988:BWB

[Alt88] N. S. Altman. Bit-wise behavior of random number generators. SIAM Journal on Scientific and Statistical Computing, 9(5):941– 949, September 1988. CODEN SIJCD4. ISSN 0196-5204.

Aluru:1996:PAL

[Alu96] S. Aluru. Parallel additive lagged Fibonacci random number generators. In ACM [ACM96], pages 102–108. ISBN 0-89791-803-7. LCCN QA76.5 I61 1996. ACM order number 415961.

Aluru:1997:LFR

[Alu97] Srinivas Aluru. Lagged Fibonacci random number generators for distributed memory parallel computers. Journal of Parallel and Distributed Computing, 45(1):1–12, August 25, 1997. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1997.1363/production;   http:/
/www.idealibrary.com/links/doi/10.1006/jpdc.1997.1363/
production/pdf;   http://www.idealibrary.com/links/doi/
10.1006/jpdc.1997.1363/production/ref.

Alvfeldt:1969:BTR

[Alv69] Olov Alvfeldt. A brain teaser related to Fibonacci numbers. Fibonacci Quarterly, 7(3):310–314, October 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/7-3/alvfeldt.pdf.

Alder:1975:IRN

[AM75a] H. L. Alder and Amin A. Muwafi. Identities relating the number of partitions into an even and odd number of parts. Fibonacci Quarterly, 13(2):147–149, April 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-2/
alder.pdf.

Alder:1975:GER

[AM75b] Henry L. Alder and Amin A. Muwafi. Generalizations of Euler’s recurrence formula for partitions. Fibonacci Quarterly, 13(4): 337–339, December 1975. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/13-4/alder.pdf.

Ainsworth:1980:PQO

[AM80] O. R. Ainsworth and Joseph E. Morris, Jr. A property of quasi-orthogonal polynomials. Fibonacci Quarterly, 18(2):163– 164, April 1980. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/18-2/ainsworth.pdf.

Andersen:2015:MEL

[AM15] Timothy D. Andersen and Michael Mascagni. Memory efficient lagged-Fibonacci random number generators for GPU supercomputing. Monte Carlo Methods and Applications, 21(2):163–174, June 2015. CODEN MCMAC6. ISSN 0929-9629 (print), 1569-

3961 (electronic). URL http://www.degruyter.com/view/

j/mcma.2015.21.issue-2/mcma-2014-0017/mcma-2014-0017.
xml.

Alder:1985:EPI

[AML85] Henry L. Alder, Amin A. Muwafi, and Jeffrey K. Lewis. Euler’s partition identity — are there any more like it? Fibonacci Quarterly, 23(2):113–119, May 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-2/
alder.pdf.

Mbirika:2022:GSC

[aMS22a] aBa Mbirika and Jürgen Spilker. GCD of sums of k consecutive squares of generalized Fibonacci numbers. Fibonacci Quarterly, 60(5):255–??, December 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/60-
5/mbirika.pdf.

Alekseyev:2022:TCR

[AMS^{+}22b] Max A. Alekseyev, Joseph Samuel Myers, Richard Schroeppel, S. R. Shannon, N. J. A. Sloane, and Paul Zimmermann. Three cousins of Recamán’s sequence. Fibonacci Quarterly, 60(3):201– ??, August 2022. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/60-3/alekseyev.pdf.

Ahuja:1966:NOP

[AN66] J. C. Ahuja and S. W. Nash. A note on orthogonal polynomials. Fibonacci Quarterly, 4(1):49–55, February 1966. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-1/ahuja.pdf.

Adikesavan:1979:MGB

[AN79] A. S. Adikesavan and S. Narayanaswami. A modification of Goka’s binary sequence. Fibonacci Quarterly, 17(3):212–219, October 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/17-3/adikesavan.pdf.

Ainsworth:1983:FPP

[AN83] O. R. Ainsworth and J. Neggers. A family of polynomials and powers of the secant. Fibonacci Quarterly, 21(2):132–138, May 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/21-2/ainsworth.pdf.

Andrews:1969:SFF

[And69] George E. Andrews. Some formulae for the Fibonacci sequence with generalizations. Fibonacci Quarterly, 7(2):113–130, April 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/7-2/andrews.pdf.

Anderson:1970:SCP

[And70] Jean H. Anderson. Spirals, checkerboards, polyominoes, and the Fibonacci sequence. Fibonacci Quarterly, 8(1):90–95, February 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-1/anderson.pdf.

Andreassian:1974:FSM

[And74a] Agnes Andreassian. Fibonacci sequences modulo m. Fibonacci Quarterly, 12(1):51–64, February 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
1/andreassian.pdf.

Andrews:1974:CAF

[And74b] George E. Andrews. Combinatorial analysis and Fibonacci numbers. Fibonacci Quarterly, 12(2):141–145, April 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/12-2/andrews.pdf.

Anderson:1978:SMP

[And78a] O. D. Anderson. Some more pattern from Pascal’s triangle. Fibonacci Quarterly, 16(4):296–301, August 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/16-4/anderson.pdf.

Anderson:1978:F

[And78b] Peter G. Anderson. On the formula π = 2 ∑ arccotf_{2k+1}. Fibonacci Quarterly, 16(2):118–??, April 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-2/
anderson.pdf.

Ando:1981:NPN

[And81] Shiro Ando. A note on the polygonal numbers. Fibonacci Quarterly, 19(2):180–183, April 1981. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/19-2/ando.pdf.

Ando:1982:SDE

[And82] Shiro Ando. On a system of Diophantine equations concerning the polygonal numbers. Fibonacci Quarterly, 20(4):349–353, November 1982. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/20-4/ando.pdf.

Ando:1984:NF

[And84] Shiro Ando. On the numbers of the form an^{2} + bn. Fibonacci Quarterly, 22(3):259–260, August 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-
3/ando.pdf.

Anderson:1990:RNG

[And90] Stuart L. Anderson. Random number generators on vector supercomputers and other advanced architectures. SIAM Review, 32(2):221–251, June 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://www.jstor.org/

stable/2030521.

Anderson:1995:FKG

[And95a] R. Anderson. On Fibonacci keystream generators. Lecture Notes in Computer Science, 1008:346–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ando:1995:SSD

[And95b] Shiro Ando. On a system of sequences defined by a recurrence relation. Fibonacci Quarterly, 33(3):279–282, June 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-3/ando.pdf.

Andaloro:2000:TST

[And00] Paul Andaloro. On total stopping times under 3x + 1 iteration. Fibonacci Quarterly, 38(1):73–78, February 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-1/andaloro.pdf.

Andaloro:2002:PDG

[And02] Paul J. Andaloro. The 3x + 1 problem and directed graphs. Fibonacci Quarterly, 40(1):43–54, February 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-1/andaloro.pdf.

Anderson:2005:BRB

[And05] Peter G. Anderson. Book review: Proofs that Really Count: The Art of Combinatorial Proof, by Arthur T. Benjamin and Jennifer J. Quinn. Fibonacci Quarterly, 43(4):326–327, November 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/43-4/paper43-4-6.pdf. See [BQ03].

Anderson:2014:MPZ

[And14] Peter G. Anderson. More properties of the Zeckendorf array. Fibonacci Quarterly, 52(5):15–??, December 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/52-5/Anderson.pdf.

Anderson:2017:NER

[And17] Peter G. Anderson. Notes & extensions for a remarkable continued fraction. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-5/anderson.pdf.

Anderson:2020:FWA

[And20] Peter G. Anderson. The Fibonacci word as a 2-adic number and its continued fraction. Fibonacci Quarterly, 58(5):21–??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/58-5/anderson.pdf.

Anisimov:1995:LFF

[Ani95] A. V. Anisimov. Linear Fibonacci forms and parallel algorithms for high dimension arithmetic. Lecture Notes in Computer Science, 964:16–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1963:CP

[Ano63a] Anonymous. Cover page. Fibonacci Quarterly, 1(1):??, February 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/1-1/editorial.pdf.

Anonymous:1963:GFS

[Ano63b] Anonymous. Generating Fibonacci series on a desk calculator. Fibonacci Quarterly, 1(2):56–??, April 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-2/
generating-a.pdf.

Anonymous:1963:HFQ

[Ano63c] Anonymous. History of the Fibonacci Quarterly. Fibonacci Quarterly, 1(4):8–??, December 1963. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/1-4/history1.
pdf.

Anonymous:1963:IV

[Ano63d] Anonymous. Index to Volume I. Fibonacci Quarterly, 1(4):??, December 1963. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/1-4/index.pdf.

Anonymous:1963:OMW

[Ano63e] Anonymous. Obituary for Morgan Ward. Fibonacci Quarterly, 1(3):32–??, October 1963. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/1-3/obit-ward.pdf.

Anonymous:1963:PDa

[Ano63f] Anonymous. Problem department. Fibonacci Quarterly, 1(1): 28–29, February 1963. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/1-1/problems1-1.pdf.

Anonymous:1963:PDb

[Ano63g] Anonymous. Problem department. Fibonacci Quarterly, 1(2): 74–??, April 1963. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/1-2/problems1-2.pdf.

Anonymous:1964:Ca

[Ano64a] Anonymous. Corrections. Fibonacci Quarterly, 2(1):32–??, February 1964. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/2-1/corrections1.pdf.

Anonymous:1964:Cb

[Ano64b] Anonymous. Corrections. Fibonacci Quarterly, 2(2):118–??, April 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/2-2/corrections.pdf.

Anonymous:1964:O

[Ano64c] Anonymous. Omissions. Fibonacci Quarterly, 2(2):122–??, April 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/2-2/omissions.pdf.

Anonymous:1964:VI

[Ano64d] Anonymous. Volume index. Fibonacci Quarterly, 2(4):331–??, December 1964. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/2-4/index.pdf.

Anonymous:1965:AC

[Ano65a] Anonymous. Acknowledgments and corrections. Fibonacci Quarterly, 3(2):160–??, April 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-2/
acknowledgments.pdf.

Anonymous:1965:Cb

[Ano65b] Anonymous. Correction. Fibonacci Quarterly, 3(2):114–??, April 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/3-2/correction.pdf.

Anonymous:1965:Ca

[Ano65c] Anonymous. Corrections. Fibonacci Quarterly, 3(1):66–??, February 1965. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/3-1/corrections.pdf.

Anonymous:1965:Cc

[Ano65d] Anonymous. Corrections. Fibonacci Quarterly, 3(3):184–??, October 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/3-3/corrections.pdf.

Anonymous:1965:O

[Ano65e] Anonymous. Omission. Fibonacci Quarterly, 3(4):257–??, December 1965. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/3-4/omission.pdf.

Anonymous:1965:VI

[Ano65f] Anonymous. Volume index. Fibonacci Quarterly, 3(4):331–??, December 1965. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/3-4/index.pdf.

Anonymous:1966:Aa

[Ano66a] Anonymous. Announcement. Fibonacci Quarterly, 4(3):201–??, October 1966. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/4-3/announcement1.pdf.

Anonymous:1966:Ab

[Ano66b] Anonymous. Announcement. Fibonacci Quarterly, 4(3):216–??, October 1966. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/4-3/announcement2.pdf.

Anonymous:1966:VI

[Ano66c] Anonymous. Volume index. Fibonacci Quarterly, 4(4):379–??, December 1966. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/4-4/index.pdf.

Anonymous:1967:C

[Ano67a] Anonymous. Corrections. Fibonacci Quarterly, 5(2):162–168, April 1967. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/5-2/corrections1.pdf.

Anonymous:1967:VI

[Ano67b] Anonymous. Volume index. Fibonacci Quarterly, 5(5):491–??, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/5-5/index.pdf.

Anonymous:1968:Ea

[Ano68a] Anonymous. Errata. Fibonacci Quarterly, 6(1):21–??, February 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/6-1/errata1.pdf.

Anonymous:1968:Eb

[Ano68b] Anonymous. Errata. Fibonacci Quarterly, 6(3):70–??, June 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-3/errata.pdf.

Anonymous:1968:VI

[Ano68c] Anonymous. Volume index. Fibonacci Quarterly, 6(6):408–??, December 1968. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/6-6/index.pdf.

Anonymous:1969:AM

[Ano69a] Anonymous. Association meeting. Fibonacci Quarterly, 7(3): 252–??, October 1969. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/7-3/meeting.pdf.

Anonymous:1969:C

[Ano69b] Anonymous. Corrections. Fibonacci Quarterly, 7(3):284–286, October 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/7-3/correction.pdf.

Anonymous:1969:E

[Ano69c] Anonymous. Errata. Fibonacci Quarterly, 7(1):13–??, February 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/7-1/errata-a.pdf.

Anonymous:1969:VI

[Ano69d] Anonymous. Volume index. Fibonacci Quarterly, 7(5):552–??, December 1969. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/7-5/index.pdf.

Anonymous:1970:AM

[Ano70a] Anonymous. Association meeting. Fibonacci Quarterly, 8(3): 336–??, April 1970. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/8-3/meeting.pdf.

Anonymous:1970:Ca

[Ano70b] Anonymous. Corrections. Fibonacci Quarterly, 8(1):87–??, February 1970. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/8-1/corrections-a.pdf.

Anonymous:1970:Cb

[Ano70c] Anonymous. Corrections. Fibonacci Quarterly, 8(3):248–??, April 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-3/corrections.pdf.

Anonymous:1970:E

[Ano70d] Anonymous. Errata. Fibonacci Quarterly, 8(5):530–??, December 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-5/errata-a.pdf.

Anonymous:1970:FRC

[Ano70e] Anonymous. Fall’s research conference. Fibonacci Quarterly, 8 (5):481–??, December 1970. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/8-5/conference.pdf.

Anonymous:1970:LE

[Ano70f] Anonymous. Letters to the editor. Fibonacci Quarterly, 8(1): 88–89, February 1970. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/8-1/letters.pdf.

Anonymous:1970:VI

[Ano70g] Anonymous. Volume index. Fibonacci Quarterly, 8(5):552–??, December 1970. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/8-5/index.pdf.

Anonymous:1971:E

[Ano71a] Anonymous. Errata. Fibonacci Quarterly, 9(1):81–??, February 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/9-1/errata.pdf.

Anonymous:1971:NAF

[Ano71b] Anonymous. Ninth Annual Fall Conference of the Fibonacci Association. Fibonacci Quarterly, 9(5):504–??, December 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-5/program-a.pdf.

Anonymous:1971:PEA

[Ano71c] Anonymous. Program of the eigth anniversary meeting. Fibonacci Quarterly, 9(4):412–??, October 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-4/
program-a.pdf.

Anonymous:1971:VI

[Ano71d] Anonymous. Volume index. Fibonacci Quarterly, 9(5):552–??, December 1971. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/9-5/index.pdf.

Anonymous:1972:A

[Ano72a] Anonymous. Announcement. Fibonacci Quarterly, 10(6):662–??, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/10-6/announcement.pdf.

Anonymous:1972:Ca

[Ano72b] Anonymous. Corrections. Fibonacci Quarterly, 10(2):198–??, February 1972. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/10-2/correction.pdf.

Anonymous:1972:Cb

[Ano72c] Anonymous. Corrigenda. Fibonacci Quarterly, 10(3):328–??, April 1972. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/10-3/corrigenda.pdf.

Anonymous:1972:Ea

[Ano72d] Anonymous. Errata. Fibonacci Quarterly, 10(4):448–453, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/10-4/errata.pdf.

Anonymous:1972:Eb

[Ano72e] Anonymous. Errata. Fibonacci Quarterly, 10(6):656–??, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/10-6/errata.pdf.

Anonymous:1972:FMS

[Ano72f] Anonymous. Fibonacci makes the sports page. Fibonacci Quarterly, 10(4):454–??, October 1972. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/10-4/sports.
pdf.

Anonymous:1972:PNA

[Ano72g] Anonymous. Program of the ninth anniversary meeting. Fibonacci Quarterly, 10(4):445–446, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-4/meeting.pdf.

Anonymous:1972:VI

[Ano72h] Anonymous. Volume index. Fibonacci Quarterly, 10(6):667–??, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/10-6/index.pdf.

Anonymous:1973:Ea

[Ano73a] Anonymous. Errata. Fibonacci Quarterly, 11(2):162–??, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/11-2/errata-a.pdf.

Anonymous:1973:Eb

[Ano73b] Anonymous. Errata. Fibonacci Quarterly, 11(3):294–??, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/11-3/errata-a.pdf.

Anonymous:1973:Ec

[Ano73c] Anonymous. Errata. Fibonacci Quarterly, 11(5):524–??, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/11-5/errata.pdf.

Anonymous:1973:FAM

[Ano73d] Anonymous. Fibonacci association meeting: Conference program. Fibonacci Quarterly, 11(1):55–??, February 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-1/program.pdf.

Anonymous:1973:LEb

[Ano73e] Anonymous. Letter to the editor. Fibonacci Quarterly, 11(5): 500–??, December 1973. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/11-5/letter.pdf.

Anonymous:1973:LEa

[Ano73f] Anonymous. Letters to the editors. Fibonacci Quarterly, 11(3): 284–??, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/11-3/letters.pdf.

Anonymous:1973:VI

[Ano73g] Anonymous. Volume index. Fibonacci Quarterly, 11(5):554–??, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/11-5/index.pdf.

Anonymous:1974:E

[Ano74a] Anonymous. Errata. Fibonacci Quarterly, 12(4):345–??, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/12-4/errata.pdf.

Anonymous:1974:FAM

[Ano74b] Anonymous. Fibonacci association meeting: Program. Fibonacci Quarterly, 12(1):66–??, February 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
1/program.pdf.

Anonymous:1974:FAR

[Ano74c] Anonymous. The Fibonacci Association Research Conference. Fibonacci Quarterly, 12(3):232–??, October 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/12-3/program.pdf.

Anonymous:1974:VI

[Ano74d] Anonymous. Volume index. Fibonacci Quarterly, 12(4):407–??, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/12-4/index.pdf.

Anonymous:1975:LE

[Ano75a] Anonymous. Letter to the editor. Fibonacci Quarterly, 13(3): 198–??, October 1975. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/13-3/letter.pdf.

Anonymous:1975:VI

[Ano75b] Anonymous. Volume index. Fibonacci Quarterly, 13(4):378–??, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/13-4/index.pdf.

Anonymous:1976:LEa

[Ano76a] Anonymous. Letter to the editor. Fibonacci Quarterly, 14(1): 12–??, February 1976. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/14-1/letter1.pdf.

Anonymous:1976:LEb

[Ano76b] Anonymous. Letter to the editor. Fibonacci Quarterly, 14(1): 48–??, February 1976. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/14-1/letter2-a.pdf.

Anonymous:1976:LEc

[Ano76c] Anonymous. Letter to the editor. Fibonacci Quarterly, 14(4): 368–??, November 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-4/letter.pdf.

Anonymous:1976:VI

[Ano76d] Anonymous. Volume index. Fibonacci Quarterly, 14(5):474–??, December 1976. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/14-5/index.pdf.

Anonymous:1977:A

[Ano77a] Anonymous. Addendum. Fibonacci Quarterly, 15(4):361–??, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/15-4/addendum.pdf.

Anonymous:1977:FAR

[Ano77b] Anonymous. Fibonacci association research conference. Fibonacci Quarterly, 15(4):342–??, December 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
4/program.pdf.

Anonymous:1977:LE

[Ano77c] Anonymous. Letter to the editor. Fibonacci Quarterly, 15(4): 346–??, December 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-4/letter.pdf.

Anonymous:1977:VI

[Ano77d] Anonymous. Volume index. Fibonacci Quarterly, 15(4):378–??, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/15-4/index.pdf.

Anonymous:1978:AAP

[Ano78a] Anonymous. Addenda to advanced problems and solutions. Fibonacci Quarterly, 16(2):154–??, April 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-2/
addenda-a.pdf.

Anonymous:1978:VI

[Ano78b] Anonymous. Volume index. Fibonacci Quarterly, 16(6):570–??, December 1978. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/16-6/index.pdf.

Anonymous:1979:E

[Ano79a] Anonymous. Errata. Fibonacci Quarterly, 17(2):188–??, April 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/17-2/errata.pdf.

Anonymous:1979:IAP

[Ano79b] Anonymous. Index of advanced problems. Fibonacci Quarterly, 17(4):379–??, December 1979. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/17-4/index-
advanced.pdf.

Anonymous:1979:VI

[Ano79c] Anonymous. Volume index. Fibonacci Quarterly, 17(4):380–??, December 1979. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/17-4/index.pdf.

Anonymous:1980:BA

[Ano80a] Anonymous. Book announcement. Fibonacci Quarterly, 18(2): 185–??, April 1980. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/18-2/announcement.pdf.

Anonymous:1980:C

[Ano80b] Anonymous. Corrections. Fibonacci Quarterly, 18(1):82–83, February 1980. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/18-1/corrections.pdf.

Anonymous:1980:VI

[Ano80c] Anonymous. Volume index. Fibonacci Quarterly, 18(4):382–??, December 1980. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/18-4/index.pdf.

Anonymous:1981:VI

[Ano81] Anonymous. Volume index. Fibonacci Quarterly, 19(5):477–??, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/19-5/index.pdf.

Anonymous:1982:A

[Ano82a] Anonymous. Acknowledgments. Fibonacci Quarterly, 20(1):88– ??, February 1982. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/20-1/acknowledgments.pdf.

Anonymous:1982:AFR

[Ano82b] Anonymous. Announcement of Fibonacci Research Conference. Fibonacci Quarterly, 20(3):248–??, August 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-3/announcement.pdf.

Anonymous:1982:E

[Ano82c] Anonymous. Errata. Fibonacci Quarterly, 20(2):192–??, May 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/20-2/errata.pdf.

Anonymous:1982:VI

[Ano82d] Anonymous. Volume index. Fibonacci Quarterly, 20(4):381–??, November 1982. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/20-4/index.pdf.

Anonymous:1983:A

[Ano83a] Anonymous. Acknowledgments. Fibonacci Quarterly, 21(1):2–??, February 1983. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/21-1/acknowledgments.pdf.

Anonymous:1983:CA

[Ano83b] Anonymous. Conference announcement. Fibonacci Quarterly, 21 (4):318–??, November 1983. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/21-4/announcement.
pdf.

Anonymous:1983:VI

[Ano83c] Anonymous. Volume index. Fibonacci Quarterly, 21(4):319–??, November 1983. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/21-4/index.pdf.

Anonymous:1984:A

[Ano84a] Anonymous. Acknowledgments. Fibonacci Quarterly, 22(1):83– ??, February 1984. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/22-1/acknowledgments.pdf.

Anonymous:1984:FRCa

[Ano84b] Anonymous. Fibonacci Research Conference announcement. Fibonacci Quarterly, 22(1):82–??, February 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/22-1/announcement.pdf.

Anonymous:1984:FRCb

[Ano84c] Anonymous. Fibonacci Research Conference announcement. Fibonacci Quarterly, 22(2):181–182, May 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-2/
announcement.pdf.

Anonymous:1984:VI

[Ano84d] Anonymous. Volume index. Fibonacci Quarterly, 22(4):382–??, November 1984. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/22-4/index.pdf.

Anonymous:1985:Aa

[Ano85a] Anonymous. Acknowledgments. Fibonacci Quarterly, 23(1):2–??, February 1985. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/23-1/acknowledgments.pdf.

Anonymous:1985:Ab

[Ano85b] Anonymous. Announcement. Fibonacci Quarterly, 23(3):276–??, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/23-3/announcement.pdf.

Anonymous:1985:Ac

[Ano85c] Anonymous. Announcement. Fibonacci Quarterly, 23(4):346– ??, November 1985. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/23-4/announcement.pdf.

Anonymous:1985:VI

[Ano85d] Anonymous. Volume index. Fibonacci Quarterly, 23(4):383–??, November 1985. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/23-4/index.pdf.

Anonymous:1986:Ab

[Ano86a] Anonymous. Acknowledgements. Fibonacci Quarterly, 24(2):98– ??, May 1986. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/24-2/acknowledgements.pdf.

Anonymous:1986:Aa

[Ano86b] Anonymous. Announcement. Fibonacci Quarterly, 24(1):83–??, February 1986. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/24-1/announcement.pdf.

Anonymous:1986:RCA

[Ano86c] Anonymous. Research conference announcement. Fibonacci Quarterly, 24(2):144–??, May 1986. CODEN FIBQAU. ISSN

0015-0517.  URL http://www.fq.math.ca/Scanned/24-2/
announcement.pdf.

Anonymous:1986:VI

[Ano86d] Anonymous. Volume index. Fibonacci Quarterly, 24(4):383–??, November 1986. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/24-4/index.pdf.

Anonymous:1987:R

[Ano87a] Anonymous. Referees. Fibonacci Quarterly, 25(1):84–??, February 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/25-1/referees.pdf.

Anonymous:1987:TAC

[Ano87b] Anonymous. Third Annual Conference announcement. Fibonacci Quarterly, 25(4):290–??, November 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-
4/announcement.pdf.

Anonymous:1987:TICa

[Ano87c] Anonymous. Third International Conference announcement. Fibonacci Quarterly, 25(2):105–??, May 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-2/
announcement.pdf.

Anonymous:1987:TICb

[Ano87d] Anonymous. Third International Conference announcement. Fibonacci Quarterly, 25(3):244–??, August 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-3/announcement.pdf.

Anonymous:1987:VI

[Ano87e] Anonymous. Volume index. Fibonacci Quarterly, 25(4):383–??, November 1987. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/25-4/index.pdf.

Anonymous:1988:BRB

[Ano88a] Anonymous. Book review: The Book of Squares. Fibonacci Quarterly, 26(4):382–??, November 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
4/review.pdf. See [FS87].

Anonymous:1988:R

[Ano88b] Anonymous. Referees. Fibonacci Quarterly, 26(1):84–??, February 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/26-1/referees.pdf.

Anonymous:1988:TEB

[Ano88c] Anonymous. Table errata: The book of squares [Academic Press, Boston, Mass., 1987, MR 88m:01021a] by L. P. Fibonacci. Mathematics of Computation, 50(182):654, April 1988. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:1988:TICa

[Ano88d] Anonymous. Third International Conference announcement. Fibonacci Quarterly, 26(1):2–??, February 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-1/
announcement.pdf.

Anonymous:1988:TICb

[Ano88e] Anonymous. Third International Conference announcement. Fibonacci Quarterly, 26(2):130–??, May 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-2/
announcement.pdf.

Anonymous:1988:TBA

[Ano88f] Anonymous. A tribute to Brother Alfred Brousseau. Fibonacci Quarterly, 26(3):194–??, August 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
3/memoriam.pdf.

Anonymous:1988:VI

[Ano88g] Anonymous. Volume index. Fibonacci Quarterly, 26(4):383–??, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/26-4/index.pdf.

Anonymous:1989:AFIb

[Ano89a] Anonymous. Announcement: Fourth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 27(3):208–??, June 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-3/
announcement.pdf.

Anonymous:1989:AFIc

[Ano89b] Anonymous. Announcement: Fourth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 27(4):316–??, August 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-
4/announcement.pdf.

Anonymous:1989:AFId

[Ano89c] Anonymous. Announcement: Fourth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 27(5):408–??, November 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-
5/announcement.pdf.

Anonymous:1989:AFIa

[Ano89d] Anonymous. Announcement on Fourth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 27(2):124–??, May 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-2/
announcement.pdf.

Anonymous:1989:R

[Ano89e] Anonymous. Referees. Fibonacci Quarterly, 27(1):31–32, February 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/27-1/referees.pdf.

Anonymous:1989:VI

[Ano89f] Anonymous. Volume index. Fibonacci Quarterly, 27(5):479–??, November 1989. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/27-5/index.pdf.

Anonymous:1990:AFId

[Ano90a] Anonymous. Announcement: Fifth International Conference. Fibonacci Quarterly, 28(4):384–??, November 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/28-4/announcement.pdf.

Anonymous:1990:AFIa

[Ano90b] Anonymous. Announcement: Fourth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 28(1):15–??, February 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
1/announcement.pdf.

Anonymous:1990:AFIc

[Ano90c] Anonymous. Announcement: Fourth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 28(3):214–??, August 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
3/announcement.pdf.

Anonymous:1990:AFIb

[Ano90d] Anonymous. Announcement of Fourth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 28(2):192–193, May 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-2/
announcement.pdf.

Anonymous:1990:BR

[Ano90e] Anonymous. Book review: A new chapter for Pythagorean triples, by A. G. Schaake and J. C. Turner. Fibonacci Quarterly, 28(2):140, 155, May 1990. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/28-2/review-a.

pdf. See [ST89].

Anonymous:1990:R

[Ano90f] Anonymous. Referees. Fibonacci Quarterly, 28(1):2–??, February 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/28-1/referees-a.pdf.

Anonymous:1990:VI

[Ano90g] Anonymous. Volume index. Fibonacci Quarterly, 28(4):383–??, November 1990. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/28-4/index.pdf.

Anonymous:1991:AFIa

[Ano91a] Anonymous. Announcement: Fifth International Conference. Fibonacci Quarterly, 29(1):23–??, February 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/29-1/announcement.pdf.

Anonymous:1991:AFIb

[Ano91b] Anonymous. Announcement: Fifth International Conference. Fibonacci Quarterly, 29(2):136–??, May 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-2/
announcement.pdf.

Anonymous:1991:AFIc

[Ano91c] Anonymous. Announcement: Fifth International Conference. Fibonacci Quarterly, 29(3):199–??, August 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/29-3/announcement.pdf.

Anonymous:1991:AFId

[Ano91d] Anonymous. Announcement: Fifth International Conference. Fibonacci Quarterly, 29(4):309–??, November 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/29-4/announcement.pdf.

Anonymous:1991:EPE

[Ano91e] Anonymous. Elementary problems editor retires. Fibonacci Quarterly, 29(2):180–??, May 1991. CODEN FIBQAU. ISSN

0015-0517.  URL http://www.fq.math.ca/Scanned/29-2/
notice.pdf.

Anonymous:1991:R

[Ano91f] Anonymous. Referees. Fibonacci Quarterly, 29(1):2–??, February 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/29-1/referees.pdf.

Anonymous:1991:VI

[Ano91g] Anonymous. Volume index. Fibonacci Quarterly, 29(4):383–??, November 1991. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/29-4/index.pdf.

Anonymous:1992:AFIa

[Ano92a] Anonymous. Announcement: Fifth International Conference. Fibonacci Quarterly, 30(1):20–??, February 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-1/announcement.pdf.

Anonymous:1992:AFIb

[Ano92b] Anonymous. Announcement: Fifth International Conference. Fibonacci Quarterly, 30(2):132–??, May 1992. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/30-2/
announcement.pdf.

Anonymous:1992:R

[Ano92c] Anonymous. Referees. Fibonacci Quarterly, 30(1):84–??, February 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/30-1/referees.pdf.

Anonymous:1992:VI

[Ano92d] Anonymous. Volume index. Fibonacci Quarterly, 30(4):383–??, November 1992. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/30-4/index.pdf.

Anonymous:1993:ASIa

[Ano93a] Anonymous. Announcement: Sixth International Conference. Fibonacci Quarterly, 31(2):145–??, May 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-2/
announcement.pdf.

Anonymous:1993:ASIb

[Ano93b] Anonymous. Announcement: Sixth International Conference. Fibonacci Quarterly, 31(3):255–??, August 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-3/announcement.pdf.

Anonymous:1993:ASIc

[Ano93c] Anonymous. Announcement: Sixth International Conference. Fibonacci Quarterly, 31(4):370–??, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-4/announcement.pdf.

Anonymous:1993:ATI

[Ano93d] Anonymous. Author and title index for sale. Fibonacci Quarterly, 31(1):40–??, February 1993. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/31-1/index.pdf.

Anonymous:1993:BABc

[Ano93e] Anonymous. Book announcement: Generalized Pascal Triangles and Pyramids, by Boris A. Bondarenko. Fibonacci Quarterly, 31(3):226–??, August 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-
3/bondarenko.pdf.

Anonymous:1993:BABa

[Ano93f] Anonymous. Book announcement: Generalized Pascal Triangles and Pyramids. Their Fractals and Graphs and Applications, by Boris A. Bondarenko. Fibonacci Quarterly, 31(1):52–??, February 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/31-1/bondarenko.pdf.

Anonymous:1993:BABb

[Ano93g] Anonymous. Book announcement: Generalized Pascal Triangles and Pyramids. Their Fractals Graphs and Applications, by Boris A. Bondarenko. Fibonacci Quarterly, 31(2):165–??, May 1993.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-2/bondarenko.pdf.

Anonymous:1993:C

[Ano93h] Anonymous. Correction. Fibonacci Quarterly, 31(2):133–??, May 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/31-2/correction.pdf.

Anonymous:1993:NBL

[Ano93i] Anonymous. A new book on Liber Abaci. Fibonacci Quarterly, 31(1):72–??, February 1993. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/31-1/announcement.

pdf. See [Lün92].

Anonymous:1993:R

[Ano93j] Anonymous. Referees. Fibonacci Quarterly, 31(1):81–??, February 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/31-1/referees-a.pdf.

Anonymous:1993:SIC

[Ano93k] Anonymous. Sixth International Conference on Fibonacci Numbers and Their Applications. Journal of Computational and Applied Mathematics, 47(3):N9, September 30, 1993. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037704279390080U.

Anonymous:1993:VI

[Ano93l] Anonymous. Volume index. Fibonacci Quarterly, 31(4):383–??, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/31-4/index.pdf.

Anonymous:1994:Aa

[Ano94a] Anonymous. Announcement. Fibonacci Quarterly, 32(3):288– 289, June 1994. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/32-3/announcement.pdf.

Anonymous:1994:Ab

[Ano94b] Anonymous. Announcement. Fibonacci Quarterly, 32(5):440– ??, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/32-5/announcement.pdf.

Anonymous:1994:ASI

[Ano94c] Anonymous. Announcement: Sixth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 32(1):14–??, February 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-
1/announcement.pdf.

Anonymous:1994:NEP

[Ano94d] Anonymous. New editorial policies. Fibonacci Quarterly, 32(5): 411–??, November 1994. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/32-5/editorial.pdf.

Anonymous:1994:R

[Ano94e] Anonymous. Referees. Fibonacci Quarterly, 32(1):58–59, February 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/32-1/referees.pdf.

Anonymous:1994:VI

[Ano94f] Anonymous. Volume index. Fibonacci Quarterly, 32(5):479–??, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/32-5/index.pdf.

Anonymous:1995:FF

[Ano95a] Anonymous. Fibonacci forgeries. Scientific American, 272(5): 82–??, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:FNS

[Ano95b] Anonymous. The Fibonacci Number Series, volume 302 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/etext95/fibns10.zip.

Anonymous:1995:MRa

[Ano95c] Anonymous. Mathematical recreations. Scientific American, 272 (1):96–??, January 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1995:R

[Ano95d] Anonymous. Referees. Fibonacci Quarterly, 33(1):2–??, February 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/33-1/referees.pdf.

Anonymous:1995:SIRa

[Ano95e] Anonymous. Seventh International Research Conference. Fibonacci Quarterly, 33(3):221–??, June 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-3/
announcement.pdf.

Anonymous:1995:SIRb

[Ano95f] Anonymous. Seventh International Research Conference. Fibonacci Quarterly, 33(4):303–??, August 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-4/announcement.pdf.

Anonymous:1995:SIRc

[Ano95g] Anonymous. Seventh International Research Conference. Fibonacci Quarterly, 33(5):391–??, November 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-5/announcement.pdf.

Anonymous:1995:VI

[Ano95h] Anonymous. Volume index. Fibonacci Quarterly, 33(5):478–??, November 1995. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/33-5/index.pdf.

Anonymous:1996:DHT

[Ano96a] Anonymous. Dedication to Herta Taussig Freitag. Fibonacci Quarterly, 34(5):467–??, November 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-
5/dedication.pdf.

Anonymous:1996:MPS

[Ano96b] Anonymous. In memory of Professor Steven Vajda. Fibonacci Quarterly, 34(2):175–??, May 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-2/
vajda.pdf.

Anonymous:1996:R

[Ano96c] Anonymous. Referees. Fibonacci Quarterly, 34(1):79–80, February 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/34-1/referees.pdf.

Anonymous:1996:SIR

[Ano96d] Anonymous. Seventh International Research Conference. Fibonacci Quarterly, 34(1):17–??, February 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-1/announcement.pdf.

Anonymous:1996:VI

[Ano96e] Anonymous. Volume index. Fibonacci Quarterly, 34(5):479–??, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/34-5/index.pdf.

Anonymous:1997:AEI

[Ano97a] Anonymous. Announcement of the Eighth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 35(1):28–??, February 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-
1/announcement.pdf.

Anonymous:1997:EAC

[Ano97b] Anonymous. Eighth Annual Conference announcement. Fibonacci Quarterly, 35(3):232–??, August 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-3/announcement.pdf.

Anonymous:1997:EICa

[Ano97c] Anonymous. Eighth International Conference announcement. Fibonacci Quarterly, 35(2):136–??, May 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-2/
announcement.pdf.

Anonymous:1997:EICb

[Ano97d] Anonymous. Eighth International Conference announcement. Fibonacci Quarterly, 35(4):317–??, November 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-4/announcement.pdf.

Anonymous:1997:NE

[Ano97e] Anonymous. New editor. Fibonacci Quarterly, 35(3):268–??, August 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/35-3/editor.pdf.

Anonymous:1997:NES

[Ano97f] Anonymous. New editor and submission of articles. Fibonacci Quarterly, 35(4):308–??, November 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-
4/editor.pdf.

Anonymous:1997:R

[Ano97g] Anonymous. Referees. Fibonacci Quarterly, 35(1):2–??, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/35-1/referees.pdf.

Anonymous:1997:RSM

[Ano97h] Anonymous. Retirement of subscription manager. Fibonacci Quarterly, 35(1):10–??, February 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-
1/retirement.pdf.

Anonymous:1997:VI

[Ano97i] Anonymous. Volume index. Fibonacci Quarterly, 35(4):383–??, November 1997. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/35-4/index.pdf.

Anonymous:1998:EIC

[Ano98a] Anonymous. Eighth International Conference announcement. Fibonacci Quarterly, 36(1):55–??, February 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-1/announcement.pdf.

Anonymous:1998:NPW

[Ano98b] Anonymous. New problem Web site. Fibonacci Quarterly, 36(5): 406–??, November 1998. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/36-5/problems.pdf.

Anonymous:1998:R

[Ano98c] Anonymous. Referees. Fibonacci Quarterly, 36(1):2–??, February 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/36-1/referees.pdf.

Anonymous:1998:TOE

[Ano98d] Anonymous. A toast to our editor. Fibonacci Quarterly, 36(4): 372–??, August 1998. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/36-4/toast.pdf.

Anonymous:1999:ANIa

[Ano99a] Anonymous. Announcement of the Ninth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 37(3):264–??, August 1999. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/37-
3/announcement.pdf.

Anonymous:1999:ANIb

[Ano99b] Anonymous. Announcement of the Ninth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 37(4):366–??, November 1999. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/37-
4/announcement.pdf.

Anonymous:1999:C

[Ano99c] Anonymous. Corrigendum. Fibonacci Quarterly, 37(1):33–??, February 1999. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/37-1/corrigendum.pdf.

Anonymous:1999:MRS

[Ano99d] Anonymous. In memoriam — Richard Spain Vine. Fibonacci Quarterly, 37(2):177–??, May 1999. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/37-2/in-
memoriam.pdf.

Anonymous:1999:R

[Ano99e] Anonymous. Referees. Fibonacci Quarterly, 37(1):2–??, February 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/37-1/referees.pdf.

Anonymous:1999:VI

[Ano99f] Anonymous. Volume index. Fibonacci Quarterly, 37(4):383–??, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/37-4/index.pdf.

Anonymous:19xx:FLL

[Anoxx] Anonymous. Fibonacci, Leonardo, or Leonardo of Pisa. Online encyclopedia article., 19xx. URL http://www.encyclopedia.

com/doc/1G2-2830901418.html;   https://www.math.utah.
edu/pub/tex/bib/fibquart.bib.

Anonymous:2000:ANIa

[Ano00a] Anonymous. Announcement: Ninth International Conference. Fibonacci Quarterly, 38(1):60–??, February 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/38-1/announcement.pdf.

Anonymous:2000:ANIb

[Ano00b] Anonymous. Announcement: Ninth International Conference. Fibonacci Quarterly, 38(2):180–??, May 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-2/
announcement.pdf.

Anonymous:2000:ANIc

[Ano00c] Anonymous. Announcement: Ninth International Conference. Fibonacci Quarterly, 38(3):216–??, June 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-3/
announcement.pdf.

Anonymous:2000:MGS

[Ano00d] Anonymous. A message of gratitude to Stanley Rabinowitz. Fibonacci Quarterly, 38(2):179–??, May 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-2/
message.pdf.

Anonymous:2000:NEP

[Ano00e] Anonymous. New elementary problem editors. Fibonacci Quarterly, 38(1):84–??, February 2000. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/38-1/editors.
pdf.

Anonymous:2000:R

[Ano00f] Anonymous. Referees. Fibonacci Quarterly, 38(1):2–??, February 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/38-1/referees.pdf.

Anonymous:2000:VI

[Ano00g] Anonymous. Volume index. Fibonacci Quarterly, 38(5):479–??, November 2000. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/38-5/index.pdf.

Anonymous:2001:ATIa

[Ano01a] Anonymous. Announcement of the Tenth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 39(4):384–385, August 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-
4/announcement.pdf.

Anonymous:2001:ATIb

[Ano01b] Anonymous. Announcement of the Tenth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 39(5):478–??, November 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-
5/announcement.pdf.

Anonymous:2001:R

[Ano01c] Anonymous. Referees. Fibonacci Quarterly, 39(1):2–??, February 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/39-1/referees.pdf.

Anonymous:2001:VI

[Ano01d] Anonymous. Volume index. Fibonacci Quarterly, 39(5):479–??, November 2001. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/39-5/index.pdf.

Anonymous:2002:ATI

[Ano02a] Anonymous. Announcement of the Tenth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 40(1):90–??, February 2002. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/40-
1/announcement.pdf.

Anonymous:2002:DCL

[Ano02b] Anonymous. Dedication to Cal Long. Fibonacci Quarterly, 40(3): 242–??, June 2002. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/40-3/tribute-a.pdf.

Anonymous:2002:PTF

[Ano02c] Anonymous. The passing of three Fibonacci Association friends. Fibonacci Quarterly, 40(5):398–??, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-5/passing.pdf.

Anonymous:2002:R

[Ano02d] Anonymous. Referees. Fibonacci Quarterly, 40(1):2–??, February 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/40-1/referees.pdf.

Anonymous:2002:RTI

[Ano02e] Anonymous. Report on the Tenth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 40(5):416–??, November 2002. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/40-5/report.
pdf.

Anonymous:2002:VI

[Ano02f] Anonymous. Volume index. Fibonacci Quarterly, 40(5):479–??, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/40-5/index.pdf.

Anonymous:2003:ICF

[Ano03a] Anonymous. 11th International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 41(4):384–385, August 2003. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/41-4/announcement.pdf.

Anonymous:2003:EIC

[Ano03b] Anonymous. The Eleventh International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 41(5): 450–??, November 2003. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/41-5/announcement.
pdf.

Anonymous:2003:R

[Ano03c] Anonymous. Referees. Fibonacci Quarterly, 41(1):2–??, February 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/41-1/referees.pdf.

Anonymous:2003:TJV

[Ano03d] Anonymous. Tribute to JoAnn Vine. Fibonacci Quarterly, 41(2): 180–??, May 2003. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/41-2/tribute.pdf.

Anonymous:2003:VI

[Ano03e] Anonymous. Volume index. Fibonacci Quarterly, 41(5):478–??, November 2003. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/41-5/index.pdf.

Anonymous:2004:EFC

[Ano04a] Anonymous. Errata for 2004: Fibonacci in complex camouflage: 10858. American Mathematical Monthly, 111(10):922, December 2004. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/

4145117. See errata [KC04].

Anonymous:2004:R

[Ano04b] Anonymous. Referees for 2003. Fibonacci Quarterly, 42(1):2– ??, February 2004. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Papers1/42-1/febreferee04.pdf.

Anonymous:2004:VI

[Ano04c] Anonymous. Volume index. Fibonacci Quarterly, 42(4):383–??, November 2004. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Papers1/42-4/volume_index_2004.pdf.

Anonymous:2005:R

[Ano05a] Anonymous. Referees. Fibonacci Quarterly, 43(1):2–??, February 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers/43-1/referees43.pdf.

Anonymous:2005:VI

[Ano05b] Anonymous. Volume index. Fibonacci Quarterly, 43(4):383–??, November 2005. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/volume_index_2005.pdf.

Anonymous:2006:R

[Ano06a] Anonymous. Referees. Fibonacci Quarterly, 44(1):2–??, February 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers/44-1/febreferee06.pdf.

Anonymous:2006:VI

[Ano06b] Anonymous. Volume index. Fibonacci Quarterly, 44(4):383–??, November 2006. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/volume_index_2006.pdf.

Anonymous:2007:MPS

[Ano07a] Anonymous. In memoriam: Patricia A. Solsaa. Fibonacci Quarterly, 45(4):290–??, November 2007. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/45-
4/quartpatty04_2007.pdf.

Anonymous:2007:R

[Ano07b] Anonymous. Referees. Fibonacci Quarterly, 45(1):2–??, February 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.

math.ca/Papers/45-1/febreferee06_for_2007_issue.pdf.

Anonymous:2007:VI

[Ano07c] Anonymous. Volume index. Fibonacci Quarterly, 45(4):383–??, November 2007. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/volume_index_2007.pdf.

Anonymous:2008:FAF

[Ano09a] Anonymous. First announcement: Fourteenth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 46/47(3):261–??, August 2008/2009. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/46_47-3/Announcement4.pdf.

Anonymous:2008:VI

[Ano09b] Anonymous. Volume index. Fibonacci Quarterly, 46/47(4): 381–??, November 2008/2009. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Volume_Index/volume_
index2008_09.pdf.

Anonymous:2010:FCA

[Ano10a] Anonymous. Fibonacci Conference announcement. Fibonacci Quarterly, 48(2):192–193, May 2010. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/48-2/
Announcement4.pdf.

Anonymous:2010:MBJ

[Ano10b] Anonymous. Marjorie Bicknell-Johnson. Fibonacci Quarterly, 48 (3):194–??, August 2010. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers/48-3/Dedication4.pdf.

Anonymous:2010:R

[Ano10c] Anonymous. Referees. Fibonacci Quarterly, 48(1):2–3, February 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers/48-1/referees10.pdf.

Anonymous:2010:VI

[Ano10d] Anonymous. Volume index. Fibonacci Quarterly, 48(4):381–??, November 2010. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/volume_index_2010.pdf.

Anonymous:2011:FCAa

[Ano11a]  Anonymous.  Fibonacci Conference announcement.  Fi-

bonacci Quarterly, 49(2):98–??, May 2011. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/49-2/
Announcement2012.pdf.

Anonymous:2011:FCAb

[Ano11b] Anonymous. Fibonacci Conference announcement. Fibonacci Quarterly, 49(3):194–??, August 2011. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/49-3/
Announcement2012.pdf.

Anonymous:2011:FCAc

[Ano11c] Anonymous. Fibonacci Conference announcement. Fibonacci Quarterly, 49(4):366–??, November 2011. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/49-4/
Announcement2012.pdf.

Anonymous:2011:R

[Ano11d] Anonymous. Referees. Fibonacci Quarterly, 49(1):2–3, February 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers/49-1/referees11.pdf.

Anonymous:2011:VI

[Ano11e] Anonymous. Volume index. Fibonacci Quarterly, 49(4):381–??, November 2011. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/volume_index_2011.pdf.

Anonymous:2012:FCAb

[Ano12a] Anonymous. Fibonacci conference announcement. Fibonacci Quarterly, 50(2):98–??, May 2012. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
Announcement2012.pdf.

Anonymous:2012:VI

[Ano12b] Anonymous. Volume index. Fibonacci Quarterly, 50(4):382–??, November 2012. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/VolumeIndex2012.pdf.

Anonymous:2013:FCA

[Ano13a] Anonymous. Fibonacci conference announcement. Fibonacci Quarterly, 51(3):288–??, August 2013. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
Announcement2014back.pdf.

Anonymous:2013:VI

[Ano13b] Anonymous. Volume index. Fibonacci Quarterly, 51(4):381–??, November 2013. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/VolumeIndex2013.pdf.

Anonymous:2014:CP

[Ano14a] Anonymous. Cover page. Fibonacci Quarterly, 52(5):i–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Papers1/52-5/introduction.pdf.

Anonymous:2014:FCA

[Ano14b] Anonymous. Fibonacci conference announcement. Fibonacci Quarterly, 52(1):2–??, February 2014. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
Announcement2014.pdf.

Anonymous:2014:F

[Ano14c] Anonymous. Foreword. Fibonacci Quarterly, 52(5):v–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Papers1/52-5/contents.pdf.

Anonymous:2014:I

[Ano14d] Anonymous. Introduction. Fibonacci Quarterly, 52(5):iii–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Papers1/52-5/foreword.pdf.

Anonymous:2014:VI

[Ano14e] Anonymous. Volume index. Fibonacci Quarterly, 52(4):381–??, November 2014. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/VolumeIndex2014.pdf.

Anonymous:2015:JRP

[Ano15a] Anonymous. The John Riordan Prize. Fibonacci Quarterly, 53 (1):96–??, February 2015. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Announcements/Riordan6.pdf.

Anonymous:2015:VI

[Ano15b] Anonymous. Volume index. Fibonacci Quarterly, 53(4):378–??, November 2015. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/VolumeIndex2015.pdf.

Anonymous:2016:BRM

[Ano16a] Anonymous. Book review: The Magic of Math: Solving for x and Figuring out Why. Fibonacci Quarterly, 54(1):94–??, February 2016. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Announcements/a-b-review.pdf.

Anonymous:2016:PBPa

[Ano16b] Anonymous. The Paul Bruckman Prize. Fibonacci Quarterly, 54(3):288–??, August 2016. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprize16.pdf.

Anonymous:2016:PBPb

[Ano16c] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 54(4):380–??, November 2016. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeNov2016.pdf.

Anonymous:2016:TW

[Ano16d] Anonymous. Thanks and welcome. Fibonacci Quarterly, 54(4):290–??, November 2016. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
ThanksWelcomeNov2016.pdf.

Anonymous:2016:VI

[Ano16e] Anonymous. Volume index. Fibonacci Quarterly, 54(4):381–??, November 2016. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/VolumeIndex2016.pdf.

Anonymous:2016:WPB

[Ano16f] Anonymous. Winners of the Paul Bruckman Prize. Fibonacci Quarterly, 54(3):287–??, August 2016. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
BruckmanWinners2016.pdf.

Anonymous:2017:CR

[Ano17a] Anonymous. Conference report. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Papers1/55-5/Foreword.pdf.

Melham:2017:CFCa

[Ano17b] Anonymous. Fibonacci Conference and Caen Proceedings announcements. Fibonacci Quarterly, 55(3):194–??, August 2017.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Announcements/ConfAnnAug2017.pdf.

Anonymous:2017:FCC

[Ano17c] Anonymous. Fibonacci Conference and Caen Proceedings announcements. Fibonacci Quarterly, 55(4):290–??, November 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Announcements/ConfAnnNov2017.pdf.

Anonymous:2017:FCA

[Ano17d] Anonymous. Fibonacci Conference announcement. Fibonacci Quarterly, 55(2):98–??, May 2017. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
ConfAnnMay2017.pdf.

Anonymous:2017:F

[Ano17e] Anonymous. Foreword. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers1/55-5/Participants.pdf.

Anonymous:2017:LP

[Ano17f] Anonymous. List of participants. Fibonacci Quarterly, 55(5): ??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Papers1/55-5/Contents.pdf.

Anonymous:2017:PBPa

[Ano17g] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 55(1): 96–??, February 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Announcements/bruckmanprize.pdf.

Anonymous:2017:PBPb

[Ano17h] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 55(2): 192–??, May 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Announcements/bruckmanprizeMay2017.
pdf.

Luca:2017:APSd

[Ano17i] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly,

55(3):288–??, August 2017.  CODEN FIBQAU.  ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprize17.pdf.

Anonymous:2017:PBP

[Ano17j] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 55(4):380–??, November 2017. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeNov2017.pdf.

Anonymous:2017:TP

[Ano17k] Anonymous. Title page. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers1/55-5/Report.pdf.

Anonymous:2017:VI

[Ano17l] Anonymous. Volume index. Fibonacci Quarterly, 55(4):381–??, November 2017. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/VolumeIndex2017.pdf.

Anonymous:2018:FCC

[Ano18a] Anonymous. Fibonacci Conference and Caen Proceedings announcements. Fibonacci Quarterly, 56(1):2–??, February 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Announcements/ConfAnnFeb2018.pdf.

Anonymous:2018:FCF

[Ano18b]  Anonymous.  Fibonacci Conference and final announce-

ment. Fibonacci Quarterly, 56(2):98–??, May 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Announcements/ConfAnnMay2018.pdf.

Anonymous:2018:PBPa

[Ano18c] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 56(1):96–??, February 2018. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeFeb2018.pdf.

Anonymous:2018:PBPb

[Ano18d] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 56(2): 192–??, May 2018. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Announcements/bruckmanprize2018.
pdf.

Anonymous:2018:PBPc

[Ano18e] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly,

56(3):288–??, August 2018.  CODEN FIBQAU.  ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeAug2018.pdf.

Anonymous:2018:PBPd

[Ano18f] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 56(4):380–??, November 2018. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeNov2018.pdf.

Anonymous:2018:VI

[Ano18g] Anonymous. Volume index. Fibonacci Quarterly, 56(4):381–??, November 2018. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/VolumeIndex2018.pdf.

Anonymous:2019:FCAa

[Ano19a] Anonymous. Fibonacci Conference announcement. Fibonacci Quarterly, 57(1):2–??, February 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
ConfAnnFeb2019.pdf.

Anonymous:2019:FCAb

[Ano19b] Anonymous. Fibonacci Conference announcement. Fibonacci Quarterly, 57(2):98–??, May 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
ConfAnnMay2019.pdf.

Anonymous:2019:FCAc

[Ano19c] Anonymous. Fibonacci conference announcement. Fibonacci Quarterly, 57(3):194–??, August 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
ConfAnnAug2019.pdf.

Anonymous:2019:FCAd

[Ano19d]  Anonymous.  Fibonacci Conference announcement.  Fi-

bonacci Quarterly, 57(4):290–??, November 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Announcements/ConfAnnNov2019.pdf.

Anonymous:2019:PBPa

[Ano19e] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 57(1):96–??, February 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeFeb2019.pdf.

Anonymous:2019:PBPb

[Ano19f] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 57(2): 192–??, May 2019. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Announcements/bruckmanprize2019.
pdf.

Anonymous:2019:PBPc

[Ano19g] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly,

57(3):288–??, August 2019.  CODEN FIBQAU.  ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeAug2019.pdf.

Anonymous:2019:PBPd

[Ano19h] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 57(4):380–??, November 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeNov2019.pdf.

Anonymous:2019:VI

[Ano19i] Anonymous. Volume index. Fibonacci Quarterly, 57(4):381–??, November 2019. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/VolumeIndex2019.pdf.

Anonymous:2020:FCAa

[Ano20a] Anonymous. Fibonacci conference announcement. Fibonacci Quarterly, 58(1):2–??, February 2020. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
ConfAnnFeb2020Rev.pdf.

Anonymous:2020:FCAb

[Ano20b] Anonymous. Fibonacci conference announcement. Fibonacci Quarterly, 58(2):98–??, May 2020. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
ConfAnnMay2020.pdf.

Anonymous:2020:PBPa

[Ano20c] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 58(1):96–??, February 2020. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeFeb2020.pdf.

Anonymous:2020:PBPb

[Ano20d] Anonymous. The Paul Bruckman Prize. Fibonacci Quarterly, 58(2):192–??, May 2020. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeMay2020.pdf.

Anonymous:2020:PBPc

[Ano20e] Anonymous. The Paul Bruckman Prize. Fibonacci Quarterly, 58(3):288–??, August 2020. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeAug2020.pdf.

Anonymous:2020:PBPd

[Ano20f] Anonymous. Paul Bruckman Prize. Fibonacci Quarterly, 58(4):380–??, November 2020. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
bruckmanprizeNov2020.pdf.

Anonymous:2020:VI

[Ano20g] Anonymous. Volume index. Fibonacci Quarterly, 58(4):381–??, November 2020. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/VolumeIndex2020.pdf.

Anonymous:2021:FCAa

[Ano21a] Anonymous. Fibonacci conference announcement. Fibonacci Quarterly, 59(1):3–??, February 2021. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
ConfAnnFeb2021.pdf.

Anonymous:2021:FCAb

[Ano21b] Anonymous. Fibonacci conference announcement. Fibonacci Quarterly, 59(4):290–??, November 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Announcements/
ConfAnn2022.pdf.

Anonymous:2021:MGB

[Ano21c] Anonymous. In memoriam — Gerald E. Bergum. Fibonacci Quarterly, 59(1):2–??, February 2021. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
gbergum.pdf.

Anonymous:2021:VI

[Ano21d] Anonymous. Volume index. Fibonacci Quarterly, 59(4):381–??, November 2021. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/VolumeIndex2021.pdf.

Anonymous:2022:FCAa

[Ano22a] Anonymous. Fibonacci conference announcement. Fibonacci Quarterly, 60(1):2–??, February 2022. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
ConfAnn2022.pdf.

Anonymous:2022:FCAb

[Ano22b] Anonymous. Fibonacci conference announcement. Fibonacci Quarterly, 60(2):98–??, May 2022. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
ConfAnn2022-May.pdf.

Anonymous:2022:LST

[Ano22c] Anonymous. List of speakers and talks. Fibonacci Quarterly, 60 (5):2–??, December 2022. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers1/60-5/talks.pdf.

Anonymous:2022:MBS

[Ano22d] Anonymous. Mathematics books for sale (as a fundraiser). Fibonacci Quarterly, 60(5):403–??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers1/60-5/booksale.pdf.

Anonymous:2023:VI

[Ano23] Anonymous. Volume index. Fibonacci Quarterly, 61(4):382–??, November 2023. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Volume_Index/VolumeIndex2023.pdf.

Anonymous:2024:FCA

[Ano24] Anonymous. Fibonacci conference announcement. Fibonacci Quarterly, 62(1):2–??, February 2024. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Announcements/
ConfAnn2024-Feb.pdf.

Antoniadis:1985:FLN

[Ant85a] Jannis A. Antoniadis. Fibonacci and Lucas numbers of the form 3z^{2} + 1. Fibonacci Quarterly, 23(4):300–307, November 1985.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-4/antoniadis.pdf.

Antoniadis:1985:GFN

[Ant85b] Jannis A. Antoniadis. Generalized Fibonacci numbers and some Diophantine equations. Fibonacci Quarterly, 23(3):199–213, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/23-3/antoniadis.pdf.

Antonini:1991:NUD

[Ant91] Rita Giuliano Antonini. On the notion of uniform distribution mod 1. Fibonacci Quarterly, 29(3):230–234, August 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-3/antonini.pdf.

Arkin:1970:RF

[AP70] Joseph Arkin and Richard Pollack. Recurrence formulas. Fibonacci Quarterly, 8(1):4–5, February 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-1/
arkin.pdf.

Andrade:1992:ONR

[AP92] Ana Andrade and S. P. Pethe. On the r th-order nonhomogeneous recurrence relation and some generalized Fibonacci sequences. Fibonacci Quarterly, 30(3):256–262, August 1992. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/30-3/andrade.pdf.

Aursukaree:2021:EDP

[AP21] Saralee Aursukaree and Prapanpong Pongsriiam. On exactly 3-deficient-perfect numbers. Fibonacci Quarterly, 59(1):33–??, February 2021. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/59-1/aursukaree.pdf.

Arce:1984:FRS

[Arc84] Gonzalo R. Arce. Fibonacci and related sequences in digital filtering. Fibonacci Quarterly, 22(3):208–217, August 1984. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-3/arce.pdf.

Ardila:2004:CFP

[Ard04] Federico Ardila. The coefficients of a Fibonacci power series. Fibonacci Quarterly, 42(3):202–204, August 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-3/quartardila03_2004.pdf.

Arkin:1965:EFN

[Ark65a] Joseph Arkin. An extension of the Fibonacci numbers. American Mathematical Monthly, 72(3):275–279, March 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Arkin:1965:LNA

[Ark65b] Joseph Arkin. Ladder network analysis using polynomials. Fibonacci Quarterly, 3(2):139–142, April 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-2/
arkin.pdf.

Arkin:1966:NTJ

[Ark66] Joseph Arkin. A note on a theorem of Jacobi. Fibonacci Quarterly, 4(4):359–362, December 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-4/
arkin.pdf.

Arkin:1969:CCR

[Ark69] Joseph Arkin. Convergence of the coefficients in a recurring power series. Fibonacci Quarterly, 7(1):41–55, February 1969. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-1/arkin.pdf.

Arkin:1970:CAP

[Ark70a] Joseph Arkin. Certain arithmetical properties of 0.5k(ak + 1). Fibonacci Quarterly, 8(5):531–537, December 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/8-5/arkin.pdf.

Arkin:1970:ETE

[Ark70b] Joseph Arkin. An extension of a theorem of Euler. Fibonacci Quarterly, 8(4):421–427, October 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-4/
arkin.pdf.

Arkin:1973:CCP

[Ark73a] Joseph Arkin. Convergence of the coefficients in the k th power of a power series. Fibonacci Quarterly, 11(1):15–24, February 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/11-1/arkin.pdf.

Arkin:1973:FSC

[Ark73b] Joseph Arkin. The first solution of the classical Eulerian magic cube problem of order ten. Fibonacci Quarterly, 11(2):174–178, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/11-2/arkin.pdf.

Arkin:1973:SCP

[Ark73c] Joseph Arkin. A solution to the classical problem of finding systems of three mutually orthogonal numbers in a cube formed by three superimposed 10×10×10 cubes. Fibonacci Quarterly, 11(5): 485–489, December 1973. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/11-5/arkin-a.pdf.

Arkin:1974:SOT

[Ark74] Joseph Arkin. A solution of orthogonal triples in four 10×10×10 superimposed Latin cubes. Fibonacci Quarterly, 12(2):133–140, April 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/12-2/arkin.pdf.

Arnault:1995:CCN

[Arn95] François Arnault. Constructing Carmichael numbers which are strong pseudoprimes to several bases. Journal of Symbolic Computation, 20(2):151–162 (or 151–161??), August 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Arnault:1997:RMT

[Arn97] F. Arnault. The Rabin–Monier theorem for Lucas pseudoprimes. Mathematics of Computation, 66(218):869–881, April 1997. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/jourcgi/jour-pbprocess?fn=110&
arg1=S0025-5718-97-00836-3&u=/mcom/1997-66-218/.

Arpaia:1994:GSS

[Arp94] Pasquale J. Arpaia. Generating solutions for a special class of Diophantine equations. Fibonacci Quarterly, 32(2):170–173, May 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/32-2/arpaia.pdf.

Arpaia:2004:GSS

[Arp04] Pasquale J. Arpaia. Generating solutions for a special class of Diophantine equations II. Fibonacci Quarterly, 42(1):36–37, February 2004. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Papers1/42-1/quartarpaia01_2004.pdf.

Arrighi:1969:TGV

[Arr69] Gino Arrighi. Il trattato di geometria e la volgarizzazione del Liber quadratorum di Leonardo Pisano del Codice Palatino 577 (sec XV) della Biblioteca Nazionale di Firenze. (Italian) [the treatise on geometry and the vulgarization of the Liber quadratorum of Leonardo Pisano in the Palatino Code 577 (XV century) the National Library of Florence]. Atti Fond. Giorgio Ronchi Contrib. Ist. Naz. Ottica, 22:760–775, 1969.

Article:2004:FNR

[Art04] Featured Article. Fibonacci numbers and the Rogers–Ramanujan identities. Fibonacci Quarterly, 42(1):3–27, February 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-1/quartgranville01_2004.pdf.

Aho:1973:SDE

[AS73] A. V. Aho and N. J. A. Sloane. Some doubly exponential sequences. Fibonacci Quarterly, 11(4):429–437, November 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-4/aho-a.pdf.

Arkin:1974:LC

[AS74] Joseph Arkin and E. G. Straus. Latin k-cubes. Fibonacci Quarterly, 12(3):288–291, October 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
3/arkin.pdf.

Arkin:1976:TMS

[AS76] Joseph Arkin and Paul Smith. Trebly-magic systems in a Latin 3-cube of order eight. Fibonacci Quarterly, 14(2):167–170, April 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/14-2/arkin.pdf.

Alladi:1977:PCF

[AS77] Krishnaswami Alladi and A. G. Shannon. On a property of consecutive Farey–Fibonacci fractions. Fibonacci Quarterly, 15(2): 153–154, April 1977. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/15-2/alladi.pdf.

Arkin:1981:NTM

[AS81a] J. Arkin and K. Singh. A new type magic Latin 3-cube of order ten. Fibonacci Quarterly, 19(1):76–81, February 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-1/arkin.pdf.

Arkin:1981:OLS

[AS81b] Joseph Arkin and E. G. Straus. Orthogonal Latin systems. Fibonacci Quarterly, 19(4):289–292, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-4/arkin.pdf.

Al-Salam:1984:SOF

[AS84] Nadhla A. Al-Salam. Some operational formulas for the q-Laguerre polynomials. Fibonacci Quarterly, 22(2):166–170, May 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/22-2/alsalam.pdf.

Ando:2000:ACT

[AS00] Shiro Ando and Daihachiro Sato. On p-adic complementary theorems between Pascal’s triangle and the modified Pascal triangle. Fibonacci Quarterly, 38(3):194–200, June 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/38-3/ando.pdf.

Ando:2005:GLD

[AS05] Shiro Ando and Daihachiro Sato. On GCD–LCM duality between Pascal’s pyramid and the modified Pascal pyramid. Fibonacci Quarterly, 43(1):15–21, February 2005. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/43-1/
paper43-1-2.pdf.

Austin:2020:GFS

[AS20] Jathan Austin and Lisa Schneider. Generalized Fibonacci sequences in Pythagorean triple preserving matrices. Fibonacci Quarterly, 58(4):340–??, November 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/58-
4/austin.pdf.

Ascher:1974:CI

[Asc74] Marcia Ascher. A combinatorial identity. Fibonacci Quarterly, 12(2):186–188, April 1974. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/12-2/ascher.pdf.

Al-Shaghay:2015:CPS

[ASD15] Abdullah Al-Shaghay and Karl Dilcher. Congruences for partial sums of reciprocals. Fibonacci Quarterly, 53(2):98–??, May 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/53-2/alshaghay.pdf.

Al-Salam:1971:FNE

[ASV71] W. A. Al-Salam and A. Verma. Fibonacci numbers and Eulerian polynomials. Fibonacci Quarterly, 9(1):18–22, February 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-1/alsalam.pdf.

Asveld:1987:AFF

[Asv87a] Peter R. J. Asveld. Another family of Fibonacci-like sequences. Fibonacci Quarterly, 25(4):361–364, November 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-4/asveld.pdf.

Asveld:1987:FFL

[Asv87b] Peter R. J. Asveld. A family of Fibonacci-like sequences. Fibonacci Quarterly, 25(1):81–83, February 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-1/asveld.pdf.

Asveld:1989:FLD

[Asv89] Peter R. J. Asveld. Fibonacci-like differential equations with a polynomial nonhomogeneous part. Fibonacci Quarterly, 27(4): 303–309, August 1989. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/27-4/asveld.pdf.

Atanassov:1986:SNG

[Ata86] Krassimir T. Atanassov. On a second new generalization of the Fibonacci sequence. Fibonacci Quarterly, 24(4):362–365, November 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/24-4/atanassov.pdf.

Atanassov:1989:GFS

[Ata89] Krassimir T. Atanassov. On a generalization of the Fibonacci sequence in the case of three sequences. Fibonacci Quarterly, 27 (1):7–10, February 1989. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/27-1/atanassov.pdf.

Atanassov:1995:RND

[Ata95] Krassimir T. Atanassov. Remark on a new direction for a generalization of the Fibonacci sequence. Fibonacci Quarterly, 33(3): 249–250, June 1995. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/33-3/atanassov.pdf.

Ahlbach:2013:EAZ

[AUFP13] Connor Ahlbach, Jeremy Usatine, Christiane Frougny, and Nicholas Pippenger. Efficient algorithms for Zeckendorf arithmetic. Fibonacci Quarterly, 51(3):249–??, August 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/51-3/ahlbach.pdf.

Austin:2023:PTP

[Aus23] Jathan Austin. On Pythagorean triple preserving matrices that contain Fibonacci numbers. Fibonacci Quarterly, 61(4):321–??, November 2023. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/61-4/austin.pdf.

Aissani:2003:MCS

[AV03] Djamil Aissani and Dominique Valerian. Mathématiques, commerce et société à Béjaïa (Bugia) au moment du séjour de Leonardo Fibonacci XII^{e}–XIII^{e} siècles. Bollettino di Storia delle Scienze Matematiche, 23(2):9–31 (2005), 2003. ISSN 0392-4432 (print), 1724-1650 (electronic). Leonardo Fibonacci. Matematica e societa nel Mediterraneo nel secolo XIII. Vol. I.

Avart:2011:CCD

[Ava11] Christian Avart. A characterization of converging Ducci sequences over Z_{2}. Fibonacci Quarterly, 49(2):155–157, May 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/49-2/avart.pdf.

Avart:2012:RDS

[Ava12] Christian Avart. Reversing Ducci sequences. Fibonacci Quarterly, 50(3):265–??, August 2012. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/50-3/avart.pdf.

Avart:2013:RAC

[Ava13] Christian Avart. A result about cycles in Ducci sequences. Fibonacci Quarterly, 51(2):137–??, May 2013. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/51-
2/avart.pdf.

Avriel:1966:OPS

[AW66] Mordecai Avriel and Douglass J. Wilde. Optimality proof for the symmetric Fibonacci search technique. Fibonacci Quarterly, 4 (3):265–269, October 1966. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/4-3/avriel.pdf.

Avdispahic:2020:ISD

[AZ20] Muharem Avdispahić and Faruk Zejnulahi. An integer sequence with a divisibility property. Fibonacci Quarterly, 58(4):321–??, November 2020. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/58-4/avdispahic.pdf.

Blau:1996:VPA

[B^{+}96] Brian Blau et al., editors. Visual proceedings: the art and interdisciplinary programs of SIGGRAPH 96: SIGGRAPH 96, August 4–9, 1996, New Orleans, LA, Computer Graphics. ACM Press, New York, NY 10036, USA, 1996. ISBN 0-89791-784-7. ISSN 1069-5419. LCCN T385 .S54 1996b. URL http://info.

acm.org/pubs/contents/proceedings/graph/.

Bruckman:1998:CDF

[BA98] Paul S. Bruckman and Peter G. Anderson. Conjectures on the Zdensities of the Fibonacci sequence. Fibonacci Quarterly, 36(3): 263–271, June 1998. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/36-3/bruckman.pdf.

Babini:1941:LSL

[Bab41] José Babini. On the linear systems of Leonardo Pisano. Archeion, 23:57–61, 1941.

Babb:1976:OST

[Bab76] Robert Gordon Babb II. On the order of systems of two simultaneous linear difference equations in two variables. Fibonacci Quarterly, 14(1):78–82, February 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
1/babb.pdf.

Backstrom:1966:DZF

[Bac66] Robert P. Backstrom. On the determination of the zeros of the Fibonacci sequence. Fibonacci Quarterly, 4(4):313–322, December 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/4-4/backstrom.pdf.

Backstrom:1981:RSR

[Bac81] Robert P. Backstrom. On reciprocal series related to Fibonacci numbers with subscripts in arithmetic progression. Fibonacci Quarterly, 19(1):14–20, February 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
1/backstrom.pdf.

Baica:1983:DFN

[Bai83] Malvina Baica. n-dimensional Fibonacci numbers and their applications. Fibonacci Quarterly, 21(4):285–301, November 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-4/baica.pdf.

Bailey:1992:MBC

[Bai92] D. F. Bailey. More binomial coefficient congruences. Fibonacci Quarterly, 30(2):121–125, May 1992. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/30-2/
bailey.pdf.

Balasubramanian:1990:FPA

[Bal90] N. Balasubramanian. On friendly-pairs of arithmetic functions. Fibonacci Quarterly, 28(1):43–47, February 1990. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/28-1/balasubramanian.pdf.

Balakrishnan:1994:SR

[Bal94] U. Balakrishnan. Some remarks on σ(θ(n)). Fibonacci Quarterly, 32(4):293–296, August 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-
4/balakrishnan.pdf.

Ballot:2013:ZBB

[Bal13] Christian Ballot. On Zeckendorf and base b digit sums. Fibonacci Quarterly, 51(4):319–??, November 2013. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/51-
4/ballot.pdf.

Ballot:2015:DFL

[Bal15] Christian Ballot. Divisibility of Fibonomials and Lucasnomials via a general Kummer rule. Fibonacci Quarterly, 53(3):194–??, August 2015. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/53-3/ballot.pdf.

Ballot:2017:DML

[Bal17] Christian Ballot. Divisibility of the middle Lucasnomial coefficient. Fibonacci Quarterly, 55(4):297–??, November 2017. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-4/ballot.pdf.

Ballot:2019:EPA

[Bal19a] Christian Ballot. Errata — The p-adic Valuation of Lucas Sequences When p Is a Special Prime. Fibonacci Quarterly, 57(4): 366–??, November 2019. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers/57-4/errata09212019.

pdf. See [Bal19b].

Ballot:2019:AVL

[Bal19b] Christian Ballot. The p-adic valuation of Lucas sequences when p is a special prime. Fibonacci Quarterly, 57(3):265–??, August 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/57-3/ballot.pdf. See errata [Bal19a].

Bankoff:1976:FC

[Ban76a] Leon Bankoff. A Fibonacci curiosity. Fibonacci Quarterly, 14(1): 17–??, February 1976. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/14-1/bankoff1.pdf.

Bankoff:1976:FP

[Ban76b] Leon Bankoff. A Fibonacci pleasantry. Fibonacci Quarterly, 14(1):29–??, February 1976. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/14-1/bankoff2-
a.pdf.

Brousseau:1973:CPU

[BAP73] Brother Alfred Brousseau, Harold Andersen, and Jerome Povse. A curious property of unit fractions of the form 1/d where (d, 10) = 1. Fibonacci Quarterly, 11(1):91–98, February 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-1/brousseau.pdf.

Barley:1971:TF

[Bar71] W. C. Barley. Triangles de Fibonacci. Fibonacci Quarterly, 9 (4):413–422, October 1971. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/9-4/barley.pdf.

Basin:1963:EPSa

[Bas63a] S. L. Basin. Elementary problems and solutions. Fibonacci Quarterly, 1(1):73–??, February 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-1/
elementary1-1.pdf.

Basin:1963:EPSb

[Bas63b] S. L. Basin. Elementary problems and solutions. Fibonacci Quarterly, 1(2):85–??, April 1963. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/1-2/elementary1-2.
pdf.

Basin:1963:EPSc

[Bas63c] S. L. Basin. Elementary problems and solutions. Fibonacci Quarterly, 1(3):75–??, October 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-3/
elementary1-3-a.pdf.

Basin:1963:EPSd

[Bas63d] S. L. Basin. Elementary problems and solutions. Fibonacci Quarterly, 1(4):73–??, December 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-4/
elementary1-4.pdf.

Basin:1963:FSI

[Bas63e] S. L. Basin. The Fibonacci sequence as it appears in nature. Fibonacci Quarterly, 1(1):53–56, February 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-1/
basin.pdf.

Basin:1963:GFS

[Bas63f] S. L. Basin. Generalized Fibonacci sequences and squared rectangles. American Mathematical Monthly, 70(4):372–379, April 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Basin:1964:NWF

[Bas64] S. L. Basin. A note on Waring’s formula for sums of like powers of roots. Fibonacci Quarterly, 2(2):119–121, April 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-2/basin.pdf.

Basmakova:1978:LQL

[Baš78a] I. G. Bašmakova. The Liber quadratorum of Leonardo of a Pisa. In History and methodology of the natural sciences, No. XX (Russian), pages 27–37. Moskov. Gos. Univ., Moscow, USSR, 1978.

Basmakova:1978:BQL

[Baš78b] I. G. Bašmakova. The Liber quadratorum of Leonardo of Pisa. (Russian). History and methodology of the natural sciences, XX (??):27–37, ???? 1978.

Bastida:1981:DCF

[Bas81] Julio R. Bastida. The determination of certain fields of invariants. Fibonacci Quarterly, 19(2):147–148, April 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-2/bastida2.pdf.

Bhargava:1993:TST

[BAS93] S. Bhargava, Chandrashekar Adiga, and D. D. Somashekara. Three-square theorem as an application of Andrews’ identity. Fibonacci Quarterly, 31(2):129–132, May 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-2/
bhargava.pdf.

Bastasz:2020:LWS

[Bas20] Bob Bastasz. Lyndon words of a second-order recurrence. Fibonacci Quarterly, 58(5):25–??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/58-5/bastasz.pdf.

Baumgart:1964:LE

[Bau64] B. G. Baumgart. Letter to the editor. Fibonacci Quarterly, 2 (4):260–??, December 1964. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/2-4/baumgart-a.pdf.

Bauer:1971:NBT

[Bau71] R. L. Bauer. Numbers that are both triangular and square their triangular roots and square roots. Fibonacci Quarterly, 9(2):196– 198, April 1971. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/9-2/bauer.pdf.

Baudert:1980:AP

[Bau80] F. R. Baudert. The Apollonius problem. Fibonacci Quarterly, 18 (1):33–??, February 1980. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/18-1/baudert.pdf.

Bayless:2008:LPP

[Bay08] Jonathan Bayless. The Lucas–Pratt primality tree. Mathematics of Computation, 77(261):495–502, January 2008. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/mcom/2008-77-261/S0025-5718-
07-02002-9/home.html;   http://www.ams.org/mcom/2008-
77-261/S0025-5718-07-02002-9/S0025-5718-07-02002-9.dvi;
 http://www.ams.org/mcom/2008-77-261/S0025-5718-07-02002-
9/S0025-5718-07-02002-9.pdf;   http://www.ams.org/mcom/
2008-77-261/S0025-5718-07-02002-9/S0025-5718-07-02002-
9.ps.

Bridger:1965:CFC

[BB65] Clyde A. Bridger and Marjorie Bicknell. Continued fraction convergents as a source of Fibonacci and Lucas identities. Fibonacci Quarterly, 3(4):304–308, December 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-4/
bridger.pdf.

Barakat:1969:TPM

[BB69] Richard Barakat and Elizabeth Baumann. M th power of an N ×N matrix and its connection with the generalized Lucas polynomials. Journal of Mathematical Physics, 10(8):1474–1476, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v10/i8/p1474_s1.

Bauer:1979:PCI

[BB79] Friedrich Ludwig Bauer and M. Broy, editors. Program construction: international summer school, volume 69 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. CO-DEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .P75117. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t0069.htm;   http://www.springerlink.com/content/978-
0-387-09251-5;   http://www.springerlink.com/openurl.

asp?genre=issue&issn=0302-9743&volume=69. “The interna-

tional summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Bange:1983:FGG

[BB83] David W. Bange and Anthony E. Barkauskas. Fibonacci graceful graphs. Fibonacci Quarterly, 21(3):174–188, August 1983. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-3/bange.pdf.

Barry:1985:FLT

[BB85] Arthur Barry and Stanley Bezuszka, S. J. A Fibonacci and Lucas Tannenbaum. Fibonacci Quarterly, 23(4):369–370, November 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/23-4/barry.pdf.

Bollinger:1990:LTS

[BB90] Richard C. Bollinger and Charles L. Burchard. Lucas’s theorem and some related results for extended Pascal triangles. American Mathematical Monthly, 97(3):198–204, March 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Berrizbeitia:2004:BRL

[BB04] Pedro Berrizbeitia and T. G. Berry. Biquadratic reciprocity and a Lucasian primality test. Mathematics of Computation, 73 (247):1559–1564, July 2004. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.

org/mcom/2004-73-247/S0025-5718-03-01575-8/home.html;
 http://www.ams.org/mcom/2004-73-247/S0025-5718-03-01575-
8/S0025-5718-03-01575-8.dvi;   http://www.ams.org/mcom/
2004-73-247/S0025-5718-03-01575-8/S0025-5718-03-01575-
8.pdf;   http://www.ams.org/mcom/2004-73-247/S0025-5718-
03-01575-8/S0025-5718-03-01575-8.ps;   http://www.ams.
org/mcom/2004-73-247/S0025-5718-03-01575-8/S0025-5718-
03-01575-8.tex.

Brown:2008:FFN

[BB08] Ezra Brown and Jason C. Brunson. Fibonacci’s forgotten number. College Mathematics Journal, 39(2):112–120, March 2008. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2008.11922284.

Bundschuh:2011:DFL

[BB11] Ralf Bundschuh and Peter Bundschuh. Distribution of Fibonacci and Lucas numbers modulo 3^{k}. Fibonacci Quarterly, 49(3):201– 210, August 2011. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/49-3/bundschuh.pdf.

BenTaher:2019:SLD

[BB19] R. Ben Taher and H. Benkhaldoun. Solving the linear difference equation with periodic coefficients via Fibonacci sequences. Linear Multilinear Algebra, 67(12):2549–2564, 2019. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Barcucci:2004:TS

[BBB04] Elena Barcucci, Luc Bélanger, and Srečko Brlek. On Tribonacci sequences. Fibonacci Quarterly, 42(4):314–319, November 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-4/quartbarcucci04_2004.pdf.

Bradley:2006:GAT

[BBB06] Sean Bradley, Patrick Brewer, and Christopher Brazfield. Generalized arithmetic triangles via convolution. Fibonacci Quarterly, 44(1):13–19, February 2006. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/44-1/bradley.pdf.

Beckwith:2013:AGD

[BBG^{+}13] Olivia Beckwith, Amanda Bower, Louis Gaudet, Rachel Insoft, Shiyu Li, Steven J. Miller, and Philip Tosteson. The average gap distribution for generalized Zeckendorf decompositions. Fibonacci Quarterly, 51(1):13–??, February 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/51-1/beckwith.pdf.

Bergum:1985:JPC

[BBHM85] G. E. Bergum, Larry Bennett, A. F. Horadam, and S. D. Moore. Jacobsthal polynomials and a conjecture concerning Fibonaccilike matrices. Fibonacci Quarterly, 23(3):240–248, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-3/bergum.pdf.

Bleichenbacher:1995:SRL

[BBL95] Daniel Bleichenbacher, Wieb Bosma, and Arjen K. Lenstra. Some remarks on Lucas-based cryptosystems. Lecture Notes in Computer Science, 963:386–396, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer-ny.com/link/service/series/0558/bibs/
0963/09630386.htm;   http://link.springer-ny.com/link/
service/series/0558/papers/0963/09630386.pdf.

Bravo:2014:CGL

[BBL14] Eric F. Bravo, Jhon J. Bravo, and Florian Luca. Coincidences in generalized Lucas sequences. Fibonacci Quarterly, 52(4):296– ??, November 2014. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/52-4/bravo.pdf.

Baron:1985:NST

[BBP^{+}85] G. Baron, F. T. Boesch, H. Prodinger, R. F. Tichy, and J. F. Wang. The number of spanning trees in the square of a cycle. Fibonacci Quarterly, 23(3):258–264, August 1985. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-3/baron.pdf.

Bange:1981:FNT

[BBS81] D. W. Bange, A. E. Barkauskas, and P. J. Slater. Fibonacci numbers in tree counts for maximal outerplane and related graphs. Fibonacci Quarterly, 19(1):28–34, February 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-1/bange.pdf.

Brualdi:1977:FSE

[BC77] R. A. Brualdi and J. Csima. Fibonacci sequence and extremal stochastic matrices. Fibonacci Quarterly, 15(4):333–335, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/15-4/brualdi.pdf.

Back:2010:GPF

[BC10] Greg Back and Mihai Caragiu. The greatest prime factor and recurrent sequences. Fibonacci Quarterly, 48(4):358–362, November 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/48-4/back.pdf.

Borade:2020:GSZ

[BCC^{+}20] Neelima Borade, Dexter Cai, David Z. Chang, Bruce Fang, Alex Liang, Steven J. Miller, and Wanqiao Xu. Gaps of summands of the Zeckendorf lattice. Fibonacci Quarterly, 58(2):143–??, May 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/58-2/borade.pdf.

Boldyriew:2020:EZT

[BCD^{+}20] Elżbieta Bołdyriew, Anna Cusenza, Linglong Dai, Pei Ding, Aidan Dunkelberg, John Haviland, Kate Huffman, Dianhui Ke, Daniel Kleber, Jason kuretski, John Lentfer, Tianhao Luo, Steven J. Miller, Clayton Mizgerd, Vashisth Tiwari, Jingkai Ye, Yunhao Zhang, Xiaoyan Zheng, and Weiduo Zhu. Extending Zeckendorf’s theorem to a non-constant recurrence and the Zeckendorf game on this non-constant recurrence relation. Fibonacci Quarterly, 58(5):55–??, December 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/58-
5/boldyriew1.pdf.

Beanland:2022:GSS

[BCFS22] Kevin Beanland, Hùng Vi{ê \atop .}t Chu, and Carrie E. Finch-Smith. Generalized Schreier sets, linear recurrence relation, and Turán graphs. Fibonacci Quarterly, 60(4):352–??, November 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-4/beanland.pdf.

Bonnin-Cadogan:2008:CFR

[BCFX09] Jose M. Bonnin-Cadogan, Christopher P. French, and Buchan Xue. Continued fractions of roots of Fibonacci-like fractions. Fibonacci Quarterly, 46/47(4):298–311, November 2008/2009. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/46_47-4/bonnin-cadogan.pdf.

Bacon:2012:EFF

[BCH12] Michael R. Bacon, Charles K. Cook, and Russell Jay Hendel. Extending Freitag’s Fibonacci-like magic square to other dimensions. Fibonacci Quarterly, 50(2):119–??, May 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/50-2/bacon.pdf.

Bateman:1991:PCM

[BCHR91] R. A. Bateman, E. A. Clark, M. L. Hancock, and C. A. Reiter. The period of convergents modulo M of reduced quadratic irrationals. Fibonacci Quarterly, 29(3):220–229, August 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-3/bateman.pdf.

Brigham:2005:DFT

[BCJ^{+}05] Robert C. Brigham, Julie R. Carrington, Dal Y. Jeong, Richard P. Vitray, and Jay Yellen. Domination in Fibonacci trees. Fibonacci Quarterly, 43(2):157–165, May 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers/43-2/paper43-2-11.pdf.

Barbero:2010:SPE

[BCM10] Stefano Barbero, Umberto Cerruti, and Nadir Murru. Solving the Pell equation via Rédei rational functions. Fibonacci Quarterly, 48(4):348–357, November 2010. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/48-
4/barbero.pdf.

Balkin:1995:SPC

[BCOR95] S. D. Balkin, D. S. Cousins, C. K. Orr, and C. A. Reiter. Short periods of continued fraction convergents modulo M: a generalization of the Fibonacci case. Fibonacci Quarterly, 33(3):222–233, June 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/33-3/balkin.pdf.

Benjamin:2007:FDC

[BCQ07] Arthur T. Benjamin, Naiomi T. Cameron, and Jennifer J. Quinn. Fibonacci determinants — a combinatorial approach. Fibonacci Quarterly, 45(1):39–55, February 2007. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/45-1/

benjamin.pdf. See errata [BP10b].

Benjamin:2019:UTP

[BCS19] Arthur T. Benjamin, Joshua Crouch, and James A. Sellers. Unified tiling proofs of a family of Fibonacci identities. Fibonacci Quarterly, 57(1):29–??, February 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/57-1/
benjamin.pdf.

Benjamin:2011:CTT

[BCSS11] Arthur T. Benjamin, Phyllis Chinn, Jacob N. Scott, and Greg Simay. Combinations of two-toned tilings. Fibonacci Quarterly, 49(4):290–??, November 2011. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/49-4/
benjamin.pdf.

Brown:1970:MOU

[BD70] J. L. Brown, Jr. and R. L. Duncan. Modulo one uniform distribution of the sequence of logarithms of certain recursive sequences. Fibonacci Quarterly, 8(5):482–486, December 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-5/brown.pdf. See remark [KsS73].

Brown:1971:LRA

[BD71] J. L. Brown, Jr. and R. L. Duncan. The least remainder algorithm. Fibonacci Quarterly, 9(4):347–350, October 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-4/brown-a.pdf.

Brown:1972:MOU

[BD72] J. L. Brown, Jr. and R. L. Duncan. Modulo one uniform distribution of certain Fibonacci-related sequences. Fibonacci Quarterly, 10(3):277–280, April 1972. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/10-3/brown-a.pdf.

Bezuszka:1977:ATN

[BD77] Stanley Bezuszka and Lou D’Angelo. An application of Tribonacci numbers. Fibonacci Quarterly, 15(2):140–144, April 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/15-2/bezuszka.pdf.

Bastida:1981:QPC

[BD81] Julio R. Bastida and M. J. DeLeon. A quadratic property of certain linearly recurrent sequences. Fibonacci Quarterly, 19(2): 144–146, April 1981. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/19-2/bastida1.pdf.

Berg:2011:FNE

[BD11] Christian Berg and Antonio J. Durán. Fibonacci numbers, Euler’s 2-periodic continued fractions and moment sequences. Fibonacci Quarterly, 49(1):66–75, February 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/49-1/berg.pdf.

Boman:2017:WDF

[BDD^{+}17] Bruce M. Boman, Thien-Nam Dinh, Keith Decker, Brooks Emerick, Christopher Raymond, and Gilberto Schleiniger. Why do Fibonacci numbers appear in patterns of growth in nature? Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/55-
5/boman.pdf.

Baily:2022:BGa

[BDD^{+}22] Benjamin Baily, Justine Dell, Irfan Durmić, Henry L. Fleischmann, Faye Jackson, Isaac Mijares, Steven J. Miller, Ethan Pesikoff, Luke Reifenberg, Alicia Smith Reina, and Yingzi Yang. The Bergman game. Fibonacci Quarterly, 60(5):18–??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/60-5/baily.pdf.

Best:2014:BBZ

[BDE^{+}14a] Andrew Best, Patrick Dynes, Xixi Edelsbrunner, Brian McDonald, Steven J. Miller, Kimsy Tor, Caroline Turnage-Butterbaugh, and Madeleine Weinstein. Benford behavior of Zeckendorf decompositions. Fibonacci Quarterly, 52(5):35–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/52-5/Best-Benford.pdf.

Best:2014:GBN

[BDE^{+}14b] Andrew Best, Patrick Dynes, Xixi Edelsbrunner, Brian McDonald, Steven J. Miller, Kimsy Tor, Caroline Turnage-Butterbaugh, and Madeleine Weinstein. Gaussian behavior of the number of summands in Zeckendorf decompositions in small intervals. Fibonacci Quarterly, 52(5):47–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers1/52-5/Best-Gaussian.pdf.

Brodal:2016:TDR

[BDL^{+}16] Gerth Stølting Brodal, Pooya Davoodi, Moshe Lewenstein, Rajeev Raman, and Srinivasa Rao Satti. Two dimensional range minimum queries and Fibonacci lattices. Theoretical Computer Science, 638(??):33–43, July 25, 2016. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397516001389.

Brown:2012:SRT

[BDM12] Jason I. Brown, Karl Dilcher, and Dante V. Manna. Series representations of theta functions in terms of a sequence of polynomials. Fibonacci Quarterly, 50(1):5–??, February 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-1/brown.pdf.

Ballot:2007:RPP

[BE07] Christian Ballot and Michele Elia. Rank and period of primes in the Fibonacci sequence. A trichotomy. Fibonacci Quarterly, 45 (1):56–63, February 2007. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/45-1/ballot.pdf.

Benjamin:2022:CHTa

[BE22] Arthur T. Benjamin and Daniela Elizondo. Counting on Hosoya’s triangle. Fibonacci Quarterly, 60(5):47–??, December 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-5/benjamin.pdf.

Beard:1950:GSF

[Bea50] Robert S. Beard. The golden section and Fibonacci numbers. Scripta Mathematica, 16(??):116–119, ???? 1950. ISSN 0036-9713.

Beard:1964:FDB

[Bea64] Colonel R. S. Beard. Fibonacci drawing board. Fibonacci Quarterly, 2(3):161–162, October 1964. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/2-3/beard.pdf.

Beard:1966:PGS

[Bea66a] R. S. Beard. Powers of the golden section. Fibonacci Quarterly, 4(2):163–167, April 1966. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/4-2/beard.pdf.

Beard:1966:SG

[Bea66b] R. S. Beard. Star geometry. Fibonacci Quarterly, 4(1):70–72, February 1966. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/4-1/beard.pdf.

Beard:1968:FDB

[Bea68] Col. R. S. Beard. The Fibonacci drawing board — design of the Great Pyramid of Gizeh. Fibonacci Quarterly, 6(1):85–87, February 1968. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/6-1/beard.pdf.

Beardon:2003:SDV

[Bea03] Alan F. Beardon. Sums and differences of values of a quadratic polynomial. Fibonacci Quarterly, 41(4):372–373, August 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-4/beardon.pdf.

Beardon:2018:FM

[Bea18] Alan F. Beardon. Fibonacci matrices. Fibonacci Quarterly, 56 (4):363–??, November 2018. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/56-4/beardon.pdf.

Beardon:2022:FNC

[Bea22a] A. F. Beardon. 106.33 Fibonacci numbers and Cassini’s identity. The Mathematical Gazette, 106(567):498–501, November 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10633-fibonacci-numbers-
and-cassinis-identity/5E7A4CD24EDA06A769F51A4989BCAEA8.

Beardon:2022:RGF

[Bea22b] A. F. Beardon. Ratios of generalized Fibonacci numbers. Fibonacci Quarterly, 60(3):235–??, August 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/60-3/beardon.pdf.

Beck:1990:POF

[Bec90] Istvan Beck. Partial orders and the Fibonacci numbers. Fibonacci Quarterly, 28(2):172–174, May 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-2/
beck.pdf.

Becvar:2001:LPF

[Beč01] Jindřich Bečvář. Leonardo of Pisa—Fibonacci. In Mathematics in medieval Europe (Czech) (Jevíčko, 1999), volume 19 of Děj. Mat./Hist. Math., pages 265–339. Prometheus, Prague, 2001.

Bedratyuk:2013:DIF

[Bed13] Leonid Bedratyuk. Derivations and identitites for Fibonacci and Lucas polynomials. Fibonacci Quarterly, 51(4):351–??, November 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/51-4/bedratyuk.pdf.

Beebe:2004:JPF

[Bee04] Nelson H. F. Beebe. Java programming: Fun with Fibonacci. World-Wide Web document., March 2004. URL https://

www.math.utah.edu/pub/tex/bib/fibquart.bib;   https:/
/www.math.utah.edu/~beebe/software/java/fibonacci/index.

html. This report summarizes the origin of the Fibonacci sequence, giving the full Latin text from the original book written in 1202 (not previously available on the Web). Computation of the Fibonacci sequence, and its term ratios, is implemented in about 50 different programming languages. The report comments on the relative difficulty of the task in some of those languages, and on their suitability for numerical computation. It also provides a complete floating-point formatted output package for Java.

Beebe:2011:CBPy

[Bee11] Nelson H. F. Beebe. A complete bibliography of publications in The Fibonacci Quarterly. Technical report, University of Utah, Department of Mathematics, 155 S 1400 E RM 233, Salt Lake City, UT 84112-0090, USA, October 21, 2011. 550

pp. URL https://www.math.utah.edu/pub/tex/bib/index-
table-f.html#fibquart.

Behrend:2012:PKE

[Beh12] Michael Behrend. Proof of Kimberling’s “even second column” conjecture. Fibonacci Quarterly, 50(2):106–??, May 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-2/behrend.pdf.

Beiter:1978:CCP

[Bei78] Marion Beiter. Coefficients of the cyclotomic polynomial F_{3qr}(x). Fibonacci Quarterly, 16(4):302–306, August 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/16-4/beiter.pdf.

Bender:1993:FTL

[Ben93] C. Bender. Fibonacci transmission lines. Fibonacci Quarterly, 31 (3):227–238, August 1993. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/31-3/bender.pdf.

Benyi:2002:OPE

[Ben02] Arpad Benyi. An Olympiad problem, Euler’s sequence, and Stirling’s formula. Fibonacci Quarterly, 40(4):295–298, August 2002.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-4/benyi.pdf.

Benjamin:2006:SAW

[Ben06] Arthur T. Benjamin. Self-avoiding walks and Fibonacci numbers. Fibonacci Quarterly, 44(4):330–334, November 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/44-4/benjamin.pdf.

Berg:1966:PGR

[Ber66] M. Berg. Phi, the golden ratio (to 4599 decimal places), and Fibonacci numbers. Fibonacci Quarterly, 4(2):157–162, April 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-2/berg.pdf.

Bernstein:1968:LDE

[Ber68] Leon Bernstein. The linear Diophantine equation in n variables and its application to generalized Fibonacci numbers. Fibonacci Quarterly, 6(3):1–63, June 1968. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/6-3/bernstein.
pdf.

Bergum:1974:NFP

[Ber74] Gerald E. Bergum. A note on the Fermat–Pellian equation. Fibonacci Quarterly, 12(2):212–??, April 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-2/
bergum.pdf.

Berzsenyi:1975:SPG

[Ber75] George Berzsenyi. Sums of products of generalized Fibonacci numbers. Fibonacci Quarterly, 13(4):343–344, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-4/berzsenyi.pdf.

Bernstein:1976:FFN

[Ber76] Leon Bernstein. A formula for Fibonacci numbers from a new approach to generalized Fibonacci numbers. Fibonacci Quarterly, 14(4):358–367, November 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
4/bernstein.pdf.

Berzsenyi:1977:GFN

[Ber77] George Berzsenyi. Gaussian Fibonacci numbers. Fibonacci Quarterly, 15(3):233–235, October 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
3/berzsenyi.pdf.

Bernstein:1978:ICI

[Ber78] Leon Bernstein. An invariant for combinatorial identities. Fibonacci Quarterly, 16(4):354–369, August 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/16-4/bernstein.pdf.

Bernstein:1982:PPT

[Ber82] Leon Bernstein. Primitive Pythagorean triples. Fibonacci Quarterly, 20(3):227–241, August 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
3/bernstein.pdf.

Bergum:1984:AGG

[Ber84] Gerald E. Bergum. Addenda to geometry of a generalized Simson’s formula. Fibonacci Quarterly, 22(1):22–28, February 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-1/bergum.pdf.

Bergum:1985:LE

[Ber85] Gerald E. Bergum. Letter from the editor. Fibonacci Quarterly, 23(2):180–??, May 1985. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/23-2/letter.pdf.

Bergum:1986:LE

[Ber86] G. E. Bergum. Letter from the editor. Fibonacci Quarterly, 24 (1):46–??, February 1986. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/24-1/letter.pdf.

Bernstein:1989:PPT

[Ber89] Leon Bernstein. On primitive Pythagorean triangles with equal perimeters. Fibonacci Quarterly, 27(1):2–6, February 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-1/bernstein.pdf.

Berthe:1996:FFS

[Ber96] Valérie Berthé. Frequences des facteurs des suites sturmiennes. (French) [Frequencies of factors of Sturmian series]. Theoretical Computer Science, 165(2):295–309, October 10, 1996. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/
store/tcs/cas_sub/browse/browse.cgi?year=1996&volume=
165&issue=2&aid=2165.

Bergart:2011:FWA

[Ber11] Jeff Bergart. A Fibonacci “whack” alongside the head: My journey into the world of book collecting. Fibonacci Quarterly, 49(2): 177–179, May 2011. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Papers/49-2/bergart.pdf.

Beslin:1991:RGM

[Bes91] Scott J. Beslin. Reciprocal GCD matrices and LCM matrices. Fibonacci Quarterly, 29(3):271–274, August 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-3/beslin.pdf.

Beverage:1970:LE

[Bev70a] David G. Beverage. Letter to the editor. Fibonacci Quarterly, 8 (5):498–??, December 1970. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/8-5/letter.pdf.

Beverage:1970:CDT

[Bev70b] David G. Beverage. On a conjecture of Dmitri Thoro. Fibonacci Quarterly, 8(5):475–??, December 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-5/
beverage-a.pdf.

Beverage:1971:PRF

[Bev71] David Beverage. A polynomial representation of Fibonacci numbers. Fibonacci Quarterly, 9(5):541–544, December 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-5/beverage.pdf.

Beverage:1976:PS

[Bev76] David G. Beverage. Polynomials P_{2n+1}(x) satisfying P_{2n+1}(F_{k}) = F_{(2n+1)k}. Fibonacci Quarterly, 14(3):197–200, October 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-3/beverage.pdf.

Beverage:1977:LE

[Bev77] D. Beverage. Letter to the editor. Fibonacci Quarterly, 15(3): 238–??, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/15-3/letter.pdf.

Bartolozzi:1990:TPA

[BF90] Margherita Bartolozzi and Raffaella Franci. The theory of proportions in abacus mathematics from Leonardo Pisano to Luca Pacioli. Bollettino di Storia delle Scienze Matematiche, 10(1): 3–28, 1990. ISSN 0392-4432 (print), 1724-1650 (electronic).

Brown:1991:SSA

[BF91] Tom C. Brown and Allen R. Freedman. Some sequences associated with the golden ratio. Fibonacci Quarterly, 29(2):157–159, May 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/29-2/brown.pdf.

Brugia:1995:PDF

[BF95] Odoardo Brugia and Piero Filipponi. Polynomial divisibility in finite fields, and recurring sequences. Fibonacci Quarterly, 33(5): 459–463, November 1995. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/33-5/brugia.pdf.

Bai:2016:FLR

[BF16] Zai-Qiao Bai and Steven R. Finch. Fibonacci and Lucas representations. Fibonacci Quarterly, 54(4):319–??, November 2016. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/54-4/bai.pdf.

Bonomo:2023:SF

[BF23] John P. Bonomo and Montana Ferita. A small fib. College Mathematics Journal, 54(4):274–289, 2023. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/07468342.2023.
2225397.

Baczkowski:2011:LSL

[BFF11] Daniel Baczkowski, Olaolu Fasoranti, and Carrie E. Finch. Lucas–Sierpiński and Lucas–Riesel numbers. Fibonacci Quarterly, 49(4):334–??, November 2011. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/49-4/
baczkowski.pdf.

Blair:2019:MHT

[BFM19] Matthew Blair, Rigoberto Flórez, and Antara Mukherjee. Matrices in the Hosoya triangle. Fibonacci Quarterly, 57(5):15–??, December 2019. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/57-5/blair.pdf.

Blair:2020:MDH

[BFMR20] Matthew Blair, Rigoberto Flórez, Antara Mukherjee, and José L. Ramírez. Matrices in the determinant Hosoya triangle. Fibonacci Quarterly, 58(5):34–??, December 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/58-
5/blair2.pdf.

Bonardo:2019:NVF

[BFS19] Pierre Bonardo, Anna E. Frid, and Jeffrey Shallit. The number of valid factorizations of Fibonacci prefixes. Theoretical Computer Science, 775(??):68–75, July 5, 2019. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397518307497.

Berg:2015:GSR

[BFT15] Kimmo Berg, János Flesch, and Frank Thuijsman. Golden and silver ratios in bargaining. Fibonacci Quarterly, 53(2):130–??, May 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/53-2/berg.pdf.

Bruckman:1976:GSD

[BG76] P. S. Bruckman and I. J. Good. A generalization of a series of De-Morgan with applications of Fibonacci type. Fibonacci Quarterly, 14(3):193–195, October 1976. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/14-3/bruckman.
pdf.

Bialek:1991:FE

[BG91] Krystyna Bialek and Aleksander Grytczuk. On Fermat’s equation. Fibonacci Quarterly, 29(1):62–65, February 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-1/bialek.pdf.

Bramham:2016:CIS

[BG16] Alex Bramham and Martin Griffiths. Combinatorial interpretations of some convolution identities. Fibonacci Quarterly, 54(4):

335–??, November 2016. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/54-4/bramham.pdf.

Broderius:2021:LSC

[BG21] Mark Broderius and John Greene. Lucas sequences containing few primes. Fibonacci Quarterly, 59(2):136–??, May 2021. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/59-2/broderius.pdf.

Becker:2011:BLC

[BGTM^{+}11] T. Becker, A. Greaves-Tunnell, S. J. Miller, R. Ronan, and F. W. Strauch. Benford’s Law and continuous dependent random variables. ArXiv e-prints, November 2011. CODEN ???? ISSN

???? URL http://adsabs.harvard.edu/abs/2011arXiv1111.
0568B;   http://arxiv.org/abs/1111.0568.

Birmajer:2014:CTF

[BGW14] Daniel Birmajer, Juan B. Gil, and Michael D. Weiner. Convolutions of Tribonacci, Fuss–Catalan, and Motzkin sequences. Fibonacci Quarterly, 52(5):54–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers1/52-5/Birmajer.pdf.

Basin:1963:PFSa

[BH63a] S. L. Basin and Verner E. Hoggatt, Jr. A primer on the Fibonacci sequence, Part I. Fibonacci Quarterly, 1(1):65–72, February 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-1/primer-1.pdf.

Basin:1963:PFSb

[BH63b] S. L. Basin and Verner E. Hoggatt, Jr. A primer on the Fibonacci sequence, Part II. Fibonacci Quarterly, 1(2):61–68, April 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-2/basin.pdf.

Bicknell:1963:FML

[BH63c] Marjorie Bicknell and Verner E. Hoggatt, Jr. Fibonacci matrices and lambda functions. Fibonacci Quarterly, 1(2):47–52, April 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/1-2/bicknell.pdf.

Bruggles:1963:PFS

[BH63d] I. D. Bruggles and V. E. Hoggatt, Jr. A primer on the Fibonacci sequence, Part IV. Fibonacci Quarterly, 1(4):65–71, December 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/1-4/hoggatt.pdf.

Bernstein:1969:EDP

[BH69a] Leon Bernstein and Helmut Hasse. Explicit determination of the Perron matrices in periodic algorithms of the Perron–Jacobi type with application to generalized Fibonacci numbers with time impulses. Fibonacci Quarterly, 7(4):394–436, November 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-4/bernstein.pdf.

Bicknell:1969:GTR

[BH69b] Marjorie Bicknell and Verner E. Hoggatt, Jr. Golden triangles, rectangles, and cuboids. Fibonacci Quarterly, 7(1):73–91, February 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/7-1/bicknell-a.pdf.

Bicknell:1971:PFNb

[BH71] Marjorie Bicknell and Verner E. Hoggatt, Jr. A primer for the Fibonacci numbers: Part IX. Fibonacci Quarterly, 9(5):529–537, December 1971. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/9-5/bicknell1.pdf.

Bicknell:1973:GFP

[BH73a] Marjorie Bicknell and V. E. Hoggatt, Jr. Generalized Fibonacci polynomials. Fibonacci Quarterly, 11(5):457–465, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/11-5/hoggatt.pdf.

Bicknell:1973:UDG

[BH73b] Marjorie Bicknell and V. E. Hoggatt, Jr. Unit determinants in generalized Pascal triangles. Fibonacci Quarterly, 11(2):131–144, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/11-2/bicknell.pdf.

Bergum:1974:ILG

[BH74] G. E. Bergum and V. E. Hoggatt, Jr. Irreducibility of Lucas and generalized Lucas polynomials. Fibonacci Quarterly, 12(1):95– 100, February 1974. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/12-1/bergum.pdf.

Bergum:1975:SPR

[BH75] G. E. Bergum and V. E. Hoggatt, Jr. Sums and products for recurring sequences. Fibonacci Quarterly, 13(2):115–120, April 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/13-2/bergum.pdf.

Bergum:1976:NPC

[BH76] G. E. Bergum and V. E. Hoggatt, Jr. Numerator polynomial coefficient array for the convolved Fibonacci sequence. Fibonacci Quarterly, 14(1):43–47, February 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
1/bergum.pdf.

Bergum:1977:ACG

[BH77a] G. E. Bergum and V. E. Hoggatt, Jr. An application of the characteristic of the generalized Fibonacci sequence. Fibonacci Quarterly, 15(3):215–219, October 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
3/bergum.pdf.

Bergum:1977:LQC

[BH77b] Gerald E. Bergum and Verner E. Hoggatt, Jr. Limits of quotients for the convolved Fibonacci sequence and related sequences. Fibonacci Quarterly, 15(2):113–115, April 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-2/
bergum.pdf.

Bergum:1978:CPI

[BH78a] G. E. Bergum and V. E. Hoggatt, Jr. A combinatorial problem involving recursive sequences and tridiagonal matrices. Fibonacci Quarterly, 16(2):113–117, April 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-2/
bergum.pdf.

Bergum:1978:FTM

[BH78b] Gerald E. Bergum and Verner E. Hoggatt, Jr. A family of tridiagonal matrices. Fibonacci Quarterly, 16(3):285–288, June 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-3/bergum.pdf.

Brown:1978:PFN

[BH78c] J. L. Brown, Jr. and V. E. Hoggatt, Jr. A primer for the Fibonacci numbers, Part XVI, The central column sequence. Fibonacci Quarterly, 16(1):41–46, February 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
1/brown2.pdf.

Bergum:1979:ISF

[BH79] Gerald E. Bergum and Verner E. Hoggatt, Jr. Infinite series with Fibonacci and Lucas polynomials. Fibonacci Quarterly, 17(2): 147–150, April 1979. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/17-2/bergum.pdf.

Babu:1980:NMT

[BH80a] A. G. T. Babu and Wei Shen Hsia. A note on the multiplication of two 3 × 3 Fibonacci-rowed matrices. Fibonacci Quarterly, 18 (1):43–??, February 1980. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/18-1/babu.pdf.

Bergum:1980:SEW

[BH80b] Gerald E. Bergum and Verner E. Hoggatt, Jr. Some extensions of Wythoff pair sequences. Fibonacci Quarterly, 18(1):28–32, February 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/18-1/bergum.pdf.

Biebighauser:2005:FDS

[BH05] Daniel P. Biebighauser and Gerald A. Heuer. Final digit strings of powers where the exponents end in 1, 3, 7 or 9. Fibonacci Quarterly, 43(4):339–350, November 2005. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/43-
4/biebighauser.pdf.

Benjamin:2014:CRF

[BH14a] Arthur T. Benjamin and Curtis R. Heberle. Counting on r-Fibonacci numbers. Fibonacci Quarterly, 52(2):121–??, May 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/52-2/benjamin.pdf.

Bode:2014:STG

[BH14b] Jens-P. Bode and Heiko Harborth. Steinhaus triangles with generalized Pascal addition. Fibonacci Quarterly, 52(5):61–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Papers1/52-5/Bode.pdf.

Bryant:2017:GRB

[BH17] Mark Bryant and David Hobill. Golden-ratio-based rectangular tilings. Fibonacci Quarterly, 55(2):137–??, May 2017. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-2/bryant.pdf.

Bhatnagar:2016:AFL

[Bha16] Gaurav Bhatnagar. Analogues of a Fibonacci–Lucas identity. Fibonacci Quarterly, 54(2):166–??, May 2016. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/54-
2/bhatnagar.pdf.

Bode:2007:CFS

[BHK07] Jens P. Bode, Heiko Harborth, and Clark Kimberling. Complementary Fibonacci sequences. Fibonacci Quarterly, 45(3):254– 264, August 2007. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/45-3/bode.pdf.

Boldyriew:2020:ICP

[BHL^{+}20] Elżbieta Bołdyriew, John Haviland, Phúc Lâm, John Lentfer, Steven J. Miller, and Fernando Trejos Suárez. An introduction to completeness of positive linear recurrence sequences. Fibonacci Quarterly, 58(5):77–??, December 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/58-
5/boldyriew2.pdf.

Bravo:2021:CVG

[BHL21] Jhon J. Bravo, Jose L. Herrera, and Florian Luca. Common values of generalized Fibonacci and Pell sequences. Journal of Number Theory, 226(??):51–71, September 2021. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X21000767.

Blanton:1992:DDS

[BHM92] Earle L. Blanton, Jr., Spencer P. Hurd, and Judson S. McCranie. On a digraph defined by squaring modulo n. Fibonacci Quarterly, 30(4):322–333, November 1992. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/30-4/blanton.
pdf.

Bergum:1992:AFN

[BHP92] Gerald E. Bergum, A. F. Horadam, and Andreas N. Philippou, editors. Applications of Fibonacci numbers: proceedings of the Fifth International Conference on Fibonacci numbers and their applications, the University of St. Andrews, Scotland, July 20–24, 1992. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1992. ISBN 0-7923-2491-9. LCCN ????

Bundschuh:2006:GMI

[BHS06] P. Bundschuh, L. C. Hsu, and P. J.-S. Shiue. Generalized Möbius inversion — theoretical and computational aspects. Fibonacci Quarterly, 44(2):109–116, May 2006. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/44-
2/bundschuh.pdf.

Barrale:2011:SID

[BHS11] Tom Barrale, R. Hendel, and Michael Sluys. Sequences of the initial digits of Fibonacci numbers. In Proceedings of the 14th international conference on Fibonacci numbers and their applications, Morelia, Mexico, July 5–9, 2010, pages 25–42. Sociedad Matemática Mexicana, México, DF, México, 2011. ISBN 607-02-2543-0 (paperback).

Butler:2010:IDT

[BHT10] Steve Butler, Paul Horn, and Eric Tressler. Intersecting domino tilings. Fibonacci Quarterly, 48(2):114–120, May 2010. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/48-2/butler.pdf.

Bialek:1995:NCR

[Bia95] Krystyna Bialek. A note on Choudhry’s results. Fibonacci Quarterly, 33(2):179–180, May 1995. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/33-2/bialek.
pdf.

Bicknell:1965:FFS

[Bic65] Marjorie Bicknell. Fibonacci fantasy: The square root of the Q matrix. Fibonacci Quarterly, 3(1):67–70, February 1965. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/3-1/bicknell.pdf.

Bicknell:1968:CPS

[Bic68a] Marjorie Bicknell. A curious property of a second fraction. Fibonacci Quarterly, 6(5):34–??, November 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/6-5/bicknell1.pdf.

Bicknell:1968:RM

[Bic68b] Marjorie Bicknell. Romance in mathematics. Fibonacci Quarterly, 6(5):43–??, November 1968. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/6-5/bicknell2.
pdf.

Bicknell:1969:NFN

[Bic69] Marjorie Bicknell. A note on Fibonacci numbers in high school algebra. Fibonacci Quarterly, 7(3):301–302, October 1969. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-3/bicknell.pdf.

Bicknell:1970:PFN

[Bic70] Marjorie Bicknell. A primer for the Fibonacci numbers — Part VII. Fibonacci Quarterly, 8(4):407–420, October 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-4/bicknell.pdf.

Bicknell:1971:FC

[Bic71a] Marjorie Bicknell. A Fibonacci crostic. Fibonacci Quarterly, 9(5): 538–540, December 1971. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/9-5/bicknell2.pdf.

Bicknell:1971:PFNa

[Bic71b] Marjorie Bicknell. A primer for Fibonacci numbers: Part VIII. Fibonacci Quarterly, 9(1):74–80, February 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-1/bicknell.pdf.

Bicknell:1972:BRB

[Bic72a] Marjorie Bicknell. Book review: A Mathematical model of life and living, by Li-kung Shaw. Fibonacci Quarterly, 10(4):444, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/10-4/review.pdf. See [kS74].

Bicknell:1972:DII

[Bic72b] Marjorie Bicknell. Determinants and identities involving Fibonacci squares. Fibonacci Quarterly, 10(2):147–156, February 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/10-2/bicknell-a.pdf.

Bicknell:1973:BRB

[Bic73a] Marjorie Bicknell. Book review: I Ching Games, by H. Y. Li and Sibley S. Morrill. Fibonacci Quarterly, 11(3):266–??, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/11-3/review.pdf. See [LM71].

Bicknell:1973:PFN

[Bic73b] Marjorie Bicknell. A primer for the Fibonacci numbers: Part XIII. Fibonacci Quarterly, 11(5):511–516, December 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-5/bicknell2.pdf.

Bicknell:1975:PPS

[Bic75] Marjorie Bicknell. A primer on the Pell sequence and related sequences. Fibonacci Quarterly, 13(4):345–349, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-4/bicknell.pdf.

Biggs:2021:LPF

[Big21] Norman Biggs. Linear programming from Fibonacci to Farkas. Annals of Science, 78(1):1–21, 2021. CODEN ANNSA8. ISSN 0003-3790 (print), 1464-505X (electronic).

Bilgici:2014:TGL

[Bil14] Goksal Bilgici. Two generalizations of Lucas sequence. Applied Mathematics and Computation, 245(??):526–538, October 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300314010807.

Bisht:1984:SCP

[Bis84] C. S. Bisht. Some congruence properties of generalized Lucas integral sequences. Fibonacci Quarterly, 22(4):290–295, November 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/22-4/bisht.pdf.

Bicknell-Johnson:1978:FCH

[BJ78a] Marjorie Bicknell-Johnson. Fibonacci chromotology or how to paint your rabbit. Fibonacci Quarterly, 16(5):426–427, October 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/16-5/bicknell.pdf.

Bicknell-Johnson:1978:GDC

[BJ78b] Marjorie Bicknell-Johnson. A golden double crostic. Fibonacci Quarterly, 16(1):67–69, February 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
1/bicknell-a.pdf.

Bicknell-Johnson:1979:APT

[BJ79a] Marjorie Bicknell-Johnson. Addenda to “Pythagorean Triples Containing Fibonacci Numbers: Solutions for F_{n}^{2} ± F_{k}^{2} = K^{2}”. Fibonacci Quarterly, 17(4):293–??, December 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/17-4/bicknell.pdf.

Bicknell-Johnson:1979:PTC

[BJ79b] Marjorie Bicknell-Johnson. Pythagorean triples containing Fibonacci numbers: Solution for F_{n}^{2} ± F_{k}^{2} = K^{2}. Fibonacci Quarterly, 17(1):1–12, February 1979. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/17-1/bicknell.
pdf.

Bicknell-Johnson:1980:MVH

[BJ80] Marjorie Bicknell-Johnson. In memoriam — Verner E. Hoggatt, Jr. Fibonacci Quarterly, 18(4):289–??, December 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-4/in-memoriam.pdf.

Bicknell-Johnson:1981:DSH

[BJ81] Marjorie Bicknell-Johnson. Diagonal sums in the harmonic triangle. Fibonacci Quarterly, 19(3):196–199, August 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-3/bicknell.pdf.

Bicknell-Johnson:1985:GWN

[BJ85] Marjorie Bicknell-Johnson. Generalized Wythoff numbers form simultaneous Fibonacci representations. Fibonacci Quarterly, 23 (4):308–318, November 1985. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/23-4/bicknell.
pdf.

Bicknell-Johnson:1986:LE

[BJ86] Marjorie Bicknell-Johnson. Letter to the editor. Fibonacci Quarterly, 24(4):309–??, November 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-
4/letter1.pdf.

Bicknell-Johnson:1987:SHF

[BJ87] Marjorie Bicknell-Johnson. A short history of The Fibonacci Quarterly. Fibonacci Quarterly, 25(1):2–5, February 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-1/bicknell.pdf.

Bicknell-Johnson:1989:RRP

[BJ89] Marjorie Bicknell-Johnson. Retrograde renegades and the Pascal connection: Repeating decimals represented by Fibonacci and other sequences appearing from right to left. Fibonacci Quarterly, 27(5):448–457, November 1989. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/27-5/bicknell.
pdf.

Bicknell-Johnson:1990:DPF

[BJ90] Marjorie Bicknell-Johnson. Divisibility properties of the Fibonacci numbers minus one, generalized to C_{n} = C_{n−1} + C_{n−2} + k. Fibonacci Quarterly, 28(2):107–112, May 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/28-2/bicknell.pdf.

Bicknell-Johnson:1993:RRP

[BJ93] Marjorie Bicknell-Johnson. Retrograde renegades and the Pascal connection II: Repeating decimals represented by sequences of diagonal sums of generalized Pascal triangles appearing from right to left. Fibonacci Quarterly, 31(4):346–353, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-4/bicknell.pdf.

Brenton:1995:SC

[BJ95] Lawrence Brenton and Mi-Kyung Joo. On the system of congruences Π_{j≠i}n_{j} ≡ 1 mod n_{i}. Fibonacci Quarterly, 33(3):258–267, June 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/33-3/brenton.pdf.

Bicknell-Johnson:1999:BRB

[BJ99] Marjorie Bicknell-Johnson. Book review: The Golden Ratio and Fibonacci Numbers, by Richard A. Dunlap. Fibonacci Quarterly, 37(2):116–??, May 1999. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/37-2/review.pdf. See

[Dun97].

Bicknell-Johnson:2002:LIH

[BJ02] Marjorie Bicknell-Johnson. The least integer having p Fibonacci representations, p prime. Fibonacci Quarterly, 40(3):260–265, June 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/40-3/bicknell.pdf.

Bicknell-Johnson:2003:SDA

[BJ03] Marjorie Bicknell-Johnson. Stern’s diatomic array applied to Fibonacci representations. Fibonacci Quarterly, 41(2):169–179, May 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/41-2/bicknell.pdf.

Bicknell-Johnson:2008:JRH

[BJ09] Marjorie Bicknell-Johnson. Julia Robinson and Hilbert’s Tenth Problem (DVD). Fibonacci Quarterly, 46/47(2):135–??, May 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Papers/46_47-2/Johnson_DVD_11-08.pdf.

Bicknell-Johnson:2014:FAH

[BJ14] Marjorie Bicknell-Johnson. The Fibonacci Association: Historical snapshots. Fibonacci Quarterly, 52(5):1–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/52-5/Johnson.pdf.

Bicknell-Johnson:2018:RIC

[BJ18] Marjorie Bicknell-Johnson. A report on the 18th International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 56(3):194–??, August 2018. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Announcements/ConferenceReport18.pdf.

Bicknell-Johnson:2019:CR

[BJ19] Marjorie Bicknell-Johnson. Conference report. Fibonacci Quarterly, 57(5):i–??, December 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/57-
5/ConferenceReport.pdf.

Bicknell-Johnson:2020:CR

[BJ20] Marjorie Bicknell-Johnson. Conference report. Fibonacci Quarterly, 58(5):1–??, December 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/58-
5/ConferenceReport.pdf.

Bicknell-Johnson:2022:CRa

[BJ22] Marjorie Bicknell-Johnson. Conference report. Fibonacci Quarterly, 60(5):1–??, December 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/60-
5/report2022.pdf.

Bicknell-Johnson:1995:VGR

[BJD95] Marjorie Bicknell-Johnson and Duane DeTemple. Visualizing golden ratio sums with tiling patterns. Fibonacci Quarterly, 33 (4):298–302, August 1995. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/33-4/bicknell.pdf.

Bicknell-Johnson:1995:GII

[BJE95] Marjorie Bicknell-Johnson and David A. Englund. Greatest integer identities for generalized Fibonacci sequences {H_{n}}, where H_{n} = H_{n−1} + H_{n−2}. Fibonacci Quarterly, 33(1):50–58, February 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/33-1/bicknell.pdf.

Bicknell-Johnson:1999:NRU

[BJF99] M. Bicknell-Johnson and D. C. Fielder. The number of representations of N using distinct Fibonacci numbers, counted by recursive formulas. Fibonacci Quarterly, 37(1):47–60, February 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/37-1/bicknell.pdf.

Bicknell-Johnson:2001:LNH

[BJF01] Marjorie Bicknell-Johnson and Daniel C. Fielder. The least number having 331 representations as a sum of distinct Fibonacci numbers. Fibonacci Quarterly, 39(5):455–461, November 2001.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-5/bicknell.pdf.

Bicknell-Johnson:1996:CIG

[BJS96] Marjorie Bicknell-Johnson and Colin Paul Spears. Classes of identities for the generalized Fibonacci numbers G_{n} = G_{n−1} + G_{n−c} from matrices with constant valued determinants. Fibonacci Quarterly, 34(2):121–128, May 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-2/
bicknell.pdf.

Bezuszka:1990:PFF

[BK90] Stanley Bezuszka and Stephen Kokoska. A polynomial formula for Fibonacci numbers. Fibonacci Quarterly, 28(2):151–155, May 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/28-2/bezuszka.pdf.

Burger:1996:SQI

[BK96] Edward B. Burger and Christopher S. Kollett. On the structure of quadratic irrationals associated with generalized Fibonacci and Lucas numbers. Fibonacci Quarterly, 34(3):200–212, June 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-3/burger.pdf.

Balakrishnan:2000:RCB

[BK00] Narayanaswamy Balakrishnan and Markos V. Koutras. Random combinations with bounded differences and cospan. Fibonacci Quarterly, 38(2):145–156, May 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-2/
balakrishnan.pdf.

Bottcher:2016:LZS

[BK16] Albrecht Böttcher and Fuad Kittaneh. The limit of the zero set of polynomials of the Fibonacci type. Journal of Number Theory, 163(??):89–100, June 2016. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X16000184.

Ballot:2017:LRO

[BKM17] Christian Ballot, Clark Kimberling, and Peter J. C. Moses. Linear recurrences originating from polynomial trees. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/55-5/ballot.
pdf.

Baril:2022:GCF

[BKV22] Jean-Luc Baril, Sergey Kirgizov, and Vincent Vajnovszki. Gray codes for Fibonacci q-decreasing words. Theoretical Computer Science, 927(??):120–132, August 26, 2022. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397522003577.

Bicknell:1965:NGR

[BL65] Marjorie Bicknell and James Leissner. A near-golden rectangle and related recursive series. Fibonacci Quarterly, 3(3):227–231, October 1965. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/3-3/bicknell.pdf.

Brillhart:1971:C

[BL71] John Brillhart and Emma Lehmer. Challenge. Fibonacci Quarterly, 9(5):525–??, December 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-5/
challenge.pdf.

Brezinski:1986:AEF

[BL86] C. Brezinski and A. Lembarki. Acceleration of extended Fibonacci sequences. Applied Numerical Mathematics: Transactions of IMACS, 2(1):1–8, February 1986. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Beslin:1992:GCS

[BL92] Scott Beslin and Steve Ligh. GCD-closed sets and the determinants of GCD matrices. Fibonacci Quarterly, 30(2):157–160, May 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/30-2/beslin.pdf.

Barry:1993:MGF

[BL93] Michael J. J. Barry and Anthony J. Lo Bello. The moment generating function of the geometric distribution of order k. Fibonacci Quarterly, 31(2):178–180, May 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-2/
barry.pdf.

Bagdasar:2013:CPC

[BL13a] Ovidiu D. Bagdasar and Peter J. Larcombe. On the characterization of periodic complex Horadam sequences. Fibonacci Quarterly, 51(1):28–??, February 2013. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/51-1/
bagdasar.pdf.

Bagdasar:2013:NCH

[BL13b] Ovidiu D. Bagdasar and Peter J. Larcombe. On the number of complex Horadam sequences with a fixed period. Fibonacci Quarterly, 51(4):339–??, November 2013. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/51-
4/bagdasar.pdf.

Bravo:2013:CGF

[BL13c] Jhon J. Bravo and Florian Luca. Coincidences in generalized Fibonacci sequences. Journal of Number Theory, 133(6):2121– 2137, June 2013. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X13000073. See corrigendum

[BL13d].

Bravo:2013:CCG

[BL13d] Jhon J. Bravo and Florian Luca. Corrigendum to “Coincidences in generalized Fibonacci sequences” [J. Number Theory 133 (6) (2013) 2121–2137]. Journal of Number Theory, 133(9):3205, September 2013. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X13000929. See [BL13c].

Bravo:2013:LPF

[BL13e] Jhon J. Bravo and Florian Luca. On the largest prime factor of the k-Fibonacci numbers. International Journal of Number Theory (IJNT), 9(5):1351–1366, August 2013. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042113500309.

Bagdasar:2017:MPH

[BL17a] Ovidiu D. Bagdasar and Peter J. Larcombe. On the masked periodicity of Horadam sequences: A generator-based approach. Fibonacci Quarterly, 55(4):332–??, November 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/55-4/bagdasar.pdf.

Bulawa:2017:IVG

[BL17b] Andrew Bulawa and Whan Ki Lee. Integer values of generating functions for the Fibonacci and related sequences. Fibonacci Quarterly, 55(1):74–??, February 2017. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/55-1/
bulawa.pdf.

Beresin:1971:CGS

[BLL71] M. Beresin, E. Levine, and D. Lubell. On the coefficients of a generating series. Fibonacci Quarterly, 9(5):467–476, December 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/9-5/beresin-a.pdf.

Benjamin:2020:CEB

[BLM20] Arthur T. Benjamin, John Lentfer, and Thomas C. Martinez. Counting on Euler and Bernoulli number identities. Fibonacci Quarterly, 58(5):30–??, December 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/58-
5/benjamin.pdf.

Block:1953:CFS

[Blo53a] D. Block. Curiosum # 330: Fibonacci summations. Scripta Mathematica, 19(2–3):191, 1953. ISSN 0036-9713.

Block:1953:SFI

[Blo53b] Daniel Block. Symbolic formulae involving Fibonacci numbers. Scripta Mathematica, 18(??):306–307, ???? 1953. ISSN 0036-9713.

Bloom:1964:CFF

[Blo64] David M. Bloom. Corrected factorizations of Fibonacci numbers. Fibonacci Quarterly, 2(3):218–??, October 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-3/bloom.pdf.

Bloom:1965:PGF

[Blo65] D. M. Bloom. On periodicity in generalized Fibonacci sequences. American Mathematical Monthly, 72(8):856–861, October 1965. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Berczes:2010:GBS

[BLP10] Attila Bérczes, Kálmán Liptai, and István Pink. On generalized balancing sequences. Fibonacci Quarterly, 48(2):121–128, May 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-2/berczes.pdf.

Behera:2011:BFP

[BLPS11] Akrur Behera, Kálmán Liptai, Gopal Krishna Panda, and László Szalay. Balancing with Fibonacci powers. Fibonacci Quarterly, 49(1):28–33, February 2011. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/49-1/behera.pdf.

Brillhart:1975:NPC

[BLS75] John Brillhart, D. H. Lehmer, and J. L. Selfridge. New primality criteria and factorizations of 2^{m} ± 1. Mathematics of Computation, 29(130):620–647, April 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brodal:2012:SFH

[BLT12] Gerth Stølting Brodal, George Lagogiannis, and Robert E. Tarjan. Strict Fibonacci heaps. In ACM [ACM12], pages 1177–1184. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

Blumenson:1972:CFN

[Blu72] Leslie E. Blumenson. A characterization of Fibonacci numbers suggested by a problem in cancer research. Fibonacci Quarterly, 10(3):262–264, April 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
3/blumenson-a.pdf.

Brown:2007:NDS

[BM07a] Ron Brown and Jonathan L. Merzel. The number of Ducci sequences with given period. Fibonacci Quarterly, 45(2):115–121, May 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/45-2/brown.pdf.

Bruckman:2007:STI

[BM07b] Paul S. Bruckman and Ray S. Melham. Some theorems involving powers of generalized Fibonacci numbers at non-equidistant points. Fibonacci Quarterly, 45(3):208–220, August 2007. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/45-3/bruckman.pdf.

Bernoussi:2001:AIG

[BMRS01] B. Bernoussi, W. Motta, M. Rachidi, and O. Saeki. Approximation of infinite generalized Fibonacci sequences and their asymptotic Binet formula. Fibonacci Quarterly, 39(2):168–180, May 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/39-2/bernoussi.pdf.

Bernoussi:2004:PGF

[BMRS04] B. Bernoussi, W. Motta, M. Rachidi, and O. Saeki. On periodic ∞-generalized Fibonacci sequences. Fibonacci Quarterly, 42(4):361–367, November 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
4/quartosamu04_2004.pdf.

Brillhart:1988:STF

[BMS88a] John Brillhart, Peter L. Montgomery, and Robert D. Silverman. Supplement to tables of Fibonacci and Lucas factorizations. Mathematics of Computation, 50(181):S1–S15, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Brillhart:1988:TFL

[BMS88b] John Brillhart, Peter L. Montgomery, and Robert D. Silverman. Tables of Fibonacci and Lucas factorizations. Mathematics of Computation, 50(181):251–260, S1–S15, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Beck:1977:FPC

[BN77a] Walter E. Beck and Rudolph M. Najar. Fixed points of certain arithmetic functions. Fibonacci Quarterly, 15(4):337–341, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/15-4/beck2.pdf.

Beck:1977:MRA

[BN77b] Walter E. Beck and Rudolph M. Najar. More reduced amicable pairs. Fibonacci Quarterly, 15(4):331–332, December 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-4/beck1.pdf.

Beck:1985:HUH

[BN85] Walter E. Beck and Rudolph M. Najar. Hyperperfect and unitary hyperperfect numbers. Fibonacci Quarterly, 23(3):270–275, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/23-3/beck.pdf.

Beck:1993:RAA

[BN93] Walter E. Beck and Rudolph M. Najar. Reduced and augmented amicable pairs to 10 8. Fibonacci Quarterly, 31(4):295– 298, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/31-4/beck.pdf.

Brison:2003:LRS

[BN03] Owen J. Brison and J. Eurico Nogueira. Linear recurring sequence subgroups in the complex field. Fibonacci Quarterly, 41 (5):397–404, November 2003. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/41-5/brison.
pdf.

Brison:2006:MLR

[BN06] Owen J. Brison and J. Eurico Nogueira. Matrices and linear recurrences in finite fields. Fibonacci Quarterly, 44(2):103–108, May 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/44-2/brison.pdf.

Booth:2008:BPP

[BN09] Robert Booth and Hieu D. Nguyen. Bernoulli polynomials and Pascal’s square. Fibonacci Quarterly, 46/47(1):38–47, February 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/46_47-1/nguyen.pdf.

Brison:2019:LRS

[BN19] Owen J. Brison and J. Eurico Nogueira. Linear recurring sequence subgroups in the complex field — II. Fibonacci Quarterly, 57(2):148–??, May 2019. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/57-2/brison.pdf.

Benjamin:2003:PVC

[BNOS03] Arthur T. Benjamin, Judson D. Neer, Daniel E. Otero, and James A. Sellers. A probabilistic view of certain weighted Fibonacci sums. Fibonacci Quarterly, 41(4):360–364, August 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-4/benjamin.pdf.

Benjamin:2017:BPD

[BO17] Arthur T. Benjamin and Joel Ornstein. A bijective proof of a derangement recurrence. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/55-5/benjamin.pdf.

Boardman:1979:NMH

[Boa79] John Boardman. The normal modes of a hanging oscillator of order N. Fibonacci Quarterly, 17(1):37–39, February 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-1/boardman.pdf.

Boase:2001:RAP

[Boa01] Mansur S. Boase. A result about the primes dividing Fibonacci numbers. Fibonacci Quarterly, 39(5):386–391, November 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-5/boase.pdf.

Bohigian:1979:EWM

[Boh79] Haig E. Bohigian. Extensions of the W. Mnich problem. Fibonacci Quarterly, 17(2):172–177, April 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-2/
bohigian.pdf.

Boisen:1969:OPT

[Boi69] Monte B. Boisen, Jr. Overlays of Pascal’s triangle. Fibonacci Quarterly, 7(2):131–139, April 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-2/
boisen.pdf.

Boklan:1984:NG

[Bok84] Kent D. Boklan. The n-number game. Fibonacci Quarterly, 22 (2):152–155, May 1984. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/22-2/boklan.pdf.

Bolker:1977:TPW

[Bol77] Ethan D. Bolker. A topological proof of a well-known fact about Fibonacci numbers. Fibonacci Quarterly, 15(3):245–??, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/15-3/bolker-a.pdf.

Bollinger:1984:FSP

[Bol84] Richard C. Bollinger. Fibonacci k-sequences, Pascal-T triangles, and k-in-a-row problems. Fibonacci Quarterly, 22(2):146–151, May 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/22-2/bollinger.pdf.

Bollinger:1986:NPT

[Bol86] Richard C. Bollinger. A note on Pascal-T triangles, multinomial coefficients, and Pascal pyramids. Fibonacci Quarterly, 24(2): 140–143, May 1986. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/24-2/bollinger.pdf.

Bollinger:1989:MSP

[Bol89] Richard C. Bollinger. The Mann–Shanks primality criterion in the Pascal-T triangle T_{3}. Fibonacci Quarterly, 27(3):272–275, June 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/27-3/bollinger.pdf.

Boman:2020:GCP

[Bom20] Bruce M. Boman. Geometric capitulum patterns based on Fibonacci p-proportions. Fibonacci Quarterly, 58(5):91–??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/58-5/boman.pdf.

Booth:1975:IRR

[Boo75] Ada Booth. Idiot’s roulette revisited. Fibonacci Quarterly, 13(2): 181–184, April 1975. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/13-2/booth.pdf.

Boothroyd:1979:CAF

[Boo79] J. Boothroyd. Certification of Algorithm 2: Fibonacci search. The Computer Journal, 22(1):86, February 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL

http://www3.oup.co.uk/computer_journal/hdb/Volume_22/

Issue_01/tiff/86.tif. Reprinted from The Computer Bulletin. See [PP79].

Boscarol:1982:PBC

[Bos82] Mauro Boscarol. A property of binomial coefficients. Fibonacci Quarterly, 20(3):249–251, August 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
3/boscarol.pdf.

Boscarol:1985:NBC

[Bos85] Mauro Boscarol. A note on binomial coefficients and Chebyshev polynomials. Fibonacci Quarterly, 23(2):166–168, May 1985. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-2/boscarol.pdf.

Botten:1982:UFR

[Bot82] L. C. Botten. On the use of Fibonacci recurrence relations in the design of long wavelength filters and interferometers. Fibonacci Quarterly, 20(1):1–6, February 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-1/
botten.pdf.

Bouazzaoui:2020:PFS

[Bou20] Zakariae Bouazzaoui. On periods of Fibonacci sequences and real quadratic p-rational fields. Fibonacci Quarterly, 58(5):103– ??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/58-5/bouazzaoui.pdf.

Bowman:1988:NGD

[Bow88] Douglas Bowman. A new generalization of Davison’s theorem. Fibonacci Quarterly, 26(1):40–45, February 1988. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/26-1/bowman.pdf.

Bowman:2014:FCC

[Bow14] Douglas Bowman. Fibonacci contractions of continued fractions. Fibonacci Quarterly, 52(3):206–??, August 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-3/bowman.pdf.

Boyd:1992:BCN

[Boy92] A. V. Boyd. Bounds for the Catalan numbers. Fibonacci Quarterly, 30(2):136–138, May 1992. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/30-2/boyd.pdf.

Boyer:1997:AAH

[Boy97] John Boyer. Algorithm alley — heaps are usually implemented via binary trees, with the property that for every subtree, the root is the minimum item. here, John describes how to implement exceptionally fast ‘Fibonacci’ heaps. Dr. Dobb’s Journal of Software Tools, 22(1):106–??, January 1997. CODEN DDJOEB. ISSN 1044-789X.

Boyer:1998:AAR

[Boy98] John Boyer. Algorithm alley: Resizable data structures. Dr. Dobb’s Journal of Software Tools, 23(1):115–116, 118, 129, January 1998. CODEN DDJOEB. ISSN 1044-789X. Discusses some deficiencies of the Java library hash table support, and compares it with his algorithm and that used in the C++ Standard Template Library. Also compares the Jenkins hash function [Jen97] with the one proposed in this paper.

Boyadzhiev:2007:DPE

[Boy07] Khristo N. Boyadzhiev. Derivative polynomials for tanh, tan, sech and sec in explicit form. Fibonacci Quarterly, 45(4):291– 303, November 2007. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/45-4/boyadzhiev.pdf.

Boyadzhiev:2008:PSI

[Boy09] Khristo N. Boyadzhiev. Power sum identities with generalized Stirling numbers. Fibonacci Quarterly, 46/47(4): 326–330, November 2008/2009. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/46_47-
4/boyadzhiev.pdf.

Behera:1999:SRT

[BP99] A. Behera and G. K. Panda. On the square roots of triangular numbers. Fibonacci Quarterly, 37(2):98–105, May 1999. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-2/behera.pdf.

Burns:2007:CNW

[BP07] Chris Burns and Benjamin Purcell. Counting the number of winning binary strings in the 1-dimensional same game. Fibonacci Quarterly, 45(3):233–238, August 2007. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/45-
3/burns.pdf.

Benjamin:2008:CAF

[BP09] Arthur T. Benjamin and Sean S. Plott. A combinatorial approach to Fibonomial coefficients. Fibonacci Quarterly, 46/47(1):7–9, February 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/46_47-1/benjamin.pdf.

Basu:2010:LRV

[BP10a] Manjusri Basu and Bandhu Prasad. Long range variations on the Fibonacci universal code. Journal of Number Theory, 130(9):1925–1931, September 2010. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X10000533.

Benjamin:2010:EFD

[BP10b] Arthur T. Benjamin and Sean S. Plott. Errata — “Fibonacci determinants — a combinatorial approach”. Fibonacci Quarterly, 48(3):276–??, August 2010. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers/48-3/Errata.pdf. See

[BCQ07].

Bodroza-Pantic:1997:FNA

[BPGC97] Olga Bodroza-Pantic, Ivan Gutman, and Sven J. Cyvin. Fibonacci numbers and algebraic structure count of some nonbenzenoid conjugated polymers. Fibonacci Quarterly, 35(1):75– 83, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/35-1/bodroza-pantic.pdf.

Bergum:1993:AFN

[BPH93] Gerald E. Bergum, Andreas N. Philippou, and A. F. Horadam, editors. Applications of Fibonacci numbers: proceedings of “The Fifth International Conference of Fibonacci Numbers and Their Applications”, The University of St. Andrews, Scotland, July 20– July 24, 1992. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. ISBN 0-7923-2491-9. LCCN QA241 .I58 1992.

Bodroza-Pantic:2007:ASC

[BPIK07] Olga Bodroža-Pantić and Angelina Ilić-Kovačević. Algebraic structure count of angular hexagonal-square chains. Fibonacci Quarterly, 45(1):3–9, February 2007. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/45-1/
bodroza.pdf.

Brinda:2014:BBW

[BPT14] Karel Břinda, Edita Pelantová, and Ondřej Turek. Balances of mbonacci words. Fundamenta Informaticae, 132(1):33–61, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Benjamin:1999:RFL

[BQ99] Arthur T. Benjamin and Jennifer J. Quinn. Recounting Fibonacci and Lucas identities. College Mathematics Journal, 30 (5):359–366, November 1999. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http://www.tandfonline.

com/doi/abs/10.1080/07468342.1999.11974086.

Benjamin:2003:PRC

[BQ03] Arthur Benjamin and Jennifer J. Quinn. Proofs that really count: the art of combinatorial proof, volume 27 of Dolciani mathematical expositions. Mathematical Association of America, Washington, DC, USA, 2003. ISBN 0-88385-333-7. xiv + 194 pp.

LCCN QA164.8 .B46 2003. URL http://www.loc.gov/catdir/
description/cam051/2003108524.html;   http://www.loc.
gov/catdir/toc/cam051/2003108524.html;   https://www.
math.utah.edu/pub/tex/bib/fibquart.bib.

Benjamin:2000:PTG

[BQS00] Arthur T. Benjamin, Jennifer J. Quinn, and Francis Edward Su. Phased tilings and generalized Fibonacci identities. Fibonacci Quarterly, 38(3):282–289, June 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-3/
benjamin.pdf.

Bruschi:1982:EFG

[BR82] Massimo Bruschi and Paolo Emilio Ricci. An explicit formula for f(A) and the generating functions of the generalized Lucas polynomials. SIAM Journal on Mathematical Analysis, 13(1): 162–165, January 1982. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Baker:1985:NIE

[BR85] I. N. Baker and P. J. Rippon. A note on infinite exponentials. Fibonacci Quarterly, 23(2):106–112, May 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-2/
baker.pdf.

Benjamin:2014:CPF

[BR14] Arthur T. Benjamin and Elizabeth Reiland. Combinatorial proofs of Fibonomial identities. Fibonacci Quarterly, 52(5):28–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Papers1/52-5/Benjamin.pdf.

Brady:1971:ASR

[Bra71] Wray G. Brady. Additions to the summation of reciprocal Fibonacci and Lucas series. Fibonacci Quarterly, 9(4):402–404, October 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/9-4/brady-a.pdf.

Brady:1972:LF

[Bra72] Wray G. Brady. The Lambert function. Fibonacci Quarterly, 10 (2):199–200, February 1972. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/10-2/brady.pdf.

Brady:1978:MBL

[Bra78] W. G. Brady. More on Benford’s law. Fibonacci Quarterly, 16 (1):51–52, February 1978. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/16-1/brady.pdf.

Bradford:1990:FST

[Bra90] Phillip G. Bradford. The Fibonacci sequence and the time complexity of generating the Conway polynomial and related topological invariants. Fibonacci Quarterly, 28(3):240–251, August 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/28-3/bradford.pdf.

Bracken:1995:DMF

[Bra95] Paul Bracken. Dynamics of the mapping f(x) = (x + 1)^{−1}. Fibonacci Quarterly, 33(4):357–358, August 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-4/bracken.pdf.

Branson:1996:ESN

[Bra96] David Branson. An extension of Stirling numbers. Fibonacci Quarterly, 34(3):213–223, June 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-3/
branson.pdf.

Bradley:2000:GCB

[Bra00] Sean Bradley. A geometric connection between generalized Fibonacci sequences and nearly golden sections. Fibonacci Quarterly, 38(2):174–178, May 2000. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/38-2/bradley.
pdf.

Branson:2006:CII

[Bra06] David Branson. A combinatorial interpretation of identities involving Stirling numbers and their generalizations. Fibonacci Quarterly, 44(2):131–140, May 2006. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/44-2/
branson.pdf.

Brennan:1964:FPPb

[Bre64a] T. A. Brennan. Fibonacci powers and Pascal’s triangle in a matrix — Part II. Fibonacci Quarterly, 2(3):177–183, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-3/brennan.pdf.

Brennan:1964:FPPa

[Bre64b] Terrence A. Brennan. Fibonacci powers and Pascal’s triangle in a matrix. Fibonacci Quarterly, 2(2):93–103, April 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-2/brennan.pdf.

Brent:1974:ERO

[Bre74a] Barry Brent. An expansion of e^{x} off roots of one. Fibonacci Quarterly, 12(2):208–??, April 1974. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/12-2/brent2.
pdf.

Brent:1974:FEP

[Bre74b] Barry Brent. Functional equations with prime roots from arithmetic expressions for G_{α}. Fibonacci Quarterly, 12(2):199–207, April 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/12-2/brent1.pdf.

Brent:1994:PGF

[Bre94] Richard P. Brent. On the periods of generalized Fibonacci recurrences. Mathematics of Computation, 63(207):389–401, July 1994. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Breuer:1998:NPG

[Bre98] Florian Breuer. A note on a paper by Glaser and Schöffl. Fibonacci Quarterly, 36(5):463–466, November 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-5/breuer.pdf.

Bautista-Ramos:2020:IPF

[BRGGGS20] César Bautista-Ramos, Carlos Guillén-Galván, and Paulino Gómez-Salgado. Independence polynomials of Fibonacci trees are log-concave. Fibonacci Quarterly, 58(1):49–??, February 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/58-1/bautista.pdf.

Brian:1964:PLO

[Bri64a] Richard Brian. The problem of the little old lady trying to cross the busy street or Fibonacci gained and Fibonacci relost. Fibonacci Quarterly, 2(4):310–313, December 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/2-4/brian.pdf.

Brillhart:1964:RSO

[Bri64b] J. Brillhart. Remarks on a second order recurring sequence. Fibonacci Quarterly, 2(3):220–222, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/2-3/brillhart.pdf.

Brillhart:1983:LE

[Bri83a] John Brillhart. Letter to the editor: “Properties of polynomials having Fibonacci numbers for coefficients” [Fibonacci Quart. 21 (1983), no. 1, 62–64; MR 84f:10018a] by D. H. Lehmer and E. Lehmer. Fibonacci Quarterly, 21(4):259, November 1983. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-4/letter.pdf.

Brillhart:1983:LEP

[Bri83b] John Brillhart. Letter to the editor: “Properties of polynomials having Fibonacci numbers for coefficients” [Fibonacci Quart. 21 (1983), no. 1, 62–64; MR 84f:10018a] by D. H. Lehmer and E. Lehmer. Fibonacci Quarterly, 21(4):259, 1983. CODEN FIBQAU. ISSN 0015-0517.

Brison:1992:CFS

[Bri92] Owen J. Brison. Complete Fibonacci sequences in finite fields. Fibonacci Quarterly, 30(4):295–304, November 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-4/brison.pdf.

Brillhart:1998:NFP

[Bri98] John Brillhart. Note on Fibonacci primality testing. Fibonacci Quarterly, 36(3):222–228, June 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-3/
brillhart.pdf.

Brietzke:2006:GIA

[Bri06] Eduardo H. M. Brietzke. Generalization of an identity of Andrews. Fibonacci Quarterly, 44(2):166–171, May 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/44-2/brietzke.pdf.

Brooke:1963:LEF

[Bro63a] Maxey Brooke. Letter to the editor: Fibonacci formulas. Fibonacci Quarterly, 1(2):60–??, April 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-2/
brooke.pdf.

Brown:1963:GSC

[Bro63b] J. L. Brown, Jr. A generalization of semi-completeness for integer sequences. Fibonacci Quarterly, 1(1):3–15, February 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-1/brown.pdf.

Brooke:1964:FNT

[Bro64a] Maxey Brooke. Fibonacci numbers: Their history through 1900. Fibonacci Quarterly, 2(2):149–153, April 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/2-2/brooke.pdf.

Brown:1964:ZTS

[Bro64b] J. L. Brown, Jr. Zeckendorf’s theorem and some applications. Fibonacci Quarterly, 2(3):163–168, October 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-3/brown.pdf.

Brown:1965:NCF

[Bro65a] J. L. Brown, Jr. A new characterization of the Fibonacci numbers. Fibonacci Quarterly, 3(1):1–8, February 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/3-1/brown.pdf.

Brown:1965:REF

[Bro65b] John L. Brown, Jr. Reply to exploring Fibonacci magic squares. Fibonacci Quarterly, 3(2):146–??, April 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-2/
brown.pdf.

Brousseau:1966:BRB

[Bro66] Brother Alfred Brousseau. Book review: Recurring Sequences, by Dov Jarden. Fibonacci Quarterly, 4(3):208–??, October 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-3/review1.pdf. See [Jar66].

Brousseau:1967:SS

[Bro67a] Brother Alfred Brousseau. 1967 as the sum of squares. Fibonacci Quarterly, 5(2):208–??, April 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-2/
brousseau.pdf.

Brousseau:1967:BSR

[Bro67b] Brother Alfred Brousseau. Recurring Sequences (Review of Book by Dov Jarden). Fibonacci Quarterly, 5(4):328–??, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/5-4/review.pdf. See [Jar66].

Brousseau:1967:E

[Bro67c] Brother Alfred Brousseau. Editorial. Fibonacci Quarterly, 5(2): 169–170, April 1967. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/5-2/editorial.pdf.

Brousseau:1967:FG

[Bro67d] Brother Alfred Brousseau. A Fibonacci generalization. Fibonacci Quarterly, 5(2):171–174, April 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-2/
brousseau.pdf.

Brousseau:1967:RCP

[Bro67e] Brother Alfred Brousseau. Recreation corner — population explosion. Fibonacci Quarterly, 5(5):444–??, December 1967. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-5/brousseau.pdf.

Brousseau:1967:SFD

[Bro67f] Brother Alfred Brousseau. Summation of ∑_{k=1}^{n} k^{m}F_{k+r} finite difference approach. Fibonacci Quarterly, 5(1):91–98, February 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/5-1/brousseau.pdf.

Brown:1967:LT

[Bro67g] J. L. Brown, Jr. On Lamé’s theorem. Fibonacci Quarterly, 5(2): 153–160, April 1967. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/5-2/brown.pdf.

Brousseau:1968:BRB

[Bro68a] Brother Alfred Brousseau. Book review: 536 Puzzles and Curious Problems. Fibonacci Quarterly, 6(1):84–??, February 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-1/review1.pdf. See [DG67].

Brousseau:1968:LRR

[Bro68b] Brother Alfred Brousseau. Linear recursion relations, lesson two. Fibonacci Quarterly, 6(6):393–399, December 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-6/brousseau.pdf.

Brousseau:1968:TCP

[Bro68c] Brother Alfred Brousseau. On the trail of the California pine. Fibonacci Quarterly, 6(1):69–76, February 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-1/
brousseau2.pdf.

Brousseau:1968:RCS

[Bro68d] Brother Alfred Brousseau. Recreation corner — solution to population explosion. Fibonacci Quarterly, 6(1):58–59, February 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-1/brousseau1.pdf.

Brousseau:1968:RSL

[Bro68e] Brother Alfred Brousseau. Recurring sequences — lesson 1. Fibonacci Quarterly, 6(4):279–285, October 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/6-4/brousseau-a.pdf.

Brousseau:1968:SPF

[Bro68f] Brother Alfred Brousseau. A sequence of power formulas. Fibonacci Quarterly, 6(1):81–83, February 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-1/
brousseau3.pdf.

Brown:1968:CPI

[Bro68g] J. L. Brown, Jr. A combinatorial problem involving Fibonacci numbers. Fibonacci Quarterly, 6(1):34–35, February 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-1/brown.pdf.

Brousseau:1969:BRB

[Bro69a] Brother Alfred Brousseau. Book reviews: Fibonacci and Lucas numbers, by Verner E. Hoggatt, Jr., and Invitation to number theory, by Øystein Øre. Fibonacci Quarterly, 7(1):105–106, February 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/7-1/reviews.pdf. See [Hog69, Øre67].

Brousseau:1969:FLI

[Bro69b] Brother Alfred Brousseau. Fibonacci–Lucas infinite series research topic. Fibonacci Quarterly, 7(2):211–217, April 1969. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-2/brousseau3.pdf.

Brousseau:1969:FSC

[Bro69c] Brother Alfred Brousseau. Fibonacci statistics in conifers. Fibonacci Quarterly, 7(5):525–532, December 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/7-5/brousseau1.pdf.

Brousseau:1969:LRRc

[Bro69d] Brother Alfred Brousseau. Linear recursion relations — lesson five recursion relations of higher order. Fibonacci Quarterly, 7(3):295–300, October 1969. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/7-3/brousseau-
a.pdf.

Brousseau:1969:LRRb

[Bro69e] Brother Alfred Brousseau. Linear recursion relations — lesson four second-order linear recursion relations. Fibonacci Quarterly, 7(2):194–200, April 1969. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/7-2/brousseau2-a.
pdf.

Brousseau:1969:LRRd

[Bro69f] Brother Alfred Brousseau. Linear recursion relations — lesson six combining linear recursion relations. Fibonacci Quarterly, 7(5): 533–537, December 1969. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/7-5/brousseau2-a.
pdf.

Brousseau:1969:LRRa

[Bro69g] Brother Alfred Brousseau. Linear recursion relations lesson three — the Binet formulas. Fibonacci Quarterly, 7(1):99–104, February 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/7-1/brousseau-a.pdf.

Brousseau:1969:SIF

[Bro69h] Brother Alfred Brousseau. Summation of infinite Fibonacci series. Fibonacci Quarterly, 7(2):143–168, April 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/7-2/brousseau1.pdf.

Brown:1969:URI

[Bro69i] J. L. Brown, Jr. Unique representations of integers as sums of distinct Lucas numbers. Fibonacci Quarterly, 7(3):243–251, October 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/7-3/brown.pdf.

Brousseau:1970:BRB

[Bro70a] Brother Alfred Brousseau. Book review: Leonard of Pisa and the new mathematics of the Middle Ages, by Joseph Gies and Frances Gies. Fibonacci Quarterly, 8(3):280–??, April 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-3/review-a.pdf. See [GG69, GGA83].

Brousseau:1970:LRRb

[Bro70b] Brother Alfred Brousseau. Linear recursion relations — lesson eight asymptotic ratios in recursion relations. Fibonacci Quarterly, 8(3):311–316, April 1970. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/8-3/brousseau-
a.pdf.

Brousseau:1970:LRRa

[Bro70c] Brother Alfred Brousseau. Linear recursion relations — lesson seven analyzing linear recursion relations. Fibonacci Quarterly, 8(1):96–101, February 1970. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/8-1/brousseau-
a.pdf.

Brousseau:1970:LA

[Bro70d] Brother Alfred Brousseau. A Lucas analogue. Fibonacci Quarterly, 8(4):439–442, October 1970. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/8-4/brousseau.
pdf.

Brousseau:1971:CFQ

[Bro71a] Brother Alfred Brousseau. Continued fractions of quadratic Fibonacci ratios. Fibonacci Quarterly, 9(4):427–436, October 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-4/brousseau.pdf.

Brown:1971:GBR

[Bro71b] J. L. Brown, Jr. On generalized bases for real numbers. Fibonacci Quarterly, 9(5):477–496, December 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-5/
brown-a.pdf.

Brousseau:1972:AAL

[Bro72a] Brother Alfred Brousseau. Algorithm for analyzing a linear recursion sequence. Fibonacci Quarterly, 10(4):429–432, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/10-4/brousseau1-a.pdf.

Brousseau:1972:FMC

[Bro72b] Brother Alfred Brousseau. Fibonacci magic cards. Fibonacci Quarterly, 10(2):197–??, February 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
2/brousseau2.pdf.

Brousseau:1972:FNG

[Bro72c] Brother Alfred Brousseau. Fibonacci numbers and geometry. Fibonacci Quarterly, 10(3):303–318, April 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-3/
brousseau2-a.pdf.

Brousseau:1972:NCN

[Bro72d] Brother Alfred Brousseau. Note on the characteristic number of a sequence of Fibonacci squares. Fibonacci Quarterly, 10(3): 247–248, April 1972. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/10-3/brousseau1.pdf.

Brousseau:1972:NNF

[Bro72e] Brother Alfred Brousseau. A note on the number of Fibonacci sequences. Fibonacci Quarterly, 10(6):657–658, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-6/brousseau2.pdf.

Brousseau:1972:PFN

[Bro72f] Brother Alfred Brousseau. A primer for the Fibonacci numbers. Part X: On the representation of integers. Fibonacci Quarterly, 10(6):647–650, December 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
6/brousseau1.pdf.

Brousseau:1972:TIF

[Bro72g] Brother Alfred Brousseau. Table of indices with a Fibonacci relation. Fibonacci Quarterly, 10(2):182–184, February 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-2/brousseau1.pdf.

Brousseau:1972:YOF

[Bro72h] Brother Alfred Brousseau. Ye Olde Fibonacci Curiosity Shoppe. Fibonacci Quarterly, 10(4):441–443, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-4/brousseau2.pdf.

Brousseau:1973:GFS

[Bro73a] Brother Alfred Brousseau. Generalized Fibonacci shift formulas. Fibonacci Quarterly, 11(2):209–223, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-2/brousseau2.pdf.

Brousseau:1973:TOE

[Bro73b] Brother Alfred Brousseau. Through the other end of the telescope. Fibonacci Quarterly, 11(2):189–194, April 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-2/brousseau1.pdf.

Brousseau:1973:YOF

[Bro73c] Brother Alfred Brousseau. Ye olde Fibonacci curiosity shoppe. Fibonacci Quarterly, 11(3):332–??, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-3/brousseau.pdf.

Brousseau:1974:ATO

[Bro74a] Brother Alfred Brousseau. Algorithms for third-order recursion sequences. Fibonacci Quarterly, 12(2):167–174, April 1974. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-2/brousseau2.pdf.

Brousseau:1974:FSI

[Bro74b] Brother Alfred Brousseau. Fibonacci summations involving a power of a rational number — summary. Fibonacci Quarterly, 12 (2):146–??, April 1974. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/12-2/brousseau1-a.pdf.

Brousseau:1975:LIS

[Bro75a] Brother Alfred Brousseau. A least integer sequence investigation. Fibonacci Quarterly, 13(2):145–146, April 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-2/brousseau-a.pdf.

Brousseau:1975:SS

[Bro75b] Brother Alfred Brousseau. Symmetric sequences. Fibonacci Quarterly, 13(1):33–41, February 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-
1/brousseau.pdf.

Brown:1975:RMC

[Bro75c] J. L. Brown, Jr. A recursive method for counting integers not representable in certain expansions. Fibonacci Quarterly, 13(4): 299–302, December 1975. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/13-4/brown.pdf.

Brousseau:1976:RRP

[Bro76] Brother Alfred Brousseau. Recursion relations of products of linear recursion sequences. Fibonacci Quarterly, 14(2):159–166, April 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/14-2/brousseau.pdf.

Brousseau:1977:FND

[Bro77] Brother Alfred Brousseau. Fibonacci numbers in diatoms? Fibonacci Quarterly, 15(4):370–??, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/15-4/brousseau.pdf.

Brousseau:1978:FDT

[Bro78a] Brother Alfred Brousseau. Formula development through finite differences. Fibonacci Quarterly, 16(1):53–66, February 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-1/brousseau.pdf.

Brown:1978:SPG

[Bro78b] J. L. Brown, Jr. Simplified proof of a greatest integer function theorem. Fibonacci Quarterly, 16(4):307–309, August 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-4/brown.pdf.

Brown:1978:SSS

[Bro78c] J. L. Brown, Jr. Some sequence-to-sequence transformations which preserve completeness. Fibonacci Quarterly, 16(1):19–22, February 1978. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/16-1/brown1.pdf.

Brooks:1989:GRR

[Bro89] Jeffrey A. Brooks. A general recurrence relation for reflections in multiple glass plates. Fibonacci Quarterly, 27(3):267–271, June 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/27-3/brooks.pdf.

Brown:1994:PDS

[Bro94] Tom C. Brown. Powers of digital sums. Fibonacci Quarterly, 32(3):207–210, June 1994. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/32-3/brown.pdf.

Broy:1996:DPD

[Bro96] M. Broy, editor. Deductive program design: Proceedings of the NATO Advanced Study Institute on Deductive Program Design, held in Marktoberdorf, Germany, July 26–August 7, 1994, volume 152 of NATO ASI series. Series F, Computer and systems sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60947-4 (hardcover). LCCN QA76.9.D5 D38 1996. URL http://www.loc.

gov/catdir/enhancements/fy0812/96010788-d.html.

Brown:1997:DGD

[Bro97] Ezra Brown. Directed graphs defined by arithmetic (mod n). Fibonacci Quarterly, 35(4):346–351, November 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-4/brown.pdf.

Broughan:2003:ATR

[Bro03] Kevin A. Broughan. Adic topologies for the rational integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 55(??):711–723, ???? 2003. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:2017:NLG

[Bro17] Christopher Brown. The natural logarithm of the golden section. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/55-5/brown.pdf.

Bernoussi:2004:FBF

[BRS04a] Benaissa Bernoussi, Mustapha Rachidi, and Osamu Saeki. Factorial Binet formula and distributional moment formulation of generalized Fibonacci sequences. Fibonacci Quarterly, 42(4):320–329, November 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
4/quartbernoussi04_2004.pdf.

Bernoussi:2004:EBE

[BRS04b] Bernaissa Bernoussi, Mustapha Rachidi, and Osamu Saeki. Extending the Bernoulli–Euler method for finding zeros of holomorphic functions. Fibonacci Quarterly, 42(1):55–65, February 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-1/quartbernoussi01_2004.pdf.

Bruckner:1970:FSM

[Bru70] Gottfried Bruckner. Fibonacci sequence modulo a prime p ≡ 3 (mod 4). Fibonacci Quarterly, 8(2):217–220, March 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-2/bruckner.pdf.

Bruckman:1972:ISN

[Bru72] Paul S. Bruckman. An interesting sequence of numbers derived from various generating functions. Fibonacci Quarterly, 10(2): 169–181, February 1972. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/10-2/bruckman.pdf.

Bruckman:1973:SGS

[Bru73] Paul S. Bruckman. Some generalizations suggested by Gould’s systematic treatment of certain binomial identities. Fibonacci Quarterly, 11(3):225–240, October 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
3/bruckman.pdf.

Bruckman:1975:FX

[Bru75a] Paul S. Bruckman. A formula for A_{n}^{2}(x). Fibonacci Quarterly, 13(2):105–106, April 1975. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/13-2/bruckman2.pdf.

Bruckman:1975:GIM

[Bru75b] Paul S. Bruckman. A general identity for multisecting generating functions. Fibonacci Quarterly, 13(2):103–104, April 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-2/bruckman1.pdf.

Bruckman:1977:CM

[Bru77a] Paul S. Bruckman. Constantly mean. Fibonacci Quarterly, 15 (3):236–??, October 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-3/bruckman2.pdf.

Bruckman:1977:GFC

[Bru77b] Paul S. Bruckman. On generating functions with composite coefficients. Fibonacci Quarterly, 15(3):269–275, October 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-3/bruckman3.pdf.

Bruckman:1977:ECI

[Bru77c] Paul S. Bruckman. On the evaluation of certain infinite series by elliptic functions. Fibonacci Quarterly, 15(4):293–310, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/15-4/bruckman.pdf.

Bruckman:1977:POM

[Bru77d] Paul S. Bruckman. Pi-oh-MY! Fibonacci Quarterly, 15(3):230– ??, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/15-3/bruckman1.pdf.

Bruckman:1979:SDP

[Bru79] Paul S. Bruckman. Some divisibility properties of generalized Fibonacci sequences. Fibonacci Quarterly, 17(1):42–48, February 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/17-1/bruckman.pdf.

Bruce:1984:MTS

[Bru84] Ian Bruce. A modified Tribonacci sequence. Fibonacci Quarterly, 22(3):244–246, August 1984. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/22-3/bruce.pdf.

Bruce:1986:SGM

[Bru86a] Ian Bruce. Sequences generated by multiple reflections. Fibonacci Quarterly, 24(3):268–272, August 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-
3/bruce.pdf.

Bruckman:1986:RPD

[Bru86b] Paul S. Bruckman. A relation for the prime distribution function. Fibonacci Quarterly, 24(3):273–276, August 1986. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-3/bruckman.pdf.

Bruckman:1989:GZT

[Bru89] Paul S. Bruckman. The generalized Zeckendorf theorems. Fibonacci Quarterly, 27(4):338–347, August 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/27-4/bruckman.pdf.

Bruce:1993:RTP

[Bru93] J. W. Bruce. A really trivial proof of the Lucas–Lehmer test. American Mathematical Monthly, 100(4):370–371, April 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bruce:1994:AIG

[Bru94a] Ian Bruce. Another instance of the golden right triangle. Fibonacci Quarterly, 32(3):232–233, June 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-3/
bruce.pdf.

Bruckman:1994:LPO

[Bru94b] Paul S. Bruckman. Lucas pseudoprimes are odd. Fibonacci Quarterly, 32(2):155–157, May 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-
2/bruckman2.pdf.

Bruckman:1994:CPF

[Bru94c] Paul S. Bruckman. On a conjecture of Di Porto and Filipponi. Fibonacci Quarterly, 32(2):158–159, May 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-2/
bruckman3.pdf.

Bruckman:1994:ILP

[Bru94d] Paul S. Bruckman. On the infinitude of Lucas pseudoprimes. Fibonacci Quarterly, 32(2):153–154, May 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-2/
bruckman1.pdf.

Bruckman:1996:SIS

[Bru96] Paul S. Bruckman. Some interesting subsequences of the Fibonacci and Lucas pseudoprimes. Fibonacci Quarterly, 34(4): 332–341, August 1996. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/34-4/bruckman.pdf.

Bruckman:2000:DCP

[Bru00] Paul S. Bruckman. On the degree of the characteristic polynomial of powers of sequences. Fibonacci Quarterly, 38(1):35–38, February 2000. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/38-1/bruckman.pdf.

Balasubrahmanyan:1987:FPM

[BS87] N. Balasubrahmanyan and R. Sivaramakrishnan. Friendly-pairs of multiplicative functions. Fibonacci Quarterly, 25(4):320–321, November 1987. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/25-4/balasubrahmanyan.pdf.

Brown:1993:RRF

[BS93] Tom C. Brown and Peter Jau-Shyong Shiue. A remark related to the Frobenius problem. Fibonacci Quarterly, 31(1):32–36, February 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/31-1/brown.pdf.

Butler:1996:ANN

[BS96] Jon T. Butler and Tsutomu Sasao. Average number of nodes in binary decision diagrams of Fibonacci functions. Fibonacci Quarterly, 34(5):413–422, November 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-
5/butler.pdf.

Butler:1997:PDR

[BS97] Jon T. Butler and Tsutomu Sasao. On the proportion of digits in redundant numeration systems. Fibonacci Quarterly, 35(2): 172–180, May 1997. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/35-2/butler.pdf.

Benjamin:2011:TFB

[BS11] Arthur T. Benjamin and Jacob N. Scott. Third and fourth binomial coefficients. Fibonacci Quarterly, 49(2):99–101, May 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/49-2/benjamin.pdf.

Baird-Smith:2019:GZG

[BSEFM19] Paul Baird-Smith, Alyssa Epstein, Kristen Flint, and Steven J. Miller. The generalized Zeckendorf game. Fibonacci Quarterly, 57(5):1–??, December 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/57-5/
bairdsmith.pdf.

Brown:1995:SSO

[BsS95] Tom C. Brown and Peter Jau shyong Shiue. Squares of secondorder linear recurrence sequences. Fibonacci Quarterly, 33(4): 352–356, August 1995. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/33-4/brown.pdf.

Bostan:2008:PSC

[BSS08] Alin Bostan, Bruno Salvy, and Éric Schost. Power series composition and change of basis. In Jeffrey [Jef08], pages 269–276. ISBN 1-59593-904-0. LCCN ????

Bihani:2005:AIF

[BSY05] Prerna Bihani, Wendy Pusser Sheppard, and Paul Thomas Young. p-adic Interpolation of the Fibonacci sequence via hypergeometric functions. Fibonacci Quarterly, 43(3):213–226, August 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers/43-3/paper43-3-4.pdf.

Bunder:1991:ZRG

[BT91] Martin Bunder and Keith Tognetti. The Zeckendorf representation and the Golden Sequence. Fibonacci Quarterly, 29(3):217– 219, August 1991. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/29-3/bunder.pdf.

Burger:2000:DAB

[BT00] Edward B. Burger and Johathan M. Todd. On Diophantine approximation below the Lagrange constant. Fibonacci Quarterly, 38(2):136–144, May 2000. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/38-2/burger.pdf.

Bozkurt:2012:DIC

[BT12] Durmus Bozkurt and Tin-Yau Tam. Determinants and inverses of circulant matrices with Jacobsthal and Jacobsthal–Lucas Numbers. Applied Mathematics and Computation, 219(2):544–551, October 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300312006388.

Belkhir:2022:SCAa

[BTD22] Amine Belkhir, Elif Tan, and Mehmet Dağlı. Some combinatorial aspects of bi-periodic incomplete Horadam sequences. Fibonacci Quarterly, 60(5):39–??, December 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/60-
5/belkhir.pdf.

Buchanan:1964:MNP

[Buc64a] Floyd Buchanan. Mathematical notes: N-th powers in the Fibonacci series. American Mathematical Monthly, 71(6):647–649, June/July 1964. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Buchanan:1964:MNT

[Buc64b] Floyd Buchanan. Mathematical notes: N-th powers in the Fibonacci series. American Mathematical Monthly, 71(6):647–649, June/July 1964. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Buchanan:1964:MNR

[Buc64c] Floyd Buchanan. Mathematical notes: Retraction of “N-th Powers in the Fibonacci Series”. American Mathematical Monthly, 71 (10):1112, December 1964. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bucknell:1967:FNS

[Buc67] R. S. Bucknell. Fibonacci numbers and some prime reciprocals. Fibonacci Quarterly, 5(3):294–297, October 1967. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-3/bucknell.pdf.

Budgor:1980:SPP

[Bud80] Aaron B. Budgor. Star polygons, Pascal’s triangle, and Fibonacci numbers. Fibonacci Quarterly, 18(3):229–230, October 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-3/budgor.pdf.

Bumby:1987:IIR

[Bum87] Richard T. Bumby. Incredible identities revisited. Fibonacci Quarterly, 25(1):62–64, February 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-
1/bumby.pdf.

Bunder:1975:HFF

[Bun75a] M. W. Bunder. On Halsey’s Fibonacci function. Fibonacci Quarterly, 13(3):209–210, October 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-
3/bunder1.pdf.

Bunder:1975:PP

[Bun75b] M. W. Bunder. Products and powers. Fibonacci Quarterly, 13 (3):279–??, October 1975. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/13-3/bunder2.pdf.

Bunder:1975:SCG

[Bun75c] M. W. Bunder. A special case of the generalized Fibonacci sequence over an arbitrary ring with identity. Fibonacci Quarterly, 13(3):280–??, October 1975. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/13-3/bunder3.pdf.

Bunder:1978:MFF

[Bun78] M. W. Bunder. More Fibonacci functions. Fibonacci Quarterly, 16(2):97–98, April 1978. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/16-2/bunder.pdf.

Bunder:1981:NDD

[Bun81] M. W. Bunder. A new definition of division in rings of quotients of Euclidean rings. Fibonacci Quarterly, 19(5):440–446, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-5/bunder.pdf.

Bunder:1992:ZRU

[Bun92] M. W. Bunder. Zeckendorf representations using negative Fibonacci numbers. Fibonacci Quarterly, 30(2):111–115, May 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/30-2/bunder.pdf.

Bunder:2006:SM

[Bun06] Martin W. Bunder. Self matching in ⌊nα⌋. Fibonacci Quarterly, 44(4):290–296, November 2006. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/44-4/bunder.
pdf.

Bunder:2012:HFP

[Bun12] Martin W. Bunder. Horadam functions and powers of irrationals. Fibonacci Quarterly, 50(4):304–??, November 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-4/bunder.pdf.

Bunder:2014:PPP

[Bun14] Martin W. Bunder. Products and powers, powers and exponentiations, …. Fibonacci Quarterly, 52(2):172–??, May 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-2/bunder.pdf.

Burr:1971:MWF

[Bur71] S. A. Burr. On moduli for which the Fibonacci sequence contains a complete system of residues. Fibonacci Quarterly, 9(5):497–503, December 1971. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/9-5/burr.pdf.

Burrage:1990:GFP

[Bur90] Kevin Burrage. Generalized Fibonacci polynomials and the functional iteration of rational functions of degree one. Fibonacci Quarterly, 28(2):175–180, May 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-2/
burrage.pdf.

Burr:1992:UEN

[Bur92] Stefan A. Burr, editor. The unreasonable effectiveness of number theory: American Mathematical Society short course, August 6–7, 1991, Orono, Maine, volume 46 of Proceedings of symposia in applied mathematics. American Mathematical Society, Providence, RI, USA, 1992. ISBN 0-8218-5501-8. LCCN QA241 .U67 1992.

Buschman:1963:FNC

[Bus63] R. G. Buschman. Fibonacci numbers, Chebyshev polynomials, generalizations and difference equations. Fibonacci Quarterly, 1 (4):1–7, December 1963. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/1-4/buschman-a.pdf.

Buschman:1965:GFF

[Bus65] R. G. Buschman. A generating function for Fibonacci numbers. Fibonacci Quarterly, 3(3):199–200, October 1965. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/3-3/buschman.pdf.

Buschman:1970:DSM

[Bus70] R. G. Buschman. The “difference series” of Madachy. Fibonacci Quarterly, 8(4):372–374, October 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-4/
buschman.pdf.

Buschman:1973:RPS

[Bus73a] R. G. Buschman. Representations as products or as sums. Fibonacci Quarterly, 11(3):295–301, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-3/buschman2.pdf.

Buschman:1973:SSS

[Bus73b] R. G. Buschman. Some simple sieves. Fibonacci Quarterly, 11 (3):247–254, October 1973. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/11-3/buschman1.pdf.

Butter:1973:ICD

[But73] D. A. Butter. An inequality in a certain Diophantine equation. Fibonacci Quarterly, 11(3):315–316, October 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-3/butter.pdf.

Butcher:1978:CCS

[But78] J. C. Butcher. On a conjecture concerning a set of sequences satisfying the Fibonacci difference equation. Fibonacci Quarterly, 16(1):81–83, February 1978. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/16-1/butcher.pdf.

Butler:1990:TNP

[But90] Jon T. Butler. On teh number of propagation paths in multilayer media. Fibonacci Quarterly, 28(4):334–339, November 1990. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/28-4/butler.pdf.

Baril:2005:MCL

[BV05a] Jean-Luc Baril and Vincent Vajnovszki. Minimal change list for Lucas strings and some graph theoretic consequences. Theoretical Computer Science, 346(2–3):189–199, November 28, 2005. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Bruin:2005:CFL

[BV05b] H. Bruin and O. Volkova. The complexity of Fibonacci-like kneading sequences. Theoretical Computer Science, 337(1–3): 379–389, June 9, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Balasubramanian:1995:SDD

[BVB95] K. Balasubramanian, Roman Viveros, and N. Balakrishnan. Some discrete distributions related to extended Pascal triangles. Fibonacci Quarterly, 33(5):415–425, November 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-5/balasubramanian.pdf.

Ballew:1979:PTT

[BW79] David W. Ballew and Ronald C. Weger. Pythagorean triples and triangular numbers. Fibonacci Quarterly, 17(2):168–171, April 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/17-2/ballew.pdf.

Baillie:1980:LP

[BW80] Robert Baillie and Samuel S. Wagstaff, Jr. Lucas pseudoprimes. Mathematics of Computation, 35(152):1391–1417, October 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Burke:1981:ABL

[BW81] John R. Burke and William A. Webb. Asymptotic behavior of linear recurrences. Fibonacci Quarterly, 19(4):318–321, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-4/burke.pdf.

Burstein:1997:CSL

[BW97] Alexander Burstein and Herbert S. Wilf. On cyclic strings without long constant blocks. Fibonacci Quarterly, 35(3):240–247, August 1997. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/35-3/burstein.pdf.

Bergum:1975:CPR

[BWH75] G. E. Bergum, W. J. Wagner, and V. E. Hoggatt, Jr. Chebyshev polynomials and related sequences. Fibonacci Quarterly, 13(1): 19–24, February 1975. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/13-1/bergum.pdf.

Brilleslyper:2016:CGP

[BWWW16] Michael A. Brilleslyper, Nathan Wakefield, A. J. Wallerstein, and Bradley Warner. Comparing the growth of the prime numbers to the natural numbers. Fibonacci Quarterly, 54(1):65–??, February 2016. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/54-1/brilleslyper.pdf.

Bahsi:2016:NAN

[BY16] Ayşe Kurt Bahşı and Salih Yalçınbaş. A new algorithm for the numerical solution of telegraph equations by using Fibonacci polynomials. Mathematical and Computational Applications, 21(2): ??, June 2016. CODEN ???? ISSN 2297-8747. URL https://

www.mdpi.com/2297-8747/21/2/15.

Boman:2019:GBP

[BYD^{+}19] Bruce M. Boman, Yihan Ye, Keith Decker, Christopher Raymond, and Gilberto Schleiniger. Geometric branching patterns based on p-Fibonacci sequences: Self-similarity across different degrees of branching and multiple dimensions. Fibonacci Quarterly, 57(5):29–??, December 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/57-
5/boman.pdf.

Byrd:1963:EAF

[Byr63] Paul F. Byrd. Expansion of analytic functions in polynomials associated with Fibonacci numbers. Fibonacci Quarterly, 1(1): 16–27, February 1963. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/1-1/byrd.pdf.

Byrd:1965:EAF

[Byr65] Paul F. Byrd. Expansion of analytic functions in terms involving Lucas numbers or similar number sequences. Fibonacci Quarterly, 3(2):101–113, April 1965. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/3-2/byrd.pdf.

Byrd:1975:NRB

[Byr75a] Paul F. Byrd. New relations between Fibonacci and Bernoulli numbers. Fibonacci Quarterly, 13(1):59–69, February 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-1/byrd.pdf.

Byrd:1975:RBE

[Byr75b] Paul F. Byrd. Relations between Euler and Lucas numbers. Fibonacci Quarterly, 13(2):111–114, April 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-2/
byrd.pdf.

Brockman:2007:ABC

[BZ07] Greg Brockman and Ryan J. Zerr. Asymptotic behavior of certain Ducci sequences. Fibonacci Quarterly, 45(2):155–163, May 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/45-2/brockman.pdf.

Cadogan:1971:POP

[Cad71] C. C. Cadogan. On partly ordered partitions of a positive integer. Fibonacci Quarterly, 9(3):329–336, May 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-3/
cadogan.pdf.

Cadilhac:2016:RGR

[Cad16] Michaël Cadilhac. Review of: The Golden Ratio and Fibonacci Numbers by Richard A. Dunlap. ACM SIGACT News, 47(4):15– 17, December 2016. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Cai:1996:NNN

[Cai96a] Tianxin Cai. On 2-Niven numbers and 3-Niven numbers. Fibonacci Quarterly, 34(2):118–120, May 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-2/
cai1.pdf.

Cai:1996:SNU

[Cai96b] Tianxin Cai. On k-self-numbers and universal generated numbers. Fibonacci Quarterly, 34(2):144–146, May 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-2/cai2.pdf.

Cakic:1991:NEN

[Cak91] Nenad Cakić. A note on Euler’s numbers. Fibonacci Quarterly, 29(3):215–216, August 1991. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/29-3/cakic.pdf.

Cakic:1998:NSN

[Cak98] Nenad P. Cakić. Note on Stirling numbers of the second kind. Fibonacci Quarterly, 36(3):204–205, June 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-3/
cakic.pdf.

Callas:1972:RAN

[Cal72] Nicholas P. Callas. Representations of automorphic numbers. Fibonacci Quarterly, 10(4):393–396, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-4/callas-a.pdf.

Callan:1997:WDD

[Cal97] David Callan. When does m − n divide f(m) − f(n) ? A look at column-finite matrices. Fibonacci Quarterly, 35(4):290–299, November 1997. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/35-4/callan.pdf.

Callan:2000:CIL

[Cal00] David Callan. Certificates of integrality for linear binomials. Fibonacci Quarterly, 38(4):317–325, August 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-4/callan.pdf.

Campbell:2022:MBR

[Cam22a] John M. Campbell. A matrix-based recursion relation for F_{F_{n}}. Fibonacci Quarterly, 60(3):256–??, August 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-3/campbellj-1.pdf.

Campbell:2022:PMI

[Cam22b] John M. Campbell. Products of multiple-index Fibonacci numbers. Fibonacci Quarterly, 60(3):267–??, August 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-3/campbellj-2.pdf.

Campbell:2023:SFN

[Cam23] John M. Campbell. Sums of Fibonacci numbers indexed by integer parts. Fibonacci Quarterly, 61(2):143–??, May 2023. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-2/campbell.pdf.

Candido:1951:RBF

[Can51] G. Candido. A relationship between the fourth powers of the terms of the Fibonacci series. Scripta Mathematica, 17(3-4):230, 1951. ISSN 0036-9713.

Canessa:2003:TAFa

[Can03a] Enrique Canessa. Theory of analogous force on number sets.

ArXiv e-prints, 2003. URL http://arxiv.org/abs/cond-mat/
0307703.

Canessa:2003:TAFb

[Can03b] Enrique Canessa. Theory of analogous force on number sets. Physica A, Statistical Mechanics and its Applications, 328(1–2):44–52, October 2003. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0378437103005260.

Capocelli:1988:NFT

[Cap88] Renato M. Capocelli. A note on Fibonacci trees and the Zeckendorf representation of integers. Fibonacci Quarterly, 26(4):318– 324, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/26-4/capocelli.pdf.

Capocelli:1989:CAR

[Cap89] Renato Capocelli. Comments and additions to ‘robust transmission of unbounded strings using Fibonacci representations’. IEEE Transactions on Information Theory, 35(1):191–193, January 1989. CODEN IETTAW. ISSN 0018-9448 (print), 1557-9654 (electronic).

Carruccio:1939:LRC

[Car39] Ettore Carruccio. L’estrazione di radice cubica, mediante inserzione di due medie proporzionali fra due segmenti dati, in Leonardo Pisano. (Italian) [The extraction of the cube root, by inserting two mean proportions between two data segments, in Leonardo of Pisa]. Period. Mat., 19:189–197, 1939.

Carlitz:1963:FA

[Car63] L. Carlitz. A Fibonacci array. Fibonacci Quarterly, 1(2):17–28, April 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/1-2/carlitz.pdf.

Carlitz:1964:NFN

[Car64a] L. Carlitz. A note on Fibonacci numbers. Fibonacci Quarterly, 2(1):15–28, February 1964. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/2-1/carlitz.pdf.

Carlitz:1964:PDE

[Car64b] L. Carlitz. A partial difference equation related to the Fibonacci numbers. Fibonacci Quarterly, 2(3):185–196, October 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-3/carlitz.pdf.

Carlitz:1965:CPC

[Car65] L. Carlitz. The characteristic polynomial of a certain matrix of binomial coefficients. Fibonacci Quarterly, 3(2):81–89, April 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/3-2/carlitz.pdf.

Carlitz:1966:SBC

[Car66a] L. Carlitz. Some binomial coefficient identities. Fibonacci Quarterly, 4(4):323–331, December 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-4/
carlitz.pdf.

Carlitz:1966:SDC

[Car66b] L. Carlitz. Some determinants containing powers of Fibonacci numbers. Fibonacci Quarterly, 4(2):129–134, April 1966. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-2/carlitz.pdf.

Carlitz:1966:SOP

[Car66c] L. Carlitz. Some orthogonal polynomials related to Fibonacci numbers. Fibonacci Quarterly, 4(1):43–48, February 1966. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-1/carlitz.pdf.

Carlitz:1968:BN

[Car68a] L. Carlitz. Bernoulli numbers. Fibonacci Quarterly, 6(3):71–84, June 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/6-3/carlitz.pdf.

Carlitz:1968:FR

[Car68b] L. Carlitz. Fibonacci representations. Fibonacci Quarterly, 6 (4):193–220, October 1968. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/6-4/carlitz.pdf.

Carlitz:1969:GF

[Car69] L. Carlitz. Generating functions. Fibonacci Quarterly, 7(4):359– 393, November 1969. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/7-4/carlitz.pdf.

Carlitz:1970:FRI

[Car70a] L. Carlitz. Fibonacci representations — II. Fibonacci Quarterly, 8(2):113–134, March 1970. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/8-2/carlitz.pdf.

Carlitz:1970:SGF

[Car70b] L. Carlitz. Some generalized Fibonacci identities. Fibonacci Quarterly, 8(3):249–254, April 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-3/
carlitz.pdf.

Carlson:1970:DHT

[Car70c] John H. Carlson. Determination of Heronian triangles. Fibonacci Quarterly, 8(5):499–506, December 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-5/
carlson-a.pdf.

Carlitz:1971:RFF

[Car71a] L. Carlitz. Reduction formulas for Fibonacci summations. Fibonacci Quarterly, 9(5):449–466, December 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/9-5/carlitz-a.pdf.

Carlitz:1971:SSF

[Car71b] L. Carlitz. Some summation formulas. Fibonacci Quarterly, 9(1): 28–33, February 1971. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/9-1/carlitz.pdf.

Carlitz:1972:CPC

[Car72a] L. Carlitz. Concavity properties of certain sequences of numbers. Fibonacci Quarterly, 10(5):523–525, November 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-5/carlitz4-a.pdf.

Carlitz:1972:CCL

[Car72b] L. Carlitz. A conjecture concerning Lucas numbers. Fibonacci Quarterly, 10(5):526–??, November 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
5/carlitz5-a.pdf.

Carlitz:1972:EA

[Car72c] L. Carlitz. Enumeration of 3 × 3 arrays. Fibonacci Quarterly, 10(5):489–498, November 1972. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/10-5/carlitz2.
pdf.

Carlitz:1972:GFP

[Car72d] L. Carlitz. A generating function for partly ordered partitions. Fibonacci Quarterly, 10(2):157–162, February 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-2/carlitz.pdf.

Carlitz:1972:NSF

[Car72e] L. Carlitz. Note on summation formulas. Fibonacci Quarterly, 10(3):281–282, April 1972. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/10-3/carlitz-a.pdf.

Carlitz:1972:SPB

[Car72f] L. Carlitz. Sequences, paths, ballot numbers. Fibonacci Quarterly, 10(5):531–550, November 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
5/carlitz7.pdf.

Carlitz:1973:ETL

[Car73] L. Carlitz. Enumeration of two-line arrays. Fibonacci Quarterly, 11(2):113–130, April 1973. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/11-2/carlitz.

pdf. See corrigendum [CH74].

Carlitz:1974:FNZ

[Car74a] L. Carlitz. Fibonacci notes — 1. zero-one sequences and Fibonacci numbers of higher order. Fibonacci Quarterly, 12(1): 1–10, February 1974. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/12-1/carlitz.pdf.

Carlitz:1974:FNM

[Car74b] L. Carlitz. Fibonacci notes — 2. multiple generating functions. Fibonacci Quarterly, 12(2):179–185, April 1974. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-2/carlitz.pdf.

Carlitz:1974:FNF

[Car74c] L. Carlitz. Fibonacci notes — 3: q-Fibonacci numbers. Fibonacci Quarterly, 12(4):317–322, December 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
4/carlitz1.pdf.

Carlitz:1974:I

[Car74d] L. Carlitz. A q-identity. Fibonacci Quarterly, 12(4):369–372, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/12-4/carlitz2.pdf.

Carlitz:1975:FNF

[Car75a] L. Carlitz. Fibonacci notes — 4: q-Fibonacci polynomials. Fibonacci Quarterly, 13(2):97–102, April 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-2/
carlitz1.pdf.

Carlitz:1975:NSG

[Car75b] L. Carlitz. Note on some generating functions. Fibonacci Quarterly, 13(2):129–133, April 1975. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/13-2/carlitz3.
pdf.

Carlitz:1975:SIB

[Car75c] L. Carlitz. Some identities of Bruckman. Fibonacci Quarterly, 13(2):121–126, April 1975. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/13-2/carlitz2.pdf.

Carlitz:1976:RC

[Car76a] L. Carlitz. Restricted compositions. Fibonacci Quarterly, 14(3): 254–264, October 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-3/carlitz2.pdf.

Carlitz:1976:ST

[Car76b] L. Carlitz. The Saalschützian theorems. Fibonacci Quarterly, 14 (1):55–63, February 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-1/carlitz.pdf.

Carlitz:1976:SP

[Car76c] L. Carlitz. Set partitions. Fibonacci Quarterly, 14(4):327–342, November 1976. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/14-4/carlitz.pdf.

Carlitz:1976:SBS

[Car76d] L. Carlitz. Some binomial sums. Fibonacci Quarterly, 14(3):249– 253, October 1976. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/14-3/carlitz1.pdf.

Carlitz:1976:SSM

[Car76e] L. Carlitz. Some sums of multinomial coefficients. Fibonacci Quarterly, 14(5):427–438, December 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
5/carlitz.pdf.

Carlitz:1977:FNZ

[Car77a] L. Carlitz. Fibonacci notes — 5. zero-one sequences again. Fibonacci Quarterly, 15(1):49–56, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/15-1/carlitz1.pdf.

Carlitz:1977:FNG

[Car77b] L. Carlitz. Fibonacci notes 6. A generating function for Halsey’s Fibonacci function. Fibonacci Quarterly, 15(3):276–280, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/15-3/carlitz2.pdf.

Carlitz:1977:SSC

[Car77c] L. Carlitz. Some sums containing the greatest integer function. Fibonacci Quarterly, 15(1):78–84, February 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-1/carlitz2.pdf.

Carlitz:1978:RTO

[Car78a] L. Carlitz. Recurrences of the third order and related combinatorial identities. Fibonacci Quarterly, 16(1):11–18, February 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/16-1/carlitz.pdf.

Carlitz:1978:ECW

[Car78b] Leonard Carlitz. Enumeration of certain weighted sequences. Fibonacci Quarterly, 16(3):249–254, June 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-3/
carlitz4.pdf.

Carlitz:1978:EPS

[Car78c] Leonard Carlitz. Enumeration of permutations by sequences. Fibonacci Quarterly, 16(3):259–268, June 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-3/
carlitz6.pdf.

Carlitz:1978:NDS

[Car78d] Leonard Carlitz. The number of derangements of a sequence with given specification. Fibonacci Quarterly, 16(3):255–258, June 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/16-3/carlitz5.pdf.

Carlitz:1978:RSC

[Car78e] Leonard Carlitz. A recurrence suggested by a combinatorial problem. Fibonacci Quarterly, 16(3):227–242, June 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-3/carlitz2.pdf.

Carlitz:1978:SCF

[Car78f] Leonard Carlitz. Some classes of Fibonacci sums. Fibonacci Quarterly, 16(5):411–425, October 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
5/carlitz.pdf.

Carlitz:1978:SPR

[Car78g] Leonard Carlitz. Some polynomials related to Fibonacci and Eulerian numbers. Fibonacci Quarterly, 16(3):216–226, June 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-3/carlitz1.pdf.

Carlitz:1978:SRC

[Car78h] Leonard Carlitz. Some remarks on a combinatorial identity. Fibonacci Quarterly, 16(3):243–248, June 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-3/
carlitz3.pdf.

Carson:1978:FNP

[Car78i] Judithlynne Carson. Fibonacci numbers and pineapple phyllotaxy. Two-Year College Mathematics Journal, 9(3):132–136, June 1978. CODEN ???? ISSN 0049-4925 (print), 2325-9116

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1978.11974562.

Carlitz:1979:RCI

[Car79a] Leonard Carlitz. Restricted compositions II. Fibonacci Quarterly, 17(4):321–327, December 1979. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/17-4/carlitz.
pdf.

Carlitz:1979:RMC

[Car79b] Leonard Carlitz. Restricted multipartite compositions. Fibonacci Quarterly, 17(3):220–227, October 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
3/carlitz.pdf.

Carlitz:1980:FNZ

[Car80a] L. Carlitz. Fibonacci notes 6. Zero-one sequence once more. Fibonacci Quarterly, 18(2):177–184, April 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-2/
carlitz2.pdf.

Carlitz:1980:NPG

[Car80b] L. Carlitz. The number of permutations with a given number of sequences. Fibonacci Quarterly, 18(4):347–352, December 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-4/carlitz.pdf.

Carlitz:1980:WSNa

[Car80c] L. Carlitz. Weighted Stirling numbers of the first and second kind — I. Fibonacci Quarterly, 18(2):147–162, April 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-2/carlitz1.pdf.

Carlitz:1980:WSNb

[Car80d] L. Carlitz. Weighted Stirling numbers of the first and second kind — II. Fibonacci Quarterly, 18(3):242–257, October 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-3/carlitz.pdf.

Carlitz:1980:SRB

[Car80e] Leonard Carlitz. Some remarks on the Bell numbers. Fibonacci Quarterly, 18(1):66–72, February 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-
1/carlitz2.pdf.

Carlitz:1980:SRM

[Car80f] Leonard Carlitz. Some restricted multiple sums. Fibonacci Quarterly, 18(1):58–65, February 1980. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/18-1/carlitz1.
pdf.

Carlitz:1981:EPS

[Car81a] L. Carlitz. Enumeration of permutations by sequences — II. Fibonacci Quarterly, 19(5):398–405, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-5/carlitz-a.pdf.

Carlitz:1981:SGB

[Car81b] L. Carlitz. Some generalizations of a binomial identity conjectured by Hoggatt. Fibonacci Quarterly, 19(3):200–207, August 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-3/carlitz.pdf.

Carlitz:1984:ART

[Car84] L. Carlitz. An application of the reciprocity theorem for Dedekind sums. Fibonacci Quarterly, 22(3):266–269, August 1984. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-3/carlitz.pdf.

Carlin:1998:LVM

[Car98] Albert V. Carlin. A Layman’s view of music of the spheres. Fibonacci Quarterly, 36(1):65–??, February 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-1/carlin.pdf.

Carson:2007:PRR

[Car07] T. Richard Carson. Periodic recurrence relations and continued fractions. Fibonacci Quarterly, 45(4):357–361, November 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/45-4/carson.pdf.

Castellanos:1986:RCE

[Cas86] Dario Castellanos. Rapidly converging expansions with Fibonacci coefficients. Fibonacci Quarterly, 24(1):70–82, February 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-1/castellanos.pdf.

Castellanos:1989:GBF

[Cas89] Dario Castellanos. A generalization of Binet’s formula and some of its consequences. Fibonacci Quarterly, 27(5):424–438, November 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/27-5/castellanos.pdf.

Castellanos:1991:GRS

[Cas91a] Dario Castellanos. A generalization of a result of Shannon and Horadam. Fibonacci Quarterly, 29(1):57–58, February 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-1/castellanos.pdf.

Castellanos:1991:NBP

[Cas91b] Dario Castellanos. A note on Bernoulli polynomials. Fibonacci Quarterly, 29(2):98–102, May 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-2/
castellanos.pdf.

Catlin:1974:LBP

[Cat74a] Paul A. Catlin. A lower bound for the period of the Fibonacci series modulo M. Fibonacci Quarterly, 12(4):349–350, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/12-4/catlin1.pdf.

Catlin:1974:DSO

[Cat74b] Paul A. Catlin. On the divisors of second-order recurrences. Fibonacci Quarterly, 12(2):175–178, April 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-2/
catlin.pdf.

Catlin:1974:MR

[Cat74c] Paul A. Catlin. On the multiplication of recurrences. Fibonacci Quarterly, 12(4):365–367, December 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
4/catlin2.pdf.

Catan:2008:LFA

[Cat09] Joanne Catan. Letter to Fibonacci Association — 2008/2009 subscription. Fibonacci Quarterly, 46/47(1):2–4, February 2008/ 2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers/46_47-1/2008membersletter.pdf.

Cavior:1968:TPS

[Cav68] Stephen R. Cavior. A theorem on power sums. Fibonacci Quarterly, 6(2):157–160, April 1968. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/6-2/cavior.
pdf.

Cavior:1977:UDR

[Cav77a] Stephan R. Cavior. Uniform distribution (mod m) of recurrent sequences. Fibonacci Quarterly, 15(3):265–267, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-3/cavior2.pdf.

Cavior:1977:UDP

[Cav77b] Stephan R. Cavior. Uniform distribution for prescribed moduli. Fibonacci Quarterly, 15(3):209–210, October 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-3/cavior1.pdf.

Church:1973:EGF

[CB73] C. A. Church and Marjorie Bicknell. Exponential generating functions for Fibonacci identities. Fibonacci Quarterly, 11(3): 275–281, October 1973. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/11-3/church.pdf.

Cook:2014:SPN

[CB14] Charles K. Cook and Michael R. Bacon. Some polygonal number summation formulas. Fibonacci Quarterly, 52(4):336–??, November 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/52-4/cook.pdf.

Cook:2010:MPS

[CBH10] Charles K. Cook, Michael R. Bacon, and Rebecca A. Hillman. The “magicness” of powers of some magic squares. Fibonacci Quarterly, 48(4):298–306, November 2010. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/48-
4/cook.pdf.

Cook:2017:HOB

[CBH17] Charles K. Cook, Michael R. Bacon, and Rebecca A. Hillman. Higher order boustrophedon transforms for certain well-known sequences. Fibonacci Quarterly, 55(3):201–??, August 2017. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-3/cook.pdf.

Capocelli:2003:RSF

[CC03] Renato M. Capocelli and Paul Cull. Rounding the solutions of Fibonacci-like difference equations. Fibonacci Quarterly, 41(2): 133–141, May 2003. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/41-2/capocelli.pdf.

Chern:2014:FNC

[CC14] Shane Chern and Alvin Cui. Fibonacci numbers close to a power of 2. Fibonacci Quarterly, 52(4):344–??, November 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-4/chern.pdf.

Chan:2016:FMP

[CC16] Eunice Y. S. Chan and Robert M. Corless. Fibonacci–Mandelbrot polynomials and matrices. ACM Communications in Computer Algebra, 50(4):155–157, December 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Chu:2020:TRI

[CC20] Hùng Vi˙ ê Chu and Lân Khánh Chu. A twist of a Ramanujan identity. Fibonacci Quarterly, 58(4):351–??, November 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/58-4/chu2.pdf.

Chuan:2000:SFW

[CCC00] Wai-Fong Chuan, Chih-Hao Chang, and Yen-Liang Chang. Suffixes of Fibonacci word patterns. Fibonacci Quarterly, 38(5): 432–439, November 2000. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/38-5/chuan2.pdf.

Cui:2022:SRRa

[CCD^{+}22] Hao Cui, Xiaoyu Cui, Sophia C. Davis, Irfan Durmić, Qingcheng Hu, Lisa Liu, Steven J. Miller, Fengping Ren, Alicia Smith Reina, and Eliel Sosis. Sums of reciprocals of recurrence relations. Fibonacci Quarterly, 60(5):111–??, December 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-5/cui.pdf.

Chen:2019:GBZ

[CCG^{+}19] Eric Chen, Robin Chen, Lucy Guo, Cindy Jiang, Steven J. Miller, Joshua M. Siktar, and Peter Yu. Gaussian behavior in Zeckendorf decompositions from lattices. Fibonacci Quarterly, 57(3):201–??, August 2019. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/57-3/chen.pdf.

Chung:1989:FCP

[CCL89] Kuo Liang Chung, Wen Chin Chen, and Ferng-Ching Lin. Fast computation of periodic continued fractions. Information Processing Letters, 33(2):67–72, November 10, 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Callaghan:2005:BFM

[CCT05] Joseph Callaghan, John J. Chew III, and Stephen M. Tanny. On the behavior of a family of meta-Fibonacci sequences. SIAM Journal on Discrete Mathematics, 18(4):794–824, 2005. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL

http://epubs.siam.org/sam-bin/dbq/article/42139.

Cai:2015:PNF

[CCZ15] Tianxin Cai, Deyi Chen, and Yong Zhang. Perfect numbers and Fibonacci primes (I). International Journal of Number Theory (IJNT), 11(1):159–169, February 2015. ISSN 1793-0421 (print),

1793-7310 (electronic). URL https://www.worldscientific.
com/doi/10.1142/S1793042115500098.

Cull:1978:KTR

[CD78] Paul Cull and Jeffery De Curtins. Knight’s tour revisited. Fibonacci Quarterly, 16(3):276–284, June 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-3/
cull.pdf.

Chazelle:1987:ICO

[CD87] B. Chazelle and D. P. Dobkin. Intersection of convex objects in two and three dimensions. Journal of the Association for Computing Machinery, 34(1):1–27, January 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http:/

/www.acm.org/pubs/toc/Abstracts/0004-5411/24036.html;
 https://www.math.utah.edu/pub/tex/bib/fibquart.bib.

Cosgrave:2013:PRQ

[CD13] John B. Cosgrave and Karl Dilcher. Pairs of reciprocal quadratic congruences involving primes. Fibonacci Quarterly, 51(2):98–??, May 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/51-2/cosgrave.pdf.

Cusenza:2021:WSM

[CDH^{+}21] Anna Cusenza, Aidan Dunkelberg, Kate Huffman, Dianhui Ke, Daniel Kleber, Steven J. Miller, Clayton Mizgerd, Vashisth Tiwari, Jingkai Ye, and Xiaoyan Zheng. Winning strategy for multiplayer and multialliance Zeckendorf games. Fibonacci Quarterly, 59(4):308–??, November 2021. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/59-4/
cusenza.pdf.

Cusenza:2022:BZG

[CDH^{+}22] Anna Cusenza, Aidan Dunkelberg, Kate Huffman, Dianhui Ke, Micah McClatchey, Steven J. Miller, Clayton Mizgerd, Vashisth Tiwari, Jingkai Ye, and Xiaoyan Zheng. Bounds on Zeckendorf games. Fibonacci Quarterly, 60(1):57–??, February 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-1/cusenza.pdf.

Cangul:2010:DE

[CDL^{+}10] Ismail Naci Cangul, Musa Demirci, Florian Luca, Ákos Pintér, and Gökhan Soydan. On the Diophantine equation x^{2} +2^{a} ·11^{b} = y^{n}. Fibonacci Quarterly, 48(1):39–46, February 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/48-1/cangul.pdf.

Colucci:2000:FLN

[CDM00] Luca Colucci, Ottavio D’Antona, and Carlo Mereghetti. Fibonacci and Lucas numbers as cumulative connection constants. Fibonacci Quarterly, 38(2):157–164, May 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-2/colucci.pdf.

Coleman:2006:PFS

[CDM^{+}06] Deidra A. Coleman, Christopher J. Dugan, Robert A. McEwen, Clifford A. Reiter, and Tran T. Tang. Periods of (q, r)-Fibonacci sequences and elliptic curves. Fibonacci Quarterly, 44(1):59–70, February 2006. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/44-1/coleman.pdf.

Cahill:2002:FD

[CDNN02] Nathan D. Cahill, John R. D’Errico, Darren A. Narayan, and Jack Y. Narayan. Fibonacci determinants. College Mathematics Journal, 33(3):221–225, May 2002. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2002.11921945.

Cahill:2003:CFF

[CDS03] Nathan D. Cahill, John R. D’Errico, and John P. Spence. Complex factorizations of the Fibonacci and Lucas numbers. Fibonacci Quarterly, 41(1):13–19, February 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/41-1/cahill.pdf.

Collins:2013:BWC

[CDW13] Alexander Collins, Charles Dedrickson, and Hua Wang. Binary words, n-color compositions and bisection of the Fibonacci numbers. Fibonacci Quarterly, 51(2):130–??, May 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/51-2/collins.pdf.

Cashwell:1966:FS

[CE66] F. D. Cashwell and C. J. Everett. Fibonacci spaces. Fibonacci Quarterly, 4(2):97–115, April 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-2/
cashwell.pdf.

Cerimele:1977:SMH

[Cer77] B. J. Cerimele. Summation of multiparameter harmonic series. Fibonacci Quarterly, 15(2):116, 144, 149, 172, April 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-2/cerimele-a.pdf.

Cerin:2017:CLI

[Cer17] Zvonko Cerin. On Candido like identities. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/55-5/cerin.pdf.

Cereceda:2021:BF

[Cer21] José Luis Cereceda. Bernoulli and Faulhaber. Fibonacci Quarterly, 59(2):145–??, May 2021. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/59-2/
cereceda.pdf.

Carlitz:1970:SFL

[CF70] L. Carlitz and H. H. Ferns. Some Fibonacci and Lucas identities. Fibonacci Quarterly, 8(1):61–73, February 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-1/carlitz.pdf.

Chatterjee:2019:NPW

[CFG19] Krishnendu Chatterjee, Hongfei Fu, and Amir Kafshdar Goharshady. Non-polynomial worst-case analysis of recursive programs. ACM Transactions on Programming Languages and Systems, 41 (4):20:1–20:??, November 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL https://dl.acm.org/

ft_gateway.cfm?id=3339984.

Catral:2014:GZT

[CFH^{+}14] Minerva Catral, Pari Ford, Pamela Harris, Steven J. Miller, and Dawn Nelson. Generalizing Zeckendorf’s theorem: The Kentucky sequence. Fibonacci Quarterly, 52(5):68–??, December 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/52-5/Catral.pdf.

Catral:2016:LDA

[CFH^{+}16] Minerva Catral, Pari L. Ford, Pamela E. Harris, Steven J. Miller, and Dawn Nelson. Legal decompositions arising from non-positive linear recurrences. Fibonacci Quarterly, 54(4):348– ??, November 2016. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/54-4/catral.pdf.

Catral:2017:NBL

[CFH^{+}17] Minerva Catral, Pari L. Ford, Pamela E. Harris, Steven J. Miller, Dawn Nelson, Zhao Pan, and Huanzhong Xu. New behavior in legal decompositions arising from non-positive linear recurrences. Fibonacci Quarterly, 55(3):252–??, August 2017. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-3/catral.pdf.

Ching:2022:PCDa

[CFL^{+}22] Hsin-Yun Ching, Rigoberto Flórez, F. Luca, Antara Mukherjee, and J. C. Saunders. Primes and composites in the determinant Hosoya triangle. Fibonacci Quarterly, 60(5):56–??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/60-5/ching.pdf.

Church:1967:LPS

[CG67] C. A. Church, Jr. and H. W. Gould. Lattice point solution of the generalized problem of Terquem and an extension of Fibonacci numbers. Fibonacci Quarterly, 5(1):59–68, February 1967. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-1/church.pdf.

Conway:1996:BN

[CG96] John H. Conway and Richard K. Guy. The Book of Numbers. Copernicus (a division of Springer-Verlag New York, Inc.), 175 Fifth Avenue, New York, NY 10010, USA, 1996. ISBN 0-387-97993-X. ix + 310 pp. LCCN QA241 .C6897 1996. US$35.00.

Chen:2003:SCB

[CG03] Zhuo Chen and John Greene. Some comments on Baillie–PSW pseudoprimes. Fibonacci Quarterly, 41(4):334–344, August 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-4/chen.pdf.

Chung:2020:SFW

[CGS20] Fan Chung, Ron Graham, and Sam Spiro. Slow Fibonacci walks. Journal of Number Theory, 210(??):142–170, May 2020. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X19303488.

Carlitz:1969:SPF

[CH69] L. Carlitz and J. A. H. Hunter. Sums of powers of Fibonacci and Lucas numbers. Fibonacci Quarterly, 7(5):467–473, December 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/7-5/carlitz.pdf.

Carlitz:1974:CET

[CH74] L. Carlitz and Margaret Hodel. Corrigendum to: “Enumeration of Two-Line Arrays”. Fibonacci Quarterly, 12(3):265–266, October 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/12-3/corrigendum.pdf. See [Car73].

Carlitz:1978:GEN

[CH78] L. Carlitz and V. E. Hoggatt, Jr. Generalized Eulerian numbers and polynomials. Fibonacci Quarterly, 16(2):138–146, April 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-2/carlitz.pdf.

Cohen:1983:ESE

[CH83] M. E. Cohen and D. L. Hudson. On exponential series expansions and convolutions. Fibonacci Quarterly, 21(2):111–117, May 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-2/cohen.pdf.

Cull:1989:CFN

[CH89] Paul Cull and James L. Holloway. Computing Fibonacci numbers quickly. Information Processing Letters, 32(3):143–149, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cassidy:1994:SPF

[CH94] Charles Cassidy and Bernard R. Hodgson. On some properties of Fibonacci diagonals in Pascal’s triangle. Fibonacci Quarterly, 32(2):145–152, May 1994. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/32-2/cassidy.pdf.

Chang:2002:EPR

[CH02] Ching-Jua Chang and Chuang-Wei Ha. Eulerian polynomials and related explicit formulas. Fibonacci Quarterly, 40(5):399– 404, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/40-5/chang.pdf.

Chuan:2005:LFI

[CH05] Wai-Fong Chuan and Hui-Ling Ho. Locating factors of the infinite Fibonacci word. Theoretical Computer Science, 349(3):429–442, December 16, 2005. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Chang:2006:IIB

[CH06] Ching-Hua Chang and Chung-Wei Ha. On identities involving Bernoulli and Euler polynomials. Fibonacci Quarterly, 44(1):39– 45, February 2006. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/44-1/chang.pdf.

Chakerian:1973:GRG

[Cha73] G. D. (Don) Chakerian. The golden ratio and a Greek crisis. Fibonacci Quarterly, 11(2):195–200, April 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-2/
chakerian.pdf.

Charalambides:1981:CFN

[Cha81] Ch. A. Charalambides. Central factorial numbers and related expansions. Fibonacci Quarterly, 19(5):451–455, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-5/charalambides.pdf.

Charalambides:1982:ECC

[Cha82] Ch. A. Charalambides. On the enumeration of certain compositions and related sequences of numbers. Fibonacci Quarterly, 20(2):135–145, May 1982. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/20-2/charalambides.
pdf.

Chang:1984:NAN

[Cha84a] Derek K. Chang. A note on Apery numbers. Fibonacci Quarterly, 22(2):178–180, May 1984. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/22-2/chang.pdf.

Charalambides:1984:WSO

[Cha84b] Ch. A. Charalambides. On weighted Stirling and other related numbers and some combinatorial applications. Fibonacci Quarterly, 22(4):296–309, November 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-
4/charalambides.pdf.

Chang:1986:HOF

[Cha86a] Derek K. Chang. Higher-order Fibonacci sequences modulo M. Fibonacci Quarterly, 24(2):138–139, May 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/24-2/chang1.pdf.

Chang:1986:FBS

[Cha86b] Derek K. Chang. On Fibonacci binary sequences. Fibonacci Quarterly, 24(2):178–179, May 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-2/
chang2.pdf.

Chang:1986:FAT

[Cha86c] Derek K. Chang. On Fibonacci k-ary trees. Fibonacci Quarterly, 24(3):258–262, August 1986. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/24-3/chang.pdf.

Chang:1990:CFB

[Cha90] Derek K. Chang. On circular Fibonacci binary sequences. Fibonacci Quarterly, 28(1):28–30, February 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/28-1/chang.pdf.

Charalambides:1991:LNP

[Cha91] Ch. A. Charalambides. Lucas numbers and polynomials of order K and the length of the longest circular success run. Fibonacci Quarterly, 29(4):290–297, November 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-
4/charalambides.pdf.

Chamberland:2000:FSC

[Cha00] Marc Chamberland. Families of solutions of a cubic Diophantine equation. Fibonacci Quarterly, 38(3):250–253, June 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-3/chamberland.pdf.

Chastkofsky:2001:SEA

[Cha01] Leonard Chastkofsky. The subtractive Euclidean algorithm and Fibonacci numbers. Fibonacci Quarterly, 39(4):320–323, August 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/39-4/chastkofsky.pdf.

Chan:2005:AGR

[Cha05] Hei-Chi Chan. The asymptotic growth rate of random Fibonacci type sequences. Fibonacci Quarterly, 43(3):243–255, August 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers/43-3/paper43-3-7.pdf.

Chan:2006:T

[Cha06a] Hei-Chi Chan. π in terms of ϕ. Fibonacci Quarterly, 44(2): 141–144, May 2006. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/44-2/chan.pdf.

Chan:2006:AGR

[Cha06b] Hei-Chi Chan. The asymptotic growth rate of random Fibonacci type sequences II. Fibonacci Quarterly, 44(1):73–84, February 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/44-1/chan.pdf.

Chan:2007:AFF

[Cha07] Hei-Chi Chan. From Andrews’ formula for the Fibonacci numbers to the Rogers–Ramanujan identities. Fibonacci Quarterly, 45(3): 221–229, August 2007. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/45-3/chan.pdf.

Chan:2008:MTF

[Cha09] Hei-Chi Chan. Machin-type formulas expressing π in terms of ϕ. Fibonacci Quarterly, 46/47(1):32–37, February 2008/2009. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/46_47-1/chan.pdf.

Chamberland:2012:FTP

[Cha12] Marc Chamberland. Finite trigonometric product and sum identities. Fibonacci Quarterly, 50(3):217–??, August 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-3/chamberland.pdf.

Chen:2004:CEN

[Che04] Kwang-Wu Chen. Congruences for Euler numbers. Fibonacci Quarterly, 42(2):128–140, May 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-2/
quartchenkwang02_2004.pdf.

Chen:2012:EAI

[Che12] Kwang-Wu Chen. Extensions of an amazing identity of Ramanujan. Fibonacci Quarterly, 50(3):227–??, August 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-3/chen.pdf.

Curchin:1985:QDP

[CHF85] Leonard Curchin and Roger Herz-Fischler. De quand date le premier rapprochement entre la suite de Fibonacci et la division en extreme et moyenne raison?. (French) [When was the first reconciliation between the Fibonacci series and division in extreme and mean reason?]. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 28(2):129–138, July 1985. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Chia:1986:NCN

[Chi86] G. L. Chia. A note concerning the number of odd-order magic squares. Fibonacci Quarterly, 24(4):328–331, November 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-4/chia.pdf.

Choi:2005:IST

[CHKS05] Gyoung-Sik Choi, Suk-Geun Hwang, Ik-Pyo Kim, and Bryan L. Shader. (±1)-invariant sequences and truncated Fibonacci sequences. Linear Algebra and its Applications, 395(1):303–312, January 15, 2005. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Cai:2019:LBL

[CHL19] Zhaodong Cai, A. J. Hildebrand, and Junxian Li. A local Benford Law for a class of arithmetic sequences. International Journal of Number Theory (IJNT), 15(3):613–638, April 2019. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042119500325.

Chong:1982:LWL

[Cho82] Poh Ken Chong. The life and work of Leonardo of Pisa. Menemui Mat., 4(2):60–66, ???? 1982. ISSN 0126-9003.

Chow:1991:NCF

[Cho91] Timothy Chow. A new characterization of the Fibonacci-free partition. Fibonacci Quarterly, 29(2):174–179, May 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-2/chow.pdf.

Chrysaphinou:1984:NCI

[Chr84] Ourania Chrysaphinou. A note on the cycle indicator. Fibonacci Quarterly, 22(4):350–353, November 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-
4/chrysaphinou.pdf.

Christensen:2012:GFN

[Chr12] Sören Christensen. Generalized Fibonacci numbers and Blackwell’s renewal theorem. Statistics & Probability Letters, 82 (9):1665–1668, September 2012. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016771521200168X.

Carlitz:1972:SFR

[CHS72] L. Carlitz, V. E. Hoggatt, and Richard Scoville. Some functions related to Fibonacci and Lucas representations. Lecture Notes in Mathematics, 251:71–102, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058786/.

Corcino:2006:AGS

[CHT06] Roberto B. Corcino, Leetsch Charles Hsu, and Evelyn L. Tan. A q-analogue of generalized Stirling numbers. Fibonacci Quarterly, 44(2):154–165, May 2006. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/44-2/corcino.pdf.

Church:1970:ECT

[Chu70] C. A. Church, Jr. On the enumeration of certain triangular arrays. Fibonacci Quarterly, 8(3):235–241, April 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-3/church.pdf.

Church:1971:CTD

[Chu71] C. A. Church, Jr. Combinations and their duals. Fibonacci Quarterly, 9(5):505–511, December 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-5/
church.pdf.

Church:1974:LPF

[Chu74] C. A. Church, Jr. Lattice paths and Fibonacci and Lucas numbers. Fibonacci Quarterly, 12(4):336–338, December 1974. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-4/church.pdf.

Church:1985:CPS

[Chu85] C. A. Church, Jr. Combinatorial proof for a sorting problem identity. Fibonacci Quarterly, 23(4):366–368, November 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-4/church.pdf.

Chu:1990:AIS

[Chu90] W. C. Chu. An algebraic identity and some partial convolutions. Fibonacci Quarterly, 28(3):252–254, August 1990. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/28-3/chu.pdf.

Chuan:1992:FW

[Chu92] Wai-Fong Chuan. Fibonacci words. Fibonacci Quarterly, 30(1): 68–76, February 1992. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/30-1/chuan.pdf.

Chun:1996:GLP

[Chu96] Shen Ze Chun. GCD and LCM power matrices. Fibonacci Quarterly, 34(4):290–297, August 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-
4/chun.pdf.

Chu:2010:FPS

[Chu10] Wenchang Chu. Fibonacci polynomials and Sylvester determinant of tridiagonal matrix. Applied Mathematics and Computation, 216(3):1018–1023, April 1, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Chu:2020:RNF

[Chu20] Hùng Vi˙ êt Chu. Representation of \frac{1}{2}(F_{n} − 1)(F_{n+1} − 1) and \frac{1}{2}(F_{n} − 1)(F_{n+2} − 1). Fibonacci Quarterly, 58(4):334–??, November 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/58-4/chu1.pdf.

Chu:2021:GZD

[Chu21a] Hùng Vi{ê \atop .}t Chu. On generalized Zeckendorf decompositions and generalized golden strings. Fibonacci Quarterly, 59(3):254–??, August 2021. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/59-3/chu.pdf.

Chu:2021:PSF

[Chu21b] Hùng Vi{ê \atop .}t Chu. Partial sums of the Fibonacci sequence. Fibonacci Quarterly, 59(2):132–??, May 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/59-
2/chu1.pdf.

Chu:2021:VSC

[Chu21c] Hùng Vi{ê \atop .}t Chu. Various sequences from counting subsets. Fibonacci Quarterly, 59(2):152–??, May 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/59-
2/chu2.pdf.

Chu:2023:NFS

[Chu23] Hùng Vi˙ êt Chu. A note on the Fibonacci sequence and Schreiertype sets. Fibonacci Quarterly, 61(3):194–??, August 2023. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-3/chu.pdf.

Cigler:2003:FP

[Cig03] Johann Cigler. q-Fibonacci polynomials. Fibonacci Quarterly, 41 (1):31–40, February 2003. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/41-1/cigler.pdf.

Ciocci:2018:BRT

[Cio18] Argante Ciocci. Book review: The Twelfth Chapter of Fibonacci’s Liber Abaci in its 1202 version, Bollettino di Storia delle Scienze Matematiche, by Enrico Giusti (ed.). Nuncius, 33(1):137–139, ???? 2018. CODEN ???? ISSN 0394-7394 (print), 1825-3911

(electronic). URL http://booksandjournals.brillonline.
com/content/journals/10.1163/18253911-03301006.

Carlip:1996:SCL

[CJ96] Walter Carlip and Eliot Jacobson. On the stability of certain Lucas sequences modulo 2^{k}. Fibonacci Quarterly, 34(4):298–305, August 1996. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/34-4/carlip.pdf.

Chow:2024:VFP

[CJ24] Sam Chow and Owen Jones. On the variance of the Fibonacci partition function. Journal of Number Theory, 257(??):341– 353, April 2024. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X23002329.

Corcino:2017:GMP

[CJCK17] Roberto B. Corcino, Hassan Jolany, Cristina B. Corcino, and Takao Komatsu. On generalized multi poly-Euler polynomials. Fibonacci Quarterly, 55(1):41–??, February 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/55-1/corcino.pdf.

Carroll:1994:DTT

[CJS94] Dana Carroll, Eliot Jacobson, and Lawrence Somer. Distribution of two-term recurrence sequences mod P^{e}. Fibonacci Quarterly, 32(3):260–265, June 1994. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/32-3/carroll.pdf.

Carlip:1998:PPN

[CJS98] W. Carlip, Eliot Jacobson, and Lawrence Somer. Pseudoprimes, perfect numbers, and a problem of Lehmer. Fibonacci Quarterly, 36(4):361–371, August 1998. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/36-4/carlip.pdf.

Cooper:1988:PAF

[CK88] Curtis N. Cooper and Robert E. Kennedy. A partial asymptotic formula for the Niven numbers. Fibonacci Quarterly, 26(2):163– 168, May 1988. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/26-2/cooper.pdf.

Cooper:1993:CNN

[CK93] Curtis Cooper and Robert E. Kennedy. On consecutive Niven numbers. Fibonacci Quarterly, 31(2):146–151, May 1993. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-2/cooper.pdf.

Cooper:1995:PRJ

[CK95] Curtis Cooper and Robert E. Kennedy. Proof of a result by Jarden by generalizing a proof by Carlitz. Fibonacci Quarterly, 33(4):304–310, August 1995. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/33-4/cooper.pdf.

Christopher:1997:BGT

[CK97] Peter R. Christopher and John W. Kennedy. Binomial graphs and their spectra. Fibonacci Quarterly, 35(1):48–53, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/35-1/christopher.pdf.

Caldwell:2010:SPC

[CK10] Chris K. Caldwell and Takao Komatsu. Some periodicities in the continued fraction expansions of Fibonacci and Lucas Dirichlet series. Fibonacci Quarterly, 48(1):47–55, February 2010. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/48-1/caldwell.pdf.

Chen:2012:HSM

[CK12] Hongwei Chen and Chris Kennedy. Harmonic series meets Fibonacci sequence. College Mathematics Journal, 43(3):237–243, May 2012. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
4169/college.math.j.43.3.237.

Cook:2016:SIS

[CK16] Charles K. Cook and Takao Komatsu. Some identities for sequences of binomial sums of generalized Fibonacci numbers. Fibonacci Quarterly, 54(2):105–??, May 2016. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/54-
2/cook.pdf.

Campbell:2024:BSI

[CK24] John M. Campbell and Emrah Kılıç. Binomial sums involving second-order linearly recurrent sequences. Fibonacci Quarterly, 62(1):57–??, February 2024. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/62-1/
campbell.pdf.

Castro:2011:DNP

[CKMR11] Aline Castro, Sandi Klavzar, Michel Mollard, and Yoomi Rho. On the domination number and the 2-packing number of Fibonacci cubes and Lucas cubes. Computers and Mathematics and Applications, 61(9):2655–2660, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0898122111001647.

Cooper:1998:CSF

[CKR98] C. Cooper, R. E. Kennedy, and M. Renberg. On certain sums of functions of base B expansions. Fibonacci Quarterly, 36(5): 407–415, November 1998. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/36-5/cooper.pdf.

Chu:2011:PSP

[CL11] Wenchang Chu and Nadia N. Li. Power sums of Pell and Pell– Lucas polynomials. Fibonacci Quarterly, 49(2):139–150, May 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/49-2/chu.pdf.

Chu:2012:SFL

[CL12] Wenchang Chu and Nadia N. Li. Subsequences of Fibonacci and Lucas polynomials with geometric subscripts. Fibonacci Quarterly, 50(1):27–??, February 2012. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/50-1/
chu.pdf.

Clark:1986:FNE

[Cla86] Dean S. Clark. Fibonacci numbers as expected values in a game of chance. Fibonacci Quarterly, 24(3):263–267, August 1986. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-3/clark.pdf.

Claussen:2008:TER

[Cla08] Jens Christian Claussen. Time evolution of the rule 150 cellular automaton activity from a Fibonacci iteration. Journal of Mathematical Physics, 49(6):062701, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v49/i6/
p062701_s1.

Chuan:2012:FWP

[CLHY12] Wai-Fong Chuan, Fang-Yi Liao, Hui-Ling Ho, and Fei Yu. Fibonacci word patterns in two-way infinite Fibonacci words. Theoretical Computer Science, 437(1):69–81, June 15, 2012. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397512001612.

Chu:2022:ZRP

[CLM22] Hùng Vi{ê \atop .}t Chu, David C. Luo, and Steven J. Miller. On Zeckendorf related partitions using the Lucas sequence. Fibonacci Quarterly, 60(2):111–??, May 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/60-
2/chu.pdf.

Chor:1986:ANT

[CLRS86] Benny Chor, Charles E. Leiserson, Ronald L. Rivest, and James B. Shearer. An application of number theory to the organization of raster-graphics memory. Journal of the Association for Computing Machinery, 33(1):86–104, January 1986. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X

(electronic). URL http://www.acm.org/pubs/toc/Abstracts/
0004-5411/4800.html;   https://www.math.utah.edu/pub/
tex/bib/fibquart.bib.

Cai:2008:HAF

[CLX08] Jin-Yi Cai, Pinyan Lu, and Mingji Xia. Holographic algorithms by Fibonacci gates and holographic reductions for hardness. In IEEE [IEE08], pages 644–653. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN QA76 .S95 2008. URL http://ieeexplore.ieee.

org/servlet/opac?punumber=4690923. IEEE Computer Soci-

ety order number P3436.

Cai:2013:HAF

[CLX13] Jin-Yi Cai, Pinyan Lu, and Mingji Xia. Holographic algorithms by Fibonacci gates. Linear Algebra and its Applications, 438(2):690–707, January 15, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0024379511001625.

Cseh:1989:RRP

[CM89] László Cseh and Imre Merényi. Recurrence relations for a power series. Fibonacci Quarterly, 27(2):153–155, May 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-2/cseh.pdf.

Cerruti:2001:CNS

[CM01] Umberto Cerruti and Gabriella Margaria. Counting the number of solutions of equations in groups by recurrences. Fibonacci Quarterly, 39(4):290–298, August 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-
4/cerruti.pdf.

Carlip:2007:CGI

[CM07] Walter Carlip and Martina Mincheva. Component growth of iteration graphs under the squaring map modulo p^{k}. Fibonacci Quarterly, 45(3):239–246, August 2007. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/45-
3/carlip.pdf.

Cureg:2010:NRS

[CM10] Edgardo Cureg and Arunava Mukherjea. Numerical results on some generalized random Fibonacci sequences. Computers and Mathematics and Applications, 59(1):233–246, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122109005136.

Chaves:2014:DER

[CM14] Ana Paula Chaves and Diego Marques. A Diophantine equation related to the sum of squares of consecutive k-generalized Fibonacci numbers. Fibonacci Quarterly, 52(1):70–??, February 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/52-1/chaves.pdf.

Chaves:2015:DER

[CM15] Ana Paula Chaves and Diego Marques. A Diophantine equation related to the sum of powers of two consecutive generalized Fibonacci numbers. Journal of Number Theory, 156(??):1–14, November 2015. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X15001298.

Cox:2017:PGS

[CM17] Danielle Cox and Karyn McLellan. A problem on generation sets containing Fibonacci numbers. Fibonacci Quarterly, 55(2): 105–??, May 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/55-2/cox.pdf.

Cooper:2017:IRH

[CMM^{+}17] Curtis Cooper, Steven Miller, Peter J. C. Moses, Murat Sahin, and Thotsaporn Thanatipanonda. On identities of Ruggles, Horadam, Howard, and Young. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/55-5/cooper.pdf.

Cuccaro:1994:TTQ

[CMP94] Steven A. Cuccaro, Michael Mascagni, and Daniel V. Pryor. Techniques for testing the quality of parallel pseudorandom number generators. Technical report SRC-TR-94-128, Supercomputing Research Center: IDA, Lanham, MD, USA, October 4, 1994. 6 pp.

Chaoui:2002:AMC

[CMR02] F. Chaoui, M. Mouline, and M. Rachidi. Application of Markov chains properties to ∞-generalized Fibonacci sequences. Fibonacci Quarterly, 40(5):453–459, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-5/chaoui.pdf.

Curtin:2012:LHF

[CMS12] Brian Curtin, Ena Salter Michael, and David Stone. Lucas’ hyperbolas for Fibonacci vectors. Fibonacci Quarterly, 50(1):51–??, February 2012. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/50-1/curtin.pdf.

Chu:2020:HOF

[CMX20] Hùng Vi{ê \atop .}t Chu, Steven J. Miller, and Zimu Xiang. Higher order Fibonacci sequences from generalized Schreier sets. Fibonacci Quarterly, 58(3):249–??, August 2020. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/58-3/
chu.pdf.

Cahill:2004:FLN

[CN04] Nathan D. Cahill and Darren A. Narayan. Fibonacci and Lucas numbers as tridiagonal matrix determinants. Fibonacci Quarterly, 42(3):216–221, August 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
3/quartcahill03_2004.pdf.

Cohn:1964:LE

[Coh64a] John H. E. Cohn. Letter to the editor. Fibonacci Quarterly, 2(2): 108–??, April 1964. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/2-2/cohn1.pdf.

Cohn:1964:SFN

[Coh64b] John H. E. Cohn. Square Fibonacci numbers, etc. Fibonacci Quarterly, 2(2):109–113, April 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-2/
cohn2.pdf.

Cohn:1967:TRM

[Coh67] John H. E. Cohn. On m-tic residues modulo n. Fibonacci Quarterly, 5(4):305–318, December 1967. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/5-4/cohn.
pdf.

Cohn:1970:CDS

[Coh70] Ernst M. Cohn. Complete Diophantine solution of the Pythagorean triple. Fibonacci Quarterly, 8(4):402–405, October 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-4/cohn.pdf.

Cohn:1972:PNT

[Coh72] Ernst M. Cohn. Pell number triples. Fibonacci Quarterly, 10 (4):403–404, October 1972. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/10-4/cohn-a.pdf.

Cohn:1973:LE

[Coh73] Ernst M. Cohn. Letter to the editor. Fibonacci Quarterly, 11(1): 62–??, February 1973. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/11-1/letter-d.pdf.

Cohen:1978:OPN

[Coh78] G. L. Cohen. On odd perfect numbers. Fibonacci Quarterly, 16 (6):523–526, December 1978. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/16-6/cohen.
pdf.

Cohn:1979:GTF

[Coh79] Harvey Cohn. Growth types of Fibonacci and Markoff. Fibonacci Quarterly, 17(2):178–183, April 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-2/
cohn.pdf.

Cohen:1982:NQN

[Coh82] Graeme L. Cohen. The nonexistence of quasiperfect numbers of certain forms. Fibonacci Quarterly, 20(1):81–84, February 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-1/cohen.pdf.

Cohn:1990:PD

[Coh90] J. H. E. Cohn. Palindromic differences. Fibonacci Quarterly, 28(2):113–120, May 1990. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/28-2/cohn.pdf.

Cohn:1991:RSI

[Coh91] J. H. E. Cohn. Recurrent sequences including N. Fibonacci Quarterly, 29(1):30–36, February 1991. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/29-1/cohn.pdf.

Cohen:1999:AVT

[Coh99] Henri Cohen. On the 2-adic valuations of the truncated polylogarithm series. Fibonacci Quarterly, 37(2):117–121, May 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-2/cohen.pdf.

Colman:1983:GMD

[Col83] W. J. A. Colman. A general method for determining a closed formula for the number of partitions of the integer n into m positive integers for small values of m. Fibonacci Quarterly, 21(4):

272–284, November 1983. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/21-4/colman.pdf.

Collings:1987:CRN

[Col87a] Bruce Jay Collings. Compound random number generators. Journal of the American Statistical Association, 82(398):525– 527, June 1987. CODEN JSTNAL. ISSN 0162-1459 (print),

1537-274X (electronic). URL http://www.jstor.org/stable/
2289456.

Colman:1987:UBG

[Col87b] W. J. A. Colman. An upper bound for the general restricted partition problem. Fibonacci Quarterly, 25(1):38–44, February 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/25-1/colman.pdf.

Colman:1988:CSP

[Col88a] W. J. A. Colman. On certain semi-perfect cuboids. Fibonacci Quarterly, 26(1):54–57, February 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
1/colman.pdf.

Colman:1988:SOC

[Col88b] W. J. A. Colman. Some observations on the classical cuboid and its parametric solutions. Fibonacci Quarterly, 26(4):338– 343, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/26-4/colman.pdf.

Colman:1994:PCG

[Col94] W. J. A. Colman. A perfect cuboid in Gaussian integers. Fibonacci Quarterly, 32(3):266–268, June 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-3/
colman.pdf.

Comtet:1979:MGB

[Com79] Louis Comtet. A multinomial generalization of a binomial identity. Fibonacci Quarterly, 17(2):108–110, April 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/17-2/comtet.pdf.

Consul:1976:SFD

[Con76] P. C. Consul. Some factorable determinants. Fibonacci Quarterly, 14(2):171–172, April 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-2/consul.pdf.

Cooper:1984:AGF

[Coo84] Curtis Cooper. Application of a generalized Fibonacci sequence. College Mathematics Journal, 15(2):145–147, March 1984. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/00494925.1984.11972766.

Cooper:2006:CFP

[Coo06] Joshua N. Cooper. Continued fractions with partial quotients bounded in average. Fibonacci Quarterly, 44(4):297–301, November 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/44-4/cooper.pdf.

Cook:2010:RFI

[Coo10] Charles K. Cook. A report on the Fourteenth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 48(3):195–196, August 2010. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/48-3/
ConfRep14.pdf.

Cook:2014:RR

[Coo14a] Charles K. Cook. RIT redux. Fibonacci Quarterly, 52(3):194–??, August 2014. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Announcements/ConferenceReport16.pdf.

Cooper:2014:ASS

[Coo14b] Curtis Cooper. Algebraic statements similar to those in Ramanujan’s “Lost Notebook”. Fibonacci Quarterly, 52(5):91–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Papers1/52-5/Cooper.pdf.

Cook:2016:CCR

[Coo16] Charles K. Cook. Caen and cheese: A report on the Seventeenth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 54(3):194–??, August 2016. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Announcements/ConferenceReport17.pdf.

Cooper:2019:SHD

[Coo19] Curtis Cooper. Some high degree generalized Fibonacci identities. Fibonacci Quarterly, 57(5):42–??, December 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-5/cooper.pdf.

Coppenbarger:2018:IMS

[Cop18] Matthew E. Coppenbarger. Iterations of a modified Sisyphus function. Fibonacci Quarterly, 56(2):130–??, May 2018. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/56-2/coppenbarger.pdf.

Corley:1989:CFE

[Cor89] H. W. Corley. The convolved Fibonacci equation. Fibonacci Quarterly, 27(3):283–284, June 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-3/
corley.pdf.

Costello:2002:NLS

[Cos02] Patrick Costello. A new largest Smith number. Fibonacci Quarterly, 40(4):369–371, August 2002. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/40-
4/costello.pdf.

Cox:1987:AAR

[Cox87] Michael Cox. An alternative approach to recursive Fibonacci sequences. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 19(3):38–40, September 1, 1987. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Crandall:1979:SLT

[CP79] R. E. Crandall and M. A. Penk. A search for large twin prime pairs. Mathematics of Computation, 33(145):383–388, January 1979. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Cacoullos:1984:MSN

[CP84] T. Cacoullos and H. Papageorgiou. Multiparameter Stirling and C-numbers: Recurrences and applications. Fibonacci Quarterly, 22(2):119–133, May 1984. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/22-2/cacoullos.pdf.

Callan:2003:IME

[CP03] David Callan and Helmut Prodinger. An involutory matrix of eigenvectors. Fibonacci Quarterly, 41(2):105–107, May 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-2/callan.pdf.

Cristea:2008:EWN

[CP09] Ligia L. Cristea and Helmut Prodinger. q-enumeration of up-down words by number of rises. Fibonacci Quarterly, 46/47(2):126–134, May 2008/2009. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/46_47-
2/cristea.pdf.

Cox:1970:SUC

[CPH70] Nannette Cox, John W. Phillips, and V. E. Hoggatt, Jr. Some universal counterexamples. Fibonacci Quarterly, 8(3):242–247, April 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/8-3/cox.pdf.

Carlitz:1972:TAS

[CR72] L. Carlitz and D. P. Roselle. Triangular arrays subject to MacMahon’s conditions. Fibonacci Quarterly, 10(6):591–598, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/10-6/carlitz-a.pdf.

Campbell:1975:NFT

[CR75] C. M. Campbell and E. F. Robertson. A note on Fibonacci type groups. Canadian Mathematical Bulletin = Bulletin canadien de mathématiques, 18(??):173–176, ???? 1975. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Castellanos:1993:RCA

[CR93] Dario Castellanos and William E. Rosenthal. Rational Chebyshev approximations of analytic functions. Fibonacci Quarterly, 31(3): 205–215, August 1993. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/31-3/castellanos.pdf.

Clark:2010:CMD

[CR10] Tyler Clark and Tom Richmond. Collections of mutually disjoint convex subsets of a totally ordered set. Fibonacci Quarterly, 48 (1):77–79, February 2010. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/48-1/clark.pdf.

Crenshaw:1970:SRO

[Cre70] Joe R. Crenshaw. Some remarks on the ordering of general Fibonacci sequences. Fibonacci Quarterly, 8(5):516–521, December 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-5/crenshaw.pdf.

Creely:1986:SCS

[Cre86] Joseph W. Creely. Some combinatorial sequences. Fibonacci Quarterly, 24(3):209–220, August 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-
3/creely.pdf.

Creely:1987:LTN

[Cre87] Joseph W. Creely. The length of a two-number game. Fibonacci Quarterly, 25(2):174–179, May 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-2/
creely.pdf.

Creely:1988:LTN

[Cre88] Joseph W. Creely. The length of a three-number game. Fibonacci Quarterly, 26(2):141–143, May 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-2/
creely.pdf.

Creely:1989:GPU

[Cre89] Joseph W. Creely. Generating partitions using a modified greedy algorithm. Fibonacci Quarterly, 27(3):257–258, June 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-3/creely.pdf.

Crilly:2007:MIY

[Cri07] A. J. (Tony) Crilly. 50 mathematical ideas you really need to know. Quercus, London, UK, thirteenth edition, 2007. ISBN 1-84724-008-9 (UK), 1-84724-147-6 (USA). 208 pp. LCCN QA39.3 .C75 2007.

Cringanu:2014:AEC

[Crî14] Jenica Crînganu. On approximating Euler’s constant. Fibonacci Quarterly, 52(4):318–??, November 2014. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/52-
4/cringanu.pdf.

Cross:1975:FMM

[Cro75] Donald C. Cross. Fibonacci multi-multigrades. Fibonacci Quarterly, 13(3):211–212, October 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-
3/cross.pdf.

Cross:1976:MGF

[Cro76] Donald Cross. A more general Fibonacci multigrade. Fibonacci Quarterly, 14(1):24–??, February 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
1/cross-a.pdf.

Carlitz:1975:ENO

[CS75a] L. Carlitz and Richard Scoville. Eulerian numbers and operators. Fibonacci Quarterly, 13(1):71–83, February 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-1/carlitz.pdf.

Carlitz:1975:NWS

[CS75b] L. Carlitz and Richard Scoville. A note on weighted sequences. Fibonacci Quarterly, 13(4):303–306, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-4/carlitz.pdf.

Carlitz:1977:ZOS

[CS77] L. Carlitz and Richard Scoville. Zero-one sequences and Fibonacci numbers. Fibonacci Quarterly, 15(3):246–253, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/15-3/carlitz1.pdf.

Cohen:1979:SEW

[CS79] M. E. Cohen and H. Sun. On some extensions of the Wang– Carlitz identity. Fibonacci Quarterly, 17(4):299–305, December 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/17-4/cohen.pdf.

Creutz:1980:CIE

[CS80] Michael Creutz and R. M. Sternheimer. On the convergence of iterated exponentiation — I. Fibonacci Quarterly, 18(4):341–346, December 1980. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/18-4/creutz.pdf.

Clarke:1981:SCF

[CS81a] J. H. Clarke and A. G. Shannon. Some constraints on Fermat’s Last Theorem. Fibonacci Quarterly, 19(4):375–376, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/19-4/clarke.pdf.

Cohen:1981:GFD

[CS81b] M. E. Cohen and H. S. Sun. On generating functions and double series expansions. Fibonacci Quarterly, 19(1):69–73, February 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-1/cohen.pdf.

Cohen:1981:SEM

[CS81c] M. E. Cohen and H. S. Sun. On some extensions of the Meixner– Weisner generating functions. Fibonacci Quarterly, 19(5):422– 425, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/19-5/cohen.pdf.

Creutz:1981:CIE

[CS81d] Michael Creutz and R. M. Sternheimer. On the convergence of iterated exponentiation — II. Fibonacci Quarterly, 19(4):326– 335, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/19-4/creutz.pdf.

Creutz:1982:CIE

[CS82] Michael Creutz and R. M. Sternheimer. On the convergence of iterated exponentiation — III. Fibonacci Quarterly, 20(1):7–11, February 1982. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/20-1/creutz.pdf.

Clarke:1985:SGL

[CS85] J. H. Clarke and A. G. Shannon. Some generalized Lucas sequences. Fibonacci Quarterly, 23(2):120–125, May 1985. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-2/clarke.pdf.

Cohen:1989:NCW

[CS89] G. L. Cohen and S. L. Segal. A note concerning those n for which Φ(n) + 1 divides n. Fibonacci Quarterly, 27(3):285–286, June 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/27-3/cohen.pdf.

Cavicchioli:1992:CMS

[CS92] Alberto Cavicchioli and Fulvia Spaggiari. The classification of 3manifolds with spines related to Fibonacci groups. Lecture Notes in Mathematics, 1509:50–78, 1992. CODEN LNMAA2. ISBN 3-540-55195-6 (print), 3-540-46772-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0087501/.

Cook:1995:SAP

[CS95] Roger Cook and David Sharpe. Sums of arithmetic progressions. Fibonacci Quarterly, 33(3):218–220, June 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-3/cook.pdf.

Cohen:1998:HS

[CS98] G. L. Cohen and R. M. Sorli. Harmonic seeds. Fibonacci Quarterly, 36(5):386–390, November 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-
5/cohen.pdf.

Cohen:2001:HSE

[CS01] G. L. Cohen and R. M. Sorli. Harmonic seeds: Errata. Fibonacci Quarterly, 39(1):4–??, February 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-1/
errata.pdf.

Carlip:2003:ESM

[CS03] Walter Carlip and Lawrence Somer. The existence of special multipliers of second-order recurrence sequences. Fibonacci Quarterly, 41(2):156–168, May 2003. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/41-2/carlip.
pdf.

Cavicchioli:2004:VFT

[CS04] Alberto Cavicchioli and Fulvia Spaggiari. Varieties of Fibonacci type. Fibonacci Quarterly, 42(3):256–265, August 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-3/quartcavicchioli03_2004.pdf.

Clark:2012:RLA

[CS12] Liat Clark and Ian Steadman. The rich legacy of Alan Turing. Wired UK Web site., June 18, 2012. URL http://www.wired.

com/wiredscience/2012/06/alan-turing-legacy/.

Chow:2021:FP

[CS21] Sam Chow and Tom Slattery. On Fibonacci partitions. Journal of Number Theory, 225(??):310–326, August 2021. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X21000846.

Carlitz:1972:APF

[CSH72a] L. Carlitz, Richard Scoville, and V. E. Hoggatt, Jr. Addendum to the paper “Fibonacci Representations”. Fibonacci Quarterly, 10(5):527–530, November 1972. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/10-5/carlitz6.
pdf.

Carlitz:1972:FR

[CSH72b] L. Carlitz, Richard Scoville, and V. E. Hoggatt, Jr. Fibonacci representations. Fibonacci Quarterly, 10(1):1–28, January 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-1/carlitz1.pdf.

Carlitz:1972:FRHa

[CSH72c] L. Carlitz, Richard Scoville, and V. E. Hoggatt, Jr. Fibonacci representations of higher order. Fibonacci Quarterly, 10(1):43– 70, January 1972. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/10-1/carlitz3-a.pdf.

Carlitz:1972:FRHb

[CSH72d] L. Carlitz, Richard Scoville, and V. E. Hoggatt, Jr. Fibonacci representations of higher order — II. Fibonacci Quarterly, 10(1): 71–80, January 1972. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/10-1/carlitz4-a.pdf.

Carlitz:1972:LR

[CSH72e] L. Carlitz, Richard Scoville, and V. E. Hoggatt, Jr. Lucas representations. Fibonacci Quarterly, 10(1):29–42, January 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-1/carlitz2-a.pdf.

Carlitz:1972:PR

[CSH72f] L. Carlitz, Richard Scoville, and V. E. Hoggatt, Jr. Pellian representations. Fibonacci Quarterly, 10(5):449–488, November 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-5/carlitz1.pdf.

Carlitz:1972:RSS

[CSH72g] L. Carlitz, Richard Scoville, and V. E. Hoggatt, Jr. Representations for a special sequence. Fibonacci Quarterly, 10(5):499–518, November 1972. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/10-5/carlitz3-a.pdf.

Curtin:2007:SFF

[CSS07] Brian Curtin, Ena Salter, and David Stone. Some formulae for the Fibonacci numbers. Fibonacci Quarterly, 45(2):171–180, May 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/45-2/curtin.pdf.

Calkin:2005:CLC

[CST05] Neil J. Calkin, John G. Stevens, and Diana M. Thomas. A characterization for the length of cycles of the N-number Ducci game. Fibonacci Quarterly, 43(1):53–59, February 2005. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/43-1/paper43-1-7.pdf.

Carlitz:1973:SAF

[CSV73] L. Carlitz, Richard Scoville, and Theresa Vaughan. Some arithmetic function related to Fibonacci numbers. Fibonacci Quarterly, 11(4):337–386, November 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
4/carlitz.pdf.

Cranston:1975:SRS

[CT75] Ben Cranston and Rick Thomas. Simplified recombination scheme for the Fibonacci buddy system. Communications of the Association for Computing Machinery, 18(6):331–332, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Coskun:2018:NBP

[CT18] Arzu Coskun and Necati Taskara. A note on the bi-periodic Fibonacci and Lucas matrix sequences. Applied Mathematics and Computation, 320(??):400–406, March 1, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300317306598.

Cullen:1976:SI

[Cul76] Theodore J. Cullen. A summation identity. Fibonacci Quarterly, 14(1):35–36, February 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-1/cullen.pdf.

Cull:2017:WTY

[Cul17] Paul Cull. What I tell you K times is true. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/55-5/cull.pdf.

Curl:1968:FNS

[Cur68] James C. Curl. Fibonacci numbers and the slow learner. Fibonacci Quarterly, 6(4):266–274, October 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/6-4/curl.pdf.

Cusick:1968:CIA

[Cus68] T. W. Cusick. On a certain integer associated with a generalized Fibonacci sequence. Fibonacci Quarterly, 6(2):117–126, April 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/6-2/cusick.pdf.

Coquet:1986:SFI

[CV86] Jean Coquet and Pascal Van Den Bosch. A summation formula involving Fibonacci digits. Journal of Number Theory, 22(2):139– 146, February 1986. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022314X8690065X.

Carlton:1988:DPS

[CV88] Mike Carlton and Peter Van Roy. Distributed Prolog system with and parallelism. IEEE Software, 5(1):43–51, January 1988. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Caragiu:2017:CSF

[CVZ17] Mihai Caragiu, Paul A. Vicol, and Mohammad Zaki. On Conway’s subprime function, a covering of N and an unexpected appearance of the Golden Ratio. Fibonacci Quarterly, 55(4):327–??, November 2017. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/55-4/caragiu.pdf.

Cohen:1985:ESR

[CW85] G. L. Cohen and R. J. Williams. Extensions of some results concerning odd perfect numbers. Fibonacci Quarterly, 23(1):70– 76, February 1985. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/23-1/cohen.pdf.

Caragiu:2003:MFS

[CW03] Mihai Caragiu and William Webb. On modular Fibonacci sets. Fibonacci Quarterly, 41(4):307–309, August 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-4/caragiu.pdf.

Campbell:2022:ZSR

[CW22] H. E. A. Campbell and D. L. Wehlau. Zigzag sequences and representations of integers. Fibonacci Quarterly, 60(3):220–??, August 2022. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/60-3/campbellH.pdf.

Carroll:1991:DCP

[CY91] Joseph E. Carroll and Ken Yanosko. The determination of a class of primitive integral triangles. Fibonacci Quarterly, 29(1): 3–6, February 1991. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/29-1/carroll.pdf.

Chuan:2000:EPP

[CY00] Wai-Fong Chuan and Fei Yu. Extraction problem of the Pell sequence. Fibonacci Quarterly, 38(5):425–431, November 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-5/chuan1.pdf.

Chuan:2007:TNE

[CY07] Wai-Fong Chuan and Fei Yu. Three new extraction formulae. Fibonacci Quarterly, 45(1):76–84, February 2007. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/45-1/chuan.pdf.

Chen:1981:FNS

[CZ81] Robert Chen and Alan Zame. Fibonacci numbers and stopping times. Fibonacci Quarterly, 19(2):127–130, April 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-2/chen.pdf.

Chu:2010:TMC

[CZ10] Wenchang Chu and Roberta R. Zhou. Two multiple convolutions on Fibonacci-like sequences. Fibonacci Quarterly, 48(1):80–84, February 2010. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/48-1/chu.pdf.

Caragiu:2011:DSA

[CZZ11] Mihai Caragiu, Alexandru Zaharescu, and Mohammad Zaki. On Ducci sequences with algebraic numbers. Fibonacci Quarterly, 49 (1):34–40, February 2011. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/49-1/caragiu.pdf.

Caragiu:2014:DSP

[CZZ14] Mihai Caragiu, Alexandru Zaharescu, and Mohammad Zaki. On Ducci sequences with primes. Fibonacci Quarterly, 52(1):32–??, February 2014. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/52-1/caragiu.pdf.

Davis:1979:GMH

[DA79] T. Antony Davis and Rudolf Altevogt. Golden mean of the human body. Fibonacci Quarterly, 17(4):340–343, December 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-4/davis-a.pdf.

Azevedo:1979:FN

[dAA79] J. C. de Almeida Azevedo. Fibonacci numbers. Fibonacci Quarterly, 17(2):162–164, April 1979. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/17-2/azevedo.
pdf.

Dafnis:2020:RBF

[Daf20] Spiros D. Dafnis. On the relation between Fibonacci and Lucas numbers. Fibonacci Quarterly, 58(5):111–??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/58-5/dafnis.pdf.

Damphousse:1989:APR

[Dam89] Pierre Damphousse. The arithmetic of powers and roots in Gl_{2}(C) and Sl_{2}(C). Fibonacci Quarterly, 27(5):386–401, November 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/27-5/damphousse.pdf.

Dantchev:1998:CFF

[Dan98] Stefan Dantchev. A closed form of the (2, F) generalizations of the Fibonacci sequence. Fibonacci Quarterly, 36(5):448–451, November 1998. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/36-5/dantchev.pdf.

Davis:1971:WFS

[Dav71] T. Antony Davis. Why Fibonacci sequence for palm leaf spirals? Fibonacci Quarterly, 9(3):237–244, May 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-3/
davis1.pdf.

Davis:1972:FNP

[Dav72] Basil Davis. Fibonacci numbers in physics. Fibonacci Quarterly, 10(6):659–660, December 1972. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/10-6/davis-a.
pdf.

Davis:1976:SRI

[Dav76] Basil Davis. Some remarks on initial digits. Fibonacci Quarterly, 14(1):13–14, February 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-1/davis.pdf.

Davis:1977:RAN

[Dav77] K. Joseph Davis. A result in analytic number theory. Fibonacci Quarterly, 15(2):164–165, April 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-2/
davis.pdf.

Davis:1982:GDP

[Dav82] K. Joseph Davis. A generalization of the Dirichlet product. Fibonacci Quarterly, 20(1):41–43, February 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/20-1/davis.pdf.

Davis:2014:PAS

[Dav14] Donald M. Davis. p-Adic Stirling numbers of the second kind. Fibonacci Quarterly, 52(3):226–??, August 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-3/davis.pdf.

Daykin:1969:CLS

[Day69a] D. E. Daykin. On the completeness of the Lucas sequence. Fibonacci Quarterly, 7(5):464–??, December 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/7-5/daykin1.pdf.

Daykin:1969:RNN

[Day69b] D. E. Daykin. Representation of natural numbers as sums of generalized Fibonacci numbers — II. Fibonacci Quarterly, 7(5): 494–510, December 1969. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/7-5/daykin2.pdf.

Daykin:1970:AAG

[Day70] D. E. Daykin. An addition algorithm for greatest common divisor. Fibonacci Quarterly, 8(4):347–349, October 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-4/daykin.pdf.

Dazheng:1999:FM

[Daz99] Lin Dazheng. Fibonacci matrices. Fibonacci Quarterly, 37(1): 14–20, February 1999. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/37-1/lin.pdf.

Dazheng:2002:FLQ

[Daz02] Lin Dazheng. Fibonacci–Lucas quasi-cyclic matrices. Fibonacci Quarterly, 40(3):280–286, June 2002. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/40-3/lin.
pdf.

Draim:1966:SPR

[DB66] N. A. Draim and Marjorie Bicknell. Sums of n-th powers of roots of a given quadratic equation. Fibonacci Quarterly, 4(2):170–178, April 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/4-2/draim.pdf.

Draim:1967:EWR

[DB67] N. A. Draim and Marjorie Bicknell. Equations whose roots as the n th powers of the roots of a given cubic equation. Fibonacci Quarterly, 5(3):267–274, October 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-3/
draim.pdf.

Draim:1970:SPR

[DB70] N. A. Draim and Marjorie Bicknell. Summation of powers of roots of special equations. Fibonacci Quarterly, 8(2):221–224, March 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-2/draim.pdf.

Davis:1971:FSA

[DB71] T. Antony Davis and T. K. Bose. Fibonacci system in aroids. Fibonacci Quarterly, 9(3):253–263, May 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-3/
davis2.pdf.

deBruyn:1995:F

[dB95] G. F. C. de Bruyn. Formulas a + a^{2}2^{p} + a^{3}3^{p} + ⋯ + a^{n}n^{p}. Fibonacci Quarterly, 33(2):98–103, May 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-2/
debruyn.pdf.

Deshpande:2001:DTP

[DB01] M. N. Deshpande and Ezra Brown. Diophantine triplets and the Pell sequence. Fibonacci Quarterly, 39(3):242–249, June 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-3/deshpande.pdf.

Dougherty-Bliss:2022:MCFa

[DB22] Robert Dougherty-Bliss. The meta-C-finite ansatz. Fibonacci Quarterly, 60(5):143–??, December 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/60-
5/dougherty-bliss.pdf.

deBruyn:1994:F

[dBdV94] G. F. C. de Bruyn and J. M. de Villiers. Formulas for 1 + 2^{p} + 3^{p} + ⋯ + n^{p}. Fibonacci Quarterly, 32(3):271–276, June 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-3/debruyn.pdf.

deBouvere:1985:FIC

[dBF85] Karel L. de Bouvère and Herta T. Freitag. First International Conference on Fibonacci Numbers and Their Applications: Two reports. Fibonacci Quarterly, 23(2):98–99, May 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/23-2/report.pdf.

deBouvere:1983:IEG

[dBL83] Karel L. de Bouvère and Regina E. Lathrop. Injectivity of extended generalized Fibonacci sequences. Fibonacci Quarterly, 21 (1):37–52, February 1983. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/21-1/debouvere.pdf.

Dutton:1993:NIS

[DCB93] R. Dutton, N. Chandrasekharam, and R. Brigham. On the number of independent sets of nodes in a tree. Fibonacci Quarterly, 31(2):98–104, May 1993. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/31-2/dutton.pdf.

Dresel:1965:FFN

[DD65] L. A. G. Dresel and D. E. Daykin. Factorization of 36 Fibonacci numbers F_{n} with n > 100. Fibonacci Quarterly, 3(3):232–??, October 1965. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/3-3/dresel.pdf.

Daykin:1967:IPF

[DD67] D. E. Daykin and L. A. G. Dresel. Identities for products of Fibonacci and Lucas numbers. Fibonacci Quarterly, 5(4):367– 369, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/5-4/daykin2.pdf.

Daykin:1970:FFN

[DD70] D. E. Daykin and L. A. G. Dresel. Factorization of Fibonacci numbers. Fibonacci Quarterly, 8(1):23–30, February 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-1/daykin-a.pdf.

Das:1987:RGF

[DD87] Sajal D. Das and Narsingh Deo. Rencontres graphs: a family of bipartitie graphs. Fibonacci Quarterly, 25(3):250–262, August 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/25-3/das.pdf.

Dimitrov:1995:FMM

[DD95] Vassil S. Dimitrov and Borislav D. Donevsky. Faster multiplication of medium large numbers via the Zeckendorf representation. Fibonacci Quarterly, 33(1):74–77, February 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-1/dimitrov.pdf.

Deshpande:2002:IPR

[DD02] M. N. Deshpande and Andrej Dujella. An interesting property of a recurrence related to the Fibonacci sequence. Fibonacci Quarterly, 40(2):157–160, May 2002. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/40-2/
deshpande.pdf.

DeKoninck:2003:NNN

[DD03] Jean-Marie DeKoninck and Nicolas Doyon. On the number of Niven numbers up to x. Fibonacci Quarterly, 41(5):431–440, November 2003. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/41-5/dekoninck.pdf.

Dence:2013:IRG

[DD13] Joseph B. Dence and Thomas P. Dence. On inverse relations for general Lucas sequences of polynomials. Fibonacci Quarterly, 51 (1):55–??, February 2013. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/51-1/dence.pdf.

Ddamulira:2020:CPE

[Dda20] Mahadi Ddamulira. On the x-coordinates of Pell equations that are products of two Lucas numbers. Fibonacci Quarterly, 58(1): 18–??, February 2020. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/58-1/ddamulira.pdf.

Demontigny:2014:GFF

[DDK^{+}14] Philippe Demontigny, Thao Do, Archit Kulkarni, Steven J. Miller, and Umang Varma. A generalization of Fibonacci fardifference representations and Gaussian behavior. Fibonacci Quarterly, 52(3):247–??, August 2014. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/52-3/
demontigny.pdf.

Rosa:1978:PPP

[de 78a] B. de La Rosa. Primes, powers, and partitions. Fibonacci Quarterly, 16(6):518–522, December 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
6/delarosa.pdf.

DeVita:1978:FIF

[De 78b] Joseph De Vita. Fibonacci, insects, and flowers. Fibonacci Quarterly, 16(4):315–317, August 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
4/devita.pdf.

DeBouvere:1981:FIG

[De 81a] Karel De Bouvère. Fibonacci induced groups and their hierarchies. Fibonacci Quarterly, 19(3):264–270, August 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-3/debouvere.pdf.

deLuca:1981:CPF

[de 81b] Aldo de Luca. A combinatorial property of the Fibonacci words. Information Processing Letters, 12(4):193–195, ???? ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deLuca:1995:DPF

[de 95] Aldo de Luca. A division property of the Fibonacci word. Information Processing Letters, 54(6):307–312, June 23, 1995. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dalenberg:2018:CFB

[DE18] Paul Dalenberg and Tom Edgar. Consecutive factorial base Niven numbers. Fibonacci Quarterly, 56(2):163–??, May 2018. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/56-2/dalenberg.pdf.

deBruijn:1974:EFS

[deB74] P. J. deBruijn. An extension of Fibonacci’s sequence. Fibonacci Quarterly, 12(3):251–258, October 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
3/debruijn.pdf.

DeCarli:1970:GFS

[DeC70] D. J. DeCarli. A generalized Fibonacci sequence over an arbitrary ring. Fibonacci Quarterly, 8(2):182–184, March 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-2/decarli-a.pdf.

DeCarli:1973:PRM

[DeC73] R. J. DeCarli. Periodicity over the ring of matrices. Fibonacci Quarterly, 11(5):466–468, December 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
5/decarli.pdf.

Deford:2013:CRG

[Def13] Daryl Deford. Counting rearrangements on generalized wheel graphs. Fibonacci Quarterly, 51(3):259–??, August 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/51-3/deford.pdf.

DeFord:2014:EDC

[DeF14] Daryl DeFord. Enumerating distinct chessboard tilings. Fibonacci Quarterly, 52(5):102–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers1/52-5/DeFord.pdf.

Deily:1966:TDC

[Dei66a] G. R. Deily. Terminal digit coincidences between Fibonacci numbers and their indices. Fibonacci Quarterly, 4(2):151–156, April 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/4-2/deily.pdf.

Deily:1966:LFF

[Dei66b] Gerard R. Deily. A logarithmic formula for Fibonacci numbers. Fibonacci Quarterly, 4(1):89–??, February 1966. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-1/deily.pdf.

Deininger:1972:FNW

[Dei72] Rolf A. Deininger. Fibonacci numbers and water pollution control. Fibonacci Quarterly, 10(3):299–300, April 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-3/deininger-a.pdf.

Dekking:2020:BPR

[Dek20a] F. Michel Dekking. Base phi representations and Golden mean beta-expansions. Fibonacci Quarterly, 58(1):38–??, February 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/58-1/dekking.pdf.

Dekking:2020:MWB

[Dek20b] Michel Dekking. Morphic words, Beatty sequences and integer images of the Fibonacci language. Theoretical Computer Science, 809(??):407–417, February 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397520300062.

Dekking:2021:HAT

[Dek21] F. Michel Dekking. How to add two natural numbers in base phi. Fibonacci Quarterly, 59(1):19–??, February 2021. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/59-1/dekking.pdf.

DeLeon:1970:FNC

[DeL70] Morris Jack DeLeon. The Fibonacci numbers considered as a Pisot sequence. Fibonacci Quarterly, 8(5):476–480, December 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-5/deleon.pdf.

DeLeon:1976:PEP

[DeL76] M. J. DeLeon. Pell’s equation and Pell number triples. Fibonacci Quarterly, 14(5):456–460, December 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
5/deleon.pdf.

DeLeon:1977:FPR

[DeL77] M. J. DeLeon. Fibonacci primitive roots and the period of the Fibonacci numbers modulo p. Fibonacci Quarterly, 15(4):353– 355, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/15-4/deleon.pdf.

DeLeon:1981:CFS

[DeL81] M. J. DeLeon. A characterization of the fundamental solutions to Pell’s equation u^{2} − Dv^{2} = C. Fibonacci Quarterly, 19(1): 4–5, February 1981. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/19-1/deleon-a.pdf.

DeLeon:1982:CW

[DeL82] M. J. DeLeon. The congruence x^{n} ≡ a (mod m), where (n, ϕ(m)) = 1. Fibonacci Quarterly, 20(2):129–134, May 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-2/deleon.pdf.

DeLeon:1984:PRC

[DeL84] M. J. DeLeon. n th power residues congruent to one. Fibonacci Quarterly, 22(4):358–364, November 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-
4/deleon.pdf.

DeLeon:1988:CFT

[DeL88] Morris Jack DeLeon. Carlitz four-tuples. Fibonacci Quarterly, 26 (3):224–232, August 1988. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/26-3/deleon.pdf.

DeLeon:1995:SRI

[DeL95] Morris Jack DeLeon. Sequences related to an infinite product expansion for the square root and cube root functions. Fibonacci Quarterly, 33(1):41–49, February 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-
1/deleon.pdf.

Demir:2008:TML

[Dem08] Ali Demir. Trisection method by k-Lucas numbers. Applied Mathematics and Computation, 198(1):339–345, April 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dence:1987:RGF

[Den87] Thomas P. Dence. Ratios of generalized Fibonacci sequences. Fibonacci Quarterly, 25(2):137–143, May 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-2/
dence.pdf.

Desmond:1971:IFS

[Des71] James E. Desmond. On iterative Fibonacci subscripts. Fibonacci Quarterly, 9(1):35–40, February 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-1/
desmond.pdf.

Desmond:1978:EPD

[Des78a] James E. Desmond. On the equality of periods of different moduli in the Fibonacci sequence. Fibonacci Quarterly, 16(1):86–87, February 1978. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/16-1/desmond-a.pdf.

Desmond:1978:ERA

[Des78b] James E. Desmond. On the existence of the rank of apparition of m in the Lucas sequence. Fibonacci Quarterly, 16(1):7–10, February 1978. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/16-1/desmond.pdf.

Desmond:1986:EFN

[Des86] James E. Desmond. Expansion of the Fibonacci numbers F_{mn+r} in the m th powers of Fibonacci or Lucas numbers. Fibonacci Quarterly, 24(3):194–208, August 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-
3/desmond.pdf.

Deshpande:1994:UEF

[Des94] M. N. Deshpande. An unexpected encounter with the Fibonacci numbers. Fibonacci Quarterly, 32(2):108–109, May 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-2/deshpande.pdf.

Davis:2017:NIB

[DES17] Chad Davis, Daniel Eloff, and Blair K. Spearman. Normal integral bases of a cyclic quintic field. Fibonacci Quarterly, 55(2): 152–??, May 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/55-2/davis.pdf.

DeTemple:1974:PA

[DeT74] Duane W. DeTemple. A pentagonal arch. Fibonacci Quarterly, 12 (3):235–236, October 1974. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/12-3/detemple.pdf.

DeTemple:1981:NAG

[DeT81] Duane W. DeTemple. A new angle on the geometry of the Fibonacci numbers. Fibonacci Quarterly, 19(1):35–38, February 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-1/detemple.pdf.

DeTemple:1992:TSP

[DeT92] Duane W. DeTemple. The triangle of smallest perimeter which circumscribes a semicircle. Fibonacci Quarterly, 30(3):274–??, August 1992. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/30-3/detemple.pdf.

Devaney:1999:MSF

[Dev99] Robert L. Devaney. The Mandelbrot set, the Farey tree, and the Fibonacci sequence. American Mathematical Monthly, 106 (4):289–302, April 1999. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Devlin:2011:MNF

[Dev11] Keith J. Devlin. The man of numbers: Fibonacci’s arithmetic revolution. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 2011. ISBN 0-8027-7812-7 (hardcover). viii + 183 + 8 pp. LCCN QA29.F5 D48 2011.

Devlin:2012:RML

[Dev12] Keith Devlin. Recreational mathematics in Leonardo of Pisa’s Liber abbaci. In Recreational Mathematics Colloquium II, pages 67–78. Assoc. Ludus, Lisbon, 2012.

Devlin:2017:FFQ

[Dev17] Keith Devlin. Finding Fibonacci: the quest to rediscover the forgotten mathematical genius who changed the world. Princeton University Press, Princeton, NJ, USA, 2017. ISBN 0-691-17486-5. vi + 241 pp. LCCN QA29.F5 D48 2017; QA29.F5 F56 2017.

Deza:1977:MNT

[Dez77] M. Deza. On minimal number of terms in representation of natural numbers as a sum of Fibonacci numbers. Fibonacci Quarterly, 15(3):237–??, October 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-3/deza.pdf.

DiPorto:1989:NFP

[DF89] Adina Di Porto and Piero Filipponi. Note on the Fibonacci pseudoprimes. Fibonacci Quarterly, 27(3):232–241, June 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-3/diporto.pdf.

DiPorto:1992:GSF

[DF92] Adina Di Porto and Piero Filipponi. Generating M-strong Fibonacci pseudoprimes. Fibonacci Quarterly, 30(4):339–343, November 1992. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/30-4/diporto.pdf.

daFonseca:2011:SOQ

[dF11] C. M. da Fonseca. Some open questions on the Fibonacci polytope. Linear Multilinear Algebra, 59(3):317–??, 2011. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

daFonseca:2014:USP

[dF14] Carlos M. da Fonseca. Unifying some Pell and Fibonacci identities. Applied Mathematics and Computation, 236(??):41–42, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300314004263.

Damir:2014:MLS

[DFLT14] Mohamed Taoufiq Damir, Bernadette Faye, Florian Luca, and Amadou Tall. Members of Lucas sequences whose Euler function is a power of 2. Fibonacci Quarterly, 52(1):3–??, February 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-1/damir.pdf.

DiPorto:1990:GFP

[DFM90] Adino Di Porto, Piero Filipponi, and Emilio Montolivo. On the generalized Fibonacci pseudoprimes. Fibonacci Quarterly, 28(4): 347–353, November 1990. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/28-4/diporto.pdf.

Dunton:1966:FEF

[DG66] M. Dunton and R. E. Grimm. Fibonacci on Egyptian fractions. Fibonacci Quarterly, 4(4):339–353, December 1966. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-4/dunton.pdf.

Dudeney:1967:PCP

[DG67] Henry Ernest Dudeney and Martin Gardner. 536 puzzles and curious problems. Scribner, New York, NY, USA, 1967. xii + 488 pp. LCCN ????

Downey:1984:FNR

[DG84] Peter J. Downey and Ralph E. Griswold. On a family of nested recurrences. Fibonacci Quarterly, 22(4):310–317, November 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-4/downey.pdf.

Drmota:1998:PSD

[DG98] Michael Drmota and Johannes Gajdosik. The parity of the sumof-digits-function of generalized Zeckendorf representations. Fibonacci Quarterly, 36(1):3–19, February 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-1/
drmota.pdf.

Dixon:2014:TTF

[DGMS14] John Dixon, Michael Goldenberg, Ben Mathes, and Justin Sukiennik. Triangular truncation of k-Fibonacci and k-Lucas circulant matrices. Linear Algebra and its Applications, 440 (??):177–187, January 1, 2014. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0024379513005818.

Dawson:1985:RFT

[DGNW85] R. Dawson, G. Gabor, R. Nowakowski, and D. Wiens. Random Fibonacci-type sequences. Fibonacci Quarterly, 23(2):169–176, May 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/23-2/dawson.pdf.

Driscoll:1988:RHA

[DGST88] James R. Driscoll, Harold N. Gabow, Ruth Shrairman, and Robert E. Tarjan. Relaxed heaps: An alternative to Fibonacci heaps with applications to parallel computation. Communications of the Association for Computing Machinery, 31(11): 1343–1354, November 1988. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0001-0782/50096.html;   https:
//www.math.utah.edu/pub/tex/bib/fibquart.bib.

Daykin:1967:BII

[DH67] D. E. Daykin and A. J. W. Hilton. Bases for infinite intervals of integers. Fibonacci Quarterly, 5(4):329–345, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-4/daykin1.pdf.

Daykin:1968:BIR

[DH68] D. E. Daykin and A. J. W. Hilton. Bases for intervals of real numbers. Fibonacci Quarterly, 6(6):335–349, December 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-6/daykin.pdf.

Davis:1976:EBF

[DH76] Bro. Basil Davis and V. E. Hoggatt, Jr. Exponentials and Bessel functions. Fibonacci Quarterly, 14(5):405, 418, 426, December 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/14-5/davis-a.pdf.

Das:2017:SNF

[DH17] Anupam Das and Bipan Hazarika. Some new Fibonacci difference spaces of non-absolute type and compact operators. Linear Multilinear Algebra, 65(12):2551–2573, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

DiPorto:1993:NEF

[Di 93] Adino Di Porto. Nonexistence of even Fibonacci pseudoprimes of the 1st kind. Fibonacci Quarterly, 31(2):173–177, May 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-2/diporto.pdf.

DiDomenico:2013:ISB

[Di 13] A. S. Di Domenico. Integer sequences that behave as Fibonacci– Lucas pairs. The Mathematical Gazette, 97(538):1–7, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Diffenderfer:2012:BBT

[Dif12] James Diffenderfer. A bijection between two classes of restricted compositions. Fibonacci Quarterly, 50(4):360–??, November 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/50-4/diffenderfer.pdf.

Dijkstra:1978:HFa

[Dij78] Edsger W. Dijkstra. In honour of Fibonacci. Circulated pri-

vately., 1978. URL http://www.cs.utexas.edu/users/EWD/
ewd06xx/EWD654.PDF.

Dijkstra:1979:HF

[Dij79] Edsger W. Dijkstra. In honour of Fibonacci. In Bauer and Broy [BB79], pages 49–50. CODEN LNCSD9. ISBN 0-387-09251-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6

.P75117. URL http://link.springer.com/chapter/10.1007/

BFb0014655. “The international summer school took place from July 26 to August 6, 1978, in Marktoberdorf … and was sponsored by the NATO Scientific Affairs Division under the 1978 Advanced Study Institutes programme.”.

Dijkstra:1981:FNL

[Dij81] Edsger W. Dijkstra. Fibonacci numbers and Leonardo numbers. Circulated privately., July 1981. URL http://www.cs.utexas.

edu/users/EWD/ewd07xx/EWD797.PDF.

Dijkstra:1990:FGC

[Dij90] Edsger W. Dijkstra. Fibonacci and the greatest common divisor. Circulated privately., April 1990. URL http://www.cs.utexas.

edu/users/EWD/ewd10xx/EWD1077.PDF.

Dijkstra:1996:FGC

[Dij96] Edsger W. Dijkstra. Fibonacci and the greatest common divisor. In Broy [Bro96], pages 7–10. ISBN 3-540-60947-4 (hardcover). LCCN QA76.9.D5 D38 1996. URL http://www.loc.

gov/catdir/enhancements/fy0812/96010788-d.html.

Dilcher:1987:GFP

[Dil87] Karl Dilcher. A generalization of Fibonacci polynomials and a representation of Gegenbauer polynomials of integer order. Fibonacci Quarterly, 25(4):300–303, November 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-4/dilcher.pdf.

Dilcher:1991:ZCC

[Dil91] Karl Dilcher. Zeros of certain cyclotomy-generated polynomials. Fibonacci Quarterly, 29(2):150–156, May 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/29-2/dilcher.pdf.

Dilcher:2000:HFF

[Dil00] Karl Dilcher. Hypergeometric functions and Fibonacci numbers. Fibonacci Quarterly, 38(4):342–363, August 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-4/dilcher.pdf.

Dov:1967:SPC

[DJ67] Dov and Moshe Jarden. Simultaneous prime and composite members in two arithmetic progressions. Fibonacci Quarterly, 5(3): 286–??, October 1967. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/5-3/jarden2.pdf.

Dov:1968:EIC

[DJ68] Dov and Moshe Jarden. On the existence of an infinitude of composite primitive divisors of second-order recurring sequences. Fibonacci Quarterly, 6(6):322–334, December 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/6-6/jarden1-a.pdf.

Djordjevic:1996:SPG

[Djo96] Gospava Djordjević. On some properties of generalized Hermite polynomials. Fibonacci Quarterly, 34(1):2–6, February 1996. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-1/djordjevic.pdf.

Djordjevic:1998:GCP

[Djo98] Gospava B. Djordjević. On a generalization of a class of polynomials. Fibonacci Quarterly, 36(2):110–117, May 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-2/djordjevic.pdf.

Djordjevic:2000:DSG

[Djo00a] Gospava B. Djordjević. Derivative sequences of generalized Jacobsthal and Jacobsthal–Lucas polynomials. Fibonacci Quarterly, 38(4):334–338, August 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-
4/djordjevic.pdf.

Djordjevic:2000:GJP

[Djo00b] Gospava B. Djordjević. Generalized Jacobthal polynomials. Fibonacci Quarterly, 38(3):239–243, June 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-3/
djordjevic.pdf.

Djordjevic:2001:GLP

[Djo01a] Gospava B. Djordjević. On the generalized Laguerre polynomials. Fibonacci Quarterly, 39(5):403–407, November 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-5/djordjevic.pdf.

Djordjevic:2001:SPP

[Djo01b] Gospava B. Djordjević. Some properties of partial derivatives of generalized Fibonacci and Lucas polynomials. Fibonacci Quarterly, 39(2):138–141, May 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-2/
djordjevic.pdf.

Djordjevic:2004:GFI

[Djo04] Gospava B. Djordjević. Generating functions of the incomplete generalized Fibonacci and generalized Lucas numbers. Fibonacci Quarterly, 42(2):106–113, May 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-2/
quartdjordjevic02_2004.pdf.

Djordjevic:2005:TOD

[Djo05a] Gospava B. Djordjević. On the K th-order derivative sequences of generalized Fibonacci and Lucas polynomials. Fibonacci Quarterly, 43(4):290–298, November 2005. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/43-
4/djordjevic.pdf.

Djordjevic:2005:SPS

[Djo05b] Gospava B. Djordjević. Some properties of the sequences C_{n,3} = C_{n−1,3} + C_{n−3,3} + r. Fibonacci Quarterly, 43(3):202–207, August 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers/43-3/paper43-3-2.pdf.

Dubner:1999:NFL

[DK99] Harvey Dubner and Wilfrid Keller. New Fibonacci and Lucas primes. Mathematics of Computation, 68(225):417– 427, January 1999. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://www.ams.

org/jourcgi/jour-pbprocess?fn=110&arg1=S0025-5718-99-
00981-3&u=/mcom/1999-68-225/.

Dekking:2017:CCF

[DK17] F. Michel Dekking and Michael S. Keane. On the conjugacy class of the Fibonacci dynamical system. Theoretical Computer Science, 668(??):59–69, March 15, 2017. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397517300506.

Debellevue:2018:FBF

[DK18] Michael Debellevue and Ekaterina Kryuchkova. Fractal behavior of the Fibonomial triangle modulo prime p, where the rank of apparition of p is p + 1. Fibonacci Quarterly, 56(2):113–??, May 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/56-2/debellevue.pdf.

Deza:1992:FPI

[DL92] Michel Deza and Monique Laurent. The Fibonacci and parachute inequalities for l_{1}-metrics. Fibonacci Quarterly, 30(1):54–61, February 1992. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/30-1/deza.pdf.

deLuca:1997:SSM

[dL97] Aldo de Luca. Standard Sturmian morphisms. Theoretical Computer Science, 178(1–2):205–224, May 30, 1997. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/
store/tcs/cas_sub/browse/browse.cgi?year=1997&volume=
178&issue=1-2&aid=2256.

Deshouillers:2011:DEF

[DL11] Jean-Marc Deshouillers and Florian Luca. On the distribution of the Euler functions with Fibonacci numbers. Fibonacci Quarterly, 49(2):102–109, May 2011. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/49-2/
deshouillers.pdf.

Dris:2016:NOP

[DL16] Jose Arnaldo B. Dris and Florian Luca. A note on odd perfect numbers. Fibonacci Quarterly, 54(4):291–??, November 2016. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/54-4/dris.pdf.

Ddamulira:2020:PPK

[DL20] Mahadi Ddamulira and Florian Luca. On the problem of Pillai with k-generalized Fibonacci numbers and powers of 3. International Journal of Number Theory (IJNT), 16(7): 1643–1666, August 2020. ISSN 1793-0421 (print), 1793-7310

(electronic). URL https://www.worldscientific.com/doi/
10.1142/S1793042120500876.

deLuca:1994:SCP

[dLM94] Aldo de Luca and Filippo Mignosi. Some combinatorial properties of Sturmian words. Theoretical Computer Science, 136(2): 361–385, December 29, 1994. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/tcs/cas_sub/browse/browse.
cgi?year=1994&volume=136&issue=2&aid=1676.

Ddamulira:2016:FNW

[DLR16] Mahadi Ddamulira, Florian Luca, and Mihaja Rakotomalala. Fibonacci numbers which are products of two Pell numbers.

Fibonacci Quarterly, 54(1):11–??, February 2016. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/54-1/ddamulira.pdf.

Dil:2008:SAH

[DM08] Ayhan Dil and István Mező. A symmetric algorithm for hyperharmonic and Fibonacci numbers. Applied Mathematics and Computation, 206(2):942–951, December 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dilcher:2010:DSS

[DM10] Karl Dilcher and Jeffrey L. Meyer. Dedekind sums and some generalized Fibonacci and Lucas sequences. Fibonacci Quarterly, 48(3):260–264, August 2010. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/48-3/dilcher.pdf.

Dafnis:2007:ROK

[DMP07] Spiros D. Dafnis, Frosso S. Makri, and Andreas N. Philippou. Restricted occupancy of s kinds of cells and generalized Pascal triangles. Fibonacci Quarterly, 45(4):347–356, November 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/45-4/dafnis.pdf.

Du:2016:DAFa

[DMR^{+}16] Chen Fei Du, Hamoon Mousavi, Eric Rowland, Luke Schaeffer, and Jeffrey Shallit. Decision algorithms for Fibonacci-automatic words, II: Related sequences and avoidability. Theoretical Computer Science, 657 (part B)(??):146–162, ???? 2016. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397516305199.

Dubeau:1997:WGF

[DMRS97] F. Dubeau, W. Motta, M. Rachidi, and O. Saeki. On weighted r-generalized Fibonacci sequences. Fibonacci Quarterly, 35(2): 102–110, May 1997. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/35-2/dubeau.pdf.

Du:2016:DAFb

[DMSS16] Chen Fei Du, Hamoon Mousavi, Luke Schaeffer, and Jeffrey Shallit. Decision algorithms for Fibonacci-automatic words, III: Enumeration and Abelian properties. International Journal of Foundations of Computer Science (IJFCS), 27(8):943–964, December 2016. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Damanik:2013:OPU

[DMY13] David Damanik, Paul Munger, and William N. Yessen. Orthogonal polynomials on the unit circle with Fibonacci Verblunsky coefficients, II. Applications. Journal of Statistical Physics, 153 (2):339–362, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-013-0830-9.

Darvasi:1996:RFB

[DN96] Gyula Darvasi and Mihály Nagy. On repetitions in frequency blocks of the generalized Fibonacci sequence u(3, 1) with u_{0} = u_{1} = 1. Fibonacci Quarterly, 34(2):176–180, May 1996. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-2/darvasi.pdf.

Dikici:2003:AFS

[DÖ03] Ramazan Dikici and Engin Özkan. An application of Fibonacci sequences in groups. Applied Mathematics and Computation, 136 (2–3):323–331, March 15, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dobson:2014:ACF

[Dob14] John Blythe Dobson. On the Andrews congruence for the Fibonacci quotient. Fibonacci Quarterly, 52(4):314–??, November 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/52-4/dobson.pdf.

Dodd:1984:NFF

[Dod84] Fred Dodd. The number field Q(\sqrt{5}) and the Fibonacci numbers. Fibonacci Quarterly, 22(2):171–177, May 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/22-2/dodd.pdf.

Acosta-de-Orozco:1996:LMP

[dOGC96] Maria T. Acosta de Orozco and Javier Gomez-Calderon. Local minimal polynomials over finite fields. Fibonacci Quarterly, 34 (2):139–143, May 1996. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/34-2/acosta.pdf.

Dordevic:1993:MFC

[Dor93] Gospava B. Dordevic. Mixed Fermat convolutions. Fibonacci Quarterly, 31(2):152–157, May 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-2/
dordevic.pdf.

Dordevic:1999:PRM

[Dor99] Gospava B. Dordević. Polynomials related to Morgan–Voyce polynomials. Fibonacci Quarterly, 37(1):61–66, February 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-1/dordevic.pdf.

Demir:2008:OLN

[DOU08a] Ali Demir, Nese Omur, and Yucel Turker Ulutas. Optimization by k-Lucas numbers. Applied Mathematics and Computation, 197 (1):366–371, March 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Demir:2008:PFS

[DOU08b] Ali Demir, Nese Omur, and Yucel Turker Ulutas. Parametrized Fibonacci search method with k-Lucas numbers. Applied Mathematics and Computation, 198(1):355–360, April 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Dressler:1978:IFT

[DP78a] Robert E. Dressler and Louis Pigno. Interpolation of Fourier transforms on sums of Fibonacci numbers. Fibonacci Quarterly, 16(3):193–194, June 1978. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/16-3/dressler1.pdf.

Dressler:1978:TMT

[DP78b] Robert E. Dressler and Louis Pigno. Topological, measure theoretic and analytic properties of the Fibonacci numbers. Fibonacci Quarterly, 16(3):195–197, June 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-3/
dressler2.pdf.

Dubeau:1995:TRN

[DP95] François Dubeau and Alain Pautasso. On triangular rectangular numbers. Fibonacci Quarterly, 33(3):244–248, June 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-3/dubeau.pdf.

Dafnis:2016:ISW

[DP16] Spiros D. Dafnis and Andreas N. Philippou. Infinite sums of weighted Fibonacci numbers of order k. Fibonacci Quarterly, 54 (2):149–??, May 2016. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/54-2/dafnis.pdf.

Dvorakova:2024:UBA

[DP24]  L’ubomíra Dvořáková and Edita Pelantová.  An upper

bound on asymptotic repetition threshold of balanced sequences via colouring of the Fibonacci sequence. Theoretical Computer Science, 995(??):??, May 12, 2024. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397524001051.

Deo:1983:PGT

[DQ83] Narsingh Deo and Michael Quinn. Pascal graphs and their properties. Fibonacci Quarterly, 21(3):203–214, August 1983. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-3/deo.pdf.

Draim:1964:EPT

[Dra64] N. A. Draim. Expansion of pi in terms of an infinite continued fraction with predictable terms. Fibonacci Quarterly, 2(4):290– ??, December 1964. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/2-4/draim.pdf.

Draim:1969:FCF

[Dra69] N. A. Draim. π in the form of a continued fraction with infinite terms. Fibonacci Quarterly, 7(3):275–276, October 1969. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-3/draim.pdf.

Dresel:1980:LE

[Dre80] L. A. G. Dresel. Letter to the editor. Fibonacci Quarterly, 18(1): 34–??, February 1980. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/18-1/letter1.pdf.

Dresel:1997:PRG

[Dre97] L. A. G. Dresel. On pseudoprimes related to generalized Lucas sequences. Fibonacci Quarterly, 35(1):35–42, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-1/dresel.pdf.

Drmota:1993:LDE

[Drm93] M. Drmota. On linear Diophantine equations and Fibonacci numbers. Journal of Number Theory, 44(3):315–327, July 1993. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X83710565.

Droubay:1995:PFW

[Dro95] Xavier Droubay. Palindromes in the Fibonacci word. Information Processing Letters, 55(4):217–221, August 25, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Drobot:2000:PFS

[Dro00] Vladimir Drobot. On primes in the Fibonacci sequences. Fibonacci Quarterly, 38(1):71–72, February 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/38-1/drobot.pdf.

Dziok:2011:CRC

[DRS11a] Jacek Dziok, Ravinder Krishna Raina, and Janusz Sokól. Certain results for a class of convex functions related to a shelllike curve connected with Fibonacci numbers. Computers and Mathematics and Applications, 61(9):2605–2613, May 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S089812211100157X.

Dziok:2011:CFR

[DRS11b] Jacek Dziok, Ravinder Krishna Raina, and Janusz Sokól. On α-convex functions related to shell-like functions connected with Fibonacci numbers. Applied Mathematics and Computation, 218(3):996–1002, October 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/

S0096300311000865. Special Issue in Honour of Hari M. Srivastava on his 70th birth anniversary.

Dalton:2011:SBG

[DRT11] Barnaby Dalton, Mustazee Rahman, and Stephen Tanny. Spotbased generations for meta-Fibonacci sequences. Experimental Mathematics, 20(2):129–137, ???? 2011. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http:/

/projecteuclid.org/euclid.em/1317924404.

DSouza:1989:TRN

[D’S89] Harry D’Souza. Trapping a real number between adjacent rationals. Fibonacci Quarterly, 27(4):369–371, August 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-4/dsouza.pdf.

Dubeau:1996:FMI

[DS96] F. Dubeau and A. G. Shannon. A Fibonacci model of infectious disease. Fibonacci Quarterly, 34(3):257–270, June 1996. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-3/dubeau.pdf.

DiScala:2004:IP

[DS04] Antonio J. Di Scala and Martín Sombra. Intrinsic palindromes. Fibonacci Quarterly, 42(1):76–81, February 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers1/42-1/quartscala01_2004.pdf.

Delo:2019:ETR

[DS19] Ben Delo and Filip Saidak. Euclid’s Theorem redux. Fibonacci Quarterly, 57(4):331–??, November 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/57-
4/delo.pdf.

Davis:2020:FCA

[DS20] Robert Davis and Greg Simay. Further combinatorics and applications of two-toned tilings. Fibonacci Quarterly, 58(4):300–??, November 2020. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/58-4/davis.pdf.

McNabb:1963:P

[dSM63] Sister Mary de Sales McNabb. Phyllotaxis. Fibonacci Quarterly, 1(4):57–60, December 1963. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/1-4/mcnabb-a.pdf.

Dudley:1971:GCD

[DT71] Underwood Dudley and Bessie Tucker. Greatest common divisors in altered Fibonacci sequences. Fibonacci Quarterly, 9(1):89–92, February 1971. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/9-1/dudley.pdf.

Duverney:2020:LII

[DT20] Daniel Duverney and Yohei Tachiya. Linear independence of infinite products generated by the Lucas numbers. Fibonacci Quarterly, 58(5):115–??, December 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/58-
5/duverney.pdf.

Du:1989:SMW

[Du89] Bau-Sen Du. A simple method which generates infinitely many congruence identities. Fibonacci Quarterly, 27(2):116–123, May 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/27-2/du.pdf.

Du:2000:OND

[Du00] Bau-Sen Du. Obtaining new dividing formulas n|Q(n) from the known ones. Fibonacci Quarterly, 38(3):217–222, June 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-3/du.pdf.

Dubeau:1989:GFN

[Dub89] François Dubeau. On r-generalized Fibonacci numbers. Fibonacci Quarterly, 27(3):221–228, June 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-3/
dubeau.pdf.

Dubeau:1993:RPR

[Dub93] François Dubeau. The rabbit problem revisited. Fibonacci Quarterly, 31(3):268–273, August 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-
3/dubeau.pdf.

Dujella:1996:GFN

[Duj96] Andrej Dujella. Generalized Fibonacci numbers and the problem of Diophantus. Fibonacci Quarterly, 34(2):164–174, May 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-2/dujella.pdf.

Dujella:1999:EOP

[Duj99] Andrej Dujella. An extension of an old problem of Diophantus and Euler. Fibonacci Quarterly, 37(4):312–314, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-4/dujella.pdf.

Dujella:2002:EOP

[Duj02] Andrej Dujella. An extension of an odd problem of Diophantus and Euler — II. Fibonacci Quarterly, 40(2):118–123, May 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-2/dujella.pdf.

Dular:2020:CSI

[Dul20] Bruno Dular. Cycles of sums of integers. Fibonacci Quarterly, 58 (2):126–??, May 2020. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/58-2/dular.pdf.

Duncan:1966:NEA

[Dun66] R. L. Duncan. Note on the Euclidean algorithm. Fibonacci Quarterly, 4(4):367–368, December 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-4/
duncan.pdf.

Duncan:1967:CEF

[Dun67a] Dewey C. Duncan. Chains of equivalent Fibonacci-wise triangle. Fibonacci Quarterly, 5(1):87–88, February 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/5-1/duncan.pdf.

Duncan:1967:AUD

[Dun67b] R. L. Duncan. An application of uniform distributions to the Fibonacci numbers. Fibonacci Quarterly, 5(2):137–140, April 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-2/duncan.pdf.

Duncan:1969:NID

[Dun69a] R. L. Duncan. Note on the initial digit problem. Fibonacci Quarterly, 7(5):474–475, December 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-5/
duncan.pdf.

Duncan:1969:DFI

[Dun69b] R. L. Duncan. On the density of the k-free integers. Fibonacci Quarterly, 7(2):140–142, April 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-2/
duncan.pdf.

Dunlap:1997:GRF

[Dun97] R. A. Dunlap. The golden ratio and Fibonacci numbers. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1997. ISBN 981-02-3264-0. vii + 162 pp. LCCN QA466 .D86 1997.

Dunne:2009:LDF

[Dun09] Edward Dunne. Long days on the Fibonacci clock. Mathematics Magazine, 82(2):127–134, 2009. CODEN MAMGA8. ISSN 0025-570X.

Duvanenko:1991:ERM

[Duv91] Victor J. Duvanenko. Efficiently raising matrices to an integer power. Dr. Dobb’s Journal of Software Tools, 16(6):86–87, 157, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Duvall:1988:PPC

[DV88] Paul Duvall and Theresa Vaughan. Pell polynomials and a conjecture of Mahon and Horadam. Fibonacci Quarterly, 26(4):344– 353, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/26-4/duvall.pdf.

Duvall:1989:RCT

[DV89] Paul Duvall and Theresa P. Vaughan. Recursions for Carlitz triples. Fibonacci Quarterly, 27(2):131–138, May 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-2/duvall.pdf.

Dekking:2023:RNN

[DvL23] Michel Dekking and Ad van Loon. On the representation of the natural numbers by powers of the golden mean. Fibonacci Quarterly, 61(2):105–??, May 2023. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/61-2/
dekking.pdf.

Darbro:1983:SGS

[DvT83] Wesley A. Darbro and Georg von Tiesenhausen. Sequences generated by self-replicating systems. Fibonacci Quarterly, 21(2): 97–106, May 1983. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/21-2/darbro.pdf.

Davis:1991:PTM

[DW91] Kenneth S. Davis and William A. Webb. Pascal’s triangle modulo 4. Fibonacci Quarterly, 29(1):79–83, February 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-1/davis.pdf.

Dresden:2022:WSF

[DX22] Greg Dresden and Yu Xiao. Weighted sums of Fibonacci and Lucas numbers through colorful tilings. Fibonacci Quarterly, 60 (2):126–??, May 2022. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/60-2/dresden.pdf.

Enneking:1976:GBN

[EA76] E. A. Enneking and J. C. Ahuja. Generalized Bell numbers. Fibonacci Quarterly, 14(1):67–73, February 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-1/enneking.pdf.

Ericksen:2012:PDB

[EA12] Larry Ericksen and Peter G. Anderson. Patterns in differences between rows in k-Zeckendorf arrays. Fibonacci Quarterly, 50(1): 11–??, February 2012. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/50-1/ericksen.pdf.

Edwards:2019:NCI

[EA19] Kenneth Edwards and Michael A. Allen. A new combinatorial interpretation of the Fibonacci numbers squared. Fibonacci Quarterly, 57(5):48–??, December 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/57-
5/edwards.pdf.

Edwards:2020:NCIb

[EA20a] Kenneth Edwards and Michael A. Allen. A new combinatorial interpretation of the Fibonacci numbers cubed. Fibonacci Quarterly, 58(5):128–??, December 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/58-
5/edwards.pdf.

Edwards:2020:NCIa

[EA20b] Kenneth Edwards and Michael A. Allen. A new combinatorial interpretation of the Fibonacci numbers squared. Part II. Fibonacci Quarterly, 58(2):169–??, May 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/58-
2/edwards.pdf.

El-Basil:1987:CPF

[EB87] Sherif El-Basil. On color polynomials of Fibonacci graphs. Journal of Computational Chemistry, 8(7):956–959, October 1987. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

El-Basil:1988:TCA

[EB88] S. El-Basil. Theory and computational applications of Fibonacci graphs. Journal of Mathematical Chemistry, 2(1):1–29, February 1988. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-

8897 (electronic). URL http://link.springer.com/article/
10.1007/BF01166466.

Ercan:2017:STG

[EB17] Sinan Ercan and Çigdem A. Bektas. Some topological and geometric properties of a new BK-space derived by using regular matrix of Fibonacci numbers. Linear Multilinear Algebra, 65(5): 909–921, 2017. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Englund:2000:MSW

[EBJ00] David A. Englund and Marjorie Bicknell-Johnson. Maximal subscripts within generalized Fibonacci sequences. Fibonacci Quarterly, 38(2):104–113, May 2000. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/38-2/englund.
pdf.

El-Basil:1989:FNT

[EBK89] Sherif El-Basil and Douglas J. Klein. Fibonacci numbers in the topological theory of benzenoid hydrocarbons and related graphs. Journal of Mathematical Chemistry, 3(1):1–23, January 1989. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01171882.

El-Basil:1985:FGP

[EBKT85] Sherif El-Basil, Pavel Křivka, and Nenad Trinajstić. Fibonacci graphs possessing identical matching polynomials. Journal of Mathematical Physics, 26(9):2396–2398, September 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

El-Desouky:1994:MNS

[ED94] B. S. El-Desouky. The multiparameter noncentral Stirling numbers. Fibonacci Quarterly, 32(3):218–225, June 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-3/eldesouky.pdf.

Eddy:1990:RNG

[Edd90] William F. Eddy. Random number generators for parallel processors. Journal of Computational and Applied Mathematics, 31(1): 63–71, July 24, 1990. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037704279090336X. See [MM65, Col87a]

for combined generators.

Edgar:1972:PS

[Edg72] Gerald Edgar. Perfect N-sequences for N, N + 1, and N + 2. Fibonacci Quarterly, 10(4):377–380, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-4/edgar-a.pdf.

Edgar:1986:LCM

[Edg86] Hugh M. Edgar. On the least common multiple of some binomial coefficients. Fibonacci Quarterly, 24(4):310–312, November 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-4/edgar.pdf.

Edgar:2016:ESF

[Edg16] Tom Edgar. Extending some Fibonacci–Lucas relations. Fibonacci Quarterly, 54(1):79–??, February 2016. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/54-1/edgar.pdf.

Edwards:2008:PLT

[Edw09] Kenneth Edwards. A Pascal-like triangle related to the Tribonacci numbers. Fibonacci Quarterly, 46/47(1):18–25, February 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/46_47-1/edwards.pdf.

Esmaeili:2010:FPB

[EE10] Mostafa Esmaeili and Morteza Esmaeili. A Fibonacci-polynomial based coding method with error detection and correction. Computers and Mathematics and Applications, 60(10):2738–2752, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122110006760.

Eggan:1982:PPP

[EES82] L. C. Eggan, Peter C. Eggan, and J. L. Selfridge. Polygonal products of polygonal numbers and the Pell equation. Fibonacci Quarterly, 20(1):24–27, February 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
1/eggan.pdf.

Epp:1980:NTG

[EF80] Robert J. Epp and Thomas S. Ferguson. A note on take-away games. Fibonacci Quarterly, 18(4):300–302, December 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-4/epp.pdf.

Elia:1998:EBJ

[EF98] Michele Elia and Piero Filipponi. Equations of the Bring–Jerrard form, the Golden Section, and the square Fibonacci numbers. Fibonacci Quarterly, 36(3):282–286, June 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-3/
elia.pdf.

Erdos:1972:SFN

[EG72] P. Erdős and R. L. Graham. On sums of Fibonacci numbers. Fibonacci Quarterly, 10(3):249–254, April 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-3/
erdos.pdf.

Erdos:1976:PFB

[EG76] P. Erdős and R. L. Graham. On the prime factors of binom(n,k). Fibonacci Quarterly, 14(4):348–352, November 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-4/erdos.pdf.

Everett:1989:ADH

[EG89] H. Everett and A. Gupta. Acyclic directed hypercubes may have exponential diameter. Information Processing Letters, 32(5):243– 245, September 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Edwards:2016:CIR

[EG16] Steven Edwards and William Griffiths. A combinatorial identity related to cross polytope numbers. Fibonacci Quarterly, 54(3): 253–??, August 2016. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/54-3/edwards.pdf.

Edwards:2017:GDC

[EG17] Steven Edwards and William Griffiths. Generalizations of Delannoy and cross polytope numbers. Fibonacci Quarterly, 55(4): 357–??, November 2017. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/55-4/edwards.pdf.

Egecioglu:1983:PCN

[Ege83] Ömer Egecioğlu. The parity of the Catalan number via lattice paths. Fibonacci Quarterly, 21(1):65–66, February 1983. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-1/egecioglu.pdf.

Etminaniesfahani:2022:AHA

[EGS22] Alireza Etminaniesfahani, Hanyu Gu, and Amir Salehipour. AB-FIA: a hybrid algorithm based on artificial bee colony and Fibonacci indicator algorithm. Journal of Computational Science, 61:??, May 2022. CODEN ???? ISSN 1877-7503 (print),

1877-7511 (electronic). URL https://www.sciencedirect.
com/science/article/pii/S1877750322000679.

Eichenauer-Herrmann:1995:PNG

[EH95] Jürgen Eichenauer-Herrmann. Pseudorandom number generation by nonlinear methods. International Statistical Review = Revue Internationale de Statistique, 63(2):247–255, August 1995. CO-DEN ISTRDP. ISSN 0306-7734 (print), 1751-5823 (electronic).

URL http://www.jstor.org/stable/1403620.

Ebeid:2007:ARI

[EH07] Nevine Maurice Ebeid and M. Anwar Hasan. On τ-adic representations of integers. Designs, Codes, and Cryptography, 45(3):271–296, December 2007. CODEN DCCREC. ISSN 0925-1022 (print), 1573-7586 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0925-
1022&volume=45&issue=3&spage=271.

Evans:2022:IDAa

[EH22] Emily Evans and Russell Jay Hendel. An infinite 2-dimensional array associated with electric circuits. Fibonacci Quarterly, 60 (5):151–??, December 2022. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/60-5/evans.pdf.

Ehrhart:1983:AHF

[Ehr83] E. Ehrhart. Associated hyperbolic and Fibonacci identities. Fibonacci Quarterly, 21(2):87–96, May 1983. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/21-
2/ehrhart.pdf.

Ehrhart:1984:EI

[Ehr84] E. Ehrhart. Euler’s integers. Fibonacci Quarterly, 22(3):218–228, August 1984. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/22-3/ehrhart.pdf.

Ehrhart:1988:PN

[Ehr88] Eugene Ehrhart. On prime numbers. Fibonacci Quarterly, 26 (3):271–274, August 1988. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/26-3/ehrhart.pdf.

Ehrlich:1989:PFS

[Ehr89] Amos Ehrlich. On the periods of the Fibonacci sequence modulo m. Fibonacci Quarterly, 27(1):11–13, February 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-1/ehrlich.pdf.

Ehrlich:1990:PDN

[Ehr90] Amos Ehrlich. Periods in Ducci’s n-number game of differences. Fibonacci Quarterly, 28(4):302–305, November 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/28-4/ehrlich.pdf.

Ehrlich:1994:CDD

[Ehr94] Amos Ehrlich. Cycles in doubling diagrams mod m. Fibonacci Quarterly, 32(1):74–78, February 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-
1/ehrlich.pdf.

Elia:2003:CFI

[EI03] Michele Elia and J. Carmelo Interlando. A class of Fibonacci ideal lattices in Z[ζ_{12}]. Fibonacci Quarterly, 41(3):279–289, June 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/41-3/elia.pdf.

Eisele:1951:BAT

[Eis51] Carolyn Eisele. The Liber Abaci through the eyes of Charles S. Peirce. Scripta Math., 17:236–259, 1951.

Eriksson:2017:LSB

[EJ17] Henrik Eriksson and Markus Jonsson. Level sizes of the Bulgarian solitaire game tree. Fibonacci Quarterly, 55(3):243–??, August 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/55-3/eriksson.pdf.

Erdos:1969:G

[EK69] P. Erdős and I. Kátai. On the growth of d_{k}(n). Fibonacci Quarterly, 7(3):267–274, October 1969. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/7-3/erdos.pdf.

Erdos:1988:LBC

[EKS88] P. Erdős, P. Kiss, and A. Sárközy. A lower bound for the counting function of Lucas pseudoprimes. Mathematics of Computation, 51(183):315–323, July 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

ElBachraoui:2008:RPP

[El 09] Mohamed El Bachraoui. Relatively prime partitions with two and three parts. Fibonacci Quarterly, 46/47(4):341–345, November 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/46_47-4/elbachraoui.pdf.

Elia:2001:DST

[Eli01] Michele Elia. Derived sequences, the Tribonacci recurrence and cubic forms. Fibonacci Quarterly, 39(2):107–115, May 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-2/elia.pdf.

Elizalde:2021:SBP

[Eli21] Sergi Elizalde. A simple bijective proof of a familiar derangement recurrence. Fibonacci Quarterly, 59(2):150–??, May 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/59-2/elizalde.pdf.

Ellia:2012:RRO

[Ell12] Ph. Ellia. A remark on the radical of odd perfect numbers. Fibonacci Quarterly, 50(3):231–??, August 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-3/ellia.pdf.

Elmore:1967:FF

[Elm67] Merritt Elmore. Fibonacci functions. Fibonacci Quarterly, 5(4): 371–382, December 1967. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/5-4/elmore.pdf.

Elsner:1996:AIN

[Els96] C. Elsner. On the approximation of irrational numbers with rationals restricted by congruence relations. Fibonacci Quarterly, 34(1):18–29, February 1996. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/34-1/elsner.pdf.

Elsner:1998:MRC

[Els98] C. Elsner. A metric result concerning the approximation of real numbers by continued fractions. Fibonacci Quarterly, 36(4):290– 294, August 1998. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/36-4/elsner.pdf.

Elsner:2000:DAR

[Els00] Carsten Elsner. On Diophantine approximations with rational restricted by arithmetical conditions. Fibonacci Quarterly, 38 (1):25–34, February 2000. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/38-1/elsner.pdf.

Elsner:2003:RAP

[Els03] Carsten Elsner. On rational approximations by Pythagorean numbers. Fibonacci Quarterly, 41(2):98–104, May 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-2/elsner.pdf.

Elsner:2005:RFS

[Els05] Carsten Elsner. On recurrence formulae for sums involving binomial coefficients. Fibonacci Quarterly, 43(1):31–45, February 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers/43-1/paper43-1-5.pdf.

Ely:1965:FF

[Ely65] Robert B. Ely III. Fibonaccious factors. Fibonacci Quarterly, 3 (3):187–198, October 1965. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/3-3/ely.pdf.

Eugeni:2001:NGF

[EM01] Franco Eugeni and Raffaele Mascella. A note on generalized Fibonacci numbers. Journal of Discrete Mathematical Sciences and Cryptography, 4(1):33–45, 2001. CODEN ???? ISSN 0972-0529.

Ellia:2015:RNT

[EM15] Philippe Ellia and Paolo Menegatti. Ramanujan–Nagell type equations and perfect numbers. Fibonacci Quarterly, 53(1):78– ??, February 2015. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/53-1/ellia.pdf.

Emerson:1966:ISE

[Eme66] Edgar I. Emerson. On the integer solution of the equation 5x^{2} + 6x + 1 = y^{2} and some related observations. Fibonacci Quarterly, 4(1):63–69, February 1966. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/4-1/emerson.pdf.

Emerson:1969:RSE

[Eme69] Edgar I. Emerson. Recurrent sequences in the equation DQ^{2} = R^{2} + N. Fibonacci Quarterly, 7(3):231–242, October 1969. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-3/emerson.pdf.

Esmaeili:2017:NCF

[EMG17] M. Esmaeili, M. Moosavi, and T. A. Gulliver. A new class of Fibonacci sequence based error correcting codes. Cryptography and Communications, 9(3):379–396, May 2017. CODEN ???? ISSN 1936-2447 (print), 1936-2455 (electronic). URL http:/

/link.springer.com/accesspage/article/10.1007/s12095-
015-0178-x.

ElWahbi:2002:SNR

[EMR02] Bouazza El Wahbi, Mehdi Mouline, and Mustapha Rachidi. Solving nonhomogeneous recurrence relations of order r by matrix methods. Fibonacci Quarterly, 40(2):106–117, May 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-2/el-wahbi.pdf.

El-Mikkawy:2010:NFF

[EMS10] Moawwad El-Mikkawy and Tomohiro Sogabe. A new family of k-Fibonacci numbers. Applied Mathematics and Computation, 215 (12):4456–4461, February 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Engstrom:1987:SGG

[Eng87] Philip G. Engstrom. Sections, golden and not so golden. Fibonacci Quarterly, 25(2):118–127, May 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-2/
engstrom.pdf.

Engel:1990:FNL

[Eng90] Konrad Engel. On the Fibonacci number of an M × N lattice. Fibonacci Quarterly, 28(1):72–78, February 1990. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/28-1/engel.pdf.

Englund:1991:EPR

[Eng91] David Englund. Entry point reciprocity of characteristic conjugate generalized Fibonacci sequences. Fibonacci Quarterly, 29 (3):197–198, August 1991. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/29-3/englund.pdf.

Englund:2001:ADS

[Eng01] David A. Englund. An algorithm for determining R(N) from the subscripts of the Zeckendorf representation of N. Fibonacci Quarterly, 39(3):250–252, June 2001. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/39-3/englund.
pdf.

Evans:2007:NOP

[EP07] Ronald Evans and Jonathan Pearlman. Nonexistence of odd perfect numbers of a certain form. Fibonacci Quarterly, 45(2):122– 127, May 2007. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/45-2/evans.pdf.

Epasinghe:1985:EAF

[Epa85] P. W. Epasinghe. Euclid’s algorithm and the Fibonacci numbers. Fibonacci Quarterly, 23(2):177–179, May 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/23-2/epasinghe.pdf.

Er:1983:CSO

[Er83a] M. C. Er. Computing sums of order-k Fibonacci numbers in log time. Information Processing Letters, 17(1):1–5, July 19, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Er:1983:FAC

[Er83b] M. C. Er. Fast algorithm for computing order-K Fibonacci numbers. The Computer Journal, 26(3):224–227, August 1983. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_26/Issue_03/tiff/224.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_26/Issue_
03/tiff/225.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_26/Issue_03/tiff/226.tif;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_26/Issue_03/tiff/227.
tif.

Er:1984:MFN

[Er84a] M. C. Er. The matrices of Fibonacci numbers. Fibonacci Quarterly, 22(2):134–139, May 1984. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/22-2/er.pdf.

Er:1984:SFN

[Er84b] M. C. Er. Sums of Fibonacci numbers by matrix methods. Fibonacci Quarterly, 22(3):204–207, August 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/22-3/er.pdf.

ElWahbi:2000:GFS

[ER00] Bouaza El Wahbi and Mustapha Rachidi. r-generalized Fibonacci sequences and the linear moment problem. Fibonacci Quarterly, 38(5):386–393, November 2000. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/38-5/el-wahbi.
pdf.

Ercolano:1972:MGC

[Erc72] Joseph L. Ercolano. Making golden cuts with a shoemaker’s knife. Fibonacci Quarterly, 10(4):439–440, October 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-4/ercolano-a.pdf.

Ercolano:1973:GTS

[Erc73] Joseph L. Ercolano. A geometric treatment of some of the algebraic properties of the golden section. Fibonacci Quarterly, 11(2): 204–208, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/11-2/ercolano.pdf.

Ercolano:1976:GSM

[Erc76] Joseph Ercolano. Golden sequences of matrices with application to Fibonacci algebra. Fibonacci Quarterly, 14(5):419–426, December 1976. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/14-5/ercolano.pdf.

Ercolano:1979:MGP

[Erc79] Joseph Ercolano. Matrix generators of Pell sequences. Fibonacci Quarterly, 17(1):71–76, February 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
1/ercolano.pdf.

Erdos:1974:LE

[Erd74] Paul Erdős. Letter to the editor. Fibonacci Quarterly, 12(4):335– ??, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/12-4/letter.pdf.

Erdos:1981:SEP

[Erd81] Paul Erdős. Some extremal problems on divisibility properties of sequences of integers. Fibonacci Quarterly, 19(3):208–213, August 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/19-3/erdos.pdf.

Ericksen:1998:PMT

[Eri98] Larry Ericksen. The Pascal–De Moivre triangles. Fibonacci Quarterly, 36(1):20–33, February 1998. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/36-1/ericksen.
pdf.

Euler:2000:EPSa

[ES00a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 38(4):372–376, August 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-4/elementary38-4.pdf.

Euler:2000:EPSb

[ES00b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 38(5):467–472, November 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-5/elementary38-5.pdf.

Euler:2001:EPSa

[ES01a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 39(1):85–90, February 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-1/elementary39-1.pdf.

Euler:2001:EPSb

[ES01b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 39(2):181–186, May 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-2/elementary39-2.pdf.

Euler:2001:EPSc

[ES01c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 39(4):373–377, August 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-4/elementary39-4.pdf.

Euler:2001:EPSd

[ES01d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 39(5):467–472, November 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-5/elementary39-5.pdf.

Euler:2002:EPSa

[ES02a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 40(1):85–89, February 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-1/elementary40-1.pdf.

Euler:2002:EPSb

[ES02b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 40(2):181–186, May 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-2/elementary40-2.pdf.

Euler:2002:EPSc

[ES02c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 40(4):372–378, August 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-4/elementary40-4.pdf.

Euler:2002:EPSd

[ES02d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 40(5):467–471, November 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-5/elementary40-5.pdf.

Euler:2003:EPSa

[ES03a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 41(1):85–90, February 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-1/elementary41-1.pdf.

Euler:2003:EPSb

[ES03b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 41(2):181–186, May 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-2/elementary41-2.pdf.

Euler:2003:EPSc

[ES03c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 41(4):374–379, August 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-4/elementary41-4.pdf.

Euler:2003:EPSd

[ES03d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 41(5):466–471, November 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-5/elementary41-5.pdf.

Euler:2004:EPSa

[ES04a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 42(1):86–91, February 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-1/February2004elementary.pdf.

Euler:2004:EPSb

[ES04b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 42(2):181–186, May 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-2/May2004elementary.pdf.

Euler:2004:EPSc

[ES04c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 42(3):277–282, August 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-3/August2004elementary.pdf.

Euler:2004:EPSd

[ES04d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 42(4):370–376, November 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-4/November2004elementary.pdf.

Euler:2005:EPSa

[ES05a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 43(1):85–90, February 2005. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/elementary43-1.pdf.

Euler:2005:EPSb

[ES05b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 43(2):181–186, May 2005. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/elementary43-2.pdf.

Euler:2005:EPSc

[ES05c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 43(3):277–282, August 2005. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/elementary43-3.pdf.

Euler:2005:EPSd

[ES05d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 43(4):371–376, November 2005. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/elementary43-4.pdf.

Euler:2006:EPSa

[ES06a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 44(1):85–90, February 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/February2006elementary.pdf.

Euler:2006:EPSb

[ES06b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 44(2):181–186, May 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/May2006elementary.pdf.

Euler:2006:EPSc

[ES06c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 44(3):277–282, August 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/August2006elementary.pdf.

Euler:2006:EPSd

[ES06d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 44(4):370–375, November 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/November2006elementary.pdf.

Euler:2007:EPSa

[ES07a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 45(1):85–90, February 2007. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/February2007elementary.pdf.

Euler:2007:EPSb

[ES07b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 45(2):181–186, May 2007. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/elementary45-2.pdf.

Euler:2007:EPSc

[ES07c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 45(3):277–282, August 2007. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/elementary45-3.pdf.

Euler:2007:EPSd

[ES07d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 45(4):368–375, November 2007. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/November2007elementary.pdf.

Euler:2008:EPSa

[ES09a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 46/47(1):85–90, February 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/Elem2008.pdf.

Euler:2008:EPSb

[ES09b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 46/47(2):181–185, May 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/Elem_Prob_May2009.pdf.

Euler:2008:EPSc

[ES09c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 46/47(3):279–282, August 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/Elem_Prob_August2009.pdf.

Euler:2008:EPSd

[ES09d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 46/47(4):369–373, November 2008/ 2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Problems/Elem_Prob_NOV2009.pdf.

Euler:2010:EPSa

[ES10a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 48(1):87–88, February 2010. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/Elem_Prob_Feb_2010.pdf.

Euler:2010:EPSb

[ES10b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 48(2):182–183, May 2010. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbMay2010.pdf.

Euler:2010:EPSc

[ES10c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 48(3):277–282, August 2010. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbAug2010.pdf.

Euler:2010:EPSd

[ES10d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 48(4):366–372, November 2010. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbNov2010.pdf.

Eustis:2010:CPS

[ES10e] Alex Eustis and Mark Shattuck. Combinatorial proofs of some formulas for L_{m}^{r}. Fibonacci Quarterly, 48(1):62–67, February 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-1/eustis.pdf.

Euler:2011:EPSa

[ES11a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 49(1):82–88, February 2011. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbFeb2011.pdf.

Euler:2011:EPSb

[ES11b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 49(2):180–186, May 2011. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbMay2011.pdf.

Euler:2011:EPSc

[ES11c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 49(3):273–280, August 2011. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbAug2011.pdf.

Euler:2011:EPSd

[ES11d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 49(4):367–??, November 2011. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbNOV2011.pdf.

Euler:2012:EPSa

[ES12a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 50(1):82–??, February 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbFeb2012.pdf.

Euler:2012:EPSb

[ES12b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 50(2):180–??, May 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Problems/ElemProbMay2012.pdf.

Euler:2012:EPSc

[ES12c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 50(3):272–??, August 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbAugust2012.pdf.

Euler:2012:EPS

[ES12d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 50(4):366–??, November 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbNov2012.pdf.

Euler:2012:CRB

[ES12e] Russell Euler and Jawad Sadek. Congruence relations from Binet forms. Fibonacci Quarterly, 50(3):246–??, August 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-3/euler.pdf.

Euler:2013:EPSa

[ES13a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 51(1):84–??, February 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbFeb2013.pdf.

Euler:2013:EPSb

[ES13b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 51(2):177–??, May 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Problems/ElemProbMay2013.pdf.

Euler:2013:EPSc

[ES13c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 51(3):274–??, August 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbAugust2013.pdf.

Euler:2013:EPS

[ES13d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 51(4):367–??, November 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbSolnNov13.pdf.

Euler:2014:EPSa

[ES14a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 52(1):80–??, February 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbSolnFeb14.pdf.

Euler:2014:EPSb

[ES14b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 52(2):178–??, May 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Problems/ElemProbSolnMay14.pdf.

Euler:2014:EPSc

[ES14c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 52(3):274–??, August 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbSolnAug14.pdf.

Euler:2014:EPS

[ES14d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 52(4):367–??, November 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbSolnNov14.pdf.

Euler:2014:GPE

[ES14e] Russell Euler and Jawad Sadek. On a generalized Pell equation and a characterization of the Fibonacci and Lucas numbers. Fibonacci Quarterly, 52(3):243–??, August 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-3/euler.pdf.

Epple:2015:CFS

[ES15a] Dennis D. A. Epple and Jason Siefken. Collapse: A Fibonacci and Sturmian game. American Mathematical Monthly, 122(6):515– 527, June 2015. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/
10.4169/amer.math.monthly.122.6.515.

Euler:2015:EPSa

[ES15b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 53(1):81–??, February 2015. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbFeb2015.pdf.

Euler:2015:EPSb

[ES15c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 53(2):180–??, May 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Problems/ElemProbSolnMay15.pdf.

Euler:2015:EPSc

[ES15d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 53(3):272–??, August 2015. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbSolnAugust2015.pdf.

Euler:2015:EPSd

[ES15e] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 53(4):365–??, November 2015. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbSolnNov2015.pdf.

Euler:2015:EPE

[ES15f] Russell Euler and Jawad Sadek. An extension of the periodicity of an extended Fibonacci family. Fibonacci Quarterly, 53(4):335– ??, November 2015. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/53-4/euler.pdf.

Euler:2016:EPSa

[ES16a] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 54(1):80–??, February 2016. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbSolnFeb2016.pdf.

Euler:2016:EPSb

[ES16b] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 54(2):178–??, May 2016. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Problems/ElemProbSolnMay16No53.pdf.

Euler:2016:EPSc

[ES16c] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 54(3):271–??, August 2016. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbSolnAug16No3.pdf.

Euler:2016:EPSd

[ES16d] Russ Euler and Jawad Sadek. Elementary problems and solutions. Fibonacci Quarterly, 54(4):366–??, November 2016. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/ElemProbSolnNov2016.pdf.

Euler:2016:DPD

[ES16e] Russell Euler and Jawad Sadek. A direct proof that F_{n} divides F_{mn} extended to divisibility properties of related numbers. Fibonacci Quarterly, 54(2):160–??, May 2016. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/54-
2/euler.pdf.

Euler:2017:RBK

[ES17] Russell Euler and Jawad Sadek. Relationships between k-gonal numbers that are centered k-gonal, and Lucas and related numbers. Fibonacci Quarterly, 55(4):315–??, November 2017. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-4/euler.pdf.

Essebbar:2003:DIF

[Ess03] Belkheir Essebbar. Double indexed Fibonacci sequences and the bivariate probability distribution. Fibonacci Quarterly, 41(4): 290–300, August 2003. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/41-4/essebbar.pdf.

Egecioglu:2020:KFC

[ESS20] Ömer Eğecioğlu, Elif Saygı, and Zülfükar Saygı. k-Fibonacci cubes: A family of subgraphs of Fibonacci cubes. International Journal of Foundations of Computer Science (IJFCS), 31(05):

639–661, August 2020. ISSN 0129-0541. URL https://www.

worldscientific.com/doi/10.1142/S0129054120500318.

Egecioglu:2021:NSC

[ESS21] Ömer Eğecioğlu, Elif Saygı, and Zülfükar Saygı. The number of short cycles in Fibonacci cubes. Theoretical Computer Science, 871(??):134–146, June 6, 2021. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397521002395.

Eswarathasan:1978:SPF

[Esw78a] A. Eswarathasan. On square pseudo-Fibonacci numbers. Fibonacci Quarterly, 16(4):310–314, August 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/16-4/eswarathasan.pdf.

Eswarathasan:1978:SPL

[Esw78b] A. Eswarathasan. On square pseudo-Lucas numbers. Canadian Mathematical Bulletin = Bulletin canadien de mathématiques, 21 (??):297–304, ???? 1978. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eswarathasan:1979:PFN

[Esw79a] A. Eswarathasan. On pseudo-Fibonacci numbers of the form 2S^{2}, where S is an integer. Fibonacci Quarterly, 17(2):142–146, April 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/17-2/eswarathasan.pdf.

Eswarathasan:1979:PLN

[Esw79b] A. Eswarathasan. On pseudo-Lucas numbers of the form 2S^{2}. Canadian Mathematical Bulletin = Bulletin canadien de mathématiques, 22(??):29–34, ???? 1979. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Eloff:2007:NFI

[ESW07] Daniel Eloff, Blair K. Spearman, and Kenneth S. Williams. A number field with infinitely many normal integral bases. Fibonacci Quarterly, 45(2):151–154, May 2007. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/45-
2/eloff.pdf.

Estes:2013:ISC

[ESW13] John Estes, William Staton, and Bing Wei. Independent sets of cardinality s of maximal outerplanar graphs. Fibonacci Quarterly, 51(2):147–??, May 2013. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/51-2/estes.
pdf.

Embree:1999:GDR

[ET99] Mark Embree and Lloyd N. Trefethen. Growth and decay of random Fibonacci sequences. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 455 (1987):2471–2485, July 8, 1999. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Erlebach:1983:EFS

[EV83] Lee Erlebach and William Yslas Vélez. Equiprobability in the Fibonacci sequence. Fibonacci Quarterly, 21(3):189–191, August 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/21-3/erlebach.pdf.

Eckert:1989:GIT

[EV89] Ernest J. Eckert and Preben Dahl Vestergaard. Groups of integral triangles. Fibonacci Quarterly, 27(5):458–464, November 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-5/eckert.pdf.

Eterevsky:2004:NPD

[EV04] Oleg Eterevsky and Maxim Vsemirnov. On the number of prime divisors of higher-order Carmichael numbers. Fibonacci Quarterly, 42(2):141–148, May 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-2/
quarteterevsky02_2004.pdf.

Evans:1977:DEI

[Eva77] Ronald Evans. Diophantine equations involving the greatest integer function. Fibonacci Quarterly, 15(2):170–172, April 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-2/evans.pdf.

Eves:1968:BRB

[Eve68] Howard Eves. Book review: Mathematical Quickies. Fibonacci Quarterly, 6(1):88–??, February 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-1/

review2.pdf. See [Tri67e].

Everett:1975:GIT

[Eve75] C. J. Everett. A greatest integer theorem for Fibonacci spaces. Fibonacci Quarterly, 13(3):260–262, October 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-3/everett.pdf.

Eves:1981:HTB

[Eve81] Howard Eves. Hail to thee, blithe spirit! Fibonacci Quarterly, 19 (3):193–195, August 1981. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/19-3/eves.pdf.

Ewell:1980:RTR

[Ewe80] John A. Ewell. Recurrences for two restricted partition function. Fibonacci Quarterly, 18(1):1–2, February 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/18-1/ewell.pdf.

Ewell:1982:CWQ

[Ewe82] John A. Ewell. Consequences of Watson’s quintuple-product identity. Fibonacci Quarterly, 20(3):256–262, August 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-3/ewell.pdf.

Ewell:1986:EST

[Ewe86a] John A. Ewell. On the enumerator for sums of three squares. Fibonacci Quarterly, 24(2):150–153, May 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-2/
ewell.pdf.

Ewell:1986:TFP

[Ewe86b] John A. Ewell. On two- and four-part partitions of numbers each part a square. Fibonacci Quarterly, 24(1):67–69, February 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-1/ewell.pdf.

Ewell:1988:STT

[Ewe88] John A. Ewell. On sums of three triangular numbers. Fibonacci Quarterly, 26(4):332–335, November 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
4/ewell.pdf.

Ewell:1992:RNS

[Ewe92] John A. Ewell. On representations of numbers by sums of two triangular numbers. Fibonacci Quarterly, 30(2):175–178, May 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/30-2/ewell.pdf.

Ewell:2001:ADE

[Ewe01] John A. Ewell. Algorithmic determination of the enumerator for sums of three triangular numbers. Fibonacci Quarterly, 39(3): 276–278, June 2001. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/39-3/ewell.pdf.

Ewell:2003:EPJ

[Ewe03] John A. Ewell. An elementary proof of Jacobi’s four-square theorem. Fibonacci Quarterly, 41(3):224–228, June 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-3/ewell.pdf.

Ewell:2007:AED

[Ewe07a] John A. Ewell. Additive evaluation of the divisor function. Fibonacci Quarterly, 45(1):22–25, February 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/45-1/ewell.pdf.

Ewell:2007:SDF

[Ewe07b] John A. Ewell. On the sum-of-divisors function. Fibonacci Quarterly, 45(3):205–207, August 2007. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/45-3/
ewell.pdf.

Edson:2004:RPI

[EZ04] Marcia Edson and Luca Q. Zamboni. On representations of positive integers in the Fibonacci base. Theoretical Computer Science, 326(1–3):241–260, October 20, 2004. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Fielder:1989:CHE

[FA89] Daniel C. Fielder and Cecil O. Alford. On a conjecture by Hoggatt with extensions to Hoggatt sums and Hoggatt triangles. Fibonacci Quarterly, 27(2):160–168, May 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-2/
fielder.pdf.

Fahssi:2016:SII

[Fah16] Nour-Eddine Fahssi. Some identities involving polynomial coefficients. Fibonacci Quarterly, 54(2):125–??, May 2016. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/54-2/fahssi.pdf.

Fang:2008:ASN

[Fan09] Jin-Hui Fang. On almost superperfect numbers. Fibonacci Quarterly, 46/47(2):111–114, May 2008/2009. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/46_
47-2/fang.pdf.

Farrell:1986:OFS

[Far86] E. J. Farrell. On the occurrences of Fibonacci sequences in the counting of matchings in linear polygonal chains. Fibonacci Quarterly, 24(3):238–246, August 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-
3/farrell.pdf.

Faulconbridge:1964:FSE

[Fau64] Albert J. Faulconbridge. Fibonacci summation economics Part I. Fibonacci Quarterly, 2(4):320–322, December 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/2-4/faulconbridge.pdf.

Faulconbridge:1965:FSE

[Fau65] Albert J. Faulconbridge. Fibonacci summation economics Part II. Fibonacci Quarterly, 3(4):309–314, December 1965. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/3-4/faulconbridge.pdf.

Filipponi:1994:ICF

[FB94] Piero Filipponi and Marco Bucci. On the integrity of certain Fibonacci sums. Fibonacci Quarterly, 32(3):245–252, June 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-3/filipponi.pdf.

Filipponi:1999:IF

[FB99] P. Filipponi and O. Brugia. On the integers of the form n(n − 1) − 1. Fibonacci Quarterly, 37(3):262–263, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-3/filipponi.pdf.

Fielder:2001:FTS

[FBJ01] D. C. Fielder and M. Bicknell-Johnson. The first 330 terms of sequence A013583. Fibonacci Quarterly, 39(1):75–84, February 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/39-1/fielder.pdf.

Chuan:1993:SFW

[fC93] Wai fong Chuan. Symmetric Fibonacci words. Fibonacci Quarterly, 31(3):251–254, August 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-
3/chuan.pdf.

Chuan:1995:EPG

[fC95a] Wai fong Chuan. Extraction property of the golden sequence. Fibonacci Quarterly, 33(2):113–122, May 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-2/
chuan2.pdf.

Chuan:1995:GFW

[fC95b] Wai fong Chuan. Generating Fibonacci words. Fibonacci Quarterly, 33(2):104–112, May 1995. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/33-2/chuan1.
pdf.

Chuan:2003:CWM

[fC03] Wai fong Chuan. Characterizations of α-words, moments, and determinants. Fibonacci Quarterly, 41(3):194–208, June 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-3/chuan.pdf.

Filipponi:1999:FFN

[FE99] Piero Filipponi and Michele Elia. The factorization of x^{5} ± p^{2}x − k and Fibonacci numbers. Fibonacci Quarterly, 37(4):290–298, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/37-4/filipponi1.pdf.

Federico:1964:MNF

[Fed64] P. J. Federico. Mathematical notes: a Fibonacci perfect squared square. American Mathematical Monthly, 71(4):404–406, April 1964. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Feeman:1967:RFL

[Fee67] G. F. Feeman. On ratios of Fibonacci and Lucas numbers. Fibonacci Quarterly, 5(1):99–106, February 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/5-1/feeman.pdf.

Feinberg:1963:FT

[Fei63] Mark Feinberg. Fibonacci–Tribonacci. Fibonacci Quarterly, 1(3): 71–74, October 1963. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/1-3/feinberg.pdf.

Feinberg:1964:NS

[Fei64] M. Feinberg. New slants. Fibonacci Quarterly, 2(3):223–227, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/2-3/feinberg.pdf.

Feinberg:1967:LT

[Fei67] Mark Feinberg. A Lucas triangle. Fibonacci Quarterly, 5(5):486– 490, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/5-5/feinberg.pdf.

Feinberg:1972:PPP

[Fei72] Andrew Feinberg. Polyhedra, pentagrams and Plato. Fibonacci Quarterly, 10(4):435–438, October 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
4/feinberg.pdf.

Fenwick:2003:ZIA

[Fen03] Peter Fenwick. Zeckendorf integer arithmetic. Fibonacci Quarterly, 41(5):405–413, November 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-
5/fenwick.pdf.

Feng:2011:FID

[Fen11a] Jishe Feng. Fibonacci identities via the determinant of tridiagonal matrix. Applied Mathematics and Computation, 217(12):5978– 5981, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Feng:2011:SNR

[Fen11b] Jishe Feng. Some new remarks about the dying rabbit problem. Fibonacci Quarterly, 49(2):171–176, May 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/49-2/feng.pdf.

Ferguson:1960:FS

[Fer60] David E. Ferguson. Fibonaccian searching. Communications of the Association for Computing Machinery, 3(12):648, December 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ferns:1965:RIS

[Fer65] H. H. Ferns. On the representation of integers as sums of distinct Fibonacci numbers. Fibonacci Quarterly, 3(1):21–30, February 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/3-1/ferns.pdf.

Ferguson:1966:EGF

[Fer66] David E. Ferguson. An expression for generalized Fibonacci numbers. Fibonacci Quarterly, 4(3):270–272, October 1966. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-3/ferguson.pdf.

Ferns:1968:PFN

[Fer68] H. H. Ferns. Pseudo-Fibonacci numbers. Fibonacci Quarterly, 6(6):305–317, December 1968. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/6-6/ferns.pdf.

Ferns:1969:PFL

[Fer69] H. H. Ferns. Products of Fibonacci and Lucas numbers. Fibonacci Quarterly, 7(1):1–12, February 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-1/
ferns.pdf.

Ferguson:1973:BNN

[Fer73] Helaman Rolfe Pratt Ferguson. Bernoulli numbers and nonstandard differentiable structures on (4k − 1)-spheres. Fibonacci Quarterly, 11(1):1–14, February 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-1/
ferguson.pdf.

Ferguson:1976:GFN

[Fer76] Helaman Rolfe Pratt Ferguson. On a generalization of the Fibonacci numbers useful in memory allocation schema; or all about the zeroes of Z^{k} − Z^{k−1} − 1, k > 0. Fibonacci Quarterly, 14(3): 233–243, October 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-3/ferguson.pdf.

Ferguson:1978:FPC

[Fer78] Helaman Rolfe Pratt Ferguson. The Fibonacci pseudogroup, characteristic polynomials and eigenvalues of tridiagonal matrices, periodic linear recurrence systems and application to quantum mechanics. Fibonacci Quarterly, 16(5):435–446, October 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/16-5/ferguson.pdf.

Ferri:1997:AFL

[Fer97] Giuseppe Ferri. The appearance of Fibonacci and Lucas numbers in the simulation of electrical power lines supplied by two sides. Fibonacci Quarterly, 35(2):149–155, May 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-2/ferri.pdf.

Ferraro:2001:FNP

[Fer01] Peter J. Ferraro. Fibonacci numbers and powers: 10765. American Mathematical Monthly, 108(10):978–979, December 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/2695429.

Freitag:1989:RIS

[FF89] Herta T. Freitag and Piero Filipponi. On the F -representation of integral sequences {F_{n}^{2}/d} and {L_{n}^{2}/d} where d is either a Fibonacci or a Lucas number. Fibonacci Quarterly, 27(3):276– 282, June 1989. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/27-3/freitag-a.pdf.

Filipponi:1994:FAS

[FF94] Piero Filipponi and Herta T. Freitag. Fibonacci autocorrelation sequences. Fibonacci Quarterly, 32(4):356–368, August 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-4/filipponi.pdf.

Filipponi:1999:S

[FF99a] Piero Filipponi and Giuseppe Fierro. On the sequences T_{n} = T_{n−1}+T_{n−2}+hn+k. Fibonacci Quarterly, 37(4):326–332, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/37-4/filipponi2.pdf.

Freitag:1999:DFN

[FF99b] Herta T. Freitag and Piero Filipponi. Division of Fibonacci numbers by k. Fibonacci Quarterly, 37(2):128–134, May 1999. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-2/freitag.pdf.

Ferri:1991:NNT

[FFD91] G. Ferri, M. Faccio, and A. D’Amico. A new numerical triangle showing links with Fibonacci numbers. Fibonacci Quarterly, 29 (4):316–321, November 1991. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/29-4/ferri.
pdf.

Ferri:1992:FNL

[FFD92] G. Ferri, M. Faccio, and A. D’Amico. Fibonacci numbers and ladder network impedance. Fibonacci Quarterly, 30(1):62–67, February 1992. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/30-1/ferri.pdf.

Fairgrieve:2005:PDF

[FG05] Steve Fairgrieve and Henry W. Gould. Product difference Fibonacci identities of Simson, Gelin–Cesàro, Tagiuri and generalizations. Fibonacci Quarterly, 43(2):137–141, May 2005. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/43-2/paper43-2-8.pdf.

Faber:2023:IWS

[FG23] Xander Faber and Jon Grantham. On integers whose sum is the reverse of their product. Fibonacci Quarterly, 61(1):28–??, February 2023. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/61-1/faber.pdf.

Fibonacci:2020:LA

[FGD20] Leonardo Fibonacci, Enrico Giusti, and Paolo D’Alessandro. Liber Abbaci. Leo S. Olschki, Firenze, Italia, 2020. ISBN 88-222-6658-7. cxvii + 822 + 22 pp. LCCN QA32 .F47 2020.

Filipponi:1988:MAC

[FH88] Piero Filipponi and A. F. Horadam. A matrix approach to certain identities. Fibonacci Quarterly, 26(2):115–126, May 1988. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/26-2/filipponi.pdf.

Filipponi:1993:SDS

[FH93] Piero Filipponi and Alwyn F. Horadam. Second derivative sequences of Fibonacci and Lucas polynomials. Fibonacci Quarterly, 31(3):194–204, August 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-
3/filipponi.pdf.

Filipponi:1994:ASD

[FH94] Piero Filipponi and Alwyn F. Horadam. Addendum to “Second Derivative Sequences of Fibonacci and Lucas Polynomials”. Fibonacci Quarterly, 32(2):110–??, May 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-2/
filipponi.pdf.

Filipponi:1998:ZDC

[FH98] Piero Filipponi and Evelyn L. Hart. The Zeckendorf decomposition of certain Fibonacci–Lucas products. Fibonacci Quarterly, 36(3):240–247, June 1998. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/36-3/filipponi.pdf.

Fibonacci:2008:FPG

[FH08] Leonardo Fibonacci and Barnabas Hughes, editors. Fibonacci’s De practica geometrie. Sources and studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-72930-5. xxxv + 408 pp. LCCN QA32 .F4613 2008. URL http://www.loc.gov/

catdir/enhancements/fy0826/2007934985-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0826/2007934985-t.
html;   https://www.math.utah.edu/pub/tex/bib/fibquart.
bib.

Florez:2020:APP

[FH20] Rigoberto Flørez and Robinson A. Higuita. Another proof for partial strong divisibility property of Lucas-type polynomials. Fibonacci Quarterly, 58(1):70–??, February 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/58-1/florez.pdf.

Filipin:2010:T

[FHT10] Alan Filipin, Bo He, and Alain Togbé. On the D(4)-triple F_{2k}, F_{2k+6}, F_{2k+4}. Fibonacci Quarterly, 48(3):219–227, August 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-3/filipin.pdf.

Fielder:1964:PEM

[Fie64] Daniel Fielder. Partition enumeration by means of simpler partitions. Fibonacci Quarterly, 2(2):115–117, April 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-2/fielder.pdf.

Fielder:1966:EPS

[Fie66] Daniel C. Fielder. Enumeration of partitions subject to limitations on size of members. Fibonacci Quarterly, 4(3):209–215, October 1966. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/4-3/fielder.pdf.

Fielder:1967:CLL

[Fie67a] Daniel C. Fielder. Certain Lucas-like sequences and their generation by partitions of numbers. Fibonacci Quarterly, 5(4):319–324, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/5-4/fielder1.pdf.

Fielder:1967:RTR

[Fie67b] Daniel C. Fielder. Remarks on two related sequences of numbers. Fibonacci Quarterly, 5(4):325–327, December 1967. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-4/fielder2.pdf.

Fielder:1968:GSN

[Fie68a] Daniel C. Fielder. Generation of Stirling numbers by means of special partitions of numbers. Fibonacci Quarterly, 6(5):1– 9, November 1968. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/6-5/fielder.pdf.

Fielder:1968:SIS

[Fie68b] Daniel C. Fielder. Special integer sequences controlled by three parameters. Fibonacci Quarterly, 6(3):64–69, June 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-3/fielder.pdf.

Fielder:1973:CCP

[Fie73a] Daniel C. Fielder. Counting of certain partitions of numbers. Fibonacci Quarterly, 11(4):441–442, November 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-4/fielder2.pdf.

Fielder:1973:DSS

[Fie73b] Daniel C. Fielder. A discussion of subscript sets with some Fibonacci counting help. Fibonacci Quarterly, 11(4):420–428, November 1973. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/11-4/fielder1.pdf.

Fielder:1974:FNT

[Fie74] Daniel C. Fielder. Fibonacci numbers in tree counts for sector and related graphs. Fibonacci Quarterly, 12(4):355–359, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/12-4/fielder.pdf.

Fifield:1970:UD

[Fif70] Dorothy Fifield. Ups and downs. Fibonacci Quarterly, 8(3):335– ??, April 1970. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/8-3/fifield.pdf.

Filaseta:1986:NMS

[Fil86a] Michael Filaseta. Newton’s method and simple continued fractions. Fibonacci Quarterly, 24(1):41–45, February 1986. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-1/filaseta.pdf.

Filipponi:1986:NRI

[Fil86b] Piero Filipponi. A note on the representation of integers as a sum of distinct Fibonacci numbers. Fibonacci Quarterly, 24(4): 336–343, November 1986. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/24-4/filipponi.pdf.

Filipponi:1991:NCL

[Fil91] Piero Filipponi. A note on a class of Lucas sequences. Fibonacci Quarterly, 29(3):256–263, August 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-
3/filipponi.pdf.

Filipponi:1992:WFB

[Fil92] Piero Filipponi. Waring’s formula, the binomial formula, and generalized Fibonacci matrices. Fibonacci Quarterly, 30(3):225– 231, August 1992. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/30-3/filipponi.pdf.

Filipponi:1993:RFL

[Fil93] Piero Filipponi. Real Fibonacci and Lucas numbers with real subscripts. Fibonacci Quarterly, 31(4):307–314, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-4/filipponi.pdf.

Filipponi:1995:SBF

[Fil95] Piero Filipponi. Some binomial Fibonacci identities. Fibonacci Quarterly, 33(3):251–257, June 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-3/
filipponi.pdf.

Filipponi:1996:FNW

[Fil96] Piero Filipponi. On the Fibonacci numbers whose subscript is a power. Fibonacci Quarterly, 34(3):271–276, June 1996. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-3/filipponi.pdf.

Filipponi:1997:FFM

[Fil97a] Piero Filipponi. A family of 4-by-4 Fibonacci matrices. Fibonacci Quarterly, 35(4):300–307, November 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-
4/filipponi.pdf.

Filipponi:1997:MDP

[Fil97b] Piero Filipponi. Modified Dickson polynomials. Fibonacci Quarterly, 35(1):11–18, February 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-
1/filipponi1.pdf.

Filipponi:1997:OSF

[Fil97c] Piero Filipponi. An observation of summation formulas for generalized sequences. Fibonacci Quarterly, 35(1):57–61, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/35-1/filipponi2.pdf.

Filipponi:1997:SFS

[Fil97d] Piero Filipponi. Summation formulas for special Lehmer numbers. Fibonacci Quarterly, 35(3):252–257, August 1997. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-3/filipponi.pdf.

Filipponi:1998:CEL

[Fil98a] Piero Filipponi. Combinatorial expressions for Lucas numbers. Fibonacci Quarterly, 36(1):63–64, February 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-1/filipponi1.pdf.

Filipponi:1998:NTT

[Fil98b] Piero Filipponi. A note on two theorems of Melham and Shannon. Fibonacci Quarterly, 36(1):66–67, February 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-1/filipponi2.pdf.

Filipponi:1998:RGL

[Fil98c] Piero Filipponi. Representing generalized Lucas numbers in terms of their α-values. Fibonacci Quarterly, 36(5):457–462, November 1998. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/36-5/filipponi.pdf.

Filipponi:2000:ECI

[Fil00] Piero Filipponi. Evaluation of certain infinite series involving terms of generalized sequences. Fibonacci Quarterly, 38(4):310– 315, August 2000. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/38-4/filipponi.pdf.

Filipin:2015:EDP

[Fil15] Alan Filipin. The extendibility of D(4)-pair {F_{2k}, 5F_{2k}}. Fibonacci Quarterly, 53(2):124–??, May 2015. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/53-
2/filipin.pdf.

Finkelstein:1975:LNW

[Fin75] Raphael Finkelstein. On Lucas numbers which are one more than a square. Fibonacci Quarterly, 13(4):340–342, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/13-4/finkelstein.pdf.

Finch:1991:CAA

[Fin91] Steven R. Finch. Conjectures about s-additive sequences. Fibonacci Quarterly, 29(3):209–214, August 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/29-3/finch.pdf.

Fischer:1976:FSE

[Fis76] Kurt Fischer. The Fibonacci sequence encountered in nerve physiology. Fibonacci Quarterly, 14(4):377–379, November 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-4/fischer.pdf.

Fischler:1981:HFG

[Fis81] Roger Fischler. How to find the “golden number” without really trying. Fibonacci Quarterly, 19(5):406–409, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-5/fischler.pdf.

Fisher:1996:FGM

[Fis96] David W. Fisher. Fibonacci and the golden mean. In Blau et al. [B^{+}96], page 183. ISBN 0-89791-784-7. ISSN 1069-5419.

LCCN T385 .S54 1996b. URL http://www.acm.org:80/pubs/
citations/proceedings/graph/253607/p183-fisher/.

Florez:2012:GPH

[FJ12] Rigoberto Flórez and Leandro Junes. GCD properties in Hosoya’s triangle. Fibonacci Quarterly, 50(2):163–??, May 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-2/florez.pdf.

Fang:2020:CLT

[FJL^{+}20] Evan Fang, Jonathan Jenkins, Zack Lee, Daniel Li, Ethan Lu, Steven J. Miller, Dilhan Salgado, and Joshua M. Siktar. Central limit theorems for compound paths on the two-dimensional lattice. Fibonacci Quarterly, 58(3):208–??, August 2020. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/58-3/fang.pdf.

Fisher:1972:GFS

[FK72] P. S. Fisher and E. E. Kohlbecker. A generalized Fibonacci sequence. Fibonacci Quarterly, 10(4):337–344, October 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-4/fisher-a.pdf.

Finkelstein:1970:ARS

[FKL70] Raphael Finkelstein, Edgar Karst, and Hymie London. Application of recursive sequences to Diophantine equations. Fibonacci Quarterly, 8(5):463–469, December 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-5/
finkelstein.pdf.

Fox:2014:SPF

[FKM^{+}14] Kyle Fox, William B. Kinnersley, Daniel McDonald, Nathan Orlow, and Gregory J. Puleo. Spanning paths in Fibonacci-sum graphs. Fibonacci Quarterly, 52(1):46–??, February 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-1/fox.pdf.

Forget:1968:PTF

[FL68] T. W. Forget and T. A. Larkin. Pythagorean triads of the form X, X + 1, Z described by recurrence sequences. Fibonacci Quarterly, 6(3):94–104, June 1968. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/6-3/forget.
pdf.

Faye:2018:XCP

[FL18] Bernadette Faye and Florian Luca. On X-coordinates of Pell equations that are repdigits. Fibonacci Quarterly, 56(1):52–??, February 2018. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/56-1/faye.pdf.

Frontczak:2022:APS

[FL22] Robert Frontczak and Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 60(4):373–??, November 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Problems/AdvProbNov2022.pdf.

Flanigan:1978:GTP

[Fla78] Jim Flanigan. Generalized two-pile Fibonacci Nim. Fibonacci Quarterly, 16(5):459–469, October 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
5/flanigan.pdf.

Flanigan:1982:OPT

[Fla82] Jim Flanigan. One-pile time and size dependent take-away games. Fibonacci Quarterly, 20(1):51–58, February 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-1/flanigan.pdf.

Flores:1967:DCG

[Flo67] Ivan Flores. Direct calculation of k-generalized Fibonacci numbers. Fibonacci Quarterly, 5(3):259–266, October 1967. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-3/flores.pdf.

Florek:2018:CGT

[Flo18] Wojciech Florek. A class of generalized Tribonacci sequences applied to counting problems. Applied Mathematics and Computation, 338(??):789–808, December 1, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300318305022.

Foundas:2008:IAC

[FLP08] E. Foundas, Ch. Lytras, and C. Patsakis. Improved algorithms for the calculation of Fibonacci numbers. Journal of Discrete Mathematical Sciences and Cryptography, 11 (1):113–120, February 2008. CODEN ???? ISSN 0972-

0529. URL http://www.connectjournals.com/achivestoc.
php?bookmark=CJ-003072&volume=11&issue_id=01.

Filipponi:1995:SPA

[FM95] Piero Filipponi and Renato Menicocci. Some probabilistic aspects of the terminal digits of Fibonacci numbers. Fibonacci Quarterly, 33(4):325–331, August 1995. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/33-4/filipponi.pdf.

Fallahpour:2015:AWB

[FM15] M. Fallahpour and D. Megias. Audio watermarking based on Fibonacci numbers. IEEE/ACM Transactions on Audio, Speech, and Language Processing, 23(8):1273–1282, August 2015. CO-DEN ???? ISSN 2329-9290.

Filipponi:1994:EDP

[FMH94] Piero Filipponi, Renato Menicocci, and Alwyn F. Horadam. Extended Dickson polynomials. Fibonacci Quarterly, 32(5):455–464, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/32-5/filipponi.pdf.

Folkerts:2004:LFK

[Fol04] Menso Folkerts. Leonardo Fibonacci’s knowledge of Euclid’s Elements and of other mathematical texts. Bollettino di Storia delle Scienze Matematiche, 24(1):93–113 (2005), 2004. ISSN 0392-4432 (print), 1724-1650 (electronic). Leonardo Fibonacci. Matematica e societa nel Mediterraneo nel secolo XIII. Vol. II.

Ford:1967:RRS

[For67a] Gary G. Ford. Recurrence relations for sequences like F_{F_{n}}. Fibonacci Quarterly, 5(2):129–136, April 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-2/
ford.pdf.

Ford:1967:SFR

[For67b] Gary G. Ford. A shift formula for recurrence relations of order m. Fibonacci Quarterly, 5(5):461–465, December 1967. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-5/ford.pdf.

Fishburn:1989:TSG

[FOR89] Peter C. Fishburn, Andrew M. Odlyzko, and Fred S. Roberts. Two-sided generalized Fibonacci sequences. Fibonacci Quarterly, 27(4):352–361, August 1989. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/27-4/fishburn.pdf.

Foster:1966:LBM

[Fos66] B. L. Foster. A lower bound for maximum zero-one determinants. Fibonacci Quarterly, 4(2):187–189, April 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/4-2/foster.pdf.

Fowler:1982:GGS

[Fow82] D. H. Fowler. A generalization of the golden section. Fibonacci Quarterly, 20(2):146–158, May 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-2/
fowler.pdf.

Fox:2001:CRR

[Fox01] Glenn J. Fox. Congruences relating rational values of Bernoulli and Euler polynomials. Fibonacci Quarterly, 39(1):50–57, February 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/39-1/fox.pdf.

Fox:2021:ISF

[Fox21] Glenn J. Fox. Integers that satisfy a Fermat’s congruence of higher power. Fibonacci Quarterly, 59(4):291–??, November 2021.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/59-4/fox.pdf.

Freitag:1986:CRC

[FP86] H. T. Freitag and G. M. Phillips. A congruence relation for certain recursive sequences. Fibonacci Quarterly, 24(4):332–335, November 1986. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/24-4/freitag.pdf.

Freitag:1996:ZF

[FP96] H. T. Freitag and G. M. Phillips. On the Zeckendorf form of F_{kn}/F_{n}. Fibonacci Quarterly, 34(5):444–446, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-5/freitag.pdf.

Falcon:2009:FNA

[FP09] Sergio Falcon and Angel Plaza. On k-Fibonacci numbers of arithmetic indexes. Applied Mathematics and Computation, 208(1): 180–185, February 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Foundas:2008:LPS

[FPC08] E. Foundas, C. Patsakis, and G. Chondrocoukis. Lucas permutations and some notes on Fibonacci permutations. Journal of Discrete Mathematical Sciences and Cryptography, 11(2):181– 190, April 2008. CODEN ???? ISSN 0972-0529. URL http:

//www.connectjournals.com/achivestoc.php?bookmark=CJ-
003072&volume=11&issue_id=02.

Federighi:1966:LE

[FR66] Enrico T. Federighi and Ronald G. Roll. A letter to the editor. Fibonacci Quarterly, 4(1):85–87, February 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-1/
federighi.pdf.

Franci:1985:THA

[FR85] R. Franci and L. Toti Rigatelli. Towards a history of algebra from Leonardo of Pisa to Luca Pacioli. Janus, 72(1–3):17–82, ???? 1985. CODEN JNUSA6. ISSN 0021-4264.

Frajese:1940:LGL

[Fra40] Attilio Frajese. L’algebra geometrica in Leonardo Pisano. (Italian) [The geometric algebra of Leonardo Pisano]. Boll. Un. Mat. Ital. (2), 2:363–365, 1940.

Fray:1967:GFA

[Fra67] Robert Fray. A generating function associated with the generalized Stirling numbers. Fibonacci Quarterly, 5(4):356–366, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/5-4/fray.pdf.

Frankel:1970:FNP

[Fra70] Edward T. Frankel. Fibonacci numbers as paths of rooks on a chessboard. Fibonacci Quarterly, 8(5):538–541, December 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-5/frankel.pdf.

Frankel:1977:BIP

[Fra77] Edward T. Frankel. Bell’s imperfect perfect numbers. Fibonacci Quarterly, 15(4):336–??, December 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
4/frankel.pdf.

Frame:1980:FBC

[Fra80] J. S. Frame. Factors of the binomial circulant determinant. Fibonacci Quarterly, 18(1):9–23, February 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-1/
frame.pdf.

Fraenkel:1984:WGC

[Fra84] A. S. Fraenkel. Wythoff games, continued fractions, cedar trees and Fibonacci searches. Theoretical Computer Science, 29(1-2): 49–73, March 1984. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Frappier:1987:DCF

[Fra87] Clement Frappier. On the derivatives of composite functions. Fibonacci Quarterly, 25(3):229–239, August 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-3/frappier.pdf.

Frappier:1991:IKE

[Fra91] Clément Frappier. Iterations of a kind of exponentials. Fibonacci Quarterly, 29(4):351–361, November 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-
4/frappier.pdf.

Frame:1994:FNC

[Fra94] J. S. Frame. Fibonacci numbers and a chaotic piecewise linear function. Fibonacci Quarterly, 32(2):167–169, May 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-2/frame.pdf.

Franco:1998:DBC

[Fra98] Zachary M. Franco. Distribution of binomial coefficients modulo three. Fibonacci Quarterly, 36(3):272–275, June 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-3/franco.pdf.

Franci:2002:BAL

[Fra02] Raffaella Franci. The Liber abaci of Leonardo Fibonacci, 1202– 2002. Bollettino della Unione matematica italiana. A., 5(2):293– 328, 2002. CODEN BLUMAM. ISSN 0392-4033.

Franci:2003:LFT

[Fra03] Raffaella Franci. Leonardo Fibonacci and treatises on the abacus in Italy in the fourteenth and fifteenth centuries. Bollettino di Storia delle Scienze Matematiche, 23(2, Leonardo Fibonacci. Matematica e societa nel Mediterraneo nel secolo XIII. Vol. I): 33–54 (2005), 2003. ISSN 0392-4432 (print), 1724-1650 (electronic).

Freitag:1968:MSI

[Fre68] Herta T. Freitag. A magic square involving Fibonacci numbers. Fibonacci Quarterly, 6(1):77–80, February 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-1/freitag.pdf.

Freitag:1973:SEF

[Fre73] Herta T. Freitag. On summations and expansions of Fibonacci numbers. Fibonacci Quarterly, 11(1):63–71, February 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-1/freitag.pdf.

Frenklach:1985:LRR

[Fre85] Michael Frenklach. Linear recurrence relations with binomial coefficients. Fibonacci Quarterly, 23(4):359–363, November 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-4/frenklach.pdf.

Freitag:1987:SIC

[Fre87] Herta T. Freitag. The Second International Conference on Fibonacci Numbers and Their Applications: a memory-laden experience. Fibonacci Quarterly, 25(2):98–99, May 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-2/freitag.pdf.

Freitag:1988:RTI

[Fre88] Herta T. Freitag. A report on the Third International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 26(4):289–??, November 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
4/report-a.pdf.

Freitag:1990:RFI

[Fre90] Herta T. Freitag. Report on the Fourth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 28(4):354–??, November 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
4/report-a.pdf.

Freguglia:1992:DFP

[Fre92a] Paolo Freguglia. The determination of π in Fibonacci’s Practica geometriae in a fifteenth-century manuscript. In Contributions to the history of mathematics (Italian) (Modena, 1990), volume 8 of Coll. Studi, pages 75–84. Accad. Naz. Sci. Lett. Arti, Modena, Italy, 1992.

Freitag:1992:FCS

[Fre92b] Herta T. Freitag. The Fibonacci Conference in Scotland. Fibonacci Quarterly, 30(4):334–??, November 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-4/report-a.pdf.

Freitag:1994:FCP

[Fre94] Herta T. Freitag. The Fibonacci Conference in Pullman. Fibonacci Quarterly, 32(5):465–466, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-5/freitag.pdf.

Freitag:1997:NSI

[Fre97] Herta T. Freitag. A note on the Seventh International Conference. Fibonacci Quarterly, 35(1):84–??, February 1997. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-1/freitag.pdf.

Freitag:1999:EIC

[Fre99] Herta T. Freitag. The Eighth International Conference. Fibonacci Quarterly, 37(1):46–??, February 1999. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/37-
1/freitag.pdf.

French:2006:FRF

[Fre06] Christopher P. French. Fifth roots of Fibonacci fractions. Fibonacci Quarterly, 44(3):209–215, August 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/44-3/french.pdf.

Freguglia:2022:NEL

[Fre22] Paolo Freguglia. A new edition of Leonardo Fibonacci’s Liber Abbaci. Nuncius, 37(1):207–213, February 2022. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic). URL https://

brill.com/view/journals/nun/37/1/article-p207_8.xml.

Fridy:1966:GBR

[Fri66] J. A. Fridy. Gerneralized bases for the real numbers. Fibonacci Quarterly, 4(3):193–200, October 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-3/
fridy.pdf.

Frontczak:2018:SFL

[Fro18] Robert Frontczak. Some Fibonacci–Lucas–Tribonacci–Lucas identities. Fibonacci Quarterly, 56(3):263–??, August 2018. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/56-3/frontczak.pdf.

Frontczak:2023:APSa

[Fro23a] Robert Frontczak. Advanced problems and solutions. Fibonacci Quarterly, 61(1):91–??, February 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbFeb2023.pdf.

Frontczak:2023:APSb

[Fro23b] Robert Frontczak. Advanced problems and solutions. Fibonacci Quarterly, 61(2):185–??, May 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbMay2023.pdf.

Frontczak:2023:APSc

[Fro23c] Robert Frontczak. Advanced problems and solutions. Fibonacci Quarterly, 61(3):282–??, August 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbAug2023.pdf.

Frontczak:2023:APSd

[Fro23d] Robert Frontczak. Advanced problems and solutions. Fibonacci Quarterly, 61(4):374–??, November 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
NovAdvProb2023.pdf.

Frontczak:2024:APS

[Fro24] Robert Frontczak. Advanced problems and solutions. Fibonacci Quarterly, 62(1):90–??, February 2024. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbFeb2024.pdf.

Flanagan:2015:SFP

[FRU15] Patrick Flanagan, Marc S. Renault, and Josh Updike. Symmetries of Fibonacci points, mod m. Fibonacci Quarterly, 53(1): 34–??, February 2015. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/53-1/flanagan.pdf.

Fibonacci:1987:BS

[FS87] Leonardo Fibonacci and L. E. Sigler. The book of squares. Academic Press, New York, USA, 1987. ISBN 0-12-643130-2. xx + 124 pp. LCCN QA32 .F4813 1987. URL http://www.loc.

gov/catdir/description/els032/86017336.html;   http:/
/www.loc.gov/catdir/toc/els031/86017336.html;   https:
//www.math.utah.edu/pub/tex/bib/fibquart.bib.  Anno-

tated translation into modern English of Liber quadratorum.

Fraenkel:1999:ENS

[FS99] Aviezri S. Fraenkel and Jamie Simpson. The exact number of squares in Fibonacci words. Theoretical Computer Science, 218 (1):95–106, April 28, 1999. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.elsevier.

com/cas/tree/store/tcs/sub/1999/218/1/3061.pdf. See cor-

rigendum [FS14].

Frey:2001:GBG

[FS01] Darrin D. Frey and James A. Sellers. Generalizing Bailey’s generalization of the Catalan numbers. Fibonacci Quarterly, 39(2): 142–148, May 2001. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/39-2/frey.pdf.

Fraenkel:2014:CEN

[FS14] Aviezri S. Fraenkel and Jamie Simpson. Corrigendum to “The exact number of squares in Fibonacci words” [Theoret. Comput. Sci. 218 (1) (1999) 95–106]. Theoretical Computer Science, 547(??):122, August 28, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397514004356. See [FS99].

Fredman:1984:FHT

[FT84] M. L. Fredman and R. E. Tarjan. Fibonacci heaps and their uses in improved network optimization algorithms. In IEEE [IEE84], pages 338–346. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

Fredman:1987:FHT

[FT87] Michael L. Fredman and Robert Endre Tarjan. Fibonacci heaps and their uses in improved network optimization algorithms. Journal of the Association for Computing Machinery, 34(3):596–615, July 1987. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/28874.html;   https:
//www.math.utah.edu/pub/tex/bib/fibquart.bib.

Firengiz:2014:SM

[FT14] M. Cetin Firengiz and Naim Tuglu. On the q-Seidel matrix. Fibonacci Quarterly, 52(5):117–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers1/52-5/Firengiz.pdf.

Fuchs:1984:PCC

[Fuc84] Eduard Fuchs. Partitions, compositions and cyclomatic number of function lattices. Fibonacci Quarterly, 22(1):42–49, February 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/22-1/fuchs.pdf.

Fults:1977:SCR

[Ful77] Douglas A. Fults. Solution of a certain recurrence relation. Fibonacci Quarterly, 15(1):41–??, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/15-1/fults-a.pdf.

Fuller:1978:VWE

[Ful78] Leonard E. Fuller. Vectors whose elements belong to a generalized Fibonacci sequence. Fibonacci Quarterly, 16(5):447–450, October 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/16-5/fuller.pdf.

Fuller:1980:GRR

[Ful80a] Leonard E. Fuller. Geometric recurrence relation. Fibonacci Quarterly, 18(2):126–128, April 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-
2/fuller1.pdf.

Fuller:1980:RRR

[Ful80b] Leonard E. Fuller. Representations for r, s recurrence relations. Fibonacci Quarterly, 18(2):129–134, April 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-2/fuller2.pdf.

Fuller:1981:GFR

[Ful81a] Leonard E. Fuller. Generating functions for recurrence relations. Fibonacci Quarterly, 19(2):106–109, April 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-2/fuller.pdf.

Fuller:1981:SGR

[Ful81b] Leonard E. Fuller. Solutions for general recurrence relations. Fibonacci Quarterly, 19(1):64–68, February 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-1/fuller2.pdf.

Fuller:1981:SOR

[Ful81c] Leonard E. Fuller. A special m th-order recurrence relation. Fibonacci Quarterly, 19(1):24–27, February 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-1/fuller1.pdf.

Finkelstein:1978:FNC

[FW78] Mark Finkelstein and Robert Whitley. Fibonacci numbers in coin tossing sequences. Fibonacci Quarterly, 16(6):539–540, December 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/16-6/finkelstein.pdf.

Franco:2001:DCC

[FZ01] B. J. O. Franco and Antônio Zumpano. Divisibility of the coefficients of Chebyshev polynomials by primes. Fibonacci Quarterly, 39(4):304–308, August 2001. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/39-4/franco.pdf.

Feng:2003:CFC

[FZ03] Hong Feng and Zhizheng Zhang. Computational formulas for convoluted generalized Fibonacci and Lucas numbers. Fibonacci Quarterly, 41(2):144–151, May 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-2/
feng.pdf.

Fu:2008:MRF

[FZ08] Xudan Fu and Xia Zhou. On matrices related with Fibonacci and Lucas numbers. Applied Mathematics and Computation, 200 (1):96–100, June 15, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Fu:2008:ELC

[FZZ08] Xu Dan Fu, Xia Zhou, and Xiao Dong Zhao. Extensions of Lucas’ congruence and Babbage’s congruence. Acta Math. Sinica (Chin. Ser.), 51(4):693–698, 2008. CODEN SHHPBO. ISSN 0583-1431.

Gabai:1970:GFS

[Gab70] Hyman Gabai. Generalized Fibonacci k-sequences. Fibonacci Quarterly, 8(1):31–38, February 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-1/
gabai.pdf.

Gale:1968:EFI

[Gal68a] Gene B. Gale. Errata for “Factorization of 2×2 Integral Matrices with Determinant ±1”. Fibonacci Quarterly, 6(5):10–??, November 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/6-5/errata-a.pdf. See [Gal68b].

Gale:1968:FIM

[Gal68b] Gene B. Gale. Factorization of 2 × 2 integral matrices with determinant ±1. Fibonacci Quarterly, 6(1):3–20, February 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-1/gale.pdf. See errata [Gal68a].

Galambos:1972:CUT

[Gal72] Janos Galambos. A constructive uniqueness theorem on representing integers. Fibonacci Quarterly, 10(6):569–570, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/10-6/galambos-a.pdf.

Gallinar:1979:FRT

[Gal79] Jean-Pierre Gallinar. Fibonacci ratio in a thermodynamical case. Fibonacci Quarterly, 17(3):239–240, October 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-3/gallinar.pdf.

Gallinar:1986:FRT

[Gal86] J.-P. Gallinar. The Fibonacci ratio in a thermodynamical problem: a combinatorial approach. Fibonacci Quarterly, 24(3):247– 250, August 1986. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/24-3/gallinar.pdf.

Galuzzi:1998:RAB

[Gal98] Massimo Galuzzi. A remark about the binomial transform. Fibonacci Quarterly, 36(3):287–289, June 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-3/
galuzzi.pdf.

Galuzzi:2008:BRA

[Gal08] Massimo Galuzzi. Book review: Alfred S. Posamentier and Ingmar Lehmann, The (Fabulous) Fibonacci Numbers. New York: Prometheus Books, 2007. 385 pp., ISBN 976-1-59102-475-0. Nuncius, 23(1):188–189, ???? 2008. CO-DEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/
10.1163/182539108x00544.

Gamkrelidze:1995:PPF

[Gam95] N. G. Gamkrelidze. On a probabilistic property of the Fibonacci sequence. Fibonacci Quarterly, 33(2):147–152, May 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-2/gamkrelidze.pdf.

Ganis:1959:CNN

[Gan59] Sam E. Ganis. Classroom notes: Notes on the Fibonacci sequence. American Mathematical Monthly, 66(2):129–130, February 1959. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Garcia:2004:HNF

[Gar04] Mariano Garcia. Hyperperfect numbers with five and six different prime factors. Fibonacci Quarterly, 42(4):292–294, November 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers1/42-4/quartgarcia04_2004.pdf.

Gauthier:1989:DF

[Gau89] N. Gauthier. Derivation of a formula for ∑ r^{k}x^{r}. Fibonacci Quarterly, 27(5):402–407, November 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-
5/gauthier.pdf.

Gauthier:1998:ICS

[Gau98] N. Gauthier. Identities for a class of sums involving Horadam’s generalized numbers {W_{n}}. Fibonacci Quarterly, 36(4):295–304, August 1998. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/36-4/gauthier.pdf.

Gauthier:2002:BRB

[Gau02] Napoleon Gauthier. Book review: Fibonacci and Lucas Numbers with Applications, by Thomas Koshy. Fibonacci Quarterly, 40 (1):55–??, February 2002. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/40-1/review.pdf. See

[Kos01].

Gauthier:2004:CPC

[Gau04] N. Gauthier. Convolving the m-th powers of the consecutive integers with the general Fibonacci sequence using Carlitz’s weighted Stirling polynomials of the second kind. Fibonacci Quarterly, 42(4):306–313, November 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
4/quartgauthier04_2004.pdf.

Gauthier:2010:TKB

[Gau10] Napoleon Gauthier. Thomas Koshy’s Catalan Numbers with Applications (book review). Fibonacci Quarterly, 48(1):85–86, February 2010. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Papers/48-1/Gauthier_Review.pdf. See

[Kos09].

Gauthier:2011:FPI

[Gau11] N. Gauthier. 95.50 Fractional-power identities for Fibonacci and Lucas polynomials and numbers. The Mathematical Gazette, 95 (534):486–494, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Gauthier:2012:PFA

[Gau12] N. Gauthier. Polynomial forms for alternating sums of products of binomial-Catalan numbers. Fibonacci Quarterly, 50(1):62–??, February 2012. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/50-1/gauthier.pdf.

Gauthier:2006:SEI

[GB06] N. Gauthier and Paul S. Bruckman. Sums of the even integral powers of the cosecant and secant. Fibonacci Quarterly, 44(3): 263–273, August 2006. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/44-3/gauthier.pdf.

Griffiths:2015:JNT

[GB15] Martin Griffiths and Alex Bramham. The Jacobsthal numbers: Two results and two questions. Fibonacci Quarterly, 53(2):147– ??, May 2015. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/53-2/griffiths.pdf.

Graves:2016:UMD

[GBC16] Rachel K. Graves, Michael R. Bacon, and Charles K. Cook. Using matrices to derive identities for recursive sequences. Fibonacci Quarterly, 54(3):204–??, August 2016. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/54-3/
graves.pdf.

Gbur:1981:GPS

[Gbu81] Mary E. Gbur. A generalization of a problem of Stolarsky. Fibonacci Quarterly, 19(2):117–120, April 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-2/
gbur.pdf.

Gutman:1990:RFR

[GC90] Ivan Gutman and Sven J. Cyvin. A result on 1-factors related to Fibonacci numbers. Fibonacci Quarterly, 28(1):81–84, February 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/28-1/gutman.pdf.

Gueganic:2019:LDG

[GCK^{+}19] Alexandre Gueganic, Granger Carty, Yujin H. Kim, Steven J. Miller, Alina Shubina, Shannon Sweitzer, Eric Winsor, and Jianing Yang. Limiting distributions in generalized Zeckendorf decompositions. Fibonacci Quarterly, 57(2):109–??, May 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-2/gueganic.pdf.

Garcia-Caballero:2014:CVV

[GCMP14a] Esther M. García-Caballero, Samuel G. Moreno, and Michael P. Prophet. A complete view of Viète-like infinite products with Fibonacci and Lucas numbers. Applied Mathematics and Computation, 247(??):703–711, November 15, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300314012363.

Garcia-Caballero:2014:NVL

[GCMP14b] Esther M. García-Caballero, Samuel G. Moreno, and Michael P. Prophet. New Viète-like infinite products of nested radicals with Fibonacci and Lucas numbers. Fibonacci Quarterly, 52(1):27–??, February 2014. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/52-1/garcia-caballero.pdf.

Gatta:2008:SWA

[GD09] F. Gatta and A. D’Amico. Sequences H_{n} for which H_{n+1}/H_{n} approaches the golden ratio. Fibonacci Quarterly, 46/47(4): 346–349, November 2008/2009. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/46_47-
4/gatta.pdf.

Gebhardt:1967:GPR

[Geb67] Friedrich Gebhardt. Generating pseudo-random numbers by shuffling a Fibonacci sequence (in Technical Notes and Short Papers). Mathematics of Computation, 21(100):708–709, October 1967. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gegenbauer:1893:BLP

[Geg93] Leopold Gegenbauer. Bemerkung über Leonardo Pisano’s liber Abaci. (German) [Remark on Leonardo Pisano’s liber Abaci]. Monatshefte für Mathematik und Physik, 4(1):402, 1893. CO-DEN MMPHA8. ISSN 0026-9255 (print), 1436-5081 (electronic).

Geller:1963:CIP

[Gel63] Stephen P. Geller. A computer investigation of a property of the Fibonacci sequence. Fibonacci Quarterly, 1(2):84–??, April 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/1-2/geller.pdf.

Geldenhuys:1981:FNM

[Gel81] G. Geldenhuys. On the Fibonacci numbers minus one. Fibonacci Quarterly, 19(5):456–457, December 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
5/geldenhuys.pdf.

Georgieva:1975:DZO

[Geo75] N. Georgieva. Distribution of the zeroes of one class of polynomials. Fibonacci Quarterly, 13(4):312–314, December 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-4/georgieva.pdf.

Georghiou:1989:SSO

[Geo89] C. Georghiou. On some second-order linear recurrences. Fibonacci Quarterly, 27(2):156–159, May 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-2/
georghiou.pdf.

Gerdes:1977:CGF

[Ger77] Walter Gerdes. Convergent generalized Fibonacci sequences. Fibonacci Quarterly, 15(2):156–160, April 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-2/
gerdes.pdf.

Gerdes:1978:GTN

[Ger78] Walter Gerdes. Generalized Tribonacci numbers and their convergent sequences. Fibonacci Quarterly, 16(3):269–275, June 1978.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-3/gerdes.pdf.

Gerdemann:2008:CPZ

[Ger09] Dale Gerdemann. Combinatorial proofs of Zeckendorf family identities. Fibonacci Quarterly, 46/47(3):249–260, August 2008/ 2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/46_47-3/gerdemann.pdf.

Gessel:1981:CBT

[Ges81] Ira Gessel. Congruences for Bell and tangent numbers. Fibonacci Quarterly, 19(2):137–143, April 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-2/
gessel.pdf.

Gueunet:2019:TBA

[GFJT19] C. Gueunet, P. Fortin, J. Jomier, and J. Tierny. Task-based augmented contour trees with Fibonacci heaps. IEEE Transactions on Parallel and Distributed Systems, 30(8):1889–1905, August 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Garcia-Fernandezsesma:2024:AZG

[GFMR^{+}24] Diego Garcia-Fernandezsesma, Steven J. Miller, Thomas Rascon, Risa Vandegrift, and Ajmain Yamin. The accelerated Zeckendorf game. Fibonacci Quarterly, 62(1):3–??, February 2024. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/62-1/garcia-fernandezsesma.pdf.

Gies:1969:LPN

[GG69] Joseph Gies and Frances Gies. Leonard of Pisa and the new mathematics of the Middle Ages. Crowell, New York, NY, USA, 1969. 127 pp. LCCN QA29.F5 G5. Illustrated by Enrico Arno.

Giuli:1979:PSDa

[GG79a] Christine Giuli and Robert Giuli. A primer on Stern’s diatomic sequence. Fibonacci Quarterly, 17(2):103–107, April 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-2/giuli.pdf.

Giuli:1979:PSDb

[GG79b] Christine Giuli and Robert Giuli. A primer on Stern’s diatomic sequence — II. Fibonacci Quarterly, 17(3):246–247, October 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/17-3/giuli.pdf.

Giuli:1979:PSDc

[GG79c] Christine Giuli and Robert Giuli. A primer on Stern’s diatomic sequence: Part III: Additional results. Fibonacci Quarterly, 17 (4):318–320, December 1979. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/17-4/giuli.
pdf.

Gould:1985:LTP

[GG85] H. W. Gould and W. E. Greig. A Lucas triangle primality criterion dual to that of Mann–Shanks. Fibonacci Quarterly, 23(1): 66–69, February 1985. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/23-1/gould.pdf.

Griffiths:2013:FRS

[GG13] Jonny Griffiths and Martin Griffiths. Fibonacci-related sequences via iterated QRT maps. Fibonacci Quarterly, 51(3):218–??, August 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/51-3/griffiths.pdf.

Gordon:2015:CPT

[GG15] Russell A. Gordon and Sara L. Graham. Comments on proofs that there are no four squares in arithmetic progression. Fibonacci Quarterly, 53(1):68–??, February 2015. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/53-1/
gordon.pdf.

Gies:1983:LPN

[GGA83] Joseph Gies, Frances Gies, and Enrico Arno. Leonard of Pisa and the new mathematics of the Middle Ages. New Classics Library, Gainesville, GA, USA, 1983. ISBN 0-317-57849-9. 127 pp. LCCN ????

Gomez:2021:TRF

[GGL21] Carlos A. Gómez, Jhonny C. Gómez, and Florian Luca. Two b-repunits in the Fibonacci sequence. Journal of Number Theory, 222(??):393–422, May 2021. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X20303309.

Garcia:2024:SSP

[GGL24] Jonathan García, Carlos A. Gómez, and Florian Luca. Solving Skolem’s problem for the k-generalized Fibonacci sequence with negative indices. Journal of Number Theory, 257(??):273– 299, April 2024. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X23002184.

Grundman:2018:SCH

[GH18] Helen G. Grundman and Pamela E. Harris. Sequences of consecutive happy numbers in negative bases. Fibonacci Quarterly, 56(3): 221–??, August 2018. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/56-3/grundman.pdf.

Gica:2008:QRF

[Gic09] Alexandru Gica. Quadratic residues in Fibonacci sequences. Fibonacci Quarterly, 46/47(1):68–72, February 2008/2009. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/46_47-1/gica.pdf.

Gica:2022:CMS

[Gic22] Alexandru Gica. Congruences modulo the square of a prime for sums containing Fibonacci numbers. Fibonacci Quarterly, 60(3): 243–??, August 2022. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/60-3/gica.pdf.

Gies:2013:LP

[Gie13] Frances Carney Gies. Leonardo Pisano. Encyclopedia Britannica online article., 2013. URL http://www.britannica.com/

EBchecked/topic/336467/Leonardo-Pisano.

Gillespie:1989:GFL

[Gil89] Frank S. Gillespie. A generalization of Fermat’s Little Theorem. Fibonacci Quarterly, 27(2):109–115, May 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/27-2/gillespie.pdf.

Gillespie:1992:GKC

[Gil92a] Frank S. Gillespie. A generalization of Kummer’s congruences and related results. Fibonacci Quarterly, 30(4):349–367, November 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/30-4/gillespie.pdf.

Gillman:1992:M

[Gil92b] Leonard Gillman. Miscellanea. Fibonacci Quarterly, 30(2):102– ??, May 1992. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/30-2/miscellanea.pdf.

Giles:2005:BLN

[Gil05] David E. Giles. Benford’s Law and naturally occurring prices in certain ebaY auctions. Econometrics Working Paper EWP0505, Department of Economics, University of Victoria, Victoria, BC,

Canada, May 2005. URL http://web.uvic.ca/econ/ewp0505.
pdf.

Giles:2007:BLN

[Gil07] D. E. Giles. Benford’s Law and naturally occurring prices in certain eBay auctions. Applied Economics Letters, 14(3):157– 161, 2007. CODEN ???? ISSN 1350-4851 (print), 1466-4291

(electronic). URL http://www.tandfonline.com/toc/rael20/
current.

Ginsburg:1954:RBC

[Gin54] J. Ginsburg. A relationship between cubes of Fibonacci numbers. Scripta Mathematica, 19(??):242, December 1954. ISSN 0036-9713.

Ginat:2002:AIP

[Gin02] David Ginat. Aha! an illuminating perspective. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(1):1–2, March 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Inroads: paving the way towards excellence in computing education.

Giordano:1994:NCP

[Gio94] George Giordano. A note on consecutive prime numbers. Fibonacci Quarterly, 32(4):352–355, August 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-4/giordano.pdf.

Girse:1981:ICP

[Gir81] Robert D. Girse. Identities for certain partition functions and their differences. Fibonacci Quarterly, 19(4):361–368, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-4/girse.pdf.

Girse:1983:NFC

[Gir83] Robert D. Girse. A note on Fibonacci cubature. Fibonacci Quarterly, 21(2):129–131, May 1983. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/21-2/girse.
pdf.

Giuli:1971:BFL

[Giu71] Robert M. Giuli. Binet forms by Laplace transform. Fibonacci Quarterly, 9(1):41–50, February 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-1/
giuli-a.pdf.

Giuli:1972:LHD

[Giu72] Robert M. Giuli. Linear homogeneous difference equations. Fibonacci Quarterly, 10(3):265–270, April 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-3/
giuli-a.pdf.

Giusti:2004:LFM

[Giu04] Enrico Giusti. Leonardo Fibonacci and the mathematical renaissance in the West. In Mathematics, culture and society. 2003 (Italian), Pubbl. Cent. Ric. Mat. Ennio Giorgi, pages 1– 17. Scuola Norm. Sup., Pisa, 2004.

Glasser:1976:SIT

[GK76] M. L. Glasser and M. S. Klamkin. On some inverse tangent summations. Fibonacci Quarterly, 14(5):385–387, December 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-5/glasser.pdf.

Govindaraju:1994:FN

[GKD94] Rama K. Govindaraju, M. S. Krishnamoorthy, and Narsingh Deo. Fibonacci networks. Fibonacci Quarterly, 32(4):329–345, August 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/32-4/govindaraju.pdf.

Gould:1977:SAA

[GKH77] H. W. Gould, J. B. Kim, and V. E. Hoggatt, Jr. Sequences associated with t-ary coding of Fibonacci’s rabbits. Fibonacci Quarterly, 15(4):311–318, December 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
4/gould.pdf.

Gilson:2021:ZTU

[GKL^{+}21] Amelia Gilson, Hadley Killen, Tamás Lengyel, Steven J. Miller, Nadia Razek, Joshua M. Siktar, and Liza Sulkin. Zeckendorf’s theorem using indices in an arithmetic progression. Fibonacci Quarterly, 59(4):327–??, November 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/59-
4/gilson.pdf.

Graham:1989:CM

[GKP89] Ronald L. Graham, Donald E. Knuth, and Oren Patashnik. Concrete Mathematics. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-14236-8. xiii + 625 pp. LCCN QA39.2 .G7331 1989.

Graham:1994:CM

[GKP94] Ronald L. Graham, Donald E. Knuth, and Oren Patashnik. Concrete Mathematics. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-55802-5. xiii + 657 pp. LCCN QA39.2 .G733 1994.

Gilbert:2001:FDQ

[GKRS01] C. L. Gilbert, J. D. Kolesar, C. A. Reiter, and J. D. Storey. Function digraphs of quadrative maps modulo p. Fibonacci Quarterly, 39(1):32–49, February 2001. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/39-1/gilbert.pdf.

Gries:1980:CFN

[GL80] David Gries and Gary Levin. Computing Fibonacci numbers (and similarly defined functions) in log time. Information Processing Letters, 11(2):68–69, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Garcia:1983:GEF

[GL83] P. G. Garcia and Steve Ligh. A generalization of Euler’s ϕfunction. Fibonacci Quarterly, 21(1):26–28, February 1983. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-1/garcia.pdf.

Gessel:2001:OSN

[GL01] Ira M. Gessel and Tamás Lengyel. On the order of Stirling numbers and alternating binomial coefficient sums. Fibonacci Quarterly, 39(5):444–454, November 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-
5/gessel.pdf.

Garcia:2016:QFN

[GL16] Stephan Ramon Garcia and Florian Luca. Quotients of Fibonacci numbers. American Mathematical Monthly, 123(10):1039–1044, December 2016. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/
10.4169/amer.math.monthly.123.10.1039.

Glaze:1970:AMP

[Gla70] Janet Waterman Glaze. Angle multisection by parallel straightedges (recreational mathematics edited by Joseph S. Madachy). Fibonacci Quarterly, 8(4):393–396, October 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/8-4/glaze.pdf.

Gladwin:1978:EFN

[Gla78] A. S. Gladwin. Expansion of the Fibonacci number F_{nm} in n th powers of Fibonacci or Lucas numbers. Fibonacci Quarterly, 16(3):213–215, June 1978. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/16-3/gladwin.pdf.

Glasson:1995:RFI

[Gla95] Alan R. Glasson. Remainder formulas involving generalized Fibonacci and Lucas polynomials. Fibonacci Quarterly, 33(3):268– 272, June 1995. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/33-3/glasson.pdf.

Gleijeses:1981:PMI

[Gle81] B. Gleijeses. Points at mutual integral distances in S^{n}. Fibonacci Quarterly, 19(2):153–159, April 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-2/
gleijeses.pdf.

Greene:2022:UFF

[GLMS22] John Greene, Junhyun Lim, Shaunak Mashalkar, and Edward F. Schaefer. Using Fibonacci factors to create Fibonacci pseudoprimes. Fibonacci Quarterly, 60(4):320–??, November 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-4/greene.pdf.

Glushkov:1976:AML

[Glu76] Stanislaw Glushkov. On approximation methods of Leonardo Fibonacci. Historia Mathematica, 3(3):291–296, August 1976. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0315086076900987.

Gregory:1978:FSS

[GM78] M. B. Gregory and J. M. Metzger. Fibonacci sine sequences. Fibonacci Quarterly, 16(2):119–120, April 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-2/
gregory.pdf.

Gill:1981:NMR

[GM81] John Gill and Gary Miller. Newton’s method and ratios of Fibonacci numbers. Fibonacci Quarterly, 19(1):1–3, February 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-1/gill.pdf.

Guy:1996:NSI

[GM96] Richard K. Guy and William O. J. Moser. Numbers of sequences without isolated odd members. Fibonacci Quarterly, 34(2):152– 155, May 1996. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/34-2/guy.pdf.

Ghalayini:2011:GPH

[GM11] Bassem Ghalayini and Joseph Malkoun. Golden proportions in higher dimensions. Fibonacci Quarterly, 49(3):267–272, August 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/49-3/ghalayini.pdf.

Givens:2020:IFP

[GM20] Berit Givens and Victor H. Moll. Integrals of Fibonacci polynomials and their valuations. Fibonacci Quarterly, 58(3):261–??, August 2020. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/58-3/givens.pdf.

Giambalvo:1991:ACB

[GMP91] V. Giambalvo, Ray Mines, and David J. Pengelley. p-adic congruences between binomial coefficients. Fibonacci Quarterly, 29(2): 114–119, May 1991. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/29-2/giambalvo.pdf.

Golic:1995:DLR

[Gol95] Jovan Dj. Golić. On decimation of linear recurring sequences. Fibonacci Quarterly, 33(5):407–411, November 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-5/golic.pdf.

Goli:2006:FND

[Gol06] Jovan Dj. Goli. Fibonacci numbers and decimation of binary sequences. Fibonacci Quarterly, 44(3):216–223, August 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/44-3/golic.pdf.

Gonnet:1989:PAI

[Gon89] Gaston H. Gonnet, editor. Proceedings of the ACM-SIGSAM 1989 International Symposium on Symbolic and Algebraic Computation: ISSAC ’89 / July 17–19, 1989, Portland, Oregon. ACM Press, New York, NY 10036, USA, 1989. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Gootherts:1968:LACa

[Goo68a] J. W. Gootherts. Linear algebra constructed from Fibonacci sequences Part I: Fundamentals and polynomial interpretations. Fibonacci Quarterly, 6(5):35–42, November 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/6-5/gootherts1.pdf.

Gootherts:1968:LACb

[Goo68b] J. W. Gootherts. Linear algebra constructed from Fibonacci sequences. Part II: Function sequences and Taylor series of function sequences. Fibonacci Quarterly, 6(5):44–54, November 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-5/gootherts2.pdf.

Good:1971:SFF

[Goo71] Robert C. Good, Jr. A series form for the Fibonacci numbers F_{12n}. Fibonacci Quarterly, 9(4):405–411, October 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-4/good.pdf.

Good:1974:RSF

[Goo74a] I. J. Good. A reciprocal series of Fibonacci numbers. Fibonacci Quarterly, 12(4):346–??, December 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
4/good.pdf.

Goodwin:1974:PSC

[Goo74b] Norris Goodwin. Power series and cyclic decimals. Fibonacci Quarterly, 12(4):347–??, December 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
4/goodwin.pdf.

Good:1975:NOC

[Goo75] I. J. Good. The number of orderings of n candidates when ties are permitted. Fibonacci Quarterly, 13(1):11–18, February 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-1/good.pdf.

Good:1986:SCTa

[Goo86a] I. J. Good. Skew circulants and the theory of numbers. Fibonacci Quarterly, 24(1):47–60, February 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-
1/good.pdf.

Good:1986:SCTb

[Goo86b] I. J. Good. Skew circulants and the theory of numbers: An addendum. Fibonacci Quarterly, 24(2):176–177, May 1986. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-2/good.pdf.

Goodman:1986:EFG

[Goo86c] A. W. Goodman. An entire function that gives the Fibonacci numbers at the integers. Fibonacci Quarterly, 24(2):145–149, May 1986. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/24-2/goodman.pdf.

Goodall:1991:EAC

[Goo91] Colin Goodall. Eigenshape analysis of a cut-grow mapping for triangles, and its application to phyllotaxis in plants. SIAM Journal on Applied Mathematics, 51(3):775–798, June 1991. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Good:1993:CTF

[Goo93a] I. J. Good. C399. Trigonometrical factorization of the Fibonacci and Lucas numbers, and sums of powers of cotangents. Journal of Statistical Computation and Simulation, 47(1–2):94–99, 1993. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1993:CFL

[Goo93b] I. J. Good. Complex Fibonacci and Lucas numbers, continued fractions, and the square root of the golden ratio. Fibonacci Quarterly, 31(1):7–20, February 1993. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/31-1/good.pdf.

See erratum [Goo93c].

Good:1993:ECF

[Goo93c] I. J. Good. Erratum for “Complex Fibonacci and Lucas Numbers, Continued Fractions, and the Square Root of the Golden Ratio”. Fibonacci Quarterly, 31(3):274–??, August 1993. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-3/good.pdf. See [Goo93b].

Good:1994:CFL

[Goo94a] I. J. Good. C416. Fibonacci and Lucas numbers and trigonometry. Journal of Statistical Computation and Simulation, 49(3–4): 237, 1994. CODEN JSCSAJ. ISSN 0094-9655 (print), 1026-7778 (electronic), 1563-5163.

Good:1994:SPA

[Goo94b] I. J. Good. A symmetry property of alternating sums of products of reciprocals. Fibonacci Quarterly, 32(3):284–287, June 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-3/good.pdf.

Gorroochurn:2012:BPB

[Gor12a] Prakash Gorroochurn. Benford and the peculiar behavior of the first significant digit (1938). In Classic Problems of Probability [Gor12b], chapter 27, pages 233–239. ISBN 1-118-06325-2 (paperback), 1-118-31432-8, 1-118-31433-6 (e-book), 1-118-31434-4 (e-book), 1-118-31431-X (e-book). LCCN QA273.A4.

Gorroochurn:2012:CPP

[Gor12b] Prakash Gorroochurn. Classic Problems of Probability. John Wiley, New York, NY, USA, 2012. ISBN 1-118-06325-2 (paperback), 1-118-31432-8, 1-118-31433-6 (e-book), 1-118-31434-4 (e-book), 1-118-31431-X (e-book). xi + 314 pp. LCCN QA273.A4.

Gould:1963:FEG

[Gou63a] H. W. Gould. Fibonacci exponentials and generalizations of Hermite polynomials. Fibonacci Quarterly, 1(4):31–34, December 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/1-4/gould.pdf.

Gould:1963:GFP

[Gou63b] H. W. Gould. Generating functions for products of powers of Fibonacci numbers. Fibonacci Quarterly, 1(2):1–16, April 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-2/gould.pdf.

Gould:1963:ORI

[Gou63c] H. W. Gould. Operational recurrences involving Fibonacci numbers. Fibonacci Quarterly, 1(1):30–33, February 1963. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-1/gould.pdf.

Gould:1964:AGS

[Gou64a] H. W. Gould. Associativity and the golden section. Fibonacci Quarterly, 2(3):203–??, October 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-3/
gould.pdf.

Gould:1964:BCB

[Gou64b] H. W. Gould. Binomial coefficients, the bracket function, and composition with relatively prime summands. Fibonacci Quarterly, 2(4):241–259, December 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-4/
gould.pdf.

Gould:1965:NFN

[Gou65a] H. W. Gould. Non-Fibonacci numbers. Fibonacci Quarterly, 3 (3):177–183, October 1965. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/3-3/gould.pdf.

Gould:1965:VPT

[Gou65b] H. W. Gould. A variant of Pascal’s triangle. Fibonacci Quarterly, 3(4):257–271, December 1965. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/3-4/gould.pdf.

Gould:1966:FCPa

[Gou66a] H. W. Gould. A Fibonacci crossword puzzle. Fibonacci Quarterly, 4(1):59–62, February 1966. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/4-1/gould.pdf.

Gould:1966:FCPb

[Gou66b] H. W. Gould. A Fibonacci crossword puzzle. Fibonacci Quarterly, 4(2):150–??, April 1966. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/4-2/gould.pdf.

Gould:1967:BFB

[Gou67a] H. W. Gould. The bracket function, q-binomial coefficients, and some new Stirling number formulas. Fibonacci Quarterly, 5(5): 401–423, December 1967. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/5-5/gould.pdf.

Gould:1967:NCI

[Gou67b] H. W. Gould. Note on a combinatorial identity in the theory of bi-colored graphs. Fibonacci Quarterly, 5(3):247–250, October 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/5-3/gould.pdf.

Gould:1968:NPP

[Gou68] H. W. Gould. Note on a paper of Paul F. Byrd, and a solution of Problem P-3. Fibonacci Quarterly, 6(6):318–321, December 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/6-6/gould.pdf.

Gould:1969:BFF

[Gou69] H. W. Gould. The bracket function and Fontené–Ward generalized binomial coefficients with application to Fibonomial coefficients. Fibonacci Quarterly, 7(1):23–40, February 1969. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-1/gould-a.pdf.

Gougenheim:1971:ALS

[Gou71a] André Gougenheim. About the linear sequence of the integers such that each term is the sum of the two preceding. Fibonacci Quarterly, 9(3):277–295, May 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-3/
gougenheim-a.pdf.

Gould:1971:EPG

[Gou71b] H. W. Gould. Equal products of generalized binomial coefficients. Fibonacci Quarterly, 9(4):337–346, October 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-4/gould.pdf.

Gould:1972:CSB

[Gou72a] H. W. Gould. The case of the strange binomial identities of Professor Moriarty. Fibonacci Quarterly, 10(4):381–392, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/10-4/gould2-a.pdf.

Gould:1972:NGC

[Gou72b] H. W. Gould. A new greatest common divisor property of the binomial coefficients. Fibonacci Quarterly, 10(6):579–584, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/10-6/gould1-a.pdf.

Gould:1972:NPC

[Gou72c] H. W. Gould. A new primality criterion of Mann and Shanks and its relation to a theorem of Hermite with extension to Fibonomials. Fibonacci Quarterly, 10(4):355–364, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-4/gould1-a.pdf.

Gould:1972:SCI

[Gou72d] H. W. Gould. Some combinatorial identities of Bruckman — a systematic treatment with relation to the older literature. Fibonacci Quarterly, 10(6):613–628, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-6/gould2.pdf.

Gould:1973:TIS

[Gou73] H. W. Gould. A triangle with integral sides and area. Fibonacci Quarterly, 11(1):27–39, February 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
1/gould.pdf.

Gould:1974:DFB

[Gou74a] H. W. Gould. The design of the four binomial identities: Moriarty intervenes. Fibonacci Quarterly, 12(3):300–308, October 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-3/gould.pdf.

Gould:1974:SSG

[Gou74b] H. W. Gould. Some sequences generated by spiral sieving methods. Fibonacci Quarterly, 12(4):393–397, December 1974. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-4/gould.pdf.

Gould:1974:GHD

[Gou74c] Henry W. Gould. Generalization of Hermite’s divisibility theorems and the Mann–Shanks primality criterion for s-Fibonomial arrays. Fibonacci Quarterly, 12(2):157–166, April 1974. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-2/gould.pdf.

Gould:1975:FPE

[Gou75] Henry W. Gould. Formal proof of equivalence of two solutions of the general Pascal recurrence. Fibonacci Quarterly, 13(2):127– 128, April 1975. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/13-2/gould.pdf.

Gould:1976:LE

[Gou76] H. W. Gould. Letter to the editor. Fibonacci Quarterly, 14(2): 143–??, April 1976. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/14-2/letter-a.pdf.

Gould:1977:FFL

[Gou77a] H. W. Gould. A Fibonacci formula of Lucas and its subsequent manifestations and rediscoveries. Fibonacci Quarterly, 15(1):25– 29, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/15-1/gould1.pdf.

Gould:1977:RSB

[Gou77b] H. W. Gould. A rearrangement of series based on a partition of the natural numbers. Fibonacci Quarterly, 15(1):67–72, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/15-1/gould2.pdf.

Gould:1978:ESC

[Gou78a] H. W. Gould. Evaluation of sum of convolved powers using Stirling and Eulerian numbers. Fibonacci Quarterly, 16(6):488–496, December 1978. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/16-6/gould1-a.pdf.

Gould:1978:OFU

[Gou78b] H. W. Gould. Operational formulas for unusual Fibonacci series. Fibonacci Quarterly, 16(6):555–560, December 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-6/gould2.pdf.

Gould:1981:HFM

[Gou81] H. W. Gould. A history of the Fibonacci Q-matrix and a higherdimensional problem. Fibonacci Quarterly, 19(3):250–256, August 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/19-3/gould.pdf.

Gould:1989:EPP

[Gou89] H. W. Gould. Equivalence of Piza’s primality criterion with that of Gould–Greig and its dual relationship to the Mann–Shanks criterion. Fibonacci Quarterly, 27(4):362–365, August 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-4/gould.pdf.

Gould:1990:STF

[Gou90] H. W. Gould. Series transformations for finding recurrences for sequences. Fibonacci Quarterly, 28(2):166–171, May 1990. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/28-2/gould.pdf.

Gould:1994:BFT

[Gou94] H. W. Gould. A bracket function transform and its inverse. Fibonacci Quarterly, 32(2):176–179, May 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-2/
gould.pdf.

Gould:1995:EHG

[Gou95] H. W. Gould. Extensions of the Hermite G.C.D. theorems for binomial coefficients. Fibonacci Quarterly, 33(5):386–390, November 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/33-5/gould.pdf.

Gould:1999:FWP

[Gou99] Henry W. Gould. The Firard–Waring power sum formulas for symmetric functions, and Fibonacci sequences. Fibonacci Quarterly, 37(2):135–140, May 1999. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/37-2/gould.
pdf.

Gould:2006:IFS

[Gou06] H. W. Gould. The inverse of a finite series and a third-order recurrent sequence. Fibonacci Quarterly, 44(4):302–315, November 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/44-4/gould.pdf.

Georghiou:1983:HSZ

[GP83] C. Georghiou and A. N. Philippou. Harmonic sums and the zeta function. Fibonacci Quarterly, 21(1):29–36, February 1983. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-1/georghiou.pdf.

Gordon:1991:DLE

[GP91] Daniel M. Gordon and Carl Pomerance. The distribution of Lucas and elliptic pseudoprimes. Mathematics of Computation, 57(196): 825–838, October 1991. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Gordon:1993:CDL

[GP93] Daniel M. Gordon and Carl Pomerance. Corrigendum: “The distribution of Lucas Pomerance and elliptic pseudoprimes” [Math. Comp. 57 (1991), no. 196, 825–838, MR 92h:11081]. Mathematics of Computation, 60(202):877, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Grabner:1994:FK

[GP94] Peter J. Grabner and Helmut Prodinger. The Fibonacci killer. Fibonacci Quarterly, 32(5):389–394, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-5/grabner.pdf.

Garth:2016:SSS

[GP16] David Garth and Joseph Palmer. Self-similar sequences and generalized Wythoff arrays. Fibonacci Quarterly, 54(1):72–??, February 2016. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/54-1/garth.pdf.

Gries:1987:HRC

[GPS87] David Gries, Adriano Pascoletti, and Luigi Sbriz. Horner’s rule and the computation of linear recurrences. Information Processing Letters, 25(4):237–240, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Georghiou:2013:MPD

[GPS13] Constantinos Georghiou, Andreas N. Philippou, and Abolfazl Saghafi. On the modes of the Poisson distribution of order k. Fibonacci Quarterly, 51(1):44–??, February 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/51-1/georghiou.pdf.

Gould:2007:PNW

[GQ07] H. W. Gould and Jocelyn Quaintance. Products of numbers which obey a Fibonacci-type recurrence. Fibonacci Quarterly, 45(4):337–346, November 2007. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/45-4/gould.
pdf.

Gould:2008:PFE

[GQ09] H. W. Gould and Jocelyn Quaintance. Partial fraction expansions and a question of Bruckman. Fibonacci Quarterly, 46/47(3):245–248, August 2008/2009. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/46_47-
3/gould.pdf.

Gould:2010:GVI

[GQ10] H. W. Gould and Jocelyn Quaintance. Generalizations of Vosmansky’s identity. Fibonacci Quarterly, 48(1):56–61, February 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-1/gould.pdf.

Gould:2011:IFS

[GQ11] H. W. Gould and Jocelyn Quaintance. Inverting a finite series with constant coefficients. Fibonacci Quarterly, 49(2):158–165, May 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/49-2/gould.pdf.

Godsil:1983:PFT

[GR83] Christopher D. Godsil and Reinhard Razen. A property of Fibonacci and Tribonacci numbers. Fibonacci Quarterly, 21(1): 13–17, February 1983. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/21-1/godsil.pdf.

Goldstein:1993:FVK

[GR93] Arthur S. Goldstein and Edward M. Reingold. A Fibonacci version of Kraft’s inequality applied to discrete unimodal search. SIAM Journal on Computing, 22(4):751–777, August 1993. CO-DEN SMJCAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Garnier:2008:FNT

[GR09] N. Garnier and O. Ramaré. Fibonacci numbers and trigonometric identities. Fibonacci Quarterly, 46/47(1):56–61, February 2008/ 2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/46_47-1/ramare.pdf.

Graham:1964:PFN

[Gra64] R. L. Graham. A property of Fibonacci numbers. Fibonacci Quarterly, 2(1):1–10, February 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-1/
graham.pdf.

Grassl:1982:SGS

[Gra82] Richard Grassl. Self-generating systems. Fibonacci Quarterly, 20 (4):299–310, November 1982. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/20-4/grassl.
pdf.

Graham:1989:UPN

[Gra89a] S. W. Graham. Unitary perfect numbers with squarefree odd part. Fibonacci Quarterly, 27(4):317–322, August 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-4/graham.pdf.

Granville:1989:CD

[Gra89b] Andrew Granville. On a class of determinants. Fibonacci Quarterly, 27(3):253–256, June 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-
3/granville.pdf.

Gray:1989:BRB

[Gra89c] C. T. Gray. Book review: Applications of Fibonacci Numbers, by A. N. Philippou; A. F. Horadam; G. E. Bergum. Journal of the Royal Statistical Society. Series D (The Statistician), 38(4):308, ???? 1989. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.org/stable/2349066.

Greenwood:1964:LPF

[Gre64] Robert E. Greenwood. Lattice paths and Fibonacci numbers. Fibonacci Quarterly, 2(1):13–14, February 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-1/
greenwood.pdf.

Green:1968:LDE

[Gre68] Thomas M. Green. Linear Diophantine equations with nonnegative parameters and solutions. Fibonacci Quarterly, 6(2): 177–184, April 1968. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/6-2/green.pdf.

Greig:1976:BRF

[Gre76] W. Elliott Greig. Bode’s rule and folded sequences. Fibonacci Quarterly, 14(2):129–133, April 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-2/
greig.pdf.

Greig:1977:FTN

[Gre77a] W. E. Greig. On Fibonacci and triangular numbers. Fibonacci Quarterly, 15(2):176–177, April 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-2/
greig.pdf.

Greig:1977:SFT

[Gre77b] W. E. Greig. On sums of Fibonacci-type reciprocals. Fibonacci Quarterly, 15(4):356–358, December 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
4/greig.pdf.

Greig:1977:RPL

[Gre77c] W. E. Greig. The reciprocal period law. Fibonacci Quarterly, 15 (1):17–20, February 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-1/greig1.pdf.

Greig:1977:SFR

[Gre77d] W. E. Greig. Sums of Fibonacci reciprocals. Fibonacci Quarterly, 15(1):46–48, February 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-1/greig2.pdf.

Greig:1978:FSB

[Gre78a] W. E. Greig. Folded sequences and Bode’s problem. Fibonacci Quarterly, 16(6):530–538, December 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
6/greig.pdf.

Greig:1978:GN

[Gre78b] W. E. Greig. On generalized G_{j,k} numbers. Fibonacci Quarterly, 16(2):166–170, April 1978. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/16-2/greig.pdf.

Greene:2002:BC

[Gre02] John Greene. The Burgstahler coincidence. Fibonacci Quarterly, 40(3):194–202, June 2002. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/40-3/greene.pdf.

Greene:2008:UFD

[Gre09] John Greene. The unboundedness of a family of difference equations over the integers. Fibonacci Quarterly, 46/47(2):146–152, May 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/46_47-2/greene.pdf.

Greene:2018:LST

[Gre18] John Greene. Lucas sequences and traces of matrix products. Fibonacci Quarterly, 56(3):200–??, August 2018. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/56-3/greene.pdf.

Gridgeman:1973:NLF

[Gri73a] N. T. Gridgeman. A new look at Fibonacci generalization. Fibonacci Quarterly, 11(1):40–54, February 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-1/gridgeman.pdf.

Grimm:1973:ALP

[Gri73b] Richard E. Grimm. The autobiography of Leonardo Pisano. Fibonacci Quarterly, 11(1):99–104, February 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-1/grimm.pdf.

Grimson:1974:ECA

[Gri74] R. C. Grimson. The evaluation of certain arithmetic sums. Fibonacci Quarterly, 12(4):373–380, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/12-4/grimson.pdf.

Griffin:1975:SEO

[Gri75] William Raymond Griffin. Significance of even-oddness of a Prime’s penultimate digit. Fibonacci Quarterly, 13(3):204–??, October 1975. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/13-3/griffin-a.pdf.

Griffin:1992:ASF

[Gri92] Peter Griffin. Acceleration of the sum of Fibonacci reciprocals. Fibonacci Quarterly, 30(2):179–181, May 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-2/griffin.pdf.

Griffiths:2008:SRE

[Gri09] Martin Griffiths. Symmetric rational expressions in the Fibonacci numbers. Fibonacci Quarterly, 46/47(3):262–267, August 2008/ 2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/46_47-3/griffiths.pdf.

Griffiths:2010:DPZ

[Gri10] Martin Griffiths. Digit proportions in Zeckendorf representations. Fibonacci Quarterly, 48(2):168–174, May 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/48-2/griffiths.pdf.

Griffiths:2011:BLF

[Gri11a] Martin Griffiths. Binet-like formulas from a simple expansion. Fibonacci Quarterly, 49(4):355–??, November 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/49-4/griffiths.pdf.

Griffiths:2011:FFL

[Gri11b] Martin Griffiths. Families of Fibonacci and Lucas sums via the moments of a random variable. Fibonacci Quarterly, 49(1):76– 81, February 2011. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/49-1/griffiths2.pdf.

Griffiths:2011:FD

[Gri11c] Martin Griffiths. Fibonacci diagonals. Fibonacci Quarterly, 49(1):51–56, February 2011. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/49-1/
griffiths1.pdf.

Griffiths:2011:FEA

[Gri11d] Martin Griffiths. Fibonacci expressions arising from a cointossing scenario involving pairs of consecutive heads. Fibonacci Quarterly, 49(3):249–254, August 2011. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/49-
3/griffiths.pdf.

Griffiths:2012:EDD

[Gri12] Martin Griffiths. Extending the domains of definition of some Fibonacci identities. Fibonacci Quarterly, 50(4):352–??, November 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/50-4/griffiths.pdf.

Griffiths:2013:GRE

[Gri13] Martin Griffiths. From golden-ratio equalities to Fibonacci and Lucas identities. The Mathematical Gazette, 97(539):234–241, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2014:FTZ

[Gri14a] Martin Griffiths. Fixed-term Zeckendorf representations. Fibonacci Quarterly, 52(4):331–??, November 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/52-4/griffiths.pdf.

Griffiths:2014:SIG

[Gri14b] Martin Griffiths. Some identities via geometric series. Fibonacci Quarterly, 52(3):218–??, August 2014. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/52-3/
griffiths.pdf.

Griffiths:2015:RWA

[Gri15a] Martin Griffiths. Random walks arising from a Fibonacci’srabbits scenario. The Mathematical Gazette, 99(544):60–67, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2015:ZRB

[Gri15b] Martin Griffiths. The Zeckendorf representation of a Beattyrelated Fibonacci sum. Fibonacci Quarterly, 53(3):230–??, August 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/53-3/griffiths.pdf.

Griffiths:2017:TFI

[Gri17a] Martin Griffiths. 101.44 on a trivariate Fibonacci identity. The Mathematical Gazette, 101(552):519–522, November 2017.

CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10144-on-a-trivariate-fibonacci-
identity/3B7FB8358D69D30E49364C7EF70D87A5.

Grimaldi:2017:ESG

[Gri17b] Ralph P. Grimaldi. Extraordinary subsets: A generalization. Fibonacci Quarterly, 55(2):114–??, May 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/55-
2/grimaldi.pdf.

Griffiths:2018:FIF

[Gri18] Martin Griffiths. A formula for an infinite family of Fibonacciword sequences. Fibonacci Quarterly, 56(1):75–??, February 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/56-1/griffiths.pdf.

Griffiths:2021:SCF

[Gri21] Martin Griffiths. 105.01 On some composite Fibonacci expressions. The Mathematical Gazette, 105(562):106–108, March 2021. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10501-on-some-composite-
fibonacci-expressions/962C6CBAE91D197480CD67D200E9942F.

Grossman:1997:FCO

[Gro97] George W. Grossman. Fractal construction by orthogonal projection using the Fibonacci sequences. Fibonacci Quarterly, 35 (3):206–224, August 1997. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/35-3/grossman.pdf.

Grundman:1994:SCN

[Gru94] H. G. Grundman. Sequences of consecutive n-Niven numbers. Fibonacci Quarterly, 32(2):174–175, May 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-2/
grundman.pdf.

Grundman:2001:ARN

[Gru01] H. G. Grundman. An analysis of n-Riven numbers. Fibonacci Quarterly, 39(3):253–255, June 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-3/
grundman.pdf.

Grundman:2007:CZN

[Gru07] Helen G. Grundman. Consecutive Zeckendorf–Niven and lazy-Fibonacci–Niven numbers. Fibonacci Quarterly, 45(3):272–276, August 2007. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/45-3/grundman.pdf.

Grundman:2008:ENS

[Gru09] H. G. Grundman. Escalator number sequences. Fibonacci Quarterly, 46/47(2):98–102, May 2008/2009. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/46_47-
2/grundman.pdf.

Grubb:2014:FCP

[Gru14] D. J. Grubb. Factoring Chebyshev polynomials. Fibonacci Quarterly, 52(4):360–??, November 2014. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/52-
4/grubb.pdf.

Grytczuk:2003:SCE

[Gry03] Krystyna Grytczuk. On some classes of effectively integrable differential equations and functional recurrences. Fibonacci Quarterly, 41(3):209–219, June 2003. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/41-3/grytczuk.
pdf.

Guy:1974:CS

[GS74] Richard Guy and Daniel Shanks. A constructed solution of σ(n) = σ(n + 1). Fibonacci Quarterly, 12(3):299–??, October 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/12-3/guy.pdf.

Garfinkel:1981:SCG

[GS81] Robert S. Garfinkel and Stanley M. Selkow. On some conjectures of Gould on the parities of the binomial coefficients. Fibonacci Quarterly, 19(1):61–63, February 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
1/garfinkel.pdf.

Glaser:1995:DSP

[GS95] Herbert Glaser and Gerd Schöffl. Ducci-sequences and Pascal’s triangle. Fibonacci Quarterly, 33(4):313–324, August 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-4/glaser.pdf.

Grassl:2017:FLS

[GS17] Markus Grassl and Andrew J. Scott. Fibonacci–Lucas SIC– POVMs. Journal of Mathematical Physics, 58(12):122201, December 2017. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Glen:2019:MPF

[GSS19] Amy Glen, Jamie Simpson, and W. F. Smyth. More properties of the Fibonacci word on an infinite alphabet. Theoretical Computer Science, 795(??):301–311, November 26, 2019. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397519304372.

Gandhi:1972:C

[GT72] J. M. Gandhi and V. S. Taneja. The coefficients of cosh x/ cos x. Fibonacci Quarterly, 10(4):349–354, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-4/gandhi.pdf.

Grundman:2001:GHN

[GT01] H. G. Grundman and E. A. Teeple. Generalized happy numbers. Fibonacci Quarterly, 39(5):462–466, November 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-5/grundman.pdf.

Grundman:2003:HHN

[GT03] H. G. Grundman and E. A. Teeple. Heights of happy numbers and cubic happy numbers. Fibonacci Quarterly, 41(4):301–306, August 2003. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/41-4/grundman.pdf.

Grabner:1996:LSD

[GTNP96] P. J. Grabner, R. F. Tichy, I. Nemes, and A. Pethö. On the least significant digit of Zeckendorf expansions. Fibonacci Quarterly, 34(2):147–151, May 1996. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/34-2/grabner.pdf.

Gulec:2013:NAG

[GTU13] H. H. Gulec, N. Taskara, and K. Uslu. A new approach to generalized Fibonacci and Lucas numbers with binomial coefficients. Applied Mathematics and Computation, 220(??):482–486, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300313005900.

Gillespie:1966:GLP

[GU66] F. S. Gillespie and W. R. Utz. A generalized Langford problem. Fibonacci Quarterly, 4(2):184–186, April 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/4-2/gillespie.pdf.

Guaraldo:1978:DISa

[Gua78a] Rosalind Guaraldo. On the density of the image sets of certain arithmetic functions — I. Fibonacci Quarterly, 16(4):318–326, August 1978. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/16-4/guaraldo.pdf.

Guaraldo:1978:DISb

[Gua78b] Rosalind Guaraldo. On the density of the image sets of certain arithmetic functions — II. Fibonacci Quarterly, 16(5):428–434, October 1978. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/16-5/guaraldo.pdf.

Guaraldo:1978:DISc

[Gua78c] Rosalind Guaraldo. On the density of the image sets of certain arithmetic functions — III. Fibonacci Quarterly, 16(6):481–487, December 1978. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/16-6/guaraldo.pdf.

Guerin:1978:MCA

[Gue78] E. E. Guerin. Matrices and convolutions of arithmetic functions. Fibonacci Quarterly, 16(4):327–334, August 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-4/guerin.pdf.

Guerin:1990:PC

[Gue90] E. E. Guerin. Partitions with “M(a) Copies of a”. Fibonacci Quarterly, 28(4):298–301, November 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
4/guerin.pdf.

Guillot:1977:P

[Gui77a] Guy A. R. Guillot. Problems. Fibonacci Quarterly, 15(3):232– ??, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/15-3/guillot1-a.pdf.

Guillot:1977:UNT

[Gui77b] Guy A. R. Guillot. The unified number theory. Fibonacci Quarterly, 15(3):254–??, October 1977. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/15-3/guillot2.
pdf.

Guillera:2020:WRM

[Gui20] Jesus Guillera. When Ramanujan meets Fibonacci and Lucas. American Mathematical Monthly, 127(2):159, 2020. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Gupta:1974:GHH

[Gup74] A. K. Gupta. Generalized hidden hexagon squares. Fibonacci Quarterly, 12(1):45–??, February 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
1/gupta.pdf.

Gupta:1976:DMO

[Gup76] Hansraj Gupta. A direct method of obtaining Farey–Fibonacci sequences. Fibonacci Quarterly, 14(5):389–390, December 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-5/gupta.pdf.

Gupta:1978:AFF

[Gup78a] Hansraj Gupta. The Andrews formula for Fibonacci numbers. Fibonacci Quarterly, 16(6):552–554, December 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/16-6/gupta2.pdf.

Gupta:1978:RVP

[Gup78b] Hansraj Gupta. The rank-vector of a partition. Fibonacci Quarterly, 16(6):548–551, December 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
6/gupta1.pdf.

Guy:1982:SCI

[Guy82] Robert Guy. Sums of consecutive integers. Fibonacci Quarterly, 20(1):36–37, February 1982. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/20-1/guy.pdf.

Gil:2019:GMM

[GW19] Juan B. Gil and Aaron Worley. Generalized metallic means. Fibonacci Quarterly, 57(1):45–??, February 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/57-1/gil.pdf.

Griffiths:2015:PFS

[GWT15] Martin Griffiths and William Wynn-Thomas. A property of a Fibonacci staircase. Fibonacci Quarterly, 53(1):61–??, February 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/53-1/griffiths.pdf.

Gil:2007:CPS

[GWZ07] Juan B. Gil, Michael D. Weiner, and Catalin Zara. Complete Padovan sequences in finite fields. Fibonacci Quarterly, 45(1): 64–75, February 2007. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/45-1/gil.pdf.

Glasser:2015:IRF

[GZ15] M. Lawrence Glasser and Yajun Zhou. An integral representation for the Fibonacci numbers and their generalization. Fibonacci Quarterly, 53(4):313–??, November 2015. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/53-
4/glasser.pdf.

Goy:2019:ONT

[GZ19] Taras Goy and Roman Zatorsky. On Oresme numbers and their connection with Fibonacci and Pell numbers. Fibonacci Quarterly, 57(3):238–??, August 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/57-3/
goy.pdf.

Guzicki:2022:BPF

[GZ22] Wojciech Guzicki and Piotr Zakrzewski. Bijective proofs of formulas with (−1)^{n}. Fibonacci Quarterly, 60(3):262–??, August 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/60-3/guzicki.pdf.

Hillman:1964:MCF

[HA64] A. P. Hillman and G. L. Alexanderson. A motivation for continued fractions. Fibonacci Quarterly, 2(2):145–148, April 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-2/hillman.pdf.

Hillman:1966:ATP

[HA66] Abraham P. Hillman and Gerald L. Alexanderson. Algebra through problem solving. The Allyn and Bacon series: topics in contemporary mathematics. Allyn and Bacon, Needham Heights, MA, USA, 1966. vii + 129 pp. LCCN QA157 .H6.

Hoggatt:1971:PMC

[HA71] V. E. Hoggatt, Jr. and G. L. Alexanderson. A property of multinomial coefficients. Fibonacci Quarterly, 9(4):351–356, October 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/9-4/hoggatt1-a.pdf.

Hoggatt:1973:PFNa

[HA73] Verner E. Hoggatt, Jr. and Janet Crump Anaya. A primer for the Fibonacci numbers. Part XI: Multisection generating functions for the columns of Pascal’s triangle. Fibonacci Quarterly, 11(1): 85–90, February 1973. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/11-1/hoggatt-a.pdf.

Hoggatt:1975:GFT

[HA75] V. E. Hoggatt, Jr. and Krishnaswami Alladi. Generalized Fibonacci tiling. Fibonacci Quarterly, 13(2):137–144, April 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-2/hoggatt1.pdf.

Hoggatt:1976:SPS

[HA76a] V. E. Hoggatt, Jr. and G. L. Alexanderson. Sums of partition sets in generalized Pascal triangles I. Fibonacci Quarterly, 14(2): 117–125, April 1976. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/14-2/hoggatt1.pdf.

Hoggatt:1976:WS

[HA76b] V. E. Hoggatt, Jr. and Krishnaswami Alladi. In-winding spirals. Fibonacci Quarterly, 14(2):144–146, April 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-2/hoggatt3.pdf.

Hoggatt:1977:CRR

[HA77a] V. E. Hoggatt, Jr. and Krishnaswami Alladi. Compositions and recurrence relations II. Fibonacci Quarterly, 15(3):239–244, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/15-3/hoggatt3.pdf.

Hoggatt:1977:LRC

[HA77b] V. E. Hoggatt, Jr. and Krishnaswami Alladi. Limiting ratios of convolved recursive sequences. Fibonacci Quarterly, 15(3):211– 214, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/15-3/hoggatt1.pdf.

Haase:1993:HFN

[Haa93] Herman Haase. Hyperspaces and Fibonacci numbers. Fibonacci Quarterly, 31(2):158–161, May 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-2/
haase.pdf.

Habsieger:2019:EBD

[Hab19] Laurent Habsieger. Explicit bounds for the Diophantine equation A!B! = C!. Fibonacci Quarterly, 57(1):21–??, February 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-1/habsieger.pdf.

Hagis:1964:APF

[Hag64] Peter Hagis, Jr. An analytic proof of the formula for F_{n}. Fibonacci Quarterly, 2(4):267–268, December 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/2-4/hagis.pdf.

Hagis:1984:LBU

[Hag84] Peter Hagis, Jr. Lower bounds for unitary multiperfect numbers. Fibonacci Quarterly, 22(2):140–143, May 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/22-2/hagis.pdf.

Hagis:1987:BUA

[Hag87a] Peter Hagis, Jr. Bi-unitary amicable and multiperfect numbers. Fibonacci Quarterly, 25(2):144–150, May 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-2/hagis.pdf.

Hagis:1987:SSU

[Hag87b] Peter Hagis, Jr. A systematic search for unitary hyperperfect numbers. Fibonacci Quarterly, 25(1):6–10, February 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-1/hagis.pdf.

Hagis:1990:ONP

[Hag90] Peter Hagis, Jr. Odd nonunitary perfect numbers. Fibonacci Quarterly, 28(1):11–14, February 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
1/hagis.pdf.

Hahn:1972:CFI

[Hah72] Hwa S. Hahn. A counting function of integral n-triples. Fibonacci Quarterly, 10(6):609–612, December 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
6/hahn-a.pdf.

Hahn:1975:APM

[Hah75] H. S. Hahn. Another property of magic squares. Fibonacci Quarterly, 13(3):205–208, October 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-
3/hahn.pdf.

Halton:1964:FR

[Hal64] J. H. Halton. On Fibonacci residues. Fibonacci Quarterly, 2(3): 217–??, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/2-3/halton.pdf.

Halsey:1965:FNW

[Hal65a] Eric Halsey. The Fibonacci number F_{u} where u is not an integer. Fibonacci Quarterly, 3(2):147–152, April 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/3-2/halsey.pdf.

Halsey:1965:LE

[Hal65b] Eric Halsey. Letter to the editor. Fibonacci Quarterly, 3(3):233– ??, October 1965. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/3-3/halsey.pdf.

Halton:1965:NFS

[Hal65c] John H. Halton. A note on Fibonacci subsequences. Fibonacci Quarterly, 3(4):321–322, December 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-4/
halton.pdf.

Halton:1965:GFI

[Hal65d] John H. Halton. On a general Fibonacci identity. Fibonacci Quarterly, 3(1):31–43, February 1965. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/3-1/halton.
pdf.

Halton:1966:DPF

[Hal66] John H. Halton. On the divisibility properties of Fibonacci numbers. Fibonacci Quarterly, 4(3):217–240, October 1966. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-3/halton.pdf.

Halton:1967:SPA

[Hal67] John H. Halton. Some properties associated with square Fibonacci numbers. Fibonacci Quarterly, 5(4):347–354, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/5-4/halton.pdf.

Halberg:1968:GFO

[Hal68] Charles J. A. Halberg, Jr. The generalized Fibonacci operator. Fibonacci Quarterly, 6(5):15–33, November 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-5/halberg.pdf.

Hale:1983:VNF

[Hal83] David R. Hale. A variant of Nim and a function defined by Fibonacci representation. Fibonacci Quarterly, 21(2):139–142, May 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/21-2/hale.pdf.

Hall:2022:BRF

[Hal22] Peter Hall. Book review: Fibonacci’s rabbits by Adam Hart-Davis, pp. 176, £12.99 (paper), ISBN 978-1-912827-03-9, Modern Books (2019). The Mathematical Gazette, 106(565):191, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/fibonaccis-rabbits-by-adam-hartdavis-pp-176-1299-
paper-isbn-9781-912827039-modern-books-2019/D1C56DFBA72389E0E70A458F2FCA27F5.

Hansen:1970:APN

[Han70] Rodney T. Hansen. Arithmetic of pentagonal numbers. Fibonacci Quarterly, 8(1):83–86, February 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-1/
hansen.pdf.

Hansen:1972:GIF

[Han72] Rodney T. Hansen. Generating identities for Fibonacci and Lucas triples. Fibonacci Quarterly, 10(6):571–578, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-6/hansen.pdf.

Hanani:1976:CI

[Han76] Haim Hanani. A combinatorial identity. Fibonacci Quarterly, 14 (1):49–51, February 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-1/hanani.pdf.

Hansen:1978:GIL

[Han78] Rodney T. Hansen. General identities for linear Fibonacci and Lucas summations. Fibonacci Quarterly, 16(2):121–127, April 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/16-2/hansen.pdf.

Hannah:2007:FPL

[Han07] John Hannah. False position in Leonardo of Pisa’s Liber abbaci. Historia Mathematica, 34(3):306–332, August 2007. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0315086006001297.

Hannah:2011:CRP

[Han11] John Hannah. Conventions for recreational problems in Fibonacci’s Liber Abbaci. Archive for History of Exact Sciences, 65(2):155–180, March 2011. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=????&
volume=65&issue=2&spage=155.

Hanlon:2012:TFT

[Han12] Michael Hanlon. Turing’s flower theory blossoms. The Times [London], October 28, 2012. URL http://www.

thesundaytimes.co.uk/sto/news/uk_news/Environment/article1156602.
ece.

Harris:1965:IIF

[Har65] V. C. Harris. On identities involving Fibonacci numbers. Fibonacci Quarterly, 3(3):214–218, October 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/3-3/harris.pdf.

Harris:1970:AFG

[Har70a] V. C. Harris. An algorithm for finding the greatest common divisor. Fibonacci Quarterly, 8(1):102–103, February 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-1/harris1.pdf.

Harris:1970:NND

[Har70b] V. C. Harris. Note on the number of divisions required in finding the greatest common divisor. Fibonacci Quarterly, 8(1):104–??, February 1970. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/8-1/harris2.pdf.

Harris:1974:DAF

[Har74] V. C. Harris. On Daykin’s algorithm for finding the G. C. D. Fibonacci Quarterly, 12(1):80–??, February 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/12-1/harris.pdf.

Harborth:1978:ANP

[Har78] Heiko Harborth. b-adic numbers in Pascal’s triangle modulo b. Fibonacci Quarterly, 16(6):497–500, December 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-6/harborth.pdf.

Harman:1981:CFN

[Har81] C. J. Harman. Complex Fibonacci numbers. Fibonacci Quarterly, 19(1):82–86, February 1981. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/19-1/harman.pdf.

Hare:1994:FNF

[Har94] E. O. Hare. Fibonacci numbers and fractional domination of P_{m} × P_{n}. Fibonacci Quarterly, 32(1):69–73, February 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-1/hare.pdf.

Hart:1998:UPB

[Har98] Evelyn Hart. On using patterns in beta-expansions to study Fibonacci–Lucas products. Fibonacci Quarterly, 36(5):396–405, November 1998. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/36-5/hart.pdf.

Hargreaves:2002:IAM

[Har02] Gareth I. Hargreaves. Interval analysis in MATLAB. Numerical analysis report 416, Manchester Centre for Computational Mathematics, Manchester, England, December 2002. 49 pp. URL

http://www.maths.man.ac.uk/~nareports/narep416.pdf.

Hartshorn:2009:BRB

[Har09] Kevin Hartshorn. Book review: The fabulous Fibonacci numbers by Alfred Posamentier and Ingmar Lehmann. Journal of Mathematics and the Arts, 3(2):113–116, ???? 2009. CODEN ???? ISSN 1751-3472 (print), 1751-3480 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/17513470902897569.

Hassin:1981:MFF

[Has81] Refael Hassin. On maximizing functions by Fibonacci search. Fibonacci Quarterly, 19(4):347–350, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-4/hassin.pdf.

Haukkanen:1988:NSM

[Hau88] Pentti Haukkanen. A note on specially multiplicative arithmetic functions. Fibonacci Quarterly, 26(4):325–327, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/26-4/haukkanen.pdf.

Haukkanen:1990:VIT

[Hau90] Pentti Haukkanen. Vinogradov’s inversion theorem for generalized arithmetical functions. Fibonacci Quarterly, 28(4):316–320, November 1990. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/28-4/haukkanen.pdf.

Haukkanen:1993:FPS

[Hau93a] Pentti Haukkanen. Formal power series for binomial sums of sequences of numbers. Fibonacci Quarterly, 31(1):28–31, February 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/31-1/haukkanen.pdf.

Haukkanen:1993:NRA

[Hau93b] Pentti Haukkanen. A note on rational arithmetic functions of order (2, 1). Fibonacci Quarterly, 31(4):302–306, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-4/haukkanen.pdf.

Haukkanen:1994:RSU

[Hau94a] Pentti Haukkanen. Roots of sequences under covolutions. Fibonacci Quarterly, 32(4):369–372, August 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-4/haukkanen.pdf.

Hauss:1994:FLC

[Hau94b] Michael Hauss. Fibonacci, Lucas, and central factorial numbers, and π. Fibonacci Quarterly, 32(5):395–396, November 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-5/hauss.pdf.

Haukkanen:1996:BSF

[Hau96] Pentti Haukkanen. On a binomial sum for the Fibonacci and Lucas pseudoprimes. Fibonacci Quarterly, 34(4):326–331, August 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/34-4/haukkanen.pdf.

Haukkanen:1997:NBF

[Hau97a] Pentti Haukkanen. A note on the bracket function transform. Fibonacci Quarterly, 35(2):156–159, May 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-2/
haukkanen.pdf.

Haukkanen:1997:RRT

[Hau97b] Pentti Haukkanen. On a recurrence relation in two variables. Fibonacci Quarterly, 35(1):32–34, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-1/haukkanen.pdf.

Haukkanen:2000:ACA

[Hau00] Pentti Haukkanen. On the k-ary convolution of arithmetical functions. Fibonacci Quarterly, 38(5):440–445, November 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-5/haukkanen.pdf.

Haukkanen:2002:NHS

[Hau02] Pentti Haukkanen. A note on Horadam’s sequence. Fibonacci Quarterly, 40(4):358–361, August 2002. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/40-
4/haukkanen.pdf.

Hayashi:1992:FNR

[Hay92] Elmer K. Hayashi. Fibonacci numbers, recursion, complexity, and induction proofs. College Mathematics Journal, 23(5):407– 410, November 1992. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1992.11973491.

Hoggatt:1963:RCS

[HB63] V. E. Hoggatt, Jr. and S. L. Basin. Representations by complete sequences — Part I (Fibonacci). Fibonacci Quarterly, 1(3):1–14, October 1963. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/1-3/hoggatt1-a.pdf.

Hoggatt:1964:FPF

[HB64a] Verner E. Hoggatt, Jr. and Marjorie Bicknell. Fourth power Fibonacci identities from Pascal’s triangle. Fibonacci Quarterly, 2(4):261–266, December 1964. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/2-4/hoggatt.
pdf.

Hoggatt:1964:SNF

[HB64b] Verner E. Hoggatt, Jr. and Marjorie Bicknell. Some new Fibonacci identities. Fibonacci Quarterly, 2(1):29–31, February 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/2-1/hoggatt1.pdf.

Hoggatt:1964:PFS

[HB64c] Verner E. Hoggatt, Jr. and I. D. Bruggles. A primer on the Fibonacci sequence, Part V. Fibonacci Quarterly, 2(1):59–66, February 1964. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/2-1/hoggatt2.pdf.

Hoggatt:1969:DSG

[HB69] V. E. Hoggatt, Jr. and Marjorie Bicknell. Diagonal sums of generalized Pascal triangles. Fibonacci Quarterly, 7(4):341–358, November 1969. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/7-4/hoggatt-a.pdf.

Hoggatt:1972:FFC

[HB72a] V. E. Hoggatt, Jr. and Marjorie Bicknell. Fun with Fibonacci at the chess match. Fibonacci Quarterly, 10(4):433–434, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/10-4/hoggatt-a.pdf.

Hoggatt:1972:CT

[HB72b] Verner E. Hoggatt, Jr. and Marjorie Bicknell. Convolution triangles. Fibonacci Quarterly, 10(6):599–608, December 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-6/hoggatt.pdf.

Hoggatt:1973:RFP

[HB73a] V. E. Hoggatt, Jr. and Marjorie Bicknell. Roots of Fibonacci polynomials. Fibonacci Quarterly, 11(3):271–274, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-3/hoggatt1.pdf.

Hoggatt:1973:SDF

[HB73b] V. E. Hoggatt, Jr. and Marjorie Bicknell. Special determinants found within generalized Pascal triangles. Fibonacci Quarterly, 11(5):469–479, December 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
5/bicknell1.pdf.

Hoggatt:1973:GFP

[HB73c] Verner E. Hoggatt, Jr. and Marjorie Bicknell. Generalized Fibonacci polynomials and Zeckendorf’s theorem. Fibonacci Quarterly, 11(4):399–419, November 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
4/hoggatt2.pdf.

Hoggatt:1974:DCR

[HB74a] V. E. Hoggatt, Jr. and Gerald E. Bergum. Divisibility and congruence relations. Fibonacci Quarterly, 12(2):189–195, April 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/12-2/hoggatt3.pdf.

Hoggatt:1974:DST

[HB74b] V. E. Hoggatt, Jr. and Marjorie Bicknell. Diagonal sums of the trinomial triangle. Fibonacci Quarterly, 12(1):47–50, February 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/12-1/hoggatt.pdf.

Hoggatt:1974:PFN

[HB74c] V. E. Hoggatt, Jr. and Marjorie Bicknell. A primer for the Fibonacci numbers: Part XIV. Fibonacci Quarterly, 12(2):147–156, April 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/12-2/hoggatt2.pdf.

Hoggatt:1974:TN

[HB74d] V. E. Hoggatt, Jr. and Marjorie Bicknell. Triangular numbers. Fibonacci Quarterly, 12(3):221–230, October 1974. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-3/hoggatt.pdf.

Hoggatt:1975:GCA

[HB75a] V. E. Hoggatt, Jr. and G. E. Bergum. Generalized convolution arrays. Fibonacci Quarterly, 13(3):193–197, October 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-3/hoggatt1.pdf.

Hoggatt:1975:PC

[HB75b] V. E. Hoggatt, Jr. and Marjorie Bicknell. Palindromic compositions. Fibonacci Quarterly, 13(4):350–355, December 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-4/hoggatt1.pdf.

Hoggatt:1975:SP

[HB75c] V. E. Hoggatt, Jr. and Marjorie Bicknell. Special partitions. Fibonacci Quarterly, 13(3):278–??, October 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-3/hoggatt2-a.pdf.

Hoggatt:1975:CT

[HB75d] V. E. Hoggatt, Jr. and Paul S. Bruckman. The H-convolution transform. Fibonacci Quarterly, 13(4):357–368, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-4/hoggatt2.pdf.

Hoggatt:1976:CRS

[HB76a] V. E. Hoggatt, Jr. and Marjorie Bicknell. Catalan and related sequences arising from inverses of Pascal’s triangle matrices. Fibonacci Quarterly, 14(5):395–404, December 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-5/hoggatt1.pdf.

Hoggatt:1976:PCG

[HB76b] V. E. Hoggatt, Jr. and Marjorie Bicknell. Pascal, Catalan, and general sequence convolution arrays in a matrix. Fibonacci Quarterly, 14(2):135–142, April 1976. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/14-2/hoggatt2.
pdf.

Hoggatt:1976:PFN

[HB76c] V. E. Hoggatt, Jr. and Marjorie Bicknell. A primer for the Fibonacci numbers, Part XV: Variations on summing a series of reciprocals of Fibonacci numbers. Fibonacci Quarterly, 14(3): 272–276, October 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-3/hoggatt2.pdf.

Hoggatt:1976:RSF

[HB76d] V. E. Hoggatt, Jr. and Marjorie Bicknell. A reciprocal series of Fibonacci numbers with subscripts 2^{n}k. Fibonacci Quarterly, 14(5):453–454, December 1976. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/14-5/hoggatt2.
pdf.

Hoggatt:1976:SMI

[HB76e] V. E. Hoggatt, Jr. and Marjorie Bicknell. Sequences of matrix inverses from Pascal, Catalan, and related convolution arrays. Fibonacci Quarterly, 14(3):224–232, October 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-3/hoggatt1.pdf.

Hoggatt:1977:PFF

[HB77a] V. E. Hoggatt, Jr. and G. E. Bergum. A problem of Fermat and the Fibonacci sequence. Fibonacci Quarterly, 15(4):323–330, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/15-4/hoggatt1.pdf.

Hoggatt:1977:PCF

[HB77b] V. E. Hoggatt, Jr. and Paul S. Bruckman. Periodic continued fraction representations of Fibonacci-type irrationals. Fibonacci Quarterly, 15(3):225–229, October 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
3/hoggatt2.pdf.

Hathaway:1997:FPF

[HB97] Dale K. Hathaway and Stephen L. Brown. Fibonacci powers and a fascinating triangle. College Mathematics Journal, 28(2): 124–128, March 1997. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/07468342.1997.11973848.

Hoggatt:1977:CPA

[HBJ77a] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Composites and primes among powers of Fibonacci numbers, increased or decreased by one. Fibonacci Quarterly, 15(1):2–??, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-1/hoggatt1.pdf.

Hoggatt:1977:DFL

[HBJ77b] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Divisibility by Fibonacci and Lucas squares. Fibonacci Quarterly, 15(1):3– 7, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/15-1/hoggatt2.pdf.

Hoggatt:1977:FCS

[HBJ77c] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Fibonacci convolution sequences. Fibonacci Quarterly, 15(2):117–121, April 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/15-2/hoggatt1.pdf.

Hoggatt:1977:NPC

[HBJ77d] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Numerator polynomial coefficient arrays for Catalan and related sequence convolution triangles. Fibonacci Quarterly, 15(1):30–34, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/15-1/hoggatt3.pdf.

Hoggatt:1977:GLS

[HBJ77e] Verner E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Generalized Lucas sequences. Fibonacci Quarterly, 15(2):131–139, April 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/15-2/hoggatt2.pdf.

Hoggatt:1978:CAJ

[HBJ78a] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Convolution arrays for Jacobsthal and Fibonacci polynomials. Fibonacci Quarterly, 16(5):385–402, October 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
5/hoggatt1.pdf.

Hoggatt:1978:DPP

[HBJ78b] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Divisibility properties of polynomials in Pascal’s triangle. Fibonacci Quarterly, 16(6):501–512, December 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
6/hoggatt.pdf.

Hoggatt:1978:PFN

[HBJ78c] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. A primer for the Fibonacci numbers XVII: Generalized Fibonacci numbers satisfying u_{n+1}u_{n−1}−u_{n}^{2} = ±1. Fibonacci Quarterly, 16(2):128–137, April 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/16-2/hoggatt.pdf.

Hoggatt:1978:PGF

[HBJ78d] Verner E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Properties of generating functions of a convolution array. Fibonacci Quarterly, 16(4):289–295, August 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
4/hoggatt.pdf.

Hoggatt:1979:RAT

[HBJ79a] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Reflections across two and three glass plates. Fibonacci Quarterly, 17(2): 118–141, April 1979. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/17-2/hoggatt.pdf.

Hoggatt:1979:RIT

[HBJ79b] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Representations of integers in terms of greatest integer functions and the golden section ratio. Fibonacci Quarterly, 17(4):306–317, December 1979. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/17-4/hoggatt.pdf.

Hoggatt:1980:MFC

[HBJ80] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Multisection of the Fibonacci convolution array and generalized Lucas squence. Fibonacci Quarterly, 18(1):51–57, February 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-1/hoggatt1.pdf.

Hoggatt:1982:CAG

[HBJ82a] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Composition arrays generated by Fibonacci numbers. Fibonacci Quarterly, 20(2):122–128, May 1982. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/20-2/hoggatt.pdf.

Hoggatt:1982:LOF

[HBJ82b] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Lexicographic ordering and Fibonacci representations. Fibonacci Quarterly, 20 (3):193–218, August 1982. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/20-3/hoggatt.pdf.

Hoggatt:1982:STR

[HBJ82c] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Sequence transforms related to representations using generalized Fibonacci numbers. Fibonacci Quarterly, 20(4):289–298, November 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-4/hoggatt.pdf.

Hoggatt:1984:APP

[HBJ84] V. E. Hoggatt, Jr. and Marjorie Bicknell-Johnson. Additive partition of the positive integers and generalized Fibonacci representations. Fibonacci Quarterly, 22(1):2–21, February 1984. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-1/hoggatt.pdf.

Hoggatt:1979:GWG

[HBJS79] V. E. Hoggatt, Jr., Marjorie Bicknell-Johnson, and Richard Sarsfield. A generalization of Wythoff’s game. Fibonacci Quarterly, 17 (3):198–211, October 1979. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/17-3/hoggatt.pdf.

Hoggatt:1972:FLT

[HBK72] V. E. Hoggatt, Jr., Marjorie Bichnell, and Ellen L. King. Fibonacci and Lucas triangles. Fibonacci Quarterly, 10(5):555–560, November 1972. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/10-5/hoggatt2.pdf.

Hoggatt:1972:STC

[HC72] V. E. Hoggatt, Jr. and Bob Chow. Some theorems on completeness. Fibonacci Quarterly, 10(5):551–554, November 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-5/hoggatt1-a.pdf.

Howard:2011:SIF

[HC11] F. T. Howard and Curtis Cooper. Some identities for r-Fibonacci numbers. Fibonacci Quarterly, 49(3):231–242, August 2011. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/49-3/howard.pdf.

Hoggatt:1973:PFNb

[HCB73] Verner E. Hoggatt, Jr., Nannette Cox, and Marjorie Bicknell. A primer for the Fibonacci numbers: Part XII. Fibonacci Quarterly, 11(3):317–331, October 1973. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/11-3/hoggatt2.
pdf.

Hsu:1997:LRN

[HCD97] Wen-Jing Hsu, Moon Jung Chung, and Amitabha Das. Linear recursive networks and their applications in distributed systems. IEEE Transactions on Parallel and Distributed Systems, 8(7):673–680, July 1997. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://dlib.

computer.org/td/books/td1997/pdf/l0673.pdf;   http://
www.computer.org/tpds/td1997/l0673abs.htm.

Heubach:2003:RLD

[HCG03] S. Heubach, P. Z. Chinn, and R. P. Grimaldi. Rises, levels, drops and “+” signs in compositions: Extensions of a paper by Alladi and Hoggatt. Fibonacci Quarterly, 41(3):229–239, June 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-3/heubach.pdf.

Hart-Davis:2019:FR

[HD19] Adam Hart-Davis. Fibonacci’s Rabbits: and 49 other discoveries that revolutionised mathematics. Modern Books, London, UK, 2019. ISBN 1-912827-03-4. LCCN ????

Hoggatt:1980:APA

[HE80] Verner E. Hoggatt, Jr. and Hugh Edgar. Another proof that ϕ(F_{n}) ≡ 0 mod 4 for all n > 4. Fibonacci Quarterly, 18(1):80– 81, February 1980. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/18-1/hoggatt2.pdf.

Hedian:1976:GSA

[Hed76] Helene Hedian. The golden section and the artist. Fibonacci Quarterly, 14(5):406–418, December 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
5/hedian.pdf.

Heed:1984:WPZ

[Hee84] Joseph J. Heed. Wieferichs and the problems z(p^{2}) = z(p). Fibonacci Quarterly, 22(2):116–118, May 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-2/
heed.pdf.

Heimer:1964:FCP

[Hei64] R. L. Heimer. Further comments on the periodicity of the digits of the Fibonacci sequence. Fibonacci Quarterly, 2(3):211–214, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/2-3/heimer.pdf.

Heimer:1967:GFF

[Hei67] Richard L. Heimer. A general Fibonacci function. Fibonacci Quarterly, 5(5):481–483, December 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-5/
heimer.pdf.

Heichelheim:1979:SPI

[Hei79] Peter Heichelheim. The study of positive integers (a, b) such that ab + 1 is a square. Fibonacci Quarterly, 17(3):269–274, October 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/17-3/heichelheim.pdf.

Henning:1967:PTR

[Hen67] H. B. Henning. Pythagorean triangles and related concepts. Fibonacci Quarterly, 5(2):185–192, April 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-2/
henning.pdf.

Hendy:1978:SDP

[Hen78a] M. D. Hendy. Stolarsky’s distribution of the positive integers. Fibonacci Quarterly, 16(1):70–80, February 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-1/hendy.pdf.

Hensley:1978:FTH

[Hen78b] Douglas Hensley. Fibonacci tiling and hyperbolas. Fibonacci Quarterly, 16(1):37–40, February 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
1/hensley.pdf.

Hensley:1982:ENU

[Hen82] Douglas Hensley. Eulerian numbers and the unit cube. Fibonacci Quarterly, 20(4):344–348, November 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
4/hensley.pdf.

Hensley:1988:WST

[Hen88] Douglas Hensley. A winning strategy at taxman. Fibonacci Quarterly, 26(3):262–270, August 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
3/hensley.pdf.

Hendel:1994:AFT

[Hen94] Russell Jay Hendel. Approaches to the formula for the n th Fibonacci number. College Mathematics Journal, 25(2):139–142, March 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1994.11973598.

Henze:1995:DSL

[Hen95] Norbert Henze. The distribution of spaces on lottery tickets. Fibonacci Quarterly, 33(5):426–431, November 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-5/henze.pdf.

Hendel:2006:LEF

[Hen06] Russell Jay Hendel. Linear equalities in Fibonacci numbers. Fibonacci Quarterly, 44(3):235–248, August 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/44-3/hendel.pdf.

Hendel:2007:F

[Hen07] Russell Jay Hendel. Factorizations of ∑_{j=i}^{n+i−1} F_{aj−b}. Fibonacci Quarterly, 45(2):128–132, May 2007. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/45-2/
hendel.pdf.

Hendel:2011:ARR

[Hen11a] Russell Jay Hendel. Almost-recursiveness of reciprocals of linearly recurrent sequences. Fibonacci Quarterly, 49(1):41–50, February 2011. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/49-1/hendel.pdf.

Hendel:2011:KF

[Hen11b] Russell Jay Hendel. Kimberling’s ⌊n^{2}α⌋ − n⌊nα⌋ function. Fibonacci Quarterly, 49(3):211–219, August 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/49-3/hendel.pdf.

Hendel:2014:CHC

[Hen14] Russell Jay Hendel. A Cayley–Hamilton and circulant approach to jump sums. Fibonacci Quarterly, 52(5):124–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers1/52-5/Hendel.pdf.

Hendel:2015:CCR

[Hen15a] Russell Jay Hendel. Coefficient convergence of recursively defined polynomials. Fibonacci Quarterly, 53(3):247–??, August 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/53-3/hendel.pdf.

Hendel:2015:QPH

[Hen15b] Russell Jay Hendel. Quasi-periods for the Hofstadter Q function. Fibonacci Quarterly, 53(2):112–??, May 2015. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/53-
2/hendel.pdf.

Hendel:2017:PGC

[Hen17] Russell Jay Hendel. Proof and generalization of the Cassini– Catalan–Tagiuri–Gould identities. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/55-5/hendel.pdf.

Hendel:2018:TFT

[Hen18] Russell Jay Hendel. Towards formulating a Tagiuri generating method conjecture. Fibonacci Quarterly, 56(2):142–??, May 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/56-2/hendel.pdf.

Hendel:2019:PTH

[Hen19] Russell Jay Hendel. Proof of the Tagiuri Histogram Conjecture. Fibonacci Quarterly, 57(5):54–??, December 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-5/hendel.pdf.

Hendel:2020:RTA

[Hen20] Russell Jay Hendel. Recursive triangles appearing embedded in recursive families. Fibonacci Quarterly, 58(5):135–??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/58-5/hendel.pdf.

Hendel:2022:MUP

[Hen22a] Russell Jay Hendel. A method for uniformly proving a family of identities. Fibonacci Quarterly, 60(2):151–??, May 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-2/hendel.pdf.

Hendel:2022:SFSa

[Hen22b] Russell Jay Hendel. A system of four simultaneous recursions: Generalization of the Ledin–Shannon–Ollerton identity. Fibonacci Quarterly, 60(5):172–??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/60-5/hendel.pdf.

Herrick:1973:PTD

[Her73] Daniel Lance Herrick. On the periodicity of the terminal digits in Fibonacci sequences. Fibonacci Quarterly, 11(5):535–538, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/11-5/herrick.pdf.

Herget:1978:MPM

[Her78] W. Herget. Minimum periods modulo n for Bernoulli numbers. Fibonacci Quarterly, 16(6):544–547, December 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/16-6/herget.pdf.

Herda:1981:TPI

[Her81] Hans Herda. Tiling the plane with incongruent regular polygons. Fibonacci Quarterly, 19(5):437–439, December 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-5/herda.pdf.

Herget:1982:MPM

[Her82] Wilfried Herget. Minimum periods modulo n for Bernoulli polynomials. Fibonacci Quarterly, 20(2):106–110, May 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-2/herget.pdf.

Herda:1985:SPS

[Her85] Hans Herda. The series of prime square reciprocals. Fibonacci Quarterly, 23(4):364–365, November 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-
4/herda.pdf.

Herrmann:2003:IFS

[Her03] Ernst Herrmann. Interval-filling sequences involving reciprocal Fibonacci numbers. Fibonacci Quarterly, 41(5):441–449, November 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/41-5/herrmann.pdf.

Herrmann:2014:GS

[Her14] Burghard Herrmann. The Golden Sequence. Fibonacci Quarterly, 52(1):66–??, February 2014. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/52-1/
herrmann.pdf.

Hernandez:2018:FNF

[Her18] Santos Hernández Hernández. The Fibonacci numbers of the form 2^{a} ± 2^{b} + 1. Fibonacci Quarterly, 56(4):354–??, November 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/56-4/hernandez.pdf.

Herrmann:2019:HIS

[Her19] Burghard Herrmann. How integer sequences find their way into areas outside pure mathematics. Fibonacci Quarterly, 57(5):67– ??, December 2019. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/57-5/herrmann.pdf.

Herrmann:2020:CFP

[Her20] Burghard Herrmann. The continued fraction pendulum. Fibonacci Quarterly, 58(5):144–??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/58-5/herrmann.pdf.

Hewgill:1977:RBP

[Hew77] Denton Hewgill. A relationship between Pascal’s triangle and Fermat’s numbers. Fibonacci Quarterly, 15(2):183–184, April 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-2/hewgill.pdf.

Heyde:1980:PAF

[Hey80] C. C. Heyde. On a probabilistic analogue of the Fibonacci sequence. Journal of Applied Probability, 17(4):1079–1082, December 1980. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072

(electronic). URL http://www.jstor.org/stable/3213217.

Heyde:1981:FLB

[Hey81] C. C. Heyde. On Fibonacci (or lagged Bienaymé–Galton– Watson) branching processes. Journal of Applied Probability, 18 (3):583–591, September 1981. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.

org/stable/3213313.

Hoggatt:1973:AVS

[HF73] Verner E. Hoggatt, Jr. and Danial C. Fielder. Analytical verification of an “at sight” transformation. Fibonacci Quarterly, 11(4): 395–398, November 1973. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/11-4/hoggatt1.pdf.

Herz-Fischler:1986:LE

[HF86] R. Herz-Fischler. Letter to the editor. Fibonacci Quarterly, 24 (4):382–??, November 1986. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/24-4/letter2.pdf.

Horadam:1991:CAM

[HF91] Alwyn F. Horadam and Piero Filipponi. Cholesky algorithm matrices of Fibonacci type and properties of generalized sequences. Fibonacci Quarterly, 29(2):164–173, May 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-2/horadam2.pdf.

Horadam:1995:RPP

[HF95] Alwyn F. Horadam and Piero Filipponi. Real Pell and Pell– Lucas numbers with real subscripts. Fibonacci Quarterly, 33(5): 398–406, November 1995. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/33-5/horadam.pdf.

Horadam:1997:DSJ

[HF97] A. F. Horadam and P. Filipponi. Derivative sequences of Jacobsthal and Jacobsthal–Lucas polynomials. Fibonacci Quarterly, 35(4):352–357, November 1997. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/35-4/horadam1.
pdf.

Horadam:2001:MVP

[HF01] Alwyn F. Horadam and Piero Filipponi. Morgan–Voyce polynomial derivative sequences. Fibonacci Quarterly, 39(2):116–122, May 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/39-2/horadam.pdf.

Hill:1981:NFS

[HG81] Robert R. Hill and Kenneth L. Goldstein. A non-Fibonacci search plan with Fibonacci-like results. Fibonacci Quarterly, 19(2):131– 136, April 1981. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/19-2/hill.pdf.

Hoggatt:1965:CPG

[HH65] V. E. Hoggatt, Jr. and A. P. Hillman. The characteristic polynomial of the generalized shift matrix. Fibonacci Quarterly, 3(2): 91–94, April 1965. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/3-2/hoggatt.pdf.

Hoggatt:1971:HHS

[HH71] V. E. Hoggatt, Jr. and Walter Hansell. The hidden hexagon squares. Fibonacci Quarterly, 9(2):120–??, April 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-2/hoggatt1-a.pdf.

Hillman:1972:PGP

[HH72] A. P. Hillman and V. E. Hoggatt, Jr. A proof of Gould’s Pascal hexagon conjecture. Fibonacci Quarterly, 10(6):565–568, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/10-6/hillman-a.pdf.

Hoggatt:1978:PWP

[HH78] V. E. Hoggatt, Jr. and A. P. Hillman. A property of Wythoff pairs. Fibonacci Quarterly, 16(5):472–??, October 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-5/hoggatt2.pdf.

Hoggatt:1979:NLF

[HH79] V. E. Hoggatt, Jr. and A. P. Hillman. Nearly linear functions. Fibonacci Quarterly, 17(1):84–89, February 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-1/hoggatt.pdf.

Hoggatt:1980:RSS

[HH80] V. E. Hoggatt, Jr. and A. P. Hillman. Recursive, spectral, and self-generating sequences. Fibonacci Quarterly, 18(2):97–103, April 1980. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/18-2/hoggatt.pdf.

Horadam:1982:RRG

[HH82] A. F. Horadam and E. M. Horadam. Roots of recurrencegenerated polynomials. Fibonacci Quarterly, 20(3):219–226, August 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/20-3/horadam.pdf.

Hoft:1984:OTR

[HH84] Hartmut Höft and Margret Höft. An order-theoretic representation of the polygonal numbers. Fibonacci Quarterly, 22(4): 318–323, November 1984. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/22-4/hoft.pdf.

Hoft:1985:FSD

[HH85] Hartmut Höft and Margret Höft. A Fibonacci sequence of distributive lattices. Fibonacci Quarterly, 23(3):232–237, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/23-3/hoft.pdf.

Howard:1989:CNR

[HH89] F. T. Howard and E. K. Hayashi. Congruences for numbers of Ramanujan. Fibonacci Quarterly, 27(1):61–69, February 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-1/howard.pdf.

Halbeisen:2000:DFC

[HH00] Lorenz Halbeisen and Norbert Hungerbühler. Dual form of combinatorial problems and Laplace techniques. Fibonacci Quarterly, 38(5):395–407, November 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-
5/halbeisen.pdf.

Hawkins:2015:FIO

[HHJM15] Kyle Hawkins, Ursula Hebert-Johnson, and Ben Mathes. The Fibonacci identities of orthogonality. Linear Algebra and its Applications, 475(??):80–89, June 15, 2015. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0024379515000531.

Hilton:2002:MVR

[HHP02] Peter Hilton, Derek Allan Holton, and Jean Pedersen. Mathematical vistas: from a room with many windows. Undergraduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95064-8. xiv + 335 pp. LCCN QA93 .H533 2002. URL http://www.loc.gov/

catdir/enhancements/fy0812/00056268-d.html;   http://
www.loc.gov/catdir/toc/fy022/00056268.html;   https:/
/www.math.utah.edu/pub/tex/bib/fibquart.bib.

Hickerson:1973:RTF

[Hic73] Dean R. Hickerson. Recursion-type formulae for partitions into distinct parts. Fibonacci Quarterly, 11(3):307–311, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-3/hickerson.pdf.

Hickerson:1978:IRN

[Hic78a] Dean R. Hickerson. Identities relating the number of partitions into an even and odd number of parts. II. Fibonacci Quarterly, 16(1):5–6, February 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-1/
hickerson1.pdf.

Hickerson:1978:IRC

[Hic78b] Dean R. Hickerson. An identity relating compositions and partitions. Fibonacci Quarterly, 16(1):23–26, February 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-1/hickerson2.pdf.

Higgins:1972:SAP

[Hig72] John C. Higgins. Subsemigroups of the additive positive integers. Fibonacci Quarterly, 10(3):225–230, April 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-3/higgins.pdf.

Higgins:1979:MTS

[Hig79] Rada Higgins. More in the theory of sequences. Fibonacci Quarterly, 17(3):193–197, October 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
3/higgins.pdf.

Higgins:1987:NPF

[Hig87] Peter M. Higgins. The naming of popes and a Fibonacci sequence in two noncommuting indeterminates. Fibonacci Quarterly, 25 (1):57–61, February 1987. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/25-1/higgins.pdf.

Hillman:1964:EPSa

[Hil64a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 2(1):72–80, February 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-1/
elementary2-1.pdf.

Hillman:1964:EPSb

[Hil64b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 2(2):154–160, April 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-2/
elementary2-2.pdf.

Hillman:1964:EPSc

[Hil64c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 2(3):231–240, October 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-3/
elementary2-3.pdf.

Hillman:1964:EPSd

[Hil64d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 2(4):323–330, December 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-4/
elementary2-4.pdf.

Hillman:1965:EPSa

[Hil65a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 3(1):74–80, February 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-1/
elementary3-1.pdf.

Hillman:1965:EPSb

[Hil65b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 3(2):153–159, April 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-2/
elementary3-2.pdf.

Hillman:1965:EPSc

[Hil65c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 3(3):235–240, October 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-3/
elementary3-3.pdf.

Hillman:1965:EPSd

[Hil65d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 3(4):323–329, December 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-4/
elementary3-4.pdf.

Hillman:1966:EPSa

[Hil66a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 4(1):90–96, February 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-1/
elementary4-1.pdf.

Hillman:1966:EPSb

[Hil66b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 4(2):190–192, April 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-2/
elementary4-2.pdf.

Hillman:1966:EPSc

[Hil66c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 4(3):285–288, October 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-3/
elementary4-3.pdf.

Hillman:1966:EPSd

[Hil66d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 4(4):373–378, December 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-4/
elementary4-4.pdf.

Hillman:1967:EPSa

[Hil67a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 5(1):107–112, February 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-1/
elementary5-1.pdf.

Hillman:1967:EPSb

[Hil67b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 5(2):201–207, April 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-2/
elementary5-2.pdf.

Hillman:1967:EPSc

[Hil67c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 5(3):287–293, October 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-3/
elementary5-3.pdf.

Hillman:1967:EPSd

[Hil67d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 5(5):466–471, December 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-5/
elementary5-5.pdf.

Hillman:1968:EPSa

[Hil68a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 6(1):89–95, February 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-1/
elementary6-1.pdf.

Hillman:1968:EPSb

[Hil68b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 6(2):185–191, April 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-2/
elementary6-2.pdf.

Hillman:1968:EPSc

[Hil68c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 6(4):288–298, October 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-4/
elementary6-4-a.pdf.

Hillman:1968:EPSd

[Hil68d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 6(6):400–406, December 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-6/
elementary6-6.pdf.

Hillman:1969:EPSa

[Hil69a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 7(1):107–112, February 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-1/
elementary7-1.pdf.

Hillman:1969:EPSb

[Hil69b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 7(2):218–224, April 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-2/
elementary7-2.pdf.

Hillman:1969:EPSc

[Hil69c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 7(3):331–336, October 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-3/
elementary7-3-a.pdf.

Hillman:1969:EPSd

[Hil69d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 7(5):545–551, December 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-5/
elementary7-5.pdf.

Hillman:1970:EPSa

[Hil70a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 8(1):105–112, February 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-1/
elementary8-1.pdf.

Hillman:1970:EPSb

[Hil70b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 8(3):325–334, April 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-3/
elementary8-3.pdf.

Hillman:1970:EPSc

[Hil70c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 8(4):443–448, October 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-4/
elementary8-4.pdf.

Hillman:1970:EPSd

[Hil70d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 8(5):542–551, December 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-5/
elementary8-5.pdf.

Hillman:1971:EPSa

[Hil71a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 9(1):106–112, February 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-1/
elementary9-1.pdf.

Hillman:1971:EPSb

[Hil71b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 9(2):217–224, April 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-2/
elementary9-2.pdf.

Hillman:1971:EPSc

[Hil71c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 9(4):438–448, October 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-4/
elementary9-4.pdf.

Hillman:1971:EPSd

[Hil71d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 9(5):545–551, December 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-5/
elementary9-5.pdf.

Hillman:1972:EPSa

[Hil72a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 10(2):218–224, February 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
2/elementary10-2.pdf.

Hillman:1972:EPSb

[Hil72b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 10(3):329–336, April 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-3/
elementary10-3.pdf.

Hillman:1972:EPSc

[Hil72c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 10(4):447–??, October 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
4/elementary10-4.pdf.

Hillman:1972:EPSd

[Hil72d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 10(6):663–666, December 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
6/elementary10-6.pdf.

Hillman:1973:EPSa

[Hil73a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 11(1):105–112, February 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
1/elementary11-1.pdf.

Hillman:1973:EPSb

[Hil73b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 11(2):224–??, April 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-2/
elementary11-2.pdf.

Hillman:1973:EPSc

[Hil73c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 11(3):333–336, October 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
3/elementary11-3.pdf.

Hillman:1973:EPSd

[Hil73d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 11(5):550–553, December 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
5/elementary11-5.pdf.

Hillman:1974:EPSa

[Hil74a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 12(1):101–106, February 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
1/elementary12-1.pdf.

Hillman:1974:EPSb

[Hil74b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 12(2):220–??, April 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-2/
elementary12-2.pdf.

Hillman:1974:EPSc

[Hil74c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 12(3):313–316, October 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
3/elementary12-3.pdf.

Hillman:1974:EPSd

[Hil74d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 12(4):403–406, December 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
4/elementary12-4.pdf.

Hilton:1974:FTE

[Hil74e] A. J. W. Hilton. On Fern’s theorem on the expansion of Fibonacci and Lucas numbers. Fibonacci Quarterly, 12(3):231–??, October 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/12-3/hilton1.pdf.

Hilton:1974:PHG

[Hil74f] A. J. W. Hilton. On the partition of Haradam’s generalized sequences into generalized Fibonacci and generalized Lucas sequences. Fibonacci Quarterly, 12(4):339–344, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-4/hilton.pdf.

Hilton:1974:STF

[Hil74g] A. J. W. Hilton. Spanning trees and Fibonacci and Lucas numbers. Fibonacci Quarterly, 12(3):259–262, October 1974. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-3/hilton2.pdf.

Hillman:1975:EPSa

[Hil75a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 13(1):94–96, February 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-
1/elementary13-1.pdf.

Hillman:1975:EPSb

[Hil75b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 13(2):190–192, April 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-2/
elementary13-2.pdf.

Hillman:1975:EPSc

[Hil75c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 13(3):285–288, October 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-
3/elementary13-3.pdf.

Hillman:1975:EPSd

[Hil75d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 13(4):373–377, December 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-
4/elementary13-4.pdf.

Hilliker:1976:IMEa

[Hil76a] David Lee Hilliker. On the infinite multinomial expansion. Fibonacci Quarterly, 14(3):203–205, October 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-3/hilliker.pdf.

Hilliker:1976:IMEb

[Hil76b] David Lee Hilliker. On the infinite multinomial expansion, II. Fibonacci Quarterly, 14(5):392–394, December 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-5/hilliker.pdf.

Hillman:1976:EPSa

[Hil76c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 14(1):93–96, February 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
1/elementary14-1.pdf.

Hillman:1976:EPSb

[Hil76d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 14(2):188–192, April 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-2/
elementary14-2.pdf.

Hillman:1976:EPSc

[Hil76e] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 14(3):286–288, October 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
3/elementary14-3.pdf.

Hillman:1976:EPSd

[Hil76f] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 14(5):470–473, December 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
5/elementary14-5.pdf.

Hilliker:1977:MT

[Hil77a] David Lee Hilliker. On the multinomial theorem. Fibonacci Quarterly, 15(1):22–24, February 1977. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/15-1/hilliker.
pdf.

Hillman:1977:EPSa

[Hil77b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 15(1):93–96, February 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
1/elementary15-1.pdf.

Hillman:1977:EPSb

[Hil77c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 15(2):189–192, April 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-2/
elementary15-2.pdf.

Hillman:1977:EPSc

[Hil77d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 15(3):285–288, October 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
3/elementary15-3.pdf.

Hillman:1977:EPSd

[Hil77e] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 15(4):375–377, December 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
4/elementary15-4.pdf.

Hillman:1978:EPSa

[Hil78a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 16(1):88–91, February 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
1/elementary16-1.pdf.

Hillman:1978:EPSb

[Hil78b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 16(2):184–187, April 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-2/
elementary16-2.pdf.

Hillman:1978:EPSc

[Hil78c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 16(5):473–476, October 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
5/elementary16-5.pdf.

Hillman:1978:EPSd

[Hil78d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 16(6):562–565, December 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
6/elementary16-6.pdf.

Hillman:1979:EPSa

[Hil79a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 17(1):90–93, February 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
1/elementary17-1.pdf.

Hillman:1979:EPSb

[Hil79b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 17(2):184–187, April 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-2/
elementary17-2.pdf.

Hillman:1979:EPSc

[Hil79c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 17(3):281–285, October 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
3/elementary17-3.pdf.

Hillman:1979:EPSd

[Hil79d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 17(4):369–373, December 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
4/elementary17-4.pdf.

Hillman:1980:EPSa

[Hil80a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 18(1):84–89, February 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-
1/elementary18-1.pdf.

Hillman:1980:EPSb

[Hil80b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 18(2):186–189, April 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-2/
elementary18-2.pdf.

Hillman:1980:EPSc

[Hil80c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 18(3):273–279, October 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-
3/elementary18-3.pdf.

Hillman:1980:EPSd

[Hil80d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 18(4):370–374, December 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-
4/elementary18-4.pdf.

Hillman:1981:EPSa

[Hil81a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 19(1):87–92, February 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
1/elementary19-1.pdf.

Hillman:1981:EPSb

[Hil81b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 19(2):184–188, April 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-2/
elementary19-2.pdf.

Hillman:1981:EPSc

[Hil81c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 19(4):377–380, October 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
4/elementary19-4.pdf.

Hillman:1981:EPSd

[Hil81d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 19(5):466–469, December 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
5/elementary19-5.pdf.

Hillman:1981:PAI

[Hil81e] A. P. Hillman. Proportional allocation in integers. Fibonacci Quarterly, 19(3):233–239, August 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
3/hillman.pdf.

Hillman:1982:EPSa

[Hil82a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 20(1):89–92, February 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
1/elementary20-1.pdf.

Hillman:1982:EPSb

[Hil82b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 20(2):179–184, May 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-2/
elementary20-2.pdf.

Hillman:1982:EPSc

[Hil82c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 20(3):279–283, August 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
3/elementary20-3.pdf.

Hillman:1982:EPSd

[Hil82d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 20(4):366–371, November 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
4/elementary20-4.pdf.

Hillman:1983:EPSa

[Hil83a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 21(1):67–73, February 1983. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/21-
1/elementary21-1.pdf.

Hillman:1983:EPSb

[Hil83b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 21(2):147–152, May 1983. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/21-2/
elementary21-2.pdf.

Hildebrandt:1987:LE

[Hil87a] Thomas H. Hildebrandt. Letter to the editor. Fibonacci Quarterly, 25(3):240–??, August 1987. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/25-3/letter.
pdf.

Hillman:1987:EPSd

[Hil87b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 25(4):370–375, November 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-
4/elementary25-4.pdf.

Hillman:1988:EPSa

[Hil88a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 26(1):85–88, February 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
1/elementary26-1.pdf.

Hillman:1988:EPSb

[Hil88b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 26(2):181–185, May 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-2/
elementary26-2.pdf.

Hillman:1988:EPSc

[Hil88c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 26(3):278–282, August 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
3/elementary26-3.pdf.

Hillman:1988:EPSd

[Hil88d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 26(4):372–376, November 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
4/elementary26-4.pdf.

Hillman:1989:EPSa

[Hil89a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 27(1):87–91, February 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-
1/elementary27-1.pdf.

Hillman:1989:EPSb

[Hil89b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 27(2):181–185, May 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-2/
elementary27-2.pdf.

Hillman:1989:EPSc

[Hil89c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 27(4):373–377, August 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-
4/elementary27-4.pdf.

Hillman:1989:EPSd

[Hil89d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 27(5):467–472, November 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-
5/elementary27-5.pdf.

Hillman:1990:EPSa

[Hil90a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 28(1):85–89, February 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
1/elementary28-1.pdf.

Hillman:1990:EPSb

[Hil90b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 28(2):182–186, May 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-2/
elementary28-2.pdf.

Hillman:1990:EPSc

[Hil90c] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 28(3):277–282, August 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
3/elementary28-3.pdf.

Hillman:1990:EPSd

[Hil90d] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 28(4):371–375, November 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
4/elementary28-4.pdf.

Hillman:1991:EPSa

[Hil91a] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 29(1):84–88, February 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-
1/elementary29-1.pdf.

Hillman:1991:EPSb

[Hil91b] A. P. Hillman. Elementary problems and solutions. Fibonacci Quarterly, 29(2):181–185, May 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-2/
elementary29-2.pdf.

Hiller:2021:FBP

[Hil21] Josh Hiller. Fibonacci’s bunny-pocalypse! American Mathematical Monthly, 128(1):78, 2021. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hinden:1977:PAR

[Hin77] Harvey J. Hinden. Phi again: a relationship between the golden ratio and the limit of a ratio of modified Bessel functions. Fibonacci Quarterly, 15(2):112, 152, April 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-2/
hinden-a.pdf.

Hindin:1978:FNC

[Hin78] Harvey J. Hindin. A figurate number curiosity: Every integer is a quadratic function of a figurate number. Fibonacci Quarterly, 16 (6):561–??, December 1978. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/16-6/hindin.pdf.

Hindin:1980:ARC

[Hin80a] Harvey J. Hindin. An alternate representation for Césaro’s Fibonacci–Lucas identity. Fibonacci Quarterly, 18(3):259–260, October 1980. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/18-3/hindin2.pdf.

Hindin:1980:TCH

[Hin80b] Harvey J. Hindin. A theorem concerning heptagonal numbers. Fibonacci Quarterly, 18(3):258–??, October 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-3/hindin1.pdf.

Hinson:1992:DPT

[Hin92] Edward K. Hinson. On the distribution of Pythagorean triples. Fibonacci Quarterly, 30(4):335–338, November 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-4/hinson.pdf.

Hindin:1997:LE

[Hin97] Harvey J. Hindin. Letter to the editor. Fibonacci Quarterly, 35 (3):229–??, August 1997. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/35-3/letter.pdf.

Hinz:2017:LS

[Hin17] Andreas M. Hinz. The Lichtenberg sequence. Fibonacci Quarterly, 55(1):2–??, February 2017. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/55-1/hinz.
pdf.

Hirschberg:1973:CDM

[Hir73] Daniel S. Hirschberg. A class of dynamic memory allocation algorithms. Communications of the Association for Computing Machinery, 16(10):615–618, October 1973. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Hirschhorn:1981:AFF

[Hir81] Michael D. Hirschhorn. The Andrews formula for Fibonacci numbers. Fibonacci Quarterly, 19(4):373–374, October 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-4/hirschhorn.pdf.

Hirose:1986:SPN

[Hir86] Shoichi Hirose. On some polygonal numbers which are, at the same time, the sums, differences, and products of two other polygonal numbers. Fibonacci Quarterly, 24(2):99–105, May 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/24-2/hirose.pdf.

Hirschhorn:2005:NTI

[Hir05] Michael D. Hirschhorn. Non-trivial intertwined second-order recurrence relations. Fibonacci Quarterly, 43(4):316–325, November 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Papers/43-4/paper43-4-5.pdf.

Hirschhorn:2006:CSO

[Hir06a] Michael D. Hirschhorn. Coupled second-order recurrences. Fibonacci Quarterly, 44(1):20–25, February 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/44-1/hirschhorn1.pdf.

Hirschhorn:2006:CTO

[Hir06b] Michael D. Hirschhorn. Coupled third-order recurrences. Fibonacci Quarterly, 44(1):26–31, February 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/44-1/hirschhorn2.pdf.

Hirschhorn:2011:AEC

[Hir11a] Michael D. Hirschhorn. Approximating Euler’s constant. Fibonacci Quarterly, 49(3):243–248, August 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/49-3/hirschhorn.pdf.

Hirschhorn:2011:FLS

[Hir11b] Michael D. Hirschhorn. Factorization of lens sequences. Fibonacci Quarterly, 49(2):110–115, May 2011. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/49-
2/hirschhorn.pdf.

Hirschhorn:2012:EAN

[Hir12a] Michael D. Hirschhorn. Estimating the Apéry numbers. Fibonacci Quarterly, 50(2):129–??, May 2012. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/50-
2/hirschhorn1.pdf.

Hirschhorn:2012:NST

[Hir12b] Michael D. Hirschhorn. The number of sequences of n tosses of a coin with k pairs of consecutive heads. Fibonacci Quarterly, 50 (2):140–??, May 2012. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/50-2/hirschhorn2.pdf.

Hirschhorn:2013:AB

[Hir13a] Michael D. Hirschhorn. The asymptotic behavior of \binom{∏_{k=0}^{n} n}{k}. Fibonacci Quarterly, 51(2):163–??, May 2013. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/51-
2/hirschhorn2.pdf.

Hirschhorn:2013:EAN

[Hir13b] Michael D. Hirschhorn. Estimating the Apéry numbers II. Fibonacci Quarterly, 51(3):215–??, August 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/51-3/hirschhorn1.pdf.

Hirschhorn:2013:NPa

[Hir13c] Michael D. Hirschhorn. A naive proof that F_{5n} ≡ 0 (mod 5). Fibonacci Quarterly, 51(3):256–??, August 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/51-3/hirschhorn2.pdf.

Hirschhorn:2013:NPb

[Hir13d] Michael D. Hirschhorn. The number of 1’s in the partitions of n. Fibonacci Quarterly, 51(4):326–??, November 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/51-4/hirschhorn.pdf.

Hirschhorn:2014:NDP

[Hir14a] Michael D. Hirschhorn. The number of different parts in the partitions of n. Fibonacci Quarterly, 52(1):10–??, February 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-1/hirschhorn.pdf.

Hirschhorn:2014:SW

[Hir14b] Michael D. Hirschhorn. Stirling without Wallis. Fibonacci Quarterly, 52(4):321–??, November 2014. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/52-
4/hirschhorn.pdf.

Hirschhorn:2015:BIC

[Hir15a] Michael D. Hirschhorn. Binomial identities and congruences for Euler numbers. Fibonacci Quarterly, 53(4):319–??, November 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/53-4/hirschhorn.pdf.

Hirschhorn:2015:CBP

[Hir15b] Michael D. Hirschhorn. A connection between π and Φ. Fibonacci Quarterly, 53(1):42–??, February 2015. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/53-1/
hirschhorn1.pdf.

Hirschhorn:2015:RLP

[Hir15c] Michael D. Hirschhorn. Ramanujan’s last problem. Fibonacci Quarterly, 53(1):48–??, February 2015. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/53-1/
HirschhornRamanujans06022014.pdf.

Hirschhorn:2018:CMP

[Hir18] Michael D. Hirschhorn. Congruences modulo 5 for partitions into at most four parts. Fibonacci Quarterly, 56(1):32–??, February 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/56-1/hirschhorn.pdf.

Hoggatt:1968:DSA

[HJ68a] Verner E. Hoggatt, Jr. and Free Jamison. Dissection of a square into n acute isosceles triangles. Fibonacci Quarterly, 6(6):390– 392, December 1968. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/6-6/hoggatt.pdf.

Holben:1968:TPP

[HJ68b] C. A. Holben and J. H. Jordan. The twin prime problem and Goldbach’s conjecture in the Gaussian integers. Fibonacci Quarterly, 6(5):81–85, November 1968. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/6-5/holben-a.
pdf.

Heed:1975:ISF

[HK75] Joseph J. Heed and Lucille Kelly. An interesting sequence of Fibonacci sequence generators. Fibonacci Quarterly, 13(1):29– 30, February 1975. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/13-1/heed.pdf.

Heed:1978:EPF

[HK78] Joseph J. Heed and Lucille Kelly. Entry points of the Fibonacci sequence and the Euler ϕ function. Fibonacci Quarterly, 16(1): 47–50, February 1978. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/16-1/heed.pdf.

Hunter:1989:ITC

[HK89] Aleck J. Hunter and M. Kovarik. Integral triangles and circles. Fibonacci Quarterly, 27(4):310–315, August 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-4/hunter.pdf.

Halter-Koch:1991:CFG

[HK91] Franz Halter-Koch. Continued fractions of given symmetric period. Fibonacci Quarterly, 29(4):298–303, November 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-4/halter-koch.pdf.

Hahn:1995:NMP

[HK95] Sang Geun Hahn and Jun Kyo Kim. A note on multiplicative partition of bipartite numbers. Fibonacci Quarterly, 33(3):283– 289, June 1995. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/33-3/hahn.pdf.

Han:2010:FNP

[HKN10] Jeong Soon Han, Hee Sik Kim, and J. Neggers. The Fibonaccinorm of a positive integer: Observations and conjectures.

International Journal of Number Theory (IJNT), 6(2):371– 385, March 2010. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL https://www.worldscientific.com/doi/10.

1142/S1793042110003009. See acknowledgement of priority [HKN11].

Han:2011:APF

[HKN11] Jeong Soon Han, Hee Sik Kim, and J. Neggers. Acknowledgment of priority: “The Fibonacci-norm of a positive integer: Observations and conjectures”. International Journal of Number Theory (IJNT), 7(3):853–854, May 2011. ISSN 1793-0421 (print),

1793-7310 (electronic). URL https://www.worldscientific.
com/doi/10.1142/S1793042111004927. See [HKN10].

Hansen:2017:HH

[HKTZ17] Thomas Dueholm Hansen, Haim Kaplan, Robert E. Tarjan, and Uri Zwick. Hollow heaps. ACM Transactions on Algorithms, 13(3):42:1–42:??, August 2017. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Hamlin:2015:KLC

[HKW15] Nathan Hamlin, Bala Krishnamoorthy, and William Webb. A knapsack-like code using recurrence sequence representations. Fibonacci Quarterly, 53(1):24–??, February 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/53-1/hamlin.pdf.

Hoggatt:1966:PIS

[HL66] V. E. Hoggatt, Jr. and D. A. Lind. A power identity for secondorder recurrent sequences. Fibonacci Quarterly, 4(3):274–284, October 1966. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/4-3/hoggatt.pdf.

Hoggatt:1967:HFP

[HL67a] V. E. Hoggatt, Jr. and D. A. Lind. The heights of Fibonacci polynomials and an associated function. Fibonacci Quarterly, 5 (2):141–152, April 1967. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/5-2/hoggatt.pdf.

Hoggatt:1967:PFN

[HL67b] V. E. Hoggatt, Jr. and D. A. Lind. A primer for the Fibonacci numbers: Part VI. Fibonacci Quarterly, 5(5):445–460, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/5-5/hoggatt.pdf.

Hoggatt:1968:SSF

[HL68] V. E. Hoggatt, Jr. and D. A. Lind. Symbolic substitutions into Fibonacci polynomials. Fibonacci Quarterly, 6(5):55–74, November 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/6-5/hoggatt.pdf.

Hoggatt:1969:CFN

[HL69a] V. E. Hoggatt, Jr. and D. A. Lind. Compositions and Fibonacci numbers. Fibonacci Quarterly, 7(3):253–266, October 1969. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-3/hoggatt.pdf.

Hoggatt:1969:DRP

[HL69b] V. E. Hoggatt, Jr. and D. A. Lind. The dying rabbit problem. Fibonacci Quarterly, 7(5):482–487, December 1969. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-5/hoggatt.pdf.

Hoggatt:1974:DPG

[HL74] Verner E. Hoggatt, Jr. and Calvin T. Long. Divisibility properties of generalized Fibonacci polynomials. Fibonacci Quarterly, 12(2): 113–120, April 1974. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/12-2/hoggatt1.pdf.

Hagis:1984:LE

[HL84] Peter Hagis, Jr. and Graham Lord. Letter to the editor. Fibonacci Quarterly, 22(4):365–??, November 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-
4/letter.pdf.

Horadam:1990:LE

[HL90] A. F. Horadam and J. Lahr. Letter to the editor. Fibonacci Quarterly, 28(1):90–??, February 1990. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/28-1/letter.
pdf.

He:2008:SCI

[HL09] Yuan He and Qunying Liao. Some congruences involving Euler numbers. Fibonacci Quarterly, 46/47(3):225–234, August 2008/ 2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/46_47-3/he.pdf.

Horadam:1979:DPS

[HLS79] A. F. Horadam, R. P. Loh, and A. G. Shannon. Divisibility properties of some Fibonacci-type sequences. Lecture Notes in Mathematics, 748:55–64, 1979. CODEN LNMAA2. ISBN 3-540-09555-1 (print), 3-540-34857-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0102684/.

Heuer:1994:FTS

[HLW94] Gerald A. Heuer and Ulrike Leopold-Wildburger. Fibonaccitype sequences and minimal solutions of discrete Silverman games. Fibonacci Quarterly, 32(1):22–43, February 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-1/heuer.pdf.

Hlynka:2017:MFD

[Hly17] Myron Hlynka. MCMC and the Fibonacci distribution. Communications in Statistics: Simulation and Computation, 46(5): 3375–3382, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hunter:1972:BBS

[HM72] J. A. H. Hunter and Joseph S. Madachy. Back-to-back: Some interesting relationships between representations of integers in various bases. Fibonacci Quarterly, 10(2):213–217, February 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-2/hunter2.pdf.

Hock:1983:ODB

[HM83] J. L. Hock and R. B. McQuistan. The occupational degeneracy for λ-bell particles on a saturated λ × N lattice space. Fibonacci Quarterly, 21(3):196–202, August 1983. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/21-
3/hock.pdf.

Horadam:1985:PPL

[HM85] A. F. Horadam and Bro. J. M. Mahon. Pell and Pell–Lucas polynomials. Fibonacci Quarterly, 23(1):7–20, February 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-1/horadam.pdf.

Horadam:1987:MPP

[HM87] A. F. Horadam and Bro. J. M. Mahon. Mixed Pell polynomials. Fibonacci Quarterly, 25(4):291–299, November 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-4/horadam.pdf.

Hendel:1994:HEC

[HM94] Russell Jay Hendel and Sandra A. Monteferrante. Hofstadter’s extraction conjecture. Fibonacci Quarterly, 32(2):98–107, May 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/32-2/hendel.pdf.

Herceg:2001:ZNI

[HML01] D. Herceg, H. Maličić, and I. Likić. The Zeckendorf numbers and the inverses of some band matrices. Fibonacci Quarterly, 39(1): 27–31, February 2001. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/39-1/herceg.pdf.

Hillman:1971:SSS

[HMM71] A. P. Hillman, P. L. Mana, and C. T. McAbee. A symmetric substitute for Stirling numbers. Fibonacci Quarterly, 9(1):51–60, February 1971. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/9-1/hillman-a.pdf.

He:2023:FVW

[HMT23] B. He, Y. L. Mao, and A. Togbé. A Fibonacci version of Wolstenholme’s harmonic series congruence. American Mathematical Monthly, 130(1):83–85, 2023. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Hoggatt:1987:HSL

[HO87] V. E. Hoggatt, Jr. and M. A. Owens. Hoggatt sequences and lexicographic ordering. Fibonacci Quarterly, 25(4):322–332, November 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/25-4/hoggatt.pdf.

Hochberg:1974:SDN

[Hoc74] Murray Hochberg. On the set of divisor of a number. Fibonacci Quarterly, 12(4):363–364, December 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
4/hochberg.pdf.

Hochberg:1979:CGT

[Hoc79] Murray Hochberg. A conjecture in game theory. Fibonacci Quarterly, 17(3):250–252, October 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
3/hochberg.pdf.

Hodel:1974:CIA

[Hod74] M. J. Hodel. Combinatorial interpretation of an analog of generalized binomial coefficients. Fibonacci Quarterly, 12(4):360–362, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/12-4/hodel.pdf.

Hodgson:1992:SNS

[Hod92] Bernard R. Hodgson. On some number sequences related to the parity of binomial coefficients. Fibonacci Quarterly, 30(1):35–47, February 1992. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/30-1/hodgson.pdf.

Hoggatt:1963:APSa

[Hog63a] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 1(1):46–48, February 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-1/
advanced1-1.pdf.

Hoggatt:1963:APSb

[Hog63b] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 1(2):53–55, April 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-2/
advanced1-2.pdf.

Hoggatt:1963:APSc

[Hog63c] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 1(3):46–52, October 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-3/
advanced1-3.pdf.

Hoggatt:1963:APSd

[Hog63d] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 1(4):47–52, December 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/1-4/advanced1-4.pdf.

Hoggatt:1964:APSa

[Hog64a] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 2(1):49–52, February 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-1/
advanced2-1.pdf.

Hoggatt:1964:APSb

[Hog64b] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 2(2):123–134, April 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-2/
advanced2-2.pdf.

Hoggatt:1964:APSc

[Hog64c] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 2(3):204–209, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/2-3/advanced2-3.pdf.

Hoggatt:1964:APSd

[Hog64d] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 2(4):303–309, December 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/2-4/advanced2-4.pdf.

Hoggatt:1964:FND

[Hog64e] Verner E. Hoggatt, Jr. Fibonacci numbers from a differential equation. Fibonacci Quarterly, 2(3):176–??, October 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-3/hoggatt1-a.pdf.

Hoggatt:1965:APSa

[Hog65a] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 3(1):44–52, February 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-1/
advanced3-1.pdf.

Hoggatt:1965:APSb

[Hog65b] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 3(2):115–128, April 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-2/
advanced3-2.pdf.

Hoggatt:1965:APSc

[Hog65c] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 3(3):201–207, October 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/3-3/advanced3-3.pdf.

Hoggatt:1965:APSd

[Hog65d] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 3(4):299–303, December 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/3-4/advanced3-4.pdf.

Hoggatt:1966:APSa

[Hog66a] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 4(1):56–58, February 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-1/
advanced4-1.pdf.

Hoggatt:1966:APSb

[Hog66b] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 4(2):148–149, April 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-2/
advanced4-2.pdf.

Hoggatt:1966:APSc

[Hog66c] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 4(3):251–258, October 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/4-3/advanced4-3.pdf.

Hoggatt:1966:APSd

[Hog66d] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 4(4):332–338, December 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/4-4/advanced4-4.pdf.

Hoggatt:1967:FNG

[Hog67a] V. E. Hoggatt, Jr. Fibonacci numbers and generalized binomial coefficients. Fibonacci Quarterly, 5(4):383–400, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-4/hoggatt.pdf.

Hoggatt:1967:APSa

[Hog67b] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 5(1):69–80, February 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-1/
advanced5-1.pdf.

Hoggatt:1967:APSb

[Hog67c] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 5(2):161–??, April 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-2/
advanced5-2.pdf.

Hoggatt:1967:APSc

[Hog67d] Verner E. Hoggatt, Jr. Advanced problems and solutions. Fibonacci Quarterly, 5(3):251–258, October 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/5-3/advanced5-3.pdf.

Hoggatt:1968:GRG

[Hog68a] V. E. Hoggatt, Jr. Generalized rabbits for generalized Fibonacci numbers. Fibonacci Quarterly, 6(3):105–192, June 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-3/hoggatt.pdf.

Hoggatt:1968:NAP

[Hog68b] V. E. Hoggatt, Jr. A new angle on Pascal’s triangle. Fibonacci Quarterly, 6(4):221–234, October 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-4/
hoggatt.pdf.

Hoggatt:1968:BA

[Hog68c] Verner E. Hoggatt, Jr. Belated acknowledgement. Fibonacci Quarterly, 6(3):85–??, June 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-3/
acknowledgement.pdf.

Hoggatt:1969:FLN

[Hog69] Verner E. Hoggatt, Jr. Fibonacci and Lucas numbers. Houghton Mifflin mathematics enrichment series; Houghton Mifflin modern mathematics series. Houghton-Mifflin, Boston, MA, USA, 1969. iv + 92 pp. LCCN QA241 .H57.

Hoggatt:1970:ALT

[Hog70a] Verner E. Hoggatt, Jr. An application of the Lucas triangle. Fibonacci Quarterly, 8(4):360–364, October 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-4/hoggatt-a.pdf.

Hoggatt:1970:CPG

[Hog70b] Verner E. Hoggatt, Jr. Combinatorial problems for generalized Fibonacci numbers. Fibonacci Quarterly, 8(5):456–462, December 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/8-5/hoggatt.pdf.

Hoggatt:1970:CTG

[Hog70c] Verner E. Hoggatt, Jr. Convolution triangles for generalized Fibonacci numbers. Fibonacci Quarterly, 8(2):158–171, March 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-2/hoggatt.pdf.

Hoggatt:1971:SMF

[Hog71a] V. E. Hoggatt, Jr. Some more Fibonacci Diophantine equations. Fibonacci Quarterly, 9(4):437–??, October 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-4/hoggatt2-a.pdf.

Hoggatt:1971:SSF

[Hog71b] Verner E. Hoggatt, Jr. Some special Fibonacci and Lucas generating functions. Fibonacci Quarterly, 9(2):121–133, April 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-2/hoggatt2.pdf.

Hoggatt:1972:GFN

[Hog72a] V. E. Hoggatt, Jr. Generalized Fibonacci numbers in Pascal’s pyramid. Fibonacci Quarterly, 10(3):271–276, April 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-3/hoggatt-a.pdf.

Hoggatt:1972:GZT

[Hog72b] V. E. Hoggatt, Jr. Generalized Zeckendorf theorem. Fibonacci Quarterly, 10(1):89–94, January 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
1/hoggatt2.pdf.

Hoggatt:1976:R

[Hog76] Verner E. Hoggatt, Jr. Response. Fibonacci Quarterly, 14(5): 455–??, December 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-5/hoggatt3.pdf.

Hoggatt:1977:AP

[Hog77a] V. E. Hoggatt, Jr. Additive partition I. Fibonacci Quarterly, 15(2):166–??, April 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-2/hoggatt3.pdf.

Hoggatt:1977:API

[Hog77b] V. E. Hoggatt, Jr. Additive partition II. Fibonacci Quarterly, 15(2):182–??, April 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-2/hoggatt4.pdf.

Hoggatt:1977:NSS

[Hog77c] V. E. Hoggatt, Jr. A note on the summation of squares. Fibonacci Quarterly, 15(4):367–369, December 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
4/hoggatt2.pdf.

Hoggatt:1980:APP

[Hog80] V. E. Hoggatt, Jr. Additive partitions of the positive integers. Fibonacci Quarterly, 18(3):220–225, October 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-3/hoggatt.pdf.

Hoit:2022:AUB

[Hoi22] Abigail Hoit. An asymptotic upper bound for counting fiven numbers. Fibonacci Quarterly, 60(1):52–??, February 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-1/hoit.pdf.

Holt:1965:MPP

[Hol65] Marvin H. Holt. Mystery puzzler and phi. Fibonacci Quarterly, 3(2):135–138, April 1965. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/3-2/holt.pdf.

Holladay:1966:CSC

[Hol66a] John C. Holladay. Corrigendum for “Some Convergent Recursive Sequences, Homomorphic Identities, and Inductively Defined Complementary Sequences”. Fibonacci Quarterly, 4(3):249–250, October 1966. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/4-3/holladay.pdf. See [Hol66b].

Holladay:1966:SCR

[Hol66b] John C. Holladay. Some convergent recursive sequences, homeomorphic identities, and inductively defined complementary sequences. Fibonacci Quarterly, 4(1):1–36, February 1966. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-1/holladay.pdf. See corrigendum [Hol66a].

Holden:1975:FT

[Hol75] Herbert L. Holden. Fibonacci tiles. Fibonacci Quarterly, 13(1): 45–49, February 1975. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/13-1/holden.pdf.

Holte:1994:LTT

[Hol94] John M. Holte. A Lucas-type theorem for Fibonomial-coefficient residues. Fibonacci Quarterly, 32(1):60–68, February 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-1/holte.pdf.

Holt:1995:APF

[Hol95] Derek F. Holt. An alternative proof that the Fibonacci group F (2, 9) is infinite. Experimental Mathematics, 4(2):97–100, ???? 1995. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://projecteuclid.org/euclid.
em/1047931620.

Holte:2000:RGB

[Hol00] John M. Holte. Residues of generalized binomial coefficients modulo a prime. Fibonacci Quarterly, 38(3):227–238, June 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-3/holte.pdf.

Holte:2006:FDA

[Hol06] John M. Holte. Fractal dimension of arithmetical structures of generalized binomial coefficients modulo a prime. Fibonacci Quarterly, 44(1):46–58, February 2006. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/44-1/
holte.pdf.

Holden:2011:RC

[Hol11] Dhiraj Holden. Results on the 3x + 1 and 3x + d conjectures. Fibonacci Quarterly, 49(2):131–133, May 2011. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/49-
2/holden.pdf.

Hong:2015:WGF

[Hon15] Dae S. Hong. When is the generating function of the Fibonacci numbers an integer? College Mathematics Journal, 46(2):110– 112, March 2015. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/abs/10.4169/college.math.j.46.2.110.

Hope:1995:EGR

[Hop95] Peter Hope. Exponential growth of random Fibonacci sequences. Fibonacci Quarterly, 33(2):164–168, May 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-2/hope.pdf.

Horadam:1961:CGF

[Hor61a] A. F. Horadam. Corrections: a generalized Fibonacci sequence. American Mathematical Monthly, 68(7):637, August/September 1961. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [Hor61c].

Horadam:1961:FNT

[Hor61b] A. F. Horadam. Fibonacci number triples. American Mathematical Monthly, 68(8):751–753, October 1961. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Horadam:1961:GFS

[Hor61c] A. F. Horadam. A generalized Fibonacci sequence. American Mathematical Monthly, 68(5):455–459, May 1961. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See corrections [Hor61a].

Horadam:1963:FAF

[Hor63a] A. F. Horadam. Further appearance of the Fibonacci sequence. Fibonacci Quarterly, 1(4):41–42, December 1963. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-4/horadam-a.pdf.

Horadam:1963:MNC

[Hor63b] A. F. Horadam. Mathematical notes: Complex Fibonacci numbers and Fibonacci quaternions. American Mathematical Monthly, 70(3):289–291, March 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Horner:1964:FP

[Hor64] W. W. Horner. Fibonacci and Pascal. Fibonacci Quarterly, 2(3): 228–??, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/2-3/horner.pdf.

Horadam:1965:BPC

[Hor65] A. F. Horadam. Basic properties of a certain generalized sequence of numbers. Fibonacci Quarterly, 3(3):161–176, October 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/3-3/horadam.pdf.

Horner:1966:FE

[Hor66] W. W. Horner. Fibonacci and Euclid. Fibonacci Quarterly, 4(2): 168–169, April 1966. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/4-2/horner.pdf.

Horadam:1967:SPS

[Hor67] A. F. Horadam. Special properties of the sequence W_{n}(a, b; p, q). Fibonacci Quarterly, 5(5):424–434, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/5-5/horadam.pdf.

Horadam:1969:TOF

[Hor69] A. F. Horadam. Tschebyscheff and other functions associated with the sequence {w_{n}(a, b; p, q)}. Fibonacci Quarterly, 7(1):14– 22, February 1969. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/7-1/horadam.pdf.

Horadam:1971:PI

[Hor71] A. F. Horadam. Pell identities. Fibonacci Quarterly, 9(3):245– 252, May 1971. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/9-3/horadam-a.pdf.

Horner:1973:FA

[Hor73] Walter W. Horner. Fibonacci and Apollonius. Fibonacci Quarterly, 11(5):541–542, December 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
5/horner.pdf.

Horadam:1974:GFP

[Hor74a] A. F. Horadam. On generating functions for powers of a generalized sequence of numbers. Fibonacci Quarterly, 12(4):348–??, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/12-4/horadam-a.pdf.

Horadam:1974:ON

[Hor74b] A. F. Horadam. Oresme numbers. Fibonacci Quarterly, 12(3): 267–270, October 1974. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/12-3/horadam.pdf.

Horadam:1975:EHY

[Hor75] A. F. Horadam. Eight hundred years young. The Australian Mathematics Teacher, 31(??):123–134, ???? 1975. ISSN 0045-0685.

Horadam:1977:GIG

[Hor77a] A. F. Horadam. Generating identities for generalized Fibonacci and Lucas sequences. Fibonacci Quarterly, 15(4):289–292, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/15-4/horadam.pdf.

Horadam:1977:PAC

[Hor77b] A. F. Horadam. Polynomials associated with Chebyshev polynomials of the first kind. Fibonacci Quarterly, 15(3):255–257, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/15-3/horadam.pdf.

Horadam:1978:DF

[Hor78a] A. F. Horadam. Diagonal functions. Fibonacci Quarterly, 16(1): 33–36, February 1978. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/16-1/horadam.pdf.

Horadam:1978:WP

[Hor78b] A. F. Horadam. Wythoff pairs. Fibonacci Quarterly, 16(2):147– 151, April 1978. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/16-2/horadam.pdf.

Horadam:1978:SSS

[Hor78c] E. M. Horadam. Solved, semi-solved, and unsolved problems in generalized integers: a survey. Fibonacci Quarterly, 16(4):370– 380, August 1978. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/16-4/horadam.pdf.

Horadam:1979:CFP

[Hor79a] A. F. Horadam. Chebyshev and Fermat polynomials for diagonal functions. Fibonacci Quarterly, 17(4):328–332, December 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-4/horadam.pdf.

Horadam:1979:SPE

[Hor79b] A. F. Horadam. Sums of products: An extension. Fibonacci Quarterly, 17(3):248–249, October 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
3/horadam.pdf.

Horadam:1980:EPD

[Hor80] A. F. Horadam. Extensions of a paper on diagonal functions. Fibonacci Quarterly, 18(1):3–8, February 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-1/
horadam.pdf.

Horadam:1982:GGS

[Hor82a] A. F. Horadam. Geometry of a generalized Simson’s formula. Fibonacci Quarterly, 20(2):164–167, May 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-2/
horadam2.pdf.

Horadam:1982:PT

[Hor82b] A. F. Horadam. Pythagorean triples. Fibonacci Quarterly, 20(2): 121–??, May 1982. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/20-2/horadam1.pdf.

Horibe:1982:EVF

[Hor82c] Yasuichi Horibe. An entropy view of Fibonacci trees. Fibonacci Quarterly, 20(2):168–178, May 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-2/
horibe.pdf.

Horibe:1983:NFT

[Hor83] Yasuichi Horibe. Notes on Fibonacci trees and their optimality. Fibonacci Quarterly, 21(2):118–128, May 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/21-2/horibe.pdf.

Horadam:1984:CCA

[Hor84a] A. F. Horadam. Coaxal circles associated with recurrencegenerated sequences. Fibonacci Quarterly, 22(3):270–272, August 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/22-3/horadam-a.pdf.

Horadam:1984:PNC

[Hor84b] A. F. Horadam. Pell numbers and coaxal circles. Fibonacci Quarterly, 22(4):324–326, November 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-
4/horadam.pdf.

Horadam:1985:GPR

[Hor85a] A. F. Horadam. Gegenbauer polynomials revisited. Fibonacci Quarterly, 23(4):294–299, November 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-
4/horadam-a.pdf.

Horadam:1985:ZP

[Hor85b] A. F. Horadam. Zigzag polynomials. Fibonacci Quarterly, 23(3): 214–220, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/23-3/horadam.pdf.

Horadam:1986:DHL

[Hor86a] A. F. Horadam. Determinantal hypersurfaces for Lucas sequences of order r, and a generalization. Fibonacci Quarterly, 24(3):227– 237, August 1986. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/24-3/horadam2.pdf.

Horadam:1986:GZP

[Hor86b] A. F. Horadam. Generalized zigzag polynomials. Fibonacci Quarterly, 24(1):8–16, February 1986. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/24-1/horadam.
pdf.

Horadam:1986:HAS

[Hor86c] A. F. Horadam. Hypersurfaces associated with Simson formula analogues. Fibonacci Quarterly, 24(3):221–226, August 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-3/horadam1.pdf.

Horadam:1988:EFL

[Hor88a] A. F. Horadam. Elliptic functions and Lambert series in the summation of reciprocals in certain recurrence-generated sequences. Fibonacci Quarterly, 26(2):98–114, May 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-2/
horadam.pdf.

Horadam:1988:JPC

[Hor88b] A. F. Horadam. Jacobsthal and Pell curves. Fibonacci Quarterly, 26(1):77–83, February 1988. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/26-1/horadam2.pdf.

Horak:1988:NTO

[Hor88c] Pavel Horák. A note on the third-order strong divisibility sequences. Fibonacci Quarterly, 26(4):366–371, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/26-4/horak.pdf.

Horibe:1988:ETD

[Hor88d] Yasuichi Horibe. Entropy of terminal distributions and the Fibonacci trees. Fibonacci Quarterly, 26(2):135–140, May 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/26-2/horibe.pdf.

Horadam:1991:FML

[Hor91] A. F. Horadam. Fibonacci’s mathematical letter to Master Theodorus. Fibonacci Quarterly, 29(2):103–107, May 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-2/horadam1.pdf.

Horadam:1992:GGP

[Hor92a] A. F. Horadam. Generation of Genocchi polynomials of first order by recurrence relations. Fibonacci Quarterly, 30(3):239– 243, August 1992. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/30-3/horadam.pdf.

Horadam:1992:NOG

[Hor92b] A. F. Horadam. Negative order Genocchi polynomials. Fibonacci Quarterly, 30(1):21–34, February 1992. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/30-
1/horadam.pdf.

Horak:1992:SDL

[Hor92c] Pavel Horák. Strong divisibility linear recurrences of the third order. Fibonacci Quarterly, 30(2):98–101, May 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-2/horak.pdf.

Horibe:1992:FTB

[Hor92d] Yasuichi Horibe. A Fibonacci theme on balanced binary trees. Fibonacci Quarterly, 30(3):244–250, August 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-3/horibe.pdf.

Horadam:1993:ASG

[Hor93] A. F. Horadam. Associated sequences of general order. Fibonacci Quarterly, 31(2):166–172, May 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-2/
horadam.pdf.

Horadam:1994:APU

[Hor94a] A. F. Horadam. An alternative proof of a unique representation theorem. Fibonacci Quarterly, 32(5):409–410, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-5/horadam1.pdf.

Horadam:1994:MRP

[Hor94b] A. F. Horadam. Maximal representations of positive integers by Pell numbers. Fibonacci Quarterly, 32(3):240–244, June 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-3/horadam2.pdf.

Horadam:1994:PSS

[Hor94c] A. F. Horadam. Partial sums for second-order recurrence sequences. Fibonacci Quarterly, 32(5):429–439, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-5/horadam2.pdf.

Horadam:1994:UMR

[Hor94d] A. F. Horadam. Unique minimal representation of integers by negatively subscripted Pell numbers. Fibonacci Quarterly, 32(3): 202–206, June 1994. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/32-3/horadam1.pdf.

Horadam:1996:ESC

[Hor96a] A. F. Horadam. Extension of a synthesis for a class of polynomial sequences. Fibonacci Quarterly, 34(1):68–74, February 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/34-1/horadam3.pdf.

Horadam:1996:JRN

[Hor96b] A. F. Horadam. Jacobsthal representation numbers. Fibonacci Quarterly, 34(1):40–54, February 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-
1/horadam2.pdf.

Horadam:1996:MP

[Hor96c] A. F. Horadam. Minmax polynomials. Fibonacci Quarterly, 34 (1):7–16, February 1996. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/34-1/horadam1.pdf.

Horadam:1996:PAG

[Hor96d] A. F. Horadam. Polynomials associated with generalized Morgan–Voyce polynomials. Fibonacci Quarterly, 34(4):342–348, August 1996. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/34-4/horadam.pdf.

Horadam:1997:CMV

[Hor97a] A. F. Horadam. A composite of Morgan–Voyce generalizations. Fibonacci Quarterly, 35(3):233–239, August 1997. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-3/horadam.pdf.

Horadam:1997:JRP

[Hor97b] A. F. Horadam. Jacobsthal representation polyomials. Fibonacci Quarterly, 35(2):137–148, May 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-2/
horadam.pdf.

Horadam:1997:RFJ

[Hor97c] A. F. Horadam. Rodriques’ formulas for Jacobsthal-type polynomials. Fibonacci Quarterly, 35(4):361–370, November 1997. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-4/horadam2.pdf.

Horadam:1998:MVT

[Hor98] A. F. Horadam. Morgan–Voyce type generalized polynomials with negative subscripts. Fibonacci Quarterly, 36(5):391–394, November 1998. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/36-5/horadam.pdf.

Horadam:1999:RGC

[Hor99] A. F. Horadam. Representation grids for certain Morgan–Voyce numbers. Fibonacci Quarterly, 37(4):320–325, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-4/horadam.pdf.

Horadam:2000:CNV

[Hor00a] A. F. Horadam. Completion of numerical values of generalized Morgan–Voyce and related polynomials. Fibonacci Quarterly, 38 (3):260–263, June 2000. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/38-3/horadam2.pdf.

Horadam:2000:CSP

[Hor00b] A. F. Horadam. Convolution summations for Pell and Pell–Lucas numbers. Fibonacci Quarterly, 38(5):451–462, November 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-5/horadam.pdf.

Horadam:2000:URM

[Hor00c] A. F. Horadam. Uniqueness of representations by Morgan–Voyce numbers. Fibonacci Quarterly, 38(3):212–215, June 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-3/horadam1.pdf.

Horadam:2002:CJT

[Hor02a] A. F. Horadam. Convolutions for Jacobsthal-type polynomials. Fibonacci Quarterly, 40(3):212–222, June 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-3/horadam1.pdf.

Horadam:2002:MVC

[Hor02b] A. F. Horadam. Morgan–Voyce convolutions. Fibonacci Quarterly, 40(2):98–105, May 2002. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/40-2/horadam.
pdf.

Horadam:2002:VP

[Hor02c] A. F. Horadam. Vieta polynomials. Fibonacci Quarterly, 40(3): 223–232, June 2002. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/40-3/horadam2.pdf.

Horibe:2002:FTC

[Hor02d] Yasuichi Horibe. Fibonacci tree is critically balanced — a note. Fibonacci Quarterly, 40(5):441–445, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-5/horibe.pdf.

Horadam:2003:UPQ

[Hor03a] A. F. Horadam. Unexpected Pell and quasi Morgan–Voyce summation connections. Fibonacci Quarterly, 41(4):352–359, August 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/41-4/horadam.pdf.

Horadam:2003:VCD

[Hor03b] A. F. Horadam. Vieta convolutions and diagonal polynomials. Fibonacci Quarterly, 41(3):240–252, June 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-3/
horadam.pdf.

Horadam:2004:FLA

[Hor04] A. F. Horadam. “Fibonacci’s Liber Abaci”: A Translation into Modern English of Leonardo Pisano’s Book of Calculation, by L. E. Sigler. Fibonacci Quarterly, 42(1):82–85, February 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-1/quarthoradam04review.pdf. See [Sig02].

Horadam:2005:CPC

[Hor05] A. F. Horadam. Chebyshev and Pell connections. Fibonacci Quarterly, 43(2):108–121, May 2005. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/43-2/
paper43-2-3.pdf.

Hosoya:1973:TIF

[Hos73] Haruo Hosoya. Topological index and Fibonacci numbers with relation to chemistry. Fibonacci Quarterly, 11(3):255–265, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/11-3/hosoya.pdf.

Hosoya:1976:FT

[Hos76] Haruo Hosoya. Fibonacci triangle. Fibonacci Quarterly, 14(2): 173–179, April 1976. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/14-2/hosoya.pdf.

Howard:1980:ASN

[How80] F. T. Howard. Associated Stirling numbers. Fibonacci Quarterly, 18(4):303–315, December 1980. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/18-4/howard.
pdf.

Howard:1984:WAS

[How84] F. T. Howard. Weighted associated Stirling numbers. Fibonacci Quarterly, 22(2):156–165, May 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-2/
howard.pdf.

Howard:1985:IRB

[How85] F. T. Howard. Integers related to the Bessel function J_{1}(z). Fibonacci Quarterly, 23(3):249–257, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/23-3/howard.pdf.

Howard:1986:EFN

[How86] F. T. Howard. Explicit formulas for numbers of Ramanujan. Fibonacci Quarterly, 24(2):168–175, May 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-2/
howard.pdf.

Howard:1987:RBF

[How87] F. T. Howard. The reciprocal of the Bessel function J_{k}(z). Fibonacci Quarterly, 25(4):304–311, November 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-4/howard.pdf.

Howard:1990:ECG

[How90] F. T. Howard. Extensions of congruences of Glaisher and Nielsen concerning Stirling numbers. Fibonacci Quarterly, 28(4):355–362, November 1990. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/28-4/howard.pdf.

Howard:1993:MBC

[How93] F. T. Howard. Multinomial and Q-binomial coefficient modulo 4 and modulo P. Fibonacci Quarterly, 31(1):53–64, February 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/31-1/howard.pdf.

Howard:1994:CRB

[How94] F. T. Howard. Congruences and recurrences for Bernoulli numbers of higher orders. Fibonacci Quarterly, 32(4):316–328, August 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/32-4/howard.pdf.

Howard:1996:SPI

[How96] F. T. Howard. Sums of powers of integers via generating functions. Fibonacci Quarterly, 34(3):244–256, June 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-3/howard.pdf.

Howard:1998:LRS

[How98] F. T. Howard. Lacunary recurrences for sums of powers of integers. Fibonacci Quarterly, 36(5):435–442, November 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-5/howard.pdf.

Howard:2000:MLC

[How00] F. T. Howard. In memoriam — Lenard Carlitz. Fibonacci Quarterly, 38(4):316–??, August 2000. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/38-4/howard.
pdf.

Howard:2001:PDC

[How01a] F. T. Howard. The power of 2 dividing the coefficients of certain power series. Fibonacci Quarterly, 39(4):358–364, August 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-4/howard2.pdf.

Howard:2001:TI

[How01b] F. T. Howard. A Tribonacci identity. Fibonacci Quarterly, 39 (4):352–357, August 2001. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/39-4/howard1.pdf.

Howard:2003:SST

[How03] F. T. Howard. The sum of the squares of two generalized Fibonacci numbers. Fibonacci Quarterly, 41(1):80–84, February 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/41-1/howard.pdf.

Hoyrup:2005:LFA

[Høy05] Jens Høyrup. Leonardo Fibonacci and Abbaco culture. A proposal to invert the roles. Revue d’Histoire des Mathématiques = Journal for the History of Mathematics, 11(1):23–56, 2005. ISSN 1262-022X (print), 1777-568X (electronic). URL http://

smf4.emath.fr/Publications/RevueHistoireMath/11/html/
smf_rhm_11_23-56.html.

Hoyrup:2009:BRBa

[Høy09] Jens Høyrup. Book review: Fibonacci’s De Practica Geometrie, Barnabas Hughes (Ed.), in: Sources and Studies in the History of Mathematics and Physical Sciences. Springer, New York (2008), ISBN 978-0-387-72930-5. Historia Mathematica, 36(1):73–77, February 2009. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0315086008000566.

Hoggatt:1972:SGR

[HP72] V. E. Hoggatt, Jr. and Brian Peterson. Some general results on representations. Fibonacci Quarterly, 10(1):81–88, January 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-1/hoggatt1.pdf.

Hoggatt:1975:FLS

[HP75a] V. E. Hoggatt, Jr. and John W. Phillips. Fibonacci and Lucas sums in the r-normal triangle — abstract. Fibonacci Quarterly, 13(2):161–??, April 1975. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/13-2/hoggatt2.pdf.

Hunsucker:1975:IP

[HP75b] J. L. Hunsucker and Carl Pomerance. On an interesting property of 11234955056179752809. Fibonacci Quarterly, 13(4):331–333, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/13-4/hunsucker.pdf.

Horadam:1981:PAG

[HP81] A. F. Horadam and S. Pethe. Polynomials associated with Gegenbauer polynomials. Fibonacci Quarterly, 19(5):393–397, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/19-5/horadam.pdf.

Hillman:1983:EPSc

[HP83] A. P. Hillman and G. C. Padilla. Elementary problems and solutions. Fibonacci Quarterly, 21(3):230–235, August 1983. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-3/elementary21-3.pdf.

Hillman:1984:EPSd

[HP84] A. P. Hillman and Gloria C. Padilla. Elementary problems and solutions. Fibonacci Quarterly, 22(4):369–373, November 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-4/elementary22-4.pdf.

Horadam:1986:ECG

[HP86] A. F. Horadam and S. Pethe. Euclidean coordinates as generalized Fibonacci numbers products. Fibonacci Quarterly, 24(4): 366–370, November 1986. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/24-4/horadam2.pdf.

Hsu:1992:EML

[HP92] Wen-Jing Hsu and C. V. Page. Embedding mesh in a large family of graphs. Parallel Processing Letters, 2(2-3):149–155, September 1992. CODEN PPLTEE. ISSN 0129-6264.

Hilton:1994:NGP

[HP94] Peter Hilton and Jean Pedersen. A note on a geometrical property of Fibonacci numbers. Fibonacci Quarterly, 32(5):386–388, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/32-5/hilton.pdf.

Hoggatt:1971:TFM

[HPL71] V. E. Hoggatt, Jr., John W. Phillips, and H. T. Leonard, Jr. Twenty-four master identities. Fibonacci Quarterly, 9(1):1–17, February 1971. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/9-1/hoggatt.pdf.

Hsu:1993:FCC

[HPL93] W.-J. Hsu, C. V. Page, and J.-S. Liu. Fibonacci cubes — a class of self-similar graphs. Fibonacci Quarterly, 31(1):65–71, February 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/31-1/hsu.pdf.

Hilton:1997:LN

[HPS97] Peter Hilton, Jean Pedersen, and Lawrence Somer. On Lucasian numbers. Fibonacci Quarterly, 35(1):43–47, February 1997. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-1/hilton.pdf.

Hare:2014:TSG

[HPS14] Kevin Hare, Helmut Prodinger, and Jeffrey Shallit. Three series for the generalized golden mean. Fibonacci Quarterly, 52(4):307– ??, November 2014. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/52-4/hare.pdf.

Hilton:1995:CAP

[HPV95] Peter Hilton, Jean Pedersen, and Luc Vrancken. On certain arithmetic properties of Fibonacci and Lucas numbers. Fibonacci Quarterly, 33(3):211–217, June 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-3/
hilton.pdf.

Hillman:1983:EPSd

[HPW83] A. P. Hillman, G. C. Padalla, and C. R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 21(4):306–311, November 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/21-4/elementary21-4.pdf.

Hillman:1984:EPSa

[HPW84a] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 22(1):84–88, February 1984. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/22-1/elementary22-1.pdf.

Hillman:1984:EPSb

[HPW84b] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 22(2):183–187, May 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/22-2/elementary22-2.pdf.

Hillman:1984:EPSc

[HPW84c] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 22(3):273–278, August 1984. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/22-3/elementary22-3.pdf.

Hillman:1985:EPSa

[HPW85a] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 23(1):85–88, February 1985. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/23-1/elementary23-1.pdf.

Hillman:1985:EPSb

[HPW85b] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 23(2):181–185, May 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/23-2/elementary23-2.pdf.

Hillman:1985:EPSc

[HPW85c] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 23(3):277–281, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/23-3/elementary23-3.pdf.

Hillman:1985:EPSd

[HPW85d] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 23(4):371–375, November 1985. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/23-4/elementary23-4.pdf.

Hillman:1986:EPSa

[HPW86a] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 24(1):84–87, February 1986. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/24-1/elementary24-1.pdf.

Hillman:1986:EPSb

[HPW86b] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 24(2):180–184, May 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/24-2/elementary24-2.pdf.

Hillman:1986:EPSc

[HPW86c] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 24(3):277–282, August 1986. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/24-3/elementary24-3.pdf.

Hillman:1986:EPSd

[HPW86d] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 24(4):371–375, November 1986. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/24-4/elementary24-4.pdf.

Hillman:1987:EPSa

[HPW87a] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 25(1):85–89, February 1987. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/25-1/elementary25-1.pdf.

Hillman:1987:EPSb

[HPW87b] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 25(2):180–184, May 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/25-2/elementary25-2.pdf.

Hillman:1987:EPSc

[HPW87c] A. P. Hillman, Gloria C. Padilla, and Charles R. Wall. Elementary problems and solutions. Fibonacci Quarterly, 25(3):279–283, August 1987. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/25-3/elementary25-3.pdf.

Haukkanen:1991:GFP

[HR91] Pentti Haukkanen and Jerzy Rutkowski. On generating functions for powers of recurrence sequences. Fibonacci Quarterly, 29(4): 329–332, November 1991. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/29-4/haukkanen.pdf.

Holzer:1996:SGF

[HR96] Markus Holzer and Peter Rossmanith. A simpler grammar for Fibonacci numbers. Fibonacci Quarterly, 34(5):465–466, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/34-5/holzer.pdf.

Holshouser:2003:DOP

[HRR03] Arthur Holshouser, Harold Reiter, and James Rudzinski. Dynamic one-pile Nim. Fibonacci Quarterly, 41(3):253–262, June 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/41-3/holshouser.pdf.

Harris:1964:GFN

[HS64] V. C. Harris and Carolyn C. Styles. A generalization of Fibonacci numbers. Fibonacci Quarterly, 2(4):277–289, December 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/2-4/harris.pdf.

Harris:1966:GFS

[HS66] V. C. Harris and Carolyn C. Styles. Generalized Fibonacci sequences associated with a generalized Pascal triangle. Fibonacci Quarterly, 4(3):241–248, October 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-3/
harris.pdf.

Hitotumatu:1975:SDT

[HS75] Sin Hitotumatu and Daihachiro Sato. Star of David Theorem (I). Fibonacci Quarterly, 13(1):70–??, February 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-1/hitotumatu.pdf.

Horadam:1981:ISG

[HS81] A. F. Horadam and A. G. Shannon. Irrational sequence-generated factors of integers. Fibonacci Quarterly, 19(3):240–249, August 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-3/horadam.pdf.

Horadam:1982:CAI

[HS82] A. F. Horadam and A. G. Shannon. Combinatorial aspects of an infinite pattern of integers. Fibonacci Quarterly, 20(1):44–50, February 1982. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/20-1/horadam.pdf.

Hopkins:1984:SIA

[HS84] Glenn Hopkins and William Staton. Some identities arising from the Fibonacci numbers of certain graphs. Fibonacci Quarterly, 22 (3):255–258, August 1984. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/22-3/hopkins.pdf.

Horak:1985:CSO

[HS85] P. Horák and L. Skula. A characterization of the second-order strong divisibility sequences. Fibonacci Quarterly, 23(2):126–132, May 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/23-2/horak.pdf.

Horadam:1988:FLC

[HS88] A. F. Horadam and A. G. Shannon. Fibonacci and Lucas curves. Fibonacci Quarterly, 26(1):3–13, February 1988. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/26-1/horadam1.pdf.

Harminc:1998:PNA

[HS98] Matúš Harminc and Roman Soták. Palindromic numbers in arithmetic progressions. Fibonacci Quarterly, 36(3):259–262, June 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/36-3/harminc.pdf.

Hart:1999:OZD

[HS99] Evelyn Hart and Laura Sanchis. On the occurrence of F_{n} in the Zeckendorf decomposition of nF_{n}. Fibonacci Quarterly, 37(1): 21–32, February 1999. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/37-1/hart.pdf.

Hajdu:2012:GCT

[HS12] L. Hajdu and M. Szikszai. On the GCD-s of k consecutive terms of Lucas sequences. Journal of Number Theory, 132 (12):3056–3069, December 2012. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X12001734.

Hirschhorn:2013:EAR

[HS13] Michael D. Hirschhorn and Vasile Sinescu. Elementary algebra in Ramanujan’s lost notebook. Fibonacci Quarterly, 51(2):123–??, May 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/51-2/hirschhorn1.pdf.

Hajdu:2015:CFS

[HS15] L. Hajdu and M. Szikszai. Common factors in series of consecutive terms of associated Lucas and Lehmer sequences. Fibonacci Quarterly, 53(3):221–??, August 2015. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/53-3/
hajdu.pdf.

Hare:2018:PRF

[HS18] Kevin G. Hare and J. C. Saunders. On (a, b) pairs in random Fibonacci sequences. Journal of Number Theory, 190(??):352–366, September 2018. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X18300891.

Haque:2019:CES

[HS19a] Sajed Haque and Jeffrey Shallit. A class of exponential sequences with shift-invariant discriminators. Fibonacci Quarterly, 57(1): 3–??, February 2019. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/57-1/haque.pdf.

Hinz:2019:DFN

[HS19b] Andreas M. Hinz and Paul K. Stockmeyer. Discovering Fibonacci numbers, Fibonacci words, and a Fibonacci fractal in the Tower of Hanoi. Fibonacci Quarterly, 57(5):72–??, December 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-5/hinz.pdf.

Hare:2022:GFS

[HS22] Kevin Hare and J. C. Saunders. Generalised Fibonacci sequences constructed from balanced words. Journal of Number Theory, 231(??):349–377, February 2022. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X21001931.

Horadam:1994:IDS

[HSF94] A. F. Horadam, B. Swita, and P. Filipponi. Integration and derivative sequences for Pell and Pell–Lucas polynomials. Fibonacci Quarterly, 32(2):130–135, May 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-2/
horadam.pdf.

Hillman:1963:FNZ

[HSG63] A. P. Hillman, M. T. Stroot, and R. M. Grassl. Fibonacci numbers and zigzag Hasse diagrams. Fibonacci Quarterly, 1(3):43–45, October 1963. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/1-3/hillman.pdf.

Hu:2001:RSS

[HSL01] Hong Hu, Zhi-Wei Sun, and Jian-Xin Liu. Reciprocal sums of second-order recurrent sequences. Fibonacci Quarterly, 39(3): 214–220, June 2001. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/39-3/hu.pdf.

He:1997:DZF

[HSR97] M. X. He, D. Simon, and P. E. Ricci. Dynamics of the zeros of Fibonacci polynomials. Fibonacci Quarterly, 35(2):160–168, May 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/35-2/he.pdf.

Hsiao:2003:MSF

[HSS03] Hung-Kuei Hsiao and Shyr-ShenYu. Mapped shuffled Fibonacci languages. Fibonacci Quarterly, 41(5):421–430, November 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-5/hsiao.pdf.

Hsu:1995:NCS

[HsSW95] Leetsch Charles Hsu, Peter Jau shyong Shiue, and Yi Wang. Notes on a conjecture of Singmaster. Fibonacci Quarterly, 33(5): 392–397, November 1995. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/33-5/hsu.pdf.

Hung:1994:USO

[HST94] W. T. Hung, A. G. Shannon, and B. S. Thornton. The use of a second-order recurrence relation in the diagnosis of breast cancer. Fibonacci Quarterly, 32(3):253–259, June 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-3/hung.pdf.

Hsu:1973:NCA

[Hsu73] L. C. Hsu. Note on a combinatorial algebraic identity and its application. Fibonacci Quarterly, 11(5):480–484, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-5/hsu.pdf.

Hsu:1987:GSN

[Hsu87] L. C. Hsu. Generalized Stirling number pairs associated with inverse relations. Fibonacci Quarterly, 25(4):346–351, November 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/25-4/hsu.pdf.

Hsu:1993:LE

[Hsu93a] L. C. Hsu. Letter to the editor. Fibonacci Quarterly, 31(4):294– ??, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/31-4/letter.pdf.

Hsu:1993:SRU

[Hsu93b] L. C. Hsu. A summation rule using Stirling numbers of the second kind. Fibonacci Quarterly, 31(3):256–262, August 1993. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-3/hsu.pdf.

Hsu:1993:FCN

[Hsu93c] Wen-Jing Hsu. Fibonacci cubes: a new interconnection topology. IEEE Transactions on Parallel and Distributed Systems, 4(1):3– 12, January 1993. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Hsu:1995:DOA

[Hsu95] L. C. Hsu. A difference-operational approach to the Möbius inversion formulas. Fibonacci Quarterly, 33(2):169–173, May 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-2/hsu.pdf.

Hsu:1997:KGA

[Hsu97] Leetsch C. Hsu. On a kind of generalized arithmetic-geometric progression. Fibonacci Quarterly, 35(1):62–67, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-1/hsu.pdf.

Horadam:1986:SFE

[HT86] A. F. Horadam and A. P. Treweek. Simson’s formula and an equation of degree 24. Fibonacci Quarterly, 24(4):344–346, November 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/24-4/horadam1.pdf.

Hochwald:1993:RFN

[HT93] Scott H. Hochwald and Jingcheng Tong. On the reciprocals of the Fibonacci numbers. Fibonacci Quarterly, 31(3):246–250, August 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/31-3/hochwald.pdf.

Hsu:2000:RBF

[HT00] Leetsch Charles Hsu and Evelyn L. Tan. A refinement of de Bruyn’s formulas for ∑ a^{k}k^{p}. Fibonacci Quarterly, 38(1):56–59, February 2000. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/38-1/hsu.pdf.

Hopkins:2018:FPR

[HT18] Brian Hopkins and Aram Tangboonduangjit. Fibonacciproducing rational polynomials. Fibonacci Quarterly, 56(4):303– ??, November 2018. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/56-4/hopkins.pdf.

Hopkins:2022:VGAa

[HT22] Brian Hopkins and Aram Tangboonduangjit. Verifying and generalizing Arndt’s compositions. Fibonacci Quarterly, 60(5):181– ??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/60-5/hopkins.pdf.

Hu:2002:LT

[Hu02] Hong Hu. On Lucas v-triangles. Fibonacci Quarterly, 40(4):290– 294, August 2002. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/40-4/hu.pdf.

Hudson:1987:CTD

[Hud87] Richard H. Hudson. Convergence of Tribonacci decimal expansions. Fibonacci Quarterly, 25(2):163–170, May 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-2/hudson.pdf.

Hughes:2010:EAF

[Hug10] Barnabas Hughes. An early abridgement of Fibonacci’s de practica geometrie. Historia Mathematica, 37(4):615–640, November 2010. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0315086009000743.

Hunter:1963:TIR

[Hun63] J. A. H. Hunter. Triangle inscribed in rectangle. Fibonacci Quarterly, 1(3):66–??, October 1963. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/1-3/hunter.
pdf.

Hunter:1964:TVS

[Hun64a] J. A. H. Hunter. Two very special numbers. Fibonacci Quarterly, 2(3):230–??, October 1964. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/2-3/hunter.pdf.

Huntley:1964:FG

[Hun64b] H. E. Huntley. Fibonacci geometry. Fibonacci Quarterly, 2(2): 104–??, April 1964. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/2-2/huntley.pdf.

Huntley:1964:GC

[Hun64c] H. E. Huntley. The golden cuboid. Fibonacci Quarterly, 2(3): 184–??, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/2-3/huntley-a.pdf.

Hunter:1966:FA

[Hun66a] J. A. H. Hunter. Fibonacci again. Fibonacci Quarterly, 4(3): 273–??, October 1966. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/4-3/hunter.pdf.

Hunter:1966:PNE

[Hun66b] J. A. H. Hunter. Perfect number “endings”. Fibonacci Quarterly, 4(1):82–??, February 1966. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/4-1/hunter.pdf.

Hunter:1967:RHT

[Hun67] J. A. H. Hunter. A result for Heronian triangles. Fibonacci Quarterly, 5(5):484–485, December 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-5/
hunter.pdf.

Huntley:1969:FA

[Hun69] H. E. Huntley. Fibonacci and the atom. Fibonacci Quarterly, 7(5):523–524, December 1969. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/7-5/huntley.
pdf.

Hunter:1970:FR

[Hun70a] J. A. H. Hunter. Fibonacci to the rescue. Fibonacci Quarterly, 8(4):406–??, October 1970. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/8-4/hunter-a.pdf.

Huntley:1970:DPS

[Hun70b] H. E. Huntley. The divine proportion: a study in mathematical beauty. Dover, New York, NY, USA, 1970. ISBN 0-486-22254-3. xii + 186 pp. LCCN QA466 .H85 1970.

URL http://www.loc.gov/catdir/description/dover032/
70093195.html;   https://www.math.utah.edu/pub/tex/bib/
fibquart.bib.

Hunter:1972:FOA

[Hun72] J. A. H. Hunter. Fibonacci once again. Fibonacci Quarterly, 10 (2):201–202, February 1972. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/10-2/hunter1.pdf.

Huntley:1974:GE

[Hun74a] H. E. Huntley. The golden ellipse. Fibonacci Quarterly, 12(1): 38–40, February 1974. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/12-1/huntley1.pdf.

Huntley:1974:PAH

[Hun74b] H. E. Huntley. Phi: Another hiding place. Fibonacci Quarterly, 12(1):65–??, February 1974. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/12-1/huntley2.pdf.

Hunter:1978:CPF

[Hun78] J. A. H. Hunter. Congruent primes of form (8r + 1). Fibonacci Quarterly, 16(5):407–410, October 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
5/hunter.pdf.

Hunsucker:1973:CSF

[HW73] J. L. Hunsucker and W. P. Wardlaw. Complete sequences of Fibonacci powers. Fibonacci Quarterly, 11(4):387–394, November 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/11-4/hunsucker.pdf.

Hudson:1981:CCD

[HW81] Richard H. Hudson and C. F. Winans. A complete characterization of the decimal fractions that can be represented as ∑ 10^{−k(i+1)}F_{ai}, where F_{ai} is the ai th Fibonacci number. Fibonacci Quarterly, 19(5):414–421, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-5/hudson.pdf.

Hillar:2008:FIG

[HW09] Christopher J. Hillar and Troels Windfeldt. Fibonacci identities and graph colorings. Fibonacci Quarterly, 46/47(3):220–224, August 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/46_47-3/hillar.pdf.

Hamlin:2012:RPI

[HW12] Nathan Hamlin and William A. Webb. Representing positive integers as a sum of linear recurrence sequences. Fibonacci Quarterly, 50(2):99–??, May 2012. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/50-2/hamlin.
pdf.

Huang:2015:SRW

[HW15] Yuke Huang and Zhiying Wen. The sequence of return words of the Fibonacci sequence. Theoretical Computer Science, 593 (??):106–116, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304397515005022.

Hadjicostas:2017:SSC

[HZ17] Petros Hadjicostas and Lingyun Zhang. Sommerville’s symmetrical cyclic compositions of a positive integer with parts avoiding multiples of an integer. Fibonacci Quarterly, 55(1):54–??, February 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/55-1/hadjicostas.pdf.

Inoue:2011:BFP

[IA11] Kiyoshi Inoue and Sigeo Aki. Bivariate Fibonacci polynomials of order k with statistical applications. Annals of the Institute of Statistical Mathematics, 63(1):197–210, February 2011. CODEN AISXAD. ISSN 0020-3157 (print), 1572-9052 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10463-008-0217-x.

Ipek:2014:HPD

[IA14] Ahmet Ipek and Kamil Ari. On Hessenberg and pentadiagonal determinants related with Fibonacci and Fibonacci-like numbers. Applied Mathematics and Computation, 229(??):433–439, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300313013544.

Iakin:1977:GQQ

[Iak77a] A. L. Iakin. Generalized quaternions with quaternion components. Fibonacci Quarterly, 15(4):350–352, December 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-4/iakin2.pdf.

Iakin:1977:GQH

[Iak77b] I. L. Iakin. Generalized quaternions of higher order. Fibonacci Quarterly, 15(4):343–345, December 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
4/iakin1.pdf.

Iakin:1981:AHO

[Iak81a] A. L. Iakin. Anomalies in higher-order conjugate quaternions: a clarification. Fibonacci Quarterly, 19(4):322–325, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-4/iakin.pdf.

Iakin:1981:EBF

[Iak81b] A. L. Iakin. Extended Binet forms for generalized quaternions of higher order. Fibonacci Quarterly, 19(5):410–413, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-5/iakin.pdf.

Ibstedt:1990:SSL

[Ibs90] Henry Ibstedt. Some sequences of large integers. Fibonacci Quarterly, 28(3):200–203, August 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
3/ibstedt.pdf.

IEEE:1981:PSC

[IEE81] IEEE, editor. Proceedings: 5th Symposium on Computer Arithmetic: May 18–19, 1981, University of Michigan, Ann Arbor, Michigan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1981. LCCN QA 76.6 S985t 1981. IEEE catalog number 81CH1630-C.

IEEE:1984:ASF

[IEE84] IEEE, editor. 25th annual Symposium on Foundations of Computer Science, October 24–26, 1984, Singer Island, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. CODEN ASFPDV. ISBN 0-8186-8591-3, 0-8186-0591-X (paperback), 0-8186-4591-1 (microfiche). ISSN 0272-5428. LCCN QA 76 S979 1984. IEEE catalog no. 84CH2085-9.

IEEE:2008:PAI

[IEE08] IEEE, editor. Proceedings of the 49th Annual IEEE Symposium on Foundations of Computer Science: October 25–23, 2008, Philadelphia, Pennsylvania, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. ISBN 0-7695-3436-8. ISSN 0272-5428. LCCN

QA76 .S95 2008. URL http://ieeexplore.ieee.org/servlet/

opac?punumber=4690923. IEEE Computer Society order number P3436.

Iguchi:1994:CNI

[Igu94] Kazumoto Iguchi. A class of new invariant surfaces under the trace maps for n ary Fibonacci lattices. Journal of Mathematical Physics, 35(2):1008–1019, February 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Indermark:1987:CFL

[IK87] K. Indermark and H. Klaeren. Compiling Fibonacci-like recursion. ACM SIGPLAN Notices, 22(6):101–108, June 1987. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Iannucci:2007:CNF

[IL07] Douglas E. Iannucci and Florian Luca. Catalan numbers, factorials, and sums of aliquot parts. Fibonacci Quarterly, 45(4):327– 336, November 2007. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/45-4/iannucci.pdf.

Ilamed:1971:PMG

[Ila71] Yehiel Ilamed. Powers of a matrix and the generalized Lucas polynomials. Journal of Mathematical Physics, 12(1):113, January 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Imada:1998:SMC

[Ima98] Naotaka Imada. A sparse matrix and the Catalan numbers. Fibonacci Quarterly, 36(1):76–84, February 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-1/imada.pdf.

Iliopoulos:1997:CSF

[IMS97] Costas S. Iliopoulos, Dennis Moore, and W. F. Smyth. A characterization of the squares in a Fibonacci string. Theoretical Computer Science, 172(1–2):281–291, February 10, 1997. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cgi-bin/cas/tree/
store/tcs/cas_sub/browse/browse.cgi?year=1997&volume=
172&issue=1-2&aid=2344.

Ipek:2011:SNC

[İpe11] Ahmet İpek. On the spectral norms of circulant matrices with classical Fibonacci and Lucas numbers entries. Applied Mathematics and Computation, 217(12):6011–6012, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ide:2012:PFS

[IR12] Joshua Ide and Marc S. Renault. Power Fibonacci sequences. Fibonacci Quarterly, 50(2):175–??, May 2012. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/50-
2/ide.pdf.

Ismail:2008:OPG

[Ism09] Mourad E. H. Ismail. One parameter generalizations of the Fibonacci and Lucas numbers. Fibonacci Quarterly, 46/47(2):167– 180, May 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/46_47-2/ismail.pdf.

Imai:1988:EC

[ISTY88] Yasuyuki Imai, Yasuo Seto, Shotaro Tanaka, and Hiroshi Yutani. An expansion of x^{m} and its coefficients. Fibonacci Quarterly, 26 (1):33–39, February 1988. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/26-1/imai.pdf.

Iochum:1991:PLG

[IT91] B. Iochum and D. Testard. Power law growth for the resistance in the Fibonacci model. Journal of Statistical Physics, 65(3–4): 715–723, November 1991. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/BF01053750.

Ivie:1969:MFS

[Ivi69] John Ivie. Multiple Fibonacci sums. Fibonacci Quarterly, 7(3): 303–309, October 1969. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/7-3/ivie.pdf.

Ivie:1972:GM

[Ivi72] John Ivie. A general Q-matrix. Fibonacci Quarterly, 10(3):255– 261, April 1972. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/10-3/ivie-a.pdf.

Iyer:1969:IIG

[Iye69a] Muthulakshmi R. Iyer. Identities involving generalized Fibonacci numbers. Fibonacci Quarterly, 7(1):66–72, February 1969. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-1/iyer1.pdf.

Iyer:1969:NFQ

[Iye69b] Muthulakshmi R. Iyer. A note on Fibonacci quaternions. Fibonacci Quarterly, 7(3):225–229, October 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/7-3/iyer.pdf.

Iyer:1969:SRF

[Iye69c] Muthulakshmi R. Iyer. Some results on Fibonacci quaternions. Fibonacci Quarterly, 7(2):201–210, April 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/7-2/iyer.pdf.

Iyer:1969:SIF

[Iye69d] Muthulakshmi R. Iyer. Sums involving Fibonacci numbers. Fibonacci Quarterly, 7(1):92–98, February 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-1/
iyer2.pdf.

Izotov:1995:NSN

[Izo95] Anatoly S. Izotov. A note on Sierpinski numbers. Fibonacci Quarterly, 33(3):206–207, June 1995. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/33-3/izotov.
pdf.

Izotov:1999:FSM

[Izo99] Anatoly S. Izotov. On the form of solutions of Martin Davis’ Diophantine equation. Fibonacci Quarterly, 37(3):258–261, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/37-3/izotov.pdf.

Izotov:2002:SOL

[Izo02] Anatoly S. Izotov. Second-order linear recurrences of composite numbers. Fibonacci Quarterly, 40(3):266–268, June 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-3/izotov.pdf.

Izotov:2005:PD

[Izo05] Anatoly S. Izotov. On the prime divisors of GCD(3^{n} − 2, 2^{n} − 3). Fibonacci Quarterly, 43(2):130–131, May 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers/43-2/paper43-2-6.pdf.

Jacobson:1989:AUD

[Jac89] Eliot Jacobson. Almost uniform distribution of the Fibonacci sequence. Fibonacci Quarterly, 27(4):335–337, August 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-4/jacobson.pdf.

Jacobson:1990:DRT

[Jac90] Eliot Jacobson. The distribution of residues of two-term recurrence sequences. Fibonacci Quarterly, 28(3):227–229, August 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/28-3/jacobson.pdf.

Jacobson:1992:DFN

[Jac92] Eliot T. Jacobson. Distribution of the Fibonacci numbers mod 2K. Fibonacci Quarterly, 30(3):211–215, August 1992. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/30-3/jacobson.pdf.

Jagannathan:2021:FQC

[Jag21] Anuradha Jagannathan. The Fibonacci quasicrystal: Case study of hidden dimensions and multifractality. Reviews of Modern Physics, 93(4):045001–??, April 2021. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic),

1539-0756. URL http://journals.aps.org/rmp/abstract/
10.1103/RevModPhys.93.045001.

Jaiswal:1969:DIG

[Jai69] D. V. Jaiswal. On determinants involving generalized Fibonacci numbers. Fibonacci Quarterly, 7(3):319–330, October 1969. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-3/jaiswal.pdf.

Jaiswal:1974:PRT

[Jai74a] D. V. Jaiswal. On polynomials related to Tchebichef polynomials of the second kind. Fibonacci Quarterly, 12(3):263–264, October 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/12-3/jaiswal.pdf.

Jaiswal:1974:SGP

[Jai74b] D. V. Jaiswal. Some geometrical properties of the generalized Fibonacci sequences. Fibonacci Quarterly, 12(1):67–70, February 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/12-1/jaiswal.pdf.

Jameson:1977:LE

[Jam77] John W. Jameson. Letter to the editor. Fibonacci Quarterly, 15 (1):8–??, February 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-1/letter-a.pdf.

Jamieson:1990:NFN

[Jam90] M. J. Jamieson. Notes: Fibonacci numbers and Aitken sequences revisited. American Mathematical Monthly, 97(9):829– 831, November 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jamieson:2002:FNC

[Jam02] M. J. Jamieson. Fibonacci numbers and cotangent sequences. American Mathematical Monthly, 109(7):655–657, August/ September 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/
3072430.

Jameson:2018:FPM

[Jam18] G. J. O. Jameson. Fibonacci periods and multiples. The Mathematical Gazette, 101(553):63–76, March 2018. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/fibonacci-periods-and-multiples/1CD89C358C7AD0CE86005E0BF340F766.

Janke:2002:PRN

[Jan02] Wolfhard Janke. Pseudo random numbers: Generation and quality checks. In Johannes Grotendorst, Dominik Marx, and Alejandro Muramatsu, editors, Quantum Simulations of Complex Many-Body Systems: From Theory to Algorithms: Winter School, 25 February–1 March 2002, Rolduc Conference Centre, Kerkrade, The Netherlands, volume 10, pages 447–?? John von Neumann Institute for Computing, Jülich, Germany, 2002. ISBN 3-00-009058-4. LCCN Q183.9 N492 v. 11. URL http://www2.

fz-juelich.de/nic-series/volume10/janke1.pdf.

Jarden:1946:TTF

[Jar46] Dov Jarden. Two theorems on Fibonacci’s sequence. American Mathematical Monthly, 53(8):425–427, October 1946. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jarden:1963:GPD

[Jar63a] Dov Jarden. On the greatest primitive divisors of Fibonacci and Lucas numbers with prime-power subscripts. Fibonacci Quarterly, 1(3):15–20, October 1963. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/1-3/jarden.
pdf.

Jarden:1963:PLD

[Jar63b] Dov Jarden. On the periodicity of the last digits of the Fibonacci numbers. Fibonacci Quarterly, 1(4):21–22, December 1963. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-4/jarden.pdf.

Jarden:1964:SIF

[Jar64] Dov Jarden. Strengthened inequalities for Fibonacci and Lucas numbers. Fibonacci Quarterly, 2(1):45–48, February 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-1/jarden.pdf.

Jarden:1966:RSC

[Jar66] Dov Jarden. Recurring sequences; a collection of papers. Riveon Lematematika, Jerusalem, Israel, second edition, 1966. 137 pp. LCCN QA246.5 .J3 1966. Il18.00. Includes numerous new factorizations of Fibonacci and Lucas numbers, by John Brillhart.

Jarden:1967:EAL

[Jar67a] Dov Jarden. Existence of arbitrarily long sequences of consecutive members in arithmetic progressions divisible by arbitrarily many different primes. Fibonacci Quarterly, 5(3):280–??, October 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-3/jarden1.pdf.

Jarden:1967:NIF

[Jar67b] Dov Jarden. A new important formula for Lucas numbers. Fibonacci Quarterly, 5(4):346–??, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/5-4/jarden.pdf.

Jarden:1968:LNP

[Jar68a] Dov Jarden. Any Lucas number L_{5p}, for any prime p > 5, has at least two distinct primitive prime divisors. Fibonacci Quarterly, 6 (6):407–??, December 1968. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/6-6/jarden2.pdf.

Jarden:1968:FDS

[Jar68b] Dov Jarden. Formulas for decomposing F_{3n}/F_{n}, F_{5n}/F_{n} and L_{5n}/L_{n} into a sum of difference of two squares. Fibonacci Quarterly, 6(1):96–??, February 1968. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/6-1/jarden.
pdf.

Jaroma:2006:PTP

[Jar06] John H. Jaroma. On primes and terms of prime or 2^{k} index in the Lehmer sequences. Fibonacci Quarterly, 44(3):202–208, August 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/44-3/jaroma.pdf.

Javaheri:2022:UBR

[Jav22] Mohammad Javaheri. Upper bound Residues of the Fibonacci sequence modulo primes. Fibonacci Quarterly, 60(1):48–??, February 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/60-1/javaheri.pdf.

Johnson:2017:ATO

[JC17] Virginia P. Johnson and Charles K. Cook. Areas of triangles and other polygons with vertices from various sequences. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/55-
5/johnson.pdf.

Javaheri:2022:DFN

[JC22] Mohammad Javaheri and Stephanie Cambrea. The distribution of Fibonacci numbers modulo primes. American Mathematical Monthly, 129(1):75–79, 2022. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jin:2022:TIH

[JD22] Ziqian (Alexa) Jin and Greg Dresden. Tetranacci identities via hexagonal tilings. Fibonacci Quarterly, 60(2):99–??, May 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-2/jin.pdf.

Jeffrey:2008:PAM

[Jef08] David Jeffrey, editor. Proceedings of the 21st annual meeting of the International Symposium on Symbolic Computation, ISSAC 2008, July 20–23, 2008, Hagenberg, Austria. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-59593-904-0. LCCN ????

Jentsch:1966:PDE

[Jen66] W. Jentsch. On a partial difference equation of L. Carlitz. Fibonacci Quarterly, 4(3):202–207, October 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/4-3/jentsch.pdf.

Jentsch:1971:IVP

[Jen71] W. Jentsch. On an initial-value problem for linear partial difference equations. Fibonacci Quarterly, 9(3):313–323, May 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-3/jentsch.pdf.

Jennings:1993:SPI

[Jen93] Derek Jennings. Some polynomial identities for the Fibonacci and Lucas numbers. Fibonacci Quarterly, 31(2):134–137, May 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/31-2/jennings.pdf.

Jennings:1994:SRF

[Jen94] Derek Jennings. On sums of reciprocals of Fibonacci and Lucas numbers. Fibonacci Quarterly, 32(1):18–21, February 1994.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-1/jennings.pdf.

Jenkins:1997:AAH

[Jen97] Bob Jenkins. Algorithm alley: Hash functions. Dr. Dobb’s Journal of Software Tools, 22(9):107–109, 115–116, September 1997. CODEN DDJOEB. ISSN 1044-789X. Describes a new hash function which is much better at producing uniform key distributions than others commonly used, yet remains acceptably fast. See [Boy98] for comparison with a related algorithm.

Jeske:1963:LRRb

[Jes63a] James A. Jeske. Linear recurrence relations — Part III. Fibonacci Quarterly, 1(4):35–40, December 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-4/
jeske-a.pdf.

Jeske:1963:LRRa

[Jes63b] James A. Jeske. Linear recurrence relations, Part I. Fibonacci Quarterly, 1(2):69–73, April 1963. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/1-2/jeske.pdf.

Jeske:1964:LRR

[Jes64] J. A. Jeske. Linear recurrence relations — Part III. Fibonacci Quarterly, 2(3):197–202, October 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-3/
jeske.pdf.

Junge:1973:PAR

[JH73] Bjarne Junge and V. E. Hoggatt, Jr. Polynomials arising from reflections across multiple plates. Fibonacci Quarterly, 11(3):285– 291, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/11-3/junge.pdf.

Jianguo:2005:RON

[JH05] Xia Jianguo and Qin Hourong. On reverse order numbers of certain sequences and the Jacobi symbol. Fibonacci Quarterly, 43(4): 351–358, November 2005. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/43-4/xia.pdf.

Jiang:2015:EIM

[JH15] Xiaoyu Jiang and Kicheon Hong. Explicit inverse matrices of Tribonacci skew circulant type matrices. Applied Mathematics and Computation, 268(??):93–102, October 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300315007341.

Jiang:1997:EGF

[JHK97] Feng-Shr Jiang, Shi-Jinn Horng, and Tzong-Wann Kao. Embedding of generalized Fibonacci cubes in hypercubes with faulty nodes. IEEE Transactions on Parallel and Distributed Systems, 8(7):727–737, July 1997. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://dlib.

computer.org/td/books/td1997/pdf/l0727.pdf;   http://
www.computer.org/tpds/td1997/l0727abs.htm.

Jones:1993:PWC

[JK93] J. P. Jones and Peter Kiss. On points whose coordinates are terms of a linear recurrence. Fibonacci Quarterly, 31(3):239–245, August 1993. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/31-3/jones.pdf.

Johnson:1994:CFA

[JK94] D. L. Johnson and A. C. Kim. Cyclic Fibonacci algebras. Fibonacci Quarterly, 32(5):441–444, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-5/johnson.pdf.

Jaroszewski:1997:DMM

[JK97] I. Jaroszewski and A. K. Kwaşniewski. Dynamics of the Möbius mapping and Fibonacci-like sequences. Fibonacci Quarterly, 35(3):258–264, August 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-
3/jaroszewski.pdf.

Jaroszewski:1998:SEP

[JK98] I. Jaroszewski and A. K. Kwaşniewski. Some extensions of properties of the sequence of reciprocal Fibonacci polynomials. Fibonacci Quarterly, 36(4):348–353, August 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-4/jaroszewski.pdf.

Jia:2007:AGF

[JLW07] C. Z. Jia, H. M. Liu, and T. M. Wang. q-analogs of generalized Fibonacci and Lucas polynomials. Fibonacci Quarterly, 45(1):

26–34, February 2007. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/45-1/jia.pdf.

Jones:2014:GCS

[JM14] Lenny Jones and Maria Markovich. Generating composite sequences by appending digits to special types of integers. Fibonacci Quarterly, 52(2):148–??, May 2014. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/52-
2/jones.pdf.

Jahannia:2022:CZL

[JMNRS22] Marieh Jahannia, Morteza Mohammad-Noori, Narad Rampersad, and Manon Stipulanti. Closed Ziv–Lempel factorization of the m-bonacci words. Theoretical Computer Science, 918(??):32– 47, May 29, 2022. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397522001621.

Jones:2024:SSS

[JO24] Michael A. Jones and Brittany C. Ohlinger. Self-similar structure of P -positions of the game Euclid. Fibonacci Quarterly, 62(1): 15–??, February 2024. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/62-1/jones.pdf.

Johnston:1940:FSA

[Joh40] L. S. Johnston. The Fibonacci sequence and allied trigonometric identities. American Mathematical Monthly, 47(2):85–89, February 1940. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Johnson:1990:HP

[Joh90] Karen Anne Johnson. A hypercube problem. Fibonacci Quarterly, 28(2):121–128, May 1990. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/28-2/johnson.
pdf.

Johnson:2004:TFS

[Joh04] L. F. Johnson. Tumble, a fast simple iteration algorithm for Fibonacci. Information Processing Letters, 89(4):187–189, February 28, 2004. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Johnson:2008:RTI

[Joh09] Marjorie Bicknell Johnson. A report on the Thirteenth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 46/47(1):5–6, February 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/46_47-1/Johnson_Report_11-08.pdf.

Johnson:2012:FQF

[Joh12a] Marjorie Bicknell Johnson. The Fibonacci Quarterly: Fifty years. Fibonacci Quarterly, 50(4):290–??, November 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Announcements/JohnsonFiftyYears.pdf.

Johnson:2012:RFI

[Joh12b] Marjorie Bicknell Johnson. A report on the Fifteenth International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 50(3):194–??, August 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Announcements/ConferenceReport15.pdf.

Johnson:2020:RIC

[Joh20] Marjorie Bicknell Johnson. A report on the 19th International Conference on Fibonacci Numbers and Their Applications. Fibonacci Quarterly, 58(4):290–??, November 2020. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Announcements/ConferenceReport08082020.pdf.

Johnson:2022:MFH

[Joh22a] Marjorie Bicknell Johnson. In memoriam — Fredric T. Howard. Fibonacci Quarterly, 60(4):290–??, November 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Announcements/fredhoward.pdf.

Johnson:2022:RIC

[Joh22b] Marjorie Bicknell Johnson. A report on the 20th international conference on Fibonacci numbers and their applications. Fibonacci Quarterly, 60(4):291–??, November 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Announcements/ConferenceReport08032022.pdf.

Jones:1975:DRF

[Jon75] James P. Jones. Diophantine representation of the Fibonacci numbers. Fibonacci Quarterly, 13(1):84–88, February 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-1/jones.pdf.

Jones:1976:OPT

[Jon76a] Bob Jones. Ode to Pascal’s triangle. Fibonacci Quarterly, 14 (5):452–??, December 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-5/jones.pdf.

Jones:1976:DRL

[Jon76b] James P. Jones. Diophantine representation of the Lucas numbers. Fibonacci Quarterly, 14(2):134–??, April 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-2/jones.pdf.

Jones:1978:SVP

[Jon78] Burton W. Jones. A second variation on a problem of Diophantus and Davenport. Fibonacci Quarterly, 16(2):155–165, April 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-2/jones.pdf.

Jones:1990:EXK

[Jon90] Patricia Jones. On the equation ϕ (x) + ϕ (k) = ϕ (x+k). Fibonacci Quarterly, 28(2):162–165, May 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-2/
jones.pdf.

Jones:1991:CPR

[Jon91a] Dixon J. Jones. Continued powers and roots. Fibonacci Quarterly, 29(1):37–46, February 1991. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/29-1/jones.
pdf.

Jones:1991:P

[Jon91b] Patricia Jones. ϕ-partitions. Fibonacci Quarterly, 29(4):347–350, November 1991. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/29-4/jones.pdf.

Jones:1996:GHS

[Jon96] Charles H. Jones. Generalized hockey stick identities and Ndimensional blockwalking. Fibonacci Quarterly, 34(3):280–289, June 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/34-3/jones.pdf.

Jordan:1964:FTC

[Jor64] J. H. Jordan. A Fibonacci test for convergence. Fibonacci Quarterly, 2(1):39–41, February 1964. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/2-1/jordan.
pdf.

Jordan:1965:GFL

[Jor65] J. H. Jordan. Gaussian Fibonacci and Lucas numbers. Fibonacci Quarterly, 3(4):315–318, December 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-4/
jordan.pdf.

Joseph:1979:MCT

[Jos79] James E. Joseph. Maximum cardinalities for topologies on finite sets. Fibonacci Quarterly, 17(2):97–102, April 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-2/joseph.pdf.

Josef:1983:LT

[Jos83] Sána Josef. Lucas triangle. Fibonacci Quarterly, 21(3):192–195, August 1983. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/21-3/josef.pdf.

Joseph:1998:CEF

[Jos98] John A. Joseph. A chaotic extension of the 3x + 1 function to Z_{2}[i]. Fibonacci Quarterly, 36(4):309–316, August 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-4/joseph.pdf.

Jaidee:2019:AFF

[JP19] Montree Jaidee and Prapanpong Pongsriiam. Arithmetic functions of Fibonacci and Lucas numbers. Fibonacci Quarterly, 57 (3):246–??, August 2019. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/57-3/jaidee.pdf.

Joye:1997:PFR

[JQ97] Marc Joye and Jean-Jacques Quisquater. Protocol failures for RSA-like functions using Lucas sequences and elliptic curves. Lecture Notes in Computer Science, 1189:93–100, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1998:MRF

[JR98] Erik Jensen and Ronald Rusay. Musical representations of the Fibonacci string and proteins. Mathematica Journal, 8(2):??, ???? 1998. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic).

Janvresse:2008:HDR

[JRdlR08] Élise Janvresse, Benoît Rittaud, and Thierry de la Rue. How do random Fibonacci sequences grow? Probability Theory and Related Fields, 142(3–4):619–648, November 2008. CODEN PTR-FEU. ISSN 0178-8051 (print), 1432-2064 (electronic). URL http:

//link.springer.com/article/10.1007/s00440-007-0117-7.

Janvresse:2010:ASG

[JRdlR10] Élise Janvresse, Benoît Rittaud, and Thierry de la Rue. Almostsure growth rate of generalized random Fibonacci sequences. Annales de l’Institut Henri Poincaré. Probabilités et Statistiques, 46(1):135–158, February 2010. CODEN AHPBAR. ISSN 0246-0203 (print), 1778-7017 (electronic). URL http:

//projecteuclid.org/euclid.aihp/1267454112;   http://
www.numdam.org/item?id=AIHPB_2010__46_1_135_0.

Johnson:2016:RPT

[JS16] Keith Johnson and Kira Scheibelhut. Rational polynomials that take integer values at the Fibonacci numbers. American Mathematical Monthly, 123(4):338–346, April 2016. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/10.4169/amer.
math.monthly.123.4.338.

Jones:2019:PSB

[JS19] Lenny Jones and Lawrence Somer. Primefree shifted binary linear recurrence sequences. Fibonacci Quarterly, 57(1):51–??, February 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/57-1/jones.pdf.

Juncosa:1961:LEC

[Jun61] M. L. Juncosa. Letter to the Editor: concerning Ferguson’s paper on Fibonaccian searching. Communications of the Association for Computing Machinery, 4(6):252, June 1961. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Juarna:2008:CIB

[JV08] A. Juarna and V. Vajnovszki. Combinatorial isomorphism between Fibonacci classes. Journal of Discrete Mathematical Sciences and Cryptography, 11(2):147–158, April 2008. CODEN ????

ISSN 0972-0529. URL http://www.connectjournals.com/
achivestoc.php?bookmark=CJ-003072&volume=11&issue_id=
02.

Jiu:2019:CCH

[JV19] Lin Jiu and Christophe Vignat. Connection coefficients for higher-order Bernoulli and Euler polynomials: A random walk approach. Fibonacci Quarterly, 57(5):84–??, December 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-5/jiu.pdf.

Karaduman:2003:GSF

[KA03] Erdal Karaduman and Hüseyin Aydın. General 2-step Fibonacci sequences in nilpotent groups of exponent p and nilpotency class 4. Applied Mathematics and Computation, 141(2–3):491–497, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kilic:2013:MIS

[KA13] Emrah Kiliç and Talha Arikan. More on the infinite sum of reciprocal Fibonacci, Pell and higher order recurrences. Applied Mathematics and Computation, 219(14):7783–7788, March 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300313001197.

Kilic:2019:NGF

[KA19] Emrah Kiliç and Talha Arikan. A nonlinear generalization of the Filbert matrix and its Lucas analogue. Linear Multilinear Algebra, 67(1):141–157, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kahan:1980:MCS

[Kah80] Steven Kahan. Mutually counting sequences. Fibonacci Quarterly, 18(1):47–50, February 1980. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/18-1/kahan.
pdf.

Kahan:1987:CCT

[Kah87] Steven Kahan. Cyclic counting trios. Fibonacci Quarterly, 25(1): 11–20, February 1987. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/25-1/kahan.pdf.

Kahya:2005:CTF

[Kah05] Emin Kahya. Comment on titled “On Fibonacci search method with k-Lucas numbers”. Applied Mathematics and Computation, 162(3):1321–1324, March 25, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [YK03].

Kahya:2006:CTI

[Kah06] Emin Kahya. Comment on titled “An improvement on Fibonacci search method in optimization theory”. Applied Mathematics and Computation, 173(2):753–756, February 15, 2006. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [SYY04].

Kahane:2008:NRI

[Kah09] Charles S. Kahane. A note on Ramus’ identity and associated recursion relations. Fibonacci Quarterly, 46/47(1):48–52, February 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/46_47-1/kahane.pdf.

Kainen:2022:FSC

[Kai22] Paul C. Kainen. Fibonacci in Somos-5 by complexification. Fibonacci Quarterly, 60(4):362–??, November 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/60-4/kainen.pdf.

Kalman:1982:GFN

[Kal82] Dan Kalman. Generalized Fibonacci numbers by matrix methods. Fibonacci Quarterly, 20(1):73–76, February 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-1/kalman.pdf.

Kalman:1990:SPM

[Kal90] Dan Kalman. Sums of powers by matrix methods. Fibonacci Quarterly, 28(1):60–71, February 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
1/kalman.pdf.

Kanovich:2014:MRF

[Kan14] Max Kanovich. Multiset rewriting over Fibonacci and Tribonacci numbers. Journal of Computer and System Sciences, 80(6):1138–1151, September 2014. CODEN JCSSBM. ISSN 0022-0000 (print), 1090-2724 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022000014000506.

Kaneda:2015:TRC

[Kan15] Masayoshi Kaneda. Two remarks on the Collatz cycle conjecture. Fibonacci Quarterly, 53(2):168–??, May 2015. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/53-
2/kaneda.pdf.

Kaplansky:1945:LTM

[Kap45] Irving Kaplansky. Lucas’s tests for Mersenne numbers. American Mathematical Monthly, 52(4):188–190, April 1945. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Kilic:2010:PCM

[KAP10] Emrah Kilic, Ilker Akkus, and Helmut Prodinger. A proof of a conjecture of Melham. Fibonacci Quarterly, 48(3):241–248, August 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/48-3/kilic.pdf.

Karchmar:1965:P

[Kar65] E. J. Karchmar. Phyllotaxis. Fibonacci Quarterly, 3(1):64–65, February 1965. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/3-1/karchmar.pdf.

Karst:1969:FSI

[Kar69] Edgar Karst. A four-step iteration algorithm to generate x^{2} in x^{2} + (x + 1)^{2} = y^{2}. Fibonacci Quarterly, 7(2):180–??, April 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-2/karst.pdf.

Karst:1970:MSC

[Kar70] Edgar Karst. Magic squares consisting of primes in A. P. Fibonacci Quarterly, 8(3):317–324, April 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-3/
karst.pdf.

Karst:1972:MAM

[Kar72] Edgar Karst. More about magic squares consisting of different primes. Fibonacci Quarterly, 10(6):651–655, December 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-6/karst.pdf.

Karst:1974:IAC

[Kar74] Edgar Karst. Iteration algorithms for certain sums of squares. Fibonacci Quarterly, 12(1):83–86, February 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/12-1/karst.pdf.

Karachik:1996:LMP

[Kar96] Valery V. Karachik. P -Latin matrices and Pascal’s triangle modulo a prime. Fibonacci Quarterly, 34(4):362–372, August 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-4/karachik.pdf.

Karaduman:2004:AFN

[Kar04a] Erdal Karaduman. An application of Fibonacci numbers in matrices. Applied Mathematics and Computation, 147(3):903–908, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Karttunen:2004:PTM

[Kar04b] Antti Karttunen. On Pascal’s triangle modulo 2 in Fibonacci representation. Fibonacci Quarterly, 42(1):38–46, February 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-1/quartkarttunen01_2004.pdf.

Karaduman:2005:DMG

[Kar05] Erdal Karaduman. On determinants of matrices with general Fibonacci numbers entries. Applied Mathematics and Computation, 167(1):670–676, August 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kravitz:1964:LTM

[KB64] Sidney Kravitz and Murray Berg. Lucas’ test for Mersenne numbers, 6000 < p < 7000 (in Technical Notes and Short Papers). Mathematics of Computation, 18(85):148–149, January 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Klein:2008:UFC

[KBN08] S. T. Klein and M. K. Ben-Nissan. Using Fibonacci compression codes as alternatives to dense codes. In Storer and Marcellin [SM08], pages 472–481. ISBN 0-7695-3121-0. ISSN 1068-0314. LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=4483325.

Klein:2010:UFC

[KBN10] Shmuel T. Klein and Miri Kopel Ben-Nissan. On the usefulness of Fibonacci compression codes. The Computer Journal, 53(6):701–716, July 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/cgi/content/abstract/53/6/
701;   http://comjnl.oxfordjournals.org/cgi/reprint/53/
6/701.

Krolak:1963:EFS

[KC63] P. Krolak and L. Cooper. An extension of Fibonaccian search to several variables. Communications of the Association for Computing Machinery, 6(10):639–641, October 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Katz:1986:FDP

[KC86] Talbot M. Katz and Daniel I. A. Cohen. The first digit property for exponential sequences is independent of the underlying distribution. Fibonacci Quarterly, 24(1):2–7, February 1986. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-1/katz.pdf.

Kennedy:1989:NR

[KC89] Robert E. Kennedy and Curtis N. Cooper. Niven repunits and 10^{n} ≡ 1 (mod n). Fibonacci Quarterly, 27(2):139–143, May 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/27-2/kennedy.pdf.

Kennedy:1991:ETC

[KC91a] Robert E. Kennedy and Curtis N. Cooper. An extension of a theorem by Cheo and Yien concerning digital sums. Fibonacci Quarterly, 29(2):145–149, May 1991. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/29-2/kennedy.
pdf.

Kennedy:1991:SSS

[KC91b] Robert E. Kennedy and Curtis N. Cooper. The statistics of the smallest space on a lottery ticket. Fibonacci Quarterly, 29(4): 367–370, November 1991. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/29-4/kennedy.pdf.

Kennedy:1993:SPD

[KC93] Robert E. Kennedy and Curtis Cooper. Sums of powers of digital sums. Fibonacci Quarterly, 31(4):341–345, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-4/kennedy.pdf.

Knuth:2004:FCC

[KC04] Donald E. Knuth and Robin Chapman. Fibonacci in complex camouflage: 10858. American Mathematical Monthly, 111(2): 166–167, February 2004. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/ stable/4145227. See errata [Ano04a].

Kimberling:2014:PP

[KE14] Clark Kimberling and Problem Editor. Problem proposals. Fibonacci Quarterly, 52(5):5–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers1/52-5/Problems.pdf.

Kelisky:1957:FIB

[Kel57] Richard P. Kelisky. On formulas involving both the Bernoulli and Fibonacci numbers. Scripta Mathematica, 23(??):27–35, ???? 1957. ISSN 0036-9713.

Kelisky:1958:FIB

[Kel58] Richard P. Kelisky. On formulas involving both the Bernoulli and Fibonacci numbers. Scripta Mathematica, 23(??):27–35, ???? 1958. ISSN 0036-9713.

Kelisky:1965:CEA

[Kel65] R. P. Kelisky. Concerning the Euclidean algorithm. Fibonacci Quarterly, 3(3):219–223, October 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-3/
kelisky.pdf.

Keller:1972:GZT

[Kel72] Timothy J. Keller. Generalizations of Zeckendorf’s theorem. Fibonacci Quarterly, 10(1):95–102, January 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-1/keller-a.pdf.

Kelly:1992:SFF

[Kel92] John B. Kelly. Schur functions and Fibonacci identities. Fibonacci Quarterly, 30(2):148–156, May 1992. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/30-2/
kelly.pdf.

Kelsey:2008:DFS

[Kel08] Michael A. Kelsey. Dethroning Fibonacci sequence (abstract only). ACM Communications in Computer Algebra, 42(1–2):78, March/June 2008. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Kerr:1982:EOL

[Ker82] John Kerr. The existence of K orthogonal Latin K-cubes of order 6. Fibonacci Quarterly, 20(4):360–362, November 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-4/kerr.pdf.

Klamkin:1973:RP

[KF73] Murray S. Klamkin and R. S. Fishman. A reliability problem. Fibonacci Quarterly, 11(2):169–173, April 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-2/
klamkin2.pdf.

Klostermeyer:2002:NPF

[KG02] W. F. Klostermeyer and J. L. Goldwasser. Nullspace-primes and Fibonacci polynomials. Fibonacci Quarterly, 40(4):323–327, August 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/40-4/klostermeyer.pdf.

Koshy:2017:PEDa

[KG17a] Thomas Koshy and Zhenguang Gao. Polynomial extensions of a Diminnie delight. Fibonacci Quarterly, 55(1):13–??, February 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/55-1/koshy.pdf.

Koshy:2017:PEDb

[KG17b] Thomas Koshy and Zhenguang Gao. Polynomial extensions of a Diminnie delight revisited: Part I. Fibonacci Quarterly, 55(4): 320–??, November 2017. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/55-4/koshy.pdf.

Koshy:2017:TWJ

[KG17c] Thomas Koshy and Ralph P. Grimaldi. Ternary words and Jacobsthal numbers. Fibonacci Quarterly, 55(2):129–??, May 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-2/koshy-grimaldi.pdf.

Koshy:2018:PED

[KG18a] Thomas Koshy and Zhenguang Gao. Polynomial extensions of a Diminnie delight revisited: Part II. Fibonacci Quarterly, 56(1): 10–??, February 2018. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/56-1/koshy.pdf.

Koshy:2018:SGC

[KG18b] Thomas Koshy and Martin Griffiths. Some Gibonacci convolutions with dividends. Fibonacci Quarterly, 56(3):237–??, August 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/56-3/koshy2.pdf.

Koshy:2019:POD

[KG19] Thomas Koshy and Zhenguang Gao. Periodicity of ones digits in Jacobsthal numbers with triangular and Jacobsthal subscripts. Fibonacci Quarterly, 57(4):322–??, November 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-4/koshy2.pdf.

Koshy:2020:EGSb

[KG20] Thomas Koshy and Zhenguang Gao. Extended Gibonacci sums of polynomial products of order 3 revisited. Fibonacci Quarterly, 58 (4):291–??, November 2020. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/58-4/koshy.pdf.

Koshy:2024:SICb

[KG24] Thomas Koshy and Zhenguang Gao. Sums involving a class of Jacobsthal polynomial squares. Fibonacci Quarterly, 62(1):40– ??, February 2024. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/62-1/koshy3.pdf.

Kramer:1972:SCF

[KH72] Judy Kramer and V. E. Hoggatt, Jr. Special cases of Fibonacci periodicity. Fibonacci Quarterly, 10(5):519–522, November 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-5/kramer-a.pdf.

Kidwell:1982:VFP

[KH82] Mark E. Kidwell and Clifford M. Hurvich. A variant of the Fibonacci polynomials which arises in the Gambler’s ruin problem. Fibonacci Quarterly, 20(1):66–72, February 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-1/kidwell.pdf.

Khan:1981:SDF

[Kha81] Rasul A. Khan. A simple derivation of a formula for ∑_{k=1}^{n} k^{r}. Fibonacci Quarterly, 19(2):177–179, April 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-2/
khan.pdf.

Kilic:2010:CBS

[KI10] Emrah Kilic and Eugen J. Ionascu. Certain binomial sums with recursive coefficients. Fibonacci Quarterly, 48(2):161–167, May 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-2/kilic.pdf.

Kara:2016:SPG

[KI16] Emrah Evren Kara and Merve Ilkhan. Some properties of generalized Fibonacci sequence spaces. Linear Multilinear Algebra, 64 (11):2208–??, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Kihel:2000:ES

[Kih00] Omar Kihel. On the extendibility of the set 1, 2, 5. Fibonacci Quarterly, 38(5):464–466, November 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-
5/kihel.pdf.

Killgrove:1976:STP

[Kil76] R. B. Killgrove. The sum of two powers is a third, sometimes. Fibonacci Quarterly, 14(3):206–209, October 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-3/killgrove.pdf.

Kilic:2007:GOF

[Kil07] Emrah Kilic. The generalized order-k Fibonacci–Pell sequence by matrix methods. Journal of Computational and Applied Mathematics, 209(2):133–145, December 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0377042706006698.

Kimberling:1976:DPR

[Kim76] Clark Kimberling. Divisibility properties of recurrent sequences. Fibonacci Quarterly, 14(4):369–376, November 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-4/kimberling.pdf.

Kimberling:1978:SDS

[Kim78] Clark Kimberling. Strong divisibility sequences with nonzero initial term. Fibonacci Quarterly, 16(6):541–543, December 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-6/kimberling.pdf.

Kimberling:1979:GCD

[Kim79a] Clark Kimberling. Greatest common divisors of sums and differences of Fibonacci, Lucas and Chebyshev polynomials. Fibonacci Quarterly, 17(1):18–22, February 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
1/kimberling2.pdf.

Kimberling:1979:SDS

[Kim79b] Clark Kimberling. Strong divisibility sequences and some conjectures. Fibonacci Quarterly, 17(1):13–17, February 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-1/kimberling1.pdf.

Kimberling:1980:FCI

[Kim80a] Clark Kimberling. Four composition identities for Chebyshev polynomials. Fibonacci Quarterly, 18(4):353–369, December 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/18-4/kimberling2.pdf.

Kimberling:1980:GCP

[Kim80b] Clark Kimberling. Generalized cyclotomic polynomials, Fibonacci cyclotomic polynomials, and Lucas cyclotomic polynomials. Fibonacci Quarterly, 18(2):108–125, April 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-2/kimberling.pdf.

Kimberling:1980:GFL

[Kim80c] Clark Kimberling. Generating functions of linear divisibility sequences. Fibonacci Quarterly, 18(3):193–207, October 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-3/kimberling.pdf.

Kimberling:1980:MPM

[Kim80d] Clark Kimberling. Mixing properties of mixed Chebyshev polynomials. Fibonacci Quarterly, 18(4):334–340, December 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-4/kimberling1.pdf.

Kimberling:1981:AAS

[Kim81a] Clark Kimberling. Almost arithmetic sequences and complementary systems. Fibonacci Quarterly, 19(5):426–432, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-5/kimberling.pdf.

Kimberling:1981:FPA

[Kim81b] Clark Kimberling. Fractional parts (nr − s), almost arithmetic sequences, and Fibonacci numbers. Fibonacci Quarterly, 19(3): 280–284, August 1981. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/19-3/kimberling.pdf.

Kimberling:1983:OFZ

[Kim83] Clark Kimberling. One-free Zeckendorf sums. Fibonacci Quarterly, 21(1):53–57, February 1983. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/21-
1/kimberling.pdf.

Kimberling:1990:FH

[Kim90] Clark Kimberling. Fibonacci hyperbolas. Fibonacci Quarterly, 28(1):22–27, February 1990. CODEN FIBQAU. ISSN

0015-0517.  URL http://www.fq.math.ca/Scanned/28-1/
kimberling.pdf.

Kimberling:1991:SOR

[Kim91a] Clark Kimberling. Second-order recurrence and iterates of [an + 1/2]. Fibonacci Quarterly, 29(3):194–196, August 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-3/kimberling1.pdf.

Kimberling:1991:SOS

[Kim91b] Clark Kimberling. Second-order Stolarsky arrays. Fibonacci Quarterly, 29(4):339–342, November 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-
4/kimberling.pdf.

Kimberling:1991:STD

[Kim91c] Clark Kimberling. Sets of terms that determine all the terms of a linear recurrence sequence. Fibonacci Quarterly, 29(3):244–248, August 1991. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/29-3/kimberling2.pdf.

Kimberling:1991:ZNS

[Kim91d] Clark Kimberling. Zeckendorf number systems and associated partitions. Fibonacci Quarterly, 29(2):120–123, May 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-2/kimberling.pdf.

Kimberling:1994:FCI

[Kim94] Clark Kimberling. The first column of an interspersion. Fibonacci Quarterly, 32(4):301–315, August 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-
4/kimberling.pdf.

Kimberling:1995:CCI

[Kim95a] Clark Kimberling. Conjectures concerning irrational numbers and integers. Fibonacci Quarterly, 33(3):208–210, June 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-3/kimberling.pdf.

Kimberling:1995:ZAE

[Kim95b] Clark Kimberling. The Zeckendorf array equals the Wythoff array. Fibonacci Quarterly, 33(1):3–8, February 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-1/kimberling.pdf.

Kimberling:1998:EZ

[Kim98a] Clark Kimberling. Edouard Zeckendorf. Fibonacci Quarterly, 36(5):416–418, November 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-
5/kimberling.pdf.

Kimberling:1998:PSI

[Kim98b] Clark Kimberling. Palindromic sequences from irrational numbers. Fibonacci Quarterly, 36(2):171–173, May 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-2/kimberling.pdf.

Kimberling:2001:EPC

[Kim01a] Clark Kimberling. Enumeration of paths, compositions of integers, and Fibonacci numbers. Fibonacci Quarterly, 39(5):430– 435, November 2001. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/39-5/kimberling.pdf.

Kimberling:2001:SNA

[Kim01b] Clark Kimberling. Symbiotic numbers associated with irrational numbers. Fibonacci Quarterly, 39(4):365–372, August 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-4/kimberling.pdf.

Kimberling:2002:PCF

[Kim02] Clark Kimberling. Path-counting and Fibonacci numbers. Fibonacci Quarterly, 40(4):328–338, August 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-4/kimberling.pdf.

Kimberling:2004:OPF

[Kim04] Clark Kimberling. Orderings of products of Fibonacci numbers. Fibonacci Quarterly, 42(1):28–35, February 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-1/quartkimberling01_2004.pdf.

Kimberling:2007:EUR

[Kim07] Clark Kimberling. The equation m^{2} − 4k = 5n^{2} and unique representations of positive integers. Fibonacci Quarterly, 45(4):304– 312, November 2007. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/45-4/kimberling.pdf.

Kimberling:2010:DIS

[Kim10a] Clark Kimberling. Doubly interspersed sequences, double interspersions and fractal sequences. Fibonacci Quarterly, 48(1):13– 20, February 2010. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/48-1/kimberling.pdf.

Kimberling:2010:PSG

[Kim10b] Clark Kimberling. Partial sums of generating functions as polynomial sequences. Fibonacci Quarterly, 48(4):327–334, November 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-4/kimberling.pdf.

Kimberling:2010:PDS

[Kim10c] Clark Kimberling. Polynomials defined by a second-order recurrence, interlacing zeros, and Gray codes. Fibonacci Quarterly, 48(3):209–218, August 2010. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/48-3/
kimberling.pdf.

Kimberling:2011:BSW

[Kim11] Clark Kimberling. Beatty sequences and Wythoff sequences, generalized. Fibonacci Quarterly, 49(3):195–200, August 2011. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/49-3/kimberling.pdf.

Kimberling:2012:LPS

[Kim12] Clark Kimberling. Limits of polynomial sequences. Fibonacci Quarterly, 50(4):294–??, November 2012. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/50-
4/kimberling.pdf.

Kimberling:2014:FFF

[Kim14] Clark Kimberling. Fusion, fission, and factors. Fibonacci Quarterly, 52(3):195–??, August 2014. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/52-3/
kimberling.pdf.

Kimberling:2017:PP

[Kim17] Clark Kimberling. Problem proposals. Fibonacci Quarterly, 55 (5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Papers1/55-5/Problems.pdf.

Kimberling:2019:PS

[Kim19] Clark Kimberling. Problem section. Fibonacci Quarterly, 57(5): 170–??, December 2019. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers1/57-5/problems.pdf.

King:1963:LF

[Kin63] Charles King. Leonardo Fibonacci. Fibonacci Quarterly, 1(4):15– 20, December 1963. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/1-4/king.pdf.

King:1968:CGF

[Kin68] Charles H. King. Conjugate generalized Fibonacci sequences. Fibonacci Quarterly, 6(1):46–49, February 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-1/
king.pdf.

King:1973:PGF

[Kin73] Bruce W. King. A polynomial with generalized Fibonacci coefficients. Fibonacci Quarterly, 11(5):527–532, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-5/king.pdf.

Kirkpatrick:1977:FSA

[Kir77] Thomas B. Kirkpatrick, Jr. Fibonacci sequences and additive triangles of higher order and degree. Fibonacci Quarterly, 15(4):319–322, December 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
4/kirkpatrick.pdf.

Kirrinnis:1994:OBP

[Kir94] Peter Kirrinnis. An optimal bound for path weights in Huffman trees. Information Processing Letters, 51(2):107–110, July 26, 1994. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kirgizov:2022:BWN

[Kir22] Sergey Kirgizov. Q-bonacci words and numbers. Fibonacci Quarterly, 60(5):187–??, December 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/60-
5/kirgizov1.pdf.

Kishore:1981:E

[Kis81] Masao Kishore. On the equation σ(m)σ(n) = (m+n)^{2}. Fibonacci Quarterly, 19(1):21–23, February 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
1/kishore.pdf.

Kiss:1988:NSD

[Kis88] Peter Kiss. On the number of solutions of the Diophantine equation binom(x, p) = binom(y, 2). Fibonacci Quarterly, 26(2):127– 129, May 1988. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/26-2/kiss.pdf.

Kisielewicz:2014:SBI

[Kis14] Andrzej P. Kisielewicz. Some binomial identities arising from a partition of an n-dimensional cube. Fibonacci Quarterly, 52(4): 325–??, November 2014. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/52-4/kisielewicz.
pdf.

Keinert:2015:SFM

[KISS15] Benjamin Keinert, Matthias Innmann, Michael Sänger, and Marc Stamminger. Spherical Fibonacci mapping. ACM Transactions on Graphics, 34(6):193:1–193:??, November 2015. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Knopfmacher:1989:RRN

[KK89] Arnold Knopfmacher and John Knopfmacher. Representations for real numbers via k th powers of integers. Fibonacci Quarterly, 27(1):49–60, February 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-
1/knopfmacher.pdf.

Khan:1995:SIM

[KK95] M. A. Khan and Y. H. Harris Kwong. Some invariant and minimum properties of Stirling numbers of the second kind. Fibonacci Quarterly, 33(3):203–205, June 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-3/
khan.pdf.

Kihel:2001:SWP

[KK01] Abdelkrim Kihel and Omar Kihel. Sets in which the product of any K elements increased by t is a k th power. Fibonacci Quarterly, 39(2):98–100, May 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-2/
kihel.pdf.

Khan:2005:APS

[KK05] M. A. Khan and Harris Kwong. Arithmetic progressions with square entries. Fibonacci Quarterly, 43(2):98–103, May 2005.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/43-2/paper43-2-1.pdf.

Khan:2011:SBI

[KK11] M. A. Khan and Harris Kwong. Some binomial identities associated with the generalized natural number sequence. Fibonacci Quarterly, 49(1):57–65, February 2011. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/49-1/
khan.pdf.

Khan:2014:SPF

[KK14] M. A. Khan and Harris Kwong. On sums of products of Fibonacci-type recurrences. Fibonacci Quarterly, 52(1):20–??, February 2014. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/52-1/khan.pdf.

Keskin:2015:GFL

[KK15] Refık Keskin and Olcay Karaatli. Generalized Fibonacci and Lucas numbers of the form 5x^{2}. International Journal of Number Theory (IJNT), 11(3):931–944, May 2015. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042115500517.

Kebli:2021:NIS

[KKLL21] Salima Kebli, Omar Kihel, Jesse Larone, and Florian Luca. On the nonnegative integer solutions to the equation F_{n} ± F_{m} = y^{a}. Journal of Number Theory, 220(??):107–127, March 2021. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X20302559.

Kimberling:2018:CBH

[KKLS18] Clark Kimberling, Takao Komatsu, Kálmán Liptai, and László Szalay. A connection between hyper-Fibonacci numbers and fissions of polynomial sequences. Fibonacci Quarterly, 56(3):195–??, August 2018. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/56-3/kimberling.pdf.

Kologlu:2011:NSZ

[KKMW11] Murat Koloğlu, Gene S. Kopp, Steven J. Miller, and Yinghui Wang. On the number of summands in Zeckendorf decompositions. Fibonacci Quarterly, 49(2):116–130, May 2011. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/49-2/kologlu.pdf.

Krall:1987:CSN

[KL87] Allan M. Krall and Lance L. Littlejohn. A curious set of numbers. Fibonacci Quarterly, 25(4):352–355, November 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-4/krall.pdf.

Konvalina:1991:CIS

[KL91a] John Konvalina and Yi-Hsin Liu. A combinatorial interpretation of the square of a Lucas number. Fibonacci Quarterly, 29(3): 268–270, August 1991. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/29-3/konvalina.pdf.

Konvalina:1991:SUS

[KL91b] John Konvalina and Yi-Hsin Liu. Subsets without unit separation and products of Fibonacci numbers. Fibonacci Quarterly, 29(2): 141–144, May 1991. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/29-2/konvalina.pdf.

Konvalina:1993:CSS

[KL93] John Konvalina and Yi-Hsin Liu. Circular subsets without qseparation and powers of Lucas numbers. Fibonacci Quarterly, 31(3):275–??, August 1993. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/31-3/konvalina.pdf.

Klarner:1965:SRC

[Kla65] David A. Klarner. Some results concerning polyominoes. Fibonacci Quarterly, 3(1):9–20, February 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-1/
klarner.pdf.

Klarner:1966:DIT

[Kla66a] D. A. Klarner. Determinants involving K th powers from second order sequences. Fibonacci Quarterly, 4(2):179–183, April 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-2/klarner.pdf.

Klarner:1966:RSD

[Kla66b] David A. Klarner. Representation of N as a sum of distinct elements from special sequences. Fibonacci Quarterly, 4(4):289– 306, December 1966. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/4-4/klarner-a.pdf.

Klamkin:1968:LDE

[Kla68a] Murray S. Klamkin. On the linear difference equation whose solutions are the products of solutions of two given linear difference equations. Fibonacci Quarterly, 6(5):86–92, November 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/6-5/klamkin.pdf.

Klarner:1968:PDF

[Kla68b] David A. Klarner. Partitions of N into distinct Fibonacci numbers. Fibonacci Quarterly, 6(4):235–243, October 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-4/klarner.pdf.

Klamkin:1972:NP

[Kla72] M. S. Klamkin. A number problem. Fibonacci Quarterly, 10(3): 324–??, April 1972. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/10-3/klamkin.pdf.

Klamkin:1973:SNH

[Kla73a] Murray S. Klamkin. On solving non-homogeneous linear difference equations. Fibonacci Quarterly, 11(2):166–168, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-2/klamkin1.pdf.

Klarner:1973:NSC

[Kla73b] David A. Klarner. The number of SDR’s in certain regular systems. Fibonacci Quarterly, 11(3):267–270, October 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-3/klarner.pdf.

Klarner:1976:MPG

[Kla76] David A. Klarner. A model for population growth. Fibonacci Quarterly, 14(3):277–281, October 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
3/klarner.pdf.

Klarner:1981:URS

[Kla81] David A. Klarner. The ubiquitous rational sequence. Fibonacci Quarterly, 19(3):219–227, August 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
3/klarner.pdf.

Klaska:2008:TWP

[Kla09] Jiří Klaška. On Tribonacci–Wieferich primes. Fibonacci Quarterly, 46/47(4):290–297, November 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/46_47-4/klaska.pdf.

Klaska:2018:DDW

[Kla18] Jiří Klaška. Donald Dines Wall’s conjecture. Fibonacci Quarterly, 56(1):43–??, February 2018. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/56-1/klaska.pdf.

Klein:1991:CRG

[Kle91] Shmuel T. Klein. Combinatorial representation of generalized Fibonacci numbers. Fibonacci Quarterly, 29(2):124–131, May 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-2/klein.pdf.

Klein:2007:FDF

[Kle07] S. T. Klein. Fast decoding of Fibonacci encoded texts. In Storer and Cohn [SC07], page 388. ISBN 0-7695-2791-4. ISSN 1068-

0314. LCCN ???? URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4148789. IEEE Computer Society

Order Number P2791.

Kliorys:1981:FNI

[Kli81] Constantine Kliorys. Fibonacci number identities from algebraic units. Fibonacci Quarterly, 19(2):149–152, April 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-2/kliorys.pdf.

Kafle:2019:CPE

[KLM^{+}19] Bir Kafle, Florian Luca, Amanda Montejano, László Szalay, and Alain Togbé. On the X-coordinates of Pell equations which are products of two Fibonacci numbers. Journal of Number Theory, 203(??):310–333, October 2019. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X19301167.

Kafle:2018:TR

[KLT18] Bir Kafle, Florian Luca, and Alain Togbé. Triangular repblocks. Fibonacci Quarterly, 56(4):325–??, November 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/56-4/kafle.pdf.

Knopfmacher:1995:PER

[KM95] Arnold Knopfmacher and M. E. Mays. Pierce expansions of ratios and Fibonacci and Lucas numbers and polynomials. Fibonacci Quarterly, 33(2):153–163, May 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-2/
knopfmacher.pdf.

Kitaev:2005:LRC

[KM05] Sergey Kitaev and Toufik Mansour. Linear recurrences and Chebyshev polynomials. Fibonacci Quarterly, 43(3):256–261, August 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Papers/43-3/paper43-3-8.pdf.

Kimberling:2014:IFT

[KM14] Clark Kimberling and Peter J. C. Moses. The infinite Fibonacci tree and other trees generated by rules. Fibonacci Quarterly, 52(5):136–??, December 2014. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/52-
5/Kimberling.pdf.

Kaddoura:2018:NIU

[KM18] Issam Kaddoura and Bassam Mourad. On a new improved unifying closed formula for all Fibonacci-type sequences and some applications. Journal of Number Theory, 182(??):271–283, January 2018. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X1730241X.

Kimberling:2019:LCE

[KM19] Clark Kimberling and Peter J. C. Moses. Linear complementary equations and systems. Fibonacci Quarterly, 57(5):96–??, December 2019. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/57-5/kimberling.pdf.

Komatsu:2014:HOI

[KMP14] Takao Komatsu, Zuzana Masáková, and Edita Pelantová. Higherorder identities for Fibonacci numbers. Fibonacci Quarterly, 52 (5):150–??, December 2014. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers1/52-5/Komatsu.pdf.

Klostermeyer:1997:PR

[KMST97] W. F. Klostermeyer, M. E. Mays, L. Soltes, and G. Trapp. A Pascal rhombus. Fibonacci Quarterly, 35(4):318–328, November 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/35-4/klostermeyer.pdf.

Kinlaw:2022:SRF

[KMT22] Paul Kinlaw, Michael Morris, and Samanthak Thiagarajan. Sums related to the Fibonacci sequence. Fibonacci Quarterly, 60(2): 136–??, May 2022. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/60-2/kinlaw.pdf.

Kim:2008:FMG

[KN08] Hee Sik Kim and J. Neggers. Fibonacci mean and golden section mean. Computers and Mathematics and Applications, 56(1):228– 232, July 2008. CODEN CMAPDK. ISSN 0898-1221 (print),

1873-7668 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0898122108000096.

Knopfmacher:1981:IDN

[Kno81] J. Knopfmacher. Initial digits in number theory. Fibonacci Quarterly, 19(2):121–126, April 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
2/knopfmacher.pdf.

Knopfmacher:1990:RPR

[Kno90] Arnold Knopfmacher. A radix product representation for real numbers. Fibonacci Quarterly, 28(4):290–297, November 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/28-4/knopfmacher.pdf.

Knopfmacher:1992:EPS

[Kno92a] Arnold Knopfmacher. Elementary properties of the subtractive Euclidean algorithm. Fibonacci Quarterly, 30(1):80–83, February 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/30-1/knopfmacher.pdf.

Knox:1992:FSF

[Kno92b] Steven W. Knox. Fibonacci sequences in finite groups. Fibonacci Quarterly, 30(2):116–120, May 1992. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/30-2/
knox.pdf.

Knuth:1964:TNB

[Knu64] Donald Knuth. Transcendental numbers based on the Fibonacci sequence. Fibonacci Quarterly, 2(1):43–44, February 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-1/knuth-a.pdf.

Knuth:1966:ALR

[Knu66a] Donald E. Knuth. An almost linear recurrence. Fibonacci Quarterly, 4(2):117–128, April 1966. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/4-2/knuth.pdf.

See correction [Knu66b].

Knuth:1966:CAL

[Knu66b] Donald E. Knuth. Correction: “An almost linear recurrence”. Fibonacci Quarterly, 4(4):354, December 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-4/
correction.pdf. See [Knu66a].

Knuth:1974:LE

[Knu74] Donald E. Knuth. Letter to the editor. Fibonacci Quarterly, 12(1):46, 79, 82, February 1974. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/12-1/letter-a.
pdf.

Knuth:1988:FM

[Knu88] Donald E. Knuth. Fibonacci multiplication. Applied Mathematics Letters, 1(1):57–60, 1988. CODEN AMLEEL. ISSN 0893-9659 (print), 1873-5452 (electronic).

Knuth:1990:FLS

[Knu90] Donald E. Knuth. A Fibonacci-like sequence of composite numbers. Mathematics Magazine, 63(1):21–25, February 1990. CO-DEN MAMGA8. ISSN 0025-570X.

Knuth:1997:FA

[Knu97] Donald E. Knuth. Fundamental Algorithms, volume 1 of The Art of Computer Programming. Addison-Wesley, Reading, MA, USA, third edition, 1997. ISBN 0-201-89683-4. xix + 650 pp. LCCN QA76.6 .K64 1997. US$49.50.

Kohlbecker:1966:GMC

[Koh66] Eugene E. Kohlbecker. On a generalization of multinomial coefficients for Fibonacci sequences. Fibonacci Quarterly, 4(4):307– 312, December 1966. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/4-4/kohlbecker.pdf.

Kohler:1985:GFF

[Köh85] Günter Köhler. Generating functions of Fibonacci-like sequences and decimal expansions of some fractions. Fibonacci Quarterly, 23(1):29–35, February 1985. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/23-1/kohler.pdf.

Kolodner:1965:GFA

[Kol65] I. I. Kolodner. On a generating function associated with generalized Fibonacci sequences. Fibonacci Quarterly, 3(4):272–278, December 1965. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/3-4/kolodner.pdf.

Komatsu:2001:CFN

[Kom01a] Takao Komatsu. Continued fractions and Newton’s approximations, II. Fibonacci Quarterly, 39(4):336–338, August 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-4/komatsu.pdf.

Komatsu:2001:PSI

[Kom01b] Takao Komatsu. On palindromic sequences from irrational numbers. Fibonacci Quarterly, 39(1):66–74, February 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-1/komatsu.pdf.

Komatsu:2003:IAF

[Kom03] Takao Komatsu. The interval associated with a Fibonacci number. Fibonacci Quarterly, 41(1):3–6, February 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/41-1/komatsu.pdf.

Komatsu:2008:CFE

[Kom09] Takao Komatsu. On continued fraction expansions of Fibonacci and Lucas Dirichlet series. Fibonacci Quarterly, 46/47 (3):268–278, August 2008/2009. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/46_47-
3/komatsu.pdf.

Komatsu:2010:SWA

[Kom10] Takao Komatsu. Sequences {H_{n}} for which H_{n+1}/H_{n} approaches an irrational number. Fibonacci Quarterly, 48(3):265–275, August 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/48-3/komatsu.pdf.

Komatsu:2012:CNC

[Kom12] Takao Komatsu. Congruent numbers and continued fractions. Fibonacci Quarterly, 50(3):222–??, August 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/50-3/komatsu.pdf.

Komatsu:2014:SGF

[Kom14] Takao Komatsu. Some generalized Fibonacci identities including powers and binomial coefficients. Fibonacci Quarterly, 52(1):50– ??, February 2014. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/52-1/komatsu.pdf.

Konvalina:1995:RUC

[Kon95] John Konvalina. Roots of unity and circular subsets without consecutive elements. Fibonacci Quarterly, 33(5):412–414, November 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/33-5/konvalina.pdf.

Konno:2010:SCM

[Kon10] Tatsuo Konno. Sequences constructed by a modified inclusionexclusion principle. Fibonacci Quarterly, 48(3):236–240, August 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-3/konno.pdf.

Korntved:1994:EGS

[Kor94] Ed Korntved. Extensions to the GCD Star of David Theorem. Fibonacci Quarterly, 32(2):160–166, May 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-2/
korntved.pdf.

Koshy:2001:FLN

[Kos01] Thomas Koshy. Fibonacci and Lucas numbers with applications. Pure and applied mathematics: a Wiley-Interscience series of texts, monographs, and tracts. John Wiley, New York, NY, USA, 2001. ISBN 0-471-39969-8 (hardcover). xvi + 652 pp. LCCN QA246.5 .K67 2001. URL http://www.

loc.gov/catdir/bios/wiley043/2001017506.html;   http:
//www.loc.gov/catdir/description/wiley036/2001017506.
html;   http://www.loc.gov/catdir/toc/onix07/2001017506.

html;   https://www.math.utah.edu/pub/tex/bib/fibquart.
bib.

Koshy:2009:CNA

[Kos09] Thomas Koshy. Catalan numbers with applications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-533454-X (hardcover). xiv + 422 pp. LCCN QA431 .K726

2009.  URL http://www.loc.gov/catdir/toc/ecip0825/
2008033798.html;   https://www.math.utah.edu/pub/tex/
bib/fibquart.bib.

Koshy:2014:GTM

[Kos14a] Thomas Koshy. Graph-theoretic models for the Fibonacci family. The Mathematical Gazette, 98(542):256–265, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Koshy:2014:PEL

[Kos14b] Thomas Koshy. Polynomial extensions of the Lucas and Ginsburg identities. Fibonacci Quarterly, 52(2):141–??, May 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-2/koshy.pdf.

Koshy:2015:DGP

[Kos15a] Thomas Koshy. Differences of Gibonacci products with the same order. Fibonacci Quarterly, 53(3):241–??, August 2015. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/53-3/koshy.pdf.

Koshy:2015:GCI

[Kos15b] Thomas Koshy. Gibonomial coefficients with interesting byproducts. Fibonacci Quarterly, 53(4):340–??, November 2015. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/53-4/koshy.pdf.

Koshy:2015:GTM

[Kos15c] Thomas Koshy. Graph-theoretic models for the univariate Fibonacci family. Fibonacci Quarterly, 53(2):135–??, May 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/53-2/koshy.pdf.

Koshy:2016:VPT

[Kos16] Thomas Koshy. Vieta polynomials and their close relatives. Fibonacci Quarterly, 54(2):141–??, May 2016. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/54-
2/koshy.pdf.

Koshy:2017:PEL

[Kos17] Thomas Koshy. Polynomial extensions of the Lucas and Ginsburg identities revisited. Fibonacci Quarterly, 55(2):147–??, May 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-2/koshy.pdf.

Koshy:2018:DGP

[Kos18a] Thomas Koshy. Differences of Gibonacci polynomial products of orders 2, 3, and 4. Fibonacci Quarterly, 56(3):212–??, August 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/56-3/koshy1.pdf.

Koshy:2018:PEL

[Kos18b] Thomas Koshy. Polynomial extensions of the Lucas and Ginsburg identities revisited: Additional dividends I. Fibonacci Quarterly, 56(2):106–??, May 2018. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/56-2/koshy.pdf.

Koshkin:2019:NCL

[Kos19a] Sergiy Koshkin. Non-classical linear divisibility sequences and cyclotomic polynomials. Fibonacci Quarterly, 57(1):68–??, February 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/57-1/koshkin.pdf.

Koshy:2019:JJL

[Kos19b] Thomas Koshy. Jacobsthal and Jacobsthal–Lucas walks. Fibonacci Quarterly, 57(2):99–??, May 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/57-
2/koshy1.pdf.

Koshy:2019:RGC

[Kos19c] Thomas Koshy. A recurrence for Gibonacci cubes with graphtheoretic confirmations. Fibonacci Quarterly, 57(2):139–??, May 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/57-2/koshy2.pdf.

Koshy:2019:SEG

[Kos19d] Thomas Koshy. Some extended Gibonacci polynomial sums with dividends. Fibonacci Quarterly, 57(4):303–??, November 2019.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-4/koshy1.pdf.

Koshy:2020:EGSa

[Kos20] Thomas Koshy. Extended Gibonacci sums of polynomial products of order 3. Fibonacci Quarterly, 58(3):241–??, August 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/58-3/koshy.pdf.

Koshy:2021:EGSa

[Kos21a] Thomas Koshy. Extended Gibonacci sums of polynomial products of orders 4 and 5. Fibonacci Quarterly, 59(1):3–??, February 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/59-1/koshy1.pdf.

Koshy:2021:EGSb

[Kos21b] Thomas Koshy. Extended Gibonacci sums of polynomial products of orders 4 and 5 revisited. Fibonacci Quarterly, 59(1):23–??, February 2021. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/59-1/koshy2.pdf.

Koshy:2021:FSGa

[Kos21c] Thomas Koshy. A family of sums of Gibonacci polynomial products of order 4. Fibonacci Quarterly, 59(2):98–??, May 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/59-2/koshy1.pdf.

Koshy:2021:FSGb

[Kos21d] Thomas Koshy. A family of sums of Gibonacci polynomial products of order 4 revisited. Fibonacci Quarterly, 59(3):225–??, August 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/59-3/koshy1.pdf.

Koshy:2021:GTCa

[Kos21e] Thomas Koshy. Graph-theoretic confirmations of four sums of Gibonacci polynomial products of order 4. Fibonacci Quarterly, 59(2):167–??, May 2021. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/59-2/koshy2.pdf.

Koshy:2021:GTCb

[Kos21f] Thomas Koshy. Graph-theoretic confirmations of four sums of Jacobsthal polynomial products of order 4. Fibonacci Quarterly, 59(4):319–??, November 2021. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/59-4/koshy1.
pdf.

Koshy:2021:ISIa

[Kos21g] Thomas Koshy. Infinite sums involving Gibonacci polynomial products. Fibonacci Quarterly, 59(3):237–??, August 2021. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/59-3/koshy2.pdf.

Koshy:2021:ISIb

[Kos21h] Thomas Koshy. Infinite sums involving Gibonacci polynomial products revisited. Fibonacci Quarterly, 59(3):262–??, August 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/59-3/koshy3.pdf.

Koshy:2021:ISIc

[Kos21i] Thomas Koshy. Infinite sums involving Jacobsthal polynomial products. Fibonacci Quarterly, 59(4):338–??, November 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/59-4/koshy2.pdf.

Koshy:2022:ISIf

[Kos22a] Thomas Koshy. Infinite sums involving extended Gibonacci polynomials. Fibonacci Quarterly, 60(4):292–??, November 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-4/koshy1.pdf.

Koshy:2022:ISIg

[Kos22b] Thomas Koshy. Infinite sums involving extended Gibonacci polynomials revisited. Fibonacci Quarterly, 60(4):302–??, November 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/60-4/koshy2.pdf.

Koshy:2022:ISIb

[Kos22c] Thomas Koshy. Infinite sums involving Gibonacci polynomials. Fibonacci Quarterly, 60(2):104–??, May 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/60-
2/koshy1.pdf.

Koshy:2022:ISIc

[Kos22d] Thomas Koshy. Infinite sums involving Gibonacci polynomials revisited. Fibonacci Quarterly, 60(2):120–??, May 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-2/koshy2.pdf.

Koshy:2022:ISIa

[Kos22e] Thomas Koshy. Infinite sums involving Jacobsthal polynomial products revisited. Fibonacci Quarterly, 60(1):3–??, February 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/60-1/koshy1.pdf.

Koshy:2022:ISId

[Kos22f] Thomas Koshy. Infinite sums involving Jacobsthal polynomials. Fibonacci Quarterly, 60(3):194–??, August 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/60-3/koshy1.pdf.

Koshy:2022:ISIe

[Kos22g] Thomas Koshy. Infinite sums involving Jacobsthal polynomials revisited. Fibonacci Quarterly, 60(3):229–??, August 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-3/koshy2.pdf.

Koshy:2022:PIRa

[Kos22h] Thomas Koshy. Products involving reciprocals of Gibonacci polynomials. Fibonacci Quarterly, 60(1):15–??, February 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-1/koshy2.pdf.

Koshy:2022:PIRb

[Kos22i] Thomas Koshy. Products involving reciprocals of Jacobsthal polynomials. Fibonacci Quarterly, 60(1):72–??, February 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-1/koshy3.pdf.

Koshy:2022:SIG

[Kos22j] Thomas Koshy. Sums involving Gibonacci polynomials. Fibonacci Quarterly, 60(4):344–??, November 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/60-4/koshy3.pdf.

Kuloglu:2022:NSFa

[KÖS22k] Bahar Kuloğlu, Engin Özkan, and Anthony G. Shannon. The Narayana sequence in finite groups. Fibonacci Quarterly, 60(5):

212–??, December 2022. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/60-5/kologlu.pdf.

Koshy:2023:ASId

[Kos23a] Thomas Koshy. Additional sums involving Gibonacci polynomial squares. Fibonacci Quarterly, 61(3):207–??, August 2023. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-3/koshy2.pdf.

Koshy:2023:ASI

[Kos23b] Thomas Koshy. Additional sums involving Gibonacci polynomial squares revisited. Fibonacci Quarterly, 61(4):334–??, November 2023. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/61-4/koshy5.pdf.

Koshy:2023:ASIa

[Kos23c] Thomas Koshy. Additional sums involving Gibonacci polynomials. Fibonacci Quarterly, 61(1):12–??, February 2023. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-1/koshy2.pdf.

Koshy:2023:ASIb

[Kos23d] Thomas Koshy. Additional sums involving Gibonacci polynomials: Graph-theoretic confirmations. Fibonacci Quarterly, 61(1): 42–??, February 2023. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/61-1/koshy3.pdf.

Koshy:2023:ASIc

[Kos23e] Thomas Koshy. Additional sums involving Gibonacci polynomials revisited. Fibonacci Quarterly, 61(1):60–??, February 2023. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-1/koshy4.pdf.

Koshy:2023:ASIe

[Kos23f] Thomas Koshy. Additional sums involving Jacobsthal polynomial squares. Fibonacci Quarterly, 61(3):222–??, August 2023. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-3/koshy3.pdf.

Koshy:2023:ISIa

[Kos23g] Thomas Koshy. Infinite sums involving Gibonacci polynomials revisited: Generalizations. Fibonacci Quarterly, 61(4):291–??, November 2023. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/61-4/koshy1.pdf.

Koshy:2023:ISIb

[Kos23h] Thomas Koshy. Infinite sums involving Jacobsthal polynomials: Generalizations. Fibonacci Quarterly, 61(4):305–??, November 2023. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/61-4/koshy2.pdf.

Koshy:2023:MSI

[Kos23i] Thomas Koshy. More sums involving Gibonacci polynomial squares. Fibonacci Quarterly, 61(4):346–??, November 2023. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-4/koshy6.pdf.

Koshy:2023:SIGa

[Kos23j] Thomas Koshy. Sums involving Gibonacci polynomial squares. Fibonacci Quarterly, 61(2):98–??, May 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/61-
2/koshy1.pdf.

Koshy:2023:SIG

[Kos23k] Thomas Koshy. Sums involving Gibonacci polynomial squares: Generalizations. Fibonacci Quarterly, 61(3):197–??, August 2023. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-3/koshy1.pdf.

Koshy:2023:SIGb

[Kos23l] Thomas Koshy. Sums involving Gibonacci polynomial squares: Graph-theoretic confirmations. Fibonacci Quarterly, 61(2):119– ??, May 2023. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/61-2/koshy2.pdf.

Koshy:2023:SIGc

[Kos23m] Thomas Koshy. Sums involving Gibonacci polynomial squares revisited. Fibonacci Quarterly, 61(2):153–??, May 2023. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-2/koshy4.pdf.

Koshy:2023:SIJb

[Kos23n] Thomas Koshy. Sums involving Jacobsthal polynomial squares. Fibonacci Quarterly, 61(2):135–??, May 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/61-
2/koshy3.pdf.

Koshy:2023:SIJa

[Kos23o] Thomas Koshy. Sums involving Jacobsthal polynomials. Fibonacci Quarterly, 61(1):2–??, February 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/61-
1/koshy1.pdf.

Koshy:2023:SITa

[Kos23p] Thomas Koshy. Sums involving two classes of Gibonacci polynomials. Fibonacci Quarterly, 61(4):312–??, November 2023. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-4/koshy3.pdf.

Koshy:2023:SITb

[Kos23q] Thomas Koshy. Sums involving two classes of Gibonacci polynomials revisited. Fibonacci Quarterly, 61(4):327–??, November 2023. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/61-4/koshy4.pdf.

Koshy:2024:MSI

[Kos24a] Thomas Koshy. More sums involving Gibonacci polynomial squares revisited. Fibonacci Quarterly, 62(1):29–??, February 2024. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/62-1/koshy1.pdf.

Koshy:2024:SICa

[Kos24b] Thomas Koshy. Sums involving a class of Gibonacci polynomial squares: Generalizations. Fibonacci Quarterly, 62(1):34–??, February 2024. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/62-1/koshy2.pdf.

Koshy:2024:SIF

[Kos24c] Thomas Koshy. Sums involving a family of Gibonacci polynomial squares: Generalizations. Fibonacci Quarterly, 62(1):75–??, February 2024. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/62-1/koshy4.pdf.

Koutras:1990:TCN

[Kou90] M. Koutras. Two classes of numbers appearing in the convolution of binomial-truncated Poisson and Poisson-truncated binomial random variables. Fibonacci Quarterly, 28(4):321–333, November 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/28-4/koutras.pdf.

Koutras:1994:ENA

[Kou94] M. V. Koutras. Eulerian numbers associated with sequences of polynomials. Fibonacci Quarterly, 32(1):44–57, February 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-1/koutras.pdf.

King:1969:FMP

[KP69] Bruce W. King and Francis D. Parker. A Fibonacci matrix and the permanent function. Fibonacci Quarterly, 7(5):539–544, December 1969. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/7-5/king.pdf.

Kiss:1977:CBR

[KP77] Peter Kiss and Bui Minh Phong. On the connection between the rank of apparition of a prime p in Fibonacci sequence and the Fibonacci primitive roots. Fibonacci Quarterly, 15(4):347– 349, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/15-4/kiss.pdf.

Knuth:1978:IPI

[KP78] Donald E. Knuth and Michael S. Paterson. Identities from partition involutions. Fibonacci Quarterly, 16(3):198–212, June 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-3/knuth.pdf.

Kilic:2010:GFM

[KP10a] Emrah Kilic and Helmut Prodinger. A generalized Filbert matrix. Fibonacci Quarterly, 48(1):29–33, February 2010. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/48-1/kilic.pdf.

Koseleff:2010:FK

[KP10b] P.-V. Koseleff and D. Pecker. On Fibonacci knots. Fibonacci Quarterly, 48(2):137–143, May 2010. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/48-2/
koseleff.pdf.

Kilic:2013:VFM

[KP13] Emrah Kílíç and Helmut Prodinger. Variants of the Filbert matrix. Fibonacci Quarterly, 51(2):153–??, May 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/51-2/kilic.pdf.

Kilic:2011:FFS

[KPAO11] Emrah Kiliç, Helmut Prodinger, Ilker Akkus, and Hideyuki Ohtsuka. Formulas for Fibonomial sums with generalized Fibonacci and Lucas coefficients. Fibonacci Quarterly, 49(4):320– ??, November 2011. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/49-4/kilic.pdf.

Kirschenhofer:1983:FNG

[KPT83] Peter Kirschenhofer, Helmut Prodinger, and Robert F. Tichy. Fibonacci numbers of graphs: II. Fibonacci Quarterly, 21(3): 219–228, August 1983. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/21-3/kirschenhofer.pdf.

Kapoor:1989:OLB

[KR89] Sanjiv Kapoor and Edward M. Reingold. Optimum lopsided binary trees. Journal of the Association for Computing Machinery, 36(3):573–590, July 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). URL http://www.acm.

org/pubs/toc/Abstracts/0004-5411/65955.html;   https:
//www.math.utah.edu/pub/tex/bib/fibquart.bib.

Knopfmacher:2004:PP

[KR04] Arnold Knopfmacher and Neville Robbins. On Pell partitions. Fibonacci Quarterly, 42(4):348–352, November 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers1/42-4/quartrobbins04_2004.pdf.

Knopfmacher:2010:SPC

[KR10] Arnold Knopfmacher and Neville Robbins. Some properties of cyclic compositions. Fibonacci Quarterly, 48(3):249–255, August 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-3/knopfmacher.pdf.

Kirgizov:2022:PGBa

[KR22] Sergey Kirgizov and José L. Ramírez. Polyominoes and graphs built from Fibonacci words. Fibonacci Quarterly, 60(5):196–??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/60-5/kirgizov2.pdf.

Kravitz:1970:LLT

[Kra70] Sidney Kravitz. The Lucas–Lehmer test for Mersenne numbers. Fibonacci Quarterly, 8(1):1–3, February 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-1/
kravitz.pdf.

Kral:1972:ENA

[Krá72a] J. Král. Erratum: “A new additive pseudorandom number generator for extremely short word-length”. Information Processing Letters, 1(5):216, October ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Krá72b].

Kral:1972:NAP

[Krá72b] J. Král. A new additive pseudorandom number generator for extremely short word-length. Information Processing Letters, 1 (4):164–167, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Krá72a].

Krattenthaler:1990:TVL

[Kra90] Christian Krattenthaler. A two-variable Lagrange-type inversion formula with applications to expansion and convolution identities. Fibonacci Quarterly, 28(3):215–222, August 1990. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/28-3/krattenthaler.pdf.

Krcadinac:2006:NGG

[Krč06] Vedran Krčadinac. A new generalization of the golden ratio. Fibonacci Quarterly, 44(4):335–340, November 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/44-4/krcadinac.pdf.

Kreweras:1980:NML

[Kre80] G. Kreweras. The number of more or less “regular” permutations. Fibonacci Quarterly, 18(3):226–228, October 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-3/kreweras.pdf.

Krolak:1968:FEF

[Kro68] Patrick D. Krolak. Further extensions of Fibonaccian search to nonlinear programming problems. SIAM Journal on Control, 6 (??):258–265, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Shaw:1972:MML

[kS74] Li kung Shaw. A mathematical model of life and living. Libreria Inglesa, Buenos Aires, Argentina, second edition, 1972/1974. ???? pp. LCCN BD431 .S545 1972.

Kuipers:1988:DCS

[KS88] L. Kuipers and J.-S. Shiue. On the L^{p}-discrepancy of certain sequences. Fibonacci Quarterly, 26(2):157–162, May 1988. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/26-2/kuipers.pdf.

Kandiraju:2002:CTB

[KS02a] Gokul B. Kandiraju and Anand Sivasubramaniam. Characterizing the d-TLB behavior of SPEC CPU2000 benchmarks. ACM SIGMETRICS Performance Evaluation Review, 30(1):129–139, June 2002. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Krafft:2002:NPW

[KS02b] O. Krafft and M. Schaefer. On the number of permutations within a given distance. Fibonacci Quarterly, 40(5):429–434, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/40-5/krafft.pdf.

Kessler:2004:CPG

[KS04] David Kessler and Jeremy Schiff. A combinatoric proof and generalization of Ferguson’s formula for k-generalized Fibonacci numbers. Fibonacci Quarterly, 42(3):266–273, August 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-3/quartschiff03_2004.pdf.

Klaska:2010:CCT

[KS10a] Jiří Klaška and Ladislav Skula. The cubic character of the Tribonacci roots. Fibonacci Quarterly, 48(1):21–28, February 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/48-1/klaska.pdf.

Klaska:2010:NCC

[KS10b] Jiří Klaška and Ladislav Skula. A note on the cubic characters of Tribonacci roots. Fibonacci Quarterly, 48(4):324–326, November 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-4/klaska.pdf.

Klaska:2010:PTS

[KS10c] Jiří Klaška and Ladislav Skula. Periods of the Tribonacci sequence modulo a prime p ≡ 1 (mod 3). Fibonacci Quarterly, 48 (3):228–235, August 2010. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/48-3/klaska.pdf.

Klaska:2011:MET

[KS11] Jiří Klaška and Ladislav Skula. Mordell’s equation and the Tribonacci family. Fibonacci Quarterly, 49(4):310–??, November 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/49-4/klaska.pdf.

Kaygisiz:2013:NMC

[KŞ13] Kenan Kaygısız and Adem Şahin. A new method to compute the terms of generalized order-k Fibonacci numbers. Journal of Number Theory, 133(9):3119–3126, September 2013. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X13000991.

Kimberling:2016:STP

[KS16] Clark Kimberling and László Szalay. t-sion of two polynomial sequences and factorization properties. Fibonacci Quarterly, 54(1): 3–??, February 2016. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/54-1/kimberling.pdf.

Kuipers:1972:DPS

[KsS72] Lawrence Kuipers and Jau shyong Shiue. A distribution property of the sequence of Fibonacci numbers. Fibonacci Quarterly, 10 (4):375–376, October 1972. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/10-4/kuipers-a.pdf.

Kuipers:1973:RPD

[KsS73] L. Kuipers and Jau shyong Shiue. Remark on a paper by Duncan and Brown on the sequence of logarithms of certain recursive sequences. Fibonacci Quarterly, 11(3):292–293, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-3/kuipers.pdf. See [BD70].

Kogan:2016:FFI

[KSSS16] Tamara Kogan, Luba Sapir, Amir Sapir, and Ariel Sapir. The Fibonacci family of iterative processes for solving nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 110(??):148–158, December 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0168927416301623.

Kafle:2020:MEP

[KST20] Bir Kafle, Anitha Srinivasan, and Alain Togbé. Markoff equation with Pell components. Fibonacci Quarterly, 58(3):226–??, August 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/58-3/kafle.pdf.

Kaplan:2008:THT

[KT08] Haim Kaplan and Robert Endre Tarjan. Thin heaps, thick heaps. ACM Transactions on Algorithms, 4(1):3:1–3:??, March 2008. CODEN ???? ISSN 1549-6325 (print), 1549-6333 (electronic).

Kuhapatanakul:2019:GFL

[KT19] K. Kuhapatanakul and K. Thongsing. Generalizations of the Fibonacci–Lucas relations. American Mathematical Monthly, 126 (1):81, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Knuth:1984:CQD

[KTW84] Donald E. Knuth, Erwin Tomash, and M. R. Williams. Comments, queries, and debate: FORTRAN implementations; the Madrid promptuary; Edouard Lucas vindicated; photo essays. Annals of the History of Computing, 6(4):402–405, October/December 1984. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1984/pdf/a4402.pdf;   http://www.computer.
org/annals/an1984/a4402abs.htm.

Kubelka:2004:SSS

[Kub04] Richard P. Kubelka. Self-similarity and symmetries of Pascal’s triangles and simplices. Fibonacci Quarterly, 42(1):70–75, February 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Papers1/42-1/quartkubelka01_2004.pdf.

Kuhapatanakul:2012:SCB

[Kuh12] Kantaphon Kuhapatanakul. Some connections between a generalized Tribonacci triangle and a generalized Fibonacci sequence. Fibonacci Quarterly, 50(1):44–??, February 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-1/kuhapatanakul.pdf.

Kuipers:1969:RPR

[Kui69] L. Kuipers. Remark on a paper by R. L. Duncan concerning the uniform distribution mod 1 of the sequence of the logarithms of the Fibonacci numbers. Fibonacci Quarterly, 7(5):465–466, December 1969. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/7-5/kuipers-a.pdf.

Kuipers:1976:LE

[Kui76] L. Kuipers. Letter to the editor. Fibonacci Quarterly, 14(3): 214–??, October 1976. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/14-3/letter-a.pdf.

Kuipers:1982:PFS

[Kui82] L. Kuipers. A property of the Fibonacci sequence (F_{m}), m = 0, 1, . . .. Fibonacci Quarterly, 20(2):112–113, May 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-2/kuipers.pdf.

Kumar:1974:FPC

[Kum74] Santosh Kumar. Fibonaccian pathological curves. Fibonacci Quarterly, 12(1):92–94, February 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
1/kumar.pdf.

Kumar:2021:IDS

[Kum21] Sachin Kumar. Image data security using Quasigroup combined with Fibonacci Q-transformation. Journal of Information Security and Applications (JISA), 61(??):??, September 2021. CO-

DEN ???? ISSN 2214-2126. URL http://www.sciencedirect.
com/science/article/pii/S2214212621001575.

Kung:1976:PTP

[Kun76] S. H. L. Kung. Parity triangles of Pascal’s triangle. Fibonacci Quarterly, 14(1):54–??, February 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
1/kung.pdf.

Kunoff:1987:FDP

[Kun87] Sharon Kunoff. N! has the first digit property. Fibonacci Quarterly, 25(4):365–367, November 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-
4/kunoff.pdf.

Kundert:1990:SCT

[Kun90] Esayas George Kundert. A Von Staudt–Clausen theorem for certain Bernoullianlike numbers and regular primes of the first and second kind. Fibonacci Quarterly, 28(1):16–21, February 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/28-1/kundert.pdf.

Kutz:2002:LBL

[Kut02] Martin Kutz. Lower bounds for Lucas chains. SIAM Journal on Computing, 31(6):1896–1908, December 2002. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic). URL http:

//epubs.siam.org/sam-bin/dbq/article/37925.

Kuhn:1992:NO

[KV92] Steven T. Kuhn and Andrew Vogt. Numbers without ones. Fibonacci Quarterly, 30(1):48–53, February 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-1/kuhn.pdf.

Kimball:1995:CFC

[KW95] William A. Kimball and William A. Webb. A congruence for Fibonomial coefficients modulo p^{3}. Fibonacci Quarterly, 33(4): 290–297, August 1995. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/33-4/kimball.pdf.

Kuhapatanakul:2013:EPS

[KW13] Kantaphon Kuhapatanakul and Rattanapol Wasutharat. Expressions for the products of the second order linear recurrences. Fibonacci Quarterly, 51(1):49–??, February 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/51-1/kuhapatanakul.pdf.

Killpatrick:2021:BFC

[KW21] Kendra Killpatrick and Jordan Weaver. A bijection for the Fibonomial coefficients. Fibonacci Quarterly, 59(1):14–??, February 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/59-1/killpatrick.pdf.

Kwong:1989:MPM

[Kwo89a] Y. H. Harris Kwong. Minimum periods of S(n, k) modulo M. Fibonacci Quarterly, 27(3):217–220, June 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-3/
kwong.pdf.

Kwong:1989:MPB

[Kwo89b] Y. H. Harris Kwong. Minimum periods of binomial coefficients modulo M. Fibonacci Quarterly, 27(4):348–351, August 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-4/kwong.pdf.

Kwong:2007:TDF

[Kwo07] Harris Kwong. Two determinants with Fibonacci and Lucas entries. Applied Mathematics and Computation, 194(2):568–571, December 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Kwong:2010:RFR

[Kwo10] Harris Kwong. On recurrences of Fahr and Ringel: An alternate approach. Fibonacci Quarterly, 48(4):363–365, November 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/48-4/kwong.pdf.

Kwong:2014:APS

[Kwo14] Harris Kwong. An alternate proof of Sury’s Fibonacci–Lucas relation. American Mathematical Monthly, 121(6):514, June 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/10.4169/amer.
math.monthly.121.06.514.

Kwong:2017:EPSa

[Kwo17a] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 55(1):82–??, February 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProbSolnFeb2017.pdf.

Kwong:2017:EPSb

[Kwo17b] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 55(2):178–??, May 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProvSolnMay2017.pdf.

Kwong:2017:EPSc

[Kwo17c] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 55(3):276–??, August 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProSolnAugust2017.pdf.

Kwong:2017:EPSd

[Kwo17d] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 55(4):367–??, November 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProbSolnNov2017.pdf.

Kwong:2018:EPSa

[Kwo18a] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 56(1):81–??, February 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProbSolnFeb2018.pdf.

Kwong:2018:EPSb

[Kwo18b] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 56(2):177–??, May 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProbMay2018.pdf.

Kwong:2018:EPSc

[Kwo18c] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 56(3):275–??, August 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProSolnAugust2018.pdf.

Kwong:2018:EPSd

[Kwo18d] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 56(4):366–??, November 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProbSolnNov2018.pdf.

Kwong:2019:EPSa

[Kwo19a] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 57(1):81–??, February 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2019Elem.pdf.

Kwong:2019:EPSb

[Kwo19b] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 57(2):176–??, May 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProbMay2019.pdf.

Kwong:2019:EPSc

[Kwo19c] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 57(3):276–??, August 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProSolnAugust2019.pdf.

Kwong:2019:EPSd

[Kwo19d] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 57(4):367–??, November 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProbSolnNov2019.pdf.

Kwong:2020:EPSa

[Kwo20a] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 58(1):81–??, February 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2020Elem.pdf.

Kwong:2020:EPSb

[Kwo20b] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 58(2):178–??, May 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProbMay2020.pdf.

Kwong:2020:EPSc

[Kwo20c] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 58(3):274–??, August 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProSolnAugust2020.pdf.

Kwong:2020:EPSd

[Kwo20d] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 58(4):367–??, November 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProbSolnNov2020.pdf.

Kwong:2021:EPSa

[Kwo21a] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 59(1):82–??, February 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
FQElemProbFeb2021.pdf.

Kwong:2021:EPSb

[Kwo21b] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 59(2):176–??, May 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
FQElemProbMay2021.pdf.

Kwong:2021:EPSc

[Kwo21c] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 59(3):273–??, August 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProSolnAugust2021.pdf.

Kwong:2021:EPSd

[Kwo21d] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 59(4):365–??, November 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProbSolnNov2021.pdf.

Kwong:2022:EPSa

[Kwo22a] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 60(1):82–??, February 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
FQElemProbFeb2022.pdf.

Kwong:2022:EPSb

[Kwo22b] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 60(2):177–??, May 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
FQElemProbMay2022.pdf.

Kwong:2022:EPSc

[Kwo22c] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 60(3):274–??, August 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProSolnAugust2022.pdf.

Kwong:2022:EPS

[Kwo22d] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 60(4):365–??, November 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProbSolnNov2022.pdf.

Kwong:2023:EPSa

[Kwo23a] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 61(1):84–??, February 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
FQElemProbFeb2023.pdf.

Kwong:2023:EPSb

[Kwo23b] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 61(2):178–??, May 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
FQElemProbMay2023.pdf.

Kwong:2023:EPSc

[Kwo23c] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 61(3):275–??, August 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
ElemProSolnAugust2023.pdf.

Kwong:2023:EPSd

[Kwo23d] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 61(4):367–??, November 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
FQElemProbNov2023.pdf.

Kwong:2024:EPS

[Kwo24] Harris Kwong. Elementary problems and solutions. Fibonacci Quarterly, 62(1):84–??, February 2024. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
FQElemProbFeb2024.pdf.

Karaduman:2003:PFS

[KY03] Erdal Karaduman and Uğur Yavuz. On the period of Fibonacci sequences in nilpotent groups. Applied Mathematics and Computation, 142(2–3):321–332, October 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Keepers:2008:HOL

[KY09] Kyle Keepers and Paul Thomas Young. On higher order Lucas– Bernoulli numbers. Fibonacci Quarterly, 46/47(1):26–31, February 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/46_47-1/young.pdf.

Komatsu:2020:CGS

[KY20] Takao Komatsu and Paul Thomas Young. Convolutions of generalized Stirling numbers and degenerate Bernoulli polynomials. Fibonacci Quarterly, 58(4):361–??, November 2020. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/58-4/komatsu.pdf.

Kyriakoussis:1985:CCE

[Kyr85] A. Kyriakoussis. A congruence for a class of exponential numbers. Fibonacci Quarterly, 23(1):45–48, February 1985. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-1/kyriakoyssis.pdf.

Kyriakoussis:1993:EGS

[Kyr93] A. G. Kyriakoussis. On extended generalized Stirling pairs. Fibonacci Quarterly, 31(1):44–51, February 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-1/kyriakoussis.pdf.

Kyriakoussis:1994:CWC

[Kyr94] A. Kyriakoussis. Congruences for a wide class of integers by using Gessel’s method. Fibonacci Quarterly, 32(1):79–84, February 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/32-1/kyriakoussis.pdf.

Kurt:2013:FCM

[KYS13] Ayşe Kurt, Salih Yalçınbaş, and Mehmet Sezer. Fibonacci collocation method for solving linear differential–difference equations. Mathematical and Computational Applications, 18(3):448– 458, December 2013. CODEN ???? ISSN 2297-8747. URL

https://www.mdpi.com/2297-8747/18/3/448.

Kiss:1992:GRS

[KZ92] Péter Kiss and Béla Zay. On a generalization of a recursive sequence. Fibonacci Quarterly, 30(2):103–109, May 1992. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/30-2/kiss.pdf.

Klavzar:2005:FCR

[KŽ05] Sandi Klavžar and Petra Žigert. Fibonacci cubes are the resonance graphs of Fibonaccenes. Fibonacci Quarterly, 43(3):269– 276, August 2005. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Papers/43-3/paper43-3-10.pdf.

Long:1992:AGG

[LA92] Calvin T. Long and Shiro Ando. Another generalization of Gould’s Star of David Theorem. Fibonacci Quarterly, 30(3):251– 255, August 1992. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/30-3/long.pdf.

LaBarbera:1971:FLE

[LaB71] Sal LaBarbera. Fibonacci, Lucas and the Egyptians. Fibonacci Quarterly, 9(2):177–187, April 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-2/
labarbera-a.pdf.

LaDue:2018:CIE

[LaD18] Mark D. LaDue. Clusters of integers with equal total stopping times in the 3X + 1 problem. Fibonacci Quarterly, 56(2):156–??, May 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/56-2/ladue.pdf.

Lafer:1964:EFR

[Laf64] Phil Lafer. Exploring the Fibonacci representation of integers. Fibonacci Quarterly, 2(2):114–??, April 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-2/
lafer-a.pdf.

Lafer:1971:DST

[Laf71] Phil Lafer. Discovering the square-triangular numbers. Fibonacci Quarterly, 9(1):93–105, February 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-1/
lafer.pdf.

LaGrange:2013:CDF

[LaG13] John D. LaGrange. A combinatorial development of Fibonacci numbers in graph spectra. Linear Algebra and its Applications, 438(11):4335–4347, June 1, 2013. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0024379513001109.

Lopez-Aguayo:2016:STN

[LAL16] Daniel López-Aguayo and Florian Luca. Sylvester’s theorem and the non-integrality of a certain binomial sum. Fibonacci Quarterly, 54(1):44–??, February 2016. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/54-1/lopez-
aguayo.pdf.

Lange:1981:HA

[Lan81] Lester H. Lange. Hommage à Archimède. Fibonacci Quarterly, 19(3):214–218, August 1981. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/19-3/lange.pdf.

Lander:1990:ESU

[Lan90] Leon J. Lander. Equal sums of unlike powers. Fibonacci Quarterly, 28(2):141–150, May 1990. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/28-2/lander.
pdf.

Lang:1992:CPF

[Lan92] Wolfdieter Lang. A combinatorial problem in the Fibonacci number system and two-variable generalizations of Chebyshev’s polynomials. Fibonacci Quarterly, 30(3):199–210, August 1992. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/30-3/lang.pdf.

Lang:2000:PRP

[Lan00] Wolfdieter Lang. On polynomials related to powers of the generating function of Catalan’s numbers. Fibonacci Quarterly, 38(5): 408–419, November 2000. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/38-5/lang.pdf.

Lang:2002:PRD

[Lan02] Wolfdieter Lang. On polynomials related to derivatives of the generating function of Catalan numbers. Fibonacci Quarterly, 40 (4):299–313, August 2002. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/40-4/lang.pdf.

Lang:2004:RMF

[Lan04] Wolfdieter Lang. Riccati meets Fibonacci. Fibonacci Quarterly, 42(3):231–244, August 2004. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers1/42-3/quartlang03_
2004.pdf.

Lan:2011:NCG

[Lan11] Yueheng Lan. Novel computation of the growth rate of generalized random Fibonacci sequences. Journal of Statistical Physics, 142(4):847–861, February 2011. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-011-0132-z.

Larson:1978:GSE

[Lar78] Paul Larson. The golden section in the earliest notated Western music. Fibonacci Quarterly, 16(6):513–514, December 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-6/larson.pdf.

Larison:1986:SS

[Lar86] Clarence B. Larison. Sidney’s series. Fibonacci Quarterly, 24(4): 313–315, November 1986. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/24-4/larison-a.pdf.

Larcher:1988:NEP

[Lar88] Gerhard Larcher. A new extremal property of the Fibonacci ratio. Fibonacci Quarterly, 26(3):247–255, August 1988. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/26-3/larcher.pdf.

Larcombe:2015:CFE

[Lar15a] Peter J. Larcombe. Closed form evaluations of some series comprising sums of exponentiated multiples of two-term and threeterm Catalan number linear combinations. Fibonacci Quarterly, 53(3):253–??, August 2015. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/53-3/
larcombe.pdf.

Larcombe:2015:NIG

[Lar15b] Peter J. Larcombe. A note on the invariance of the general 2 × 2 matrix anti-diagonals ratio with increasing matrix power: Four proofs. Fibonacci Quarterly, 53(4):360–??, November 2015. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/53-4/larcombe.pdf.

Larcombe:2016:ESE

[Lar16] Peter J. Larcombe. On the evaluation of sums of exponentiated multiples of generalized Catalan number linear combinations using a hypergeometric approach. Fibonacci Quarterly, 54(3):259– ??, August 2016. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/54-3/larcombe2.pdf.

Larsen:2020:FSS

[Lar20] Daniel Larsen. Focusing sequences and self-similarity. Fibonacci Quarterly, 58(3):231–??, August 2020. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/58-3/
larsen.pdf.

Laszlo:1982:SPD

[Lás82] Geröcs László. Some properties of divisibility of higher-ordered linear recursive sequences. Fibonacci Quarterly, 20(4):354–359, November 1982. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/20-4/laszlo.pdf.

Lau:1977:PGS

[Lau77] Chi-Leung Lau. The periodic generating sequence. Fibonacci Quarterly, 15(2):178–181, April 1977. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/15-2/lau.
pdf.

Laughlin:2010:IMA

[Lau10] James Mc Laughlin. An identity motivated by an amazing identity of Ramanujan. Fibonacci Quarterly, 48(1):34–38, February 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-1/mclaughlin.pdf.

Lawton:1983:KTR

[Law83] Wayne M. Lawton. Kronecker’s theorem and rational approximation of algebraic numbers. Fibonacci Quarterly, 21(2):143–146, May 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/21-2/lawton.pdf.

Laxton:1974:PMW

[Lax74] R. R. Laxton. On a problem of M. Ward. Fibonacci Quarterly, 12(1):41–44, February 1974. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/12-1/laxton.pdf.

Layman:1966:RPS

[Lay66] J. W. Layman. Relatively prime sequence solutions of non-linear difference equations. Fibonacci Quarterly, 4(2):116–??, April 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/4-2/layman-a.pdf.

Layman:1975:SIN

[Lay75] John W. Layman. Some interesting necessary conditions for (a − 1)^{n} + (b − 1)^{n} − (c − 1)^{n} = 0. Fibonacci Quarterly, 13(1):42– 44, February 1975. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/13-1/layman.pdf.

Layman:1977:CGB

[Lay77] J. W. Layman. Certain general binomial–Fibonacci sums. Fibonacci Quarterly, 15(4):362–366, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/15-4/layman.pdf.

Long:1991:SOR

[LB91] Calvin Long and John Bradshaw. Second-order recurrences and the Schröder–Bernstein theorem. Fibonacci Quarterly, 29(3):239– 243, August 1991. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/29-3/long.pdf.

Larcombe:2013:RBI

[LB13] Peter J. Larcombe and Ovidiu D. Bagdasar. On a result of Bunder involving Horadam sequences: a proof and generalization. Fibonacci Quarterly, 51(2):174–??, May 2013. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/51-
2/larcombe.pdf.

Larcombe:2014:RBI

[LBF14] Peter J. Larcombe, Ovidiu D. Bagdasar, and Eric J. Fennessey. On a result of Bunder involving Horadam sequences: A new proof. Fibonacci Quarterly, 52(2):175–??, May 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/52-2/larcombe.pdf.

LEcuyer:1994:URN

[L’E94] Pierre L’Ecuyer. Uniform random number generation. Annals of Operations Research, 53(1):77–120, 1994. CODEN AOREEV. ISSN 0254-5330 (print), 1572-9338 (electronic).

Ledin:1967:CKF

[Led67] George Ledin, Jr. On a certain kind of Fibonacci sums. Fibonacci Quarterly, 5(1):45–58, February 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-1/
ledin.pdf.

Ledin:1968:E

[Led68] George Ledin, Jr. Is Eratosthenes out? Fibonacci Quarterly, 6 (4):261–265, October 1968. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/6-4/ledin.pdf.

Lee:1987:TOA

[Lee87] Jia-Sheng Lee. The K th-order analog of a result of L. Carlitz. Fibonacci Quarterly, 25(4):368–369, November 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-4/lee2.pdf.

Lee:1992:GFE

[Lee92] Jack Y. Lee. The golden-Fibonacci equivalence. Fibonacci Quarterly, 30(3):216–220, August 1992. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/30-3/lee.
pdf.

Lee:1994:NNP

[Lee94] Jack Y. Lee. A note on the negative Pascal triangle. Fibonacci Quarterly, 32(3):269–270, June 1994. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/32-3/lee.
pdf.

Lee:1997:SBL

[Lee97] Jack Y. Lee. On some basic linear properties of the secondorder inhomogeneous line-sequence. Fibonacci Quarterly, 35(2): 111–121, May 1997. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/35-2/lee.pdf.

Lee:2000:LNA

[Lee00] Gwang-Yeon Lee. k-Lucas numbers and associated bipartite graphs. Linear Algebra and its Applications, 320(1–3):51–61, November 15, 2000. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.elsevier.

nl/gej-ng/10/30/19/139/24/28/abstract.html;   http://
www.elsevier.nl/gej-ng/10/30/19/139/24/28/article.pdf.

Lee:2001:SBL

[Lee01a] Jack Y. Lee. Some basic line-sequential properties of polynomials line-sequences. Fibonacci Quarterly, 39(3):194–205, June 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-3/lee.pdf.

Lee:2001:SGF

[Lee01b] Jack Y. Lee. Some general formulas associated with the secondorder homogeneous polynomial line-sequences. Fibonacci Quarterly, 39(5):419–429, November 2001. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/39-5/lee.
pdf.

Lee:2002:MVP

[Lee02a] Jack Y. Lee. On the Morgan–Voyce polynomial generalization of the first kind. Fibonacci Quarterly, 40(1):59–65, February 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-1/lee.pdf.

Lee:2002:PLS

[Lee02b] Jack Y. Lee. On the product of line-sequences. Fibonacci Quarterly, 40(5):438–440, November 2002. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/40-5/lee.
pdf.

Lee:2010:SPF

[Lee10] K. S. Enoch Lee. On the sum of a prime and a Fibonacci number. International Journal of Number Theory (IJNT), 6 (7):1669–1676, November 2010. ISSN 1793-0421 (print), 1793-

7310 (electronic). URL https://www.worldscientific.com/
doi/10.1142/S1793042110003708.

Lefton:1982:TDF

[Lef82] Phyllis Lefton. A trinomial discriminant formula. Fibonacci Quarterly, 20(4):363–365, November 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
4/lefton.pdf.

Legendre:2015:LFT

[Leg15] Stéphane Legendre. Labeled Fibonacci trees. Fibonacci Quarterly, 53(2):152–??, May 2015. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/53-2/
legendre.pdf.

Lehmer:1964:IFP

[Leh64] E. Lehmer. On the infinitude of Fibonacci pseudo-primes. Fibonacci Quarterly, 2(3):229–??, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/2-3/lehmer.pdf.

Lehmer:1966:QCF

[Leh66] E. Lehmer. On the quadratic character of the Fibonacci root. Fibonacci Quarterly, 4(2):135–138, April 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-2/
lehmer.pdf.

Lehmer:1975:FRS

[Leh75] D. H. Lehmer. Fibonacci and related sequences in periodic tridiagonal matrices. Fibonacci Quarterly, 13(2):150–158, April 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-2/lehmer.pdf.

Lengyel:1994:CAO

[Len94a] T. Lengyel. Characterizing the 2-adic order of the logarithm. Fibonacci Quarterly, 32(5):397–401, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-5/lengyel.pdf.

Lengyel:1994:DSN

[Len94b] T. Lengyel. On the divisibility by 2 of the Stirling numbers of the second kind. Fibonacci Quarterly, 32(3):194–201, June 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-3/lengyel.pdf.

Lengyel:1995:OFL

[Len95] T. Lengyel. The order of the Fibonacci and Lucas numbers. Fibonacci Quarterly, 33(3):234–239, June 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-3/
lengyel.pdf.

Lengyel:2003:DPM

[Len03a] Tamás Lengyel. Divisibility properties by multisection. Fibonacci Quarterly, 41(1):72–79, February 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-
1/lengyel.pdf.

Lengyel:2003:NTG

[Len03b] Tamás Lengyel. A Nim-type game and continued fractions. Fibonacci Quarterly, 41(4):310–320, August 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/41-4/lengyel.pdf.

Lengyel:2006:CBP

[Len06a] Tamás Lengyel. A counting based proof of the generalized Zeckendorf’s theorem. Fibonacci Quarterly, 44(4):324–325, November 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/44-4/lengyel.pdf.

Lengyel:2006:IGF

[Len06b] Tamás Lengyel. Identities for the generating function of the multiset ⌊nΦ^{m}⌋ for m = −1, 1, 2. Fibonacci Quarterly, 44(3):274–276, August 2006. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/44-3/lengyel.pdf.

Lengyel:2017:RPA

[Len17] Tamás Lengyel. On the rate of p-adic convergence of alternating sums of powers of binomial coefficients. Fibonacci Quarterly, 55 (5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/55-5/lengyel.pdf.

Leung:2023:PER

[Leu23] King Shun Leung. ϕ-expansions of rationals. Fibonacci Quarterly, 61(2):162–??, May 2023. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/61-2/leung.pdf.

Levine:1968:EPS

[Lev68a] Eugene Levine. The existence of perfect 3-sequences. Fibonacci Quarterly, 6(5):108–304, November 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-5/
levine2.pdf.

Levine:1968:FSI

[Lev68b] Eugene Levine. Fibonacci sequences with identical characteristic values. Fibonacci Quarterly, 6(5):75–80, November 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-5/levine1.pdf.

Levine:1968:GLP

[Lev68c] Eugene Levine. On the generalized Langford problem. Fibonacci Quarterly, 6(2):135–137, April 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-2/
levine.pdf.

Levine:1984:GLG

[Lev84] Naomi Levine. The goose that laid the golden egg. Fibonacci Quarterly, 22(3):252–254, August 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-
3/levine.pdf.

Levesque:1985:OLR

[Lev85] Claude Levesque. On m-th order linear recurrences. Fibonacci Quarterly, 23(4):290–293, November 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-
4/levesque.pdf.

Levine:1988:SMR

[Lev88] Shari Lynn Levine. Suppose more rabbits are Born. Fibonacci Quarterly, 26(4):306–311, November 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
4/levine.pdf.

Levy:2001:IFF

[Lev01] Dan Levy. The irreducible factorization of Fibonacci polynomials over Q. Fibonacci Quarterly, 39(4):309–319, August 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-4/levy.pdf.

Lewis:1970:EVM

[Lew70] John B. Lewis. The educational value in mathematics. Fibonacci Quarterly, 8(5):522–529, December 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-5/
lewis.pdf.

Lewin:1980:SCI

[Lew80] Mordechai Lewin. Some combinatorial identities. Fibonacci Quarterly, 18(3):214–219, October 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-
3/lewin.pdf.

Lewin:1991:PFL

[Lew91] Mordechai Lewin. Periodic Fibonacci and Lucas sequences. Fibonacci Quarterly, 29(4):310–315, November 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/29-4/lewin.pdf.

Lewis:1995:ADP

[Lew95] Richard Lewis. Antisocial dinner parties. Fibonacci Quarterly, 33 (4):368–370, August 1995. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/33-4/lewis.pdf.

London:1969:FLN

[LF69] Hymie London and Raphael Finkelstein. On Fibonacci and Lucas numbers which are perfect powers. Fibonacci Quarterly, 7(5): 476–481, December 1969. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/7-5/london-a.pdf.

Larcombe:2014:CGC

[LF14a] Peter J. Larcombe and Eric J. Fennessey. Conditions governing cross-family member equality in a particular class of polynomial families. Fibonacci Quarterly, 52(4):349–??, November 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/52-4/larcombe.pdf.

Larcombe:2014:NLI

[LF14b] Peter J. Larcombe and Eric J. Fennessey. A non-linear identity for a particular class of polynomial families. Fibonacci Quarterly, 52(1):75–??, February 2014. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/52-1/
larcombe.pdf.

Larcombe:2015:CAD

[LF15] Peter J. Larcombe and Eric J. Fennessey. A condition for antidiagonals product invariance across powers of 2 × 2 matrix sets characterizing a particular class of polynomial families. Fibonacci Quarterly, 53(2):175–??, May 2015. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/53-
2/larcombe.pdf.

Larcombe:2016:SBP

[LF16] Peter J. Larcombe and Eric J. Fennessey. On a scaled balancedpower product recurrence. Fibonacci Quarterly, 54(3):242–??, August 2016. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/54-3/larcombe1.pdf.

Larcombe:2017:CFF

[LF17] Peter J. Larcombe and Eric J. Fennessey. A closed form formulation for the general term of a scaled triple power product recurrence sequence. Fibonacci Quarterly, 55(2):168–??, May 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-2/larcombe.pdf.

Larcombe:2019:NLR

[LF19] Peter J. Larcombe and Eric J. Fennessey. A non-linear recurrence identity class for terms of a generalized linear recurrence sequence of degree three. Fibonacci Quarterly, 57(1):10–??, February 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-1/larcombe.pdf.

Landman:1991:MPB

[LG91] Bruce M. Landman and Raymond N. Greenwell. Multiplicative partitions of bipartite numbers. Fibonacci Quarterly, 29(3):264– 267, August 1991. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/29-3/landman.pdf.

Long:1974:SBC

[LH74] Calvin T. Long and V. E. Hoggatt, Jr. Sets of binomial coefficients with equal products. Fibonacci Quarterly, 12(1):71–79, February 1974. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/12-1/long.pdf.

Luca:2008:FRF

[LH09] Florian Luca and V. Janitzio Mejía Huguet. Fibonacci–Riesel and Fibonacci–Sierpinski numbers. Fibonacci Quarterly, 46/47(3): 216–219, August 2008/2009. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/46_47-3/luca.pdf.

Luchins:1995:LRD

[LHLT95] Edith H. Luchins, Russell Hendel, Paul Lemke, and David Tuller. Linear recurrences in difference triangles. Fibonacci Quarterly, 33(5):441–452, November 1995. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/33-5/luchins.
pdf.

Loh:1996:HNT

[LHWS96] Peter Kok Keong Loh, Wen Jing Hsu, Cai Wentong, and Nadarajah Sriskanthan. How network topology affects dynamic load balancing. IEEE Parallel and Distributed Technology: Systems and Applications, 4(3):25–35, Fall 1996. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic). URL http://dlib.

computer.org/pd/books/pd1996/pdf/p3025.pdf;   http://
www.computer.org/concurrency/pd1996/p3025abs.htm.

Li:1999:FPR

[Li99a] Hua-Chieh Li. Fibonacci primitive roots and Wall’s question. Fibonacci Quarterly, 37(1):77–84, February 1999. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-1/li.pdf.

Li:1999:SOL

[Li99b] Hua-Chieh Li. On second-order linear recurrence sequences: Wall and Wyler revisited. Fibonacci Quarterly, 37(4):342–349, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/37-4/li.pdf.

Li:2000:CRR

[Li00a] Hua-Chieh Li. Complete and reduced residue systems of secondorder recurrences modulo p. Fibonacci Quarterly, 38(3):272–281, June 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/38-3/li2.pdf.

Li:2000:CEG

[Li00b] Hua-Chieh Li. Conditions for the existence of generalized Fibonacci primitive roots. Fibonacci Quarterly, 38(3):244–249, June 2000. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/38-3/li1.pdf.

Li:2012:FHM

[Li12] Hsuan-Chu Li. On Fibonacci–Hessenberg matrices and the Pell and Perrin numbers. Applied Mathematics and Computation, 218(17):8353–8358, May 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300312001087.

Li:2021:DPF

[Li21] Yao-Qiang Li. Divisibility properties of factors of the discriminant of generalized Fibonacci numbers. Fibonacci Quarterly, 59 (1):65–??, February 2021. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/59-1/li.pdf.

Light:1973:PEL

[Lig73] F. W. Light, Jr. A procedure for the enumeration of 4 × n Latin rectangles. Fibonacci Quarterly, 11(3):241–246, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-3/light.pdf.

Light:1979:ETL

[Lig79] F. W. Light, Jr. Enumeration of truncated Latin rectangles. Fibonacci Quarterly, 17(1):34–36, February 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-1/light.pdf.

Lind:1965:CNB

[Lin65] D. A. Lind. On a class of nonlinear binomial sums. Fibonacci Quarterly, 3(4):292–298, December 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-4/
lind.pdf.

Lind:1967:ECT

[Lin67a] D. A. Lind. Extended computations of terminal digit coincidences. Fibonacci Quarterly, 5(2):183–184, April 1967. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-2/lind.pdf.

Lind:1967:IFL

[Lin67b] D. A. Lind. Iterated Fibonacci and Lucas subscripts. Fibonacci Quarterly, 5(1):89–90, February 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-1/
lind2-a.pdf.

Lind:1967:MCG

[Lin67c] D. A. Lind. The Q matrix as a counterexample in group theory. Fibonacci Quarterly, 5(1):44–??, February 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/5-1/lind1-a.pdf.

Lind:1968:QFC

[Lin68] D. A. Lind. The quadratic field Q(\sqrt{5}) and a certain Diophantine equation. Fibonacci Quarterly, 6(3):86–93, June 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-3/lind.pdf.

Lind:1970:FC

[Lin70] D. A. Lind. A Fibonacci circulant. Fibonacci Quarterly, 8(5): 449–455, December 1970. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/8-5/lind.pdf.

Lind:1971:DIG

[Lin71] D. A. Lind. A determinant involving generalized binomial coefficients. Fibonacci Quarterly, 9(2):113–119, April 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-2/lind-a.pdf.

Lindle:1975:EMI

[Lin75] Sam Lindle. Exponential modular identity elements and the generalized last digit problem. Fibonacci Quarterly, 13(2):162–170, April 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/13-2/lindle.pdf.

Lin:1984:GSD

[Lin84] Pin-Yen Lin. The general solution to the decimal fraction of Fibonacci series. Fibonacci Quarterly, 22(3):229–234, August 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-3/lin.pdf.

Lin:1988:MTI

[Lin88] Pin-Yen Lin. De Moivre-type identities for the Tribonacci numbers. Fibonacci Quarterly, 26(2):131–134, May 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/26-2/lin.pdf.

Lin:1990:RDR

[Lin90] Pin-Yen Lin. Repeating decimals represented by Tribonacci sequences appearing from left to right or from right to left. Fibonacci Quarterly, 28(2):129–139, May 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-2/
lin.pdf.

Lin:1996:TBM

[Lin96] Chyi-Lung Lin. On triangular and Baker’s maps with golden mean as the parameter value. Fibonacci Quarterly, 34(5):423– 435, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/34-5/lin.pdf.

Lin:1998:ODF

[Lin98] Chyi-Lung Lin. Obtaining dividing formulas n|Q(n) from iterated maps. Fibonacci Quarterly, 36(2):118–124, May 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-2/lin.pdf.

Lin:2021:VBE

[Lin21] Xianzu Lin. Values of Bernoulli and Euler polynomials at rational points. Fibonacci Quarterly, 59(1):78–??, February 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/59-1/lin.pdf.

Liptai:2004:FBN

[Lip04] Kálmán Liptai. Fibonacci balancing numbers. Fibonacci Quarterly, 42(4):330–340, November 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
4/quartliptai04_2004.pdf.

LitadaSilva:2020:SLR

[Lit20] João Lita da Silva. On some linear recurrences. Fibonacci Quarterly, 58(1):73–??, February 2020. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/58-1/
dasilva.pdf.

Liu:1992:MMS

[Liu92] Bolian Liu. A matrix method to solve linear recurrences with constant coefficients. Fibonacci Quarterly, 30(1):2–8, February 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/30-1/liu.pdf.

Liu:1996:VPG

[Liu96] Bolian Liu. Varol’s permutation and its generalization. Fibonacci Quarterly, 34(2):108–117, May 1996. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/34-2/liu.
pdf.

Liu:2001:ICI

[Liu01] Guodong Liu. Identities and congruences involving higher-order Euler–Bernoulli numbers and polynomials. Fibonacci Quarterly, 39(3):279–284, June 2001. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/39-3/liu.pdf.

Liu:2002:FCF

[Liu02] Guodong Liu. Formulas for convolution Fibonacci numbers and polynomials. Fibonacci Quarterly, 40(4):352–357, August 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-4/liu.pdf.

Liu:2005:CEN

[Liu05] Guodong Liu. On congruences of Euler numbers modulo an odd square. Fibonacci Quarterly, 43(2):132–136, May 2005. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/43-2/paper43-2-7.pdf.

Liu:2007:SCF

[Liu07] Guodong Liu. Some computational formulas for Nörlund numbers. Fibonacci Quarterly, 45(2):133–137, May 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/45-2/liu.pdf.

Liu:2008:IIL

[Liu09] Guodong Liu. An identity involving the Lucas numbers and Stirling numbers. Fibonacci Quarterly, 46/47(2):136–139, May 2008/ 2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/46_47-2/liu.pdf.

Long:1967:LAS

[LJ67] C. T. Long and J. H. Jordan. A limited arithmetic on simple continued fractions. Fibonacci Quarterly, 5(2):113–128, April 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-2/long.pdf.

Long:1970:LAS

[LJ70] C. T. Long and J. H. Jordan. A limited arithmetic on simple continued fractions — II. Fibonacci Quarterly, 8(2):135–157, March 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-2/long.pdf.

Ligh:1982:GFM

[LJ82] Steve Ligh and Pat Jones. Generalized Fermat and Mersenne numbers. Fibonacci Quarterly, 20(1):12–15, February 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-1/ligh.pdf.

Loh:1997:DLB

[LJ97] Peter K. K. Loh and Hsu Wen Jing. Dynamic load balancing on multiprocessor networks. International Journal of Computer Systems Science and Engineering, 12(6):369–372, November 1997. CODEN CSSEEI. ISSN 0267-6192.

Lee:2003:LAG

[LK03a] Gwang-Yeon Lee and Jin-Soo Kim. The linear algebra of the generalized Fibonacci matrices. Fibonacci Quarterly, 41(5):451– 465, November 2003. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/41-5/lee-kim.pdf.

Lee:2003:LAF

[LK03b] Gwang-Yeon Lee and Jin-Soo Kim. The linear algebra of the k-Fibonacci matrix. Linear Algebra and its Applications, 373(1): 75–87, November 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Laohakosol:2008:RSG

[LK09] Vichian Laohakosol and Kantaphon Kuhapatanakul. Reciprocal sums of generalized second order recurrence sequences. Fibonacci Quarterly, 46/47(4):316–325, November 2008/2009. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/46_47-4/laohakosol.pdf.

Long:2010:EGS

[LK10] Calvin Long and Edward Korntved. Extension of the GCD Star of David Theorem to more than two GCDs. Fibonacci Quarterly, 48(4):312–316, November 2010. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/48-4/long.
pdf.

Lee:2003:GFF

[LKC03] Gwang-Yeon Lee, Jin-Soo Kim, and Tae Ho Cho. Generalized Fibonacci functions and sequences of generalized Fibonacci functions. Fibonacci Quarterly, 41(2):108–121, May 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/41-2/lee.pdf.

Lee:2002:FEF

[LKL02] Gwang-Yeon Lee, Jin-Soo Kim, and Sang-Gu Lee. Factorizations and eigenvalues of Fibonacci and symmetric Fibonacci matrices. Fibonacci Quarterly, 40(3):203–211, June 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-3/lee.pdf.

Lehmer:1983:PPH

[LL83] D. H. Lehmer and Emma Lehmer. Properties of polynomials having Fibonacci numbers for coefficients. Fibonacci Quarterly, 21(1):62–64, February 1983. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/21-1/lehmer.pdf.

Lombardi:1984:GMS

[LL84] Oreste W. Lombardi and Margaret A. Lombardi. The golden mean in the solar system. Fibonacci Quarterly, 22(1):70–75, February 1984. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/22-1/lombardi.pdf.

Lavertu:1985:BCI

[LL85] Marie-Louis Lavertu and Claude Levesque. On Bernstein’s combinatorial identities. Fibonacci Quarterly, 23(4):347–355, November 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/23-4/lavertu.pdf.

Lee:1987:CCP

[LL87a] Jin-Zai Lee and Jia-Sheng Lee. A complete characterization of B-power fractions that can be represented as series of general n-Bonacci numbers. Fibonacci Quarterly, 25(1):72–75, February 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/25-1/lee.pdf.

Lee:1987:SPB

[LL87b] Jin-Zai Lee and Jia-Sheng Lee. Some properties of binomial coefficients. Fibonacci Quarterly, 25(4):339–342, November 1987.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-4/lee1.pdf.

Lee:1987:SPG

[LL87c] Jin-Zai Lee and Jia-Sheng Lee. Some properties of the generalization of the Fibonacci sequence. Fibonacci Quarterly, 25(2): 111–117, May 1987. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/25-2/lee.pdf.

Lee:1987:SPS

[LL87d] Jin-Zai Lee and Jia-Sheng Lee. Some properties of the sequence {W_{n}(a, b; p, q)}. Fibonacci Quarterly, 25(3):268–278, August 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/25-3/lee-a.pdf.

Lee:1988:NGF

[LL88] Jin-Zai Lee and Jia-Sheng Lee. A note on the generalized Fibonacci numbers. Fibonacci Quarterly, 26(1):14–19, February 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/26-1/lee.pdf.

Lee:1995:NGF

[LL95] Gwang-Yeon Lee and Sang-Gu Lee. A note on generalized Fibonacci numbers. Fibonacci Quarterly, 33(3):273–278, June 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-3/lee.pdf.

Liu:2005:SII

[LL05] Guodong Liu and Hui Luo. Some identities involving Bernoulli numbers. Fibonacci Quarterly, 43(3):208–212, August 2005. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/43-3/paper43-3-3.pdf.

Lang:2013:FNI

[LL13] Cheng Lien Lang and Mong Lung Lang. Fibonacci numbers and identities. Fibonacci Quarterly, 51(4):330–??, November 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/51-4/lang.pdf.

Laishram:2014:FNF

[LL14] Shanta Laishram and Florian Luca. Fibonacci numbers of the form xa ± xb ± 1. Fibonacci Quarterly, 52(4):290–??, November 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/52-4/laishram.pdf.

Lee:2001:BFR

[LLKS01] Gwang-Yeon Lee, Sang-Gu Lee, Jin-Soo Kim, and Hang-Kyun Shin. The Binet formula and representations of k-generalized Fibonacci numbers. Fibonacci Quarterly, 39(2):158–164, May 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-2/lee.pdf.

Li:2020:DZG

[LLM^{+}20] Ruoci Li, Xiaonan Li, Steven J. Miller, Clayton Mizgerd, Chenyang Sun, Dong Xia, and Zhyi Zhou. Deterministic Zeckendorf games. Fibonacci Quarterly, 58(5):152–??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/58-5/li.pdf.

Lai:2018:EQK

[LLP^{+}18] Hong Lai, Mingxing Luo, Josef Pieprzyk, Zhiguo Qu, and Mehmet A. Orgun. Efficient quantum key distribution using Fibonacci-number coding with a biased basis choice. Information Processing Letters, 134(??):24–30, June 2018. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0020019018300267.

Lee:1997:GFM

[LLS97] G.-Y. Lee, S.-G. Lee, and H.-G. Shin. On the K-generalized Fibonacci matrix Q_{k}^{∗}. Linear Algebra and its Applications, 251(1– 3):73–88, January 15, 1997. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.elsevier.

com/cgi-bin/cas/tree/store/laa/cas_sub/browse/browse.
cgi?year=1997&volume=251&issue=1-3&aid=9500553.

Li:1971:CGD

[LM71] H. Y. Li and Sibley S. Morrill. I ching games of Duke Tan of Chou and C. C. T’ung. Cadleon Press, San Francisco, CA, USA, 1971. ISBN 0-9600310-2-2. 138 pp. LCCN GV1507.T3 L5.

Lord:1979:DTC

[LM79] Graham Lord and Herve G. Morin. Degeneracy of transformed complete sequences. Fibonacci Quarterly, 17(4):358–360, December 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/17-4/lord.pdf.

Laywine:1985:LCH

[LM85] Charles Laywine and Gary L. Mullen. Latin cubes and hypercubes of prime order. Fibonacci Quarterly, 23(2):139–145, May 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/23-2/laywine.pdf.

Lidl:1993:PTL

[LM93] R. Lidl and W. B. Muller. Primality testing with Lucas functions. Lecture Notes in Computer Science, 718:539–??, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:2010:ECF

[LM10] Yan Li and Lianrong Ma. On the elements of the continued fractions of quadratic irrationals. Fibonacci Quarterly, 48(2):129– 136, May 2010. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/48-2/li.pdf.

Li:2017:CCL

[LM17] Ray Li and Steven J. Miller. A collection of central limit type results in generalized Zeckendorf decompositions. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/55-5/li.pdf.

Li:2019:CLT

[LM19a] Ray Li and Steven J. Miller. Central limit theorems for gaps of generalized Zeckendorf decompositions. Fibonacci Quarterly, 57 (3):213–??, August 2019. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/57-3/li.pdf.

Luca:2019:DER

[LM19b] Florian Luca and Sibusiso Mabaso. Diophantine equations with the Ramanujan τ function of factorials, Fibonacci numbers, and Catalan numbers. Fibonacci Quarterly, 57(3):255–??, August 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/57-3/luca.pdf.

Lucheta:1996:DPM

[LMR96] Caroline Lucheta, Eli Miller, and Clifford Reiter. Digraphs from powers modulo p. Fibonacci Quarterly, 34(3):226–239, June 1996.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-3/lucheta.pdf.

Laskari:2009:ANI

[LMV09] E. C. Laskari, G. C. Meletiou, and M. N. Vrahatis. Aitken and Neville inverse interpolation methods for the Lucas logarithm problem. Applied Mathematics and Computation, 209(1):52–56, March 1, 2009. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ligomenides:1981:CRF

[LN81] P. Ligomenides and R. Newcomb. Complement representations in the Fibonacci computer. In IEEE [IEE81], pages 6–9.

LCCN QA 76.6 S985t 1981. URL http://www.acsel-lab.com/
arithmetic/arith5/papers/ARITH5_Ligomenides_Newcomb.pdf.

IEEE catalog number 81CH1630-C.

Ligomenides:1984:MFC

[LN84] P. Ligomenides and R. Newcomb. Multilevel Fibonacci conversion and addition. Fibonacci Quarterly, 22(3):196–203, August 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-3/ligomenides.pdf.

Lim:2007:MCA

[LN07] Chjan Lim and Joseph Nebus. The Monte Carlo Approach, chapter 4, pages 51–65. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 0-387-35075-6, 0-387-49431-6 (e-book). LCCN QA911 .L466 2007.

Luis:2014:PIM

[LO14] Rafael Luís and Henrique M. Oliveira. Products of 2×2 matrices related to non autonomous Fibonacci difference equations. Applied Mathematics and Computation, 226(??):101–116, January 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300313010552.

Larcombe:2018:GFA

[LO18] Peter J. Larcombe and Sam T. O’Neill. A generating function approach to the automated evaluation of sums of exponentiated multiples of generalized Catalan number linear combinations. Fibonacci Quarterly, 56(2):121–??, May 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/56-
2/larcombe.pdf.

Larcombe:2014:CSE

[LOF14] Peter J. Larcombe, Sam T. O’Neill, and Eric J. Fennessey. On certain series expansions of the sine function: Catalan numbers and convergence. Fibonacci Quarterly, 52(3):236–??, August 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/52-3/larcombe.pdf.

Logothetti:1981:ITN

[Log81] Dave Logothetti. An implicit triangle of numbers. Fibonacci Quarterly, 19(3):276–279, August 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
3/logothetti.pdf.

Long:1973:ABC

[Lon73] Calvin T. Long. Arrays of binomial coefficients whose products are squares. Fibonacci Quarterly, 11(5):449–456, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-5/long.pdf.

Long:1981:DER

[Lon81a] C. T. Long. The decimal expansion of 1/89 and related results. Fibonacci Quarterly, 19(1):53–55, February 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-1/long.pdf.

Long:1981:LAS

[Lon81b] Calvin T. Long. A limited arithmetic on simple continued fractions — III. Fibonacci Quarterly, 19(2):163–174, April 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-2/long.pdf.

Long:1981:PTM

[Lon81c] Calvin T. Long. Pascal’s triangle modulo p. Fibonacci Quarterly, 19(5):458–462, December 1981. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/19-5/long.pdf.

Long:1981:SDP

[Lon81d] Calvin T. Long. Some divisibility properties of Pascal’s triangle. Fibonacci Quarterly, 19(3):257–263, August 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-3/long.pdf.

Long:1985:FAT

[Lon85] Calvin T. Long. On a Fibonacci arithmetical trick. Fibonacci Quarterly, 23(3):221–231, August 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-
3/long.pdf.

Long:1986:DFI

[Lon86a] Calvin T. Long. Discovering Fibonacci identities. Fibonacci Quarterly, 24(2):160–167, May 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-
2/long.pdf.

Long:1986:NMP

[Lon86b] Calvin T. Long. A note on Moessner’s process. Fibonacci Quarterly, 24(4):349–355, November 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-
4/long.pdf.

Loria:1929:LF

[Lor29] G. Loria. Leonardo Fibonacci. Storia delle mathematiche I (Turin), ??(??):379–410, ???? 1929.

Lord:1973:COV

[Lor73] Graham Lord. Counting omitted values. Fibonacci Quarterly, 11(4):443–448, November 1973. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/11-4/lord.pdf.

Luca:2023:TNP

[LOT23] Florian Luca, Japhet Odjoumani, and Alain Togbé. Tribonacci numbers that are products of two Fibonacci numbers. Fibonacci Quarterly, 61(4):298–??, November 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/61-
4/luca.pdf.

Lowman:1971:SSP

[Low71a] Edward A. Lowman. Some striking proportions in the music of Bela Bartók. Fibonacci Quarterly, 9(5):527–528, December 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-5/lowman-a.pdf.

Lowman:1971:EFN

[Low71b] Edward L. Lowman. An example of Fibonacci numbers used to generate rhythmic values in modern music. Fibonacci Quarterly, 9(4):423–426, October 1971. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/9-4/lowman-a.pdf.

Liverance:1996:DBM

[LP96] Eric Liverance and John Pitsenberger. Diagonalization of the binomial matrix. Fibonacci Quarterly, 34(1):55–67, February 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-1/liverance.pdf.

Lee:2003:SP

[LP03a] HoKyu Lee and SeungKyung Park. The r-subcomplete partitions. Fibonacci Quarterly, 41(5):386–396, November 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-5/lee-park.pdf.

Luca:2003:MGG

[LP03b] Florian Luca and Štefan Porubský. The multiplicative group generated by the Lehmer numbers. Fibonacci Quarterly, 41(2): 122–132, May 2003. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/41-2/luca.pdf.

Luca:2014:LBO

[LP14] Florian Luca and Carl Pomerance. On the local behavior of the order of appearance in the Fibonacci sequence. International Journal of Number Theory (IJNT), 10(4):915–933, June 2014. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://

www.worldscientific.com/doi/10.1142/S1793042114500079.

Luca:2018:PPS

[LP18] Florian Luca and Vandita Patel. On perfect powers that are sums of two Fibonacci numbers. Journal of Number Theory, 189(??): 90–96, August 2018. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X18300520.

Liptai:2016:BPB

[LPS16] K. Liptai, G. K. Panda, and L. Szalay. A balancing problem on a binary recurrence and its associate. Fibonacci Quarterly, 54(3):

235–??, August 2016. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/54-3/liptai.pdf.

Luca:2011:FI

[LPW11] Florian Luca, Carl Pomerance, and Stephan Wagner. Fibonacci integers. Journal of Number Theory, 131(3):440–457, March 2011. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X10002507.

Laohakosol:1985:TOA

[LR85] Vichian Laohakosol and Nit Roenrom. A third-order analog of a result of L. Carlitz. Fibonacci Quarterly, 23(3):194–198, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/23-3/laohakosol.pdf.

Luca:2006:LSW

[LS06] Florian Luca and Lawrence Somer. Lucas sequences for which 4|ϕ(|u_{n}|) for almost all n. Fibonacci Quarterly, 44(3):249–262, August 2006. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/44-3/luca.pdf.

Luca:2007:FNF

[LS07] Florian Luca and Laszlo Szalay. Fibonacci numbers of the form p^{a} ± p^{b} + 1. Fibonacci Quarterly, 45(2):98–103, May 2007. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/45-2/luca.pdf.

Luca:2013:CFT

[LS13] Florian Luca and Laszlo Szalay. On the counting function of triples whose pairwise products are close to Fibonacci numbers. Fibonacci Quarterly, 51(3):228–??, August 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/51-3/luca.pdf.

Luca:2014:FNW

[LS14a] Florian Luca and Pantelimon Stănică. On Fibonacci numbers which are elliptic Korselt numbers. Fibonacci Quarterly, 52(5): 164–??, December 2014. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers1/52-5/Luca.pdf.

Luca:2014:FDF

[LS14b] Florian Luca and Pantelimon Stănică. On the first digits of the Fibonacci numbers and their Euler function. Uniform Distribution Theory, 9(1):21–25, 2014. ISSN 1336-913X (print), 2309-

5377 (electronic). URL https://math.boku.ac.at/udt/vol09/
no1/03LucaStanica.pdf.

Luca:2018:MEF

[LS18a] Florian Luca and Anitha Srinivasan. Markov equation with Fibonacci components. Fibonacci Quarterly, 56(2):126–??, May 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/56-2/luca.pdf.

Lynd:2018:SGP

[LS18b] Chris D. Lynd and James Wright Sharpe. Sequences generated by powers of the k-th-order Fibonacci recurrence relation. American Mathematical Monthly, 125(5):443–446, 2018. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Long:2007:EGS

[LSA07] Calvin Long, William C. Schulz, and Shiro Ando. An extension of the GCD Star of David theorem. Fibonacci Quarterly, 45(3): 194–201, August 2007. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/45-3/long.pdf.

Lewis:1976:FSM

[LSS76] T. G. Lewis, B. J. Smith, and M. Z. Smith. Fibonacci sequences and memory management. Fibonacci Quarterly, 14(1):37–41, February 1976. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/14-1/lewis.pdf.

Leiserson:2012:DPR

[LSS12] Charles E. Leiserson, Tao B. Schardl, and Jim Sukha. Deterministic parallel random-number generation for dynamicmultithreading platforms. ACM SIGPLAN Notices, 47(8):193– 204, August 2012. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Lindroos:2014:OFN

[LSW14] Linus Lindroos, Andrew Sills, and Hua Wang. Odd Fibbinary numbers and the Golden Ratio. Fibonacci Quarterly, 52(1):61– ??, February 2014. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/52-1/lindroos.pdf.

Luca:2007:DE

[LT07] Florian Luca and Alain Togbé. On the Diophantine equation x^{2} + 7^{2k} = y^{n}. Fibonacci Quarterly, 45(4):322–326, November 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/45-4/luca.pdf.

Laih:1995:SLF

[LTT95] Chi Sung Laih, Fu Kuan Tu, and Wen Chung Tai. On the security of the Lucas function. Information Processing Letters, 53(5):243– 247, March 10, 1995. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lu:1970:SND

[Lu70] K. U. Lu. The smallest number with divisors a product of distinct primes. Fibonacci Quarterly, 8(4):380–382, October 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-4/lu.pdf.

Latushkin:2005:RRS

[LU05] Yaroslav Latushkin and Vladimir Ushakov. A representation of regular subsequences of recurrent sequences. Fibonacci Quarterly, 43(1):70–84, February 2005. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Papers/43-1/paper43-1-
9.pdf.

Lucas:1973:NCT

[Luc73] Dianne Smith Lucas. Numbers common to two polygonal sequences. Fibonacci Quarterly, 11(1):78–84, February 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-1/lucas.pdf.

Luca:1999:AFF

[Luc99] Florian Luca. Arithmetic functions of Fibonacci numbers. Fibonacci Quarterly, 37(3):265–268, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-3/luca.pdf.

Luca:2000:EIA

[Luc00] Florian Luca. Equations involving arithmetic functions of Fibonacci and Lucas numbers. Fibonacci Quarterly, 38(1):49–55, February 2000. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/38-1/luca.pdf.

Luca:2002:CBC

[Luc02] Florian Luca. Consecutive binomial coefficients in Pythagorean triples and squares in the Fibonacci sequences. Fibonacci Quarterly, 40(1):76–78, February 2002. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/40-1/luca.pdf.

Luca:2003:APSa

[Luc03a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 41(2):187–193, May 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-
2/advanced41-2.pdf.

Luca:2003:APSb

[Luc03b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 41(4):380–383, August 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-
4/advanced41-4.pdf.

Luca:2003:APSc

[Luc03c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 41(5):472–477, November 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-
5/advanced41-5.pdf.

Luca:2003:PNW

[Luc03d] Florian Luca. On positive numbers n for which Ω(n) divides F_{n}. Fibonacci Quarterly, 41(4):365–371, August 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-4/luca.pdf.

Luca:2004:APSa

[Luc04a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 42(1):92–97, February 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
1/February2004advanced.pdf.

Luca:2004:APSb

[Luc04b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 42(2):187–193, May 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
2/May2004advancednew.pdf.

Luca:2004:APSc

[Luc04c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 42(3):283–289, August 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
3/August2004advanced.pdf.

Luca:2004:APSd

[Luc04d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 42(4):377–382, November 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
4/November2004advanced.pdf.

Luca:2005:APSa

[Luc05a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 43(1):91–97, February 2005. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Problems/advanced43-1.
pdf.

Luca:2005:APSb

[Luc05b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 43(2):187–193, May 2005. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Problems/advanced43-2.
pdf.

Luca:2005:APSc

[Luc05c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 43(3):283–289, August 2005. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
advanced43-3.pdf.

Luca:2005:APSd

[Luc05d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 43(4):377–382, November 2005. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
advanced43-4.pdf.

Luca:2006:APSa

[Luc06a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 44(1):91–97, February 2006. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2006advanced.pdf.

Luca:2006:APSb

[Luc06b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 44(2):187–193, May 2006. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
May2006advanced.pdf.

Luca:2006:APSc

[Luc06c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 44(3):283–289, August 2006. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
August2006advanced.pdf.

Luca:2006:APSd

[Luc06d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 44(4):376–382, November 2006. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
November2006advanced.pdf.

Luca:2007:APSa

[Luc07a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 45(1):91–97, February 2007. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2007advancednewversion.pdf.

Luca:2007:APSb

[Luc07b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 45(2):187–193, May 2007. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Problems/advanced45-2.
pdf.

Luca:2007:APSc

[Luc07c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 45(3):283–289, August 2007. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
advanced45-3.pdf.

Luca:2007:APSd

[Luc07d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 45(4):376–382, November 2007. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
November2007advanced.pdf.

Luca:2008:APSa

[Luc09a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 46/47(1):91–97, February 2008/2009. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
Feb2008advanced.pdf.

Luca:2008:APSb

[Luc09b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 46/47(2):186–193, May 2008/2009. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
May2008advanced.pdf.

Luca:2008:APSc

[Luc09c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 46/47(3):283–289, August 2008/2009. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
Aug2009advanced.pdf.

Luca:2008:APSd

[Luc09d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 46/47(4):374–380, November 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Problems/Advanced_Problems_NOV2009.pdf.

Luca:2010:APSa

[Luc10a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 48(1):89–97, February 2010. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Problems/Adv_Prob_Feb_
2010.pdf.

Luca:2010:APSb

[Luc10b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 48(2):184–191, May 2010. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
May2010advanced.pdf.

Luca:2010:APSc

[Luc10c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 48(3):283–289, August 2010. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
August2010advanced.pdf.

Luca:2010:APSd

[Luc10d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 48(4):373–380, November 2010. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
November2010advanced.pdf.

Luca:2011:APSa

[Luc11a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 49(1):89–97, February 2011. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2011advanced.pdf.

Luca:2011:APSb

[Luc11b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 49(2):187–193, May 2011. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
May2011advanced.pdf.

Luca:2011:APSc

[Luc11c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 49(3):281–??, August 2011. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
August2011advanced.pdf.

Luca:2011:APSd

[Luc11d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 49(4):374–??, November 2011. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
November2011advanced.pdf.

Luca:2012:APSa

[Luc12a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 50(1):89–??, February 2012. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2012advanced.pdf.

Luca:2012:APSb

[Luc12b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 50(2):187–??, May 2012. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
May2012advanced.pdf.

Luca:2012:APSc

[Luc12c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 50(3):280–??, August 2012. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
2012AugustAdvanced.pdf.

Luca:2012:APSd

[Luc12d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 50(4):374–??, November 2012. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
November2012advanced.pdf.

Luca:2013:APSa

[Luc13a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 51(1):91–??, February 2013. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2013advanced.pdf.

Luca:2013:APSb

[Luc13b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 51(2):186–??, May 2013. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
May2013advanced.pdf.

Luca:2013:APSc

[Luc13c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 51(3):282–??, August 2013. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
August2013advanced.pdf.

Luca:2013:APS

[Luc13d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 51(4):374–??, November 2013. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
November2013advanced.pdf.

Luca:2014:APSa

[Luc14a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 52(1):87–??, February 2014. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2014advanced.pdf.

Luca:2014:APSb

[Luc14b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 52(2):185–??, May 2014. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
MayAdvanced2014.pdf.

Luca:2014:APSc

[Luc14c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 52(3):281–??, August 2014. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvancedAug2014.pdf.

Luca:2014:APS

[Luc14d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 52(4):374–??, November 2014. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
Adv2014Nov.pdf.

Luca:2015:APSa

[Luc15a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 53(1):88–??, February 2015. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Problems/AdvFeb2015.
pdf.

Luca:2015:APSb

[Luc15b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 53(2):186–??, May 2015. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
May2015advanced.pdf.

Luca:2015:APSc

[Luc15c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 53(3):279–??, August 2015. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AugustAdv2015.pdf.

Luca:2015:APSd

[Luc15d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 53(4):372–??, November 2015. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
November2015advanced.pdf.

Luca:2016:APSa

[Luc16a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 54(1):87–??, February 2016. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
FebAdvProbSoln2016.pdf.

Luca:2016:APSb

[Luc16b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 54(2):185–??, May 2016. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbSlnMay2016.pdf.

Luca:2016:APSc

[Luc16c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 54(3):280–??, August 2016. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
August2016AdvProbSln.pdf.

Luca:2016:APSd

[Luc16d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 54(4):373–??, November 2016. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
November2016advanced.pdf.

Luca:2017:APSa

[Luc17a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 55(1):89–??, February 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2017advanced.pdf.

Luca:2017:APSb

[Luc17b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 55(2):184–??, May 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
May2017advanced.pdf.

Luca:2017:APSc

[Luc17c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 55(3):282–??, August 2017. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Problems/AdvPSAug2017.
pdf.

Luca:2017:APS

[Luc17d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 55(4):374–??, November 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
November2017advanced.pdf.

Luca:2018:APSa

[Luc18a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 56(1):89–??, February 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2018AdvancedProblemREV.pdf.

Luca:2018:APSb

[Luc18b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 56(2):185–??, May 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbMay2018.pdf.

Luca:2018:APSc

[Luc18c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 56(3):283–??, August 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
August2018AdvancedREV.pdf.

Luca:2018:APSd

[Luc18d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 56(4):373–??, November 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
Nov2018AdvProbSectREV.pdf.

Luca:2019:APSa

[Luc19a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 57(1):89–??, February 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2019AdvancedREV.pdf.

Luca:2019:APSb

[Luc19b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 57(2):184–??, May 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbMay2019.pdf.

Luca:2019:APSc

[Luc19c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 57(3):283–??, August 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
August2019AdvancedRevRev.pdf.

Luca:2019:APSd

[Luc19d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 57(4):373–??, November 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
Nov2019AdvProb.pdf.

Luca:2020:APSa

[Luc20a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 58(1):89–??, February 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
February2020AdvancedRev.pdf.

Luca:2020:APSb

[Luc20b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 58(2):185–??, May 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbMay2020.pdf.

Luca:2020:APSc

[Luc20c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 58(3):281–??, August 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
August2020AdvancedProblems.pdf.

Luca:2020:APSd

[Luc20d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 58(4):375–??, November 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbNov2020.pdf.

Luca:2021:APSa

[Luc21a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 59(1):89–??, February 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
Feb2021AdvProb.pdf.

Luca:2021:APSb

[Luc21b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 59(2):185–??, May 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbMay2021.pdf.

Luca:2021:APSc

[Luc21c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 59(3):280–??, August 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
August2021AdvancedProblems.pdf.

Luca:2021:APSd

[Luc21d] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 59(4):373–??, November 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbNov2021.pdf.

Luca:2021:DGF

[Luc21e] Florian Luca. On the discriminant of the k-generalized Fibonacci polynomial. Fibonacci Quarterly, 59(4):298–??, November 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/59-4/luca.pdf.

Luca:2022:APSa

[Luc22a] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 60(1):90–??, February 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
FebAdvProb2022.pdf.

Luca:2022:APSb

[Luc22b] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 60(2):185–??, May 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbMay2022.pdf.

Luca:2022:APSc

[Luc22c] Florian Luca. Advanced problems and solutions. Fibonacci Quarterly, 60(3):281–??, August 2022. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Problems/
AdvProbAug2022.pdf.

Ludington:1987:TIA

[Lud87] Anne L. Ludington. Transposable integers in arbitrary bases. Fibonacci Quarterly, 25(3):263–267, August 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-3/ludington.pdf.

Ludington:1988:GTI

[Lud88a] Anne L. Ludington. Generalized transposable integers. Fibonacci Quarterly, 26(1):58–63, February 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
1/ludington.pdf.

Ludington:1988:LNG

[Lud88b] Anne L. Ludington. Length of the 7-number game. Fibonacci Quarterly, 26(3):195–204, August 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
3/ludington.pdf.

Lunnon:1977:PM

[Lun77] W. Fred Lunnon. The Pascal matrix. Fibonacci Quarterly, 15 (3):201–204, October 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-3/lunnon.pdf.

Luneburg:1991:FAKa

[Lün91a] H. Lüneburg. Fibonaccis aufsteigende Kettenbrüche, ein elegantes Werkzeug mittelalterlicher Rechenkunst. (German) [Fibonacci’s ascending continued fractions, an elegant tool of medieval arithmetic]. In ????, editor, Séminaire Lotharingien de Combinatoire (Strasbourg, 1991), pages 135–149. ????, ????, 1991. ISBN ???? LCCN ????

Luneburg:1991:FAKb

[Lün91b] H. Lüneburg. Fibonaccis aufsteigende Kettenbrüche, ein elegantes Werkzeug mittelalterlicher Rechenkunst. (German) [Fibonacci’s ascending continued fractions, an elegant tool of medieval arithmetic]. Sudhoffs Arch., 75(2):129–139, ???? 1991.

Luneburg:1992:LPL

[Lün92] Heinz Lüneburg. Leonardi Pisani Liber Abbaci oder Lesevergnügen eines Mathematikers. (German) [Leonardi Pisani Liber Abbaci or a mathematician’s reading pleasure]. BI-Wiss.- Verl., Mannheim, Germany, 1992. ISBN 3-411-15461-6. 340

pp. LCCN ???? URL http://www.gbv.de/dms/hbz/toc/
ht004250215.PDF;   https://www.math.utah.edu/pub/tex/
bib/fibquart.bib;   http://www.zentralblattmath.org/zmath/
en/search/?an=0755.01015.

Luo:1996:DEX

[Luo96] Ming Luo. On the Diophantine equation (x(x − 1)/2)^{2} = (y(y − 1)/2)^{2}. Fibonacci Quarterly, 34(3):277–279, June 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-3/luo.pdf.

Lagarias:1981:FLC

[LW81] J. C. Lagarias and D. P. Weisser. Fibonacci and Lucas cubes. Fibonacci Quarterly, 19(1):39–42, February 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-1/lagarias.pdf.

Ligh:1987:FNU

[LW87] Steve Ligh and Charles R. Wall. Functions of non-unitary divisors. Fibonacci Quarterly, 25(4):333–338, November 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-4/ligh.pdf.

Liu:2021:SPF

[LX21] Zhixin Liu and Mengyuan Xue. The sum of a prime and a Fibonacci number. International Journal of Number Theory (IJNT), 17(08):1815–1823, September 2021. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042121500640.

Ludington-Young:1990:LNG

[LY90] Anne Ludington-Young. Length of the n-number game. Fibonacci Quarterly, 28(3):259–265, August 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
3/ludington.pdf.

Ludington-Young:1999:EDS

[LY99] Anne L. Ludington-Young. Even Ducci-sequences. Fibonacci Quarterly, 37(2):145–153, May 1999. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/37-
2/ludington.pdf.

Leida:2002:PDA

[LY02] Johann Leida and Yongzhi (Peter) Yang. Pascal decompositions of arithmetic and convolution arrays in matrices. Fibonacci Quarterly, 40(2):136–145, May 2002. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/40-2/
leida.pdf.

Yang:2008:GFN

[lY08] Sheng liang Yang. On the k-generalized Fibonacci numbers and high-order linear recurrence relations. Applied Mathematics and Computation, 196(2):850–857, March 1, 2008. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Luca:2013:FEC

[LY13] Florian Luca and Aynur Yalçiner. L-functions of elliptic curves and Fibonacci numbers. Fibonacci Quarterly, 51(2):112–??, May 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/51-2/luca.pdf.

Lynch:1970:FNP

[Lyn70] W. C. Lynch. The t-Fibonacci numbers and polyphase sorting. Fibonacci Quarterly, 8(1):6–22, February 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/8-1/lynch.pdf.

Lynch:1973:FRS

[Lyn73] Michael F. Lynch. A Fibonacci-related series in an aspect of information retrieval. Fibonacci Quarterly, 11(5):495–499, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/11-5/lynch.pdf.

Ma:1998:GKI

[Ma98] Xinrong Ma. A generalization of the Kummer identity and its application to Fibonacci–Lucas sequences. Fibonacci Quarterly, 36(4):339–347, August 1998. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/36-4/ma.pdf.

Ma:2011:IIG

[Ma11] Shi-Mei Ma. Identities involving generalized Fibonacci-type polynomials. Applied Mathematics and Computation, 217(22):9297– 9301, July 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300311005674.

Mabry:2019:FNI

[Mab19] Rick Mabry. Fibonacci numbers, integer compositions, and nets of antiprisms. American Mathematical Monthly, 126(9):786–801, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

MacLeod:1988:GRL

[Mac88] R. A. MacLeod. Generalization of a result of E. Lucas. Canadian Mathematical Bulletin = Bulletin canadien de mathématiques, 31(??):95–98, ???? 1988. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

MacHenry:2000:FF

[Mac00a] T. MacHenry. Fibonacci fields. Fibonacci Quarterly, 38(1):17– 24, February 2000. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/38-1/machenry.pdf.

MacHenry:2000:GFL

[Mac00b] Trueman MacHenry. Generalized Fibonacci and Lucas polynomials and multiplicative arithmetic functions. Fibonacci Quarterly, 38(2):167–173, May 2000. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/38-2/machenry.pdf.

Macfarlane:2010:UDP

[Mac10] A. J. Macfarlane. Use of determinants to present identities involving Fibonacci and related numbers. Fibonacci Quarterly, 48(1): 68–76, February 2010. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/48-1/macfarlane.pdf.

Madachy:1968:RMa

[Mad68a] Joseph S. Madachy. Recreational mathematics. Fibonacci Quarterly, 6(1):60–68, February 1968. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/6-1/madachy.
pdf.

Madachy:1968:RMb

[Mad68b] Joseph S. Madachy. Recreational mathematics. Fibonacci Quarterly, 6(2):162–166, April 1968. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/6-2/madachy.
pdf.

Madachy:1968:RMc

[Mad68c] Joseph S. Madachy. Recreational mathematics. Fibonacci Quarterly, 6(4):299–??, October 1968. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/6-4/madachy.
pdf.

Madachy:1968:RMd

[Mad68d] Joseph S. Madachy. Recreational mathematics. Fibonacci Quarterly, 6(6):385–389, December 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-6/
madachy.pdf.

Madachy:1969:RMD

[Mad69] Joseph S. Madachy. Recreational mathematics — “difference series” resulting from sieving primes. Fibonacci Quarterly, 7(3): 315–318, October 1969. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/7-3/madachy.pdf.

Madachy:1972:SNN

[Mad72] Joseph S. Madachy. Some new narcissistic numbers. Fibonacci Quarterly, 10(3):295–298, April 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-3/
madachy.pdf.

Maeder:1991:MPFa

[Mae92a] Roman E. Maeder. The Mathematica programmer: Fibonacci on the fast track. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (elec-

tronic). URL http://www.mathematica-journal.com/issue/
v1i3/tutorials/maeder/index.html;   https://www.math.
utah.edu/pub/tex/bib/fibquart.bib.

Maeder:1992:MPFa

[Mae92b] Roman E. Maeder. The Mathematica programmer: Fibonacci on the fast track. Mathematica Journal, 1(3):??, Winter 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (elec-

tronic). URL http://www.mathematica-journal.com/issue/
v1i3/tutorials/maeder/index.html;   https://www.math.
utah.edu/pub/tex/bib/fibquart.bib.

Magazine:1981:NSC

[Mag81] Michael J. Magazine. The number of states in a class of serial queueing systems. Fibonacci Quarterly, 19(1):43–44, February 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-1/magazine.pdf.

Magli:2008:IIB

[Mag09] Pierluigi Magli. Identities involving Bernoulli numbers related to sums of powers of integers. Fibonacci Quarterly, 46/47(2):140– 145, May 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/46_47-2/magli.pdf.

Mahanthappa:1991:ASF

[Mah91] Mahesh K. Mahanthappa. Arithmetic sequences and Fibonacci quadratics. Fibonacci Quarterly, 29(4):343–346, November 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-4/mahanthappa.pdf.

Maier:1970:OOC

[Mai70] Eugene A. Maier. One-one correspondences between the set N of positive integers and the sets N^{n} and ∪_{n∈N}N^{n}. Fibonacci Quarterly, 8(4):365–370, October 1970. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/8-4/maier.pdf.

Makowski:1988:SEF

[Mak88] Andrzej Makowski. Stroeker’s equation and Fibonacci numbers. Fibonacci Quarterly, 26(4):336–337, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/26-4/makowski.pdf.

Makino:1994:LFR

[Mak94] Jun Makino. Lagged-Fibonacci random number generators on parallel computers. Parallel Computing, 20(9):1357–1367, September 12, 1994. CODEN PACOEJ. ISSN 0167-8191 (print),

1872-7336 (electronic). URL http://www.elsevier.com/cgi-
bin/cas/tree/store/parco/cas_sub/browse/browse.cgi?year=
1994&volume=20&issue=9&aid=893.

Mak:2012:ADS

[Mak12] Kit-Ho Mak. An analogue of the Ducci sequences over function fields. Fibonacci Quarterly, 50(4):326–??, November 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-4/mak.pdf.

Malik:1983:SEO

[Mal83] H. N. Malik. On the solution of {E^{2} + (λp − 2)E + (1 − λp − λ^{2}q)}^{m}G_{n} = n^{k} by expansions and operators. Fibonacci Quarterly, 21(4):260–265, November 1983. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/21-
4/malik.pdf.

Mamangakis:1961:MNR

[Mam61] S. E. Mamangakis. Mathematical notes: Remarks on the Fibonacci series modulo m. American Mathematical Monthly, 68 (7):648–649, August/September 1961. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mandelson:1964:AIF

[Man64] Joseph Mandelson. Amateur interests in the Fibonacci series — prime numbers. Fibonacci Quarterly, 2(2):139–144, April 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-2/mandelson.pdf.

Mandelson:1967:AIF

[Man67] Joseph Mandelson. Amateur interests in the Fibonacci series II. calculation of Fibonacci numbers and sums from the binomial. Fibonacci Quarterly, 5(3):275–279, October 1967. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-3/mandelson.pdf.

Mandelson:1968:AIF

[Man68] Joseph Mandelson. Amateur interests in the Fibonacci series III. Residues of u_{n} with respect to any modules. Fibonacci Quarterly, 6(4):275–278, October 1968. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/6-4/mandelson.pdf.

Mann:1975:SAP

[Man75] James M. Mann. Signed b-adic partitions. Fibonacci Quarterly, 13(2):174–180, April 1975. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/13-2/mann.pdf.

Mandelson:1977:AIF

[Man77] Joseph Mandelson. Amateur interests in the Fibonacci series IV calculation of group sizes of residues of moduli. Fibonacci Quarterly, 15(2):145–149, April 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-2/
mandelson.pdf.

Manolopoulos:2005:NRC

[Man05a] Yannis Manolopoulos. On the number of recursive calls of recursive functions. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(2):61–64, June 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (elec-

tronic). URL ftp://ftp.math.utah.edu/pub/mirrors/ftp.
ira.uka.de/bibliography/Misc/DBLP/2005.bib.

Mansour:2005:GSI

[Man05b] Toufik Mansour. Generalizations of some identities involving the Fibonacci numbers. Fibonacci Quarterly, 43(4):307–315, November 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Papers/43-4/paper43-4-4.pdf.

Maniscalco:2006:IN

[Man06] C. Maniscalco. Isodecimal numbers. Fibonacci Quarterly, 44(4):341–346, November 2006. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/44-4/
maniscalco.pdf.

Mansuy:2017:FWC

[Man17] Frédéric Mansuy. Fibonacci words and the construction of a “quasicrystalline” fivefold structure. Fibonacci Quarterly, 55(5): ??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/55-5/mansuy.pdf.

Maracchia:1974:REL

[Mar74] Silvio Maracchia. La regola “Elchatayn” in Leonardo Pisano. (Italian) [The Elchatayn rule in Leonardo Pisano]. Archimede. Rivista per gli Insegnanti e i Cultori di Matematiche Pure e Applicate, 26(5–6):321–329, 1974. ISSN 0390-5543.

Marsaglia:1992:MRN

[Mar92] George Marsaglia. The mathematics of random number generators. In Burr [Bur92], pages 73–90. ISBN 0-8218-5501-8. LCCN QA241 .U67 1992.

Marchisotto:1993:CMI

[Mar93] E. A. Marchisotto. Connections in mathematics: An Introduction to Fibonacci via Pythagoras. Fibonacci Quarterly, 31(1):21–27, February 1993. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/31-1/marchisotto.pdf.

Maruszewski:2009:FFN

[Mar09] Richard Maruszewski. Fibonacci’s forgotten number revisited. College Mathematics Journal, 40(4):248–251, September 2009. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2009.11922371.

Marques:2012:FPO

[Mar12a] Diego Marques. Fixed points of the order of appearance in the Fibonacci sequence. Fibonacci Quarterly, 50(4):346–??, November 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/50-4/marques.pdf.

Marques:2012:OAI

[Mar12b] Diego Marques. The order of appearance of integers at most one away from Fibonacci numbers. Fibonacci Quarterly, 50(1):36– ??, February 2012. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/50-1/marques.pdf.

Marques:2012:OAPb

[Mar12c] Diego Marques. The order of appearance of powers of Fibonacci and Lucas numbers. Fibonacci Quarterly, 50(3):239–??, August 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/50-3/marques.pdf.

Marques:2012:OAPa

[Mar12d] Diego Marques. The order of appearance of product of consecutive Fibonacci numbers. Fibonacci Quarterly, 50(2):132–??, May 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/50-2/marques1.pdf.

Marques:2013:OAP

[Mar13a] Diego Marques. The order of appearance of the product of consecutive Lucas numbers. Fibonacci Quarterly, 51(1):38–??, February 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/51-1/marques1.pdf.

Marques:2013:SUB

[Mar13b] Diego Marques. Sharper upper bounds for the order of appearance in the Fibonacci sequence. Fibonacci Quarterly, 51(3):233– ??, August 2013. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/51-3/marques.pdf.

Marques:2015:NFL

[Mar15] Diego Marques. A new Fibonacci–Lucas relation. American Mathematical Monthly, 122(7):683, August 2015. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.jstor.org/stable/10.4169/amer.
math.monthly.122.7.683.

Mascagni:1994:FHQ

[Mas94a] Michael Mascagni. A fast, high quality, and reproducible parallel lagged-Fibonacci pseudorandom number generator. Technical report SRC-TR-94-115, Supercomputing Research Center: IDA, Lanham, MD, USA, March 15, 1994. 18 pp.

Mascagni:1994:PPN

[Mas94b] Michael Mascagni. Parallel pseudorandom number generation using additive lagged-Fibonacci recursions. Technical report SRC-TR-94-133, Supercomputing Research Center: IDA, Lanham, MD, USA, December 1, 1994. 15 pp.

Masulovic:2008:NFH

[Maš09] Dragan Mašulović. The number of finite homomorphismhomogeneous tournaments with loops. Fibonacci Quarterly, 46/47(3):241–244, August 2008/2009. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/46_47-
3/masulovic.pdf.

Mathis:1973:TFS

[Mat73] William L. Mathis. The Z transform and the Fibonacci sequence. Fibonacci Quarterly, 11(5):545–546, December 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-5/mathis.pdf.

Matsui:1990:OOI

[Mat90] T. Matsui. Object-oriented implementation of EusLisp. Denshi Gijutsu Sogo Kenkyusho Iho/Bulletin of the Electrotechnical Laboratory, 54(8):18–40, 1990. CODEN DGSKAR. ISSN 0366-9092.

Matstoms:1995:PSF

[Mat95] Pontus Matstoms. Parallel sparse QR factorization on shared memory architectures. Parallel Computing, 21(3):473–486, March 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print),

1872-7336 (electronic). URL http://www.elsevier.com/cgi-
bin/cas/tree/store/parco/cas_sub/browse/browse.cgi?year=
1995&volume=21&issue=3&aid=958.

May:1968:CFS

[May68] Donna B. May. On a characterization of the Fibonacci sequence. Fibonacci Quarterly, 6(5):11–14, November 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-5/may.pdf.

Mays:1982:NFP

[May82] Michael E. Mays. A note on Fibonacci primitive roots. Fibonacci Quarterly, 20(2):111–??, May 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-2/
mays.pdf.

Mays:1987:IDA

[May87] Michael E. Mays. Iterating the division algorithm. Fibonacci Quarterly, 25(3):204–213, August 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-
3/mays.pdf.

Mootha:1989:CES

[MB89] Vamsi Krishna Mootha and George Berzsenyi. Characterizations and extendibility of P_{t}-sets. Fibonacci Quarterly, 27(3):287–289, June 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/27-3/mootha.pdf.

Marinho:2019:HPF

[MB19] Andre A. Marinho and Francisco A. Brito. Hermite polynomials and Fibonacci oscillators. Journal of Mathematical Physics, 60 (1):012101, January 2019. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marques:2021:ESF

[MBBB21] R. Marques, C. Bouville, K. Bouatouch, and J. Blat. Extensible spherical Fibonacci grids. IEEE Transactions on Visualization and Computer Graphics, 27(4):2341–2354, 2021. CODEN ITVGEA. ISSN 1077-2626.

Marques:2013:SFP

[MBR^{+}13] R. Marques, C. Bouville, M. Ribardière, L. P. Santos, and K. Bouatouch. Spherical Fibonacci point sets for illumination integrals. Computer Graphics Forum, 32(8):134–143, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Majumder:1976:VNR

[MC76] Partha Pratim Majumder and Aravinda Chakravati. Variation in the number of ray- and disc-florets in four species of compositae. Fibonacci Quarterly, 14(2):97–100, April 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-2/majumder.pdf.

Melham:2000:ECM

[MC00] R. S. Melham and Curtis Cooper. The eigenvectors of a certain matrix of binomial coefficients. Fibonacci Quarterly, 38(2):123– 126, May 2000. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/38-2/melham.pdf.

Mollin:2004:PLC

[MC04] R. A. Mollin and K. Cheng. Period lengths of continued fractions involving Fibonacci numbers. Fibonacci Quarterly, 42(2):161– 169, May 2004. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Papers1/42-2/quartmollin02_2004.pdf.

McClenon:1919:LPH

[McC19] R. B. McClenon. Leonardo of Pisa and his Liber Quadratorum. American Mathematical Monthly, 26(1):1–8, 1919. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McCausland:1972:SOC

[McC72] I. McCausland. A simple optimal control sequence in terms of Fibonacci numbers. Fibonacci Quarterly, 10(6):561–564, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/10-6/mccausland-a.pdf.

McCarty:1981:FTN

[McC81] Carl P. McCarty. A formula for Tribonacci numbers. Fibonacci Quarterly, 19(5):391–392, December 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
5/mccarty.pdf.

McClung:1985:GUA

[McC85] O. W. McClung. Generators of unitary amicable numbers. Fibonacci Quarterly, 23(2):158–165, May 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-2/
mcclung.pdf.

McClung:1986:LE

[McC86] O. William McClung. Letter to the editor. Fibonacci Quarterly, 24(2):106–??, May 1986. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/24-2/letter.pdf.

McCarthy:1988:GMI

[McC88] P. J. McCarthy. A generalization of Metrod’s identity. Fibonacci Quarterly, 26(3):275–277, August 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
3/mccarthy.pdf.

McDaniel:1982:REI

[McD82] Wayne L. McDaniel. Representations of every integer as the difference of powerful numbers. Fibonacci Quarterly, 20(1):85–87, February 1982. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/20-1/mcdaniel.pdf.

McDaniel:1987:EIM

[McD87a] Wayne L. McDaniel. The existence of infinitely many k-Smith numbers. Fibonacci Quarterly, 25(1):76–80, February 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-1/mcdaniel.pdf.

McDaniel:1987:PSN

[McD87b] Wayne L. McDaniel. Powerful k-Smith numbers. Fibonacci Quarterly, 25(3):225–228, August 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-
3/mcdaniel.pdf.

McDaniel:1991:GCD

[McD91] Wayne L. McDaniel. The G.C.D. in Lucas sequences and Lehmer number sequences. Fibonacci Quarterly, 29(1):24–29, February 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/29-1/mcdaniel.pdf.

McDaniel:1994:ICS

[McD94a] Wayne L. McDaniel. The irrationality of certain series whose terms are reciprocals of Lucas sequence terms. Fibonacci Quarterly, 32(4):346–351, August 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-
4/mcdaniel2.pdf.

McDaniel:1994:GIF

[McD94b] Wayne L. McDaniel. On the greatest integer function and Lucas sequences. Fibonacci Quarterly, 32(4):297–300, August 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-4/mcdaniel1.pdf.

McDaniel:1995:DRL

[McD95] Wayne L. McDaniel. Diophantine representation of Lucas sequences. Fibonacci Quarterly, 33(1):59–63, February 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-1/mcdaniel.pdf.

McDaniel:1996:TNP

[McD96] Wayne L. McDaniel. Triangular numbers in the Pell sequence. Fibonacci Quarterly, 34(2):105–107, May 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-2/
mcdaniel.pdf.

McDaniel:1998:PFN

[McD98a] Wayne L. McDaniel. Pronic Fibonacci numbers. Fibonacci Quarterly, 36(1):56–59, February 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-
1/mcdaniel1.pdf.

McDaniel:1998:PLN

[McD98b] Wayne L. McDaniel. Pronic Lucas numbers. Fibonacci Quarterly, 36(1):60–62, February 1998. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/36-1/mcdaniel2.pdf.

McDaniel:2001:FLN

[McD01] Wayne L. McDaniel. On the factorization of Lucas numbers. Fibonacci Quarterly, 39(3):206–210, June 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-3/
mcdaniel.pdf.

McDaniel:2002:FPN

[McD02] Wayne L. McDaniel. On Fibonacci and Pell numbers of the form kx 2. Fibonacci Quarterly, 40(1):41–42, February 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-1/mcdaniel.pdf.

McDaniel:2011:EDF

[McD11] Wayne L. McDaniel. An easy determination of the Fibonacci and Pell squares. Fibonacci Quarterly, 49(2):166–170, May 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/49-2/mcdaniel.pdf.

McGuire:2008:PCR

[McG09] Trevor McGuire. On the periodicity of certain recursive sequences. Fibonacci Quarterly, 46/47(4):350–355, November 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/46_47-4/mcguire.pdf.

McGuire:2010:GPC

[McG10] Trevor McGuire. Generalizations of the periodicity of certain recursive sequences. Fibonacci Quarterly, 48(2):175–181, May 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-2/mcguire.pdf.

McHugh:1982:CS

[McH82] Joseph McHugh. Characterization of a sequence. Fibonacci Quarterly, 20(3):252–255, August 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
3/mchugh.pdf.

McIlroy:1982:NSD

[McI82] M. D. McIlroy. The number of states of a dynamic storage allocation system. The Computer Journal, 25(3):388–392, August 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/25/3/388.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_25/Issue_03/tiff/388.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_25/
Issue_03/tiff/389.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_25/Issue_03/tiff/390.tif;   http:/
/www3.oup.co.uk/computer_journal/hdb/Volume_25/Issue_
03/tiff/391.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_25/Issue_03/tiff/392.tif.

McIntosh:1992:GCP

[McI92] Richard J. McIntosh. A generalization of a congruential property of Lucas. American Mathematical Monthly, 99(3):231–238, March 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

McIntosh:2020:CIE

[McI20] Richard J. McIntosh. Congruences involving Euler numbers and power sums. Fibonacci Quarterly, 58(4):328–??, November 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/58-4/mcintosh.pdf.

McLaughlin:1979:NTA

[McL79] William I. McLaughlin. Note on a Tetranacci alternative to Bode’s law. Fibonacci Quarterly, 17(2):116–117, April 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-2/mclaughlin.pdf.

McNeill:1987:NDS

[McN87] R. B. McNeill. A note on divisibility sequences. Fibonacci Quarterly, 25(3):214–215, August 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-
3/mcneill.pdf.

McNeill:1988:CDS

[McN88] R. B. McNeill. On certain divisibility sequences. Fibonacci Quarterly, 26(2):169–171, May 1988. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/26-2/mcneill.
pdf.

McNeill:1992:TMC

[McN92] R. B. McNeill. On a theorem of Monzingo characterizing the prime divisors of certain sequences of integers. Fibonacci Quarterly, 30(2):110–??, May 1992. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/30-2/mcneill.
pdf.

Mascagni:1995:FHQ

[MCPR95] Michael Mascagni, Steven A. Cuccaro, Daniel V. Pryor, and M. L. Robinson. A fast, high quality, and reproducible parallel lagged-Fibonacci pseudorandom number generator. Journal of Computational Physics, 119(2):211–219, July 1995. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999185711308.

McQuistan:1976:MNN

[McQ76] R. B. McQuistan. Mixed nearest neighbor degeneracy for particles on a one-dimensional lattice space. Fibonacci Quarterly, 14(4):353–357, November 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
4/mcquistan.pdf.

Munarini:2001:LC

[MCS01] E. Munarini, C. P. Cippo, and N. Z. Salvi. On the Lucas cubes. Fibonacci Quarterly, 39(1):12–21, February 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-1/munarini.pdf.

Milovanovic:1987:SPH

[MD87] G. V. Milovanović and G. Djordjević. On some properties of Humbert’s polynomials. Fibonacci Quarterly, 25(4):356–360, November 1987. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/25-4/milovanovic.pdf.

Means:1956:CNP

[Mea56] J. H. Means. Classroom notes: Predetermined determinant values using Fibonacci numbers. American Mathematical Monthly, 63(9):657–658, November 1956. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mead:1965:EMS

[Mea65] D. G. Mead. An elementary method of summation. Fibonacci Quarterly, 3(3):209–213, October 1965. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/3-3/mead.
pdf.

Melham:1997:CWC

[Mel97] Ray Melham. Conics which characterize certain Lucas sequences. Fibonacci Quarterly, 35(3):248–251, August 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-3/melham.pdf.

Melham:1998:GTP

[Mel98] Ray Melham. Generalized triple products. Fibonacci Quarterly, 36(5):452–456, November 1998. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/36-5/melham.
pdf.

Melham:1999:FII

[Mel99a] R. S. Melham. Families of identities involving sums of powers of the Fibonacci and Lucas numbers. Fibonacci Quarterly, 37(4):

315–319, November 1999. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/37-4/melham2.pdf.

Melham:1999:LSE

[Mel99b] R. S. Melham. Lambert series and elliptic functions and certain reciprocal sums. Fibonacci Quarterly, 37(3):208–212, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/37-3/melham1.pdf.

Melham:1999:LSFa

[Mel99c] R. S. Melham. Lucas sequences and functions of a 3-by-3 matrix. Fibonacci Quarterly, 37(2):111–115, May 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-2/melham2.pdf.

Melham:1999:LSFb

[Mel99d] R. S. Melham. Lucas sequences and functions of a 4-by-4 matrix. Fibonacci Quarterly, 37(3):269–276, August 1999. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-3/melham3.pdf.

Melham:1999:SAI

[Mel99e] R. S. Melham. Some analogs of the identity F_{n}^{2} + F_{n+1}^{2} = F_{2n+1}. Fibonacci Quarterly, 37(4):305–311, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-4/melham1.pdf.

Melham:1999:SCP

[Mel99f] R. S. Melham. Sums of certain products of Fibonacci and Lucas numbers. Fibonacci Quarterly, 37(3):248–251, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-3/melham2.pdf.

Melham:1999:GSI

[Mel99g] Ray Melham. Generalizations of some identities of long. Fibonacci Quarterly, 37(2):106–110, May 1999. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/37-2/
melham1.pdf.

Melham:2000:ASF

[Mel00a] R. S. Melham. Alternating sums of fourth powers of Fibonacci and Lucas numbers. Fibonacci Quarterly, 38(3):254–259, June 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/38-3/melham.pdf.

Melham:2000:ODC

[Mel00b] R. S. Melham. On an observation of D’Ocagne concerning the fundamental sequence. Fibonacci Quarterly, 38(5):446–450, November 2000. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/38-5/melham.pdf.

Melham:2000:SRW

[Mel00c] R. S. Melham. Summation of reciprocals which involve products of terms from generalized Fibonacci sequences. Fibonacci Quarterly, 38(4):294–298, August 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-
4/melham.pdf.

Melham:2000:SCP

[Mel00d] R. S. Melham. Sums of certain products of Fibonacci and Lucas numbers — Part II. Fibonacci Quarterly, 38(1):3–7, February 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/38-1/melham.pdf.

Melham:2001:SRW

[Mel01] R. S. Melham. Summation of reciprocals which involve products of terms from generalized Fibonacci sequences — Part II. Fibonacci Quarterly, 39(3):264–267, June 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-3/
melham.pdf.

Melham:2002:RFS

[Mel02] R. S. Melham. Reduction formulas for the summation of reciprocals in certain second-order recurring sequences. Fibonacci Quarterly, 40(1):71–75, February 2002. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/40-1/melham.
pdf.

Melham:2003:FIS

[Mel03a] R. S. Melham. A Fibonacci identity in the spirit of Simson and Gelin–Cesàro. Fibonacci Quarterly, 41(2):142–143, May 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-2/melham.pdf.

Melham:2003:SRS

[Mel03b] R. S. Melham. On some reciprocal sums of Brousseau: An alternative approach to that of Carlitz. Fibonacci Quarterly, 41(1): 59–62, February 2003. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/41-1/melham.pdf.

Melham:2003:TVI

[Mel03c] R. S. Melham. A three-variable identity involving cubes of Fibonacci numbers. Fibonacci Quarterly, 41(3):220–223, June 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-3/melham.pdf.

Melham:2004:CCF

[Mel04a] R. S. Melham. Certain classes on finite sums that involve generalized Fibonacci and Lucas numbers. Fibonacci Quarterly, 42 (1):47–54, February 2004. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers1/42-1/quartmelham01_
2004.pdf.

Melham:2004:YOF

[Mel04b] R. S. Melham. Ye olde Fibonacci curiosity shoppe revisited. Fibonacci Quarterly, 42(2):155–160, May 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-2/
quartmelham02_2004.pdf.

Melham:2008:SCC

[Mel09] R. S. Melham. Some conjectures concerning sums of odd powers of Fibonacci and Lucas numbers. Fibonacci Quarterly, 46/47 (4):312–315, November 2008/2009. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/46_47-
4/melham.pdf.

Melham:2010:MCH

[Mel10a] R. S. Melham. More on combinations of higher powers of Fibonacci numbers. Fibonacci Quarterly, 48(4):307–311, November 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-4/melham.pdf.

Melham:2010:CCH

[Mel10b] R. S. Melham. On certain combinations of higher powers of Fibonacci numbers. Fibonacci Quarterly, 48(3):256–259, August 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/48-3/melham.pdf.

Melham:2010:RCR

[Mel10c] R. S. Melham. On the representation of certain reals via the golden ratio. Fibonacci Quarterly, 48(2):150–160, May 2010. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/48-2/melham.pdf.

Melham:2015:CFF

[Mel15] R. S. Melham. On certain families of finite reciprocal sums that involve generalized Fibonacci numbers. Fibonacci Quarterly, 53 (4):323–??, November 2015. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/53-4/melham.pdf.

Melham:2016:CFF

[Mel16a] R. S. Melham. Closed forms for finite sums in which the denominator of the summand is a product of trigonometric functions. Fibonacci Quarterly, 54(3):196–??, August 2016. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/54-3/melham1.pdf.

Melham:2016:MNA

[Mel16b] R. S. Melham. More new algebraic identities and the Fibonacci summations derived from them. Fibonacci Quarterly, 54(1):31– ??, February 2016. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/54-1/melham2.pdf.

Melham:2016:NIS

[Mel16c] R. S. Melham. New identities satisfied by powers of Fibonacci and Lucas numbers. Fibonacci Quarterly, 54(4):296–??, November 2016. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/54-4/melham.pdf.

Melham:2016:CFI

[Mel16d] R. S. Melham. On a classical Fibonacci identity of Aurifeuille. Fibonacci Quarterly, 54(1):19–??, February 2016. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/54-1/melham1.pdf.

Melham:2016:GPE

[Mel16e] R. S. Melham. On a generalized Pell equation studied by Euler and Sadek. Fibonacci Quarterly, 54(1):49–??, February 2016.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/54-1/melham3.pdf.

Melham:2016:PIP

[Mel16f] R. S. Melham. On the positive integer points of certain two parameter families of hyperbolas. Fibonacci Quarterly, 54(3): 247–??, August 2016. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/54-3/melham2.pdf.

Melham:2016:TAI

[Mel16g] R. S. Melham. Two algebraic identities and the alternating Fibonacci sums produced by them. Fibonacci Quarterly, 54(2): 154–??, May 2016. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/54-2/melham2.pdf.

Melham:2016:TPP

[Mel16h] R. S. Melham. A two parameter Pell Diophantine equation that generalizes a Fibonacci classic. Fibonacci Quarterly, 54(2):112– ??, May 2016. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/54-2/melham1.pdf.

Melham:2017:CFCb

[Mel17a] R. S. Melham. Closed forms for certain Fibonacci type sums that involve second order products. Fibonacci Quarterly, 55(3): 195–??, August 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/55-3/melham1.pdf.

Melham:2017:CFFa

[Mel17b] R. S. Melham. Closed forms for finite sums of weighted products of generalized Fibonacci numbers. Fibonacci Quarterly, 55(2): 99–??, May 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/55-2/melham1.pdf.

Melham:2017:CFFb

[Mel17c] R. S. Melham. Closed forms for finite sums of weighted products of the sine and cosine functions. Fibonacci Quarterly, 55(2):123– ??, May 2017. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/55-2/melham2.pdf.

Melham:2017:SCP

[Mel17d] R. S. Melham. Sums of certain products of Fibonacci and Lucas numbers — Part III. Fibonacci Quarterly, 55(3):229–??, August 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/55-3/melham2.pdf.

Melham:2018:TPF

[Mel18a] R. S. Melham. 12 two-parameter families of reciprocal sums of products of the sine and cosine functions. Fibonacci Quarterly, 56 (4):329–??, November 2018. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/56-4/melham2.pdf.

Melham:2018:CFFb

[Mel18b] R. S. Melham. Closed forms for 10 families of finite sums of fractional generalized Fibonacci products. Fibonacci Quarterly, 56 (4):290–??, November 2018. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/56-4/melham1.pdf.

Melham:2018:CFFc

[Mel18c] R. S. Melham. Closed formulas for finite sums of fractional expressions that involve the sine and cosine functions. Fibonacci Quarterly, 56(4):360–??, November 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/56-
4/melham3.pdf.

Melham:2018:CFFa

[Mel18d] R. S. Melham. Closed formulas for finite sums of weighted fractional generalized Fibonacci products. Fibonacci Quarterly, 56 (2):167–??, May 2018. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/56-2/melham2.pdf.

Melham:2018:FCFa

[Mel18e] R. S. Melham. Further closed forms for finite sums of weighted products of generalized Fibonacci numbers. Fibonacci Quarterly, 56(1):3–??, February 2018. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/56-1/melham1.pdf.

Melham:2018:FCFb

[Mel18f] R. S. Melham. Further closed forms for finite sums of weighted products of the sine and cosine functions. Fibonacci Quarterly, 56(1):38–??, February 2018. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/56-1/melham2.pdf.

Melham:2018:SRW

[Mel18g] R. S. Melham. Sums of reciprocals of weighted products of the sine and and cosine functions. Fibonacci Quarterly, 56(2):99–??, May 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/56-2/melham1.pdf.

Mennicke:1989:FGS

[Men89] Jens L. Mennicke. On Fibonacci groups and some other groups. Lecture Notes in Mathematics, 1398:117–123, 1989. CODEN LNMAA2. ISBN 3-540-51695-6 (print), 3-540-46756-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0086248/.

Mestrovic:2013:LTT

[Meš13] Romeo Meštrović. A Lucas type theorem modulo prime powers. Fibonacci Quarterly, 51(2):142–??, May 2013. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/51-
2/mestrovic.pdf.

Metz:1997:GSG

[Met97] James Metz. The golden staircase and the golden line. Fibonacci Quarterly, 35(3):194–197, August 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-
3/metz.pdf.

Metsankyla:2014:NKC

[Met14] Tauno Metsänkylä. Note on Kummer’s congruences for Euler numbers. Fibonacci Quarterly, 52(2):160–??, May 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-2/metsankyla.pdf.

Meyer:2005:SAD

[Mey05] Jeffrey L. Meyer. Symmetric arguments in the Dedekind sum. Fibonacci Quarterly, 43(2):122–123, May 2005. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/43-2/
paper43-2-4.pdf.

MacKinnon:2002:FTP

[MG02] Nick MacKinnon and Stephen M. Gagola, Jr. Fibonacci twin primes: 10844. American Mathematical Monthly, 109(1):78, January 2002. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/
2695779.

McDaniel:1975:SRC

[MH75a] W. L. McDaniel and Peter Hagis, Jr. Some results concerning the non-existence of odd perfect numbers of the form p^{α}M^{2β}. Fibonacci Quarterly, 13(1):25–28, February 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-1/mcdaniel.pdf.

Mowshowitz:1975:EEL

[MH75b] Abbe Mowshowitz and Frank Harary. Enumeration of endlabeled trees. Fibonacci Quarterly, 13(3):252–254, October 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-3/mowshowitz.pdf.

Mahon:1985:ITS

[MH85] Bro. J. M. Mahon and A. F. Horadam. Inverse trigonometrical summation formulas involving Pell polynomials. Fibonacci Quarterly, 23(4):319–324, November 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-
4/mahon.pdf.

Mahon:1986:MOS

[MH86] Bro. J. M. Mahon and A. F. Horadam. Matrix and other summation techniques for Pell polynomials. Fibonacci Quarterly, 24(4): 290–308, November 1986. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/24-4/mahon.pdf.

Mahon:1987:CSG

[MH87a] Br. J. Mahon and A. F. Horadam. A constellation of sequences of generalized Pell polynomials. Fibonacci Quarterly, 25(2):106– 110, May 1987. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/25-2/mahon.pdf.

Mahon:1987:EGF

[MH87b] Bro. J. M. Mahon and A. F. Horadam. Exponential generating functions for Pell polynomials. Fibonacci Quarterly, 25(3):194– 203, August 1987. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/25-3/mahon.pdf.

Mahon:1987:OGF

[MH87c] Bro. J. M. Mahon and A. F. Horadam. Ordinary generating functions for Pell polynomials. Fibonacci Quarterly, 25(1):45–56, February 1987. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/25-1/mahon2.pdf.

Mahon:1987:PPM

[MH87d] Bro. J. M. Mahon and A. F. Horadam. Pell polynomial matrices. Fibonacci Quarterly, 25(1):21–28, February 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-1/mahon1.pdf.

Mahon:1990:TOD

[MH90] Br. J. M. Mahon and A. F. Horadam. Third-order diagonal functions of Pell polynomials. Fibonacci Quarterly, 28(1):3–10, February 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/28-1/mahon.pdf.

Mirzaee:2016:AFC

[MH16] Farshid Mirzaee and Seyede Fatemeh Hoseini. Application of Fibonacci collocation method for solving Volterra–Fredholm integral equations. Applied Mathematics and Computation, 273 (??):637–644, January 15, 2016. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300315013776.

Michael:1964:NPO

[Mic64] Glen Michael. A new proof for an old property. Fibonacci Quarterly, 2(1):57–58, February 1964. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/2-1/michael.
pdf.

Midttun:1979:CCF

[Mid79] Norvald Midttun. Congruences for certain Fibonacci numbers. Fibonacci Quarterly, 17(1):40–41, February 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-1/midttun.pdf.

Mignotte:1977:AWS

[Mig77] Maurice Mignotte. An application of W. Schmidt’s theorem transcendental numbers and golden number. Fibonacci Quarterly, 15 (1):15–16, February 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-1/mignotte.pdf.

Miles:1960:GFN

[Mil60] E. P. Miles, Jr. Generalized Fibonacci numbers and associated matrices. American Mathematical Monthly, 67(8):745–752, October 1960. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Miller:1971:MNG

[Mil71] M. D. Miller. Mathematical notes: On generalized Fibonacci numbers. American Mathematical Monthly, 78(10):1108–1109, December 1971. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Miller:1975:FSN

[Mil75a] J. C. P. Miller. On factorisation, with a suggested new approach. Mathematics of Computation, 29(129):155–172, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miller:1975:RDD

[Mil75b] Michael D. Miller. A recursively defined divisor function. Fibonacci Quarterly, 13(3):199–203, October 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-3/miller.pdf.

Miller:1991:SFL

[Mil91] Allen R. Miller. Solutions of Fermat’s Last Equation in terms of Wright’s hypergeometric function. Fibonacci Quarterly, 29(1): 52–56, February 1991. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/29-1/miller.pdf.

Milatovic:2011:POT

[Mil11] Ognjen Milatovic. 95.21 A property of odd terms of a generalised Fibonacci sequence. The Mathematical Gazette, 95(533):262–264, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Miller:2020:PS

[Mil20] Steven J. Miller. Problem section. Fibonacci Quarterly, 58(5): 236–??, December 2020. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers1/58-5/problems.pdf.

Editor:2022:PS

[Mil22] Steven J. Miller. Problem section. Fibonacci Quarterly, 60(5): 401–??, December 2022. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers1/60-5/problems.pdf.

Minoli:1981:SIH

[Min81a] Daniel Minoli. Structural issues for hyperperfect numbers. Fibonacci Quarterly, 19(1):6–13, February 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-1/
minoli.pdf.

Mintz:1981:SBM

[Min81b] Donald J. Mintz. 2, 3 sequence as binary mixture. Fibonacci Quarterly, 19(4):351–360, October 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
4/mintz.pdf.

Ming:1989:TFN

[Min89] Luo Ming. On triangular Fibonacci numbers. Fibonacci Quarterly, 27(2):98–108, May 1989. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/27-2/ming.pdf.

Mitzenmacher:2003:BHG

[Mit03] Michael Mitzenmacher. A brief history of generative models for power law and lognormal distributions. Internet Mathematics, 1(2):226–251, December 2003. CODEN ???? ISSN 1542-7951 (print), 1944-9488 (electronic). URL http://

projecteuclid.org/euclid.im/1089229510;   http://www.
eecs.harvard.edu/~michaelm/CS223/powerlaw.pdf;   http:
//www.tandfonline.com/toc/uinm20/1/2.

Miura:1981:ALA

[Miu81] Nobuo Miura. The algebra in the Liber Abaci of Leonardo Pisano. Historia Scientiarum = International journal of the History of Science Society of Japan, 21(??):57–65, 1981. CODEN HIS-

CDU. ISSN 0285-4821. URL https://ci.nii.ac.jp/naid/
110009837266.

Miura:2002:LPC

[Miu02]  Nobuo Miura.  Leonardo of Pisa and cubic equations.

Sūrikaisekikenkyūsho Kōkyūroku, (1257):37–47, 2002. Study of the history of mathematics (Japanese) (Kyoto, 2001).

Melham:1995:GLR

[MJ95] Ray Melham and Derek Jennings. On the general linear recurrence relation. Fibonacci Quarterly, 33(2):142–146, May 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-2/melham3.pdf.

Jiang:2010:AFS

[mJ10] Dong mei Jiang. Application of Fibonacci sine and cosine function to a nonlinear differential–difference equation. Applied Mathematics and Computation, 217(8):3898–3902, December 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Metzger:1987:NPE

[MK87] J. M. Metzger and S. P. Kaler. A note on the Pell equation. Fibonacci Quarterly, 25(3):216–220, August 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-3/metzger.pdf.

Miller:2002:FLE

[MK02] Carl Miller and Donald E. Knuth. A Fibonacci–Lucas extremum: 10825. American Mathematical Monthly, 109(8):762–763, October 2002. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://www.jstor.org/stable/3072414.

Moura:2023:RRL

[MKL^{+}23] Guilherme Zeus Dantas E Moura, Andrew Keisling, Astrid Lilly, Annika Mauro, Steven J. Miller, Matthew Phang, and Santiago Velazquez Iannuzzelli. Recurrence relations for S-legal index difference sequences. Fibonacci Quarterly, 61(3):257–??, August 2023. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/61-3/moura.pdf.

McLaughlin:1984:DFI

[ML84] William I. McLaughlin and Sylvia A. Lundy. Digit functions of integer sequences. Fibonacci Quarterly, 22(2):105–115, May 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/22-2/mclaughlin.pdf.

MacLaren:1965:URN

[MM65] M. Donald MacLaren and George Marsaglia. Uniform random number generators. Journal of the Association for Computing Machinery, 12(1):83–89, January 1965. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Merkes:1973:LEA

[MM73] E. P. Merkes and David Meyers. On the length of the Euclidean algorithm. Fibonacci Quarterly, 11(1):56–61, February 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-1/merkes.pdf.

Mitchell:1975:NTF

[MM75] A. R. Mitchell and R. W. Mitchell. A note on topologies on finite sets. Fibonacci Quarterly, 13(4):356–??, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-4/mitchell-a.pdf.

Mohanty:1990:PN

[MM90] Supriya Mohanty and S. P. Mohanty. Pythagorean numbers. Fibonacci Quarterly, 28(1):31–42, February 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/28-1/mohanty.pdf.

Makover:2006:EPR

[MM06] Eran Makover and Jeffrey McGowan. An elementary proof that random Fibonacci sequences grow exponentially. Journal of Number Theory, 121(1):40–44, November 2006. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X06000229.

Martinez:2022:GZTb

[MMM^{+}22] Thomas C. Martinez, Steven J. Miller, Clayton Mizgerd, Jack Murphy, and Chenyang Sun. Generalizing Zeckendorf’s theorem to homogeneous linear recurrences, II. Fibonacci Quarterly, 60(5):231–??, December 2022. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/60-5/
martinez2.pdf.

Martinez:2022:GZTa

[MMMS22] Thomas C. Martinez, Steven J. Miller, Clayton Mizgerd, and Chenyang Sun. Generalizing Zeckendorf’s theorem to homogeneous linear recurrences, I. Fibonacci Quarterly, 60(5):222–??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/60-5/martinez1.pdf.

Matsui:2010:PLT

[MMYA10] H. Matsui, D. Minematsu, T. Yamauchi, and R. Miy Adera. Pascal-like triangles and Fibonacci-like sequences. The Mathematical Gazette, 94(529):27–41, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Monteiro:1976:MMF

[MN76] P. Monteiro and R. W. Newcomb. Minimal and maximal Fibonacci representations: Boolean generation. Fibonacci Quarterly, 14(1):9–11, February 1976. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/14-1/monteiro.
pdf.

Miller:2020:FQG

[MN20] Steven J. Miller and Alexandra Newlon. The Fibonacci quilt game. Fibonacci Quarterly, 58(2):157–??, May 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/58-2/miller.pdf.

Miller:2017:ABV

[MNPX17] Steven J. Miller, Dawn Nelson, Zhao Pan, and Huanzhong Xu. On the asymptotic behavior of variance of PLRS decompositions. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/55-5/miller.pdf.

Marsaglia:1990:RNG

[MNZ90] George Marsaglia, B. Narasimhan, and Arif Zaman. A random number generator for PC’s. Computer Physics Communications, 60(3):345–349, October 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
001046559090033W.

McKenzie:2008:PPS

[MO09] Thomas McKenzie and Shannon Overbay. Purely periodic second order linear recurrences. Fibonacci Quarterly, 46/47(2):160–166, May 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/46_47-2/mckenzie.pdf.

McKenzie:2010:SSO

[MO10] Thomas McKenzie and Shannon Overbay. Sums of second order linear recurrences. Fibonacci Quarterly, 48(4):335–342, November 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/48-4/mckenzie.pdf.

Mohanty:1967:RC

[Moh67] S. G. Mohanty. Restricted compositions. Fibonacci Quarterly, 5 (3):223–234, October 1967. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/5-3/mohanty.pdf.

Mohanty:1968:PGF

[Moh68] S. G. Mohanty. On a partition of generalized Fibonacci numbers. Fibonacci Quarterly, 6(1):22–33, February 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-1/mohanty.pdf.

Mohanty:1976:IPL

[Moh76] N. C. Mohanty. Interesting properties of Laguerre polynomials. Fibonacci Quarterly, 14(1):42–??, February 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-1/mohanty.pdf.

Mohanty:1978:NPI

[Moh78] S. P. Mohanty. The number of primes is infinite. Fibonacci Quarterly, 16(4):381–384, August 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
4/mohanty.pdf.

Mohanty:1981:PJB

[Moh81] S. P. Mohanty. On a problem of S. J. Bezuszka and M. J. Kenney on cyclic difference of pairs of integers. Fibonacci Quarterly, 19 (4):314–317, October 1981. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/19-4/mohanty.pdf.

Mohanty:1991:MS

[Moh91] Supriya Mohanty. On multi-sets. Fibonacci Quarterly, 29(2): 108–113, May 1991. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/29-2/mohanty.pdf.

Mollin:1988:GFP

[Mol88] R. A. Mollin. Generalized Fibonacci primitive roots, and class numbers of real quadratic fields. Fibonacci Quarterly, 26(1):46– 53, February 1988. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/26-1/mollin.pdf.

Mollard:2018:EPC

[Mol18] Michel Mollard. The existence of perfect codes in a family of generalized Fibonacci cubes. Information Processing Letters, 140(??):ii, December 2018. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0020019018301571.

Momiyama:2001:NRF

[Mom01] Harunobu Momiyama. A new recurrence formula for Bernoulli numbers. Fibonacci Quarterly, 39(3):285–289, June 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-3/momiyama.pdf.

Monzingo:1974:EFN

[Mon74] M. G. Monzingo. On extending the Fibonacci numbers to the negative integers. Fibonacci Quarterly, 12(3):292–??, October 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/12-3/monzingo-a.pdf.

Monzingo:1976:NGE

[Mon76a] M. C. Monzingo. A note on the golden ellipse. Fibonacci Quarterly, 14(5):388–??, December 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
5/monzingo1.pdf.

Monzingo:1976:CMP

[Mon76b] M. G. Monzingo. On congruences modulo a power of a prime. Fibonacci Quarterly, 14(1):23–??, February 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-1/monzingo.pdf.

Monzingo:1976:CPR

[Mon76c] M. G. Monzingo. On consecutive primitive roots. Fibonacci Quarterly, 14(5):391–??, December 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
5/monzingo2-a.pdf.

Monzingo:1983:WSP

[Mon83] M. G. Monzingo. Why are 8:18 and 10:09 such pleasant times? Fibonacci Quarterly, 21(2):107–110, May 1983. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-2/monzingo.pdf.

Monzingo:1985:DCT

[Mon85] M. G. Monzingo. On the distribution of consecutive triples of quadratic residues and quadratic nonresidues and related topics. Fibonacci Quarterly, 23(2):133–138, May 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/23-2/monzingo.pdf.

Monzingo:1988:PDS

[Mon88] M. G. Monzingo. On prime divisors of sequences of integers involving squares. Fibonacci Quarterly, 26(1):31–32, February 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/26-1/monzingo.pdf.

Monzingo:1990:CDS

[Mon90a] M. G. Monzingo. On certain divisibility sequences. Fibonacci Quarterly, 28(2):181–??, May 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-2/
monzingo.pdf.

Monzingo:1990:S

[Mon90b] M. G. Monzingo. On the sum ∑_{a=1}^{p−1}(a/p)a. Fibonacci Quarterly, 28(1):56–59, February 1990. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/28-1/monzingo.pdf.

Moore:1970:MUP

[Moo70] Richard E. M. Moore. Mosaic units: Pattern sizes in ancient mosaics. Fibonacci Quarterly, 8(3):281–310, April 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-3/moore-a.pdf.

Moore:1973:MHH

[Moo73] Carl F. Moore. More hidden hexagon squares. Fibonacci Quarterly, 11(5):525–??, December 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
5/moore.pdf.

Moore:1989:EAL

[Moo89] Thomas E. Moore. Euclid’s algorithm and Lamé’s theorem on a microcomputer. Fibonacci Quarterly, 27(4):290–295, August 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/27-4/moore.pdf.

Moore:1992:LAR

[Moo92] Thomas E. Moore. On the least absolute remainder Euclidean algorithm. Fibonacci Quarterly, 30(2):161–165, May 1992. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/30-2/moore.pdf.

Moore:1993:FPS

[Moo93a] Gregory A. Moore. A Fibonacci polynomial sequence defined by multidimensional continued fractions; and higher-order golden ratios. Fibonacci Quarterly, 31(4):354–364, November 1993. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-4/moore.pdf.

Mootha:1993:UFN

[Moo93b] Vamsi K. Mootha. Unary Fibonacci numbers are contextsensitive. Fibonacci Quarterly, 31(1):41–43, February 1993. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-1/mootha.pdf.

Moore:1994:LGN

[Moo94] Gregory A. Moore. The limit of the golden numbers is 3/2. Fibonacci Quarterly, 32(3):211–217, June 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-3/
moore.pdf.

Mangeron:1971:GFN

[MOP71] D. Mangeron, M. N. Oguztorelli, and V. E. Poterasu. On the generation of Fibonacci numbers and the “polyvibrating” extension of these numbers. Fibonacci Quarterly, 9(3):324–328, May 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/9-3/mangeron-a.pdf.

Morrison:1975:NPT

[Mor75] Michael A. Morrison. A note on primality testing using Lucas sequences. Mathematics of Computation, 29(129):181–182, January 1975. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Morton:1977:FSP

[Mor77] Avery A. Morton. The Fibonacci series and the periodic table of elements. Fibonacci Quarterly, 15(2):173–175, April 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-2/morton.pdf.

Morgan:1983:FSM

[Mor83] Karolyn A. Morgan. The Fibonacci sequence F_{n} modulo L_{m}. Fibonacci Quarterly, 21(4):304–305, November 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/21-4/morgan.pdf.

Morrison:1989:FLM

[Mor89] John F. Morrison. Fibonacci-like matrices. Fibonacci Quarterly, 27(1):47–48, February 1989. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/27-1/morrison.pdf.

Morton:1995:FLS

[Mor95] H. R. Morton. Fibonacci-like sequences and greatest common divisors. American Mathematical Monthly, 102(8):731–734, October 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Mortici:2010:RCS

[Mor10] Cristinel Mortici. Remarks on complementary sequences. Fibonacci Quarterly, 48(4):343–347, November 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/48-4/mortici.pdf.

Moser:1993:CBS

[Mos93] W. Moser. Cyclic binary strings without long runs of like (alternating) bits. Fibonacci Quarterly, 31(1):2–6, February 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-1/moser.pdf.

Mosley:2008:BRS

[Mos08] Adam Mosley. Book reviews: Stephen Clucas (ed.) John Dee: Interdisciplinary Studies in English Renaissance Thought. International Archives of the History of Ideas, 193. Dordrecht: Springer, 2006. Pp. xvii + 366. ISBN: 1-4020-4245-0. EUR144.00, $189.00, £111.00 (hardback). ISBN: 1-4020-4246-9 (e-book). British Journal for the History of Science, 41(1):132–134, March 2008. CO-DEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

McKnight:1966:SFM

[MP66] Curtis McKnight and Dean Priest. A singular Fibonacci matrix and its related lambda function. Fibonacci Quarterly, 4(3):259– 261, October 1966. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/4-3/mcknight.pdf.

McCabe:1985:ASG

[MP85] J. H. McCabe and G. M. Phillips. Aitken sequences and generalized Fibonacci numbers. Mathematics of Computation, 45(172): 553–558, October 1985. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Miller:2022:WIFa

[MPPW22] Steven J. Miller, Fei Peng, Tudor Popescu, and Nawapan Wattanawanichkul. Walking to infinity on the Fibonacci sequence. Fibonacci Quarterly, 60(5):293–??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/60-5/miller2.pdf.

Martin:1984:PFA

[MR84] Alain J. Martin and Martin Rem. A presentation of the Fibonacci algorithm. Information Processing Letters, 19(2):67–68, August 31, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mohanty:1985:S

[MR85] S. P. Mohanty and A. M. S. Ramasamy. On P_{r,k} sequences. Fibonacci Quarterly, 23(1):36–44, February 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/23-1/mohanty.pdf.

Moree:1995:AFR

[MR95] Pieter Moree and Hans Roskam. On an arithmetical function related to Euler’s totient and the discriminator. Fibonacci Quarterly, 33(4):332–340, August 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-
4/moree.pdf.

Mouline:1999:AMC

[MR99] Mehdi Mouline and Mustapha Rachidi. Application of Markov chains properties to r-generalized Fibonacci sequences. Fibonacci Quarterly, 37(1):34–38, February 1999. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/37-
1/mouline.pdf.

Mouline:2000:GFS

[MR00] Mehdi Mouline and Mustapha Rachidi. ∞-generalized Fibonacci sequences and Markov chains. Fibonacci Quarterly, 38(4):364– 371, August 2000. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/38-4/mouline.pdf.

McIntosh:2007:SFW

[MR07] Richard J. McIntosh and Eric L. Roettger. A search for Fibonacci–Wieferich and Wolstenholme primes. Mathematics of Computation, 76(260):2087–2094, October 2007. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

URL http://www.ams.org/mcom/2007-76-260/S0025-5718-
07-01955-2/home.html;   http://www.ams.org/mcom/2007-
76-260/S0025-5718-07-01955-2/S0025-5718-07-01955-2.dvi;
 http://www.ams.org/mcom/2007-76-260/S0025-5718-07-01955-
2/S0025-5718-07-01955-2.pdf;   http://www.ams.org/mcom/
2007-76-260/S0025-5718-07-01955-2/S0025-5718-07-01955-
2.ps.

Medina:2015:PRP

[MR15] Luis A. Medina and Eric Rowland. p-regularity of the p-adic valuation of the Fibonacci sequence. Fibonacci Quarterly, 53(3): 265–??, August 2015. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/53-3/medina.pdf.

Mascagni:1995:PPNa

[MRPC95a] Michael Mascagni, M. L. Robinson, Daniel V. Pryor, and Steven A. Cuccaro. Parallel pseudorandom number generation using additive lagged-Fibonacci recursions. Report ????, Supercomputing Research Center, IDA, ????, ???? 1995. 15 pp. URL

http://www.cs.fsu.edu/~mascagni/papers/RIJP1995_1.pdf.

Mascagni:1995:PPNb

[MRPC95b] Michael Mascagni, M. L. Robinson, Daniel V. Pryor, and Steven A. Cuccaro. Parallel pseudorandom number generation using additive lagged-Fibonacci recursions. In Niederreiter and Shiue [NS95], pages 263–277. ISBN 0-387-94577-6 (softcover). LCCN Q183.9 .M66 1995.

Motta:1999:GFS

[MRS99] W. Motta, M. Rachidi, and O. Saeki. On ∞-generalized Fibonacci sequences. Fibonacci Quarterly, 37(3):223–232, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/37-3/motta.pdf.

Motta:2000:CGF

[MRS00] W. Motta, M. Rachidi, and O. Saeki. Convergent ∞-generalized Fibonacci sequences. Fibonacci Quarterly, 38(4):326–333, August 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/38-4/motta.pdf.

Morito:1979:FGS

[MS79] Susumu Morito and Harvey M. Salkin. Finding the general solution of a linear Diophantine equation. Fibonacci Quarterly, 17(4): 361–368, December 1979. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/17-4/morito.pdf.

McCarthy:1990:GFS

[MS90] P. J. McCarthy and R. Sivaramakrishnan. Generalized Fibonacci sequences via arithmetical functions. Fibonacci Quarterly, 28(4): 363–370, November 1990. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/28-4/mccarthy.pdf.

Miller:1992:GIS

[MS92] Allen R. Miller and H. M. Srivastava. On Glaisher’s infinite sums involving the inverse tangent function. Fibonacci Quarterly, 30 (4):290–294, November 1992. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/30-4/miller.
pdf.

Melham:1994:SCP

[MS94] R. S. Melham and A. G. Shannon. Some congruence properties of generalized second-order integer sequences. Fibonacci Quarterly, 32(5):424–428, November 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-
5/melham.pdf.

Melham:1995:GRD

[MS95a] R. S. Melham and A. G. Shannon. A generalization of a result of D’Ocagne. Fibonacci Quarterly, 33(2):135–138, May 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-2/melham2.pdf.

Melham:1995:GCI

[MS95b] R. S. Melham and A. G. Shannon. A generalization of the Catalan identity and some consequences. Fibonacci Quarterly, 33(1):82– 84, February 1995. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/33-1/melham4.pdf.

Melham:1995:GSS

[MS95c] R. S. Melham and A. G. Shannon. Generalizations of some simple congruences. Fibonacci Quarterly, 33(2):126–130, May 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-2/melham1.pdf.

Melham:1995:ITH

[MS95d] R. S. Melham and A. G. Shannon. Inverse trigonometric hyperbolic summation formulas involving generalized Fibonacci numbers. Fibonacci Quarterly, 33(1):32–40, February 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-1/melham2.pdf.

Melham:1995:RSC

[MS95e] R. S. Melham and A. G. Shannon. On reciprocal sums of Chebyshev related sequences. Fibonacci Quarterly, 33(3):194–202, June 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/33-3/melham.pdf.

Melham:1995:SIS

[MS95f] R. S. Melham and A. G. Shannon. Some infinite series summations using power series evaluated at a matrix. Fibonacci Quarterly, 33(1):13–20, February 1995. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/33-1/melham1.
pdf.

Melham:1995:SSI

[MS95g] R. S. Melham and A. G. Shannon. Some summation identities using generalized Q-matrices. Fibonacci Quarterly, 33(1):64–73, February 1995. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/33-1/melham3.pdf.

Mascagni:2000:ASS

[MS00a] Michael Mascagni and Ashok Srinivasan. Algorithm 806: SPRNG: a scalable library for pseudorandom number generation. ACM Transactions on Mathematical Software, 26(3):436– 461, September 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://doi.acm.org/10. 1145/358407.358427. See correction [MS00b].

Mascagni:2000:CAS

[MS00b] Michael Mascagni and Ashok Srinivasan. Corrigendum: Algorithm 806: SPRNG: a scalable library for pseudorandom number generation. ACM Transactions on Mathematical Software, 26(4): 618–619, December 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://doi.acm.org/10.

1145/365723.365738. See [MS00a].

Margenstern:2003:FTC

[MS03] M. Margenstern and G. Skordev. Fibonacci type coding for the regular rectangular tilings of the hyperbolic plane. J.UCS: The Journal of Universal Computer Science, 9(5):398–422, May 28, 2003. CODEN ???? ISSN 0948-6968. URL http://www.jucs.

org/jucs_9_5/fibonacci_type_coding_for.

Mascagni:2004:PPM

[MS04] Michael Mascagni and Ashok Srinivasan. Parameterizing parallel multiplicative lagged-Fibonacci generators. Parallel Computing, 30(7):899–916, July 2004. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Martinsen:2015:ABG

[MS15a] Corey Martinsen and Pantelimon Stanica. Asymptotic behavior of gaps between roots of weighted factorials. Fibonacci Quarterly, 53(3):213–??, August 2015. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/53-3/
martinsen.pdf.

Moyon:2015:LFO

[MS15b] Marc Moyon and Maryvonne Spiesser. L’arithmétique des fractions dans l’oeuvre de Fibonacci: fondements & usages. Archive for History of Exact Sciences, 69(4):391–427, July 2015. CO-DEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00407-015-0155-y.

Mahanta:2020:FLR

[MS20] Pankaj Jyoti Mahanta and Manjil P. Saikia. A family of lacunary recurrences for Lucas numbers. Fibonacci Quarterly, 58(4):356– ??, November 2020. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/58-4/mahanta.pdf.

Martin-Sanchez:1995:GIA

[MSPF95] Óscar Martín-Sánchez and Cristóbal Pareja-Flores. A gentle introduction to algorithm complexity for CS1 with nine variations on a theme by Fibonacci. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 27(2):49–56, June 1995. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Moghaddamfar:2008:CMR

[MSS08] Ali Reza Moghaddamfar, S. Navid Salehy, and S. Nima Salehy. Certain matrices related to the Fibonacci sequence having recursive entries. Electronic Journal of Linear Algebra, 17(1):543–576, 2008. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (elec-

tronic). URL http://repository.uwyo.edu/ela/vol17/iss1/
40.

Marques:2013:DPC

[MST13] Diego Marques, James A. Sellers, and Pavel Trojovský. On divisibility properties of certain Fibonomial coefficients by a prime. Fibonacci Quarterly, 51(1):78–??, February 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/51-1/marques2.pdf.

Miller:2022:WSGa

[MSY22] Steven J. Miller, Eliel Sosis, and Jingkai Ye. Winning strategies for generalized Zeckendorf games. Fibonacci Quarterly, 60(5): 270–??, December 2022. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/60-5/miller1.pdf.

Marsaglia:1985:MSR

[MT85] George Marsaglia and Liang-Huei Tsay. Matrices and the structure of random number sequences. Linear Algebra and its Applications, 67:147–156, 1985. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Marsaglia:2002:SDP

[MT02] George Marsaglia and Wai Wan Tsang. Some difficult-to-pass tests of randomness. Journal of Statistical Software, 7(3):1–8, ???? 2002. CODEN JSSOBK. ISSN 1548-

7660. URL http://www.jstatsoft.org/v07/i03;   http:
//www.jstatsoft.org/v07/i03/tuftests.c;   http://www.
jstatsoft.org/v07/i03/tuftests.pdf;   http://www.jstatsoft.
org/v07/i03/updates.

Marques:2012:SNS

[MT12] Diego Marques and Pavel Trojovský. On some new sums of Fibonomial coefficients. Fibonacci Quarterly, 50(2):155–??, May 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/50-2/marques2.pdf.

Motegi:2017:GTE

[MT17] Kimihiko Motegi and Masakazu Teragaito. Generalized torsion elements and bi-orderability of 3-manifold groups. Canadian Mathematical Bulletin = Bulletin canadien de mathématiques, 60 (4):830–??, December 2017. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Mullen:1980:LPPb

[Mul80a] Gary L. Mullen. Local permutation polynomials in three variables over Z_{p}. Fibonacci Quarterly, 18(3):208–213, October 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-3/mullen.pdf.

Mullen:1980:LPPa

[Mul80b] Gary L. Mullen. Local permutation polynomials over Z_{p}. Fibonacci Quarterly, 18(2):104–107, April 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-2/
mullen.pdf.

Muller:2006:SRW

[Mül06] Siguna Müller. Some remarks on Williams’ public-key crypto functions. Fibonacci Quarterly, 44(3):224–234, August 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/44-3/muller.pdf.

Munzenrider:1971:NA

[Mun71] Joseph P. Munzenrider. A new anthesis. Fibonacci Quarterly, 9(2):163–176, April 1971. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/9-2/munzenrider.pdf.

Munarini:2005:GFN

[Mun05] Emanuele Munarini. Generalized q-Fibonacci numbers. Fibonacci Quarterly, 43(3):234–242, August 2005. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/43-3/
paper43-3-6.pdf.

Munagi:2007:CSF

[Mun07] Augustine O. Munagi. Combinations with successions and Fibonacci numbers. Fibonacci Quarterly, 45(2):104–114, May 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/45-2/munagi.pdf.

Munagi:2017:TAB

[Mun17] Augustine O. Munagi. Two applications of the bijection on Fibonacci set partitions. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/55-5/munagi.pdf.

Munagi:2020:IIC

[Mun20] Augustine O. Munagi. An identity for inverse-conjugate compositions. Fibonacci Quarterly, 58(5):161–??, December 2020. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/58-5/munagi.pdf.

Murthy:1982:FCN

[Mur82a] P. V. Satyanarayana Murthy. Fibonacci–Cayley numbers. Fibonacci Quarterly, 20(1):59–64, February 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/20-1/murthy1.pdf.

Murthy:1982:GSP

[Mur82b] P. V. Satyanarayana Murthy. Generalizations of some problems on Fibonacci numbers. Fibonacci Quarterly, 20(1):65–??, February 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/20-1/murthy2.pdf.

Muskat:1993:GFL

[Mus93] Joseph B. Muskat. Generalized Fibonacci and Lucas sequences and rootfinding methods. Mathematics of Computation, 61(203): 365–372, July 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Muwafi:1981:RTF

[Muw81] Amin A. Muwafi. A recursion-type formula for some partitions. Fibonacci Quarterly, 19(5):447–448, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-5/muwafi.pdf.

Meek:1984:SIR

[MV84] D. S. Meek and G. H. J. Van Rees. The solution of an iterated recurrence. Fibonacci Quarterly, 22(2):101–104, May 1984. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-2/meek.pdf.

Moll:1991:FNC

[MV91] Richard J. Moll and Shankar M. Venkatesan. Fibonacci numbers are not context-free. Fibonacci Quarterly, 29(1):59–61, February 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/29-1/moll.pdf.

Moll:2015:GBN

[MV15] Victor H. Moll and Christophe Vignat. Generalized Bernoulli numbers and a formula of Lucas. Fibonacci Quarterly, 53(4): 349–??, November 2015. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/53-4/moll.pdf.

Moser:1973:MR

[MW73] Leo Moser and Max Wyman. Multiple reflections. Fibonacci Quarterly, 11(3):302–306, October 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
3/moser.pdf.

Mollin:1987:NPN

[MW87] R. A. Mollin and P. G. Walsh. On nonsquare powerful numbers. Fibonacci Quarterly, 25(1):34–37, February 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-1/mollin.pdf.

Miller:1992:AN

[MW92] Gordon L. Miller and Mary T. Whalen. Armstrong numbers: 153 = 1^{3} + 5^{3} + 3^{3}. Fibonacci Quarterly, 30(3):221–224, August 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/30-3/miller.pdf.

Moss:2022:FAE

[MW22] Patrick Moss and Tom Ward. Fibonacci along even powers is (almost) realizable. Fibonacci Quarterly, 60(1):40–??, February 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/60-1/moss.pdf.

Myers:1975:STW

[Mye75] B. R. Myers. Spanning trees, weighted compositions, Fibonacci numbers, and resistor networks. SIAM Review, 17(3):465–474, ???? 1975. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Marsaglia:1991:NCR

[MZ91] George Marsaglia and Arif Zaman. A new class of random number generators. Annals of Applied Probability, 1(3):462–480, August 1991. CODEN ???? ISSN 1050-5164. URL http:/

/projecteuclid.org/euclid.aoap/1177005878. See popular

description in [Pet91]. See remarks in [EH95, TLC93] about the extremely bad lattice structure in high dimensions of the generators proposed in this paper.

Minqiang:1992:PML

[MZ92] Huang Minqiang and Dai Zongduo. Projective maps of linear recurring sequences with maximal p-adic periods. Fibonacci Quarterly, 30(2):139–143, May 1992. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/30-2/minqiang.
pdf.

Marsaglia:1993:MTR

[MZ93] George Marsaglia and Arif Zaman. Monkey tests for random number generators. Computers and Mathematics and Applications, 26(9):1–10, November 1993. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). See also [PW95].

Ma:2007:SII

[MZ07] Rong Ma and Wenpeng Zhang. Several identities involving the Fibonacci numbers and Lucas numbers. Fibonacci Quarterly, 45 (2):164–170, May 2007. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/45-2/ma.pdf.

Marsaglia:1990:TUR

[MZT90] George Marsaglia, Arif Zaman, and Wai Wan Tsang. Toward a universal random number generator. Statistics & Probability Letters, 9(1):35–39, January 1990. CODEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

Nagasaka:1984:DPR

[Nag84] Kenji Nagasaka. Distribution property of recursive sequences defined by u_{n+1} ≡ u_{n} + u_{n}^{−1} (mod m). Fibonacci Quarterly, 22 (1):76–81, February 1984. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/22-1/nagasaka.pdf.

Nagy:2001:RPC

[Nag01] Judit Nagy. Rational points in Cantor sets. Fibonacci Quarterly, 39(3):238–241, June 2001. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/39-3/nagy.pdf.

Najar:1989:OGU

[Naj89] Rudolph M. Najar. Operations on generators of unitary amicable pairs. Fibonacci Quarterly, 27(2):144–152, May 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-2/najar.pdf.

Najar:1993:RPT

[Naj93] Rudolph M. Najar. Recursions and Pascal-type triangles. Fibonacci Quarterly, 31(4):290–293, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-4/najar.pdf.

Nakata:2014:APP

[Nak14] Toshio Nakata. Another probabilistic proof of a binomial identity. Fibonacci Quarterly, 52(2):139–??, May 2014. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/52-
2/nakata.pdf.

Nakagawa:2020:TSL

[Nak20] Kouichi Nakagawa. A triangle with sides lengths of a rational power of the plastic constant. Fibonacci Quarterly, 58(5):166– ??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/58-5/nakagawa.pdf.

Naor:1965:LE

[Nao65] P. Naor. Letter to the editor. Fibonacci Quarterly, 3(1):71–73, February 1965. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/3-1/naor.pdf.

Nash:1976:SOF

[Nas76] Humphrey Nash. Some operational formulas. Fibonacci Quarterly, 14(1):1–8, February 1976. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/14-1/nash.pdf.

Navas:2001:ACF

[Nav01] Luis Navas. Analytic continuation of the Fibonacci Dirichlet series. Fibonacci Quarterly, 39(5):409–418, November 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-5/navas.pdf.

Naysmith:1965:OA

[Nay65] Betty Naysmith. Op art. Fibonacci Quarterly, 3(4):330–??, December 1965. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/3-4/naysmith.pdf.

Necochea:1989:CSQ

[NB89] Alejandro Necochea and Larry Bennett. On a certain sequence of quotients of a sequence. Fibonacci Quarterly, 27(1):70–75, February 1989. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/27-1/necochea.pdf.

Neufeld:1985:LSS

[NC85] Eric M. Neufeld and Charles J. Colbourn. Lucas sequences in subgraph counts of series-parallel and related graphs. Fibonacci Quarterly, 23(4):330–337, November 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-
4/neufeld.pdf.

Nandi:1987:AGL

[ND87] S. B. Nandi and S. K. Dutta. On associated and generalized Lah numbers and applications to discrete distributions. Fibonacci Quarterly, 25(2):128–136, May 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-2/
nandi.pdf.

Neal:2007:FFH

[Nea07] David K. Neal. Fibonacci fractions from Heron’s square root approximation of the golden ratio. Fibonacci Quarterly, 45(1): 35–38, February 2007. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/45-1/neal.pdf.

Neff:1967:MLP

[Nef67] John D. Neff. A Markov limit process involving Fibonacci numbers. Fibonacci Quarterly, 5(2):179–182, April 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/5-2/neff.pdf.

Neunhauserer:2017:RFR

[Neu17] Jörg Neunhäuserer. Return of Fibonacci random walks. Statistics & Probability Letters, 121(??):51–53, February 2017. CO-DEN SPLTDC. ISSN 0167-7152 (print), 1879-2103 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0167715215301917.

Niederreiter:1972:DFN

[Nie72] Harald Niederreiter. Distribution of Fibonacci numbers mod 5^{k}. Fibonacci Quarterly, 10(4):373–374, October 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-4/niederreiter.pdf.

Noppakaew:2023:FNK

[NKW23] Passawan Noppakaew, Pavita Kanwarunyu, and Parit Wanitchatchawan. k-Fibonacci numbers and k-Lucas numbers in Beatty sequences generated by powers of metallic means. Fibonacci Quarterly, 61(2):167–??, May 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/61-
2/noppakaew.pdf.

Nishihara:1987:BSR

[NN87] Seiichi Nishihara and Hiroji Nishino. Binary search revisited: Another advantage of Fibonacci search. IEEE Transactions on Computers, C-36(9):1132–1135, September 1987. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=5009548.

Nalli:2015:TOV

[NO15] Ayse Nalli and Cagla Ozyilmaz. The third order variations on the Fibonacci universal code. Journal of Number Theory, 149(??): 15–32, April 2015. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X14002492.

Nodine:1990:NRT

[Nod90] Mark Nodine. Note on the resistance through a static carry lookahead gate. Fibonacci Quarterly, 28(2):102–106, May 1990. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/28-2/nodine.pdf.

Norden:1964:PM

[Nor64] H. Norden. Proportions in music. Fibonacci Quarterly, 2(3):219– ??, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/2-3/norden.pdf.

Norden:1972:PC

[Nor72] Hugo Norden. Proportions and the composer. Fibonacci Quarterly, 10(3):319–323, April 1972. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/10-3/norden.
pdf.

Norden:1976:PNC

[Nor76] Hugo Norden. Per Nørgård’s “Canon”. Fibonacci Quarterly, 14 (2):126–128, April 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-2/norden.pdf.

Norfleet:2005:CSO

[Nor05] Mark Norfleet. Characterization of second-order strong divisibility sequences of polynomials. Fibonacci Quarterly, 43(2):166–169, May 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Papers/43-2/paper43-2-12.pdf.

Northshield:2014:TAS

[Nor14] Sam Northshield. Three analogues of Stern’s diatomic sequence. Fibonacci Quarterly, 52(5):168–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers1/52-5/Northshield.pdf.

Nordgren:2016:FCM

[Nor16] Ronald P. Nordgren. On Franklin and complete magic square matrices. Fibonacci Quarterly, 54(4):304–??, November 2016. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/54-4/nordgren.pdf.

Northshield:2019:RCR

[Nor19] Sam Northshield. ℜ^{3} counting the rationals. Fibonacci Quarterly, 57(5):111–129, December 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/57-5/
northshield.pdf.

Northshield:2020:TCO

[Nor20] Sam Northshield. Topographs; Conway and otherwise. Fibonacci Quarterly, 58(5):172–??, December 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/58-
5/northshield.pdf.

Northshield:2022:SGFa

[Nor22] Sam Northshield. Some generalizations of a formula of Reznick. Fibonacci Quarterly, 60(5):299–??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/60-5/northshield.pdf.

Nowak:1986:MTC

[Now86] Werner Georg Nowak. On the minimum of a ternary cubic form. Fibonacci Quarterly, 24(2):129–132, May 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/24-2/nowak.pdf.

Nyul:2015:EVW

[NR15] Gábor Nyul and Bettina Rauf. On the existence of van der Waerden type numbers for linear recurrence sequences with constant coefficients. Fibonacci Quarterly, 53(1):53–??, February 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/53-1/nyul.pdf.

Nyul:2017:LSH

[NR17] Gábor Nyul and Gabriella Rácz. Lucas sequences and the Hosoya index of graphs. Fibonacci Quarterly, 55(4):340–??, November 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/55-4/nyul.pdf.

Nagasaka:1990:APL

[NS90] Kenji Nagasaka and Jau-Shyong Shiue. Asymptotic positiveness of linear recurrence sequences. Fibonacci Quarterly, 28(4):340– 346, November 1990. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/28-4/nagasaka.pdf.

Niederreiter:1994:INF

[NS94] Harald Niederreiter and Ian H. Sloan. Integration of nonperiodic functions of two variables by Fibonacci lattice rules. Journal of Computational and Applied Mathematics, 51(1):57–70, May 30, 1994. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037704279200004S.

Niederreiter:1995:MCQ

[NS95] Harald Niederreiter and Peter Jau-Shyong Shiue, editors. Monte Carlo and quasi-Monte Carlo methods in scientific computing: proceedings of a conference at the University of Nevada, Las Vegas, Nevada, USA, June 23–25, 1994, volume 106 of Lecture notes in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94577-6 (softcover). LCCN Q183.9 .M66 1995.

Nymann:1999:ENI

[NS99] J. E. Nymann and Ricardo A. Sáenz. Eulerian numbers: Inversion formulas and congruences modulo a prime. Fibonacci Quarterly, 37(2):154–161, May 1999. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/37-2/nymann.pdf.

Nyblom:2001:SFD

[NS01] M. A. Nyblom and B. G. Sloss. On the solvability of a family of Diophantine equations. Fibonacci Quarterly, 39(1):58–65, February 2001. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/39-1/nyblom.pdf.

Newell:2005:PF

[NS05a] Alan C. Newell and Patrick D. Shipman. Plants and Fibonacci. Journal of Statistical Physics, 121(5–6):937–968, December 2005. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10955-005-8665-7.

Nkwanta:2005:PWR

[NS05b] Asamoah Nkwanta and Louis W. Shapiro. Pell walks and Riordan matrices. Fibonacci Quarterly, 43(2):170–180, May 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/43-2/paper43-2-13.pdf.

Neumann:1979:SSL

[NW79] B. H. Neumann and L. G. Wilson. Some sequences like Fibonacci’s. Fibonacci Quarterly, 17(1):80–83, February 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-1/neumann.pdf. See corrigenda [NW83].

Neumann:1983:CSS

[NW83] B. H. Neumann and L. G. Wilson. Corrigenda to “Some Sequences Like Fibonacci’s”. Fibonacci Quarterly, 21(3):229–??, August 1983. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/21-3/neumann.pdf. See [NW79].

Nyblom:1998:NSA

[Nyb98] M. A. Nyblom. A note on the set of almost-isosceles right-angled triangles. Fibonacci Quarterly, 36(4):319–322, August 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-4/nyblom.pdf.

Nyblom:1999:GBT

[Nyb99a] M. A. Nyblom. On a generalization of the binomial theorem. Fibonacci Quarterly, 37(1):3–13, February 1999. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/37-1/
nyblom.pdf.

Nyblom:1999:IVS

[Nyb99b] M. A. Nyblom. On irrational valued series involving generalized Fibonacci numbers. Fibonacci Quarterly, 37(4):299–304, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/37-4/nyblom.pdf.

Nyblom:2001:IVS

[Nyb01a] M. A. Nyblom. On irrational valued series involving generalized Fibonacci numbers II. Fibonacci Quarterly, 39(2):149–157, May 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/39-2/nyblom.pdf.

Nyblom:2001:RID

[Nyb01b] M. A. Nyblom. On the representation of the integers as a difference of nonconsecutive triangular numbers. Fibonacci Quarterly, 39(3):256–263, June 2001. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/39-3/nyblom.pdf.

Nyblom:2002:RID

[Nyb02] M. A. Nyblom. On the representation of the integers as a difference of squares. Fibonacci Quarterly, 40(3):243–246, June 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-3/nyblom.pdf.

Nyblom:2003:NIP

[Nyb03] M. A. Nyblom. A non-integer property of elementary symmetric functions in reciprocals of generalized Fibonacci numbers. Fibonacci Quarterly, 41(2):152–155, May 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-2/
nyblom.pdf.

Nyblom:2004:CFT

[Nyb04] M. A. Nyblom. On the construction of a family of transcendental valued infinite products. Fibonacci Quarterly, 42(4):353–358, November 2004. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Papers1/42-4/quartnyblom04_2004.pdf.

Nyblom:2005:SRN

[Nyb05] M. A. Nyblom. On the spectrum of real numbers revisited. Fibonacci Quarterly, 43(4):299–301, November 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers/43-4/paper43-4-2.pdf.

Nyblom:2008:CFA

[Nyb09] Michael A. Nyblom. On the construction of a family of almost power free sequences. Fibonacci Quarterly, 46/47(4):366–368, November 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/46_47-4/nyblom.pdf.

Nyblom:2012:SGP

[Nyb12] Michael A. Nyblom. Sophie Germain primes and the exceptional values of the equal-sum-and-product problem. Fibonacci Quarterly, 50(1):58–??, February 2012. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/50-1/nyblom.
pdf.

Nyblom:2023:DTD

[Nyb23] Michael Nyblom. Deleting terms of the divergent p-series and reciprocals of primes series using the Thue–Morse sequence. Fibonacci Quarterly, 61(4):339–??, November 2023. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/61-4/nyblom.pdf.

Ozkan:2003:SFSa

[ÖAD03a] Engin Özkan, Hüseyin Aydın, and Ramazan Dikici. 3-step Fibonacci series modulo m. Applied Mathematics and Computation, 143(1):165–172, October 20, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Ozkan:2003:AFS

[ÖAD03b] Engin Özkan, Hüseyin Aydın, and Ramazan Dikici. Applications of Fibonacci sequences in a finite nilpotent group. Applied Mathematics and Computation, 141(2–3):565–578, September 5, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Oaks:2009:RBB

[Oak09] Jeffrey A. Oaks. Reviews: Fibonacci’s De Practica Geometrie — by Barnabas Hughes. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 51(2):168–169, May 2009. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

OConnell:1972:RIS

[O’C72] Roger O’Connell. Representations of integers as sums of Fibonacci squares. Fibonacci Quarterly, 10(1):103–112, January 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/10-1/oconnell.pdf.

ODonnell:1979:NPT

[O’D79a] William J. O’Donnell. A note on a Pell-type sequence. Fibonacci Quarterly, 17(1):49–50, February 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
1/odonnell1.pdf.

ODonnell:1979:TTC

[O’D79b] William J. O’Donnell. Two theorems concerning hexagonal numbers. Fibonacci Quarterly, 17(1):77–79, February 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-1/odonnell2.pdf.

Ogur:2020:SOS

[Ogu20] Oguz Ogur. Superposition operators on sequence spaces ℓ_{p}(F) derived by using matrix of Fibonacci numbers. Linear Multilinear Algebra, 68(10):2087–2098, 2020. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Olagunju:2000:RSD

[Ola00] Amos O. Olagunju. The role of scientific discovery in teaching and learning of computer science. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 32(4):28– 31, December 2000. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Oliverio:1996:SGP

[Oli96] Paul Oliverio. Self-generating Pythagorean quadruples and Ntuples. Fibonacci Quarterly, 34(2):98–101, May 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-2/oliverio.pdf.

Ollerton:2022:CNN

[Oll22] R. L. Ollerton. Catalan numbers and non-intersecting lattice paths. Fibonacci Quarterly, 60(3):238–??, August 2022. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/60-3/ollerton.pdf.

Ohtsuka:2008:SRF

[ON09] Hideyuki Ohtsuka and Shigeru Nakamura. On the sum of reciprocal Fibonacci numbers. Fibonacci Quarterly, 46/47(2):153–159, May 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/46_47-2/ohtsuka.pdf.

Onderdonk:1970:PFN

[Ond70] Philip B. Onderdonk. Pineapples and Fibonacci numbers. Fibonacci Quarterly, 8(5):507–508, December 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/8-5/onderdonk.pdf.

Onphaeng:2014:SDP

[OP14] Kritkhajohn Onphaeng and Prapanpong Pongsriiam. Subsequences and divisibility by powers of the Fibonacci numbers. Fibonacci Quarterly, 52(2):163–??, May 2014. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/52-
2/onphaeng.pdf.

Onphaeng:2018:CED

[OP18] Kritkhajohn Onphaeng and Prapanpong Pongsriiam. The converse of exact divisibility by powers of the Fibonacci and Lucas numbers. Fibonacci Quarterly, 56(4):296–??, November 2018.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/56-4/onphaeng.pdf.

OConnor:1999:RF

[OR99] John J. O’Connor and Edmund F. Robertson. References for

Fibonacci, 1999. URL http://www-groups.dcs.st-and.ac.
uk/~history/References/Fibonacci.html;   https://www.
math.utah.edu/pub/tex/bib/fibquart.bib.

Ochem:2014:ARR

[OR14] Pascal Ochem and Michaël Rao. Another remark on the radical of an odd perfect number. Fibonacci Quarterly, 52(3):215–??, August 2014. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/52-3/ochem.pdf.

Ordower:1993:CPG

[Ord93] Marc S. Ordower. The connectivity of a particular graph. Fibonacci Quarterly, 31(3):276–??, August 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-3/ordower.pdf.

Ore:1967:INT

[Øre67] Øystein Øre. Invitation to number theory, volume 20 of New mathematical library. Random House, New York, NY, USA, 1967. viii + 129 pp. LCCN QA241 .O68.

Ottmann:1984:BST

[ORSW84] Thomas Ottmann, Arnold L. Rosenberg, Hans-Werner Six, and Derick Wood. Binary search trees with binary comparison cost. International Journal of Computer and Information Sciences, 13 (2):77–101, April 1984. CODEN IJCIAH. ISSN 0091-7036.

Ollerton:1998:SPG

[OS98] Richard L. Ollerton and Anthony G. Shannon. Some properties of generalized Pascal squares and triangles. Fibonacci Quarterly, 36 (2):98–109, May 1998. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/36-2/ollerton.pdf. See er-

ratum [OS21].

Ollerton:2005:FPG

[OS05] R. L. Ollerton and A. G. Shannon. Further properties of generalized binomial coefficient k-extensions. Fibonacci Quarterly, 43(2):

124–129, May 2005. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Papers/43-2/paper43-2-5.pdf.

Ollerton:2020:NBS

[OS20] R. L. Ollerton and A. G. Shannon. A note on Brousseau’s summation problem. Fibonacci Quarterly, 58(5):190–??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/58-5/ollerton.pdf.

Ollerton:2021:ESP

[OS21] Richard L. Ollerton and Anthony G. Shannon. Erratum — some properties of generalized Pascal squares and triangles. Fibonacci Quarterly, 59(3):272–??, August 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/59-3/
errata05012021rev.pdf. See [OS98].

Osler:2007:VLP

[Osl07] Thomas J. Osler. Vieta-like products of nested radicals with Fibonacci and Lucas numbers. Fibonacci Quarterly, 45(3):202– 204, August 2007. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/45-3/osler.pdf.

Ocal:2005:RGF

[ÖTA05] Ahmet Ali Öcal, Naim Tuglu, and Ercan Altinişik. On the representation of k-generalized Fibonacci and Lucas numbers. Applied Mathematics and Computation, 170(1):584–596, November 1, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Overholt:1965:IFG

[Ove65] K. J. Overholt. An instability in the Fibonacci and Golden Section search methods. Nordisk Tidskrift for Informationsbehandling, 5(4):284–286, 1965. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Overholt:1967:NAF

[Ove67] K. J. Overholt. Note on Algorithm 2: Fibonacci search. The Computer Journal, 9(4):414, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
04/tiff/414.tif. See [PP79].

Overholt:1973:EFS

[Ove73] K. J. Overholt. Efficiency of the Fibonacci search method. BIT (Nordisk tidskrift for informationsbehandling), 13(1):92– 96, March 1973. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0006-
3835&volume=13&issue=1&spage=92.

Oliver:1964:SSM

[OW64] L. T. Oliver and D. J. Wilde. Symmetric sequential minimax search for a maximum. Fibonacci Quarterly, 2(3):169–175, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/2-3/oliver.pdf.

Owings:1979:STF

[Owi79] James C. Owings, Jr. Solution of binom(y + 1, x) = binom(y, x + 1) in terms of Fibonacci numbers. Fibonacci Quarterly, 17(1): 67–68, February 1979. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/17-1/owings.pdf.

Owings:1987:SS

[Owi87] James C. Owings, Jr. Solution of the system a^{2} ≡ −1 (mod b), b^{2} ≡ −1 (mod a). Fibonacci Quarterly, 25(3):245–249, August 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/25-3/owings.pdf.

Ozbolt:2012:NSD

[Ozb12] Joseph S. Ozbolt. A new sequence derived from a combination of cubes with volume F_{n}^{3}. Fibonacci Quarterly, 50(1):19–??, February 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/50-1/ozbolt.pdf.

Ozeki:2005:WFL

[Oze05] Kiyota Ozeki. On weighted Fibonacci and Lucas sums. Fibonacci Quarterly, 43(2):104–107, May 2005. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/43-2/
paper43-2-2.pdf.

Ozeki:2008:MS

[Oze09] Kiyota Ozeki. On Melham’s sum. Fibonacci Quarterly, 46/47(2):107–110, May 2008/2009. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/46_47-
2/ozeki.pdf.

Ozkan:2003:SFSb

[Özk03] Engin Özkan. 3-step Fibonacci sequences in nilpotent groups. Applied Mathematics and Computation, 144(2–3):517–527, December 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Philippou:1991:GMF

[PA91] Andreas N. Philippou and Demetris L. Antzoulakos. Generalized multivariate Fibonacci polynomials of order K and the multivariate negative binomial distributions of the same order. Fibonacci Quarterly, 29(4):322–328, November 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-
4/philippou.pdf.

Padmavathamma:1989:AGF

[Pad89] Padmavathamma. On Andrews’ generalized Frobenius partitions. Fibonacci Quarterly, 27(2):125–130, May 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/27-2/padmavathamma.pdf.

Padmakumar:1994:JPM

[Pad94] T. V. Padmakumar. A juxtaposition property for the 4×4 magic square. Fibonacci Quarterly, 32(4):290–292, August 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-4/padmakumar.pdf.

Padmakumar:1997:SMS

[Pad97] T. V. Padmakumar. Strongly magic squares. Fibonacci Quarterly, 35(3):198–205, August 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-
3/padmakumar.pdf.

Page:1974:PQM

[Pag74] Warren Page. P.Q.M.-cycles, a generalized number problem. Fibonacci Quarterly, 12(4):323–325, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/12-4/page.pdf.

Panda:2007:SBC

[Pan07] G. K. Panda. Sequence balancing and cobalancing numbers. Fibonacci Quarterly, 45(3):265–271, August 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/45-3/panda.pdf.

Parker:1964:GTR

[Par64] Francis D. Parker. On the general term of a recursive sequence. Fibonacci Quarterly, 2(1):67–71, February 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-1/parker.pdf.

Parker:1968:FF

[Par68] Francis D. Parker. A Fibonacci function. Fibonacci Quarterly, 6(1):1–2, February 1968. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/6-1/parker.pdf.

Parberry:1969:RRP

[Par69] Edward A. Parberry. A recursion relation for populations of diatoms. Fibonacci Quarterly, 7(5):449–456, December 1969. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/7-5/parberry-a.pdf.

Parberry:1970:PPP

[Par70] Edward A. Parberry. On primes and pseudo-primes related to the Fibonacci sequence. Fibonacci Quarterly, 8(1):49–60, February 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-1/parberry.pdf.

Parberry:1977:TRR

[Par77a] E. A. Parberry. Two recursion relations for F (F (n)). Fibonacci Quarterly, 15(2):122–??, April 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-2/
parberry-a.pdf.

Park:1977:ZOSb

[Par77b] C. J. Park. Zero-one sequences and Stirling numbers of the first kind. Fibonacci Quarterly, 15(3):231–??, October 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-3/park2.pdf.

Park:1977:ZOSa

[Par77c] C. J. Park. Zero-one sequences and Stirling numbers of the second kind. Fibonacci Quarterly, 15(3):205–206, October 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-3/park1.pdf.

Park:1998:CP

[Par98] SeungKyung Park. Complete partitions. Fibonacci Quarterly, 36 (4):354–360, August 1998. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/36-4/park.pdf.

Parshall:2017:PAA

[Par17] Karen Hunger Parshall. A plurality of algebras, 1200–1600: algebraic Europe from Fibonacci to Clavius. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(1): 2–16, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17498430.2016.1225340.

Pasles:2010:FMM

[Pas10] Paul C. Pasles. Fibonacci matrices and modular forms. Fibonacci Quarterly, 48(4):317–323, November 2010. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/48-
4/pasles.pdf.

Paulsen:2002:PNM

[Pau02] William Paulsen. The prime number maze. Fibonacci Quarterly, 40(3):272–279, June 2002. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/40-3/paulsen.pdf.

Paulsen:2023:CWS

[Pau23] William Paulsen. Calkin–Wilf sequences for irrational numbers. Fibonacci Quarterly, 61(1):51–??, February 2023. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-1/paulsen.pdf.

Prasad:1989:SDF

[PB89] V. Siva Rama Prasad and M. V. S. Bhramarambica. On the Schnirelmann density of M-free integers. Fibonacci Quarterly, 27 (4):366–368, August 1989. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/27-4/prasad.pdf.

Pisano:2015:FAS

[PB15] Raffaele Pisano and Paolo Bussotti. Fibonacci and the Abacus Schools in Italy. Mathematical conceptual streams — education and its changing relationship with society. Almagest: International Journal for the History of Scientific Ideas, 6(2):126–164, 2015. CODEN ???? ISSN 1792-2593.

URL https://www.brepolsonline.net/doi/epdf/10.1484/J.
ALMAGEST.5.109664.

Porpodas:2013:LLA

[PC13] Vasileios Porpodas and Marcelo Cintra. LUCAS: latencyadaptive unified cluster assignment and instruction scheduling. ACM SIGPLAN Notices, 48(5):45–54, May 2013. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Panda:2015:PBN

[PD15] G. K. Panda and Ravi Kumar Davala. Perfect balancing numbers. Fibonacci Quarterly, 53(3):261–??, August 2015. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/53-3/panda.pdf.

Philippou:2018:SPI

[PD18] Andreas N. Philippou and Spiros D. Dafnis. A simple proof of an identity generalizing Fibonacci–Lucas identities. Fibonacci Quarterly, 56(4):334–??, November 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/56-
4/philippou.pdf.

Peck:1965:ESF

[Pec65] C. B. A. Peck. Exploring scalene Fibonacci polygons. Fibonacci Quarterly, 3(1):57–58, February 1965. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/3-1/peck.
pdf.

Pedersen:1971:AET

[Ped71] Jean J. Pedersen. Asymptotic Euclidean type constructions without Euclidean tools (recreational mathematics edited by Joseph S. Madachy). Fibonacci Quarterly, 9(2):199–216, April 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-2/pedersen.pdf.

Peleg:1972:OFF

[Pel72] Reuven Peleg. An old Fibonacci formula and stopping rules. Fibonacci Quarterly, 10(6):661–??, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-6/peleg.pdf.

Peng:2020:AIT

[Pen20] Wayne Peng. ABC implies there are infinitely many non– Fibonacci–Wieferich primes. Journal of Number Theory, 212(??): 354–375, July 2020. CODEN JNUTA9. ISSN 0022-314X (print),

1096-1658 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022314X1930410X.

Pepe:2003:LFE

[Pep03] Luigi Pepe. Leonardo Fibonacci in the eighteenth century. Bollettino di Storia delle Scienze Matematiche, 23(2):75–86 (2005), 2003. ISSN 0392-4432 (print), 1724-1650 (electronic). Leonardo Fibonacci. Matematica e societa nel Mediterraneo nel secolo XIII. Vol. I.

Perrine:2016:SPE

[Per16] Serge Perrine. Some properties of the equation x^{2} = 5y^{2} − 4. Fibonacci Quarterly, 54(2):172–??, May 2016. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/54-
2/perrine.pdf.

Pettorossi:1980:DAC

[Pet80] Alberto Pettorossi. Derivation of an O(k^{2} log n) algorithm for computing order-k Fibonacci numbers from the O(k^{3} log n) matrix multiplication method. Information Processing Letters, 11 (4–5):172–179, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peters:1981:EFB

[Pet81] James V. Peters. An equivalent form of Benford’s Law. Fibonacci Quarterly, 19(1):74–75, February 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
1/peters.pdf.

Pethe:1984:LPS

[Pet84] S. Pethe. On Lucas polynomials and some summation formulas for Chebychev polynomial sequences via them. Fibonacci Quarterly, 22(1):61–69, February 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-
1/pethe.pdf.

Pethe:1985:LFF

[Pet85] S. Pethe. On Lucas fundamental functions and Chebychev polynomial sequences. Fibonacci Quarterly, 23(1):57–65, February 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/23-1/pethe.pdf.

Peters:1986:TRG

[Pet86] J. M. H. Peters. Tenth roots and the golden ratio. Fibonacci Quarterly, 24(4):323–327, November 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-
4/peters.pdf.

Pethe:1988:SIT

[Pet88] S. Pethe. Some identities for Tribonacci sequences. Fibonacci Quarterly, 26(2):144–151, May 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-2/
pethe.pdf.

Peterson:1991:NRN

[Pet91] Ivars Peterson. Numbers at random: Number theory supplies a superior random-number generator. Science News (Washington, DC), 140(19):300–301, November 9, 1991. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic). URL http://

www.jstor.org/stable/3975915.

Petersen:1994:LFS

[Pet94] W. P. Petersen. Lagged Fibonacci series random number generators for the NEC SX-3. International Journal of High Speed Computing (IJHSC), 6(3):387–??, 1994. CODEN IHSCEZ. ISSN 0129-0533.

Peters:1996:TPF

[Pet96] J. M. H. Peters. A ten point FFT calculation which features the golden ratio. Fibonacci Quarterly, 34(4):323–325, August 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-4/peters.pdf.

Peterson:1999:FRU

[Pet99] Ivars Peterson. Fibonacci at random: Uncovering a new mathematical constant. Science News (Washington, DC), 155(24): 376–377, June 12, 1999. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic). URL http://www.jstor.org/

stable/4011459.

Petronilho:2012:GFS

[Pet12] J. Petronilho. Generalized Fibonacci sequences via orthogonal polynomials. Applied Mathematics and Computation, 218 (19):9819–9824, June 1, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300312002925.

Philippou:1989:CFT

[PG89] Andreas N. Philippou and Costas Georghiou. Convolutions of Fibonacci-type polynomials of order K and the negative binomial distributions of the same order. Fibonacci Quarterly, 27(3):209– 216, June 1989. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/27-3/philippou.pdf.

Panda:2013:DFM

[PG13] Geetanjali Panda and Suman Ghosh. A derivative free multidimensional optimal search method using Lucas number. Applied Mathematics and Computation, 219(12):6536–6541, February 15, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300312010922.

Philippou:1985:FTP

[PGP85] Andreas N. Philippou, Costas Georghiou, and George N. Philippou. Fibonacci-type polynomials of order K with probability applications. Fibonacci Quarterly, 23(2):100–105, May 1985. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-2/philippou.pdf.

Pond:1965:MFN

[PH65] Jeremy C. Pond and Donald F. Howells. More on Fibonacci Nim. Fibonacci Quarterly, 3(1):61–62, February 1965. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/3-1/pond.pdf.

Peterson:1975:SDE

[PH75] Brian Peterson and V. E. Hoggatt, Jr. On the solutions to the Diophantine equation x^{2} + xy − y^{2} = ±D, or the number of Fibonacci-type sequences with a given characteristic. Fibonacci Quarterly, 13(3):243–251, October 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-
3/peterson.pdf.

Pethe:1988:GGL

[PH88] S. Pethe and A. F. Horadam. Generalized Gaussian Lucas primordial functions. Fibonacci Quarterly, 26(1):20–30, February 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/26-1/pethe.pdf.

Phares:1984:GSF

[Pha84] Alain J. Phares. General solution of a Fibonacci-like recursion relation and applications. Fibonacci Quarterly, 22(1):29–41, February 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/22-1/phares.pdf.

Philippou:1983:NFS

[Phi83] Andreas N. Philippou. A note on the Fibonacci sequence of order K and the multinomial coefficients. Fibonacci Quarterly, 21(2): 82–86, May 1983. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/21-2/philippou.pdf.

Phillips:1984:ASF

[Phi84] G. M. Phillips. Aitken sequences and Fibonacci numbers. American Mathematical Monthly, 91(6):354–357, June/July 1984. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Phillips:2001:RNI

[Phi01] George M. Phillips. Report on the Ninth International Conference. Fibonacci Quarterly, 39(1):3–??, February 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-1/phillips.pdf.

Phillips:2004:REI

[Phi04] George M. Phillips. Report on the Eleventh International Conference on Fibonacci Numbers and their Applications. Fibonacci Quarterly, 42(4):359–360, November 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
4/quartreport04_2004.pdf.

Phillips:2007:IDS

[Phi07a] George M. Phillips. Index-doubling in sequences by Aitken acceleration. Fibonacci Quarterly, 45(4):313–321, November 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/45-4/phillips.pdf.

Phillips:2007:RTI

[Phi07b] George M. Phillips. Report on the Twelfth International Conference. Fibonacci Quarterly, 45(4):366–367, November 2007. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/45-4/quartphillips04_2007.pdf.

Philippou:2014:NMP

[Phi14] Andreas N. Philippou. A note on the modes of the Poisson distribution of order k. Fibonacci Quarterly, 52(3):203–??, August 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/52-3/philippou.pdf.

Pike:1967:NAF

[PHJ67] M. C. Pike, I. D. Hill, and F. D. James. Note on Algorithm 2: Fibonacci search, and on Algorithm 7: MINX. The Computer Journal, 9(4):416–417, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
04/tiff/416.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_09/Issue_04/tiff/417.tif. See [PP79].

Phong:1991:LPR

[Pho91] Bui Minh Phong. Lucas primitive roots. Fibonacci Quarterly, 29 (1):66–71, February 1991. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/29-1/phong.pdf.

Picutti:1979:BSL

[Pic79] Ettore Picutti. The Book of squares of Leonardo of Pisa and the problems of indeterminate analysis in the Palatine Codex 557 of the National Library in Florence. Introduction and comments. Physis: Rivista Internazionale di Storia della Scienza, 21(1–4): 195–339 (1980), 1979. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Picutti:1981:SNC

[Pic81] Ettore Picutti. Sui numeri congruo-congruenti di Leonardo Pisano. (Italian) [Leonardo of Pisa’s congruous-congruent numbers]. Physis: Rivista Internazionale di Storia della Scienza, 23 (2):141–170, 1981. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Picutti:1983:FLP

[Pic83a] Ettore Picutti. The “Flos” of Leonardo Pisano (dal codice E. 75 P. sup. della Biblioteca Ambrosiana di Milano). Physis: Rivista Internazionale di Storia della Scienza, 25(2):293–387, 1983. CO-DEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Picutti:1983:BLP

[Pic83b] Ettore Picutti. Il Flos di Leonardo Pisano dal codice E.75.P. sup. della Biblioteca Ambrosiana di Milano. (Italian) [The “Flos” of Leonardo Pisano (from the Codex E.75.P. sup. of the Ambrosia Library of Milan)]. Physis: Rivista Internazionale di Storia della Scienza, 25(2):293–387, 1983. CODEN PYSSA3. ISSN 0031-9414 (print), 2038-6265 (electronic).

Pihko:1988:FLR

[Pih88] Jukka Pihko. On Fibonacci and Lucas representations and a theorem of Lekkerkerker. Fibonacci Quarterly, 26(3):256–261, August 1988. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/26-3/pihko.pdf.

Pihko:1991:NTS

[Pih91] Jukka Pihko. A note on a theorem of Schinzel. Fibonacci Quarterly, 29(4):333–338, November 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-
4/pihko.pdf.

Pihko:1992:SHS

[Pih92] Jukka Pihko. On sequences having same minimal elements in the Lemoine–Kátai algorithm. Fibonacci Quarterly, 30(4):344– 348, November 1992. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/30-4/pihko.pdf.

Pihko:2001:RGO

[Pih01] Jukka Pihko. Remarks on the “greedy odd” Egyptian fraction algorithm. Fibonacci Quarterly, 39(3):221–227, June 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-3/pihko.pdf.

Pihko:2010:RGO

[Pih10] Jukka Pihko. Remarks on the “greedy odd” Egyptian fraction algorithm II. Fibonacci Quarterly, 48(3):202–208, August 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/48-3/pihko.pdf.

Pillichshammer:2012:DVC

[Pil12] Friedrich Pillichshammer. On the discrepancy of the Van Der Corput sequence indexed by Fibonacci numbers. Fibonacci Quarterly, 50(3):235–??, August 2012. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/50-3/
pillichshammer.pdf.

Ping:2002:PFF

[Pin02] Sun Ping. Partition forms of Fibonacci numbers. Fibonacci Quarterly, 40(3):287–289, June 2002. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/40-3/sun.pdf.

Pisano:2002:FBA

[Pis02] Leonardo Pisano. Fibonacci’s Liber abaci. Sources and Studies in the History of Mathematics and Physical Sciences. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95419-8. viii + 636 pp. A translation into modern English of Leonardo Pisano’s ıt Book of calculation, Translated from the Latin and with an introduction, notes and bibliography by L. E. Sigler.

Podani:2018:HFD

[PKS18] János Podani, Ádám Kun, and András Szilágyi. How fast does Darwin’s elephant population grow? Journal of the History of Biology, 51(2):259–281, June 2018. CODEN JHBIA9. ISSN 0022-5010 (print), 1573-0387 (electronic). URL http:/

/link.springer.com/article/10.1007/s10739-017-9488-5;
 http://link.springer.com/content/pdf/10.1007/s10739-
017-9488-5.pdf.

Posamentier:2007:FFN

[PL07] Alfred S. Posamentier and Ingmar Lehmann. The Fabulous Fibonacci Numbers. Prometheus Books, Amherst, NY, USA, 2007.

ISBN 1-59102-475-7. 385 pp. LCCN QA241 .P665 2007. URL

http://www.loc.gov/catdir/toc/ecip073/2006035406.html.

Pla:1994:AND

[Pla94] Juan Pla. An “all or none” divisibility property for a class of Fibonacci-like sequences of integers. Fibonacci Quarterly, 32(3): 226–227, June 1994. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/32-3/pla.pdf.

Pla:1995:SCA

[Pla95] Juan Pla. Some conditions for “all or none” divisibility of a class of Fibonacci-like sequences. Fibonacci Quarterly, 33(5):464–465, November 1995. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/33-5/pla.pdf.

Pla:1996:ECS

[Pla96a] Juan Pla. On the existence of couples of second-order linear recurrences with reciprocal representation properties for their Fibonacci sequences. Fibonacci Quarterly, 34(5):409–412, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/34-5/pla1.pdf.

Pla:1996:PPG

[Pla96b] Juan Pla. On the possibility of programming the general 2-by-2 matrix on the complex field. Fibonacci Quarterly, 34(5):440–443, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/34-5/pla2.pdf.

Pla:1997:PMP

[Pla97a] Juan Pla. On periods modulo a prime of some classes of sequences of integers. Fibonacci Quarterly, 35(1):54–56, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-1/pla.pdf.

Pla:1997:SIB

[Pla97b] Juan Pla. The sum of inverses of binomial coefficients revisited. Fibonacci Quarterly, 35(4):342–345, November 1997. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-4/pla.pdf.

Prichett:1981:DAD

[PLL81] G. D. Prichett, A. L. Ludington, and J. F. Lapenta. The determination of all decadic Kaprekar constants. Fibonacci Quarterly, 19(1):45–52, February 1981. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/19-1/prichett.pdf.

Philippou:1982:WTC

[PM82] A. N. Philippou and A. A. Muwafi. Waiting for the K th consecutive success and the Fibonacci sequence of order K. Fibonacci Quarterly, 20(1):28–32, February 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
1/philippou.pdf.

Philippou:1985:LSR

[PM85] Andreas N. Philippou and Frosso S. Makri. Longest success runs and Fibonacci-type polynomials. Fibonacci Quarterly, 23(4):338– 345, November 1985. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/23-4/philippou.pdf.

Peterson:1977:BS

[PN77] James L. Peterson and Theodore A. Norman. Buddy systems. Communications of the Association for Computing Machinery, 20(6):421–431, June 1977. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Pond:1968:GFS

[Pon68] Jeremy C. Pond. Generalized Fibonacci summations. Fibonacci Quarterly, 6(2):97–108, April 1968. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/6-2/pond.
pdf.

Pongsriiam:2017:FLN

[Pon17a] Prapanpong Pongsriiam. Fibonacci and Lucas numbers which are one away from their products. Fibonacci Quarterly, 55(1): 29–??, February 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/55-1/pongsriiam.pdf.

Pongsriiam:2017:IVG

[Pon17b] Prapanpong Pongsriiam. Integral values of the generating functions of Fibonacci and Lucas numbers. College Mathematics Journal, 48(2):97–101, March 2017. CODEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic). URL http:

//www.tandfonline.com/doi/abs/10.4169/college.math.j.
48.2.97.

Pongsriiam:2019:FLN

[Pon19] Prapanpong Pongsriiam. Fibonacci and Lucas numbers which have exactly three prime factors and some unique properties of F_{18} and L_{18}. Fibonacci Quarterly, 57(5):130–??, December 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-5/pongsriiam.pdf.

Poonen:1988:PCS

[Poo88] Bjorn Poonen. Periodicity of a combinatorial sequence. Fibonacci Quarterly, 26(1):70–76, February 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
1/poonen.pdf.

Popov:1977:GFP

[Pop77] Blagoj S. Popov. Generating functions for powers of certain second-order recurrence sequences. Fibonacci Quarterly, 15(3): 221–223, October 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-3/popov.pdf.

Popov:1984:CSR

[Pop84] Blagoj S. Popov. On certain series of reciprocals of Fibonacci numbers. Fibonacci Quarterly, 22(3):261–265, August 1984. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-3/popov.pdf.

Popov:1985:NSF

[Pop85] Blagoj S. Popov. A note on the sums of Fibonacci and Lucas polynomials. Fibonacci Quarterly, 23(3):238–239, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-3/popov.pdf.

Popov:1986:SRS

[Pop86] Blagoj S. Popov. Summation of reciprocal series of numerical functions of second order. Fibonacci Quarterly, 24(1):17–21, February 1986. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/24-1/popov.pdf.

Powell:1977:PSC

[Pow77] Barry Powell. Proof of a special case of Dirichlet’s theorem. Fibonacci Quarterly, 15(2):167–169, April 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-2/
powell.pdf.

Powell:1996:URP

[Pow96] Corey Powell. On the uniqueness of reduced phi-partitions. Fibonacci Quarterly, 34(3):194–199, June 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-3/
powell.pdf.

Pike:1979:AFS

[PP79] M. C. Pike and J. Pixner. Algorithm 2: Fibonacci search. The Computer Journal, 22(1):86, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_

01/tiff/86.tif. Reprinted from The Computer Bulletin, vol. 8, p. 147, 1965. See [Boo79, Ove67, PHJ67].

Packard:1994:OPS

[PP94] Robert W. Packard and Erik S. Packard. The order of a perfect k-shuffle. Fibonacci Quarterly, 32(2):136–144, May 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-2/packard.pdf.

Panholzer:2000:TPF

[PP00] Alois Panholzer and Helmut Prodinger. Two proofs of Filipponi’s formula for odd-subscripted Lucas numbers. Fibonacci Quarterly, 38(2):165–166, May 2000. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/38-2/panholzer.pdf.

Panda:2014:BLS

[PP14] G. K. Panda and A. K. Panda. Balancing-like sequences associated with integral standard deviations of consecutive natural numbers. Fibonacci Quarterly, 52(5):187–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/52-5/Panda.pdf.

Panda:2017:CBN

[PP17] A. K. Panda and G. K. Panda. Circular balancing numbers. Fibonacci Quarterly, 55(4):309–??, November 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/55-4/panda.pdf.

Panda:2019:TLN

[PP19] Gopal Krishna Panda and Sushree Sangeeta Pradhan. Triangularlike numbers that are triangular. Fibonacci Quarterly, 57(4):356– ??, November 2019. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/57-4/panda.pdf.

Patra:2020:PIS

[PP20] Asim Patra and G. K. Panda. Positive integer solutions of some Diophantine equations involving Lucas-balancing numbers. Fibonacci Quarterly, 58(1):3–??, February 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/58-
1/patra.pdf.

Patra:2021:EDP

[PPK21] Asim Patra, Gopal Krishna Panda, and Tammatada Khemaratchatakumthorn. Exact divisibility by powers of the balancing and Lucas-balancing numbers. Fibonacci Quarterly, 59 (1):57–??, February 2021. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/59-1/patra.pdf.

Prasad:2001:PNA

[PR01] V. Siva Rama Prasad and B. Srinivasa Rao. Pentagonal numbers in the associated Pell sequences and Diophantine equations x^{2}(3x−1)^{2} = 8y^{2}±4. Fibonacci Quarterly, 39(4):299–303, August 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/39-4/prasad.pdf.

Prasad:2002:PNP

[PR02] V. Siva Rama Prasad and B. Srinivasa Rao. Pentagonal numbers in the Pell sequence and Diophantine equations 2x^{2} = y^{2}(3y − 1)^{2} ± 2. Fibonacci Quarterly, 40(3):233–241, June 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-3/prasad.pdf.

Panda:2013:GBN

[PR13] G. K. Panda and S. S. Rout. Gap balancing numbers. Fibonacci Quarterly, 51(3):239–??, August 2013. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/51-3/
panda.pdf.

Prasad:1982:NFF

[Pra82] K. C. Prasad. A note on the Farey–Fibonacci sequence. Fibonacci Quarterly, 20(3):242–244, August 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
3/prasad.pdf.

Preziosi:1968:HDM

[Pre68] Donald A. Preziosi. Harmonic design in Minoan architecture. Fibonacci Quarterly, 6(6):370–384, December 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/6-6/preziosi-a.pdf.

Prielipp:1968:EPN

[Pri68] Robert W. Prielipp. Even perfect numbers and seven. Fibonacci Quarterly, 6(4):286–287, October 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-4/
prielipp-a.pdf.

Price:2005:PEC

[Pri05] Thomas E. Price. Products of elliptical chord lengths and the Fibonacci numbers. Fibonacci Quarterly, 43(2):149–156, May 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/43-2/paper43-2-10.pdf.

Price:2018:PWS

[Pri18] Tim Price. Proof without words: Sum of squares of consecutive Fibonacci numbers. College Mathematics Journal, 49(2): 121, March 2018. CODEN ???? ISSN 0746-8342 (print),

1931-1346 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/07468342.2018.1424425.

Prodinger:1981:NFP

[Pro81] Helmut Prodinger. On the number of Fibonacci partitions of a set. Fibonacci Quarterly, 19(5):463–465, December 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-5/prodinger.pdf.

Prodinger:1983:OFP

[Pro83] Helmut Prodinger. Ordered Fibonacci partitions. Canadian Mathematical Bulletin = Bulletin canadien de mathématiques, 26(??):312–316, ???? 1983. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Prodinger:1994:SIA

[Pro94] Helmut Prodinger. Some information about the binomial transform. Fibonacci Quarterly, 32(5):412–415, November 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-5/prodinger.pdf.

Prodinger:1995:GDF

[Pro95] Helmut Prodinger. Geometric distributions and forbidden subwords. Fibonacci Quarterly, 33(2):139–141, May 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-2/prodinger.pdf.

Prodinger:1996:ABG

[Pro96] Helmut Prodinger. The asymptotic behavior of the golden numbers. Fibonacci Quarterly, 34(3):224–225, June 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-3/prodinger.pdf.

Prodinger:1997:QCK

[Pro97] Helmut Prodinger. On a question of Cooper and Kennedy. Fibonacci Quarterly, 35(2):135–??, May 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-2/
prodinger.pdf.

Prodinger:2004:NPG

[Pro04] Helmut Prodinger. A note on a paper of G. H. Weiss and M. Dishon. Fibonacci Quarterly, 42(4):290–291, November 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-4/quartprodinger04_2004.pdf.

Prodinger:2008:SMV

[Pro09] Helmut Prodinger. On a sum of Melham and its variants. Fibonacci Quarterly, 46/47(3):207–215, August 2008/2009. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/46_47-3/prodinger.pdf.

Prodinger:2015:TFS

[Pro15] Helmut Prodinger. Two families of series for the generalized golden ratio. Fibonacci Quarterly, 53(1):74–??, February 2015. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/53-1/prodinger.pdf.

Prodinger:2016:ARS

[Pro16a] Helmut Prodinger. Additional results on some recent infinite sums. Fibonacci Quarterly, 54(4):344–??, November 2016. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/54-4/prodinger.pdf.

Prodinger:2016:DCR

[Pro16b] Helmut Prodinger. Determinants containing rising powers of Fibonacci numbers. Fibonacci Quarterly, 54(2):137–??, May 2016. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/54-2/prodinger.pdf.

Prodinger:2021:NDP

[Pro21] Helmut Prodinger. Nondecreasing Deutsch paths. Fibonacci Quarterly, 59(3):232–??, August 2021. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/59-3/
prodinger.pdf.

Prodinger:2022:MSC

[Pro22] Helmut Prodinger. A matrix with sums of Catalan numbers — LU-decomposition and determinant. Fibonacci Quarterly, 60(3): 226–??, August 2022. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/60-3/prodinger.pdf.

Plandowski:1996:PTC

[PRS96] Wojciech Plandowski, Wojciech Rytter, and Tomasz Szymacha. Parallel tree-contraction and Fibonacci numbers. Information Processing Letters, 59(5):267–271, September 9, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pruitt:1967:FLN

[Pru67] Robert Pruitt. Fibonacci and Lucas numbers in the sequence of golden numbers. Fibonacci Quarterly, 5(2):175–178, April 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-2/pruitt.pdf.

Peele:1993:CPI

[PRW93] Rhodes Peele, A. J. Radcliffe, and Herbert S. Wilf. Congruence problems involving Stirling numbers on the first kind. Fibonacci Quarterly, 31(1):73–80, February 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-
1/peele.pdf.

Pryor:1994:IUP

[Pry94] Daniel V. Pryor. Implementation and usage of a portable and reproducible parallel pseudorandom number generator. Technical report SRC-TR-94-116, Supercomputing Research Center: IDA, Lanham, MD, USA, March 15, 1994. 16 pp.

Priest:1976:CGC

[PS76] Dean B. Priest and Stephen W. Smith. Column generators for coefficients of Fibonacci and Fibonacci-related polynomials. Fibonacci Quarterly, 14(1):30–34, February 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-1/priest.pdf.

Pollin:1980:MAF

[PS80] Jack M. Pollin and I. J. Schoenberg. On the matrix approach to Fibonacci numbers and the Fibonacci pseudoprimes. Fibonacci Quarterly, 18(3):261–267, October 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-
3/pollin.pdf.

Page:1992:ABP

[PS92] Warren Page and K. R. S. Sastry. Area-bisecting polygonal paths. Fibonacci Quarterly, 30(3):263–273, August 1992. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/30-3/page.pdf.

Peele:2002:MPC

[PS02] Rhodes Peele and Pantelimon Stanica. Matrix powers of column-justified Pascal triangles and Fibonacci sequences. Fibonacci Quarterly, 40(2):146–152, May 2002. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/40-2/
peele.pdf.

Prodinger:2019:SST

[PS19] Helmut Prodinger and Sarah J. Selkirk. Sums of squares of Tetranacci numbers: A generating function approach. Fibonacci Quarterly, 57(4):313–??, November 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/57-
4/prodinger.pdf.

Praton:2021:AHT

[PS21] Iwan Praton and Nart Shalqini. Amicable Heron triangles. Fibonacci Quarterly, 59(4):362–??, November 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/59-4/praton.pdf.

Pomerance:1980:P

[PSW80] Carl Pomerance, J. L. Selfridge, and Samuel S. Wagstaff, Jr. The pseudoprimes to 25 · 10^{9}. Mathematics of Computation, 35 (151):1003–1026, July 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Prodinger:1982:FNG

[PT82] Helmut Prodinger and Robert F. Tichy. Fibonacci numbers of graphs. Fibonacci Quarterly, 20(1):16–20, February 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-1/prodinger.pdf.

Protasi:1989:NAO

[PT89] Marco Protasi and Maurizio Talamo. On the number of arithmetical operations for finding Fibonacci numbers. Theoretical Computer Science, 64(1):119–124, April 1989. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Panraksa:2017:SAP

[PT17] Chatchawan Panraksa and Aram Tangboonduangjit. On some arithmetic properties of a sequence related to the quotient of Fibonacci numbers. Fibonacci Quarterly, 55(1):21–??, February 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/55-1/panraksa.pdf.

Panraksa:2018:AVL

[PT18] Chatchawan Panraksa and Aram Tangboonduangjit. p-adic valuation of Lucas iteration sequences. Fibonacci Quarterly, 56(4):348–??, November 2018. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/56-4/
panraksa.pdf.

Panraksa:2013:EDP

[PTW13] Chatchawan Panraksa, Aram Tangboonduangjit, and Keng Wiboonton. Exact divisibility properties of some subsequences of Fibonacci numbers. Fibonacci Quarterly, 51(4):307–??, November 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/51-4/panraksa.pdf.

Puckette:1962:RPP

[Puc62] Stephen Puckette. Recent publications: Popular Lectures in Mathematics Series, Vol. 1 The Method of Mathematical Induction, by I. S. Sominskii; Popular Lectures in Mathematics Series, Vol. 2 Fibonacci Numbers, by N. N. Vorob’ev; Popular Lectures in Mathematics Series, Vol. 3 Some Applications of Mechanics to Mathematics, by V. A. Uspenskii; Popular Lectures in Mathematics Series, Vol. 4 Geometrical Constructions Using Compasses Only, by A. N. Kostovskii; Popular Lectures in Mathematics Series, Vol. 5 The Ruler in Geometrical Constructions, by A. S. Smogorzhevskii; Popular Lectures in Mathematics Series, Vol. 6 Inequalities, by P. P. Korovkin. American Mathematical Monthly, 69(10):1020, December 1962. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Puchta:2001:NDF

[Puc01] Jan-Christoph Puchta. The number of k-digit Fibonacci numbers. Fibonacci Quarterly, 39(4):334–335, August 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-4/puchta.pdf.

Puchta:2002:DT

[Puc02a] Jan-Christoph Puchta. On the distribution of totients. Fibonacci Quarterly, 40(1):68–70, February 2002. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/40-
1/puchta2.pdf.

Puchta:2002:RNN

[Puc02b] Jan-Christoph Puchta. Representation of numbers with negative digits and multiplication of small integers. Fibonacci Quarterly, 40(1):66–67, February 2002. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/40-1/puchta1.pdf.

Puly:1994:CJN

[Pul94] Kanakku Puly. Cross-jump numbers. Fibonacci Quarterly, 32(1): 17–??, February 1994. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/32-1/puly.pdf.

Puttaswamy:1976:NTW

[Put76] T. K. Puttaswamy. A note on a theorem of W. B. Ford. Fibonacci Quarterly, 14(1):74–??, February 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
1/puttaswamy.pdf.

Paradis:1998:AQI

[PVB98] Jaume Paradís, Pelegrí Viader, and Lluís Bibiloni. Approximation of quadratic irrationals and their Pierce expansions. Fibonacci Quarterly, 36(2):146–153, May 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-2/
paradis.pdf.

Peart:1993:TFS

[PW93] Paul Peart and Leon Woodson. Triple factorization of some Riordan matrices. Fibonacci Quarterly, 31(2):121–128, May 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-2/peart.pdf.

Percus:1995:TAM

[PW95] Ora E. Percus and Paula A. Whitlock. Theory and application of Marsaglia’s monkey test for pseudorandom number generators. ACM Transactions on Modeling and Computer Simulation, 5(2): 87–100, April 1995. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See [MZ93].

Puri:2001:DPU

[PW01] Yash Puri and Thomas Ward. A dynamical property unique to the Lucas sequence. Fibonacci Quarterly, 39(5):398–402, November 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/39-5/puri.pdf.

Prodinger:2015:ILF

[PW15] Helmut Prodinger and Stephan Wagner. On identities by Larcombe–Fennessey and Cassini. Fibonacci Quarterly, 53(3): 219–??, August 2015. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/53-3/prodinger.pdf.

Parks:2023:SKB

[PW23] Harold R. Parks and Dean C. Wills. Sums of k-Bonacci numbers. Fibonacci Quarterly, 61(2):129–??, May 2023. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/61-
2/parks.pdf.

Pinter:2012:APR

[PZ12] Ákos Pintér and Volker Ziegler. On arithmetic progressions in recurrences — a new characterization of the Fibonacci sequence. Journal of Number Theory, 132(8):1686–1706, August 2012. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X12000728.

Qiu:1990:NDA

[QM90] Ke Qiu and Henk Meijer. A note on diameter of acyclic directed hypercubes. Information Processing Letters, 36(1):51–52, October 1, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Queen:1993:P

[Que93] Clifford S. Queen. Palindromes. Fibonacci Quarterly, 31(3):216– 225, August 1993. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/31-3/queen.pdf.

Qian:1996:EFC

[QW96] Haifeng Qian and Jie Wu. Enhanced Fibonacci cubes. The Computer Journal, 39(4):331–345, ???? 1996. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www.oup.co.uk/jnls/list/comjnl/
hdb/Volume_39/Issue_04/390331.sgm.abs.html;   http://
www3.oup.co.uk/computer_journal/Volume_39/Issue_04/Vol39_
04.body.html#AbstractQian.

Raab:1963:GCB

[Raa63] Joseph A. Raab. A generalization of the connection between the Fibonacci sequence and Pascal’s triangle. Fibonacci Quarterly, 1(3):21–31, October 1963. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/1-3/raab.pdf.

Rabinowitz:1991:EPSa

[Rab91a] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 29(3):277–282, August 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/29-3/elementary29-3.pdf.

Rabinowitz:1991:EPSb

[Rab91b] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 29(4):371–376, November 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/29-4/elementary29-4.pdf.

Rabinowitz:1992:EPSa FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-1/elementary30-1.pdf.

Rabinowitz:1992:EPSb

[Rab92b] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 30(2):182–186, May 1992. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/30-2/
elementary30-2.pdf.

Rabinowitz:1992:EPSc

[Rab92c] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 30(3):275–281, August 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-3/elementary30-3.pdf.

Rabinowitz:1992:EPSd

[Rab92d] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 30(4):368–375, November 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-4/elementary30-4.pdf.

Rabinowitz:1993:EPSa

[Rab93a] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 31(1):82–89, February 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-1/elementary31-1.pdf.

Rabinowitz:1993:EPSb

[Rab93b] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 31(2):181–186, May 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-2/
elementary31-2.pdf.

Rabinowitz:1993:EPSc

[Rab93c] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 31(3):277–282, August 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-3/elementary31-3.pdf.

Rabinowitz:1993:EPSd FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-4/elementary31-4.pdf.

Rabinowitz:1994:EPSa

[Rab94a] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 32(1):85–90, February 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-1/elementary32-1.pdf.

Rabinowitz:1994:EPSb

[Rab94b] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 32(2):180–186, May 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-2/
elementary32-2.pdf.

Rabinowitz:1994:EPSc

[Rab94c] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 32(4):373–378, August 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-4/elementary32-4.pdf.

Rabinowitz:1994:EPSd

[Rab94d] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 32(5):467–472, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-5/elementary32-5.pdf.

Rabinowitz:1995:EPSa

[Rab95a] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 33(1):85–90, February 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-1/elementary33-1.pdf.

Rabinowitz:1995:EPSb

[Rab95b] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 33(2):181–186, May 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-2/
elementary33-2.pdf.

Rabinowitz:1995:EPSc

[Rab95c] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 33(4):371–377, August 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-4/elementary33-4.pdf.

Rabinowitz:1995:EPSd

[Rab95d] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 33(5):466–471, November 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-5/elementary33-5.pdf.

Rabinowitz:1996:AMT

[Rab96a] Stanley Rabinowitz. Algorithmic manipulation of third-order linear recurrences. Fibonacci Quarterly, 34(5):447–464, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/34-5/rabinowitz.pdf.

Rabinowitz:1996:EPSa

[Rab96b] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 34(1):81–88, February 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-1/elementary34-1.pdf.

Rabinowitz:1996:EPSb

[Rab96c] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 34(2):181–186, May 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-2/
elementary34-2.pdf.

Rabinowitz:1996:EPSc

[Rab96d] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 34(4):373–378, August 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-4/elementary34-4.pdf.

Rabinowitz:1996:EPSd

[Rab96e] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 34(5):468–472, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-5/elementary34-5.pdf.

Rabinowitz:1997:EPSa FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-1/elementary35-1.pdf.

Rabinowitz:1997:EPSb

[Rab97b] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 35(2):181–186, May 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-2/
elementary35-2.pdf.

Rabinowitz:1997:EPSc

[Rab97c] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 35(3):277–282, August 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-3/elementary35-3.pdf.

Rabinowitz:1997:EPSd

[Rab97d] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 35(4):371–376, November 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-4/elementary35-4.pdf.

Rabinowitz:1998:EPSa

[Rab98a] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 36(1):85–90, February 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-1/elementary36-1.pdf.

Rabinowitz:1998:EPSb

[Rab98b] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 36(2):181–186, May 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-2/
elementary36-2.pdf.

Rabinowitz:1998:EPSc

[Rab98c] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 36(4):373–378, August 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-4/elementary36-4.pdf.

Rabinowitz:1998:EPSd FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-5/elementary36-5.pdf.

Rabinowitz:1999:AMS

[Rab99a] Stanley Rabinowitz. Algorithmic manipulation of second-order linear recurrences. Fibonacci Quarterly, 37(2):162–176, May 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/37-2/rabinowitz2.pdf.

Rabinowitz:1999:ASR

[Rab99b] Stanley Rabinowitz. Algorithmic summation of reciprocals of products of Fibonacci numbers. Fibonacci Quarterly, 37(2):122– 127, May 1999. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/37-2/rabinowitz1.pdf.

Rabinowitz:1999:EPSa

[Rab99c] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 37(1):85–90, February 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-1/elementary37-1.pdf.

Rabinowitz:1999:EPSb

[Rab99d] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 37(2):180–184, May 1999. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/37-2/
elementary37-2.pdf.

Rabinowitz:1999:EPSc

[Rab99e] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 37(3):277–281, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-3/elementary37-3.pdf.

Rabinowitz:1999:EPSd

[Rab99f] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 37(4):371–376, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-4/elementary37-4.pdf.

Rabinowitz:2000:EPSa FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/38-1/elementary38-1.pdf.

Rabinowitz:2000:EPSb

[Rab00b] Stanley Rabinowitz. Elementary problems and solutions. Fibonacci Quarterly, 38(2):181–185, May 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-2/
elementary38-2.pdf.

Rachmilevitch:2019:AOB

[Rac19] Shiran Rachmilevitch. Alternating offers bargaining and the golden ratio. Fibonacci Quarterly, 57(4):299–??, November 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-4/rachmilevitch.pdf.

Raine:1948:PTF

[Rai48] Charles W. Raine. Pythagorean triangles from the Fibonacci series. Scripta Mathematica, 14(??):164–??, ???? 1948. ISSN 0036-9713.

Raine:1954:FET

[Rai54] Charles W. Raine. Fibonacci equiareal triangles. Scripta Mathematica, 20(??):96–98, ???? 1954. ISSN 0036-9713.

Ralston:1985:ECC

[Ral85] Blake Ralston. Elemental complete composite number generators. Fibonacci Quarterly, 23(2):149–150, May 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/23-2/ralston.pdf.

Ramirez:2014:CGF

[Ram14] José L. Ramírez. On convolved generalized Fibonacci and Lucas polynomials. Applied Mathematics and Computation, 229 (??):208–213, February 25, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0096300313013325.

Ramirez:2016:PRG

[Ram16] José L. Ramírez. The Pascal rhombus and the generalized grand Motzkin paths. Fibonacci Quarterly, 54(2):99–??, May 2016. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/54-2/ramirez.pdf.

Raney:1966:GFS

[Ran66] George N. Raney. Generalization of the Fibonacci sequence to n dimensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 18(??):332–349, ???? 1966. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ranum:1995:SAF

[Ran95] David L. Ranum. On some applications of Fibonacci numbers. American Mathematical Monthly, 102(7):640–645, August/ September 1995. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rao:1953:SPF

[Rao53] K. Subba Rao. Some properties of Fibonacci numbers. American Mathematical Monthly, 60(10):680–684, December 1953. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rao:1955:SSF

[Rao55] K. Subba Rao. Some summation formulae involving Fibonacci numbers. Scripta Mathematica, 21(??):214–217, ???? 1955. ISSN 0036-9713.

Rao:2002:HNL

[Rao02] B. Srinivasa Rao. Heptagonal numbers in the Lucas sequence and Diophantine equations x^{2}(5x − 3)^{2} = 20y^{2} ± 16. Fibonacci Quarterly, 40(4):319–322, August 2002. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/40-4/rao.
pdf.

Rao:2003:HNF

[Rao03] B. Srinivasa Rao. Heptagonal numbers in Fibonacci sequence and Diophantine equations 4x^{2} = 5y^{2}(5y − 3)^{2} ± 16. Fibonacci Quarterly, 41(5):414–420, November 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-
5/rao.pdf.

Rao:2005:HNA

[Rao05a] B. Srinivasa Rao. Heptagonal numbers in the associated Pell sequence and Diophantine equations x^{2}(5x − 3)^{2} = 8y^{2} ± 4. Fibonacci Quarterly, 43(4):302–306, November 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers/43-4/paper43-4-3.pdf.

Rao:2005:HNP

[Rao05b] B. Srinivasa Rao. Heptagonal numbers in the Pell sequence and Diophantine equations 2x^{2} = y^{2}(5y−3)^{2}±2. Fibonacci Quarterly, 43(3):194–201, August 2005. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Papers/43-3/paper43-3-
1.pdf.

Raphael:1970:SRT

[Rap70] Brother L. Raphael. Some results in trigonometry. Fibonacci Quarterly, 8(4):371–??, October 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-4/
raphael-a.pdf.

Raphael:1974:LRS

[Rap74] Brother L. Raphael. Linearly recursive sequences of integers. Fibonacci Quarterly, 12(1):11–37, February 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/12-1/raphael.pdf.

Ratliff:1993:MFI

[Rat93] Michael I. Ratliff. Mandelbrot’s functional iteration and continued fractions. Fibonacci Quarterly, 31(3):263–267, August 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-3/ratliff.pdf.

Rawsthorne:1987:CGS

[Raw87] Daniel A. Rawsthorne. Counting the “good” sequences. Fibonacci Quarterly, 25(2):161–162, May 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-2/
rawsthorne.pdf.

Rawsthorne:1989:HMN

[Raw89] Daniel A. Rawsthorne. How many 1’s are needed? Fibonacci Quarterly, 27(1):14–17, February 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-
1/rawsthorne.pdf.

Ray:2022:DENa

[RB22] Prasanta Kumar Ray and Kisan Bhoi. On the Diophantine equation N_{n} = x^{a} ± x^{b} + 1. Fibonacci Quarterly, 60(5):316–??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/60-5/ray.pdf.

Read:1970:FSS

[Rea70] B. A. Read. Fibonacci series in the solar system. Fibonacci Quarterly, 8(4):428–438, October 1970. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/8-4/read-
a.pdf.

Read:1980:NTR

[Rea80] Ronald C. Read. A note on tiling rectangles with dominoes. Fibonacci Quarterly, 18(1):24–27, February 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-1/read.pdf.

Rebman:1975:SC

[Reb75] Kenneth R. Rebman. The sequence: 151645121320 — in combinatorics. Fibonacci Quarterly, 13(1):51–55, February 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-1/rebman.pdf.

Recski:1975:GFN

[Rec75] Andras Recski. On the generalization of the Fibonacci numbers. Fibonacci Quarterly, 13(4):315–317, December 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-4/recski.pdf.

Redmond:1994:IPF

[Red94] Don Redmond. Infinite products and Fibonacci numbers. Fibonacci Quarterly, 32(3):234–239, June 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-3/
redmond.pdf.

Reichman:1954:SFI

[Rei54] Raphael I. Reichman. A summation formula involving Fibonacci numbers. Scripta Mathematica, 20(??):111–112, ???? 1954. ISSN 0036-9713.

Reingold:1979:NT

[Rei79] Edward M. Reingold. A note on 3-2 trees. Fibonacci Quarterly, 17(2):151–157, April 1979. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/17-2/reingold.pdf.

Reiter:1992:FFN

[Rei92] Clifford Reiter. Fast Fibonacci numbers. Mathematica Journal, 2(3):??, Summer 1992. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http://www.mathematica-

journal.com/issue/v2i3/article/reiter/index.html;   https:
//www.math.utah.edu/pub/tex/bib/fibquart.bib.

Reiter:1993:DDF

[Rei93a] Ashley Melia Reiter. Determining the dimension of fractals generated by Pascal’s triangle. Fibonacci Quarterly, 31(2):112–120, May 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/31-2/reiter.pdf.

Reiter:1993:FNR

[Rei93b] Clifford A. Reiter. Fibonacci numbers: Reduction formulas and short periods. Fibonacci Quarterly, 31(4):315–324, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/31-4/reiter.pdf.

Reiter:2004:VFD

[Rei04] Clifford A. Reiter. Views of Fibonacci dynamics. Computers and Graphics, 28(2):297–300, April 2004. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Remis:2011:RBF

[Rem11] Rob F. Remis. On the relation between FDTD and Fibonacci polynomials. Journal of Computational Physics, 230 (4):1382–1386, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911000611X.

Raphael:1972:IPP

[RF72] Brother L. Raphael and FSC. Introduction to Patton polygons. Fibonacci Quarterly, 10(4):423–428, October 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-4/raphael.pdf.

Rihane:2020:PTG

[RFLT20] Salah Eddine Rihane, Bernadette Faye, Florian Luca, and Alain Togbé. Powers of two in generalized Lucas sequences. Fibonacci Quarterly, 58(3):254–??, August 2020. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/58-3/
rihane.pdf.

Rajagopal:2014:FNF

[RG14] Surajit Rajagopal and Martin Griffiths. 98.01 On Fibonacci numbers that are factorials. The Mathematical Gazette, 98(541):104– 107, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ruggles:1963:PFS

[RH63] I. D. Ruggles and Verner E. Hoggatt, Jr. A primer on the Fibonacci sequence, Part III. Fibonacci Quarterly, 1(3):61–65, October 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/1-3/hoggatt2.pdf.

Rabung:1982:ABS

[RH82] John Rabung and Jim Hyland. Analysis of a betting system. Fibonacci Quarterly, 20(3):263–278, August 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-3/rabung.pdf.

Rihane:2018:DDT

[RHT18] Salah Eddine Rihane, Mohand Ouamar Hernane, and Alain Togbé. On the D(4)-Diophantine triples of Fibonacci numbers. Fibonacci Quarterly, 56(1):63–??, February 2018. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/56-1/rihane.pdf.

Ribble:2000:MHT

[Rib00] Margie Ribble. In memoriam — Herta Taussig Freitag. Fibonacci Quarterly, 38(5):394–??, November 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-
5/in-memoriam.pdf.

Ribenboim:2005:FFF

[Rib05] Paulo Ribenboim. FFF: (Favorite Fibonacci Flowers). Fibonacci Quarterly, 43(1):3–14, February 2005. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/43-1/
paper43-1-1.pdf.

Ricci:1993:SRZ

[Ric93] Paolo Emilio Ricci. Sum rules for zeros of polynomials and generalized Lucas polynomials. Journal of Mathematical Physics, 34 (10):4884–4891, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Richardson:2001:FM

[Ric01] Thomas M. Richardson. The Filbert matrix. Fibonacci Quarterly, 39(3):268–275, June 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-
3/richardson.pdf.

Richter:2013:NPT

[Ric13] Christian Richter. A note on perfect tilings of rectangles with rectangles. Fibonacci Quarterly, 51(4):348–??, November 2013. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/51-4/richter.pdf.

Riesel:1969:LCP

[Rie69] Hans Riesel. Lucasian criteria for the primality of N = ⟨·∈^{\} −∞. Mathematics of Computation, 23(108):869–875, October 1969. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rieger:1976:HME

[Rie76] G. J. Rieger. On the Harris modification of the Euclidean algorithm. Fibonacci Quarterly, 14(3):196, October 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-3/rieger-a.pdf.

Rieger:1993:SUF

[Rie93] G. J. Rieger. Sums of unit fractions having long continued fractions. Fibonacci Quarterly, 31(4):338–340, November 1993. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-4/rieger.pdf.

Riede:1998:AES

[Rie98] Harald Riede. Asymptotic estimation of a sum of digits. Fibonacci Quarterly, 36(1):72–75, February 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-
1/riede.pdf.

Rieger:1999:FNH

[Rie99a] Georg J. Rieger. Fibonacci numbers and harmonic quadruples. Fibonacci Quarterly, 37(3):252–253, August 1999. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-3/rieger.pdf.

Rieger:1999:GSN

[Rie99b] Georg Johann Rieger. The golden section and Newton approximation. Fibonacci Quarterly, 37(2):178–179, May 1999. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-2/rieger.pdf.

Risk:1982:TEL

[Ris82] William P. Risk. Thevenin equivalents of ladder networks. Fibonacci Quarterly, 20(3):245–247, August 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/20-3/risk.pdf.

Rispoli:2005:FPT

[Ris05] Fred J. Rispoli. Fibonacci polytopes and their applications. Fibonacci Quarterly, 43(3):227–233, August 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers/43-3/paper43-3-5.pdf.

Rittaud:2023:FLS

[Rit23] Benoît Rittaud. Fibonacci-like sequences for variants of the Tower of Hanoi, with corresponding graphs and Gray codes. Fibonacci Quarterly, 61(3):240–??, August 2023. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/61-3/rittaud.pdf.

Ravenscroft:1989:SSG

[RL89] R. A. Ravenscroft, Jr. and E. A. Lamagna. Symbolic summation with generating functions. In Gonnet [Gon89], pages 228–233. ISBN 0-89791-325-6. LCCN QA76.95.I59 1989. US$29.00. ACM order number: 505890. English and French.

Ruiz:2015:LPF

[RL15] Carlos Alexis Gómez Ruiz and Florian Luca. On the largest prime factor of the ratio of two generalized Fibonacci numbers. Journal of Number Theory, 152(??):182–203, July 2015. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022314X15000426.

Ruiz:2018:PFS

[RL18] Carlos Alexis Gómez Ruiz and Florian Luca. On prime factors of the sum of two k-Fibonacci numbers. International Journal of Number Theory (IJNT), 14(4):1171–1195, May 2018. ISSN 1793-0421 (print), 1793-7310 (electronic). URL https://www.

worldscientific.com/doi/10.1142/S1793042118500720.

Ridley:2004:CUG

[RM04] J. N. Ridley and M. E. Mays. Compositions of unions of graphs. Fibonacci Quarterly, 42(3):222–230, August 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-3/quartridley03_2004.pdf.

Romano:2011:NLR

[RM11] Paul K. Romano and Harry McLaughlin. On non-linear recursive sequences and Benford’s Law. Fibonacci Quarterly, 49(2):134– 138, May 2011. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/49-2/romano.pdf.

Randic:1996:HOF

[RMA96] Milan Randić, Daniel A. Morales, and Oswaldo Araujo. Higherorder Fibonacci numbers. Journal of Mathematical Chemistry, 20(1):79–94, March 1996. CODEN JMCHEG. ISSN 0259-9791 (print), 1572-8897 (electronic). URL http://link.springer.

com/article/10.1007/BF01165157.

Robinson:1963:FMM

[Rob63] D. W. Robinson. The Fibonacci matrix modulo m. Fibonacci Quarterly, 1(2):29–36, April 1963. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/1-2/robinson.
pdf.

Robinson:1976:RPL

[Rob76] Donald W. Robinson. The rank and period of a linear recurrent sequence over a ring. Fibonacci Quarterly, 14(3):210–213, October 1976. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/14-3/robinson.pdf.

Robbins:1978:FNW

[Rob78a] Neville Robbins. On Fibonacci numbers which are powers. Fibonacci Quarterly, 16(6):515–517, December 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/16-6/robbins.pdf.

Roberts:1978:OFP

[Rob78b] David W. Roberts. The origin of Fibonacci phyllotaxis — an analysis of Adler’s contact pressure model and Mitchison’s expanding apex model. Journal of Theoretical Biology, 74 (2):217–233, September 21, 1978. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519378900735.

Robbins:1980:CSE

[Rob80] Neville Robbins. A class of solutions of the equation σ(n) = 2n + t. Fibonacci Quarterly, 18(2):137–146, April 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-2/robbins.pdf.

Robbins:1981:FLN

[Rob81] Neville Robbins. Fibonacci and Lucas numbers of the forms w^{2} − 1, w^{3} ± 1. Fibonacci Quarterly, 19(4):369–372, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-4/robbins.pdf.

Robbins:1982:SID

[Rob82a] Neville Robbins. Some identities and divisibility properties of linear second-order recursion sequences. Fibonacci Quarterly, 20 (1):21–23, February 1982. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/20-1/robbins.pdf.

Robinson:1982:LE

[Rob82b] Elmer D. Robinson. Letter to the editor. Fibonacci Quarterly, 20 (4):343–??, November 1982. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/20-4/letter.pdf.

Robbins:1983:FLN

[Rob83a] Neville Robbins. On Fibonacci and Lucas numbers which are sums of precisely four squares. Fibonacci Quarterly, 21(1):3–5, February 1983. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/21-1/robbins.pdf.

Robbins:1983:FNF

[Rob83b] Neville Robbins. On Fibonacci numbers of the form P X^{2}, where P is prime. Fibonacci Quarterly, 21(4):266–271, November 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-4/robbins.pdf.

Robbins:1983:FNW

[Rob83c] Neville Robbins. On Fibonacci numbers which are powers: II. Fibonacci Quarterly, 21(3):215–218, August 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/21-3/robbins.pdf.

Robbins:1984:PNF

[Rob84] Neville Robbins. On Pell numbers of the form P X^{2}, where P is prime. Fibonacci Quarterly, 22(4):340–348, November 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-4/robbins.pdf.

Robbins:1987:RBS

[Rob87] Neville Robbins. Representing binom(2n, n) as a sum of squares. Fibonacci Quarterly, 25(1):29–33, February 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-1/robbins.pdf.

Robbins:1990:FNF

[Rob90] Neville Robbins. Fibonacci numbers of the form CX^{2}, where 1 ≤ C ≤ 1000. Fibonacci Quarterly, 28(4):306–315, November 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/28-4/robbins.pdf.

Robbins:1991:NFL

[Rob91a] Neville Robbins. A new formula for Lucas numbers. Fibonacci Quarterly, 29(4):362–363, November 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-
4/robbins.pdf.

Robbins:1991:SCT

[Rob91b] Neville Robbins. Some convolution-type and combinatorial identities pertaining to binary linear recurrences. Fibonacci Quarterly, 29(3):249–255, August 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-
3/robbins.pdf.

Robbins:1994:FNP

[Rob94] Neville Robbins. On Fibonacci numbers and primes of the form 4k + 1. Fibonacci Quarterly, 32(1):15–16, February 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-1/robbins.pdf.

Robbins:1995:NRC

[Rob95] Neville Robbins. A note regarding continued fractions. Fibonacci Quarterly, 33(4):311–312, August 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-
4/robbins.pdf.

Robbins:1996:FP

[Rob96] Neville Robbins. Fibonacci partitions. Fibonacci Quarterly, 34 (4):306–313, August 1996. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/34-4/robbins.pdf.

Robbins:1998:WTE

[Rob98] Neville Robbins. Wilson’s theorem via Eulerian numbers. Fibonacci Quarterly, 36(4):317–318, August 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-4/robbins.pdf.

Robertson:1999:HMR

[Rob99] John S. Robertson. How many recursive calls does a recursive function make? SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 31(2):60–61, June 1999. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Robbins:2000:CP

[Rob00] Neville Robbins. On t-core partitions. Fibonacci Quarterly, 38 (1):39–48, February 2000. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/38-1/robbins.pdf.

Robbins:2002:APT

[Rob02a] Neville Robbins. An alternate proof of a theorem of J. Ewell. Fibonacci Quarterly, 40(1):56–??, February 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-1/robbins1.pdf.

Robbins:2002:DPB

[Rob02b] Neville Robbins. A divisibility property of binary linear recurrences. Fibonacci Quarterly, 40(3):269–271, June 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-3/robbins.pdf.

Robbins:2002:NPE

[Rob02c] Neville Robbins. On the number of partitions into an even and odd number of parts. Fibonacci Quarterly, 40(1):57–58, February 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/40-1/robbins2.pdf.

Robbins:2002:SCG

[Rob02d] Neville Robbins. Some consequences of Gauss’ triangular number theorem. Fibonacci Quarterly, 40(4):365–368, August 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-4/robbins.pdf.

Robbins:2004:DSL

[Rob04a] Neville Robbins. On Dedekind sums and linear recurrences of order two. Fibonacci Quarterly, 42(3):274–276, August 2004. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/42-3/quartrobbins03_2004.pdf.

Robbins:2004:PPF

[Rob04b] Neville Robbins. On the parity of the partition function. Fibonacci Quarterly, 42(4):368–369, November 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers1/42-4/quartneville04_2004.pdf.

Robbins:2005:LTR

[Rob05a] Neville Robbins. The Lucas triangle revisited. Fibonacci Quarterly, 43(2):142–148, May 2005. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Papers/43-2/paper43-2-
9.pdf.

Robbins:2005:IPF

[Rob05b] Neville Robbins. On the infinitude of primes of the form 3k + 1. Fibonacci Quarterly, 43(1):29–30, February 2005. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/43-1/paper43-1-4.pdf.

Robbins:2006:STS

[Rob06a] Neville Robbins. On sums of three squares. Fibonacci Quarterly, 44(1):71–72, February 2006. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/44-1/robbins.pdf.

Robbins:2006:NPP

[Rob06b] Neville Robbins. On the number of primitive Pythagorean triangles with a given inradius. Fibonacci Quarterly, 44(4):368–369, November 2006. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/44-4/robbins.pdf.

Robbins:2007:RFS

[Rob07] Neville Robbins. A recursive formula for sums of squares. Fibonacci Quarterly, 45(3):230–232, August 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/45-3/robbins.pdf.

Robbins:2014:TNR

[Rob14] Neville Robbins. On Tribonacci numbers and 3-regular compositions. Fibonacci Quarterly, 52(1):16–??, February 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-1/robbins.pdf.

Rockett:1981:SIB

[Roc81] Andrew M. Rockett. Sums of the inverses of binomial coefficients. Fibonacci Quarterly, 19(5):433–436, December 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-5/rockett.pdf.

Rodseth:1994:NBS

[Röd94] Öystein J. Rödseth. A note on Brown and Shiue’s paper on a remark related to the Frobenius problem. Fibonacci Quarterly, 32(5):407–408, November 1994. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/32-5/rodseth.
pdf.

Rokach:1996:OCC

[Rok96] Arie Rokach. Optimal computation, by computer, of Fibonacci numbers. Fibonacci Quarterly, 34(5):436–439, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-5/rokach.pdf.

Rolfe:2003:SAN

[Rol03] Timothy J. Rolfe. Spreadsheet-aided numerical experimentation: analytic formula for Fibonacci numbers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 35(2): 117–119, June 2003. CODEN SIGSD3. ISSN 0097-8418 (print),

2331-3927 (electronic). URL ftp://ftp.math.utah.edu/pub/
mirrors/ftp.ira.uka.de/bibliography/Misc/DBLP/2003.bib.

Ronn:1991:LER

[Ron91] Stefan Ronn. On the logarithmic evaluation of recurrence relations. Information Processing Letters, 40(4):197–199, November 25, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roselle:1967:ECT

[Ros67] D. P. Roselle. Enumeration of certain triangular arrays. Fibonacci Quarterly, 5(3):235–246, October 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-3/
roselle.pdf.

Ross:1972:LNC

[Ros72] Beverly Ross. A Lucas number counting problem. Fibonacci Quarterly, 10(3):325–327, April 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-3/
ross.pdf.

Rosenberg:1979:PN

[Ros79] Arnold L. Rosenberg. Profile numbers. Fibonacci Quarterly, 17 (3):259–263, October 1979. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/17-3/rosenberg.pdf.

Rosenberger:1983:SDP

[Ros83] Gerhard Rosenberger. On some divisibility properties of Fibonacci and related numbers. Fibonacci Quarterly, 21(4):253– 258, November 1983. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/21-4/rosenberger.pdf.

Rose:1985:NFI

[Ros85] Nicholas J. Rose. A note on a Fibonacci identity. Fibonacci Quarterly, 23(2):151–152, May 1985. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/23-2/rose.pdf.

Rosen:1988:NPL

[Ros88] Michael I. Rosen. Notes: a proof of the Lucas–Lehmer test. American Mathematical Monthly, 95(9):855–856, November 1988. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rothstein:1973:MCS

[Rot73] Jerome Rothstein. A method for constructing singly even magic squares. Fibonacci Quarterly, 11(5):543–544, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-5/rothstein.pdf.

Routledge:2011:GFP

[Rou11] Norman Routledge. 95.49 Generalised Fibonacci properties. The Mathematical Gazette, 95(534):483–486, November 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Roura:2013:FBN

[Rou13] Salvador Roura. Fibonacci BSTs: a new balancing method for binary search trees. Theoretical Computer Science, 482(??):48– 59, April 22, 2013. CODEN TCSCDI. ISSN 0304-3975 (print),

1879-2294 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304397512010572.

Rumney:1971:RBS

[RP71] Max Rumney and E. J. F. Primrose. Relations between a sequence of Fibonacci type and the sequence of its partial sums. Fibonacci Quarterly, 9(3):296–298, May 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-3/
rumney.pdf.

Rayaguru:2018:RPC

[RP18] S. G. Rayaguru and G. K. Panda. Repdigits as products of consecutive balancing or Lucas-balancing numbers. Fibonacci Quarterly, 56(4):319–??, November 2018. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/56-
4/rayaguru.pdf.

Rayaguru:2019:RPB

[RP19] Sai Gopal Rayaguru and Gopal Krishna Panda. Repdigits as products of balancing and Lucas-balancing numbers with indices in arithmetic progression. Fibonacci Quarterly, 57(3):231–??, August 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/57-3/rayaguru.pdf.

Ramirez:2014:PGF

[RR14] José L. Ramírez and Gustavo N. Rubiano. Properties and generalizations of the Fibonacci word fractal: Exploring fractal curves. Mathematica Journal, 16(??):??, ???? 2014. CODEN ???? ISSN 1047-5974 (print), 1097-1610 (electronic). URL http:

//www.mathematica-journal.com/2014/02/properties-and-
generalizations-of-the-fibonacci-word-fractal/.

Ramirez:2014:GFW

[RRD14] José L. Ramírez, Gustavo N. Rubiano, and Rodrigo De Castro. A generalization of the Fibonacci word fractal and the Fibonacci snowflake. Theoretical Computer Science, 528(??):40–56, April 3, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0304397514000851.

Reuben:1998:ECP

[RS98a] A. J. Reuben and A. G. Shannon. Ellipses, cardioides, and Penrose tiles. Fibonacci Quarterly, 36(1):45–54, February 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-1/reuben.pdf.

Rivolo:1998:CDR

[RS98b] M. T. Rivolo and A. Simi. Il calcolo delle radici quadrate e cubiche in Italia da Fibonacci a Bombelli. (Italian) [The calculation of square and cube roots in Italy from Fibonacci to Bombelli]. Archive for History of Exact Sciences, 52 (2):161–193, February 1998. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0003-
9519&volume=52&issue=2&spage=161.

Rangarajan:2002:BPT

[RS02] R. Rangarajan and H. S. Sudheer. The Brahmagupta polynomials in two complex variables and their conjugates. Fibonacci Quarterly, 40(2):161–169, May 2002. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/40-2/
rangarajan.pdf.

Rubio-Sanchez:2008:IPE

[RS08] Manuel Rubio-Sánchez. An introduction to problem equivalence with combinatorics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(3):313, September 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of ITiCSE ’08.

Radziejewska:2011:APF

[RS11a] Mirosława Radziejewska and Tomasz Schoen. Additive properties of the Fibonacci sequence. Fibonacci Quarterly, 49(1):22–27, February 2011. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/49-1/radziejewska.pdf.

Russo:2011:BSF

[RS11b] Vincent Russo and Loren Schwiebert. Beatty sequences, Fibonacci numbers, and the golden ratio. Fibonacci Quarterly, 49 (2):151–154, May 2011. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/49-2/russo.pdf.

Rubio-Sanchez:2007:ERF

[RSHL07] Manuel Rubio-Sánchez and Isidoro Hernán-Losada. Exploring recursion with Fibonacci numbers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 39(3):359, September 2007. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). Proceedings of the 12th Annual SIGCSE Conference on Innovation and Technology in Computer Science Education (ITiCSE’07).

Raghavacharyulu:1972:SDE

[RT72] I. V. V. Raghavacharyulu and A. R. Tekumalla. Solution of the difference equations of generalized Lucas polynomials. Journal of Mathematical Physics, 13(3):321–324, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rossi:2002:WFS

[RT02] Corinna Rossi and Christopher A. Tout. Were the Fibonacci series and the Golden Section known in Ancient Egypt? Historia Mathematica, 29(2):101–113, May 2002. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0315086001923341.

Rippon:2004:EOP

[RT04] P. J. Rippon and H. Taylor. Even and odd periods in continued fractions of square roots. Fibonacci Quarterly, 42(2):170–180, May 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Papers1/42-2/quartrippon02_2004.pdf.

Rybolowicz:2018:GTG

[RT18] Bernard Rybolowicz and Agnieszka Tereszkiewicz. Generalized tricobsthal and generalized tribonacci polynomials. Applied Mathematics and Computation, 325(??):297–308, May 15, 2018. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300318300018.

Ruggles:1963:SFR

[Rug63] I. Dale Ruggles. Some Fibonacci results using Fibonacci-type sequences. Fibonacci Quarterly, 1(2):75–79, April 1963. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/1-2/ruggles.pdf.

Ruiz:2017:WSS

[Rui17] Claudio Pita Ruiz. Weighted sums of squares via generalized Eulerian polynomials. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/55-5/pitaruiz.pdf.

Russell:1980:LE

[Rus80] David L. Russell. Letter to the editor. Fibonacci Quarterly, 18 (1):82–??, February 1980. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/18-1/letter2.pdf.

Russell:1981:SSO

[Rus81] David L. Russell. Summation of second-order recurrence terms and their squares. Fibonacci Quarterly, 19(4):336–339, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-4/russell.pdf.

Russell:1982:NSP

[Rus82] David L. Russell. Notes on sums of products of generalized Fibonacci numbers. Fibonacci Quarterly, 20(2):114–117, May 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-2/russell.pdf.

Ruskey:2011:FMH

[Rus11] Frank Ruskey. Fibonacci meets Hofstadter. Fibonacci Quarterly, 49(3):227–230, August 2011. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/49-3/ruskey.pdf.

Rush:2012:DRG

[Rus12] David E. Rush. Degree n relatives of the golden ratio and resultants of the corresponding polynomials. Fibonacci Quarterly, 50 (4):313–??, November 2012. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/50-4/rush.pdf.

Rice:1981:BWM

[RW81] Bart F. Rice and Robert Ward. Binary words with minimal autocorrelation at offset one. Fibonacci Quarterly, 19(4):297–312, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/19-4/rice.pdf.

Roettger:2021:SPT

[RW21] E. L. Roettger and H. C. Williams. Some primality tests constructed from a cubic extension of the Lucas functions. Fibonacci Quarterly, 59(3):194–??, August 2021. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/59-3/
roettger.pdf.

Ryavec:1975:AST

[Rya75] C. Ryavec. An application of spectral theory to Fibonacci numbers. Fibonacci Quarterly, 13(4):307–308, December 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-4/ryavec.pdf.

Ryder:1996:EFN

[Ryd96] Jack Ryder. Exploring Fibonacci numbers mod M. College Mathematics Journal, 27(2):122–124, March 1996. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.1996.11973763.

Rytter:2006:SSG

[Ryt06] Wojciech Rytter. The structure of subword graphs and suffix trees of Fibonacci words. Theoretical Computer Science, 363(2): 211–223, October 28, 2006. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Rotkiewicz:1995:EP

[RZ95] A. Rotkiewicz and K. Ziemak. On even pseudoprimes. Fibonacci Quarterly, 33(2):123–125, May 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-2/
rotkiewicz.pdf.

Safran:1992:FN

[Saf92] Charles Safran. The Fibonacci numbers. Chance, 5(1–2):43–46, Winter–Spring 1992. CODEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Saidak:2017:PNS

[Sai17] Filip Saidak. The prime numbers without the Sieve of Eratosthenes. Fibonacci Quarterly, 55(4):352–??, November 2017. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-4/saidak.pdf.

Salle:1975:MVR

[Sal75] H. J. A. Sallé. A maximum value for the rank of apparition of integers in recursive sequences. Fibonacci Quarterly, 13(2): 159–160, April 1975. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/13-2/salle.pdf.

Salomaa:2012:SOW

[Sal12] Arto Salomaa. Subword occurrences, weighted automata and iterated morphisms, especially the Fibonacci morphism. Theoretical Computer Science, 432(1):85–93, May 11, 2012. CO-DEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0304397512000539.

Sander:1989:RTW

[San89] J. W. Sander. A remark on a theorem of Weinstein. Fibonacci Quarterly, 27(3):242–246, June 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-3/
sander.pdf.

Sander:1990:FPR

[San90a] J. W. Sander. On Fibonacci primitive roots. Fibonacci Quarterly, 28(1):79–80, February 1990. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/28-1/sander.pdf.

Sanders:1990:PGT

[San90b] Lee Knisley Sanders. A proof from graph theory for a Fibonacci identity. Fibonacci Quarterly, 28(1):48–55, February 1990. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/28-1/sanders.pdf.

Sanna:2016:AVL

[San16] Carlo Sanna. The p-adic valuation of Lucas sequences. Fibonacci Quarterly, 54(2):118–??, May 2016. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/54-
2/sanna.pdf.

Sanford:2023:RDS

[San23] Chance Sanford. A remark on Dedekind sums and palindromic continued fractions. Fibonacci Quarterly, 61(4):357–??, November 2023. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/61-4/sanford.pdf.

Sarkar:1990:IRF

[Sar90] Vivek Sarkar. Instruction reordering for fork-join parallelism. ACM SIGPLAN Notices, 25(6):322–336, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http://www.acm.org:

80/pubs/citations/proceedings/pldi/93542/p322-sarkar/.

Sayer:1976:SCS

[Say76] F. P. Sayer. The sums of certain series containing hyperbolic functions. Fibonacci Quarterly, 14(3):215–223, October 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-3/sayer.pdf.

Sayer:1979:RR

[Say79] F. P. Sayer. The recurrence relation (r + 1)f_{r+1} = xf′_{r} + (K − r + 1)x^{2}f_{r−1}. Fibonacci Quarterly, 17(3):228–238, October 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-3/sayer.pdf.

Sundheim:1993:FNT

[SB93] R. Sundheim and B. Busta. Fibonacci numbers tend to obey Benford’s Law: an extension of Wlodarski and Sentance. Working paper, St. Cloud State University, St. Cloud, MN, USA, ???? 1993.

Sarkar:1997:LSF

[SBM97] Jayanta Sarkar, Manas Banerjee, and Asok K. Mukherjee. Lucas sequences and Fibonacci triads of graphs in PMO calculation on the charge-transfer bands of a series of EDA complexes: Correlation with experimental and AM1 results. International Journal of Quantum Chemistry, 63(4):817–825, ???? 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract?ID=42632;   http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=42632&PLACEBO=IE.pdf.

Sburlati:2002:GFS

[Sbu02] Giovanni Sburlati. Generalized Fibonacci sequences and linear congruences. Fibonacci Quarterly, 40(5):446–452, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/40-5/sburlati.pdf.

Sun:1984:EPG

[SC84] Hugo S. Sun and M. E. Cohen. An easy proof of the Greenwood– Gleason evaluation of the Ramsey number R(3, 3, 3). Fibonacci Quarterly, 22(3):235–238, August 1984. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/22-3/sun.
pdf.

Spickerman:1997:GFS

[SC97] W. R. Spickerman and R. L. Creech. The (2, T) generalized Fibonacci sequences. Fibonacci Quarterly, 35(4):358–360, November 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/35-4/spickerman.pdf.

Sofo:1998:FRS

[SC98] A. Sofo and P. Cerone. On a Fibonacci related series. Fibonacci Quarterly, 36(3):211–215, June 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-3/
sofo.pdf.

Shattuck:2001:DRS

[SC01] Steven Shattuck and Curtis Cooper. Divergent RATS sequences. Fibonacci Quarterly, 39(2):101–106, May 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-2/shattuck.pdf.

Somer:2004:SDS

[SC04] Lawrence Somer and Walter Carlip. p-stability of degenerate second-order recurrences. Fibonacci Quarterly, 42(2):114–127, May 2004. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Papers1/42-2/quartcarlip02_2004.pdf.

Shiu:2005:DFN

[SC05] Wai Chee Shiu and Chuan I. Chu. Distribution of the Fibonacci numbers modulo 3^{k}. Fibonacci Quarterly, 43(1):22–28, February 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers/43-1/paper43-1-3.pdf.

Storer:2007:DPD

[SC07] James A. (James Andrew) Storer and Martin Cohn, editors. DCC 2007: Proceedings: Data Compression Conference: 27– 29 March 2007, Snowbird, Utah. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007. ISBN 0-7695-2791-4. ISSN 1068-0314. LCCN ????

URL http://ieeexplore.ieee.org/servlet/opac?punumber=

4148731. IEEE Computer Society Order Number P2791.

Shen:2010:BNC

[SC10a] Shouqiang Shen and Jianmiao Cen. On the bounds for the norms of r-circulant matrices with the Fibonacci and Lucas numbers. Applied Mathematics and Computation, 216(10):2891–2897, July 15, 2010. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Somer:2010:LP

[SC10b] Lawrence Somer and Curtis Cooper. Lucas (a_{1}, a_{2}, …, a_{k} = ±1) pseudoprimes. Fibonacci Quarterly, 48(2):98–113, May 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/48-2/somer.pdf.

Scarano:1999:SEF

[Sca99] V. Scarano. On the sizes of extended Fibonacci cubes. IEEE Transactions on Parallel and Distributed Systems, 10(7):764–??, July 1999. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-

2183 (electronic). URL http://dlib.computer.org/td/books/
td1999/pdf/l0764.pdf;   http://www.computer.org/tpds/
td1999/l0764abs.htm.

Schub:1950:MFC

[Sch50] P. Schub. A minor Fibonacci curiosity. Scripta Mathematica, 16 (3):214, ???? 1950. ISSN 0036-9713.

Schwenk:1970:TG

[Sch70] Allen J. Schwenk. Take-away games. Fibonacci Quarterly, 8(3): 225–234, April 1970. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/8-3/schwenk-a.pdf.

Schoen:1981:MCR

[Sch81] Robert Schoen. Means, circles, right triangles, and the Fibonacci ratio. Fibonacci Quarterly, 19(2):160–162, April 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-2/schoen.pdf.

Schoen:1982:FSS

[Sch82] Robert Schoen. The Fibonacci sequence in successive partitions of a golden triangle. Fibonacci Quarterly, 20(2):159–163, May 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/20-2/schoen.pdf.

Schoen:1984:HGA

[Sch84] Robert Schoen. Harmonic, geometric, and arithmetic means in generalized Fibonacci sequences. Fibonacci Quarterly, 22(4):354– 357, November 1984. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/22-4/schoen.pdf.

Schrage:1985:TDG

[Sch85] Georg Schrage. A two-dimensional generalization of Grundy’s game. Fibonacci Quarterly, 23(4):325–329, November 1985. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-4/schrage.pdf.

Schissel:1989:CTT

[Sch89] Eric Schissel. Characterizations of three types of completeness. Fibonacci Quarterly, 27(5):409–419, November 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/27-5/schissel.pdf.

Schroeder:1991:FCP

[Sch91] Manfred Schroeder. Fractals, Chaos, Power Laws. W. H. Freeman and Company, New York, NY, USA, 1991. ISBN 0-671-74217-5, 0-7167-2136-8, 0-7167-2357-3. xviii + 429 pp. LCCN QD921 .S3 1990.

Schmerling:1994:BRBb

[Sch94] Siegfried Schmerling. Book review: Applications of Fibonacci numbers — volume 5: G. E. Bergum, A. N. Philippou, A. F. Horadam, eds. (1993): Dordrecht, Boston, London: Kluwer Academic Publishers, xxxiv + 625 p., 35 Fig., 30 Tab., ISBN 0-7923-2491-9. Computational Statistics & Data Analysis, 17(5): 603, June 1994. CODEN CSDADW. ISSN 0167-9473 (print),

1872-7352 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0167947394901538.

Schoffl:1997:DPT

[Sch97] Gerd Schöffl. Ducci-processes of 4-tuples. Fibonacci Quarterly, 35 (3):269–276, August 1997. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/35-3/schoffl.pdf.

Schmerl:2000:RPS

[Sch00] James H. Schmerl. A remark on parity sequences. Fibonacci Quarterly, 38(3):264–271, June 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-3/
schmerl.pdf.

Shen:2011:DIC

[SCH11] Shou-Qiang Shen, Jian-Miao Cen, and Yong Hao. On the determinants and inverses of circulant matrices with Fibonacci and Lucas numbers. Applied Mathematics and Computation, 217(23):9790–9797, August 1, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300311006424.

Schinzel:2013:DKB

[Sch13a] A. Schinzel. On divisibility by (a^{k} − b^{k})/(a − b). Fibonacci Quarterly, 51(1):72–??, February 2013. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/51-1/
schinzel.pdf.

Schinzel:2013:PLN

[Sch13b] A. Schinzel. A property of Lehmer numbers. Fibonacci Quarterly, 51(2):119–??, May 2013. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/51-2/
schinzel.pdf.

Schumacher:2017:SCI

[Sch17] Raphael Schumacher. The self-counting identity. Fibonacci Quarterly, 55(2):157–??, May 2017. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/55-2/
schumacher.pdf.

Schumacher:2019:HSS

[Sch19] Raphael Schumacher. How to sum the squares of the Tetranacci numbers and the Fibonacci m-step numbers. Fibonacci Quarterly, 57(2):168–??, May 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/57-2/
schumacher.pdf.

Schumacher:2020:EFS

[Sch20] Raphael Schumacher. Explicit formulas for sums involving the squares of the first n Tribonacci numbers. Fibonacci Quarterly, 58(3):194–??, August 2020. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/58-3/
schumacher.pdf.

Schumacher:2022:SCFa

[Sch22] Raphael Schumacher. The self-counting flow. Fibonacci Quarterly, 60(5):324–??, December 2022. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/60-5/
schumacher.pdf.

Spickerman:1993:SSD

[SCJ93] W. R. Spickerman, R. L. Creech, and R. N. Joyner. On the structure of the set of the difference systems defining (3, F) generalized Fibonacci sequences. Fibonacci Quarterly, 31(4):333– 337, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/31-4/spickerman.pdf.

Spickerman:1995:FGF

[SCJ95] W. R. Spickerman, R. L. Creech, and R. N. Joyner. On the (3, F) generalizations of the Fibonacci sequence. Fibonacci Quarterly, 33(1):9–12, February 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-
1/spickerman.pdf.

Scott:1968:CEF

[Sco68a] Alan M. Scott. Continuous extensions of Fibonacci identities. Fibonacci Quarterly, 6(4):245–249, October 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-4/scott.pdf.

Scott:1968:FIL

[Sco68b] Allan Scott. Fibonacci illustration of L’Hôpital’s rule. Fibonacci Quarterly, 6(2):138–??, April 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-2/
scott1-a.pdf.

Scott:1968:SFS

[Sco68c] Allan Scott. Scott’s Fibonacci scrapbook. Fibonacci Quarterly, 6(2):176–??, April 1968. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/6-2/scott2-a.pdf.

Scott:2018:SFR

[Sco18] J. A. Scott. 102.04 on the series of Fibonacci reciprocals. The Mathematical Gazette, 101(553):100–102, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10204-on-the-series-of-fibonacci-reciprocals/
874B3D016DC21BAC874EE884E26D4D96.

Shi:2023:DSE

[SCY^{+}23] Zhengyuan Shi, Cheng Chen, Gangqiang Yang, Hailiang Xiong, Fudong Li, Honggang Hu, and Zhiguo Wan. Design space exploration of Galois and Fibonacci configuration based on Espresso stream cipher. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 16(3):43:1–43:??, September 2023. CO-DEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL

https://dl.acm.org/doi/10.1145/3567428.

Shtefan:2018:SCF

[SD18] Dmitriy Shtefan and Irina Dobrovolska. The sums of the consecutive Fibonacci numbers. Fibonacci Quarterly, 56(3):229–??, August 2018. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/56-3/shtefan.pdf.

Scott:1977:TS

[SDH77] April Scott, Tom Delaney, and V. E. Hoggatt, Jr. The Tribonacci sequence. Fibonacci Quarterly, 15(3):193–200, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/15-3/scott.pdf.

Sadek:2013:DPF

[SE13] Jawad Sadek and Russell Euler. Divisibility properties for Fibonacci and related numbers. The Mathematical Gazette, 97 (540):461–464, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sadek:2020:PFN

[SE20a] Jawad Sadek and Russell Euler. 104.03 On periods of Fibonacci numbers using modular arithmetic on the Binet formula. The Mathematical Gazette, 104(559):150–154, March 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.cambridge.org/core/journals/mathematical-
gazette/article/10403-on-periods-of-fibonacci-numbers-
using-modular-arithmetic-on-the-binet-formula/30FE1AC16B389F17D16C5CBB420D9448.

Sadek:2020:UFL

[SE20b] Jawad Sadek and Russell Euler. 104.26 Unusual Fibonacci, Lucas and Pell congruence relations. The Mathematical Gazette, 104(561):507–509, November 2020. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/10426-unusual-fibonacci-lucas-and-pell-congruence-
relations/E22427749D42B30385C1C8CCD8D8E902.

Sedaghat:2014:ZAS

[Sed14] H. Sedaghat. Zero-avoiding solutions of the Fibonacci recurrence modulo a prime. Fibonacci Quarterly, 52(1):39–??, February 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/52-1/sedaghat.pdf.

Seebold:1991:FMS

[See91] P. Seebold. Fibonacci morphisms and Sturmian words. Theoretical Computer Science, 88(2):365–384, October 07, 1991. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Seebold:1998:CSM

[Séé98] Patrice Séébold. On the conjugation of standard morphisms. Theoretical Computer Science, 195(1):91–109, March 20, 1998. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (elec-

tronic). URL http://www.elsevier.com/cas/tree/store/
tcs/sub/1998/195/1/2695.pdf.

Selleck:1983:PSC

[Sel83] John H. Selleck. Powers of T and Soddy circles. Fibonacci Quarterly, 21(4):250–252, November 1983. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/21-
4/selleck.pdf.

Selmer:1984:NSP

[Sel84] Ernst S. Selmer. A note on Somer’s paper on linear recurrences. Fibonacci Quarterly, 22(3):194–195, August 1984. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-3/selmer.pdf.

Sentance:1973:FAB

[Sen73] W. A. Sentance. A further analysis of Benford’s Law. Fibonacci Quarterly, 11(5):490–494, December 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
5/sentance.pdf.

Serkland:1974:GIP

[Ser74] Carl Serkland. Generating identities for Pell triples. Fibonacci Quarterly, 12(2):121–128, April 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-2/
serkland.pdf.

Sesiano:2014:LM

[Ses14] Jacques Sesiano. Liber Mahameleth. Sources and Studies in the History of Mathematics and Physical Sciences. Springer International Publishing AG, Cham, Switzerland, 2014. ISBN 3-319-03939-3 (hardcover). ???? pp. LCCN QA32 .S47 2014.

Silber:1976:AFR

[SG76] Robert Silber and Ralph Gellar. The algebra of Fibonacci representations. Fibonacci Quarterly, 14(4):289–326, November 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-4/silber1.pdf.

Stolarsky:1985:RAX

[SG85] Kenneth B. Stolarsky and Steven Greenbaum. A ratio associated with Φ(x) = n. Fibonacci Quarterly, 23(3):265–269, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-3/stolarsky.pdf.

Sinha:1986:FNP

[SGBS86] B. P. Sinha, S. Ghose, B. B. Bhattacharya, and P. K. Srimani. A further note on Pascal graphs. Fibonacci Quarterly, 24(3):251– 257, August 1986. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/24-3/sinha.pdf.

Suryanarayana:1970:TCO

[SH70] D. Suryanarayana and Peter Hagis, Jr. A theorem concerning odd perfect numbers. Fibonacci Quarterly, 8(4):337–346, October 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-4/suryanarayana-a.pdf.

Shannon:1971:GPT

[SH71a] A. G. Shannon and A. F. Horadam. A generalized Pythagorean theorem. Fibonacci Quarterly, 9(3):307–312, May 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-3/shannon2.pdf.

Shannon:1971:RGF

[SH71b] A. G. Shannon and A. F. Horadam. Reciprocals of generalized Fibonacci numbers. Fibonacci Quarterly, 9(3):299–306, May 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-3/shannon1-a.pdf.

Shannon:1972:SPT

[SH72] A. G. Shannon and A. F. Horadam. Some properties of thirdorder recurrence relations. Fibonacci Quarterly, 10(2):135–146, February 1972. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/10-2/shannon.pdf.

Shannon:1973:MNG

[SH73] A. G. Shannon and A. F. Horadam. Mathematical notes: Generalized Fibonacci number triples. American Mathematical Monthly, 80(2):187–190, February 1973. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Smith:1976:PPG

[SH76] Claudia Smith and Verner E. Hoggatt, Jr. Primitive periods of generalized Fibonacci sequences. Fibonacci Quarterly, 14(4):343– 347, November 1976. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/14-4/smith.pdf.

Spears:1977:PGR

[SH77] W. D. Spears and T. F. Higginbotham. On powers of the golden ratio. Fibonacci Quarterly, 15(3):207–208, October 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-3/spears.pdf.

Shannon:1979:SRR

[SH79a] A. G. Shannon and A. F. Horadam. Special recurrence relations associated with the sequence {w_{n}(a, b; p, q)}. Fibonacci Quarterly, 17(4):294–298, December 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
4/shannon1.pdf.

Smith:1979:GFC

[SH79b] Claudia Smith and Verner E. Hoggatt, Jr. Generating functions of central values in generalized Pascal triangles. Fibonacci Quarterly, 17(1):58–66, February 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-
1/smith.pdf.

Smith:1979:SMV

[SH79c] Claudia Smith and Verner E. Hoggatt, Jr. A study of the maximal values in Pascal’s quadrinomial triangle. Fibonacci Quarterly, 17 (3):264–268, October 1979. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/17-3/smith.pdf.

Silva:1980:GFN

[SH80a] Anne Silva and Verner E. Hoggatt, Jr. Generalized Fibonacci number. Fibonacci Quarterly, 18(4):290–299, December 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-4/silva.pdf.

Smith:1980:RRE

[SH80b] Claudia Smith and Verner E. Hoggatt, Jr. Roots of (H − L)/15 recurrence equations in generalized Pascal triangles. Fibonacci Quarterly, 18(1):36–42, February 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-
1/smith.pdf.

Shannon:1982:CPG

[SH82] A. G. Shannon and A. F. Horadam. Concerning a paper by L. G. Wilson. Fibonacci Quarterly, 20(1):38–40, February 1982.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-1/shannon.pdf.

Shannon:1988:GFC

[SH88] A. G. Shannon and A. F. Horadam. Generalized Fibonacci continued fractions. Fibonacci Quarterly, 26(3):219–223, August 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/26-3/shannon.pdf.

Shannon:1991:GSS

[SH91] A. G. Shannon and A. F. Horadam. Generalized staggered sums. Fibonacci Quarterly, 29(1):47–51, February 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-1/shannon.pdf.

Shannon:2002:RLT

[SH02] A. G. Shannon and A. F. Horadam. Reflections on the lambda triangle. Fibonacci Quarterly, 40(5):405–415, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-5/shannon1.pdf.

Shapiro:1960:MNS

[Sha60] George Shapiro. Mathematical notes: Some properties of the Fibonacci numbers. American Mathematical Monthly, 67(9):876, November 1960. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sharpe:1964:FPN

[Sha64a] B. B. Sharpe. A Fibonacci — prime number relation. Fibonacci Quarterly, 2(4):317–??, December 1964. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/2-4/
sharpe.pdf.

Sharpe:1964:VS

[Sha64b] B. B. Sharpe. The vanishing square. Fibonacci Quarterly, 2(3): 215–??, October 1964. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/2-3/sharpe.pdf.

Sharpe:1965:S

[Sha65] Benjamin Sharpe. On sums F_{x}^{2} ± F_{y}^{2}. Fibonacci Quarterly, 3(1): 63–??, February 1965. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/3-1/sharpe.pdf.

Shah:1968:FSM

[Sha68] A. P. Shah. Fibonacci sequence modulo m. Fibonacci Quarterly, 6(2):139–141, April 1968. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/6-2/shah.pdf.

Shanks:1972:FPR

[Sha72a] Daniel Shanks. Fibonacci primitive roots. Fibonacci Quarterly, 10(2):163–168, February 1972. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/10-2/shanks-a.
pdf.

Sharp:1972:FDP

[Sha72b] W. E. Sharp. Fibonacci drainage patterns. Fibonacci Quarterly, 10(6):643–646, December 1972. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/10-6/sharp-a.
pdf.

Shane:1973:FPF

[Sha73] Harold D. Shane. A Fibonacci probability function. Fibonacci Quarterly, 11(5):517–522, December 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
5/shane.pdf.

Shanks:1974:II

[Sha74a] Daniel Shanks. Incredible identities. Fibonacci Quarterly, 12(3): 271–??, October 1974. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/12-3/shanks-a.pdf.

Shannon:1974:EEP

[Sha74b] A. G. Shannon. Explicit expressions for powers of linear recursive sequences. Fibonacci Quarterly, 12(3):281–287, October 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-3/shannon2.pdf.

Shannon:1974:GHF

[Sha74c] A. G. Shannon. A generalization of the Hilton–Fern theorem on the expansion of Fibonacci and Lucas numbers. Fibonacci Quarterly, 12(3):237–240, October 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
3/shannon1.pdf.

Shannon:1974:MCA

[Sha74d] A. G. Shannon. A method of Carlitz applied to the k th power generating function for Fibonacci numbers. Fibonacci Quarterly, 12(3):293–298, October 1974. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/12-3/shannon3.
pdf.

Shannon:1974:SPF

[Sha74e] A. G. Shannon. Some properties of a fundamental recursive sequence of arbitrary order. Fibonacci Quarterly, 12(4):327–334, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/12-4/shannon1.pdf.

Shallit:1975:SPP

[Sha75a] Jeffrey Shallit. A simple proof that phi is irrational. Fibonacci Quarterly, 13(1):32–??, February 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-
1/shallit.pdf.

Shanks:1975:NHN

[Sha75b] Daniel Shanks. Non-hypotenuse numbers. Fibonacci Quarterly, 13(4):319–321, December 1975. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/13-4/shanks.
pdf.

Shapiro:1976:FNU

[Sha76] Louis Shapiro. Fibonacci numbers and upper triangular groups. Fibonacci Quarterly, 14(3):201–202, October 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-3/shapiro.pdf.

Shannon:1977:TNP

[Sha77] A. G. Shannon. Tribonacci numbers and Pascal’s pyramid. Fibonacci Quarterly, 15(3):268–??, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/15-3/shannon-a.pdf.

Shannon:1978:FLN

[Sha78a] A. G. Shannon. Fibonacci and Lucas numbers and the complexity of a graph. Fibonacci Quarterly, 16(1):1–3, February 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-1/shannon1.pdf.

Shannon:1978:MRS

[Sha78b] A. G. Shannon. On the multiplication of recursive sequences. Fibonacci Quarterly, 16(1):27–32, February 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/16-1/shannon2.pdf.

Shannon:1978:PDS

[Sha78c] A. G. Shannon. Pellian Diophantine sequences. Fibonacci Quarterly, 16(2):99–102, April 1978. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/16-2/shannon.
pdf.

Shannon:1979:GHP

[Sha79a] A. G. Shannon. A generalization of Hilton’s partition of Horadam’s sequences. Fibonacci Quarterly, 17(4):349–357, December 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/17-4/shannon4.pdf.

Shannon:1979:GFN

[Sha79b] A. G. Shannon. Generalized Fibonacci numbers as elements of ideals. Fibonacci Quarterly, 17(4):347–348, December 1979. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-4/shannon3.pdf.

Shannon:1979:RRG

[Sha79c] A. G. Shannon. A recurrence relation for generalized multinomial coefficients. Fibonacci Quarterly, 17(4):344–346, December 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-4/shannon2.pdf.

Shapiro:1979:CS

[Sha79d] Louis W. Shapiro. The cycle of six. Fibonacci Quarterly, 17 (3):253–258, October 1979. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/17-3/shapiro.pdf.

Shannon:1980:SLR

[Sha80] A. G. Shannon. Some lacunary recurrence relations. Fibonacci Quarterly, 18(1):73–79, February 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-
1/shannon.pdf.

Shallit:1982:EDS

[Sha82] J. O. Shallit. Explicit descriptions of some continued fractions. Fibonacci Quarterly, 20(1):77–80, February 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-1/shallit.pdf.

Shannon:1983:ISO

[Sha83] A. G. Shannon. Intersections of second-order linear recursive sequences. Fibonacci Quarterly, 21(1):6–12, February 1983. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/21-1/shannon.pdf.

Shallit:1984:SPP

[Sha84a] J. O. Shallit. Some predictable Pierce expansions. Fibonacci Quarterly, 22(4):332–335, November 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-
4/shallit1.pdf.

Shannon:1984:SAP

[Sha84b] A. G. Shannon. Some asymptotic properties of generalized Fibonacci numbers. Fibonacci Quarterly, 22(3):239–243, August 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/22-3/shannon1.pdf.

Shallit:1986:MTP

[Sha86] J. O. Shallit. Metric theory of Pierce expansions. Fibonacci Quarterly, 24(1):22–40, February 1986. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/24-1/shallit.
pdf.

Shannon:1987:FHA

[Sha87] A. G. Shannon. A. F. Horadam — ad multos annos. Fibonacci Quarterly, 25(2):100–104, May 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-2/
shannon.pdf.

Shalhub:1988:CAA

[Sha88a] Sāmī Shalhūb. The calculations and algebra of abū Kāmil Shujā ibn Aslam and his effects on the work of al-Karajī and on the work of Leonardo Fibonacci. In Deuxième Colloque Maghrebin sur l’Histoire des Mathématiques Arabes (Tunis, 1988), pages A23– A39. Maghreb, Tunis, 1988.

Shallit:1988:GAS

[Sha88b] J. Shallit. A generalization of automatic sequences. Theoretical Computer Science, 61(1):1–16, October 1988. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Shallit:1989:LE

[Sha89] Jeffrey Shallit. Letter to the editor. Fibonacci Quarterly, 27(2): 186–??, May 1989. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/27-2/letter.pdf.

Shalhub:1990:CAA

[Sha90] S. Shalhub. The calculations and algebra of abu Kamil Shuja-ibn Aslam and his effects on the work of al-Karaji and on the work of Leonardo Fibonacci. (Arabic). In ????, editor, Deuxième Colloque Maghrebin sur l’Histoire des Mathématiques Arabes (Tunis, 1990), pages A23–A39. ????, ????, 1990. ISBN ???? LCCN ????

Shallit:1993:RNN

[Sha93] Jeffrey Shallit. Rational numbers with non-terminating, nonperiodic modified Engel-type expansions. Fibonacci Quarterly, 31(1):37–39, February 1993. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/31-1/shallit.pdf.

Shallit:1994:PER

[Sha94] Jeffrey Shallit. Pierce expansions and rules for the determination of leap years. Fibonacci Quarterly, 32(5):416–423, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/32-5/shallit.pdf.

Shallit:2005:MPN

[Sha05] Jeffrey Shallit. The mathematics of Per Nørgård’s rhythmic infinity system. Fibonacci Quarterly, 43(3):262–268, August 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/43-3/paper43-3-9.pdf.

Shaw:2006:PNG

[Sha06] Douglas J. Shaw. The pure numbers generated by the Collatz sequence. Fibonacci Quarterly, 44(3):194–201, August 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/44-3/shaw.pdf.

Sha:2011:CSR

[Sha11] Min Sha. On the cycle structure of repeated exponentiation modulo a prime power. Fibonacci Quarterly, 49(4):340–??, November 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/49-4/sha.pdf.

Shapcott:2012:CCP

[Sha12] Caroline Shapcott. C-color compositions and palindromes. Fibonacci Quarterly, 50(4):297–??, November 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/50-4/shapcott.pdf.

Shattuck:2013:CPD

[Sha13] Mark Shattuck. Combinatorial proofs of determinant formulas for the Fibonacci and Lucas polynomials. Fibonacci Quarterly, 51(1): 63–??, February 2013. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/51-1/shattuck.pdf.

Shallit:2023:STC

[Sha23] Jeffrey Shallit. Some Tribonacci conjectures. Fibonacci Quarterly, 61(3):214–??, August 2023. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/61-3/
shallit.pdf.

Shannon:1984:ICS

[SHB84] A. G. Shannon, A. F. Horadam, and Gerald E. Bergum. Infinite classes of sequence-generated circles. Fibonacci Quarterly, 22(3): 247–251, August 1984. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/22-3/shannon2.pdf.

Shannon:1974:SCF

[SHC74] A. G. Shannon, A. F. Horadam, and S. N. Collings. Some congruences for Fibonacci numbers. Fibonacci Quarterly, 12(4):351– 354, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/12-4/shannon2.pdf.

Shenton:1968:PDM

[She68] L. R. Shenton. Periodicity and density of modified Fibonacci sequences. Fibonacci Quarterly, 6(2):109–116, April 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-2/shenton.pdf.

Shea:1969:NDN

[She69] Dale D. Shea. On the number of divisions needed in finding the greatest common divisor. Fibonacci Quarterly, 7(4):337–340, November 1969. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/7-4/shea.pdf.

Shenton:1972:LDE

[She72] L. R. Shenton. Linear difference equations and generalized continuants Part I: Algebraic developments. Fibonacci Quarterly, 10(6): 585–590, December 1972. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/10-6/shenton-a.pdf.

Shea:1973:NDN

[She73] Dale D. Shea. On the number of divisions needed in finding the greatest common divisor. Fibonacci Quarterly, 11(5):508– 510, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/11-5/shea.pdf.

Shephard:2011:MIB

[She11] G. C. Shephard. Multiplicative identities for binomial coefficients. Fibonacci Quarterly, 49(1):10–21, February 2011. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/49-1/shephard.pdf.

Shi:1995:CRS

[Shi95] Xiquan Shi. Concerning the recursive sequences A_{n+k} = ∑_{i=1}^{k} a_{i}A_{n+i−1}^{a_{i}}. Fibonacci Quarterly, 33(3):240–243, June 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-3/shi.pdf.

Shibukawa:2020:NIS

[Shi20] Genki Shibukawa. New identities for some symmetric polynomials, and a higher order analogue of the Fibonacci and Lucas numbers. Fibonacci Quarterly, 58(5):200–??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/58-5/shibukawa.pdf.

Shortt:1978:IPC

[Sho78] Joseph Shortt. An iterative program to calculate Fibonacci numbers in O(log n) arithmetic operations. Information Processing Letters, 7(6):299–303, October ??, 1978. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shonhiwa:1999:GEJ

[Sho99a] Temba Shonhiwa. A generalization of the Euler and Jordan totient functions. Fibonacci Quarterly, 37(1):67–76, February 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-1/shonhiwa.pdf.

Shonhiwa:1999:GBF

[Sho99b] Temba Shonhiwa. Generalized bracket function inverse pairs. Fibonacci Quarterly, 37(3):233–239, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-3/shonhiwa.pdf.

Shonhiwa:2006:CPR

[Sho06] Temba Shonhiwa. Compositions with pairwise relatively prime summands within a restricted setting. Fibonacci Quarterly, 44(4):316–323, November 2006. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/44-4/
shonhiwa.pdf.

Shudde:1972:GSS

[Shu72] Rex H. Shudde. Golden section search problem. Fibonacci Quarterly, 10(4):422–??, October 1972. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/10-4/shudde.
pdf.

Shutov:2020:SDZ

[Shu20] Anton Shutov. On the sum of digits of the Zeckendorf representations of two consecutive numbers. Fibonacci Quarterly, 58(3): 203–??, August 2020. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/58-3/shutov.pdf.

Santos:2003:FNP

[SI03] José Plínio O. Santos and Miloš Ivković. Fibonacci numbers and partitions. Fibonacci Quarterly, 41(3):263–278, June 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-3/santos.pdf.

Santos:2005:PGP

[SI05] José Plínio O. Santos and Miloš Ivković. Polynomial generalizations of the Pell sequences and the Fibonacci sequence. Fibonacci Quarterly, 43(4):328–338, November 2005. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/43-
4/ivkovic.pdf.

Sigler:2002:FLA

[Sig02] L. E. (Laurence E.) Sigler. Fibonacci’s Liber Abaci: a Translation into Modern English of Leonardo Pisano’s Book of Calculation. Sources and studies in the history of mathematics and physical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 0-387-95419-8. viii + 636 pp. LCCN QA32 .F4713 2002. US$99.00. This historicallyimportant book is the first English translation of the original Latin edition of 1202, on the 800th anniversary of the book that introduced to Europe the Hindu numerals 0 through 9, the word zero, the notion of an algorithm, and the subject of algebra.

Siler:1963:FS

[Sil63] Ken Siler. Fibonacci summations. Fibonacci Quarterly, 1(3):67– 70, October 1963. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/1-3/siler.pdf.

Silber:1976:FPW

[Sil76] Robert Silber. A Fibonacci property of Wythoff pairs. Fibonacci Quarterly, 14(4):380–384, November 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
4/silber2.pdf.

Silber:1977:CFR

[Sil77a] Robert Silber. On the N-canonical Fibonacci representations of order N. Fibonacci Quarterly, 15(1):57–66, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-1/silber1-a.pdf.

Silber:1977:WNF

[Sil77b] Robert Silber. Wythoff’s Nim and Fibonacci representations. Fibonacci Quarterly, 15(1):85–88, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/15-1/silber2.pdf.

Silverman:1991:SIS

[Sil91] Herb Silverman. Summing infinite series with sex. Fibonacci Quarterly, 29(3):275–276, August 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-
3/silverman.pdf.

Siler:1998:MCP

[Sil98] Joseph R. Siler. Mean crowds and Pythagorean triples. Fibonacci Quarterly, 36(4):323–326, August 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-
4/siler.pdf.

Sills:2011:CPF

[Sil11] Andrew V. Sills. Compositions, partitions, and Fibonacci numbers. Fibonacci Quarterly, 49(4):348–??, November 2011. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/49-4/sills.pdf.

Simalarides:1998:CBN

[Sim98] A. Simalarides. Congruences modp^{n} for the Bernoulli numbers. Fibonacci Quarterly, 36(3):276–281, June 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-3/simalarides.pdf.

Simpson:2018:BRF

[Sim18] Richard Simpson. Book review: Finding Fibonacci: the quest to rediscover the forgotten mathematical genius who changed the world, by Keith Devlin. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(2):136–137, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/17498430.2017.1396782.

Singmaster:1970:SCP

[Sin70] David Singmaster. Some counterexamples and problems on linear recurrence relations. Fibonacci Quarterly, 8(3):264–267, April 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-3/singmaster-a.pdf.

Singmaster:1973:NBC

[Sin73a] David Singmaster. Notes on binomial coefficients: IV — proof of a conjecture of Gould on the GCD’s of two triples of binomial coefficients. Fibonacci Quarterly, 11(3):282–283, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-3/singmaster.pdf.

Singmaster:1973:SCC

[Sin73b] David Singmaster. Some corrections to Carlson’s “Determination of Heronian Triangles”. Fibonacci Quarterly, 11(2):157–158, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/11-2/singmaster.pdf.

Singh:1974:SFF

[Sin74] Sahib Singh. Stufe of a finite field. Fibonacci Quarterly, 12(1): 81–82, February 1974. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/12-1/singh.pdf.

Singmaster:1975:RBC

[Sin75] David Singmaster. Repeated binomial coefficients and Fibonacci numbers. Fibonacci Quarterly, 13(4):295–298, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-4/singmaster.pdf.

Singh:1980:TCA

[Sin80] Sahib Singh. Thoro’s conjecture and allied divisibility property of Lucas numbers. Fibonacci Quarterly, 18(2):135–136, April 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-2/singh.pdf.

Singh:1985:CFN

[Sin85] Parmanand Singh. The so-called Fibonacci numbers in Ancient and Medieval India. Historia Mathematica, 12(3):229– 244, August 1985. CODEN HIMADS. ISSN 0315-0860 (print),

1090-249X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0315086085900217.

Singh:1989:DEG

[Sin89] Sahib Singh. A Diophantine equation with generalization. Fibonacci Quarterly, 27(4):333–334, August 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/27-4/singh.pdf.

Singh:1990:HBP

[Sin90] Sahib Singh. On a Hoggatt–Bergum paper with totient function approach for divisibility and congruence relations. Fibonacci Quarterly, 28(3):273–276, August 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-
3/singh.pdf.

Sirvent:1997:SAB

[Sir97] Victor F. Sirvent. A semigroup associated with the k-Bonacci numbers with dynamic interpretation. Fibonacci Quarterly, 35 (4):335–340, November 1997. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/35-4/sirvent.
pdf.

Sitgreaves:1970:SPS

[Sit70] Rosedith Sitgreaves. Some properties of Stirling numbers of the second kind. Fibonacci Quarterly, 8(2):172–181, March 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-2/sitgreaves.pdf.

Siurys:2011:TLS

[Šiu11] Jonas Šiurys. A Tribonacci-like sequence of composite numbers. Fibonacci Quarterly, 49(4):298–??, November 2011. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/49-4/siurys.pdf.

Spickerman:1984:BFR

[SJ84] W. R. Spickerman and R. N. Joyner. Binet’s formula for the recursive sequence of order K. Fibonacci Quarterly, 22(4):327– 331, November 1984. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/22-4/spickerman.pdf.

Spickerman:1992:FGF

[SJC92] W. R. Spickerman, R. N. Joyner, and R. L. Creech. On the (2, F) generalizations of the Fibonacci sequence. Fibonacci Quarterly, 30(4):310–314, November 1992. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/30-
4/spickerman.pdf.

Smolarski:1981:RAF

[SK81] Dennis C. Smolarski and Leonard F. Klosinski. Recognition algorithms for Fibonacci numbers. Fibonacci Quarterly, 19(1):57–60, February 1981. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/19-1/smolarski.pdf.

Somer:2012:PDM

[SK12] Lawrence Somer and Michal Křížek. Power digraphs modulo n are symmetric of order M if and only if M is square free. Fibonacci Quarterly, 50(3):196–??, August 2012. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-3/somer.pdf.

Somer:2013:ECD

[SK13a] Lawrence Somer and Michal Krízek. Easy criteria to determine if a prime divides certain second-order recurrences. Fibonacci Quarterly, 51(1):3–??, February 2013. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/51-1/somer.
pdf.

Somer:2013:FPU

[SK13b] Lawrence Somer and Michal Křížek. Fixed points and upper bounds for the rank of appearance in Lucas sequences. Fibonacci Quarterly, 51(4):291–??, November 2013. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/51-
4/somer.pdf.

Somer:2013:PLL

[SK13c] Lawrence Somer and Michal Krízek. Prime Lehmer and Lucas numbers with composite indices. Fibonacci Quarterly, 51(3):194– ??, August 2013. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/51-3/somer.pdf.

Somer:2015:IDS

[SK15a] Lawrence Somer and Michal Krízek. Identically distributed second-order linear recurrences modulo p. Fibonacci Quarterly, 53(4):290–??, November 2015. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/53-4/somer.
pdf.

Somer:2015:LS

[SK15b] Lawrence Somer and Michal Krízek. On Lehmer superpseudoprimes. Fibonacci Quarterly, 53(3):206–??, August 2015. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/53-3/somer.pdf.

Somer:2015:PLS

[SK15c] Lawrence Somer and Michal Křížek. On primes in Lucas sequences. Fibonacci Quarterly, 53(1):2–??, February 2015. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/53-1/somer.pdf.

Somer:2016:IDS

[SK16] Lawrence Somer and Michal Křížek. Identically distributed second-order linear recurrences modulo p, II. Fibonacci Quarterly, 54(3):217–??, August 2016. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/54-3/somer.
pdf.

Somer:2017:MWC

[SK17] Lawrence Somer and Michal Krízek. On moduli for which certain second-order linear recurrences contain a complete system of Residues modulo m. Fibonacci Quarterly, 55(3):209–??, August 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/55-3/somer.pdf.

Somer:2020:ICA

[SK20] Lawrence Somer and Michal Křížek. Iteration of certain arithmetical functions of particular Lucas sequences. Fibonacci Quarterly, 58(1):55–??, February 2020. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/58-1/somer.
pdf.

Somer:2021:SOL

[SK21] Lawrence Somer and Michal Křížek. Second-order linear recurrences having arbitrarily large defect modulo p. Fibonacci Quarterly, 59(2):108–??, May 2021. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/59-2/somer.
pdf.

Somer:2022:FLD

[SK22a] Lawrence Somer and Michal Křížek. Frobenius, Lucas, and Dickson pseudoprimes. Fibonacci Quarterly, 60(4):325–??, November 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/60-4/somer1.pdf.

Somer:2022:GTB

[SK22b] Lawrence Somer and Michal Křížek. Generalization of a theorem of Bruckman on Dickson pseudoprimes. Fibonacci Quarterly, 60 (4):357–??, November 2022. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/60-4/somer2.pdf.

Skees:1965:PPC

[Ske65] W. D. Skees. A permutative property of certain multiples of the natural numbers. Fibonacci Quarterly, 3(4):279–291, December 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/3-4/skees.pdf.

Slater:1977:FNC

[Sla77] Peter J. Slater. Fibonacci numbers in the count of spanning trees. Fibonacci Quarterly, 15(1):11–13, February 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-1/slater.pdf.

Slavutskii:2000:RPS

[Sla00] I. Slavutskii. A remark on the paper of A. Simalarides: “Congruences Mod p” for the Bernoulli Numbers. Fibonacci Quarterly, 38(4):339–341, August 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-
4/slavutskii.pdf.

Slanina:2016:GFP

[Sla16] Piotr Slanina. Generalizations of Fibonacci polynomials and free linear groups. Linear Multilinear Algebra, 64(2):187–??, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Slijepcevic:1998:NID

[Sli98] Siniša Slijepčević. A note on initial digits of recurrence sequences. Fibonacci Quarterly, 36(4):305–308, August 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-4/slijepcevic.pdf.

Sloane:2014:A

[Slo14] N. J. A. Sloane. 2178 and all that. Fibonacci Quarterly, 52(2): 99–??, May 2014. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/52-2/sloane.pdf.

Sloane:2015:JRP

[Slo15] Neil J. A. Sloane. The John Riordan Prize. Fibonacci Quarterly, 53(2):192–??, May 2015. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Announcements/Riordan6.pdf.

Shannon:1993:CGL

[SM93] A. G. Shannon and R. S. Melham. Carlitz generalizations of Lucas and Lehmer sequences. Fibonacci Quarterly, 31(2):105– 111, May 1993. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/31-2/shannon.pdf.

Storer:2008:DDC

[SM08] James A. (James Andrew) Storer and Michael W. Marcellin, editors. DCC 2008: 2008 Data Compression Conference: March 25–27, 2008, Snowbird, Utah: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. ISBN 0-7695-3121-0. ISSN 1068-0314. LCCN QA76.9.D33 D372 2008; QA76.9.D33 INTER-

NET. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=4483271.

Small:1977:MSA

[Sma77] Donald B. Small. A matrix sequence associated with a continued fraction expansion of a number. Fibonacci Quarterly, 15(2):123– 130, April 1977. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/15-2/small.pdf.

Smith:1968:CGA

[Smi68] David A. Smith. Circularly generated Abelian groups. Fibonacci Quarterly, 6(1):36–45, February 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-1/
smith.pdf.

Smith:1978:E

[Smi78] Pamilla Graves Smith. Expansion. Fibonacci Quarterly, 16(2): 112–??, April 1978. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/16-2/smith.pdf.

Smith:1993:LPE

[Smi93] Peter Smith. LUC public-key encryption: a secure alternative to RSA. Dr. Dobb’s Journal of Software Tools, 18(1):44, 46, 48–49, 90–92, January 1993. CODEN DDJOEB. ISSN 1044-789X.

Smith:1994:CE

[Smi94] Peter Smith. Cryptography without exponentiation. Dr. Dobb’s Journal of Software Tools, 19(4):26, 28, 30, April 1994. CODEN DDJOEB. ISSN 1044-789X.

Smith:1996:CSN

[Smi96] Michael Smith. Cousins of Smith numbers: Monica and Suzanne sets. Fibonacci Quarterly, 34(2):102–104, May 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-2/smith.pdf.

Smith:2008:UFI

[Smi09] Neal O. Smith. On an ‘uncounted’ Fibonacci identity and its qanalogue. Fibonacci Quarterly, 46/47(1):73–78, February 2008/ 2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/46_47-1/smith.pdf.

Stakhov:1999:IFC

[SMS99] Alexei Stakhov, Vinancio Massingue, and Anna Sluchenkova. Introduction into Fibonacci coding and cryptography. Osnova, Kharkov State University, Kharkov, Russia, 1999. ISBN ????

???? pp. LCCN ???? URL http://www.goldenmuseum.com/
1502EMU_engl.html;   https://www.math.utah.edu/pub/tex/
bib/fibquart.bib.

Shum:2001:DEF

[SMSM01] A. Shum, P. M. Melliar-Smith, and L. E. Moser. Design and evaluation of the Fibonacci optical ATM switch. IEEE Transactions on Computers, 50(5):466–481, May 2001. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=926160.

Shannon:2002:CML

[SO02] A. G. Shannon and R. L. Ollerton. Combinatorial matrices and linear recursive sequences. Fibonacci Quarterly, 40(5):417–423, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/40-5/shannon2.pdf.

Stakhov:2009:MHE

[SO09] A. P. (Alekseí Petrovich) Stakhov and Scott Anthony Olsen. The mathematics of harmony: from Euclid to contemporary mathematics and computer science, volume 22 of K and E series on knots and everything. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2009. ISBN 981-277-582-X (hardcover). xlix + 694 pp. LCCN QA246.5 .S73 2009.

Shannon:2021:NLS

[SO21] A. G. Shannon and R. L. Ollerton. A note on Ledin’s summation problem. Fibonacci Quarterly, 59(1):47–??, February 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/59-1/shannon.pdf.

Sobolewski:2017:AVG

[Sob17] Bartosz Sobolewski. The 2-adic valuation of generalized Fibonacci sequences with an application to certain Diophantine equations. Journal of Number Theory, 180(??):730–742, November 2017. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022314X17302354.

Sofo:2012:FSP

[Sof12] A. Sofo. Finite sums in Pascal’s Triangle. Fibonacci Quarterly, 50 (4):337–??, November 2012. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/50-4/sofo.pdf.

Solomon:1976:DPC

[Sol76] Ronald Solomon. Divisibility properties of certain recurring sequences. Fibonacci Quarterly, 14(2):153–158, April 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-2/solomon.pdf.

Solak:2005:NCM

[Sol05] Süleyman Solak. On the norms of circulant matrices with the Fibonacci and Lucas numbers. Applied Mathematics and Computation, 160(1):125–132, January 5, 2005. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Solak:2007:ENC

[Sol07] Süleyman Solak. Erratum to “On the norms of circulant matrices with the Fibonacci and Lucas numbers” [Appl. Math. Comput. 160 (2005) 125–132]. Applied Mathematics and Computation, 190 (2):1855–1856, July 15, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [UNS07].

Somer:1972:FGN

[Som72] Lawrence F. Somer. The Fibonacci group and a new proof that F_{p−(5/p)} ≡ 0 (mod p). Fibonacci Quarterly, 10(4):345–348, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/10-4/somer-a.pdf.

Somer:1975:FRM

[Som75] Lawrence Somer. The Fibonacci ratios F_{k+1}/F_{k} modulo p. Fibonacci Quarterly, 13(4):322–324, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-4/somer.pdf.

Somer:1977:FLG

[Som77] Lawrence Somer. Fibonacci-like groups and periods of Fibonaccilike sequences. Fibonacci Quarterly, 15(1):35–40, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-1/somer.pdf.

Somer:1979:WSO

[Som79] Lawrence Somer. Which second-order linear integral recurrences have almost all primes as divisors? Fibonacci Quarterly, 17(2): 111–115, April 1979. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/17-2/somer.pdf.

Somer:1980:DPP

[Som80] Lawrence Somer. The divisibility properties of primary Lucas recurrences with respect to primes. Fibonacci Quarterly, 18(4): 316–333, December 1980. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/18-4/somer.pdf.

Somer:1981:RM

[Som81] Lawrence Somer. The residues of n^{n} modulo p. Fibonacci Quarterly, 19(2):110–116, April 1981. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/19-2/somer.
pdf.

Somer:1982:PPP

[Som82] Lawrence Somer. Possible periods of primary Fibonacci-like sequences with respect to a fixed odd prime. Fibonacci Quarterly, 20(4):311–333, November 1982. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/20-4/somer.
pdf.

Somer:1984:GHO

[Som84] Lawrence Somer. The generation of higher-order linear recurrences from second-order linear recurrences. Fibonacci Quarterly, 22(2):98–100, May 1984. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/22-2/somer.pdf.

Somer:1987:OR

[Som87] Lawrence Somer. On r th-order recurrences. Fibonacci Quarterly, 25(3):221–224, August 1987. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/25-3/somer.pdf.

Somer:1989:CRO

[Som89] Lawrence Somer. Congruence relations for k th-order linear recurrences. Fibonacci Quarterly, 27(1):25–30, February 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-1/somer.pdf.

Somer:1991:DRC

[Som91] Lawrence Somer. Distribution of residues of certain second-order linear recurrences modulo p — II. Fibonacci Quarterly, 29(1): 72–78, February 1991. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/29-1/somer.pdf.

Somer:2002:GTD

[Som02] Lawrence Somer. Generalization of a theorem of Drobot. Fibonacci Quarterly, 40(5):435–437, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-5/somer.pdf.

Somer:2006:LSW

[Som06a] Lawrence Somer. Lucas sequences U_{k} for which U_{2p} and U_{p} are pseudoprimes for almost all primes p. Fibonacci Quarterly, 44(1): 7–12, February 2006. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/44-1/somer.pdf.

Somer:2006:SOL

[Som06b] Lawrence Somer. Second-order linear recurrences of composite numbers. Fibonacci Quarterly, 44(4):358–361, November 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/44-4/somer.pdf.

Somer:2007:SMO

[Som07] Lawrence Somer. Special multipliers of k th-order linear recurrences modulo p^{r}. Fibonacci Quarterly, 45(1):10–21, February 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/45-1/somer.pdf.

Somer:2008:LPS

[Som09] Lawrence Somer. Lucas pseudoprimes of special types. Fibonacci Quarterly, 46/47(3):198–206, August 2008/2009. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/46_47-3/somer.pdf.

Smith:1977:RRD

[SP77] Stephen W. Smith and Dean B. Priest. Row and rising diagonal sums for a type of Pascal triangle. Fibonacci Quarterly, 15(4): 359–360, December 1977. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/15-4/smith.pdf.

Sanchez-Peregrino:2002:CFP

[SP02] Roberto Sánchez-Peregrino. Closed formula for poly-Bernoulli numbers. Fibonacci Quarterly, 40(4):362–364, August 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-4/sanchez.pdf.

Sahukar:2018:AFB

[SP18] Manasi Kumari Sahukar and G. K. Panda. Arithmetic functions of balancing numbers. Fibonacci Quarterly, 56(3):246–??, August 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/56-3/sahukar.pdf.

Sahukar:2019:REF

[SP19] M. K. Sahukar and G. K. Panda. Repdigits in Euler functions of Pell numbers. Fibonacci Quarterly, 57(2):134–??, May 2019. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-2/sahukar.pdf.

Spencer:1977:EPK

[Spe77] J. Spencer. An elementary proof of Kronecker’s theorem. Fibonacci Quarterly, 15(1):9–10, February 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-1/
spencer.pdf.

Spickerman:1970:NFF

[Spi70] W. R. Spickerman. A note on Fibonacci functions. Fibonacci Quarterly, 8(4):397–401, October 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-4/
spickerman.pdf.

Spickerman:1982:BFT

[Spi82] W. R. Spickerman. Binets’s formula for the Tribonacci sequence. Fibonacci Quarterly, 20(2):118–120, May 1982. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/20-2/spickerman.pdf.

Spilker:1997:IVH

[Spi97] Jürgen Spilker. Initial values for homogeneous linear recurrences of second order. Fibonacci Quarterly, 35(1):24–27, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-1/spilker.pdf.

Spiesser:2000:PLB

[Spi00a] Maryvonne Spiesser. Problèmes linéaires dans le Compendy de la praticque des nombres de Barthélemy de Romans et Mathieu Préhoude (1471): une approche nouvelle basée sur des sources proches du Liber abbaci de Léonard de Pise. (French) [Linear problems in the Compendium of the number practice by Barthélemy de Romans and Mathieu Préhoude (1471): an new approach based on sources near Liber abbaci of Leonardo of Pisa]. Historia Mathematica, 27(4):362–383, 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Spiesser:2000:PLC

[Spi00b] Maryvonne Spiesser. Problèmes linéaires dans le Compendy de la praticque des nombres de Barthélemy de Romans et Mathieu Préhoude (1471): une approche nouvelle basée sur des sources proches du Liber abbaci de Léonard de Pise. (French) [Linear problems in the Compendium of the number practice by Barthélemy de Romans and Mathieu Préhoude (1471): an new approach based on sources near Liber abbaci of Leonardo of Pisa]. Historia Mathematica, 27(4):362–383, 2000. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Spivey:2006:FID

[Spi06] Michael Z. Spivey. Fibonacci identities via the determinant sum property. College Mathematics Journal, 37(4):286–289, September 2006. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.tandfonline.com/doi/abs/10.
1080/07468342.2006.11922196.

Sporn:2022:FLH

[Spo22a] Howard Sporn. Fibonacci–Lucas hyperbolas. The Mathematical Gazette, 106(566):242–246, July 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/fibonaccilucas-hyperbolas/0B51B37A75BC9B51020FDF631AA6169C.

Sporn:2022:FFC

[Spo22b] Howard Sporn. Fibonacci fraction circles. The Mathematical Gazette, 106(565):1–8, March 2022. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/fibonacci-fraction-circles/40C40BE93CB0477893321A2F62FCA713.

Squire:1981:FC

[Squ81] William Squire. Fibonacci cubature. Fibonacci Quarterly, 19(4): 313–??, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/19-4/squire.pdf.

Singh:1983:PSE

[SR83] S. N. Singh and B. K. Rai. Properties of some extended Bernoulli and Euler polynomials. Fibonacci Quarterly, 21(3):162–172, August 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/21-3/singh.pdf.

Seyfarth:2012:CMU

[SR12] Ulrich Seyfarth and Kedar S. Ranade. Cyclic mutually unbiased bases, Fibonacci polynomials and Wiedemann’s conjecture. Journal of Mathematical Physics, 53(6):062201, June 2012. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v53/i6/
p062201_s1.

Spreafico:2019:FFS

[SR19] Elen Viviani Pereira Spreafico and Mustapha Rachidi. Fibonacci fundamental system and generalized Cassini identity. Fibonacci Quarterly, 57(2):155–??, May 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/57-
2/spreafico.pdf.

Srinivasan:2020:MFN

[Sri20] Anitha Srinivasan. The Markoff–Fibonacci numbers. Fibonacci Quarterly, 58(5):222–??, December 2020. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/58-
5/srinivasan.pdf.

Slivka:1991:MSP

[SS91] John Slivka and Norman C. Severo. Measures of sets partitioning Borel’s simple normal numbers to base 2 in [0, 1]. Fibonacci Quarterly, 29(1):19–22, February 1991. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/29-1/slivka.
pdf.

Smith:1995:PCD

[SS95a] P. Smith and C. Skinner. A public-key cryptosystem and a digital signature system based on the Lucas function analogue to discrete logarithms. Lecture Notes in Computer Science, 917: 357–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1995:PKC

[SS95b] P. Smith and C. Skinner. A public-key cryptosystem and a digital signature system based on the Lucas function analogue to discrete logarithms. Lecture Notes in Computer Science, 917: 357–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sanchis:2001:FOE

[SS01] Gabriela R. Sanchis and Laura A. Sanchis. On the frequency of occurrence of α^{i} in the α-expansions of the positive integers. Fibonacci Quarterly, 39(2):123–137, May 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-2/
sanchis.pdf.

Stanica:2017:REP

[SS17] Gabriela N. Stanica and Pantelimon Stanica. Recurrences for entries of powers of matrices. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/55-5/stanica.pdf.

Smajlovic:2022:GTZa

[SŠŠ22a] Lejla Smajlović, Zenan Šabanac, and Lamija Šćeta. On the generalized Tribonacci zeta function. Fibonacci Quarterly, 60(5):344– ??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/60-5/smajlovic1.pdf.

Smajlovic:2022:HTZa

[SŠŠ22b] Lejla Smajlović, Zenan Šabanac, and Lamija Šćeta. On the Hurwitz-type zeta function associated to the Lucas sequence. Fibonacci Quarterly, 60(5):355–??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/60-5/smajlovic2.pdf.

Stanica:2024:ERW

[SSS24] Andreea M. Stănică, Gabriela N. Stănică, and Pantelimon Stănică. An evaluation of rationally weighted binomial sums via some differential and integral operators. Fibonacci Quarterly, 62 (1):45–??, February 2024. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/62-1/stanica.pdf.

Shanks:1973:OFP

[ST73] Daniel Shanks and Larry Taylor. An observation of Fibonacci primitive roots. Fibonacci Quarterly, 11(2):159–160, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-2/shanks.pdf.

StJohn:1984:APZ

[St.84] Peter H. St. John. On the asymptotic proportions of zeros and ones in Fibonacci sequences. Fibonacci Quarterly, 22(2):144–145, May 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/22-2/stjohn.pdf.

Schaake:1989:NCP

[ST89] A. G. (Albertus Georg) Schaake and J. C. Turner. A new chapter for Pythagorean triples. Dept. of Mathematics and Statistics, University of Waikato, Hamilton, NZ, 1989. ISBN 0-908830-03-3. 155 pp. LCCN QA460.P8 S33 1989.

Sandor:1990:CNT

[ST90] J. Sándor and L. Tóth. On certain number-theoretic inequalities. Fibonacci Quarterly, 28(3):255–258, August 1990. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/28-3/sandor.pdf.

Seibert:2006:SCP

[ST06] Jaroslav Seibert and Pavel Trojovský. On sums of certain products of Lucas numbers. Fibonacci Quarterly, 44(2):172–180, May 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/44-2/seibert.pdf.

Sahin:2018:CSD

[ST18] Murat Sahin and Elif Tan. Conditional (strong) divisibility sequences. Fibonacci Quarterly, 56(1):18–??, February 2018. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/56-1/sahin.pdf.

Stanley:1975:FL

[Sta75] Richard P. Stanley. The Fibonacci lattice. Fibonacci Quarterly, 13(3):215–232, October 1975. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/13-3/stanley.
pdf.

Stanley:1976:SRP

[Sta76] T. E. Stanley. Some remarks on the periodicity of the sequence of Fibonacci numbers. Fibonacci Quarterly, 14(1):52–53, February 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/14-1/stanley.pdf.

Stanley:1980:PPF

[Sta80a] T. E. Stanley. Powers of the period function for the sequence of Fibonacci numbers. Fibonacci Quarterly, 18(1):44–??, February 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/18-1/stanley1.pdf.

Stanley:1980:SRP

[Sta80b] T. E. Stanley. Some remarks on the periodicity of the sequence of Fibonacci numbers — II. Fibonacci Quarterly, 18(1):45–46, February 1980. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/18-1/stanley2.pdf.

Stankovic:1982:CPT

[Sta82] Miomir S. Stanković. On a convolution product for the transform which maps derivatives into differences. Fibonacci Quarterly, 20(4):334–342, November 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
4/stankovic.pdf.

Stam:1985:RPR

[Sta85] A. J. Stam. Regeneration points in random permutations. Fibonacci Quarterly, 23(1):49–56, February 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/23-1/stam.pdf.

Stachowiak:2000:FCN

[Sta00] Grzegorz Stachowiak. Fibonacci correction networks. Lecture Notes in Computer Science, 1851:535–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/service/series/0558/
bibs/1851/18510535.htm;   http://link.springer-ny.com/
link/service/series/0558/papers/1851/18510535.pdf.

Stanica:2003:GFW

[Sta03] Pantelimon Stanica. Generating functions, weighted and nonweighted sums for powers of second-order recurrence sequences. Fibonacci Quarterly, 41(4):321–333, August 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/41-4/stanica.pdf.

Stnic:2014:NCF

[Stă14] Pantelimon Stănică. Normic continued fractions in totally and tamely ramified extensions of local fields. Fibonacci Quarterly, 52 (5):193–??, December 2014. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers1/52-5/Stanica.pdf.

Stampfli:2017:BGC

[Sta17] Marx Stampfli. Bridged graphs, circuits and Fibonacci numbers. Applied Mathematics and Computation, 302(??):68–79, June 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300316307640.

Stern:1971:SFP

[Ste71] Frederick Stern. The sum of the first n positive numbers — geometrically. Fibonacci Quarterly, 9(5):526–??, December 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-5/stern.pdf.

Stein:1973:DPA

[Ste73a] S. K. Stein. The density of the product of arithmetic progression. Fibonacci Quarterly, 11(2):145–152, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-2/stein.pdf.

Stern:1973:ILC

[Ste73b] Frederick Stern. Intersections of lines connecting two parallel lines. Fibonacci Quarterly, 11(2):201–203, April 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-2/stern.pdf.

Stern:1976:IBN

[Ste76] Samuel T. Stern. On isomorphisms between the naturals and integers. Fibonacci Quarterly, 14(1):15–16, February 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-1/stern.pdf.

Steiner:1978:PNC

[Ste78a] Ray Steiner. On k th-power numerical centers. Fibonacci Quarterly, 16(5):470–471, October 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
5/steiner2.pdf.

Steiner:1978:TPL

[Ste78b] Ray Steiner. On N th powers in the Lucas and Fibonacci series. Fibonacci Quarterly, 16(5):451–458, October 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-5/steiner1.pdf.

Stern:1979:AS

[Ste79] Frederick Stern. Absorption sequences. Fibonacci Quarterly, 17 (3):275–280, October 1979. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/17-3/stern.pdf.

Steiner:1981:PO

[Ste81a] Ray Steiner. On the “QX + 1 Problem,” Q odd. Fibonacci Quarterly, 19(3):285–288, August 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
3/steiner.pdf.

Steiner:1981:POI

[Ste81b] Ray Steiner. On the “QX + 1 Problem,” Q odd — II. Fibonacci Quarterly, 19(4):293–296, October 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
4/steiner.pdf.

Sternheimer:1985:CIE

[Ste85] R. M. Sternheimer. A corollary to iterated exponentiation. Fibonacci Quarterly, 23(2):146–148, May 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-2/
sternheimer.pdf.

Sternheimer:1986:SRC

[Ste86a] R. M. Sternheimer. Some results concerning Pythagorean triplets. Fibonacci Quarterly, 24(2):107–128, May 1986. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-2/sternheimer.pdf.

Stevens:1986:BNK

[Ste86b] Harlan R. Stevens. Bernoulli numbers and Kummer’s criterion. Fibonacci Quarterly, 24(2):154–159, May 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/24-2/stevens.pdf.

Sternheimer:1988:RII

[Ste88] R. M. Sternheimer. On a result involving iterated exponentiation. Fibonacci Quarterly, 26(2):178–180, May 1988. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/26-2/sternheimer.pdf.

Stewart:1995:MRF

[Ste95] Ian Stewart. Mathematical recreations: Fibonacci forgeries. Scientific American, 272(5):102–105, May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/
v272/n5/pdf/scientificamerican0595-102.pdf.

Stevanovic:2001:NMI

[Ste01] Dragan Stevanović. In the number of maximal independent set of vertices in star-like ladders. Fibonacci Quarterly, 39(3):211–213, June 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/39-3/stevanovic.pdf.

Steiner:2005:JDG

[Ste05] Wolfgang Steiner. The joint distribution of greedy and lazy Fibonacci expansions. Fibonacci Quarterly, 43(1):60–69, February 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers/43-1/paper43-1-8.pdf.

Steuding:2011:WFN

[Ste11a] Jörn Steuding. What Fibonacci numbers have to do with congruent numbers? Fibonacci Quarterly, 49(4):330–??, November 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/49-4/steuding.pdf.

Stewart:2011:ML

[Ste11b] Ian Stewart. Mathematics of life. Basic Books, New York, NY, USA, 2011. ISBN 0-465-02238-3, 0-465-02440-8 (e-book). viii + 358 pp. LCCN QH323.5 .S742 2011. URL http://www.loc.gov/

catdir/enhancements/fy1115/2011925563-b.html;   http:
//www.loc.gov/catdir/enhancements/fy1115/2011925563-d.
html.

Stewart:2021:SSI

[Ste21] Seán M. Stewart. Some series involving products between the harmonic numbers and the Fibonacci numbers. Fibonacci Quarterly, 59(3):214–??, August 2021. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/59-3/stewart.pdf.

Stewart:2023:SIR

[Ste23] Seán M. Stewart. 107.01 A simple integral representation of the Fibonacci numbers. The Mathematical Gazette, 107(568): 120–123, March 2023. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10701-
a-simple-integral-representation-of-the-fibonacci-numbers/
ADCAC1613B42D6660E37932E748B1A28.

Stinchcombe:1998:RRP

[Sti98] Adam M. Stinchcombe. Recurrence relations for powers of recursion sequences. Fibonacci Quarterly, 36(5):443–447, November 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/36-5/stinchcombe.pdf.

Stinchcombe:2000:LE

[Sti00] Adam Stinchcombe. Letter to the editor. Fibonacci Quarterly, 38 (5):463–??, November 2000. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/38-5/stinchcombe.
pdf.

Sonrod:2022:SPFa

[STL22] Earth Sonrod, Kate Tanner, and Colin Leyner. Some properties of the Fibonacci–Pascal triangle. Fibonacci Quarterly, 60(5):372– ??, December 2022. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/60-5/sonrod.pdf.

Stocks:1965:CLP

[Sto65] Douglas R. Stocks, Jr. Concerning lattice paths and Fibonacci numbers. Fibonacci Quarterly, 3(2):143–145, April 1965. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/3-2/stocks.pdf.

Stocks:1967:RIL

[Sto67] Douglas R. Stocks, Jr. Relations involving lattice paths and certain sequences of integers. Fibonacci Quarterly, 5(1):81–86, February 1967. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/5-1/stocks.pdf.

Stolarsky:1970:IMG

[Sto70] Kenneth B. Stolarsky. Infinitely many generalizations of Abel’s partial summation identity. Fibonacci Quarterly, 8(4):375–379, October 1970. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/8-4/stolarsky-a.pdf.

Stone:1975:GIF

[Sto75] Richard R. Stone. General identities for Fibonacci and Lucas numbers with polynomial subscripts in several variables. Fibonacci Quarterly, 13(4):289–294, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-4/stone.pdf.

Stolarsky:1977:SGF

[Sto77] Kenneth B. Stolarsky. A set of generalized Fibonacci sequences such that each natural number belongs to exactly one. Fibonacci Quarterly, 15(3):224–??, October 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-
3/stolarsky.pdf.

Stockmeyer:2008:STU

[Sto09a] Paul K. Stockmeyer. A smooth tight upper bound for the Fibonacci representation function R(n). Fibonacci Quarterly, 46/47(2):103–106, May 2008/2009. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/46_47-
2/stockmeyer.pdf.

Stoll:2008:HMF

[Sto09b] Thomas Stoll. On Hofstadter’s married functions. Fibonacci Quarterly, 46/47(1):62–67, February 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/46_47-1/stoll.pdf.

Stockmeyer:2017:ESA

[Sto17] Paul K. Stockmeyer. An exploration of sequence a000975. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/55-
5/stockmeyer.pdf.

Straus:1973:GCD

[Str73] E. G. Straus. On the greatest common divisor of some binomial coefficients. Fibonacci Quarterly, 11(1):25–26, February 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-1/straus.pdf.

Stringall:1976:DRB

[Str76] Robert W. Stringall. A density relationship between ax + b and [x/c]. Fibonacci Quarterly, 14(1):64–66, February 1976. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-1/stringall.pdf.

Strazdins:1999:PFL

[Str99a] Indulis Strazdins. Partial Fibonacci and Lucas numbers. Fibonacci Quarterly, 37(3):240–247, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-3/strazdins.pdf.

Strazdins:1999:SFG

[Str99b] Indulis Strazdins. Sieve formulas for the generalized Fibonacci and Lucas numbers. Fibonacci Quarterly, 37(4):361–365, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/37-4/strazdins.pdf.

Styles:1966:ECC

[Sty66] C. C. Styles. On evaluating certain coefficients. Fibonacci Quarterly, 4(2):139–147, April 1966. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/4-2/styles.
pdf.

Subramanian:1990:SHP

[Sub90] P. R. Subramanian. Springs of the Hermite polynomials. Fibonacci Quarterly, 28(2):156–161, May 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-2/
subramanian.pdf.

Subramanian:1995:NZH

[Sub95] P. R. Subramanian. Nonzero zeros of the Hermite polynomials are irrational. Fibonacci Quarterly, 33(2):131–134, May 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-2/subramanian.pdf.

Subramaniam:1999:AST

[Sub99] K. B. Subramaniam. Almost square triangular numbers. Fibonacci Quarterly, 37(3):194–197, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-3/subramaniam.pdf.

Subasi:2007:RCK

[Sub07] Murat Subaşı. Reply to the Comment of E. Kahya on “An improvement on Fibonacci search method in optimization theory” by Murat Subaşı, Necmettin Yildirim, Bünyamin Yildiz, [An improvement on Fibonacci search method in optimization theory, Applied Mathematics and Computation 147 (2004) 893–901]. Applied Mathematics and Computation, 188(1):1059–1060, May 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0096300306013464.

Subramaniam:2019:LBT

[Sub19] K. B. Subramaniam. 103.28 On a link between Triangular and Fibonacci numbers. The Mathematical Gazette, 103(558): 489, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/10328-
on-a-link-between-triangular-and-fibonacci-numbers/BAE56357F78DCA5BBE2CCBF593B6610A.

Sugiyama:1993:DCD

[Sug93] Kozo Sugiyama. Drawing compound digraphs and its application to an idea organizer (abstract). ACM SIGACT News, 24(1):58, Winter 1993. CODEN SIGNDM. ISSN 0163-5700 (print), 1943-5827 (electronic).

Sulanke:1989:RRD

[Sul89] Robert A. Sulanke. A recurrence restricted by a diagonal condition: Generalized Catalan arrays. Fibonacci Quarterly, 27(1): 33–46, February 1989. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/27-1/sulanke.pdf.

Sun:1973:GTP

[Sun73] Hugo S. Sun. A group-theoretical proof of a theorem in elementary number theory. Fibonacci Quarterly, 11(2):161–??, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/11-2/sun.pdf.

Sun:1975:ESR

[Sun75a] Hugo S. Sun. Embedding a semigroup in a ring. Fibonacci Quarterly, 13(1):50–??, February 1975. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/13-1/sun.pdf.

Sun:1975:SRR

[Sun75b] Hugo S. Sun. Structure of the reduced residue system with composite modulus. Fibonacci Quarterly, 13(4):329–330, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/13-4/sun.pdf.

Sun:1978:EGP

[Sun78] Hugo S. Sun. Embedding a group in the p th powers. Fibonacci Quarterly, 16(1):4–??, February 1978. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/16-1/sun.
pdf.

Sun:1979:GGS

[Sun79] H. S. Sun. On groups generated by the squares. Fibonacci Quarterly, 17(3):241–245, October 1979. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/17-3/sun.
pdf.

Sun:2001:ISU

[Sun01a] Zhi-Hong Sun. Invariant sequences under binomial transformation. Fibonacci Quarterly, 39(4):324–333, August 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-4/sun1.pdf.

Sun:2001:LRS

[Sun01b] Zhi-Hong Sun. Linear recursive sequences and the powers of matrices. Fibonacci Quarterly, 39(4):339–351, August 2001. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/39-4/sun2.pdf.

Sun:2002:FCP

[Sun02] Zhi-Hong Sun. Five congruences for primes. Fibonacci Quarterly, 40(4):345–351, August 2002. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/40-4/sun.pdf.

Sun:2005:NTS

[Sun05] Yidong Sun. Numerical triangles and several classical sequences. Fibonacci Quarterly, 43(4):359–370, November 2005. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/43-4/sun.pdf.

Sun:2006:CPC

[Sun06a] Zhi-Hong Sun. A criterion for polynomials to be congruent to the product of linear polynomials (mod p). Fibonacci Quarterly, 44(4):326–329, November 2006. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/44-
4/sun.pdf.

Sun:2006:EIC

[Sun06b] Zhi-Hong Sun. Expansions and identities concerning Lucas sequences. Fibonacci Quarterly, 44(2):145–153, May 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/44-2/sun2.pdf.

Sun:2006:PTN

[Sun06c] Zhi-Hong Sun. Primality tests for numbers of the form k · 2^{m} ± 1. Fibonacci Quarterly, 44(2):121–130, May 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/44-2/sun1.pdf.

Suryanarayan:1996:BP

[Sur96] E. R. Suryanarayan. The Brahmagupta polynomials. Fibonacci Quarterly, 34(1):30–39, February 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-
1/suryanarayan.pdf.

Suryanarayan:1998:BPT

[Sur98] E. R. Suryanarayan. The Brahmagupta polynomials in two complex variables. Fibonacci Quarterly, 36(1):34–42, February 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-1/suryanarayan.pdf.

Sury:2014:PPF

[Sur14] B. Sury. A polynomial parent to a Fibonacci–Lucas relation. American Mathematical Monthly, 121(3):236, March 2014. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/stable/pdfplus/10.4169/amer.
math.monthly.121.03.236.pdf.

Suttenfield:1978:NS

[Sut78] James M. Suttenfield, Jr. A new series. Fibonacci Quarterly, 16 (4):335–343, August 1978. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/16-4/suttenfield.
pdf.

Suto:1989:SCS

[Süt89] András Sütö. Singular continuous spectrum on a Cantor set of zero Lebesgue measure for the Fibonacci Hamiltonian. Journal of Statistical Physics, 56(3–4):525–531, August 1989. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/article/10.1007/BF01044450.

Spearman:1998:F

[SW98] Blair K. Spearman and Kenneth S. Williams. The factorization of x^{5} ± x^{3} + n. Fibonacci Quarterly, 36(2):158–170, May 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-2/spearman.pdf.

Sellers:2002:ICN

[SW02] James A. Sellers and Hugh Williams. On the infinitude of composite NSW numbers. Fibonacci Quarterly, 40(3):253–259, June 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/40-3/sellers.pdf.

Shattuck:2010:PSB

[SW10] Mark Shattuck and Tamás Waldhauser. Proofs of some binomial identities using the method of last squares. Fibonacci Quarterly, 48(4):290–297, November 2010. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/48-
4/shattuck.pdf.

Swamy:1966:MFI

[Swa66a] M. N. S. Swamy. More Fibonacci identities. Fibonacci Quarterly, 4(4):369–372, December 1966. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/4-4/swamy.pdf.

Swamy:1966:PPD

[Swa66b] M. N. S. Swamy. Properties of the polynomials defined by Morgan–Voyce. Fibonacci Quarterly, 4(1):73–81, February 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-1/swamy.pdf.

Swamy:1968:FPM

[Swa68] M. N. S. Swamy. Further properties of Morgan–Voyce polynomials. Fibonacci Quarterly, 6(2):167–175, April 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-2/swamy.pdf.

Swamy:1973:GFQ

[Swa73] M. N. S. Swamy. On generalized Fibonacci quaternions. Fibonacci Quarterly, 11(5):547–549, December 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-
5/swamy.pdf.

Swamy:1977:FGB

[Swa77] M. N. S. Swamy. A formula for ∑_{1}^{n} F_{k}(x)y^{n−k} and its generalization to r-bonacci polynomials. Fibonacci Quarterly, 15(1): 73–77, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/15-1/swamy.pdf.

Swain:1992:PWS

[Swa92] James J. Swain, editor. Proceedings of the Winter Simulation Conference. Crystal Gateway Marriott Hotel, Arlington, Virginia, December 13–16, 1992, volume 24. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-7803-0797-6 (softbound), 0-7803-0798-4 (casebound), 0-7803-0799-2 (microfiche). LCCN T57.62 .W787 1992. IEEE catalog number 92CH3202-9.

Swan:1994:AAa

[Swa94] Tom R. Swan. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 19(1):111–??, January 1994. CODEN DDJOEB. ISSN 1044-789X.

Swamy:1997:CGP

[Swa97a] M. N. S. Swamy. On a class of generalized polynomials. Fibonacci Quarterly, 35(4):329–334, November 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-
4/swamy.pdf.

Swamy:1997:CII

[Swa97b] M. N. S. Swamy. On certain identities involving Fibonacci and Lucas numbers. Fibonacci Quarterly, 35(3):230–231, August 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/35-3/swamy.pdf.

Swamy:1998:BTR

[Swa98] M. N. S. Swamy. Brahmagupta’s theorems and recurrence relations. Fibonacci Quarterly, 36(2):125–128, May 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-2/swamy.pdf.

Swamy:1999:GFL

[Swa99a] M. N. Swamy. Generalized Fibonacci and Lucas polynomials, and their associated diagonal polynomials. Fibonacci Quarterly, 37(3):213–222, August 1999. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/37-3/swamy.pdf.

Swamy:1999:GJP

[Swa99b] M. N. S. Swamy. A generalization of Jacobsthal polynomials. Fibonacci Quarterly, 37(2):141–144, May 1999. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/37-2/
swamy.pdf.

Swamy:1999:NPP

[Swa99c] M. N. S. Swamy. Network properties of a pair of generalized polynomials. Fibonacci Quarterly, 37(4):350–360, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-4/swamy.pdf.

Swamy:2000:GMM

[Swa00a] M. N. S. Swamy. Generalizations of modified Morgan–Voyce polynomials. Fibonacci Quarterly, 38(1):8–16, February 2000. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-1/swamy1.pdf.

Swamy:2000:RDP

[Swa00b] M. N. S. Swamy. Rising diagonal polynomials associated with Morgan–Voyce polynomials. Fibonacci Quarterly, 38(1):61–70, February 2000. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/38-1/swamy2.pdf.

Swamy:2000:SFP

[Swa00c] M. N. S. Swamy. Some further properties of André–Jeannin and their companion polynomials. Fibonacci Quarterly, 38(2):114– 122, May 2000. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/38-2/swamy.pdf.

Swensen:1964:AFN

[Swe64] Ben L. Swensen. Application of Fibonacci numbers to solutions of systems of linear equations. Fibonacci Quarterly, 2(4):314– 316, December 1964. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/2-4/swensen.pdf.

Swinton:2004:WDG

[Swi04] Jonathan Swinton. Watching the daisies grow: Turing and Fibonacci phyllotaxis. In Teuscher [Teu04], pages 477–498. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Shallit:1984:LRD

[SY84] J. O. Shallit and J. P. Yamron. On linear recurrences and divisibility by primes. Fibonacci Quarterly, 22(4):366–368, November 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/22-4/shallit2.pdf.

Sagan:1989:PAT

[SY89] Bruce E. Sagan and Yeong-Nan Yeh. Probabilistic algorithms for trees. Fibonacci Quarterly, 27(3):201–207, June 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-3/sagan.pdf.

Subasi:2004:IFS

[SYY04] Murat Subasi, Necmettin Yildirim, and Bünyamin Yildiz. An improvement on Fibonacci search method in optimization theory. Applied Mathematics and Computation, 147(3):893–901, January 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [Kah06].

Szalay:2002:REU

[Sza02] László Szalay. On the resolution of the equation U_{n} = binom(x, 3) and V_{n} = binom(x, 3). Fibonacci Quarterly, 40(1):9–12, February 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/40-1/szalay.pdf.

Szakacs:2017:KOL

[Sza17] Tamás Szakács. k-order linear recursive sequences and the golden ratio. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/55-5/szakacs.pdf.

Szikszai:2017:DPL

[Szi17] Márton Szikszai. Distinct products in Lucas sequences — on a problem of Kimberling. Fibonacci Quarterly, 55(4):291–??, November 2017. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/55-4/szikszai.pdf.

Tasci:2002:AHS

[TA02] Dursun Taşci and Ercan Altinişik. On the almost Hilbert–Smith matrices. Fibonacci Quarterly, 40(4):339–344, August 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-4/tasci.pdf.

Tadlock:1965:PO

[Tad65] Sheryl B. Tadlock. Products of odds. Fibonacci Quarterly, 3(1): 54–56, February 1965. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/3-1/tadlock.pdf.

Takahashi:2000:FAC

[Tak00] Daisuke Takahashi. A fast algorithm for computing large Fibonacci numbers. Information Processing Letters, 75(6):243– 246, November 30, 2000. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL http://www.elsevier.nl/

gej-ng/10/23/20/64/31/25/abstract.html;   http://www.
elsevier.nl/gej-ng/10/23/20/64/31/25/article.pdf.

Tanny:1975:ASI

[Tan75] S. M. Tanny. On alternating subsets of integers. Fibonacci Quarterly, 13(4):325–328, December 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-
4/tanny.pdf.

Tang:1998:SLE

[Tan98] I. Tang. Solving linear equations using an optimization-based iterative scheme. Fibonacci Quarterly, 36(3):248–251, June 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-3/tang.pdf.

Tanton:2000:FNG

[Tan00] James S. Tanton. Fibonacci numbers, generating sets, and hexagonal properties. Fibonacci Quarterly, 38(4):299–309, August 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/38-4/tanton.pdf.

Taous:2017:CGW

[Tao17] Mohammed Taous. On the 2-class group of Q( ¯ 5pF_{p}) where F_{p} is a prime Fibonacci number. Fibonacci Quarterly, 55(5):192– 200, December 2017. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Abstracts/55-5/taous.pdf.

Tauber:1965:SFM

[Tau65] Selmo Tauber. Summation formulae for multinomial coefficients. Fibonacci Quarterly, 3(2):95–100, April 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-2/
tauber.pdf.

Tauber:1968:LNF

[Tau68a] Selmo Tauber. Lah numbers for Fibonacci and Lucas polynomials. Fibonacci Quarterly, 6(5):93–99, November 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-5/tauber1.pdf.

Tauber:1968:LNP

[Tau68b] Selmo Tauber. Lah numbers for R-polynomials. Fibonacci Quarterly, 6(5):100–107, November 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-5/
tauber2.pdf.

Tauber:1968:FP

[Tau68c] Selmo Tauber. On Q-Fibonacci polynomials. Fibonacci Quarterly, 6(2):127–134, April 1968. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/6-2/tauber.
pdf.

Tauber:1973:FP

[Tau73a] Selmo Tauber. n-Fibonacci products. Fibonacci Quarterly, 11(2): 153–156, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/11-2/tauber1.pdf.

Tauber:1973:N

[Tau73b] Selmo Tauber. On K-numbers. Fibonacci Quarterly, 11(2):179– 183, April 1973. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/11-2/tauber2.pdf.

Tauber:1976:CNC

[Tau76] S. Tauber. Combinatorial numbers in C^{n}. Fibonacci Quarterly, 14(2):101–110, April 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-2/tauber.pdf.

Tauber:1989:SNR

[Tau89] Selmo Tauber. Some new results on quasi-orthogonal numbers. Fibonacci Quarterly, 27(3):194–200, June 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/27-3/tauber.pdf.

Taylor:1967:RFL

[Tay67] Laurence Taylor. Residues of Fibonacci-like sequences. Fibonacci Quarterly, 5(3):298–304, October 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-3/
taylor.pdf.

Taylor:1975:GLQ

[Tay75] Larry Taylor. The general law of quadratic reciprocity. Fibonacci Quarterly, 13(4):318–??, December 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-
4/taylor-a.pdf.

Taylor:1976:CRQ

[Tay76] Larry Taylor. A conjecture relating quartic reciprocity and quartic residuacity to primitive Pythagorean triples. Fibonacci Quarterly, 14(2):180–181, April 1976. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/14-2/taylor.
pdf.

Tosic:1991:AEN

[TB91] Ratko Tosic and Olga Bodroza. An algebraic expression for the number of Kekulé structures of benzenoid chains. Fibonacci Quarterly, 29(1):7–12, February 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-1/
tosic.pdf.

Tsang:2004:TCT

[TCC^{+}04] W. W. Tsang, C. F. Chong, K. P. Chow, L. C. K. Hui, and C. W. Tso. Tuning the collision test for power. In Vladimir Estivill-Castro, editor, Proceedings of the Twenty-Seventh Australasian Computer Science Conference (ACSC2004) Dunedin, NZ, January 2004, volume 26 of Conferences in research and practice in information technology, pages 23–30. Australian Computer Society, Sydney, Australia, 2004. ISBN 1-920682-05-8. LCCN QA75.5 .A88 2004.

Talebi:2016:AUB

[TD16] Gholamreza Talebi and Mohammad Ali Dehghan. Approximation of upper bound for matrix operators on the Fibonacci weighted sequence spaces. Linear Multilinear Algebra, 64(2):196– ??, 2016. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Talebi:2019:AUB

[TD19] Gholamreza Talebi and Mohammad Ali Dehghan. Approximation of upper bound for matrix operators on the Fibonacci weighted sequence spaces. Linear Multilinear Algebra, 67(8): 1717–1718, 2019. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Tiebekabe:2022:DE

[TD22] Pagdame Tiebekabe and Ismaïla Diouf. On the Diophantine equation F_{n_{1}} + F_{n_{2}} + F_{n_{3}} + F_{n_{4}} + F_{n_{5}} = 2^{a}. Fibonacci Quarterly, 60(5):384–??, December 2022. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/60-5/
tiebekabe.pdf.

Thoro:1981:EA

[TE81] Dmitri Thoro and Hugh Edgar. Exploring an algorithm. Fibonacci Quarterly, 19(3):271–275, August 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-3/thoro.pdf.

Tedford:2019:CIP

[Ted19a] Steven J. Tedford. Combinatorial identities for the Padovan numbers. Fibonacci Quarterly, 57(4):291–??, November 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-4/tedford1.pdf.

Tedford:2019:FG

[Ted19b] Steven J. Tedford. Fibonacci graphs. Fibonacci Quarterly, 57(4):347–??, November 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/57-4/
tedford2.pdf.

Tee:1994:PPZ

[Tee94] Garry J. Tee. Prime powers of zeros of monic polynomials with integer coefficients. Fibonacci Quarterly, 32(3):277–283, June 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-3/tee.pdf.

Temlyakov:1993:UFC

[Tem93] V. N. Temlyakov. Universality of the Fibonacci cubature formulas. Lecture Notes in Mathematics, 1550:178–184, 1993. CODEN LNMAA2. ISBN 3-540-56931-6 (print), 3-540-47792-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0117486/.

Tengely:2008:FGN

[Ten09] Szabolcs Tengely. Finding g-gonal numbers in recurrence sequences. Fibonacci Quarterly, 46/47(3):235–240, August 2008/ 2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/46_47-3/tengely.pdf.

Tepper:1974:CSP

[Tep74] Myron Tepper. Combinations and sums of powers. Fibonacci Quarterly, 12(2):196–198, April 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-2/
tepper.pdf.

Tepper:1976:SCP

[Tep76] Myron Tepper. Sums of combination products. Fibonacci Quarterly, 14(3):265–271, October 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
3/tepper.pdf.

Terr:1996:FEA

[Ter96a] David C. Terr. Fibonacci expansions and “F -adic Integers”. Fibonacci Quarterly, 34(2):156–163, May 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-2/
terr.pdf.

Terr:1996:SDF

[Ter96b] David C. Terr. On the sums of digits of Fibonacci numbers. Fibonacci Quarterly, 34(4):349–355, August 1996. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-4/terr.pdf.

Terr:2012:SII

[Ter12] David Terr. Some interesting infinite families of primitive Pythagorean triples. Fibonacci Quarterly, 50(1):68–??, February 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/50-1/terr.pdf.

Teuscher:2004:ATL

[Teu04] Christof Teuscher, editor. Alan Turing, life and legacy of a great thinker. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-20020-7 (hardcover), 3-642-05744-6 (print), 3-662-05642-9 (e-book). LCCN QA29.T8 A57 2004. Foreword by Douglas Hofstadter. Papers from the Conference “Turing Day: Computing Science 90 Years from the Birth of Alan Mathison Turing” held at the École Polytechnique Fédérale de Lausanne, Lausanne, June 28, 2002.

Torretto:1964:GBC

[TF64] Roseanna F. Torretto and J. Allen Fuchs. Generalized binomial coefficients. Fibonacci Quarterly, 2(4):296–302, December 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-4/torretto.pdf.

Tezuka:1992:FGL

[TF92] S. Tezuka and M. Fushimi. Fast generation of low discrepancy points based on Fibonacci polynomials. In Swain [Swa92], pages 433–437. ISBN 0-7803-0797-6 (softbound), 0-7803-0798-4 (casebound), 0-7803-0799-2 (microfiche). LCCN T57.62 .W787 1992. IEEE catalog number 92CH3202-9.

Tezuka:1993:CFP

[TF93] Shu Tezuka and Masanori Fushimi. Calculation of Fibonacci polynomials for GFSR sequences with low discrepancies. Mathematics of Computation, 60(202):763–770, April 1993. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Thanatipanonda:2019:SDT

[Tha19] Thotsaporn “Aek” Thanatipanonda. Statistics of domino tilings on a rectangular board. Fibonacci Quarterly, 57(5):145–??, December 2019. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Abstracts/57-5/thanatipanonda.pdf.

Theusch:1971:CM

[The71] Sister Christelle Theusch. Composition of Φ_{2}(x) modulo m. Fibonacci Quarterly, 9(1):23–27, February 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-1/
theusch.pdf.

Turner:2023:LSD

[THKH23] Holly-Anne Turner, Matthew Humpage, Hans Kerp, and Alexander J. Hetherington. Leaves and sporangia developed in rare non-Fibonacci spirals in early leafy plants. Science, 380(6650):1188– 1192, June 16, 2023. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL https://www.science.org/doi/
10.1126/science.adg4014.

Thoro:1963:BCa

[Tho63a] Dmitri Thoro. Beginners’ corner. Fibonacci Quarterly, 1(1):49– 52, February 1963. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/1-1/thoro.pdf.

Thoro:1963:BCb

[Tho63b] Dmitri Thoro. Beginners’ corner. Fibonacci Quarterly, 1(2):57– 59, April 1963. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/1-2/thoro.pdf.

Thoro:1963:BCc

[Tho63c] Dmitri Thoro. Beginners’ corner. Fibonacci Quarterly, 1(3):53– 59, October 1963. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/1-3/thoro.pdf.

Thoro:1963:CNR

[Tho63d] Dmitri Thoro. Classroom notes: Regula falsi and the Fibonacci numbers. American Mathematical Monthly, 70(8):869, October 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Thoro:1963:HBC

[Tho63e] Dmitri Thoro. Hints to beginner’s corner problems. Fibonacci Quarterly, 1(2):80–??, April 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-2/
thoro2.pdf.

Thoro:1963:RGF

[Tho63f] Dmitri Thoro. Reciprocals of generalized Fibonacci numbers. Fibonacci Quarterly, 1(4):30–??, December 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/1-4/thoro.pdf.

Thoro:1963:SBC

[Tho63g] Dmitri Thoro. Solutions, beginner’s corner. Fibonacci Quarterly, 1(1):64–??, February 1963. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/1-1/thoro-
solutions.pdf.

Thoro:1964:AUT

[Tho64a] Dmitri Thoro. An application of unimodular transformations. Fibonacci Quarterly, 2(4):291–295, December 1964. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/2-4/thoro.pdf.

Thoro:1964:BCa

[Tho64b] Dmitri Thoro. Beginners’ corner. Fibonacci Quarterly, 2(1):53– 56, February 1964. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/2-1/thoro.pdf.

Thoro:1964:BCb

[Tho64c] Dmitri Thoro. Beginners’ corner. Fibonacci Quarterly, 2(2):135– 137, April 1964. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/2-2/thoro.pdf.

Thoro:1965:TFC

[Tho65] Dmitri Thoro. Two Fibonacci conjectures. Fibonacci Quarterly, 3(3):184–186, October 1965. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/3-3/thoro.pdf.

Tirman:1988:IDF

[TJ88] Alvin Tirman and T. Henry Jablonski, Jr. Identities derived on a Fibonacci multiplication table. Fibonacci Quarterly, 26(4):328– 331, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/26-4/tirman.pdf.

Tsangaris:1992:OTG

[TJ92] P. G. Tsangaris and J. P. Jones. And old theorem on the GCD and its application to primes. Fibonacci Quarterly, 30(3):194– 198, August 1992. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/30-3/tsangaris.pdf.

Tasci:2004:OGL

[TK04] Dursun Tasci and Emrah Kilic. On the order-k generalized Lucas numbers. Applied Mathematics and Computation, 155(3): 637–641, August 16, 2004. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Tuglu:2011:BFL

[TKS11] Naim Tuglu, E. Gokcen Kocer, and Alexey Stakhov. Bivariate Fibonacci like p-polynomials. Applied Mathematics and Computation, 217(24):10239–10246, August 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300311006771.

Tauraso:2008:FNT

[TL08] Roberto Tauraso and O. P. Lossers. Fibonacci numbers and tilling a board with cuts: 11241. American Mathematical Monthly, 115(9):858–859, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.

jstor.org/stable/27642624.

Tezuka:1993:LSA

[TLC93] Shu Tezuka, Pierre L’Ecuyer, and Raymond Couture. On the lattice structure of the add-with-carry and subtract-with-borrow random number generators. ACM Transactions on Modeling and Computer Simulation, 3(4):315–331, October 1993. CODEN ATMCEZ. ISSN 1049-3301 (print), 1558-1195 (electronic). See remark in [EH95, page 248], and [MZ91] for the original work analyzed in this paper.

Taher:2002:CGF

[TMR02a] R. B. Taher, M. Mouline, and M. Rachidi. Convergence of rgeneralized Fibonacci sequences and an extension of Ostrowski’s condition. Fibonacci Quarterly, 40(5):386–393, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-5/ben-taher.pdf.

Taher:2002:SSG

[TMR02b] R. B. Taher, M. Mouline, and M. Rachidi. Solving some general nonhomogeneous recurrence relations of order r by a linearization method and an application to polynomial and factorial polynomial cases. Fibonacci Quarterly, 40(1):79–84, February 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-1/ben-taher.pdf.

Taher:2006:FHD

[TMR06] R. Ben Taher, M. Mouline, and Mustapha Rachidi. Fibonacci– Horner decomposition of the matrix exponential and the fundamental system of solutions. Electronic Journal of Linear Algebra, 15(1):178–190, 2006. CODEN ???? ISSN 1081-3810 (print), 1537-9582 (electronic). URL http://repository.uwyo.

edu/ela/vol15/iss1/13.

Toller:2017:BRFa

[Tol17a] Owen Toller. Book review: Finding Fibonacci by Keith Devlin, pp. 241, £24.95, ISBN 978-0-691-17486-0, Princeton University Press (2017). The Mathematical Gazette, 101 (552):569–570, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/finding-fibonacci-by-keith-devlin-pp-241-2495-
isbn-9780691174860-princeton-university-press-2017/263E75E701900510EA12D5BE5458A741.

Toller:2017:BRF

[Tol17b] Owen Toller. Book review: Finding Fibonacci by Keith Devlin, pp. 241, £24.95, ISBN 978-0-691-17486-0, Princeton University Press (2017). The Mathematical Gazette, 101 (552):569–570, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://

www.cambridge.org/core/journals/mathematical-gazette/
article/finding-fibonacci-by-keith-devlin-pp-241-2495-
isbn-9780691174860-princeton-university-press-2017/263E75E701900510EA12D5BE5458A741.

Tong:1997:CFD

[Ton97] Jingcheng Tong. The constant for finite Diophantine approximation. Fibonacci Quarterly, 35(1):29–31, February 1997. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-1/tong.pdf.

Tonien:2024:FPS

[Ton24] Joseph Tonien. Fibonacci and pi squared. Fibonacci Quarterly, 62(1):65–74, February 2024. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/62-1/tonien.pdf.

Torrence:2019:PBF

[Tor19] Bruce Torrence. Passing the buck and firing Fibonacci: Adventures with the stochastic abacus. American Mathematical Monthly, 126(5):387–399, 2019. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Toscano:1978:SRG

[Tos78] L. Toscano. Some results for generalized Bernoulli, Euler, Stirling numbers. Fibonacci Quarterly, 16(2):103–111, April 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-2/toscano.pdf.

Toth:1987:NGE

[Tót87] László Tóth. A note on a generalization of Euler’s Φ function. Fibonacci Quarterly, 25(3):241–243, August 1987. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/25-3/toth.pdf.

Toth:2002:PPI

[Tót02] László Tóth. The probability that k positive integers are pairwise relatively prime. Fibonacci Quarterly, 40(1):13–18, February 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/40-1/toth.pdf.

Tovey:1985:MOB

[Tov85] Craig A. Tovey. Multiple occurrences of binomial coefficients. Fibonacci Quarterly, 23(4):356–358, November 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/23-4/tovey.pdf.

Taillie:1986:FDR

[TP86] C. Taillie and G. P. Patil. The Fibonacci distribution revisited. Communications in Statistics: Theory and Methods, 15(3):951– 959, 1986. CODEN CSTMDC. ISSN 0361-0926 (print), 1532-415X (electronic).

teRiele:1984:RCH

[tR84] H. J. J. te Riele. Rules for constructing hyperperfect numbers. Fibonacci Quarterly, 22(1):50–60, February 1984. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-1/teriele.pdf.

Turner:1990:GDZ

[TR90] J. C. Turner and T. D. Robb. Generalizations of the dual Zeckendorf integer representation theorems — discovery by Fibonacci trees and word patterns. Fibonacci Quarterly, 28(3):230–239, August 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/28-3/turner.pdf.

Taher:2001:AAR

[TR01] Rajae Ben Taher and Mustapha Rachidi. Application of the β-algorithm to the ratios of r-generalized Fibonacci sequences. Fibonacci Quarterly, 39(1):22–26, February 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/39-1/ben-taher.pdf.

Taher:2003:MPE

[TR03] Rajae Ben Taher and Mustapha Rachidi. On the matrix powers and exponential by the r-generalized Fibonacci sequences methods: the companion matrix case. Linear Algebra and its Applications, 370(1):341–353, September 1, 2003. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Taher:2004:APR

[TR04] Rajae Ben Taher and Mustapha Rachidi. On the E-algorithm for polynomial roots and linear recurrence relations. Fibonacci Quarterly, 42(2):98–105, May 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-2/
quartrachidi02_2004.pdf.

Tracy:1988:UCF

[Tra88] Craig A. Tracy. Universality class of a Fibonacci Ising model. Journal of Statistical Physics, 51(3–4):481–490, May 1988. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01028467.

Tootill:1973:ART

[TRE73] J. P. R. Tootill, W. D. Robinson, and D. J. Eagle. An asymptotically random Tausworthe sequence. Journal of the Association for Computing Machinery, 20(3):469–481, July 1973. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Trevisan:1996:NMU

[Tre96a] L. Trevisan. A note on minimum-area upward drawing of complete and Fibonacci trees. Information Processing Letters, 57 (5):231–??, March 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Trevisan:1996:NMA

[Tre96b] Luca Trevisan. A note on minimum-area upward drawing of complete and Fibonacci trees. Information Processing Letters, 57(5): 231–236, March 11, 1996. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Treeby:2016:FPD

[Tre16a] David Treeby. Further physical derivations of Fibonacci summations. Fibonacci Quarterly, 54(4):327–??, November 2016. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/54-4/treeby.pdf.

Treeby:2016:HFF

[Tre16b] David Treeby. Hidden formulas in Fibonacci tilings. Fibonacci Quarterly, 54(1):23–??, February 2016. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/54-1/
treeby.pdf.

Trigg:1965:ROT

[Tri65] Charles W. Trigg. A recursive operation on two-digit integers. Fibonacci Quarterly, 3(2):90–??, April 1965. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/3-2/
trigg.pdf.

Trigg:1967:C

[Tri67a] Charles W. Trigg. Curiosa in 1967. Fibonacci Quarterly, 5(5): 474–476, December 1967. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/5-5/trigg2.pdf.

Trigg:1967:DB

[Tri67b] Charles W. Trigg. A digital bracelet for 1967. Fibonacci Quarterly, 5(5):477–480, December 1967. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/5-5/
trigg3.pdf.

Trigg:1967:GP

[Tri67c] Charles W. Trigg. Getting primed for 1967. Fibonacci Quarterly, 5(5):472–473, December 1967. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/5-5/trigg1-a.
pdf.

Trigg:1967:LE

[Tri67d] Charles W. Trigg. A letter to the editors. Fibonacci Quarterly, 5 (4):370–??, December 1967. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/5-4/letter.pdf.

Trigg:1967:MQ

[Tri67e] Charles W. Trigg. Mathematical quickies. McGraw-Hill, New York, NY, USA, 1967. xi + 210 pp. LCCN QA95 .T75.

Trigg:1967:P

[Tri67f] Charles W. Trigg. Picking away at 1967. Fibonacci Quarterly, 5 (4):355–??, December 1967. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/5-4/trigg.pdf.

Trigg:1969:AAD

[Tri69] Charles W. Trigg. Associated additive decimal digital bracelets. Fibonacci Quarterly, 7(3):287–294, October 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/7-3/trigg.pdf.

Trigg:1971:KRT

[Tri71] Charles W. Trigg. Kaprekar’s routine with two-digit integers. Fibonacci Quarterly, 9(2):189–194, April 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-2/
trigg.pdf.

Trigg:1973:GPR

[Tri73] Charles W. Trigg. Geometric proof of a result of Lehmer’s. Fibonacci Quarterly, 11(5):539–540, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-5/trigg.pdf.

Trigg:1974:ASD

[Tri74a] Charles W. Trigg. Antimagic squares derived from the thirdorder magic square. Fibonacci Quarterly, 12(4):387–390, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/12-4/trigg2.pdf.

Trigg:1974:AP

[Tri74b] Charles W. Trigg. The Apollonius problem. Fibonacci Quarterly, 12(4):326–??, December 1974. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/12-4/trigg1.
pdf.

Trigg:1974:ISP

[Tri74c] Charles W. Trigg. Infinite sequences of palindromic triangular numbers. Fibonacci Quarterly, 12(2):209–211, April 1974. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-2/trigg.pdf.

Tripathi:1989:BFP

[Tri89] Amitabha Tripathi. A box filling problem. Fibonacci Quarterly, 27(5):465–466, November 1989. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/27-5/tripathi.
pdf.

Trif:2000:CSS

[Tri00a] Tiberiu Trif. Combinatorial sums and series involving inverses of binomial coefficients. Fibonacci Quarterly, 38(1):79–83, February 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/38-1/trif.pdf.

Tripathi:2000:NS

[Tri00b] Amitabha Tripathi. The number of solutions to ax + by = n. Fibonacci Quarterly, 38(4):290–293, August 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/38-4/tripathi.pdf.

Tripathi:2008:PTC

[Tri09] Amitabha Tripathi. On Pythagorean triples containing a fixed integer. Fibonacci Quarterly, 46/47(4):331–340, November 2008/ 2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/46_47-4/tripathi.pdf.

Tripathi:2010:NPP

[Tri10] Amitabha Tripathi. A note on products of primes that differ by a fixed integer. Fibonacci Quarterly, 48(2):144–149, May 2010. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/48-2/tripathi.pdf.

Trumper:1973:SGF

[Tru73] Frank J. D. Trumper. Some general Fibonacci shift formulae. Fibonacci Quarterly, 11(5):523–??, December 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-5/trumper.pdf.

Trzaska:1994:MNT

[Trz94] Z. W. Trzaska. Modified numerical triangle and the Fibonacci sequences. Fibonacci Quarterly, 32(2):124–129, May 1994. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/32-2/trzaska.pdf.

Trzaska:1996:FHT

[Trz96] Zdzislaw W. Trzaska. On Fibonacci hyperbolic trigonometry and modified Numerical triangles. Fibonacci Quarterly, 34(2):129– 138, May 1996. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/34-2/trzaska.pdf.

Toth:1989:AFC

[TS89a] L. Tóth and J. Sándor. An asymptotic formula concerning a generalized Euler function. Fibonacci Quarterly, 27(2):176–180, May 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/27-2/toth.pdf.

Turner:1989:TOC

[TS89b] J. C. Turner and A. G. Shannon. On K th-order colored convolution trees and a generalized Zeckendorf integer representation theorem. Fibonacci Quarterly, 27(5):439–447, November 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-5/turner.pdf.

Tosic:1992:FNN

[TS92] Ratko Tošić and Ivan Stojmenović. Fibonacci numbers and the numbers of perfect matchings of square, pentagonal, and hexagonal chains. Fibonacci Quarterly, 30(4):315–321, November 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/30-4/tosic.pdf.

Turner:2000:FSG

[TS00] J. C. Turner and A. G. Shannon. On Fibonacci sequences, geometry, and the m-square equation. Fibonacci Quarterly, 38(2): 98–103, May 2000. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/38-2/turner.pdf.

Trickovic:2004:PSC

[TS04] Slobodan B. Tričković and Miomir S. Stanković. On periodic solutions of a certain difference equation. Fibonacci Quarterly, 42(4):300–305, November 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
4/quarttrickovic04_2004.pdf.

Tsuno:2021:ERI

[Tsu21] Yuji Tsuno. Extended results on integer values of the generating functions for sequences given by Pell’s equations. Fibonacci Quarterly, 59(2):158–??, May 2021. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/59-2/tsuno.
pdf.

Tangboonduangjit:2016:DRP

[TT16] Aram Tangboonduangjit and Thotsaporn Thanatipanonda. Determinants of rising powers of second order linear recurrence entries by means of the Desnanot–Jacobi identity. Fibonacci Quarterly, 54(4):340–??, November 2016. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/54-
4/tangboonduangjit.pdf.

Tengely:2020:DEF

[TU20] Szabolcs Tengely and MacIej Ulas. The Diophantine equation F_{n} = P (x). International Journal of Number Theory (IJNT), 16 (09):2095–2111, October 2020. ISSN 1793-0421 (print), 1793-

7310 (electronic). URL https://www.worldscientific.com/
doi/10.1142/S1793042120501079.

Tuenter:2002:WAS

[Tue02] Hans J. H. Tuenter. Walking into an absolute sum. Fibonacci Quarterly, 40(2):175–180, May 2002. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/40-2/
tuenter.pdf.

Tuenter:2021:IRL

[Tue21] Hans J. H. Tuenter. Inverse relations for Lucas sequences. Fibonacci Quarterly, 59(3):246–??, August 2021. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/59-3/tuenter.pdf.

Tuenter:2022:FSI

[Tue22] Hans J. H. Tuenter. Fibonacci summation identities arising from Catalan’s identity. Fibonacci Quarterly, 60(4):312–??, November 2022. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/60-4/tuenter.pdf.

Tuenter:2023:SCA

[Tue23a] Hans J. H. Tuenter. In search of Comrade Agronomof: Some Tribonacci history. American Mathematical Monthly, 130(8):708– 719, 2023. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Tuenter:2023:TFS

[Tue23b] Hans J. H. Tuenter. A triple Fibonacci sum and some historical notes. Fibonacci Quarterly, 61(4):361–??, November 2023. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-4/tuenter.pdf.

Turner:1974:CCP

[Tur74] Michael R. Turner. Certain congruence properties (modulo 100) of Fibonacci numbers. Fibonacci Quarterly, 12(1):87–91, February 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/12-1/turner.pdf.

Turner:1979:PTO

[Tur79] Stephen John Turner. Probability via the N th order Fibonacci-T sequence. Fibonacci Quarterly, 17(1):23–28, February 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/17-1/turner.pdf.

Turner:1986:CKF

[Tur86] J. C. Turner. On a class of knots with Fibonacci invariant numbers. Fibonacci Quarterly, 24(1):61–66, February 1986. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/24-1/turner.pdf.

Turner:1988:FWP

[Tur88a] J. C. Turner. Fibonacci word patterns and binary sequences. Fibonacci Quarterly, 26(3):233–246, August 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/26-3/turner2.pdf.

Turner:1988:CTP

[Tur88b] John C. Turner. Convolution trees and Pascal-T triangles. Fibonacci Quarterly, 26(4):354–365, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/26-4/turner.pdf.

Turner:1988:FF

[Tur88c] John C. Turner. On Folyominoes and Feudominoes. Fibonacci Quarterly, 26(3):205–218, August 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
3/turner1.pdf.

Turner:1989:AOW

[Tur89a] J. C. Turner. The alpha and the omega of the Wythoff pairs. Fibonacci Quarterly, 27(1):76–86, February 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-1/turner.pdf.

Turner:1989:NFF

[Tur89b] John C. Turner. Note on a family of Fibonacci-like sequences. Fibonacci Quarterly, 27(3):229–231, June 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-3/
turner.pdf.

Turner:2003:SFG

[Tur03] J. C. Turner. Some fractals in goldpoint geometry. Fibonacci Quarterly, 41(1):63–71, February 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-
1/turner.pdf.

Taranenko:2007:FRF

[TV07] Andrej Taranenko and Aleksander Vesel. Fast recognition of Fibonacci cubes. Algorithmica, 49(2):81–93, October 2007. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0178-4617&volume=49&issue=2&spage=
81.

Tianming:1996:RSN

[TZ96] Wang Tianming and Zhang Zhizheng. Recurrence sequences and Nörlund–Euler polynomials. Fibonacci Quarterly, 34(4):314–319, August 1996. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/34-4/wang.pdf.

Uppuluri:1969:NGF

[UC69] V. R. Rao Uppuluri and John A. Carpenter. Numbers generated by the function exp(1 − e^{x}). Fibonacci Quarterly, 7(4):437–448, November 1969. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/7-4/uppuluri.pdf.

Udrea:1998:DE

[Udr98] Gheorghe Udrea. The Diophantine equations x^{2}−k = T_{n}(a^{2}±1). Fibonacci Quarterly, 36(4):335–338, August 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-4/udrea.pdf.

Ulas:2023:RGD

[Ula23] Maciej Ulas. Representing generalized derangements as sums of three squares. Fibonacci Quarterly, 61(3):231–??, August 2023. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/61-3/ulas.pdf.

Ulivi:2011:LFM

[Uli11] Elisabetta Ulivi. On Leonardo Fibonacci and on masters of the abacus in Pisa from the thirteenth to the fifteenth century. Bollettino di Storia delle Scienze Matematiche, 31(2):247–286, 2011. ISSN 0392-4432 (print), 1724-1650 (electronic).

Umansky:1971:PR

[Uma71] Harlan Umansky. Pythagoras revisited. Fibonacci Quarterly, 9 (1):83–86, February 1971. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/9-1/umansky.pdf.

Umansky:1972:NPT

[Uma72] Harlan L. Umansky. A note on Pythagorean triplets. Fibonacci Quarterly, 10(2):203–206, February 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
2/umansky-a.pdf.

Uslu:2007:ENC

[UNS07] K. Uslu, A. Nalli, and M. Sen. Erratum to “On the norms of circulant matrices with the Fibonacci and Lucas numbers” [Appl. Math. Comput. 160 (1) (2005) 125–132]. Applied Mathematics and Computation, 190(1):997–998, July 1, 2007. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See [Sol07].

Uppuluri:1983:WTG

[UP83] V. R. R. Uppuluri and S. A. Patil. Waiting times and generalized Fibonacci sequences. Fibonacci Quarterly, 21(4):242–249, November 1983. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/21-4/uppuluri.pdf.

Utz:1977:DE

[Utz77] W. R. Utz. The Diophantine equation (x_{1} + x_{2} + ⋯ + x_{n})^{2} = x_{1}^{3} + x_{2}^{3} + ⋯ + x_{n}^{3}. Fibonacci Quarterly, 15(1):14–??, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/15-1/utz-a.pdf.

Vajda:1989:FLN

[Vaj89] S. Vajda. Fibonacci and Lucas numbers, and the golden section: theory and applications. Ellis Horwood series in mathematics and its applications. Ellis Horwood, New York, NY, USA, 1989. ISBN 0-7458-0715-1, 0-470-21508-9 (Halsted Press). 189 pp. LCCN QA241 .V24 1989. US$25.00.

Vajda:2008:FLN

[Vaj08] S. Vajda. Fibonacci and Lucas numbers, and the golden section: theory and applications. Dover, New York, NY, USA, Dover edition, 2008. ISBN 0-486-46276-5 (paperback). 189 pp. LCCN QA241 .V24 2008. URL http://www.loc.gov/

catdir/enhancements/fy0739/2007031399-d.html;   http:
//www.loc.gov/catdir/toc/fy1001/2007031399.html;   https:
//www.math.utah.edu/pub/tex/bib/fibquart.bib.

VandenEynden:1986:DBS

[Van86] Charles Vanden Eynden. Differences between squares and powerful numbers. Fibonacci Quarterly, 24(4):347–348, November 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/24-4/vandeneynden.pdf.

Vantieghem:1996:SRE

[Van96] E. Vantieghem. On sequences related to expansions of real numbers. Fibonacci Quarterly, 34(4):356–361, August 1996. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/34-4/vantieghem.pdf.

Vandervelde:2012:DFS

[Van12] Sam Vandervelde. On the divisibility of Fibonacci sequences by primes of index two. Fibonacci Quarterly, 50(3):207–??, August 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/50-3/vandervelde.pdf.

Vardavas:1989:FST

[Var89] Ilias Mihail Vardavas. A Fibonacci search technique for model parameter selection. Ecological Modelling, 48(1–2):65–81, October 1989. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0304380089900604.

Vaughan:1976:NSA

[Vau76] Theresa P. Vaughan. A note on some arithmetic functions connected with the Fibonacci numbers. Fibonacci Quarterly, 14(3): 244–248, October 1976. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/14-3/vaughan.pdf.

Vegh:1969:RTW

[Veg69] Emanuel Vegh. Remark on a theorem by Waksman. Fibonacci Quarterly, 7(3):230–??, October 1969. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/7-3/vegh.
pdf.

Vegh:1974:CER

[Veg74] Emanuel Vegh. Concerning an equivalence relation for matrices. Fibonacci Quarterly, 12(4):391–392, December 1974. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-4/vegh.pdf.

Vesel:2015:LRE

[Ves15] Aleksander Vesel. Linear recognition and embedding of Fibonacci cubes. Algorithmica, 71(4):1021–1034, April 2015. CODEN ALGOEJ. ISSN 0178-4617 (print), 1432-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s00453-013-9839-3.

Vinson:1963:RPM

[Vin63] John Vinson. The relation of the period modulo m to the rank of apparition of m in the Fibonacci sequence. Fibonacci Quarterly, 1 (2):37–46, April 1963. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/1-2/vinson.pdf.

Vince:1978:FSM

[Vin78] Andrew Vince. The Fibonacci sequence modulo N. Fibonacci Quarterly, 16(5):403–406, October 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-
5/vince.pdf.

Viswanath:2000:RFS

[Vis00] Divakar Viswanath. Random Fibonacci sequences and the number 1.13198824 . . .. Mathematics of Computation, 69 (231):1131–1155, July 2000. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http:/

/www.ams.org/journal-getitem?pii=S0025-5718-99-01145-
X;   http://www.ams.org/mcom/2000-69-231/S0025-5718-99-
01145-X/S0025-5718-99-01145-X.dvi;   http://www.ams.org/
mcom/2000-69-231/S0025-5718-99-01145-X/S0025-5718-99-
01145-X.pdf;   http://www.ams.org/mcom/2000-69-231/S0025-
5718-99-01145-X/S0025-5718-99-01145-X.ps;   http://www.
ams.org/mcom/2000-69-231/S0025-5718-99-01145-X/S0025-
5718-99-01145-X.tex.

Vattulainen:1993:IIP

[VKSAN93] I. Vattulainen, K. Kankaala, J. Saarinen, and T. Ala-Nissila. Influence of implementation on the properties of pseudorandom number generators with a carry bit. arxiv.org, ??(??):??, June

8, 1993. URL http://arxiv.org/abs/hep-lat/9306008.

Volodin:1994:NMC

[Vol94] Nikolai A. Volodin. Number of multinomial coefficients not divisible by a prime. Fibonacci Quarterly, 32(5):402–406, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/32-5/volodin.pdf.

Voll:2010:CIR

[Vol10] Nils Gaute Voll. The Cassini identity and its relatives. Fibonacci Quarterly, 48(3):197–201, August 2010. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/48-
3/voll.pdf.

Voll:2013:SIF

[Vol13] Nils Gaute Voll. Some identities for four term recurrence relations. Fibonacci Quarterly, 51(3):268–??, August 2013. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/51-3/voll.pdf.

Vorobev:1961:FN

[Vor61] Nikolai N. Vorobev. Fibonacci numbers. In Ian Naismith Sneddon, editor, Fibonacci Numbers, volume 2 of Popular lectures in mathematics series, pages viii + 66. Pergamon Press, Oxford, UK, 1961. ISBN 0-932750-03-6. LCCN QA241 .V613 1961. Translation by Halina Moss. Reprinted in [Vor83].

Vorobev:1983:FN

[Vor83] Nikolai N. Vorobev. Fibonacci numbers. In Ian Naismith Sneddon, editor, Fibonacci Numbers, page ?? New Classics Library, ????, 1983. Translation by Halina Moss. Reprint of [Vor61].

Vornicescu:2008:DCH

[Vor09] Neculae Vornicescu. Disjoint covering of N by a homogeneous linear recurrence. Fibonacci Quarterly, 46/47(1):79–84, February 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/46_47-1/vornicescu.pdf.

Voss:1998:DT

[Vos98] James E. Voss. Divisibility tests in N. Fibonacci Quarterly, 36 (1):43–44, February 1998. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/36-1/voss.pdf.

Voutier:1995:PDL

[Vou95] Paul M. Voutier. Primitive divisors of Lucas and Lehmer sequences. Mathematics of Computation, 64(210):869–888, April 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Viader:1999:NPE

[VPB99] P. Viader, J. Paradis, and Bibiloni. Note on the Pierce expansion of a logarithm. Fibonacci Quarterly, 37(3):198–202, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-3/viader.pdf.

vanRavenstein:1989:OSP

[vR89] Tony van Ravenstein. Optimal spacing of points on a circle. Fibonacci Quarterly, 27(1):18–24, February 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/27-1/vanravenstein.pdf.

vanRavenstein:1990:CTG

[vRWT90] Tony van Ravenstein, Graham Winley, and Keith Tognetti. Characteristics and the three gap theorem. Fibonacci Quarterly, 28 (3):204–213, August 1990. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/28-3/vanravenstein.
pdf.

Vsemirnov:2011:QIC

[Vse11] Maxim Vsemirnov. Quadratic identities for a class of Fibonaccilike polynomials. Fibonacci Quarterly, 49(3):220–226, August 2011. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/49-3/vsemirnov.pdf.

VanRavenstein:1985:PCG

[VWT85] Tony Van Ravenstein, Graham Winley, and Keith Tognetti. A property of convergents to golden mean. Fibonacci Quarterly, 23(2):155–157, May 1985. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/23-2/vanravenstein.
pdf.

Waddill:1974:MGF

[Wad74] Marcellus E. Waddill. Matrices and generalized Fibonacci sequences. Fibonacci Quarterly, 12(4):381–386, December 1974.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/12-4/waddill.pdf.

Waddill:1978:SPG

[Wad78] Marcellus E. Waddill. Some properties of a generalized Fibonacci sequence modulo m. Fibonacci Quarterly, 16(4):344–353, August 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/16-4/waddill.pdf.

Waddill:1992:SPT

[Wad92a] Marcellus E. Waddill. Some properties of the Tetranacci sequence modulo m. Fibonacci Quarterly, 30(3):232–238, August 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/30-3/waddill.pdf.

Waddill:1992:TSG

[Wad92b] Marcellus E. Waddill. The Tetranacci sequence and generalizations. Fibonacci Quarterly, 30(1):9–19, February 1992. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/30-1/waddill.pdf.

Wagstaff:1981:IPS

[Wag81] Samuel S. Wagstaff, Jr. Iterating the product of shifted digits. Fibonacci Quarterly, 19(4):340–346, October 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-4/wagstaff.pdf.

Wagner:2006:FNG

[Wag06] Stephan G. Wagner. The Fibonacci number of generalized Petersen graphs. Fibonacci Quarterly, 44(4):362–367, November 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/44-4/wagner.pdf.

Wagner:2007:FNF

[Wag07] Stephan G. Wagner. The Fibonacci number of Fibonacci trees and a related family of polynomial recurrence systems. Fibonacci Quarterly, 45(3):247–253, August 2007. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/45-
3/wagner.pdf.

Wall:1960:FSM

[Wal60] D. D. Wall. Fibonacci series modulo m. American Mathematical Monthly, 67(6):525–532, June/July 1960. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Wall:1963:SRC

[Wal63] Charles R. Wall. Some remarks on Carlitz’s Fibonacci array. Fibonacci Quarterly, 1(4):23–29, December 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/1-4/wall.pdf.

Walker:1976:CIP

[Wal76a] David T. Walker. Consecutive integer pairs of powerful numbers and related Diophantine equations. Fibonacci Quarterly, 14(2): 111–116, April 1976. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/14-2/walker.pdf.

Walton:1976:LPC

[Wal76b] J. E. Walton. Lucas polynomials and certain circular functions of matrices. Fibonacci Quarterly, 14(1):83–87, February 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/14-1/walton.pdf.

Wall:1979:SCI

[Wal79] Charles R. Wall. Some congruences involving generalized Fibonacci numbers. Fibonacci Quarterly, 17(1):29–33, February 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/17-1/wall.pdf.

Walther:1980:FGF

[Wal80] G. Walther. Free group and Fibonacci sequence. Fibonacci Quarterly, 18(3):268–272, October 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-
3/walther.pdf.

Wall:1983:UHN

[Wal83] Charles R. Wall. Unitary harmonic numbers. Fibonacci Quarterly, 21(1):18–25, February 1983. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/21-1/wall.pdf.

Wall:1984:FLS

[Wal84a] Charles R. Wall. A Fibonacci-like sequence of abundant numbers. Fibonacci Quarterly, 22(4):349–??, November 1984. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/22-4/wall.pdf.

Wall:1984:R

[Wal84b] Charles R. Wall. Response. Fibonacci Quarterly, 22(4):365–??, November 1984. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/22-4/response.pdf.

Walker:1985:GCS

[Wal85a] Marshall Walker. Golden cuboid sequences. Fibonacci Quarterly, 23(2):153–154, May 1985. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/23-2/walker.pdf.

Wall:1985:GES

[Wal85b] Charles R. Wall. Guessing exact solutions. Fibonacci Quarterly, 23(1):80–??, February 1985. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/23-1/wall2.pdf.

Wall:1985:TFN

[Wal85c] Charles R. Wall. On triangular Fibonacci numbers. Fibonacci Quarterly, 23(1):77–79, February 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-
1/wall1.pdf.

Wall:1987:ASD

[Wal87a] Charles R. Wall. Analogs of Smith’s determinant. Fibonacci Quarterly, 25(4):343–345, November 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-
4/wall2.pdf.

Wall:1987:LOC

[Wal87b] Charles R. Wall. On the largest odd component of a unitary perfect number. Fibonacci Quarterly, 25(4):312–316, November 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/25-4/wall1.pdf.

Wall:1988:NUP

[Wal88] Charles R. Wall. New unitary perfect numbers have at least nine odd components. Fibonacci Quarterly, 26(4):312–317, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/26-4/wall.pdf.

Wall:1990:TDC

[Wal90] Charles R. Wall. Terminating decimals in the Cantor ternary set. Fibonacci Quarterly, 28(2):98–101, May 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-2/
wall.pdf.

Walczak:2010:SRS

[Wal10] Bartosz Walczak. A simple representation of subwords of the Fibonacci word. Information Processing Letters, 110(21):956–960, October 15, 2010. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1993:RPP

[Wan93] Jun Wang. Reduced ϕ-partitions of positive integers. Fibonacci Quarterly, 31(4):365–369, November 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-
4/wang.pdf.

Wang:1995:DSF

[Wan95] Jun Wang. On the k th derivative sequences of Fibonacci and Lucas polynomials. Fibonacci Quarterly, 33(2):174–178, May 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-2/wang.pdf.

Wang:2000:SBF

[Wan00] Guo-Jun Wang. Separable Boolean functions and generalized Fibonacci sequences. Computers and Mathematics and Applications, 39(3–4):205–216, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0898122199003466.

Wang:2005:SIS

[Wan05] Yi Wang. Self-inverse sequences related to a binomial inverse pair. Fibonacci Quarterly, 43(1):46–52, February 2005. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers/43-1/paper43-1-6.pdf.

Wang:2020:GWT

[Wan20] Kai Wang. Girard–Waring type formula for a generalized Fibonacci sequence. Fibonacci Quarterly, 58(5):229–??, December 2020. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/58-5/wang.pdf.

Washington:1981:BLF

[Was81] Lawrence C. Washington. Benford’s Law for Fibonacci and Lucas numbers. Fibonacci Quarterly, 19(2):175–177, April 1981. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/19-2/washington.pdf.

Washington:1999:SRF

[Was99] Lawrence C. Washington. Some remarks on Fibonacci matrices. Fibonacci Quarterly, 37(4):333–341, November 1999. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-4/washington.pdf.

Watson:1986:RCB

[Wat86] J. F. Watson. Recursion: a choice between readability and execution speed. ACM SIGPLAN Notices, 21(5):82–90, May 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Waterhouse:1988:ISC

[Wat88] William C. Waterhouse. Integral 4 by 4 skew circulants. Fibonacci Quarterly, 26(2):172–177, May 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-2/
waterhouse.pdf.

Waterhouse:1992:CFP

[Wat92] William C. Waterhouse. Continued fractions and Pythagorean triples. Fibonacci Quarterly, 30(2):144–147, May 1992. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/30-2/waterhouse.pdf.

Waller:1980:SS

[WB80] W. G. Waller and Mahadev Banerjee. Summation of the series y^{n}+(y+1)^{n}+⋯ +x^{n}. Fibonacci Quarterly, 18(1):35–??, February 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/18-1/waller.pdf.

Wassell:2012:ELR

[WB12] Stephen R. Wassell and Samantha Benito. Edge-length ratios between dual platonic solids: a surprisingly new result involving the Golden Ratio. Fibonacci Quarterly, 50(2):144–??, May 2012. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/50-2/wassell.pdf.

Wang:1999:MCL

[WCL99] Ching-Te Wang, Chin-Chen Chang, and Chu-Hsing Lin. A method for computing Lucas sequences. Computers and Mathematics and Applications, 38(11–12):187–196, December 1999. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0898122199002977.

Weiss:1976:MEC

[WD76] George H. Weiss and Menachem Dishon. A method for the evaluation of certain sums involving binomial coefficients. Fibonacci Quarterly, 14(1):75–77, February 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
1/weiss.pdf.

Webb:1975:DFD

[Web75] William Webb. Distribution of the first digits of Fibonacci numbers. Fibonacci Quarterly, 13(4):334–336, December 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-4/webb.pdf.

Webb:1982:LFN

[Web82] William A. Webb. The length of the four-number game. Fibonacci Quarterly, 20(1):33–35, February 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-
1/webb.pdf.

Webber:1995:OFP

[Web95a] Adam Webber. Optimization of functional programs by grammar thinning. ACM Transactions on Programming Languages and Systems, 17(2):293–330, March 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/toc/Abstracts/0164-0925/201067.html;
 https://www.math.utah.edu/pub/tex/bib/fibquart.bib.

Webster:1995:CPF

[Web95b] Roger Webster. A combinatorial problem with a Fibonacci solution. Fibonacci Quarterly, 33(1):26–31, February 1995. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/33-1/webster.pdf.

Wegener:1975:PPT

[Weg75] Delano P. Wegener. Primitive Pythagorean triples with sum of difference of legs equal to a prime. Fibonacci Quarterly, 13(3): 263–277, October 1975. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/13-3/wegener.pdf.

Wegener:1977:SAP

[Weg77] Delano P. Wegener. Semi-associates in Z[\sqrt{2}] and primitive Pythagorean triples. Fibonacci Quarterly, 15(3):258–264, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/15-3/wegener.pdf.

Wegener:1981:APE

[Weg81a] Delano P. Wegener. An application of Pell’s equation. Fibonacci Quarterly, 19(5):450–??, December 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-
5/wegener2.pdf.

Wegener:1981:PPT

[Weg81b] Delano P. Wegener. Primitive Pythagorean triples and the infinitude of primes. Fibonacci Quarterly, 19(5):449–??, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/19-5/wegener1.pdf.

Weinstein:1966:DPF

[Wei66a] Lenard Weinstein. A divisibility property of Fibonacci numbers. Fibonacci Quarterly, 4(1):83–84, February 1966. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/4-1/weinstein1.pdf.

Weinstein:1966:LE

[Wei66b] Lenard Weinstein. Letter to the editor. Fibonacci Quarterly, 4 (1):88–??, February 1966. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/4-1/weinstein2.pdf.

Weintraub:1973:SAL

[Wei73] Sol Weintraub. On storing and analyzing large strings of primes. Fibonacci Quarterly, 11(4):438–440, November 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-4/weintraub.pdf.

Weinstein:1979:APC

[Wei79] Gerald Weinstein. An algorithm for packing complements of finite sets of integers. Fibonacci Quarterly, 17(4):289–292, December 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/17-4/weinstein.pdf.

Weiss:1988:BOP

[Wei88] Eric A. Weiss. Biographies: Oh, pioneers! Annals of the History of Computing, 10(4):348–361, October/December 1988. CODEN

AHCOE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1988/pdf/a4348.pdf;   http://www.computer.
org/annals/an1988/a4348abs.htm.

Weinstein:2016:NFP

[Wei16] F. V. Weinstein. Notes on Fibonacci partitions. Experimental Mathematics, 25(4):482–499, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic).

Weland:1967:SRP

[Wel67] Kathleen Weland. Some rabbit production results involving Fibonacci numbers. Fibonacci Quarterly, 5(2):195–200, April 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-2/weland.pdf.

Wells:1994:FLT

[Wel94] Diana L. Wells. The Fibonacci and Lucas triangles modulo 2. Fibonacci Quarterly, 32(2):111–123, May 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-2/
wells.pdf.

Wenchang:2002:SBC

[Wen02] Chu Wenchang. Some binomial convolution formulas. Fibonacci Quarterly, 40(1):19–32, February 2002. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/40-1/chu.
pdf.

Werner:1983:CPH

[Wer83] Wilhelm Werner. Comment on problem H-315. Fibonacci Quarterly, 21(3):173–??, August 1983. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/21-3/werner-a.
pdf.

Wessner:1966:BSF

[Wes66] John Wessner. Binomial sums of Fibonacci powers. Fibonacci Quarterly, 4(4):355–358, December 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-4/
wessner.pdf.

Weszely:1980:FLP

[Wes80] Julia Weszely. Fibonacci, Leonardo Pisano (c. 1170–c. 1240). (Romanian). Gaz. Mat. Mat. Inform., 1(3):124–126, ???? 1980. ISSN 0016-5433.

Weinshenk:1970:SEO

[WH70] R. J. Weinshenk and V. E. Hoggatt, Jr. On solving C_{n+2} = C_{n+1} + C_{n} + n^{m} by expansions and operators. Fibonacci Quarterly, 8(1):39–48, February 1970. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/8-1/weinshenk-
a.pdf.

Walton:1971:SPC

[WH71] J. E. Walton and A. F. Horadam. Some properties of certain generalized Fibonacci matrices. Fibonacci Quarterly, 9(3):264– 276, May 1971. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/9-3/walton.pdf.

Walton:1974:SFI

[WH74a] J. E. Walton and A. F. Horadam. Some further identities for the generalized Fibonacci sequence H_{n}. Fibonacci Quarterly, 12 (3):272–280, October 1974. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/12-3/walton2.pdf.

Walton:1974:SAG

[WH74b] J. W. Walton and A. F. Horadam. Some aspects of generalized Fibonacci numbers. Fibonacci Quarterly, 12(3):241–250, October 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/12-3/walton1.pdf.

Walton:1984:GPP

[WH84] J. E. Walton and A. F. Horadam. Generalized Pell polynomials and other polynomials. Fibonacci Quarterly, 22(4):336–339, November 1984. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/22-4/walton.pdf.

Wang:2003:SFP

[WH03] Xinghua Wang and Leetsch C. Hsu. A summation formula for power series using Eulerian fractions. Fibonacci Quarterly, 41(1): 23–30, February 2003. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/41-1/wang.pdf.

Wang:2004:ZCF

[WH04] Yi Wang and Mingfeng He. Zeros of a class of Fibonaccitype polynomials. Fibonacci Quarterly, 42(4):341–347, November 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Papers1/42-4/quartwang04_2004.pdf.

Webb:2014:CR

[WH14] William Webb and Nathan Hamlin. Compositions and recurrences. Fibonacci Quarterly, 52(5):201–??, December 2014. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Papers1/52-5/Webb.pdf.

Whinihan:1963:FN

[Whi63] Michael J. Whinihan. Fibonacci Nim. Fibonacci Quarterly, 1(4): 9–14, December 1963. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/1-4/whinihan.pdf.

Whitney:1966:CRF

[Whi66a] Raymond E. Whitney. Composition of recursive formulae. Fibonacci Quarterly, 4(4):363–366, December 1966. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/4-4/whitney.pdf.

Whitney:1966:ERR

[Whi66b] Raymond E. Whitney. Extensions of recurrence relations. Fibonacci Quarterly, 4(1):37–42, February 1966. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/4-1/
whitney.pdf.

Whitney:1967:APS

[Whi67a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 5(5):435–443, December 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/5-5/advanced5-5.pdf.

Whitney:1967:PLR

[Whi67b] Raymond E. Whitney. A property of linear recursion relations. Fibonacci Quarterly, 5(3):281–285, October 1967. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-3/whitney.pdf.

Whitney:1968:APSa

[Whi68a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 6(1):50–57, February 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-1/
advanced6-1.pdf.

Whitney:1968:APSb

[Whi68b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 6(2):142–150, April 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-2/
advanced6-2.pdf.

Whitney:1968:APSc

[Whi68c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 6(4):250–260, October 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/6-4/advanced6-4.pdf.

Whitney:1968:APSd

[Whi68d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 6(6):350–359, December 1968. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/6-6/advanced6-6.pdf.

Whitney:1969:APSa

[Whi69a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 7(1):56–65, February 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-1/
advanced7-1.pdf.

Whitney:1969:APSb

[Whi69b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 7(2):169–179, April 1969. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/7-2/
advanced7-2.pdf.

Whitney:1969:APSc

[Whi69c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 7(3):277–283, October 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/7-3/advanced7-3.pdf.

Whitney:1969:APSd

[Whi69d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 7(5):511–522, December 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/7-5/advanced7-5.pdf.

Whitney:1970:APSa

[Whi70a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 8(1):74–82, February 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-1/
advanced8-1.pdf.

Whitney:1970:APSb

[Whi70b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 8(3):268–279, April 1970. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/8-3/
advanced8-3.pdf.

Whitney:1970:APSc

[Whi70c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 8(4):383–392, October 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/8-4/advanced8-4.pdf.

Whitney:1970:APSd

[Whi70d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 8(5):487–497, December 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/8-5/advanced8-5.pdf.

Whitney:1970:CDE

[Whi70e] Raymond E. Whitney. On a class of difference equations. Fibonacci Quarterly, 8(5):470–474, December 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/8-5/whitney.pdf.

Whitney:1971:APSa

[Whi71a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 9(1):61–73, February 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-1/
advanced9-1.pdf.

Whitney:1971:APSb

[Whi71b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 9(2):134–145, April 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-2/
advanced9-2-a.pdf.

Whitney:1971:APSc

[Whi71c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 9(4):389–401, October 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/9-4/advanced9-4.pdf.

Whitney:1971:APSd

[Whi71d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 9(5):512–524, December 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/9-5/advanced9-5.pdf.

Whitney:1972:APSa

[Whi72a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 10(2):185–196, February 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-2/advanced10-2-a.pdf.

Whitney:1972:APSb

[Whi72b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 10(3):283–294, April 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-3/
advanced10-3.pdf.

Whitney:1972:APSc

[Whi72c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 10(4):413–421, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-4/advanced10-4.pdf.

Whitney:1972:APSd

[Whi72d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 10(6):629–642, December 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/10-6/advanced10-6.pdf.

Whitney:1973:APSa

[Whi73a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 11(1):72–77, February 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-1/advanced11-1.pdf.

Whitney:1973:APSb

[Whi73b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 11(2):184–188, April 1973. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/11-2/
advanced11-2.pdf.

Whitney:1973:APSc

[Whi73c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 11(3):312–314, October 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-3/advanced11-3.pdf.

Whitney:1973:APSd

[Whi73d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 11(5):501–507, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-5/advanced11-5.pdf.

Whitney:1974:APSa

[Whi74a] R. E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 12(1):107–112, February 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
1/advanced12-1.pdf.

Whitney:1974:APSb

[Whi74b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 12(2):213–219, April 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-2/
advanced12-2.pdf.

Whitney:1974:APSc

[Whi74c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 12(3):309–312, October 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/12-3/advanced12-3.pdf.

Whitney:1974:APSd

[Whi74d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 12(4):398–402, December 1974. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/12-4/advanced12-4.pdf.

Whitney:1975:APSa

[Whi75a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 13(1):89–93, February 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-1/advanced13-1.pdf.

Whitney:1975:APSb

[Whi75b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 13(2):185–189, April 1975. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/13-2/
advanced13-2.pdf.

Whitney:1975:APSc

[Whi75c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 13(3):281–284, October 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-3/advanced13-3.pdf.

Whitney:1975:APSd

[Whi75d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 13(4):369–372, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-4/advanced13-4.pdf.

Whitney:1976:APSa

[Whi76a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 14(1):88–92, February 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-1/advanced14-1.pdf.

Whitney:1976:APSb

[Whi76b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 14(2):182–187, April 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-2/
advanced14-2.pdf.

Whitney:1976:APSc

[Whi76c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 14(3):282–285, October 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-3/advanced14-3.pdf.

Whitney:1976:APSd

[Whi76d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 14(5):466–469, December 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/14-5/advanced14-5-a.pdf.

Whitford:1977:BFG

[Whi77a] A. K. Whitford. Binet’s formula generalized. Fibonacci Quarterly, 15(1):21–??, February 1977. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/15-1/whitford-
a.pdf.

Whitney:1977:APSa

[Whi77b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 15(1):89–92, February 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/15-1/advanced15-1-a.pdf.

Whitney:1977:APSb

[Whi77c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 15(2):185–188, April 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-2/
advanced15-2-a.pdf.

Whitney:1977:APSc

[Whi77d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 15(3):281–284, October 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/15-3/advanced15-3-a.pdf.

Whitney:1977:APSd

[Whi77e] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 15(4):371–374, December 1977. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/15-4/advanced15-4.pdf.

Whitney:1978:APSa

[Whi78a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 16(1):92–96, February 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/16-1/advanced16-1.pdf.

Whitney:1978:APSb

[Whi78b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 16(2):188–192, April 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-2/
advanced16-2.pdf.

Whitney:1978:APSc

[Whi78c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 16(5):477–480, October 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/16-5/advanced16-5.pdf.

Whitney:1978:APSd

[Whi78d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 16(6):566–569, December 1978. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/16-6/advanced16-6.pdf.

Whitney:1978:GSI

[Whi78e] Raymond E. Whitney. Geometric sequences and the initial digit problem. Fibonacci Quarterly, 16(2):152–153, April 1978. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/16-2/whitney.pdf.

Whitney:1979:APSa

[Whi79a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 17(1):94–96, February 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/17-1/advanced17-1.pdf.

Whitney:1979:APSb

[Whi79b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 17(2):189–192, April 1979. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/17-2/
advanced17-2.pdf.

Whitney:1979:APSc

[Whi79c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 17(3):286–288, October 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/17-3/advanced17-3.pdf.

Whitney:1979:APSd

[Whi79d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 17(4):374–378, December 1979. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/17-4/advanced17-4.pdf.

Whitney:1980:APSa

[Whi80a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 18(1):90–96, February 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/18-1/advanced18-1.pdf.

Whitney:1980:APSb

[Whi80b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 18(2):190–192, April 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-2/
advanced18-2.pdf.

Whitney:1980:APSc

[Whi80c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 18(3):280–288, October 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/18-3/advanced18-3.pdf.

Whitney:1980:APSd

[Whi80d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 18(4):375–381, December 1980. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/18-4/advanced18-4.pdf.

Whitney:1981:APSa

[Whi81a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 19(1):93–96, February 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-1/advanced19-1.pdf.

Whitney:1981:APSb

[Whi81b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 19(2):189–192, April 1981. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/19-2/
advanced19-2.pdf.

Whitney:1981:APSc

[Whi81c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 19(4):381–384, October 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-4/advanced19-4.pdf.

Whitney:1981:APSd

[Whi81d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 19(5):470–476, December 1981. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/19-5/advanced19-5.pdf.

Whitney:1982:APSa

[Whi82a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 20(1):93–96, February 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/20-1/advanced20-1.pdf.

Whitney:1982:APSb

[Whi82b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 20(2):185–191, May 1982. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/20-2/
advanced20-2.pdf.

Whitney:1982:APSc

[Whi82c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 20(3):284–288, August 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/20-3/advanced20-3.pdf.

Whitney:1982:APSd

[Whi82d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 20(4):372–380, November 1982. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/20-4/advanced20-4.pdf.

Whitney:1983:APSa

[Whi83a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 21(1):74–81, February 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/21-1/advanced21-1.pdf.

Whitney:1983:APSb

[Whi83b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 21(2):153–161, May 1983. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/21-2/
advanced21-2.pdf.

Whitney:1983:APSc

[Whi83c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 21(3):236–241, August 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/21-3/advanced21-3.pdf.

Whitney:1983:APSd

[Whi83d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 21(4):312–317, November 1983. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/21-4/advanced21-4.pdf.

Whitney:1984:APSa

[Whi84a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 22(1):89–97, February 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/22-1/advanced22-1.pdf.

Whitney:1984:APSb

[Whi84b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 22(2):188–193, May 1984. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/22-2/
advanced22-2.pdf.

Whitney:1984:APSc

[Whi84c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 22(3):279–289, August 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/22-3/advanced22-3.pdf.

Whitney:1984:APSd

[Whi84d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 22(4):374–381, November 1984. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/22-4/advanced22-4.pdf.

Whitney:1985:APSa

[Whi85a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 23(1):89–97, February 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/23-1/advanced23-1.pdf.

Whitney:1985:APSb

[Whi85b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 23(2):186–193, May 1985. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/23-2/
advanced23-2.pdf.

Whitney:1985:APSc

[Whi85c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 23(3):282–289, August 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/23-3/advanced23-3.pdf.

Whitney:1985:APSd

[Whi85d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 23(4):376–382, November 1985. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/23-4/advanced23-4.pdf.

White:1986:CRR

[Whi86a] Tad White. On the coefficients of a recursion relation for the Fibonacci partition function. Fibonacci Quarterly, 24(2):133–137, May 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/24-2/white.pdf.

Whitney:1986:APSa

[Whi86b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 24(1):88–97, February 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/24-1/advanced24-1.pdf.

Whitney:1986:APSb

[Whi86c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 24(2):185–193, May 1986. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/24-2/
advanced24-2.pdf.

Whitney:1986:APSc

[Whi86d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 24(3):283–289, August 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/24-3/advanced24-3.pdf.

Whitney:1986:APSd

[Whi86e] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 24(4):376–381, November 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/24-4/advanced24-4.pdf.

Whitney:1987:APSa

[Whi87a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 25(1):90–97, February 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-1/advanced25-1.pdf.

Whitney:1987:APSb

[Whi87b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 25(2):185–193, May 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-2/
advanced25-2.pdf.

Whitney:1987:APSc

[Whi87c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 25(3):284–289, August 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-3/advanced25-3.pdf.

Whitney:1987:APSd

[Whi87d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 25(4):376–382, November 1987. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/25-4/advanced25-4.pdf.

Whitney:1988:APSa

[Whi88a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 26(1):89–97, February 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/26-1/advanced26-1.pdf.

Whitney:1988:APSb

[Whi88b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 26(2):186–193, May 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-2/
advanced26-2.pdf.

Whitney:1988:APSc

[Whi88c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 26(3):283–288, August 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/26-3/advanced26-3.pdf.

Whitney:1988:APSd

[Whi88d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 26(4):377–381, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/26-4/advanced26-4.pdf.

Whitney:1989:APSa

[Whi89a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 27(1):92–97, February 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/27-1/advanced27-1.pdf.

Whitney:1989:APSb

[Whi89b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 27(2):187–193, May 1989. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/27-2/
advanced27-2.pdf.

Whitney:1989:APSc

[Whi89c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 27(4):378–385, August 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/27-4/advanced27-4.pdf.

Whitney:1989:APSd

[Whi89d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 27(5):473–478, November 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/27-5/advanced27-5.pdf.

Whitney:1990:APSa

[Whi90a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 28(1):91–97, February 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/28-1/advanced28-1.pdf.

Whitney:1990:APSb

[Whi90b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 28(2):187–191, May 1990. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/28-2/
advanced28-2.pdf.

Whitney:1990:APSc

[Whi90c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 28(3):283–289, August 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/28-3/advanced28-3-a.pdf.

Whitney:1990:APSd

[Whi90d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 28(4):376–382, November 1990. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/28-4/advanced28-4.pdf.

Whitney:1991:APSa

[Whi91a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 29(1):89–97, February 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/29-1/advanced29-1.pdf.

Whitney:1991:APSb

[Whi91b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 29(2):186–193, May 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-2/
advanced29-2.pdf.

Whitney:1991:APSc

[Whi91c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 29(3):283–289, August 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/29-3/advanced29-3.pdf.

Whitney:1991:APSd

[Whi91d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 29(4):377–382, November 1991. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/29-4/advanced29-4.pdf.

Whitney:1992:APSa

[Whi92a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 30(1):90–97, February 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-1/advanced30-1.pdf.

Whitney:1992:APSb

[Whi92b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 30(2):187–193, May 1992. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/30-2/
advanced30-2.pdf.

Whitney:1992:APSc

[Whi92c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 30(3):282–289, August 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-3/advanced30-3.pdf.

Whitney:1992:APSd

[Whi92d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 30(4):376–382, November 1992. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/30-4/advanced30-4.pdf.

Whitney:1993:APSa

[Whi93a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 31(1):90–97, February 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-1/advanced31-1.pdf.

Whitney:1993:APSb

[Whi93b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 31(2):187–193, May 1993. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/31-2/
advanced31-2.pdf.

Whitney:1993:APSc

[Whi93c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 31(3):283–289, August 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-3/advanced31-3.pdf.

Whitney:1993:APSd

[Whi93d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 31(4):377–382, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-4/advanced31-4.pdf.

Whitney:1994:APSa

[Whi94a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 32(1):91–97, February 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-1/advanced32-1.pdf.

Whitney:1994:APSb

[Whi94b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 32(2):187–193, May 1994. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/32-2/
advanced32-2.pdf.

Whitney:1994:APSc

[Whi94c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 32(4):379–385, August 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-4/advanced32-4.pdf.

Whitney:1994:APSd

[Whi94d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 32(5):473–478, November 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/32-5/advanced32-5.pdf.

Whitney:1995:APSa

[Whi95a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 33(1):91–97, February 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-1/advanced33-1.pdf.

Whitney:1995:APSb

[Whi95b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 33(2):187–193, May 1995. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/33-2/
advanced33-2.pdf.

Whitney:1995:APSc

[Whi95c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 33(4):378–385, August 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-4/advanced33-4.pdf.

Whitney:1995:APSd

[Whi95d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 33(5):472–477, November 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-5/advanced33-5.pdf.

Whitney:1996:APSa

[Whi96a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 34(1):89–97, February 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-1/advanced34-1.pdf.

Whitney:1996:APSb

[Whi96b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 34(2):187–193, May 1996. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/34-2/
advanced34-2.pdf.

Whitney:1996:APSc

[Whi96c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 34(4):379–385, August 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-4/advanced34-4.pdf.

Whitney:1996:APSd

[Whi96d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 34(5):473–478, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-5/advanced34-5.pdf.

Whitney:1997:APSa

[Whi97a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 35(1):91–97, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-1/advanced35-1.pdf.

Whitney:1997:APSb

[Whi97b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 35(2):187–193, May 1997. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/35-2/
advanced35-2.pdf.

Whitney:1997:APSc

[Whi97c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 35(3):283–289, August 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-3/advanced35-3.pdf.

Whitney:1997:APSd

[Whi97d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 35(4):377–382, November 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-4/advanced35-4.pdf.

Whitney:1998:APSa

[Whi98a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 36(1):91–97, February 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-1/advanced36-1.pdf.

Whitney:1998:APSb

[Whi98b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 36(2):187–193, May 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-2/
advanced36-2.pdf.

Whitney:1998:APSc

[Whi98c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 36(4):379–385, August 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-4/advanced36-4.pdf.

Whitney:1998:APSd

[Whi98d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 36(5):473–??, November 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-5/advanced36-5.pdf.

Whitney:1999:APSa

[Whi99a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 37(1):91–97, February 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-1/advanced37-1.pdf.

Whitney:1999:APSb

[Whi99b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 37(2):185–193, May 1999. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/37-2/
advanced37-2.pdf.

Whitney:1999:APSc

[Whi99c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 37(3):282–289, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-3/advanced37-3.pdf.

Whitney:1999:APSd

[Whi99d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 37(4):377–382, November 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/37-4/advanced37-4.pdf.

Whitney:2000:APSa

[Whi00a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 38(1):91–97, February 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/38-1/advanced38-1.pdf.

Whitney:2000:APSb

[Whi00b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 38(2):186–193, May 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-2/
advanced38-2.pdf.

Whitney:2000:APSc

[Whi00c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 38(4):377–385, August 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/38-4/advanced38-4.pdf.

Whitney:2000:APSd

[Whi00d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 38(5):473–478, November 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/38-5/advanced38-5.pdf.

Whitney:2001:APSa

[Whi01a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 39(1):91–97, February 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/39-1/advanced39-1.pdf.

Whitney:2001:APSb

[Whi01b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 39(2):187–193, May 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-2/
advanced39-2.pdf.

Whitney:2001:APSc

[Whi01c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 39(4):378–383, August 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/39-4/advanced39-4.pdf.

Whitney:2001:APSd

[Whi01d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 39(5):473–477, November 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/39-5/advanced39-5.pdf.

Whitney:2002:APSa

[Whi02a] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 40(1):91–97, February 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-1/advanced40-1.pdf.

Whitney:2002:APSb

[Whi02b] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 40(2):187–193, May 2002. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/40-2/
advanced40-2.pdf.

Whitney:2002:APSc

[Whi02c] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 40(4):379–385, August 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-4/advanced40-4.pdf.

Whitney:2002:APSd

[Whi02d] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 40(5):472–478, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-5/advanced40-5.pdf.

Whitney:2003:APS

[Whi03] Raymond E. Whitney. Advanced problems and solutions. Fibonacci Quarterly, 41(1):91–97, February 2003. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/41-1/advanced41-1.pdf.

WiEckowski:1980:SSD

[Wiȩ80] Andrzej Wiȩckowski. On some systems of Diophantine equations including the algebraic sum of triangular numbers. Fibonacci Quarterly, 18(2):165–169, April 1980. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/18-2/
wieckowski.pdf.

Wiedemann:1988:IQE

[Wie88] Doug Wiedemann. An iterated quadratic extension of GF(2). Fibonacci Quarterly, 26(4):290–295, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/26-4/wiedemann.pdf.

Williams:1970:PT

[Wil70] J. M. Williams, Jr. The powers of three. Fibonacci Quarterly, 8(5):509–515, December 1970. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/8-5/williams.
pdf.

Williams:1972:FNO

[Wil72] H. C. Williams. Fibonacci numbers obtained from Pascal’s triangle with generalizations. Fibonacci Quarterly, 10(4):405–412, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/10-4/williams.pdf.

Williams:1975:FNF

[Wil75a] H. C. Williams. On Fibonacci numbers of the form k^{2} + 1. Fibonacci Quarterly, 13(3):213–214, October 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-3/williams1.pdf.

Williams:1975:RAG

[Wil75b] H. C. Williams. The rank of apparition of a generalized Fibonacci sequences. Fibonacci Quarterly, 13(3):240–242, October 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-3/williams2.pdf.

Willett:1976:TK

[Wil76] Michael Willett. On a theorem of Kronecker. Fibonacci Quarterly, 14(1):27–28, February 1976. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/14-1/willett.
pdf.

Williams:1977:PSF

[Wil77] H. C. Williams. Properties of some functions similar to Lucas functions. Fibonacci Quarterly, 15(2):97–111, April 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-2/williams.pdf.

Williams:1982:NFQ

[Wil82] H. C. Williams. A note on the Fibonacci quotient F_{p−ε}/p. Canadian Mathematical Bulletin = Bulletin canadien de mathématiques, 25(??):366–370, ???? 1982. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Williams:1983:NLU

[Wil83] M. R. Williams. From Napier to Lucas: The use of Napier’s bones in calculating instruments. Annals of the History of Computing, 5(3):279–296, July/September 1983. CODEN AH-

COE5. ISSN 0164-1239. URL http://dlib.computer.org/
an/books/an1983/pdf/a3279.pdf;   http://www.computer.
org/annals/an1983/a3279abs.htm.

Wilcox:1986:FSP

[Wil86] Howard J. Wilcox. Fibonacci sequences of period n in groups. Fibonacci Quarterly, 24(4):356–361, November 1986. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/24-4/wilcox.pdf.

Williams:1987:EPT

[Wil87] H. C. Williams. Effective primality tests for some integers of the forms A5^{n} − 1 and A7^{n} − 1. Mathematics of Computation, 48 (177):385–403, January 1987. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Williams:1988:NP

[Wil88] H. C. Williams. A note on the primality of 6^{2^{n}} + 1 and 10^{2^{n}} + 1. Fibonacci Quarterly, 26(4):296–305, November 1988. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/26-4/williams.pdf.

Wilson:1996:CSC

[Wil96] Brad Wilson. Construction of small consecutive Niven numbers. Fibonacci Quarterly, 34(3):240–243, June 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/34-3/wilson.pdf.

Wilson:1997:CCN

[Wil97] Brad Wilson. Construction of 2 ∗ n consecutive n-Niven numbers. Fibonacci Quarterly, 35(2):122–128, May 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-2/wilson.pdf.

Williams:1998:ELP

[Wil98a] Hugh C. Williams. Édouard Lucas and primality testing, volume 22 of Canadian Mathematical Society series of monographs and advanced texts. John Wiley, New York, NY, USA, 1998. ISBN 0-471-14852-0 (hardcover). xviii + 525 pp.

LCCN QA246 .W43 1998. URL http://www.loc.gov/catdir/
description/wiley033/97044760.html;   http://www.loc.
gov/catdir/toc/onix04/97044760.html;   https://www.math.
utah.edu/pub/tex/bib/fibquart.bib.

Wilson:1998:AP

[Wil98b] Brad Wilson. Acknowledgment of priority. Fibonacci Quarterly, 36(5):395–??, November 1998. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/36-5/priority.
pdf.

Wilson:1998:FTM

[Wil98c] Brad Wilson. The Fibonacci triangle modulo p. Fibonacci Quarterly, 36(3):194–203, June 1998. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/36-3/wilson1.
pdf.

Wilson:1998:PDM

[Wil98d] Brad Wilson. Power digraphs modulo n. Fibonacci Quarterly, 36(3):229–239, June 1998. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/36-3/wilson2.pdf.

Williams:2000:NTF

[Wil00] H. C. Williams. A number theoretic function arising from continued fractions. Fibonacci Quarterly, 38(3):201–211, June 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/38-3/williams.pdf.

Winthrop:1965:CTG

[Win65a] Henry Winthrop. Comments on “Time Generated, Compositions Yield Fibonacci Numbers”. Fibonacci Quarterly, 3(3):234–??, October 1965. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/3-3/winthrop-a.pdf.

Winthrop:1965:TGC

[Win65b] Henry Winthrop. Time generated compositions yield Fibonacci numbers. Fibonacci Quarterly, 3(2):131–134, April 1965. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/3-2/winthrop-a.pdf.

Winthrop:1968:MMS

[Win68] Henry Winthrop. Mathematical models for the study of the propagation of novel social behavior. Fibonacci Quarterly, 6(2):151– 155, April 1968. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/6-2/winthrop.pdf.

Winkler:2011:LBPe

[Win11] Peter Winkler. Last byte: Puzzled: Uncommon divisors. Communications of the Association for Computing Machinery, 54(8): 120, August 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wirth:2008:IRP

[Wir08] Michael Wirth. Introducing recursion by parking cars. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 40(4):52–55, December 2008. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Witzgall:1972:FSA

[Wit72] Christoph Witzgall. Fibonacci search with arbitrary first evaluation. Fibonacci Quarterly, 10(2):113–134, February 1972. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/10-2/witzgall-a.pdf.

Witno:2016:NRG

[Wit16] Amin Witno. Niven repunits in general bases. Fibonacci Quarterly, 54(1):59–??, February 2016. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Abstracts/54-1/witno.
pdf.

Williams:1976:DPU

[WJ76] H. C. Williams and J. S. Judd. Determination of the primality of N by using factors of N^{2} ± 1. Mathematics of Computation, 30(133):157–172, January 1976. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Woltman:2008:DKM

[WK09] George Woltman and Scott Kurowski. On the discovery of the 45th and 46th known Mersenne primes. Fibonacci Quarterly, 46/47(3):194–197, August 2008/2009. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/46_47-
3/woltman.pdf.

Wlodarski:1963:GRF

[Wlo63] J. Wlodarski. The “Golden Ratio” and the Fibonacci numbers in the world of atoms. Fibonacci Quarterly, 1(4):61–64, December 1963. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/1-4/wlodarski.pdf.

Wlodarski:1965:FNM

[Wlo65] J. Wlodarski. The Fibonacci numbers and the “magic” numbers. Fibonacci Quarterly, 3(3):208–??, October 1965. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/3-3/wlodarski.pdf.

Wlodarski:1967:AGR

[Wlo67] J. Wlodarski. Achieving the “Golden Ratio” by grouping the “elementary” principle. Fibonacci Quarterly, 5(2):193–194, April 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/5-2/wlodarski.pdf.

Wlodarski:1968:CPO

[Wlo68a] J. Wlodarski. Curious property of one fraction. Fibonacci Quarterly, 6(2):156–??, April 1968. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/6-2/
wlodarski1.pdf.

Wlodarski:1968:MAG

[Wlo68b] J. Wlodarski. More about the “Golden Ratio” in the world of atoms. Fibonacci Quarterly, 6(4):244–??, October 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-4/wlodarski-a.pdf.

Wlodarski:1968:NP

[Wlo68c] J. Wlodarski. A number problem. Fibonacci Quarterly, 6(2): 161–??, April 1968. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/6-2/wlodarski2-a.pdf.

Wlodarski:1968:PTS

[Wlo68d] J. Wlodarski. Pascal’s triangle and some famous number sequences. Fibonacci Quarterly, 6(2):192–??, April 1968. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/6-2/wlodarski3.pdf.

Wlodarski:1971:FLN

[Wlo71a] J. Wlodarski. Fibonacci and Lucas numbers tend to obey Benford’s Law. Fibonacci Quarterly, 9(1):87–88, February 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-1/wlodarski2.pdf.

Wlodarski:1971:GRE

[Wlo71b] J. Wlodarski. The golden ratio in an electrical network. Fibonacci Quarterly, 9(2):188–??, April 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-2/
wlodarski2-a.pdf.

Wlodarski:1971:NP

[Wlo71c] J. Wlodarski. A number problem. Fibonacci Quarterly, 9(2): 195–??, April 1971. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/9-2/wlodarski3-a.pdf.

Wlodarski:1971:PEP

[Wlo71d] J. Wlodarski. The possible end of the periodic table of elements and the “Golden Ratio”. Fibonacci Quarterly, 9(1):82–??, February 1971. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/9-1/wlodarski1-a.pdf.

Wlodarski:1971:RPP

[Wlo71e] J. Wlodarski. Regular polyhedrons and Pascal’s triangle. Fibonacci Quarterly, 9(2):146–??, April 1971. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/9-2/
wlodarski1.pdf.

Wlodarski:1972:NG

[Wlo72] J. Wlodarski. A number game. Fibonacci Quarterly, 10(3):301– 302, April 1972. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/10-3/wlodarski.pdf.

Wlodarski:1973:BSF

[Wlo73] J. Wlodarski. The Balmer series and the Fibonacci numbers. Fibonacci Quarterly, 11(5):526–??, December 1973. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/11-5/wlodarski-a.pdf.

Wloch:2013:SIG

[Wlo13] Andrzej Wloch. Some identities for the generalized Fibonacci numbers and the generalized Lucas numbers. Applied Mathematics and Computation, 219(10):5564–5568, January 15, 2013. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300312011976.

Wong:1975:GPT

[WM75] C. K. Wong and T. W. Maddocks. A generalized Pascal’s triangle. Fibonacci Quarterly, 13(2):134–136, April 1975. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-2/wong.pdf.

Woan:2002:CPR

[Woa02] Wen-Jin Woan. A combinatorial proof of a recursive relation of the Motzkin sequence by lattice paths. Fibonacci Quarterly, 40(1):3–8, February 2002. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/40-1/woan.pdf.

Wollan:1992:PRN

[Wol92] Peter C. Wollan. A portable random number generator for parallel computers. Communications in Statistics: Simulation and Computation, 21(4):1247–1254, ???? 1992. CODEN CSSCDB. ISSN 0361-0918.

Wolfram:1998:SGF

[Wol98] D. A. Wolfram. Solving generalized Fibonacci recurrences. Fibonacci Quarterly, 36(2):129–145, May 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-2/
wolfram.pdf.

Woltman:1999:DKM

[Wol99] George Woltman. On the discovery of the 38 th known Mersenne prime. Fibonacci Quarterly, 37(4):367–370, November 1999. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/37-4/woltman.pdf.

Wong:1982:DP

[Won82] Fook-Bun Wong. Ducci processes. Fibonacci Quarterly, 20(2): 97–105, May 1982. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/20-2/wong.pdf.

Woo:1975:NBT

[Woo75] Norman Woo. Non-basic triples. Fibonacci Quarterly, 13(1):56– 58, February 1975. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/13-1/woo.pdf.

Woo:1979:NBT

[Woo79] Norman Woo. A note on basic M-tuples. Fibonacci Quarterly, 17(2):165–167, April 1979. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/17-2/woo.pdf.

Woodson:1997:PLV

[Woo97] Leon C. Woodson. Professor Lucas visits the Putnam Examination. Fibonacci Quarterly, 35(4):341–??, November 1997. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-4/woodson.pdf.

Wood:2007:BPF

[Woo07] Philip Matchett Wood. Bijective proofs for Fibonacci identities related to Zeckendorf’s theorem. Fibonacci Quarterly, 45(2):138– 145, May 2007. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Abstracts/45-2/wood.pdf.

Webb:1969:DPF

[WP69] W. A. Webb and E. A. Parberry. Divisibility properties of Fibonacci polynomials. Fibonacci Quarterly, 7(5):457–463, December 1969. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/7-5/webb.pdf.

Wu:2007:SPB

[WP07] Ming Wu and Hao Pan. Sums of products of Bernoulli numbers of the second kind. Fibonacci Quarterly, 45(2):146–150, May 2007.

CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/45-2/wu.pdf.

Wadsanthat:2019:DCL

[WP19] Atsanon Wadsanthat and Chatchawan Panraksa. Distribution of cycle lengths of a quadratic map over finite fields of characteristic 2. Fibonacci Quarterly, 57(1):35–??, February 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-1/wadsanthat.pdf.

Wahbi:2001:GFS

[WR01] Bouazza Ell Wahbi and Mustapha Rachidi. On r-generalized Fibonacci sequences and Hausdorff moment problems. Fibonacci Quarterly, 39(1):5–11, February 2001. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/39-1/el-
wahbi.pdf.

Waddill:1967:AGF

[WS67] Marcellus E. Waddill and Louis Sacks. Another generalized Fibonacci sequence. Fibonacci Quarterly, 5(3):209–222, October 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/5-3/waddill.pdf.

Wilson:1980:ACG

[WS80] Thomas C. Wilson and Joseph Shortt. An O(log n) algorithm for computing general order-k Fibonacci numbers. Information Processing Letters, 10(2):68–75, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Whitaker:1991:SRA

[WS91] R. N. Whitaker and A. G. Shannon. Some recursive asymptotes. Fibonacci Quarterly, 29(3):235–238, August 1991. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/29-3/whitaker.pdf.

Wunderlich:1974:ADE

[WSH74] F. J. Wunderlich, D. E. Shaw, and M. J. Hones. Argand diagrams of extended Fibonacci and Lucas numbers. Fibonacci Quarterly, 12(3):233–234, October 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
3/wunderlich.pdf.

Wu:2004:SIB

[WSP04] Ke-Jian Wu, Zhi-Wei Sun, and Hao Pan. Some identities for Bernoulli and Euler polynomials. Fibonacci Quarterly, 42(4):295–299, November 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
4/quartsun04_2004.pdf.

Winley:1987:PNE

[WTvR87] Graham Winley, Keith Tognetti, and Tony van Ravenstein. A property of numbers equivalent to the golden mean. Fibonacci Quarterly, 25(2):171–173, May 1987. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/25-2/
winley.pdf.

Winley:1989:HTC

[WTvR89] Graham Winley, Keith Tognetti, and Tony van Ravenstein. Hurwitz’s theorem and the continued fraction with constant terms. Fibonacci Quarterly, 27(5):420–423, November 1989. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/27-5/winley.pdf.

Wu:1983:CPD

[Wu83] T. C. Wu. Counting the profiles in domino tiling. Fibonacci Quarterly, 21(4):302–303, November 1983. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/21-
4/wu.pdf.

Wu:1997:EFC

[Wu97] Jie Wu. Extended Fibonacci cubes. IEEE Transactions on Parallel and Distributed Systems, 8(12):1203–1210, December 1997. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (elec-

tronic). URL http://dlib.computer.org/td/books/td1997/
pdf/l1203.pdf;   http://www.computer.org/tpds/td1997/
l1203abs.htm.

Wulczyn:1975:MSX

[Wul75] Gregory Wulczyn. Minimum solutions to x^{2} − Dy^{2} = ±1. Fibonacci Quarterly, 13(4):309–311, December 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/13-4/wulczyn.pdf.

Wulczyn:1976:CFE

[Wul76] Gregory Wulczyn. On continued fraction expansions whose elements are all ones. Fibonacci Quarterly, 14(1):18–22, February 1976. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/14-1/wulczyn.pdf.

Wulczyn:1981:GES

[Wul81] Gregory Wulczyn. A generalized extension of some Fibonacci– Lucas identities to primitive unit identities. Fibonacci Quarterly, 19(5):385–390, December 1981. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/19-5/wulczyn.
pdf.

Wunderlich:1963:NEF

[Wun63] M. Wunderlich. On the non-existence of Fibonacci squares (in Technical Notes and Short Papers). Mathematics of Computation, 17(84):455–457, October 1963. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Wegener:1975:PTA

[WW75] Delano P. Wegener and Joseph A. Wehlen. Pythagorean triangles — abstract. Fibonacci Quarterly, 13(2):110–??, April 1975. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/13-2/wegener-a.pdf.

Wang:2006:SRP

[WW06] Jun Wang and Xin Wang. On the set of reduced ϕ-partitions of a positive integer. Fibonacci Quarterly, 44(2):98–102, May 2006. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/44-2/wang.pdf.

Wang:2017:GHP

[WW17] Weiping Wang and Hui Wang. Generalized Humbert polynomials via generalized Fibonacci polynomials. Applied Mathematics and Computation, 307(??):204–216, August 15, 2017. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300317301613.

Wenpeng:2000:FND

[WY00] Zhang Wenpeng and Yi Yuan. On the Fibonacci numbers and the Dedekind sums. Fibonacci Quarterly, 38(3):223–226, June 2000. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/38-3/zhang.pdf.

Xie:2002:SGA

[Xie02] Ziqing Xie. On the summation of generalized arithmeticgeometric trigonometric series. Fibonacci Quarterly, 40(2):128– 135, May 2002. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/40-2/xie.pdf.

Xiong:1999:LPK

[XLP99] Jin Tao Xiong, Hong Xiu Liu, and De Zhong Pi. The Lucas public-key cryptosystem and its security. Dianzi Keji Daxue Xuebao, 28(4):397–401, 1999. CODEN DKDAEM. ISSN 1001-0548.

Xekalaki:1987:SMD

[XPP87] Evdokia Xekalaki, John Panaretos, and Andreas Philippou. On some mixtures of distributions of order k. Fibonacci Quarterly, 25(2):151–160, May 1987. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/25-2/xekalaki.pdf.

Yabuta:2001:SPC

[Yab01] Minoru Yabuta. A simple proof of Carmichael’s theorem on primitive divisors. Fibonacci Quarterly, 39(5):439–443, November 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/39-5/yabuta.pdf.

Yabuta:2002:PSL

[Yab02] Minoru Yabuta. Perfect squares in the Lucas numbers. Fibonacci Quarterly, 40(5):460–466, November 2002. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/40-
5/yabuta.pdf.

Yabuta:2007:ACP

[Yab07] Minoru Yabuta. The ABC-conjecture and the powerful numbers in Lucas sequences. Fibonacci Quarterly, 45(4):362–365, November 2007. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Abstracts/45-4/yabuta.pdf.

Yalavigi:1972:PTN

[Yal72] C. C. Yalavigi. Properties of Tribonacci numbers. Fibonacci Quarterly, 10(3):231–246, April 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
3/yalavigi.pdf.

Yalavigi:1973:PST

[Yal73] C. C. Yalavigi. Periodicity of second- and third-order recurring sequences. Fibonacci Quarterly, 11(2):163–165, April 1973. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/11-2/yalavigi.pdf.

Yalavigi:1977:RGF

[Yal77] C. C. Yalavigi. Residues of generalized Fibonacci sequences. Fibonacci Quarterly, 15(1):1–??, February 1977. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/15-1/
yalavigi.pdf.

Yamada:1980:CPA

[Yam80] Masaji Yamada. A convergence proof about an integral sequence. Fibonacci Quarterly, 18(3):231–241, October 1980. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/18-3/yamada.pdf.

Yamagishi:2019:SPC

[Yam19] Masakazu Yamagishi. A short proof of congruences for Lucas sequences. Fibonacci Quarterly, 57(3):260–??, August 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-3/yamagishi.pdf.

Yang:1988:LPC

[Yan88] Kung-Wei Yang. Limits of q-polynomial coefficients. Fibonacci Quarterly, 26(1):64–69, February 1988. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/26-
1/yang.pdf.

Yang:1991:DP

[Yan91] Kung-Wei Yang. q-determinants and permutations. Fibonacci Quarterly, 29(2):160–163, May 1991. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/29-2/
yang.pdf.

Yan:1995:PTL

[Yan95] S. Y. Yan. Primality testing of large numbers in Maple. Computers and Mathematics and Applications, 29(12):1–8, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Yap:1970:SRR

[Yap70] H. P. Yap. A simple recurrence relation on finite Abelian groups. Fibonacci Quarterly, 8(3):255–263, April 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-3/yap.pdf.

Yayenie:2011:NGF

[Yay11] Omer Yayenie. A note on generalized Fibonacci sequences. Applied Mathematics and Computation, 217(12):5603–5611, February 15, 2011. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Yang:2018:PRR

[YG18] Sheng-Liang Yang and Yuan-Yuan Gao. The Pascal rhombus and Riordan arrays. Fibonacci Quarterly, 56(4):337–??, November 2018. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Abstracts/56-4/yang.pdf.

Yalavigi:1977:PLG

[YK77] C. C. Yalavigi and H. V. Krishna. Periodic lengths of the generalized Fibonacci sequence modulo p. Fibonacci Quarterly, 15(2): 150–152, April 1977. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/15-2/yalavigi.pdf.

Yildiz:2003:FSM

[YK03] Bünyamin Yildiz and Erdal Karaduman. On Fibonacci search method with k-Lucas numbers. Applied Mathematics and Computation, 143(2–3):523–531, November 10, 2003. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). See comment [Kah05].

Yu:1997:IIP

[YL97] Hongquan Yu and Chuanguang Liang. Identities involving partial derivatives of bivariate Fibonacci and Lucas polynomials. Fibonacci Quarterly, 35(1):19–23, February 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/35-1/yu.pdf.

Yang:2004:PDG

[YL04] Yongzhi (Peter) Yang and Johann Leida. Pascal decompositions of geometric arrays in matrices. Fibonacci Quarterly, 42(3): 205–215, August 2004. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers1/42-3/quartyang03_
2004.pdf.

Yoon:2021:ACB

[Yoo21] Soowhan Yoon. Analytic connection between the Fibonacci sequence and diagonal sums of binomial coefficients. Fibonacci Quarterly, 59(4):349–??, November 2021. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/59-
4/yoon.pdf.

Yordzhev:2014:FSB

[Yor14] Krasimir Yordzhev. Factor-set of binary matrices and Fibonacci numbers. Applied Mathematics and Computation, 236(??):235– 238, June 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300314004354.

Young:1992:RM

[You92a] Anne Ludington Young. k-reverse multiples. Fibonacci Quarterly, 30(2):126–131, May 1992. CODEN FIBQAU. ISSN

0015-0517.  URL http://www.fq.math.ca/Scanned/30-2/
ludington1.pdf.

Young:1992:TRM

[You92b] Anne Ludington Young. Trees for k-reverse multiples. Fibonacci Quarterly, 30(2):166–174, May 1992. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/30-2/
ludington2.pdf.

Young:1993:VTD

[You93] Anne Ludington Young. A variation on the two-digit Kaprekar routine. Fibonacci Quarterly, 31(2):138–144, May 1993. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-2/young.pdf.

Young:1994:ACG

[You94] Paul Thomas Young. p-adic congruences for generalized Fibonacci sequences. Fibonacci Quarterly, 32(1):2–10, February 1994. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/32-1/young.pdf.

Young:1995:SSR

[You95a] Anne Ludington Young. The switch, subtract, reorder routine. Fibonacci Quarterly, 33(5):432–440, November 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-5/ludington.pdf.

Young:1995:QRL

[You95b] Paul Thomas Young. Quadratic reciprocity via Lucas sequences. Fibonacci Quarterly, 33(1):78–81, February 1995. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/33-1/young.pdf.

Young:1998:DPT

[You98] Anne Ludington Young. Ducci-processes of 5-tuples. Fibonacci Quarterly, 36(5):419–433, November 1998. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/36-
5/ludington.pdf.

Young:2002:MDP

[You02] Paul Thomas Young. On modified Dickson polynomials. Fibonacci Quarterly, 40(1):33–40, February 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-1/young.pdf.

Young:2003:LR

[You03] Paul Thomas Young. On lacunary recurrences. Fibonacci Quarterly, 41(1):41–47, February 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-
1/young.pdf.

Young:2006:LBN

[You06] Paul Thomas Young. On Lucas–Bernoulli numbers. Fibonacci Quarterly, 44(4):347–357, November 2006. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/44-
4/young.pdf.

Young:2014:SSN

[You14] Paul Thomas Young. Symmetries of Stirling number series. Fibonacci Quarterly, 52(5):205–??, December 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Papers1/52-5/Young.pdf.

Young:2017:CBL

[You17] Paul Thomas Young. Congruences for Bernoulli–Lucas sums. Fibonacci Quarterly, 55(5):??, December 2017. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/55-
5/young.pdf.

Young:2019:GSZ

[You19] Paul Thomas Young. Global series for zeta functions. Fibonacci Quarterly, 57(5):154–??, December 2019. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Abstracts/57-
5/young.pdf.

Yamagishi:2008:NFH

[YS08] Michel E. Beleza Yamagishi and Alex Itiro Shimabukuro. Nucleotide frequencies in human genome and Fibonacci numbers. Bulletin of Mathematical Biology, 70(3):643–653, April 2008. CO-DEN BMTBAP. ISSN 0092-8240 (print), 1522-9602 (electronic).

URL http://link.springer.com/article/10.1007/s11538-
007-9261-6;   http://link.springer.com/content/pdf/10.
1007/s11538-007-9261-6.pdf.

Yu:1998:GSN

[Yu98] Hongquan Yu. A generalization of Stirling numbers. Fibonacci Quarterly, 36(3):252–258, June 1998. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Scanned/36-3/yu.
pdf.

Yuan:2002:NDG

[Yua02] Pingzhi Yuan. A note on the divisibility of the generalized Lucas sequences. Fibonacci Quarterly, 40(2):153–156, May 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-2/yuan.pdf.

Yu:1996:LGG

[YWH96] Hongquan Yu, Yi Wang, and Mingfeng He. On the limit of generalized golden numbers. Fibonacci Quarterly, 34(4):320–322, August 1996. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/34-4/yu.pdf.

Yang:2019:PPR

[YY19] Lin Yang and Sheng-Liang Yang. The parametric Pascal rhombus. Fibonacci Quarterly, 57(4):337–??, November 2019. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/57-4/yang.pdf.

Yuan:2002:SII

[YZ02] Yi Yuan and Wenpeng Zhang. Some identities involving the Fibonacci polynomials. Fibonacci Quarterly, 40(4):314–318, August 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/40-4/yuan.pdf.

Ye:2017:CGC

[YZ17] Xiaoli Ye and Zhizheng Zhang. A common generalization of convolved generalized Fibonacci and Lucas polynomials and its applications. Applied Mathematics and Computation, 306(??):31– 37, August 1, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300317301200.

Yang:2018:SIG

[YZ18] Jizhen Yang and Zhizheng Zhang. Some identities of the generalized Fibonacci and Lucas sequences. Applied Mathematics and Computation, 339(??):441–450, December 15, 2018. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300318306209.

Zaheer:1982:LSV

[Zah82] Neyamat Zaheer. On Lucas-sets for vector-valued abstract polynomials in K-inner product spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 34(??):832–852, ???? 1982. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zaks:1983:GPN

[Zak83] Shmuel Zaks. Generalized profile numbers. Fibonacci Quarterly, 21(1):58–61, February 1983. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/21-1/zaks.pdf.

Zaremba:1970:RLG

[Zar70] S. K. Zaremba. A remarkable lattice generated by Fibonacci numbers. Fibonacci Quarterly, 8(2):185–198, March 1970. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/8-2/zaremba.pdf.

Zarzycki:1997:MS

[Zar97] Piotr Zarzycki. On multiplicity sequences. Fibonacci Quarterly, 35(1):9–??, February 1997. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/35-1/zarzycki.

pdf. See corrigendum [Zar98].

Zarzycki:1998:CMS

[Zar98] Piotr Zarzycki. Corrigendum to “On Multiplicity Sequences”. Fibonacci Quarterly, 36(5):434–??, November 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-5/corrigendum.pdf. See [Zar97].

Zarnowski:2008:CSM

[Zar09] Roger E. Zarnowski. The congruence structure of the 3x + 1 map. Fibonacci Quarterly, 46/47(2):115–125, May 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/46_47-2/zarnowski.pdf.

Zhao:2002:SRR

[ZD02] Xiqiang Zhao and Shuangshuang Dings. Sequences related to Riordan arrays. Fibonacci Quarterly, 40(3):247–252, June 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-3/zhao.pdf.

Zhao:2004:GSR

[ZD04] Xiqiang Zhao and Shuangshuang Ding. A generalized summation rule related to Stirling numbers. Fibonacci Quarterly, 42(3):194–201, August 2004. CODEN FIBQAU. ISSN

0015-0517.  URL http://www.fq.math.ca/Papers1/42-3/
quartxiqiang03_2004.pdf.

Zeckendorf:1972:GFN

[Zec72a] E. Zeckendorf. A generalized Fibonacci numeration. Fibonacci Quarterly, 10(4):365–372, October 1972. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/10-
4/zeckendorf.pdf.

Zeckendorf:1972:RNN

[Zec72b] E. Zeckendorf. Représentation des nombres naturels par une somme de nombres de Fibonacci ou de nombres de Lucas. (French) [Representation of the natural numbers by a sum of Fibonacci or Lucas numbers]. Bulletin de la Société royale des sciences de Liège, 41:179–182, 1972. CODEN BSRSA6. ISSN 0037-9565.

Zeitlin:1963:CNI

[Zei63] David Zeitlin. Classroom notes: On identities for Fibonacci numbers. American Mathematical Monthly, 70(9):987–991, November 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Zeitlin:1964:SFF

[Zei64] David Zeitlin. On summation formulas for Fibonacci and Lucas numbers. Fibonacci Quarterly, 2(2):105–107, April 1964. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/2-2/zeitlin.pdf.

Zeitlin:1965:PIS

[Zei65] David Zeitlin. Power identities for sequences defined by W_{n+2} = dW_{n+1} − dW_{n}. Fibonacci Quarterly, 3(4):241–256, December 1965. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/3-4/zeitlin.pdf.

Zeitlin:1967:SFI

[Zei67] David Zeitlin. On summation formulas and identities for Fibonacci numbers. Fibonacci Quarterly, 5(1):1–43, February 1967. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/5-1/zeitlin.pdf.

Zeitlin:1970:DWE

[Zei70] David Zeitlin. On determinant whose elements are products of recursive sequences. Fibonacci Quarterly, 8(4):350–359, October 1970. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/8-4/zeitlin.pdf.

Zeitlin:1971:GIR

[Zei71a] David Zeitlin. General identities for recurrent sequences of order two. Fibonacci Quarterly, 9(4):357–388, October 1971. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/9-4/zeitlin-a.pdf.

Zeitlin:1971:LE

[Zei71b] David Zeitlin. Letter to the editor. Fibonacci Quarterly, 9(1): 34–??, February 1971. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/9-1/zeitlin-a.pdf.

Zeitlin:1972:PIS

[Zei72] David Zeitlin. A product identity for sequences defined by W_{n+2} = dW_{n+1} − cW_{n}. Fibonacci Quarterly, 10(4):397–402, October 1972. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/10-4/zeitlin-a.pdf.

Zeitlin:1973:LE

[Zei73] David Zeitlin. Letter to the editor. Fibonacci Quarterly, 11(1): 62–??, February 1973. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/11-1/letter-a.pdf.

Zeitlin:1974:NFL

[Zei74] David Zeitlin. A note on Fermat’s Last Theorem. Fibonacci Quarterly, 12(4):368–??, December 1974. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/12-
4/zeitlin-a.pdf.

Zeilberger:1977:SPI

[Zei77] Doron Zeilberger. Sums of products involving Fibonacci sequences. Fibonacci Quarterly, 15(2):155–??, April 1977. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/15-2/zeilberger.pdf.

Zeitlin:1978:ICP

[Zei78] David Zeitlin. An inequality for a class of polynomials. Fibonacci Quarterly, 16(2):128–??, April 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-2/
zeitlin-a.pdf.

Zeilberger:1989:CPA

[Zei89] Doron Zeilberger. A combinatorial problem that arose in biophysics. Fibonacci Quarterly, 27(4):372–??, August 1989. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/27-4/zeilberger.pdf.

Zeilberger:1996:SAW

[Zei96] Doron Zeilberger. Self-avoiding walks, the language of science, and Fibonacci numbers. Journal of Statistical Planning and Inference, 54(1):135–138, September 1996. CO-DEN JSPIDN. ISSN 0378-3758 (print), 1873-1171 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
037837589500162X.

Zelator:1992:DEX

[Zel92] Konstantine Dabmian Zelator. The Diophantine equation x^{2} + a^{2}y^{m} = z^{2}n with (x, ay) = 1. Fibonacci Quarterly, 30(4):305– 309, November 1992. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/30-4/zelator.pdf.

Zelator:1993:IPN

[Zel93] Konstantine Zelator. On independent Pythagorean numbers. Fibonacci Quarterly, 31(4):299–301, November 1993. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/31-4/zelator.pdf.

Zelator:1997:CNC

[Zel97] Konstantine Dabmian Zelator. On a class of non-congruent and non-Pythagorean numbers. Fibonacci Quarterly, 35(2):98–101, May 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/35-2/zelator.pdf.

Zenz:1978:FMB

[Zen78] F. A. Zenz. The fluid mechanics of bubbling beds. Fibonacci Quarterly, 16(2):171–183, April 1978. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/16-2/
zenz.pdf.

Zhang:2014:DTE

[ZG14] Yifan Zhang and George Grossman. Diophantine triples and the extendibility of {1, 2, 5}, {1, 5, 10}. Fibonacci Quarterly, 52(5): 212–??, December 2014. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Papers1/52-5/Zhang.pdf.

Zhang:2017:CPG

[ZG17] Yifan Zhang and George Grossman. A combinatorial proof for the generating function of powers of the Fibonacci sequence. Fibonacci Quarterly, 55(3):235–??, August 2017. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/55-3/zhang.pdf.

Zhou:2003:OFI

[ZH03] Chizhong Zhou and Fredric T. Howard. On the k th-order F-L identity. Fibonacci Quarterly, 41(4):345–351, August 2003. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/41-4/zhou.pdf.

Zhang:1997:SIIa

[Zha97a] Wenpeng Zhang. Some identities involving the Fibonacci numbers. Fibonacci Quarterly, 35(3):225–228, August 1997. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-3/zhang-wenpeng.pdf.

Zhang:1997:SIIb

[Zha97b] Zhizheng Zhang. Some identities involving generalized secondorder integer sequences. Fibonacci Quarterly, 35(3):265–267, August 1997. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/35-3/zhang-zhizheng.pdf.

Zhang:1997:SPG

[Zha97c] Zhizheng Zhang. Some properties of the generalized Fibonacci sequences C_{n} = C_{n−1} + C_{n−2} + r. Fibonacci Quarterly, 35(2): 169–171, May 1997. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/35-2/zhang.pdf.

Zhang:1998:SIIa

[Zha98a] Wenpeng Zhang. Some identities involving the Euler and the central factorial numbers. Fibonacci Quarterly, 36(2):154–157, May 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Scanned/36-2/zhang.pdf.

Zhang:1998:CSA

[Zha98b] Zhizheng Zhang. A class of sequences and the Aitken transformation. Fibonacci Quarterly, 36(1):68–71, February 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-1/zhang.pdf.

Zhang:1999:GFS

[Zha99a] Zhizheng Zhang. Generalized Fibonacci sequences and a generalization of the Q-matrix. Fibonacci Quarterly, 37(3):203–207, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Scanned/37-3/zhang.pdf.

Zhao:1999:NRS

[Zha99b] Feng-Zhen Zhao. Notes on reciprocal series related to Fibonacci and Lucas numbers. Fibonacci Quarterly, 37(3):254–257, August 1999. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/37-3/zhao.pdf.

Zhao:2000:ISI

[Zha00] Feng-Zhen Zhao. The integrity of some infinite series. Fibonacci Quarterly, 38(5):420–424, November 2000. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/38-
5/zhao.pdf.

Zhang:2001:ULS

[Zha01a] Zhenxiang Zhang. Using Lucas sequences to factor large integers near group orders. Fibonacci Quarterly, 39(3):228–237, June 2001. CODEN FIBQAU. ISSN 0015-0517. URL http://www.

fq.math.ca/Scanned/39-3/zhang.pdf.

Zhao:2001:SCR

[Zha01b] Fengzhen Zhao. Summation of certain reciprocal series related to the generalized Fibonacci and Lucas numbers. Fibonacci Quarterly, 39(5):392–397, November 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-
5/zhao1.pdf.

Zhang:2002:CPF

[Zha02] Wenpeng Zhang. On Chebyshev polynomials and Fibonacci numbers. Fibonacci Quarterly, 40(5):424–428, November 2002. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-5/zhang-wenpeng.pdf.

Zhang:2004:SIIa

[Zha04] Wenpeng Zhang. Some identities involving the Fibonacci numbers and Lucas numbers. Fibonacci Quarterly, 42(2):149–154, May 2004. CODEN FIBQAU. ISSN 0015-0517. URL http://

www.fq.math.ca/Papers1/42-2/quartwenpeng02_2004.pdf.

Zhao:2008:CTF

[Zha09] Yufei Zhao. The coefficients of a truncated Fibonacci power series. Fibonacci Quarterly, 46/47(1):53–55, February 2008/2009. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/46_47-1/zhao.pdf.

Zhou:1996:ODS

[Zho96] Chizhong Zhou. On the K th-order derivative sequences of Fibonacci and Lucas polynomials. Fibonacci Quarterly, 34(5):394– 408, November 1996. CODEN FIBQAU. ISSN 0015-0517. URL

http://www.fq.math.ca/Scanned/34-5/zhou.pdf.

Zhou:1997:GAN

[Zho97] Chizhong Zhou. A generalization of the “all or none” divisibility property. Fibonacci Quarterly, 35(2):129–134, May 1997. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/35-2/zhou.pdf.

Zhou:1999:GCL

[Zho99] Chizhong Zhou. A general conclusion on Lucas numbers of the form px^{2}, where p is prime. Fibonacci Quarterly, 37(1):39–45, February 1999. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Scanned/37-1/zhou.pdf.

Zhou:2003:AMT

[Zho03] Chizhong Zhou. Applications of matrix theory to congruence properties of k th-order F-L sequences. Fibonacci Quarterly, 41 (1):48–58, February 2003. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/41-1/zhou.pdf.

Ziegenfus:1963:LET

[Zie63] Charles Ziegenfus. Letter to the editor: Twin primes. Fibonacci Quarterly, 1(3):42–??, October 1963. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/1-3/
ziegenfus.pdf.

Zielinski:2019:FB

[Zie19] Ryan Zielinski. Faulhaber and Bernoulli. Fibonacci Quarterly, 57(1):32–??, February 2019. CODEN FIBQAU. ISSN

0015-0517. URL http://www.fq.math.ca/Abstracts/57-1/
zielinski.pdf.

Zhang:1998:SIIb

[ZJ98] Zhizheng Zhang and Jingyu Jin. Some identities involving generalized Genocchi polynomials and generalized Fibonacci–Lucas sequences. Fibonacci Quarterly, 36(4):329–334, August 1998. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/36-4/zhang2.pdf.

Zhou:2014:SNG

[ZJ14] Jianwei Zhou and Zhaolin Jiang. The spectral norms of gcirculant matrices with classical Fibonacci and Lucas numbers entries. Applied Mathematics and Computation, 233(??):582– 587, May 1, 2014. CODEN AMHCBQ. ISSN 0096-3003 (print),

1873-5649 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0096300314002598.

Zhizheng:1995:RSB

[ZL95] Zhang Zhizheng and Guo Lizhou. Recurrence sequences and Bernoulli polynomials of higher order. Fibonacci Quarterly, 33 (4):359–362, August 1995. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Scanned/33-4/zhang.pdf.

Zhang:1998:GSI

[ZL98] Zhizheng Zhang and Maixue Liu. Generalizations of some identities involving generalized second-order integer sequences. Fibonacci Quarterly, 36(4):327–328, August 1998. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/36-4/zhang1.pdf.

Zhao:2006:FNT

[ZL06] Haixing Zhao and Xueliang Li. On the Fibonacci numbers of trees. Fibonacci Quarterly, 44(1):32–38, February 2006. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/44-1/zhao.pdf.

Zollner:1993:DSL

[Zöl93] Joachim Zöllner. A disjoint system of linear recurring sequences generated by u_{n+2} = u_{n+1} + u_{n} which contains every natural number. Fibonacci Quarterly, 31(2):162–164, May 1993. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/31-2/zollner.pdf.

Zhao:2020:FRE

[ZQZ20] Xiao-Xin Zhao, Wen-Feng Qi, and Jia-Min Zhang. Further results on the equivalence between Galois NFSRs and Fibonacci NFSRs. Designs, Codes, and Cryptography, 88(1):153–171, January 2020. CODEN DCCREC. ISSN 0925-1022 (print), 1573-

7586 (electronic). URL https://link.springer.com/article/
10.1007/s10623-019-00677-y.

Zikan:1985:PCP

[ZS85] Karel Zikan and Edward Schmeichel. A path counting problem in digraphs. Fibonacci Quarterly, 23(1):3–6, February 1985. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/23-1/zikan.pdf.

Zulauf:1988:FSS

[ZT88] A. Zulauf and J. C. Turner. Fibonacci sequences of sets and their duals. Fibonacci Quarterly, 26(2):152–156, May 1988. CO-DEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/26-2/zulauf.pdf.

Zukerman:1976:FRE

[Zuk76] L. G. Zukerman. Fibonacci ratio in electric wave filters. Fibonacci Quarterly, 14(1):25–26, February 1976. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/14-
1/zukerman.pdf.

Zhao:2001:EGS

[ZW01a] Fengzhen Zhao and Tianming Wang. Errata for “Generalizations of Some Identities Involving the Fibonacci Numbers”. Fibonacci Quarterly, 39(5):408–??, November 2001. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/39-
5/errata.pdf. See [ZW01b].

Zhao:2001:GSI

[ZW01b] Fengzhen Zhao and Tianming Wang. Generalizations of some identities involving the Fibonacci numbers. Fibonacci Quarterly, 39(2):165–167, May 2001. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/39-2/zhao.pdf.

See errata [ZW01a].

Zhao:2001:SIG

[ZW01c] Fengzhen Zhao and Tianming Wang. Some identities for the generalized Fibonacci and Lucas functions. Fibonacci Quarterly, 39 (5):436–438, November 2001. CODEN FIBQAU. ISSN 0015-

0517. URL http://www.fq.math.ca/Scanned/39-5/zhao2.
pdf.

Zhang:2002:NCC

[ZW02] Zhizheng Zhang and Xin Wang. A note on a class of computational formulas involving the multiple sum of recurrence sequences. Fibonacci Quarterly, 40(5):394–397, November 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Scanned/40-5/zhang-zhizheng.pdf.

Zhao:2003:SII

[ZW03] Feng-Zhen Zhao and Tianming Wang. Some identities involving the powers of the generalized Fibonacci numbers. Fibonacci Quarterly, 41(1):7–12, February 2003. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Scanned/41-1/

zhao.pdf. See errata [ZW06].

Zhang:2004:SIIb

[ZW04a] Zhizheng Zhang and Jun Wang. On some identities involving the Chebyshev polynomials. Fibonacci Quarterly, 42(3):245–249, August 2004. CODEN FIBQAU. ISSN 0015-0517. URL http:/

/www.fq.math.ca/Papers1/42-3/quartwang03_2004.pdf.

Zhao:2004:NSC

[ZW04b] Feng-Zhen Zhao and Tianming Wang. A note on summation of certain reciprocal series. Fibonacci Quarterly, 42(1):66–69, February 2004. CODEN FIBQAU. ISSN 0015-0517. URL http:

//www.fq.math.ca/Papers1/42-1/quartzhao01_2004.pdf.

Zhao:2004:SRG

[ZW04c] Feng-Zhen Zhao and Tianming Wang. Some results on generalized Fibonacci and Lucas numbers and Dedekind sums. Fibonacci Quarterly, 42(3):250–255, August 2004. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers1/42-
3/quartzhao03_2004.pdf.

Zhao:2006:ESI

[ZW06] Feng-Zhen Zhao and Tianming Wang. Errata for “Some identities involving the powers of the generalized Fibonacci numbers”. Fibonacci Quarterly, 44(1):3–6, February 2006. CODEN FIBQAU.

ISSN 0015-0517. URL http://www.fq.math.ca/Papers/44-1/
quartzhao01_errata_2006.pdf. See [ZW03].

Zhang:2002:FSD

[ZY02] Wenpeng Zhang and Yuan Yi. On Farey series and Dedekind sums. Fibonacci Quarterly, 40(2):170–174, May 2002. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Scanned/40-2/zhang.pdf.

Zeng:2006:AWF

[ZZ06] Jiang Zeng and Jin Zhou. Applications of Waring’s formula to some identities of Chebyshev polynomials. Fibonacci Quarterly, 44(2):117–120, May 2006. CODEN FIBQAU. ISSN 0015-0517.

URL http://www.fq.math.ca/Abstracts/44-2/zeng.pdf.